The GNU C Library Reference Manual

The GNU C Library Reference Manual

Sandra Loosemore

with

Richard M. Stallman, Roland McGrath, Andrew Oram, and Ulrich Drepper

Edition 0.07 DRAFT last updated 4 Oct 1996 for version 2.00 Beta Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993, ’94, ’95, ’96, ’97 Free Software Foundation, Inc.

Published by the Free Software Foundation 59 Temple Place – Suite 330, Boston, MA 02111-1307 USA Printed copies are available for $50 each. ISBN 1-882114-53-1

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided also that the section entitled “GNU Library General Public License” is included exactly as in the original, and provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that the text of the translation of the section entitled “GNU Library General Public License” must be approved for accuracy by the Foundation.

1 Introduction

The C language provides no built-in facilities for performing such common operations as input/output, memory management, string manipulation, and the like. Instead, these facilities are defined in a standard library, which you compile and link with your programs.

The GNU C library, described in this document, defines all of the library functions that are specified by the ISO C standard, as well as additional features specific to POSIX and other derivatives of the Unix operating system, and extensions specific to the GNU system.

The purpose of this manual is to tell you how to use the facilities of the GNU library. We have mentioned which features belong to which standards to help you identify things that are potentially nonportable to other systems. But the emphasis in this manual is not on strict portability.

1.1 Getting Started

This manual is written with the assumption that you are at least somewhat familiar with the C programming language and basic programming concepts. Specifically, familiarity with ISO standard C (see Section 1.2.1 [ISO C], page 2), rather than “traditional” pre-ISO C dialects, is assumed.

The GNU C library includes several header files, each of which provides definitions and declarations for a group of related facilities; this information is used by the C compiler when processing your program. For example, the header file ‘stdio.h’ declares facilities for performing input and output, and the header file ‘string.h’ declares string processing utilities. The organization of this manual generally follows the same division as the header files.

If you are reading this manual for the first time, you should read all of the introductory material and skim the remaining chapters. There are a lot of functions in the GNU C library and it’s not realistic to expect that you will be able to remember exactly how to use each and every one of them. It’s more important to become generally familiar with the kinds of facilities that the library provides, so that when you are writing your programs you can recognize when to make use of library functions, and where in this manual you can find more specific information about them.

1.2 Standards and Portability

This section discusses the various standards and other sources that the GNU C library is based upon. These sources include the ISO C and POSIX standards, and the System V and Berkeley Unix implementations.

The primary focus of this manual is to tell you how to make effective use of the GNU library facilities. But if you are concerned about making your programs compatible with these standards, or portable to operating systems other than GNU, this can affect how you use the library. This section gives you an overview of these standards, so that you will know what they are when they are mentioned in other parts of the manual.

See Appendix B [Summary of Library Facilities], page 573, for an alphabetical list of the functions and other symbols provided by the library. This list also states which standards each function or symbol comes from.

1.2.1 ISO C

The GNU C library is compatible with the C standard adopted by the American National Standards Institute (ANSI): American National Standard X3.159-1989—“ANSI C” and later by the International Standardization Organization (ISO): ISO/IEC 9899:1990, “Programming languages—C”. We here refer to the standard as ISO C since this is the more general standard in respect of ratification. The header files and library facilities that make up the GNU library are a superset of those specified by the ISO C standard.

If you are concerned about strict adherence to the ISO C standard, you should use the ‘-ansi’ option when you compile your programs with the GNU C compiler. This tells the compiler to define only ISO standard features from the library header files, unless you explicitly ask for additional features. See Section 1.3.4 [Feature Test Macros], page 8, for information on how to do this.

Being able to restrict the library to include only ISO C features is important because ISO C puts limitations on what names can be defined by the library implementation, and the GNU extensions don’t fit these limitations. See Section 1.3.3 [Reserved Names], page 6, for more information about these restrictions.

This manual does not attempt to give you complete details on the differences between ISO C and older dialects. It gives advice on how to write programs to work portably under multiple C dialects, but does not aim for completeness.

1.2.2 POSIX (The Portable Operating System Interface)

The GNU library is also compatible with the IEEE POSIX family of standards, known more formally as the Portable Operating System Interface for Computer Environments. POSIX is derived mostly from various versions of the Unix operating system.

The library facilities specified by the POSIX standards are a superset of those required by ISO C; POSIX specifies additional features for ISO C functions, as well as specifying new additional functions. In general, the additional requirements and functionality defined by the POSIX standards are aimed at providing lower-level support for a particular kind of operating system environment, rather than general programming language support which can run in many diverse operating system environments.

The GNU C library implements all of the functions specified in IEEE Std 1003.1-1990, the POSIX System Application Program Interface, commonly referred to as POSIX.1. The primary extensions to the ISO C facilities specified by this standard include file system interface primitives (see Chapter 9 [File System Interface], page 181), device-specific terminal control functions (see Chapter 12 [Low-Level Terminal Interface], page 271), and process control functions (see Chapter 23 [Processes], page 467).

Some facilities from IEEE Std 1003.2-1992, the POSIX Shell and Utilities standard (POSIX.2) are also implemented in the GNU library. These include utilities for dealing with regular expressions and other pattern matching facilities (see Chapter 16 [Pattern Matching], page 323).

1.2.3 Berkeley Unix

The GNU C library defines facilities from some versions of Unix which are not formally standardized, specifically from the 4.2 BSD, 4.3 BSD, and 4.4 BSD Unix systems (also known as Berkeley Unix) and from SunOS (a popular 4.2 BSD derivative that includes some Unix System V functionality). These systems support most of the ISO C and POSIX facilities, and 4.4 BSD and newer releases of SunOS in fact support them all.

The BSD facilities include symbolic links (see Section 9.4 [Symbolic Links], page 190), the select function (see Section 8.6 [Waiting for Input or Output], page 162), the BSD signal functions (see Section 21.10 [BSD Signal Handling], page 444), and sockets (see Chapter 11 [Sockets], page 221).

1.2.4 SVID (The System V Interface Description)

The System V Interface Description (SVID) is a document describing the AT&T Unix System V operating system. It is to some extent a superset of the POSIX standard (see Section 1.2.2 [POSIX (The Portable Operating System Interface)], page 2).

The GNU C library defines some of the facilities required by the SVID that are not also required by the ISO C or POSIX standards, for compatibility with System V Unix and other Unix systems (such as SunOS) which include these facilities. However, many of the more obscure and less generally useful facilities required by the SVID are not included. (In fact, Unix System V itself does not provide them all.)

1.3 Using the Library

This section describes some of the practical issues involved in using the GNU C library.

1.3.1 Header Files

Libraries for use by C programs really consist of two parts: header files that define types and macros and declare variables and functions; and the actual library or archive that contains the definitions of the variables and functions.

(Recall that in C, a declaration merely provides information that a function or variable exists and gives its type. For a function declaration, information about the types of its arguments might be provided as well. The purpose of declarations is to allow the compiler to correctly process references to the declared variables and functions. A definition, on the other hand, actually allocates storage for a variable or says what a function does.)

In order to use the facilities in the GNU C library, you should be sure that your program source files include the appropriate header files. This is so that the compiler has declarations of these facilities available and can correctly process references to them. Once your program has been compiled, the linker resolves these references to the actual definitions provided in the archive file.

Header files are included into a program source file by the ‘#include’ preprocessor directive. The C language supports two forms of this directive; the first,

     #include "header"

is typically used to include a header file header that you write yourself; this would contain definitions and declarations describing the interfaces between the different parts of your particular application. By contrast,

     #include <file.h>

is typically used to include a header file ‘file.h’ that contains definitions and declarations for a standard library. This file would normally be installed in a standard place by your system administrator. You should use this second form for the C library header files.

Typically, ‘#include’ directives are placed at the top of the C source file, before any other code. If you begin your source files with some comments explaining what the code in the file does (a good idea), put the ‘#include’ directives immediately afterwards, following the feature test macro definition (see Section 1.3.4 [Feature Test Macros], page 8).

For more information about the use of header files and ‘#include’ directives, see section “Header Files” in The GNU C Preprocessor Manual.

The GNU C library provides several header files, each of which contains the type and macro definitions and variable and function declarations for a group of related facilities. This means that your programs may need to include several header files, depending on exactly which facilities you are using.

Some library header files include other library header files automatically. However, as a matter of programming style, you should not rely on this; it is better to explicitly include all the header files required for the library facilities you are using. The GNU C library header files have been written in such a way that it doesn’t matter if a header file is accidentally included more than once; including a header file a second time has no effect. Likewise, if your program needs to include multiple header files, the order in which they are included doesn’t matter.

Compatibility Note: Inclusion of standard header files in any order and any number of times works in any ISO C implementation. However, this has traditionally not been the case in many older C implementations.

Strictly speaking, you don’t have to include a header file to use a function it declares; you could declare the function explicitly yourself, according to the specifications in this manual. But it is usually better to include the header file because it may define types and macros that are not otherwise available and because it may define more efficient macro replacements for some functions. It is also a sure way to have the correct declaration.

1.3.2 Macro Definitions of Functions

If we describe something as a function in this manual, it may have a macro definition as well. This normally has no effect on how your program runs— the macro definition does the same thing as the function would. In particular, macro equivalents for library functions evaluate arguments exactly once, in the same way that a function call would. The main reason for these macro definitions is that sometimes they can produce an inline expansion that is considerably faster than an actual function call.

Taking the address of a library function works even if it is also defined as a macro. This is because, in this context, the name of the function isn’t followed by the left parenthesis that is syntactically necessary to recognize a macro call.

You might occasionally want to avoid using the macro definition of a function—perhaps to make your program easier to debug. There are two ways you can do this:

• You can avoid a macro definition in a specific use by enclosing the name of the function in parentheses. This works because the name of the function doesn’t appear in a syntactic context where it is recognizable as a macro call.

• You can suppress any macro definition for a whole source file by using the ‘#undef’ preprocessor directive, unless otherwise stated explicitly in the description of that facility. For example, suppose the header file ‘stdlib.h’ declares a function

named abs with

     extern int abs (int);

and also provides a macro definition for abs. Then, in:

     #include <stdlib.h>
     int f (int *i) { return (abs (++*i)); }

the reference to abs might refer to either a macro or a function. On the other hand, in each of the following examples the reference is to a function and not a macro.

     #include <stdlib.h>
     int g (int *i) { return ((abs)(++*i)); }
     #undef abs
     int h (int *i) { return (abs (++*i)); }

Since macro definitions that double for a function behave in exactly the same way as the actual function version, there is usually no need for any of these methods. In fact, removing macro definitions usually just makes your program slower.

1.3.3 Reserved Names

The names of all library types, macros, variables and functions that come from the ISO C standard are reserved unconditionally; your program may not redefine these names. All other library names are reserved if your program explicitly includes the header file that defines or declares them. There are several reasons for these restrictions:

• Other people reading your code could get very confused if you were using a function named exit to do something completely different from what the standard exit function does, for example. Preventing this situation helps to make your programs easier to understand and contributes to modularity and maintainability.

• It avoids the possibility of a user accidentally redefining a library function that is called by other library functions. If redefinition were allowed, those other functions would not work properly.

• It allows the compiler to do whatever special optimizations it pleases on calls to these functions, without the possibility that they may have been redefined by the user. Some library facilities, such as those for dealing with variadic arguments (see Section A.2 [Variadic Functions], page 555) and non-local exits (see Chapter 20 [Non-Local Exits], page 389), actually require a considerable amount of cooperation on the part of the C compiler, and implementationally it might be easier for the compiler to treat these as built-in parts of the language.

In addition to the names documented in this manual, reserved names include all external identifiers (global functions and variables) that begin with an underscore (‘_’) and all identifiers regardless of use that begin with either two underscores or an underscore followed by a capital letter are reserved names. This is so that the library and header files can define functions, variables, and macros for internal purposes without risk of conflict with names in user programs.

Some additional classes of identifier names are reserved for future extensions to the C language or the POSIX.1 environment. While using these names for your own purposes right now might not cause a problem, they do raise the possibility of conflict with future versions of the C or POSIX standards, so you should avoid these names.

• Names beginning with a capital ‘E’ followed a digit or uppercase letter may be used for additional error code names. See Chapter 2 [Error Reporting], page 15.

• Names that begin with either ‘is’ or ‘to’ followed by a lowercase letter may be used for additional character testing and conversion functions. See Chapter 4 [Character Handling], page 61.

• Names that begin with ‘LC_’ followed by an uppercase letter may be used for additional macros specifying locale attributes. See Chapter 19 [Locales and Internationalization], page 379.

• Names of all existing mathematics functions (see Chapter 13 [Mathematics], page 297) suffixed with ‘f’ or ‘l’ are reserved for corresponding functions that operate on float and long double arguments, respectively.

• Names that begin with ‘SIG’ followed by an uppercase letter are reserved for additional signal names. See Section 21.2 [Standard Signals], page 395.

• Names that begin with ‘SIG_’ followed by an uppercase letter are reserved for additional signal actions. See Section 21.3.1 [Basic Signal Handling], page 406.

• Names beginning with ‘str’, ‘mem’, or ‘wcs’ followed by a lowercase letter are reserved for additional string and array functions. See Chapter 5 [String and Array Utilities], page 65.

• Names that end with ‘_t’ are reserved for additional type names.

In addition, some individual header files reserve names beyond those that they actually define. You only need to worry about these restrictions if your program includes that particular header file.

• The header file ‘dirent.h’ reserves names prefixed with ‘d_’.

• The header file ‘fcntl.h’ reserves names prefixed with ‘l_’, ‘F_’, ‘O_’, and ‘S_’.

• The header file ‘grp.h’ reserves names prefixed with ‘gr_’.

• The header file ‘limits.h’ reserves names suffixed with ‘_MAX’.

• The header file ‘pwd.h’ reserves names prefixed with ‘pw_’.

• The header file ‘signal.h’ reserves names prefixed with ‘sa_’ and ‘SA_’.

• The header file ‘sys/stat.h’ reserves names prefixed with ‘st_’ and ‘S_’.

• The header file ‘sys/times.h’ reserves names prefixed with ‘tms_’.

• The header file ‘termios.h’ reserves names prefixed with ‘c_’, ‘V’, ‘I’, ‘O’, and ‘TC’; and names prefixed with ‘B’ followed by a digit.

1.3.4 Feature Test Macros

The exact set of features available when you compile a source file is controlled by which feature test macros you define.

If you compile your programs using ‘gcc -ansi’, you get only the ISO C library features, unless you explicitly request additional features by defining one or more of the feature macros. See section “GNU CC Command Options” in The GNU CC Manual, for more information about GCC options.

You should define these macros by using ‘#define’ preprocessor directives at the top of your source code files. These directives must come before any #include of a system header file. It is best to make them the very first thing in the file, preceded only by comments. You could also use the ‘-D’ option to GCC, but it’s better if you make the source files indicate their own meaning in a self-contained way.

Macro POSIX SOURCE

If you define this macro, then the functionality from the POSIX.1 standard (IEEE Standard 1003.1) is available, as well as all of the ISO C facilities.

Macro POSIX C SOURCE

If you define this macro with a value of 1, then the functionality from the POSIX.1 standard (IEEE Standard 1003.1) is made available. If you define this macro with a value of 2, then both the functionality from the POSIX.1 standard and the functionality from the POSIX.2 standard (IEEE Standard 1003.2) are made available. This is in addition to the ISO C facilities.

Macro BSD SOURCE

If you define this macro, functionality derived from 4.3 BSD Unix is included as well as the ISO C, POSIX.1, and POSIX.2 material.

Some of the features derived from 4.3 BSD Unix conflict with the corresponding features specified by the POSIX.1 standard. If this macro is defined, the 4.3 BSD definitions take precedence over the POSIX definitions. Due to the nature of some of the conflicts between 4.3 BSD and POSIX.1, you need to use a special BSD compatibility library when linking programs compiled for BSD compatibility. This is because some functions must be defined in two different ways, one of them in the normal C library, and one of them in the compatibility library. If your program defines _BSD_SOURCE, you must give the option ‘-lbsd-compat’ to the compiler or linker when linking the program, to tell it to find functions in this special compatibility library before looking for them in the normal C library.

Macro SVID SOURCE

If you define this macro, functionality derived from SVID is included as well as the ISO C, POSIX.1, POSIX.2, and X/Open material.

Macro XOPEN SOURCE

If you define this macro, functionality described in the X/Open Portability Guide is included. This is a superset of the POSIX.1 and POSIX.2 functionality and in fact _POSIX_SOURCE and _POSIX_C_ SOURCE are automatically defined. As the unification of all Unices, functionality only available in BSD and SVID is also included. If the macro _XOPEN_SOURCE_EXTENDED is also defined, even more functionality is available. The extra functions will make all functions available which are necessary for the X/Open Unix brand.

Macro GNU SOURCE

If you define this macro, everything is included: ISO C, POSIX.1, POSIX.2, BSD, SVID, X/Open, and GNU extensions. In the cases where POSIX.1 conflicts with BSD, the POSIX definitions take precedence. If you want to get the full effect of _GNU_SOURCE but make the BSD definitions take precedence over the POSIX definitions, use this sequence of definitions:

     #define _GNU_SOURCE
     #define _BSD_SOURCE
     #define _SVID_SOURCE

Note that if you do this, you must link your program with the BSD compatibility library by passing the ‘-lbsd-compat’ option to the compiler or linker. Note: If you forget to do this, you may get very strange errors at run time.

Macro REENTRANT Macro THREAD SAFE

If you define one of these macros, reentrant versions of several functions get declared. Some of the functions are specified in POSIX.1c but many others are only available on a few other systems or are unique to GNU libc. The problem is that the standardization of the thread safe C library interface still is behind. Unlike on some other systems no special version of the C library must be used for linking. There is only one version but while compiling this it must have been specified to compile as thread safe.

We recommend you use _GNU_SOURCE in new programs. If you don’t specify the ‘-ansi’ option to GCC and don’t define any of these macros explicitly, the effect is the same as defining _POSIX_C_SOURCE to 2 and _

POSIX_SOURCE, _SVID_SOURCE, and _BSD_SOURCE to 1.

When you define a feature test macro to request a larger class of features, it is harmless to define in addition a feature test macro for a subset of those features. For example, if you define _POSIX_C_SOURCE, then defining _POSIX_SOURCE as well has no effect. Likewise, if you define _GNU_SOURCE,

then defining either _POSIX_SOURCE or _POSIX_C_SOURCE or _SVID_SOURCE

as well has no effect.

Note, however, that the features of _BSD_SOURCE are not a subset of any of the other feature test macros supported. This is because it defines BSD features that take precedence over the POSIX features that are requested by the other macros. For this reason, defining _BSD_SOURCE in addition to the other feature test macros does have an effect: it causes the BSD features to take priority over the conflicting POSIX features.

1.4 Roadmap to the Manual

Here is an overview of the contents of the remaining chapters of this manual.

• Chapter 2 [Error Reporting], page 15, describes how errors detected by the library are reported.

• Appendix A [C Language Facilities in the Library], page 553, contains information about library support for standard parts of the C language, including things like the sizeof operator and the symbolic constant NULL, how to write functions accepting variable numbers of arguments, and constants describing the ranges and other properties of the numerical types. There is also a simple debugging mechanism which allows you to put assertions in your code, and have diagnostic messages printed if the tests fail.

• Chapter 3 [Memory Allocation], page 31, describes the GNU library’s facilities for dynamic allocation of storage. If you do not know in advance how much storage your program needs, you can allocate it dynamically instead, and manipulate it via pointers.

• Chapter 4 [Character Handling], page 61, contains information about character classification functions (such as isspace) and functions for performing case conversion.

• Chapter 5 [String and Array Utilities], page 65, has descriptions of functions for manipulating strings (null-terminated character arrays) and general byte arrays, including operations such as copying and comparison.

• Chapter 6 [Input/Output Overview], page 85, gives an overall look at the input and output facilities in the library, and contains information about basic concepts such as file names.

• Chapter 7 [Input/Output on Streams], page 93, describes I/O operations involving streams (or FILE * objects). These are the normal C library functions from ‘stdio.h’.

• Chapter 8 [Low-Level Input/Output], page 151, contains information about I/O operations on file descriptors. File descriptors are a lowerlevel mechanism specific to the Unix family of operating systems.

• Chapter 9 [File System Interface], page 181, has descriptions of operations on entire files, such as functions for deleting and renaming them and for creating new directories. This chapter also contains information about how you can access the attributes of a file, such as its owner and file protection modes.

• Chapter 10 [Pipes and FIFOs], page 215, contains information about simple interprocess communication mechanisms. Pipes allow communication between two related processes (such as between a parent and child), while FIFOs allow communication between processes sharing a common file system on the same machine.

• Chapter 11 [Sockets], page 221, describes a more complicated interprocess communication mechanism that allows processes running on different machines to communicate over a network. This chapter also contains information about Internet host addressing and how to use the system network databases.

• Chapter 12 [Low-Level Terminal Interface], page 271, describes how you can change the attributes of a terminal device. If you want to disable echo of characters typed by the user, for example, read this chapter.

• Chapter 13 [Mathematics], page 297, contains information about the math library functions. These include things like random-number generators and remainder functions on integers as well as the usual trigonometric and exponential functions on floating-point numbers.

• Chapter 14 [Low-Level Arithmetic Functions], page 305, describes functions for simple arithmetic, analysis of floating-point values, and reading numbers from strings.

• Chapter 15 [Searching and Sorting], page 317, contains information about functions for searching and sorting arrays. You can use these functions on any kind of array by providing an appropriate comparison function.

• Chapter 16 [Pattern Matching], page 323, presents functions for matching regular expressions and shell file name patterns, and for expanding words as the shell does.

• Chapter 17 [Date and Time], page 339, describes functions for measuring both calendar time and CPU time, as well as functions for setting alarms and timers.

• Chapter 18 [Extended Characters], page 367, contains information about manipulating characters and strings using character sets larger than will fit in the usual char data type.

• Chapter 19 [Locales and Internationalization], page 379, describes how selecting a particular country or language affects the behavior of the library. For example, the locale affects collation sequences for strings and how monetary values are formatted.

• Chapter 20 [Non-Local Exits], page 389, contains descriptions of the setjmp and longjmp functions. These functions provide a facility for goto-like jumps which can jump from one function to another.

• Chapter 21 [Signal Handling], page 393, tells you all about signals— what they are, how to establish a handler that is called when a particular kind of signal is delivered, and how to prevent signals from arriving during critical sections of your program.

• Chapter 22 [Process Startup and Termination], page 447, tells how your programs can access their command-line arguments and environment variables.

• Chapter 23 [Processes], page 467, contains information about how to start new processes and run programs.

• Chapter 24 [Job Control], page 479, describes functions for manipulating process groups and the controlling terminal. This material is probably only of interest if you are writing a shell or other program which handles job control specially.

• Chapter 25 [System Databases and Name Service Switch], page 501, describes the services which are available for looking up names in the system databases, how to determine which service is used for which database, and how these services are implemented so that contributors can design their own services.

• Section 26.12 [User Database], page 521, and Section 26.13 [Group Database], page 525, tell you how to access the system user and group databases.

• Chapter 27 [System Information], page 533, describes functions for getting information about the hardware and software configuration your program is executing under.

• Chapter 28 [System Configuration Parameters], page 537, tells you how you can get information about various operating system limits. Most of these parameters are provided for compatibility with POSIX.

• Appendix B [Summary of Library Facilities], page 573, gives a summary of all the functions, variables, and macros in the library, with complete data types and function prototypes, and says what standard or system each is derived from.

• Appendix C [Library Maintenance], page 667, explains how to build and install the GNU C library on your system, how to report any bugs you might find, and how to add new functions or port the library to a new system. If you already know the name of the facility you are interested in, you can

look it up in Appendix B [Summary of Library Facilities], page 573. This gives you a summary of its syntax and a pointer to where you can find a more detailed description. This appendix is particularly useful if you just want to verify the order and type of arguments to a function, for example. It also tells you what standard or system each function, variable, or macro is derived from.

14 The GNU C Library

2 Error Reporting

Many functions in the GNU C library detect and report error conditions, and sometimes your programs need to check for these error conditions. For example, when you open an input file, you should verify that the file was actually opened correctly, and print an error message or take other appropriate action if the call to the library function failed.

This chapter describes how the error reporting facility works. Your program should include the header file ‘errno.h’ to use this facility.

2.1 Checking for Errors

Most library functions return a special value to indicate that they have failed. The special value is typically -1, a null pointer, or a constant such as EOF that is defined for that purpose. But this return value tells you only that an error has occurred. To find out what kind of error it was, you need to look at the error code stored in the variable errno. This variable is declared in the header file ‘errno.h’.

Variable volatile int errno

The variable errno contains the system error number. You can change the value of errno. Since errno is declared volatile, it might be changed asynchronously by a signal handler; see Section 21.4 [Defining Signal Handlers], page 413. However, a properly written signal handler saves and restores the value of errno, so you generally do not need to worry about this possibility except when writing signal handlers. The initial value of errno at program startup is zero. Many library functions are guaranteed to set it to certain nonzero values when they encounter certain kinds of errors. These error conditions are listed for each function. These functions do not change errno when they succeed; thus, the value of errno after a successful call is not necessarily zero, and you should not use errno to determine whether a call failed. The proper way to do that is documented for each function. If the call the failed, you can examine errno. Many library functions can set errno to a nonzero value as a result of calling other library functions which might fail. You should assume that any library function might alter errno when the function returns an error. Portability Note: ISO C specifies errno as a “modifiable lvalue” rather than as a variable, permitting it to be implemented as a macro. For example, its expansion might involve a function call, like *_errno (). In fact, that is what it is on the GNU system itself. The GNU library, on non-GNU systems, does whatever is right for the particular system.

There are a few library functions, like sqrt and atan, that return a perfectly legitimate value in case of an error, but also set errno. For these functions, if you want to check to see whether an error occurred, the recommended method is to set errno to zero before calling the function, and then check its value afterward.

All the error codes have symbolic names; they are macros defined in ‘errno.h’. The names start with ‘E’ and an upper-case letter or digit; you should consider names of this form to be reserved names. See Section 1.3.3 [Reserved Names], page 6.

The error code values are all positive integers and are all distinct, with one exception: EWOULDBLOCK and EAGAIN are the same. Since the values are distinct, you can use them as labels in a switch statement; just don’t use both EWOULDBLOCK and EAGAIN. Your program should not make any other assumptions about the specific values of these symbolic constants.

The value of errno doesn’t necessarily have to correspond to any of these macros, since some library functions might return other error codes of their own for other situations. The only values that are guaranteed to be meaningful for a particular library function are the ones that this manual lists for that function.

On non-GNU systems, almost any system call can return EFAULT if it is given an invalid pointer as an argument. Since this could only happen as a result of a bug in your program, and since it will not happen on the GNU system, we have saved space by not mentioning EFAULT in the descriptions of individual functions.

In some Unix systems, many system calls can also return EFAULT if given as an argument a pointer into the stack, and the kernel for some obscure reason fails in its attempt to extend the stack. If this ever happens, you should probably try using statically or dynamically allocated memory instead of stack memory on that system.

2.2 Error Codes

The error code macros are defined in the header file ‘errno.h’. All of them expand into integer constant values. Some of these error codes can’t occur on the GNU system, but they can occur using the GNU library on other systems.

Macro int EPERM

Operation not permitted; only the owner of the file (or other resource) or processes with special privileges can perform the operation.

Macro int ENOENT

No such file or directory. This is a “file doesn’t exist” error for ordinary files that are referenced in contexts where they are expected to already exist.

Macro int ESRCH

No process matches the specified process ID.

Macro int EINTR

Interrupted function call; an asynchronous signal occurred and prevented completion of the call. When this happens, you should try the call again. You can choose to have functions resume after a signal that is handled, rather than failing with EINTR; see Section 21.5 [Primitives Interrupted by Signals], page 425.

Macro int EIO

Input/output error; usually used for physical read or write errors.

Macro int ENXIO

No such device or address. The system tried to use the device represented by a file you specified, and it couldn’t find the device. This can mean that the device file was installed incorrectly, or that the physical device is missing or not correctly attached to the computer.

Macro int E2BIG

Argument list too long; used when the arguments passed to a new program being executed with one of the exec functions (see Section 23.5 [Executing a File], page 470) occupy too much memory space. This condition never arises in the GNU system.

Macro int ENOEXEC

Invalid executable file format. This condition is detected by the exec functions; see Section 23.5 [Executing a File], page 470.

Macro int EBADF

Bad file descriptor; for example, I/O on a descriptor that has been closed or reading from a descriptor open only for writing (or vice versa).

Macro int ECHILD

There are no child processes. This error happens on operations that are supposed to manipulate child processes, when there aren’t any processes to manipulate.

Macro int EDEADLK

Deadlock avoided; allocating a system resource would have resulted in a deadlock situation. The system does not guarantee that it will notice all such situations. This error means you got lucky and the system noticed; it might just hang. See Section 8.11 [File Locks], page 175, for an example.

Macro int ENOMEM

No memory available. The system cannot allocate more virtual memory because its capacity is full.

Macro int EACCES

Permission denied; the file permissions do not allow the attempted operation.

Macro int EFAULT

Bad address; an invalid pointer was detected. In the GNU system, this error never happens; you get a signal instead.

Macro int ENOTBLK

A file that isn’t a block special file was given in a situation that requires one. For example, trying to mount an ordinary file as a file system in Unix gives this error.

Macro int EBUSY

Resource busy; a system resource that can’t be shared is already in use. For example, if you try to delete a file that is the root of a currently mounted filesystem, you get this error.

Macro int EEXIST

File exists; an existing file was specified in a context where it only makes sense to specify a new file.

Macro int EXDEV

An attempt to make an improper link across file systems was detected. This happens not only when you use link (see Section 9.3 [Hard Links], page 189) but also when you rename a file with rename (see Section 9.6 [Renaming Files], page 193).

Macro int ENODEV

The wrong type of device was given to a function that expects a particular sort of device.

Macro int ENOTDIR

A file that isn’t a directory was specified when a directory is required.

Macro int EISDIR

File is a directory; you cannot open a directory for writing, or create or remove hard links to it.

Macro int EINVAL

Invalid argument. This is used to indicate various kinds of problems with passing the wrong argument to a library function.

Macro int EMFILE

The current process has too many files open and can’t open any more. Duplicate descriptors do count toward this limit. In BSD and GNU, the number of open files is controlled by a resource limit that can usually be increased. If you get this error, you might want to increase the RLIMIT_NOFILE limit or make it unlimited; see Section 17.6 [Limiting Resource Usage], page 362.

Macro int ENFILE

There are too many distinct file openings in the entire system. Note that any number of linked channels count as just one file opening; see Section 8.5.1 [Linked Channels], page 160. This error never occurs in the GNU system.

Macro int ENOTTY

Inappropriate I/O control operation, such as trying to set terminal modes on an ordinary file.

Macro int ETXTBSY

An attempt to execute a file that is currently open for writing, or write to a file that is currently being executed. Often using a debugger to run a program is considered having it open for writing and will cause this error. (The name stands for “text file busy”.) This is not an error in the GNU system; the text is copied as necessary.

Macro int EFBIG

File too big; the size of a file would be larger than allowed by the system.

Macro int ENOSPC

No space left on device; write operation on a file failed because the disk is full.

Macro int ESPIPE

Invalid seek operation (such as on a pipe).

Macro int EROFS

An attempt was made to modify something on a read-only file system.

Macro int EMLINK

Too many links; the link count of a single file would become too large. rename can cause this error if the file being renamed already has as many links as it can take (see Section 9.6 [Renaming Files], page 193).

Macro int EPIPE

Broken pipe; there is no process reading from the other end of a pipe. Every library function that returns this error code also generates a SIGPIPE signal; this signal terminates the program if not handled or blocked. Thus, your program will never actually see EPIPE unless it has handled or blocked SIGPIPE.

Macro int EDOM

Domain error; used by mathematical functions when an argument value does not fall into the domain over which the function is defined.

Macro int ERANGE

Range error; used by mathematical functions when the result value is not representable because of overflow or underflow.

Macro int EAGAIN

Resource temporarily unavailable; the call might work if you try again later. The macro EWOULDBLOCK is another name for EAGAIN; they are always the same in the GNU C library. This error can happen in a few different situations:

• An operation that would block was attempted on an object that has non-blocking mode selected. Trying the same operation again will block until some external condition makes it possible to read, write, or connect (whatever the operation). You can use select to find out when the operation will be possible; see Section 8.6 [Waiting for Input or Output], page 162. Portability Note: In many older Unix systems, this condition was indicated by EWOULDBLOCK, which was a distinct error code different from EAGAIN. To make your program portable, you should check for both codes and treat them the same.

• A temporary resource shortage made an operation impossible. fork can return this error. It indicates that the shortage is expected to pass, so your program can try the call again later and it may succeed. It is probably a good idea to delay for a few seconds before trying it again, to allow time for other processes to release scarce resources. Such shortages are usually fairly serious and affect the whole system, so usually an interactive program should report the error to the user and return to its command loop.

Macro int EWOULDBLOCK

In the GNU C library, this is another name for EAGAIN (above). The values are always the same, on every operating system. C libraries in many older Unix systems have EWOULDBLOCK as a separate error code.

Macro int EINPROGRESS

An operation that cannot complete immediately was initiated on an object that has non-blocking mode selected. Some functions that must always block (such as connect; see Section 11.8.1 [Making a Connection], page 246) never return EAGAIN. Instead, they return EINPROGRESS to indicate that the operation has begun and will take some time. Attempts to manipulate the object before the call completes return EALREADY. You can use the select function to find out when the pending operation has completed; see Section 8.6 [Waiting for Input or Output], page 162.

Macro int EALREADY

An operation is already in progress on an object that has non-blocking mode selected.

Macro int ENOTSOCK

A file that isn’t a socket was specified when a socket is required.

Macro int EMSGSIZE

The size of a message sent on a socket was larger than the supported maximum size.

Macro int EPROTOTYPE

The socket type does not support the requested communications protocol.

Macro int ENOPROTOOPT

You specified a socket option that doesn’t make sense for the particular protocol being used by the socket. See Section 11.11 [Socket Options], page 266.

Macro int EPROTONOSUPPORT

The socket domain does not support the requested communications protocol (perhaps because the requested protocol is completely invalid.) See Section 11.7.1 [Creating a Socket], page 244.

Macro int ESOCKTNOSUPPORT

The socket type is not supported.

Macro int EOPNOTSUPP

The operation you requested is not supported. Some socket functions don’t make sense for all types of sockets, and others may not be implemented for all communications protocols. In the GNU system, this error can happen for many calls when the object does not support the particular operation; it is a generic indication that the server knows nothing to do for that call.

Macro int EPFNOSUPPORT

The socket communications protocol family you requested is not supported.

Macro int EAFNOSUPPORT

The address family specified for a socket is not supported; it is inconsistent with the protocol being used on the socket. See Chapter 11 [Sockets], page 221.

Macro int EADDRINUSE

The requested socket address is already in use. See Section 11.3 [Socket Addresses], page 223.

Macro int EADDRNOTAVAIL

The requested socket address is not available; for example, you tried to give a socket a name that doesn’t match the local host name. See Section 11.3 [Socket Addresses], page 223.

Macro int ENETDOWN

A socket operation failed because the network was down.

Macro int ENETUNREACH

A socket operation failed because the subnet containing the remote host was unreachable.

Macro int ENETRESET

A network connection was reset because the remote host crashed.

Macro int ECONNABORTED

A network connection was aborted locally.

Macro int ECONNRESET

A network connection was closed for reasons outside the control of the local host, such as by the remote machine rebooting or an unrecoverable protocol violation.

Macro int ENOBUFS

The kernel’s buffers for I/O operations are all in use. In GNU, this error is always synonymous with ENOMEM; you may get one or the other from network operations.

Macro int EISCONN

You tried to connect a socket that is already connected. See Section 11.8.1 [Making a Connection], page 246.

Macro int ENOTCONN

The socket is not connected to anything. You get this error when you try to transmit data over a socket, without first specifying a destination for the data. For a connectionless socket (for datagram protocols, such as UDP), you get EDESTADDRREQ instead.

Macro int EDESTADDRREQ

No default destination address was set for the socket. You get this error when you try to transmit data over a connectionless socket, without first specifying a destination for the data with connect.

Macro int ESHUTDOWN

The socket has already been shut down.

Macro int ETOOMANYREFS

???

Macro int ETIMEDOUT

A socket operation with a specified timeout received no response during the timeout period.

Macro int ECONNREFUSED

A remote host refused to allow the network connection (typically because it is not running the requested service).

Macro int ELOOP

Too many levels of symbolic links were encountered in looking up a file name. This often indicates a cycle of symbolic links.

Macro int ENAMETOOLONG

Filename too long (longer than PATH_MAX; see Section 28.6 [Limits on File System Capacity], page 545) or host name too long (in gethostname or sethostname; see Section 27.1 [Host Identification], page 533).

Macro int EHOSTDOWN

The remote host for a requested network connection is down.

Macro int EHOSTUNREACH

The remote host for a requested network connection is not reachable.

Macro int ENOTEMPTY

Directory not empty, where an empty directory was expected. Typically, this error occurs when you are trying to delete a directory.

Macro int EPROCLIM

This means that the per-user limit on new process would be exceeded by an attempted fork. See Section 17.6 [Limiting Resource Usage], page 362, for details on the RLIMIT_NPROC limit.

Macro int EUSERS

The file quota system is confused because there are too many users.

Macro int EDQUOT

The user’s disk quota was exceeded.

Macro int ESTALE

Stale NFS file handle. This indicates an internal confusion in the NFS system which is due to file system rearrangements on the server host. Repairing this condition usually requires unmounting and remounting the NFS file system on the local host.

Macro int EREMOTE

An attempt was made to NFS-mount a remote file system with a file name that already specifies an NFS-mounted file. (This is an error on some operating systems, but we expect it to work properly on the GNU system, making this error code impossible.)

Macro int EBADRPC

???

Macro int ERPCMISMATCH

???

Macro int EPROGUNAVAIL

???

Macro int EPROGMISMATCH

???

Macro int EPROCUNAVAIL

???

Macro int ENOLCK

No locks available. This is used by the file locking facilities; see Section 8.11 [File Locks], page 175. This error is never generated by the GNU system, but it can result from an operation to an NFS server running another operating system.

Macro int EFTYPE

Inappropriate file type or format. The file was the wrong type for the operation, or a data file had the wrong format. On some systems chmod returns this error if you try to set the sticky bit on a non-directory file; see Section 9.8.7 [Assigning File Permissions], page 204.

Macro int EAUTH

???

Macro int ENEEDAUTH

???

Macro int ENOSYS

Function not implemented. Some functions have commands or options defined that might not be supported in all implementations, and this is the kind of error you get if you request them and they are not supported.

Macro int EILSEQ

While decoding a multibyte character the function came along an invalid or an incomplete sequence of bytes or the given wide character is invalid.

Macro int EBACKGROUND

In the GNU system, servers supporting the term protocol return this error for certain operations when the caller is not in the foreground process group of the terminal. Users do not usually see this error because functions such as read and write translate it into a SIGTTIN or SIGTTOU signal. See Chapter 24 [Job Control], page 479, for information on process groups and these signals.

Macro int EDIED

In the GNU system, opening a file returns this error when the file is translated by a program and the translator program dies while starting up, before it has connected to the file.

Macro int ED

The experienced user will know what is wrong.

Macro int EGREGIOUS

You did what?

Macro int EIEIO

Go home and have a glass of warm, dairy-fresh milk.

Macro int EGRATUITOUS

This error code has no purpose.

Macro int EBADMSG Macro int EIDRM Macro int EMULTIHOP Macro int ENODATA Macro int ENOLINK Macro int ENOMSG Macro int ENOSR Macro int ENOSTR Macro int EOVERFLOW Macro int EPROTO Macro int ETIME

The following error codes are defined by the Linux/i386 kernel. They are not yet documented.

Macro int ERESTART Macro int ECHRNG Macro int EL2NSYNC Macro int EL3HLT Macro int EL3RST Macro int ELNRNG Macro int EUNATCH Macro int ENOCSI Macro int EL2HLT Macro int EBADE Macro int EBADR Macro int EXFULL Macro int ENOANO Macro int EBADRQC Macro int EBADSLT Macro int EDEADLOCK Macro int EBFONT Macro int ENONET Macro int ENOPKG Macro int EADV Macro int ESRMNT Macro int ECOMM Macro int EDOTDOT Macro int ENOTUNIQ Macro int EBADFD Macro int EREMCHG Macro int ELIBACC Macro int ELIBBAD Macro int ELIBSCN Macro int ELIBMAX Macro int ELIBEXEC Macro int ESTRPIPE Macro int EUCLEAN Macro int ENOTNAM Macro int ENAVAIL Macro int EISNAM Macro int EREMOTEIO Macro int ENOMEDIUM Macro int EMEDIUMTYPE

2.3 Error Messages

The library has functions and variables designed to make it easy for your program to report informative error messages in the customary format about the failure of a library call. The functions strerror and perror give you the standard error message for a given error code; the variable program_invocation_short_name gives you convenient access to the name of the program that encountered the error.

Function char * strerror (int errnum)

The strerror function maps the error code (see Section 2.1 [Checking for Errors], page 15) specified by the errnum argument to a descriptive error message string. The return value is a pointer to this string. The value errnum normally comes from the variable errno. You should not modify the string returned by strerror. Also, if you make subsequent calls to strerror, the string might be overwritten. (But it’s guaranteed that no library function ever calls strerror behind your back.) The function strerror is declared in ‘string.h’.

Function char * strerror r (int errnum, char *buf, size_t n)

The strerror_r function works like strerror but instead of returning the error message in a statically allocated buffer shared by all threads in the process, it writes the message string in the user supplied buffer starting at buf with the length of n bytes. At most n characters are written (including the NUL byte) so it is up to the user to select the buffer large enough.

This function should always be used in multi-threaded programs since there is no way to guarantee the string returned by strerror really belongs to the last call of the current thread. This function strerror_r is a GNU extension and it is declared in

‘string.h’.

Function void perror (const char *message)

This function prints an error message to the stream stderr; see Section 7.2 [Standard Streams], page 93. If you call perror with a message that is either a null pointer or an empty string, perror just prints the error message corresponding to errno, adding a trailing newline. If you supply a non-null message argument, then perror prefixes its output with this string. It adds a colon and a space character to separate the message from the error string corresponding to errno. The function perror is declared in ‘stdio.h’.

strerror and perror produce the exact same message for any given error code; the precise text varies from system to system. On the GNU system, the messages are fairly short; there are no multi-line messages or embedded newlines. Each error message begins with a capital letter and does not include any terminating punctuation.

Compatibility Note: The strerror function is a new feature of ISO C. Many older C systems do not support this function yet.

Many programs that don’t read input from the terminal are designed to exit if any system call fails. By convention, the error message from such a program should start with the program’s name, sans directories. You can find that name in the variable program_invocation_short_name; the full file name is stored the variable program_invocation_name:

Variable char * program invocation name

This variable’s value is the name that was used to invoke the program running in the current process. It is the same as argv[0]. Note that this is not necessarily a useful file name; often it contains no directory names. See Section 22.1 [Program Arguments], page 447.

Variable char * program invocation short name

This variable’s value is the name that was used to invoke the program running in the current process, with directory names removed. (That is to say, it is the same as program_invocation_name minus everything up to the last slash, if any.)

The library initialization code sets up both of these variables before call-

ing main.

Portability Note: These two variables are GNU extensions. If you want your program to work with non-GNU libraries, you must save the value of argv[0] in main, and then strip off the directory names yourself. We added these extensions to make it possible to write self-contained error-reporting subroutines that require no explicit cooperation from main.

Here is an example showing how to handle failure to open a file correctly. The function open_sesame tries to open the named file for reading and returns a stream if successful. The fopen library function returns a null pointer if it couldn’t open the file for some reason. In that situation, open_sesame constructs an appropriate error message using the strerror function, and terminates the program. If we were going to make some other library calls before passing the error code to strerror, we’d have to save it in a local variable instead, because those other library functions might overwrite errno in the meantime.

     #include <errno.h>
     #include <stdio.h>
     #include <stdlib.h>
     #include <string.h>
     FILE *
     open_sesame (char *name)
     {
       FILE *stream;
       errno = 0;
       stream = fopen (name, "r");
       if (stream == NULL)
         {
           fprintf (stderr, "%s: Couldn't open file %s; %s\n",
                    program_invocation_short_name, name, strerror (errno));
           exit (EXIT_FAILURE);
         }
       else
         return stream;
     }

30 The GNU C Library

3 Memory Allocation

The GNU system provides several methods for allocating memory space under explicit program control. They vary in generality and in efficiency.

• The malloc facility allows fully general dynamic allocation. See Section 3.3 [Unconstrained Allocation], page 32.

• Obstacks are another facility, less general than malloc but more efficient and convenient for stacklike allocation. See Section 3.4 [Obstacks], page 43.

• The function alloca lets you allocate storage dynamically that will be freed automatically. See Section 3.5 [Automatic Storage with Variable Size], page 55.

3.1 Dynamic Memory Allocation Concepts

Dynamic memory allocation is a technique in which programs determine as they are running where to store some information. You need dynamic allocation when the number of memory blocks you need, or how long you continue to need them, depends on the data you are working on.

For example, you may need a block to store a line read from an input file; since there is no limit to how long a line can be, you must allocate the storage dynamically and make it dynamically larger as you read more of the line.

Or, you may need a block for each record or each definition in the input data; since you can’t know in advance how many there will be, you must allocate a new block for each record or definition as you read it.

When you use dynamic allocation, the allocation of a block of memory is an action that the program requests explicitly. You call a function or macro when you want to allocate space, and specify the size with an argument. If you want to free the space, you do so by calling another function or macro. You can do these things whenever you want, as often as you want.

3.2 Dynamic Allocation and C

The C language supports two kinds of memory allocation through the variables in C programs:

• Static allocation is what happens when you declare a static or global variable. Each static or global variable defines one block of space, of a fixed size. The space is allocated once, when your program is started, and is never freed.

• Automatic allocation happens when you declare an automatic variable, such as a function argument or a local variable. The space for an automatic variable is allocated when the compound statement containing the declaration is entered, and is freed when that compound statement is exited. In GNU C, the length of the automatic storage can be an expression that varies. In other C implementations, it must be a constant.

Dynamic allocation is not supported by C variables; there is no storage class “dynamic”, and there can never be a C variable whose value is stored in dynamically allocated space. The only way to refer to dynamically allocated space is through a pointer. Because it is less convenient, and because the actual process of dynamic allocation requires more computation time, programmers generally use dynamic allocation only when neither static nor automatic allocation will serve.

For example, if you want to allocate dynamically some space to hold a struct foobar, you cannot declare a variable of type struct foobar whose contents are the dynamically allocated space. But you can declare a variable of pointer type struct foobar * and assign it the address of the space. Then you can use the operators ‘*’ and ‘->’ on this pointer variable to refer to the contents of the space:

     {
       struct foobar *ptr
          = (struct foobar *) malloc (sizeof (struct foobar));
       ptr->name = x;
       ptr->next = current_foobar;
       current_foobar = ptr;
     }

3.3 Unconstrained Allocation

The most general dynamic allocation facility is malloc. It allows you to allocate blocks of memory of any size at any time, make them bigger or smaller at any time, and free the blocks individually at any time (or never).

3.3.1 Basic Storage Allocation

To allocate a block of memory, call malloc. The prototype for this function is in ‘stdlib.h’.

Function void * malloc (size_t size)

This function returns a pointer to a newly allocated block size bytes long, or a null pointer if the block could not be allocated.

The contents of the block are undefined; you must initialize it yourself (or use calloc instead; see Section 3.3.5 [Allocating Cleared Space], page 36). Normally you would cast the value as a pointer to the kind of object that you want to store in the block. Here we show an example of doing so, and of initializing the space with zeros using the library function memset (see Section 5.4 [Copying and Concatenation], page 67):

     struct foo *ptr;

…

     ptr = (struct foo *) malloc (sizeof (struct foo));
     if (ptr == 0) abort ();
     memset (ptr, 0, sizeof (struct foo));

You can store the result of malloc into any pointer variable without a cast, because ISO C automatically converts the type void * to another type of pointer when necessary. But the cast is necessary in contexts other than assignment operators or if you might want your code to run in traditional C.

Remember that when allocating space for a string, the argument to malloc must be one plus the length of the string. This is because a string is terminated with a null character that doesn’t count in the “length” of the string but does need space. For example:

     char *ptr;

…

     ptr = (char *) malloc (length + 1);

See Section 5.1 [Representation of Strings], page 65, for more information about this.

3.3.2 Examples of malloc

If no more space is available, malloc returns a null pointer. You should check the value of every call to malloc. It is useful to write a subroutine that calls malloc and reports an error if the value is a null pointer, returning only if the value is nonzero. This function is conventionally called xmalloc. Here it is:

     void *
     xmalloc (size_t size)
     {
       register void *value = malloc (size);
       if (value == 0)
         fatal ("virtual memory exhausted");
       return value;
     }

Here is a real example of using malloc (by way of xmalloc). The function savestring will copy a sequence of characters into a newly allocated nullterminated string:

     char *
     savestring (const char *ptr, size_t len)
     {
       register char *value = (char *) xmalloc (len + 1);
       memcpy (value, ptr, len);
       value[len] = '\0';
       return value;
     }

The block that malloc gives you is guaranteed to be aligned so that it can hold any type of data. In the GNU system, the address is always a multiple of eight on most systems, and a multiple of 16 on 64-bit systems. Only rarely is any higher boundary (such as a page boundary) necessary; for those cases, use memalign or valloc (see Section 3.3.7 [Allocating Aligned Memory Blocks], page 37).

Note that the memory located after the end of the block is likely to be in use for something else; perhaps a block already allocated by another call to malloc. If you attempt to treat the block as longer than you asked for it to be, you are liable to destroy the data that malloc uses to keep track of its blocks, or you may destroy the contents of another block. If you have already allocated a block and discover you want it to be bigger, use realloc (see Section 3.3.4 [Changing the Size of a Block], page 35).

3.3.3 Freeing Memory Allocated with malloc

When you no longer need a block that you got with malloc, use the function free to make the block available to be allocated again. The prototype for this function is in ‘stdlib.h’.

Function void free (void *ptr)

The free function deallocates the block of storage pointed at by ptr.

Function void cfree (void *ptr)

This function does the same thing as free. It’s provided for backward compatibility with SunOS; you should use free instead.

Freeing a block alters the contents of the block. Do not expect to find any data (such as a pointer to the next block in a chain of blocks) in the block after freeing it. Copy whatever you need out of the block before freeing it! Here is an example of the proper way to free all the blocks in a chain, and the strings that they point to:

     struct chain
       {
         struct chain *next;
         char *name;
       }

     void
     free_chain (struct chain *chain)
     {
       while (chain != 0)
         {
           struct chain *next = chain->next;
           free (chain->name);
           free (chain);
           chain = next;
         }
     }

Occasionally, free can actually return memory to the operating system and make the process smaller. Usually, all it can do is allow a later call to malloc to reuse the space. In the meantime, the space remains in your program as part of a free-list used internally by malloc.

There is no point in freeing blocks at the end of a program, because all of the program’s space is given back to the system when the process terminates.

3.3.4 Changing the Size of a Block

Often you do not know for certain how big a block you will ultimately need at the time you must begin to use the block. For example, the block might be a buffer that you use to hold a line being read from a file; no matter how long you make the buffer initially, you may encounter a line that is longer.

You can make the block longer by calling realloc. This function is declared in ‘stdlib.h’.

Function void * realloc (void *ptr, size_t newsize)

The realloc function changes the size of the block whose address is ptr to be newsize. Since the space after the end of the block may be in use, realloc may find it necessary to copy the block to a new address where more free space is available. The value of realloc is the new address of the block. If the block needs to be moved, realloc copies the old contents. If you pass a null pointer for ptr, realloc behaves just like ‘malloc (newsize)’. This can be convenient, but beware that older implementations (before ISO C) may not support this behavior, and will probably crash when realloc is passed a null pointer.

Like malloc, realloc may return a null pointer if no memory space is available to make the block bigger. When this happens, the original block is untouched; it has not been modified or relocated.

In most cases it makes no difference what happens to the original block when realloc fails, because the application program cannot continue when it is out of memory, and the only thing to do is to give a fatal error message. Often it is convenient to write and use a subroutine, conventionally called xrealloc, that takes care of the error message as xmalloc does for malloc:

     void *
     xrealloc (void *ptr, size_t size)
     {
       register void *value = realloc (ptr, size);
       if (value == 0)
         fatal ("Virtual memory exhausted");
       return value;
     }

You can also use realloc to make a block smaller. The reason you is needed. In several allocation implementations, making a block smaller sometimes necessitates copying it, so it can fail if no other space is available.

If the new size you specify is the same as the old size, realloc is guaranteed to change nothing and return the same address that you gave.

3.3.5 Allocating Cleared Space

The function calloc allocates memory and clears it to zero. It is declared

in ‘stdlib.h’.

Function void * calloc (size_t count, size_t eltsize)

This function allocates a block long enough to contain a vector of count elements, each of size eltsize. Its contents are cleared to zero before calloc returns.

You could define calloc as follows:

     void *
     calloc (size_t count, size_t eltsize)
     {
       size_t size = count * eltsize;
       void *value = malloc (size);
       if (value != 0)
         memset (value, 0, size);
       return value;
     }

But in general, it is not guaranteed that calloc calls malloc internally. Therefore, if an application provides its own malloc/realloc/free outside the C library, it should always define calloc, too.

3.3.6 Efficiency Considerations for malloc

As apposed to other versions, the malloc in GNU libc does not round up block sizes to powers of two, neither for large nor for small sizes. Neighboring chunks can be coalesced on a free no matter what their size is. This makes the implementation suitable for all kinds of allocation patterns without generally incurring high memory waste through fragmentation.

Very large blocks (much larger than a page) are allocated with mmap (anonymous or via /dev/zero) by this implementation. This has the great advantage that these chunks are returned to the system immediately when they are freed. Therefore, it cannot happen that a large chunk becomes “locked” in between smaller ones and even after calling free wastes memory. The size threshold for mmap to be used can be adjusted with mallopt. The use of mmap can also be disabled completely.

3.3.7 Allocating Aligned Memory Blocks

The address of a block returned by malloc or realloc in the GNU system is always a multiple of eight (or sixteen on 64-bit systems). If you need a block whose address is a multiple of a higher power of two than that, use memalign or valloc. These functions are declared in ‘stdlib.h’.

With the GNU library, you can use free to free the blocks that memalign and valloc return. That does not work in BSD, however—BSD does not provide any way to free such blocks.

Function void * memalign (size_t boundary, size_t size)

The memalign function allocates a block of size bytes whose address is a multiple of boundary. The boundary must be a power of two! The function memalign works by allocating a somewhat larger block, and then returning an address within the block that is on the specified boundary.

Function void * valloc (size_t size)

Using valloc is like using memalign and passing the page size as the value of the second argument. It is implemented like this:

        void *
        valloc (size_t size)
        {
          return memalign (getpagesize (), size);
        }

3.3.8 Malloc Tunable Parameters

You can adjust some parameters for dynamic memory allocation with the mallopt function. This function is the general SVID/XPG interface, defined

in ‘malloc.h’.

Function int mallopt (int param, int value)

When calling mallopt, the param argument specifies the parameter to be set, and value the new value to be set. Possible choices for param, as defined in ‘malloc.h’, are:

M_TRIM_THRESHOLD

This is the minimum size (in bytes) of the top-most, releaseable chunk that will cause sbrk to be called with a negative argument in order to return memory to the system.

M_TOP_PAD

This parameter determines the amount of extra memory to obtain from the system when a call to sbrk is required. It also specifies the number of bytes to retain when shrinking the heap by calling sbrk with a negative argument. This provides the necessary hysteresis in heap size such that excessive amounts of system calls can be avoided.

M_MMAP_THRESHOLD

All chunks larger than this value are allocated outside the normal heap, using the mmap system call. This way it is guaranteed that the memory for these chunks can be returned to the system on free.

M_MMAP_MAX

The maximum number of chunks to allocate with mmap. Setting this to zero disables all use of mmap.

3.3.9 Heap Consistency Checking

You can ask malloc to check the consistency of dynamic storage by using the mcheck function. This function is a GNU extension, declared in

‘malloc.h’.
Function int mcheck (void (*abortfn) (enum mcheck_status

status)) Calling mcheck tells malloc to perform occasional consistency checks. These will catch things such as writing past the end of a block that was allocated with malloc. The abortfn argument is the function to call when an inconsistency is found. If you supply a null pointer, then mcheck uses a default function which prints a message and calls abort (see Section 22.3.4 [Aborting a Program], page 465). The function you supply is called with one argument, which says what sort of inconsistency was detected; its type is described below. It is too late to begin allocation checking once you have allocated anything with malloc. So mcheck does nothing in that case. The function returns -1 if you call it too late, and 0 otherwise (when it is successful). The easiest way to arrange to call mcheck early enough is to use the option ‘-lmcheck’ when you link your program; then you don’t need to modify your program source at all.

Function enum mcheck_status mprobe (void *pointer)

The mprobe function lets you explicitly check for inconsistencies in a particular allocated block. You must have already called mcheck at the beginning of the program, to do its occasional checks; calling mprobe requests an additional consistency check to be done at the time of the call. The argument pointer must be a pointer returned by malloc or realloc. mprobe returns a value that says what inconsistency, if any, was found. The values are described below.

Data Type enum mcheck status

This enumerated type describes what kind of inconsistency was detected in an allocated block, if any. Here are the possible values:

MCHECK_DISABLED

mcheck was not called before the first allocation. No consistency checking can be done.

MCHECK_OK

No inconsistency detected.

MCHECK_HEAD

The data immediately before the block was modified. This commonly happens when an array index or pointer is decremented too far.

MCHECK_TAIL

The data immediately after the block was modified. This commonly happens when an array index or pointer is incremented too far.

MCHECK_FREE

The block was already freed.

3.3.10 Storage Allocation Hooks

The GNU C library lets you modify the behavior of malloc, realloc, and free by specifying appropriate hook functions. You can use these hooks to help you debug programs that use dynamic storage allocation, for example.

The hook variables are declared in ‘malloc.h’.

Variable malloc hook

The value of this variable is a pointer to function that malloc uses whenever it is called. You should define this function to look like malloc; that is, like:

     void *function (size_t size)

Variable realloc hook

The value of this variable is a pointer to function that realloc uses whenever it is called. You should define this function to look like realloc; that is, like:

     void *function (void *ptr, size_t size)

Variable free hook

The value of this variable is a pointer to function that free uses whenever it is called. You should define this function to look like free; that is, like:

     void function (void *ptr)

You must make sure that the function you install as a hook for one of these functions does not call that function recursively without restoring the old value of the hook first! Otherwise, your program will get stuck in an infinite recursion.

Here is an example showing how to use __malloc_hook properly. It installs a function that prints out information every time malloc is called.

  static void *(*old_malloc_hook) (size_t);
  static void *
  my_malloc_hook (size_t size)
  {
    void *result;
    __malloc_hook = old_malloc_hook;
    result = malloc (size);

/* printf might call malloc, so protect it too. */

    printf ("malloc (%u) returns %p\n", (unsigned int) size, result);
    __malloc_hook = my_malloc_hook;
    return result;
  }
  main ()

{
...
old_malloc_hook = __malloc_hook;
__malloc_hook = my_malloc_hook;
...
}

The mcheck function (see Section 3.3.9 [Heap Consistency Checking], page 38) works by installing such hooks.

3.3.11 Statistics for Storage Allocation with malloc

You can get information about dynamic storage allocation by calling the mallinfo function. This function and its associated data type are declared in ‘malloc.h’; they are an extension of the standard SVID/XPG version.

Data Type struct mallinfo

This structure type is used to return information about the dynamic storage allocator. It contains the following members: int arena This is the total size of memory allocated with sbrk by

malloc, in bytes.

int ordblks

This is the number of chunks not in use. (The storage allocator internally gets chunks of memory from the operating system, and then carves them up to satisfy individual malloc requests; see Section 3.3.6 [Efficiency Considerations for “codemalloc], page 37.)

int smblks

This field is unused.

int hblks This is the total number of chunks allocated with mmap.

int hblkhd

This is the total size of memory allocated with mmap, in bytes.

int usmblks

This field is unused.

int fsmblks

This field is unused.

int uordblks

This is the total size of memory occupied by chunks handed out by malloc.

int fordblks

This is the total size of memory occupied by free (not in use) chunks.

int keepcost

This is the size of the top-most, releaseable chunk that normally borders the end of the heap (i.e. the “brk” of the process).

Function struct mallinfo mallinfo (void)

This function returns information about the current dynamic memory usage in a structure of type struct mallinfo.

3.3.12 Summary of malloc-Related Functions

Here is a summary of the functions that work with malloc:

void *malloc (size_t size)

Allocate a block of size bytes. See Section 3.3.1 [Basic Storage Allocation], page 32.

void free (void *addr)

Free a block previously allocated by malloc. See Section 3.3.3 [Freeing Memory Allocated with “codemalloc], page 34.

void *realloc (void *addr, size_t size)

Make a block previously allocated by malloc larger or smaller, possibly by copying it to a new location. See Section 3.3.4 [Changing the Size of a Block], page 35.

void *calloc (size_t count, size_t eltsize)

Allocate a block of count * eltsize bytes using malloc, and set its contents to zero. See Section 3.3.5 [Allocating Cleared Space], page 36.

void *valloc (size_t size)

Allocate a block of size bytes, starting on a page boundary. See Section 3.3.7 [Allocating Aligned Memory Blocks], page 37.

void *memalign (size_t size, size_t boundary)

Allocate a block of size bytes, starting on an address that is a multiple of boundary. See Section 3.3.7 [Allocating Aligned Memory Blocks], page 37.

int mallopt (int param, int value)

Adjust a tunable parameter. See Section 3.3.8 [Malloc Tunable Parameters], page 38

int mcheck (void (*abortfn) (void))

Tell malloc to perform occasional consistency checks on dynamically allocated memory, and to call abortfn when an inconsistency is found. See Section 3.3.9 [Heap Consistency Checking], page 38.

void *(*__malloc_hook) (size_t size)

A pointer to a function that malloc uses whenever it is called.

void *(*__realloc_hook) (void *ptr, size_t size)

A pointer to a function that realloc uses whenever it is called.

void (*__free_hook) (void *ptr)

A pointer to a function that free uses whenever it is called.

struct mallinfo mallinfo (void)

Return information about the current dynamic memory usage. See Section 3.3.11 [Statistics for Storage Allocation with “codemalloc], page 41.

3.4 Obstacks

An obstack is a pool of memory containing a stack of objects. You can create any number of separate obstacks, and then allocate objects in specified obstacks. Within each obstack, the last object allocated must always be the first one freed, but distinct obstacks are independent of each other.

Aside from this one constraint of order of freeing, obstacks are totally general: an obstack can contain any number of objects of any size. They are implemented with macros, so allocation is usually very fast as long as the objects are usually small. And the only space overhead per object is the padding needed to start each object on a suitable boundary.

3.4.1 Creating Obstacks

The utilities for manipulating obstacks are declared in the header file

‘obstack.h’.

Data Type struct obstack

An obstack is represented by a data structure of type struct obstack. This structure has a small fixed size; it records the status of the obstack and how to find the space in which objects are allocated. It does not contain any of the objects themselves. You should not try to access the contents of the structure directly; use only the functions described in this chapter.

You can declare variables of type struct obstack and use them as obstacks, or you can allocate obstacks dynamically like any other kind of object. Dynamic allocation of obstacks allows your program to have a variable number of different stacks. (You can even allocate an obstack structure in another obstack, but this is rarely useful.)

All the functions that work with obstacks require you to specify which obstack to use. You do this with a pointer of type struct obstack *. In the following, we often say “an obstack” when strictly speaking the object at hand is such a pointer.

The objects in the obstack are packed into large blocks called chunks. The struct obstack structure points to a chain of the chunks currently in use.

The obstack library obtains a new chunk whenever you allocate an object that won’t fit in the previous chunk. Since the obstack library manages chunks automatically, you don’t need to pay much attention to them, but you do need to supply a function which the obstack library should use to get a chunk. Usually you supply a function which uses malloc directly or indirectly. You must also supply a function to free a chunk. These matters are described in the following section.

3.4.2 Preparing for Using Obstacks

Each source file in which you plan to use the obstack functions must include the header file ‘obstack.h’, like this:

     #include <obstack.h>

Also, if the source file uses the macro obstack_init, it must declare or define two functions or macros that will be called by the obstack library. One, obstack_chunk_alloc, is used to allocate the chunks of memory into which objects are packed. The other, obstack_chunk_free, is used to return chunks when the objects in them are freed. These macros should appear before any use of obstacks in the source file.

Usually these are defined to use malloc via the intermediary xmalloc (see Section 3.3 [Unconstrained Allocation], page 32). This is done with the following pair of macro definitions:

     #define obstack_chunk_alloc xmalloc
     #define obstack_chunk_free free

Though the storage you get using obstacks really comes from malloc, using obstacks is faster because malloc is called less often, for larger blocks of memory. See Section 3.4.10 [Obstack Chunks], page 52, for full details.

At run time, before the program can use a struct obstack object as an obstack, it must initialize the obstack by calling obstack_init.

Function int obstack init (struct obstack *obstack-ptr)

Initialize obstack obstack-ptr for allocation of objects. This function calls the obstack’s obstack_chunk_alloc function. It returns 0 if obstack_chunk_alloc returns a null pointer, meaning that it is out of memory. Otherwise, it returns 1. If you supply an obstack_ chunk_alloc function that calls exit (see Section 22.3 [Program Termination], page 462) or longjmp (see Chapter 20 [Non-Local Exits], page 389) when out of memory, you can safely ignore the value that

  obstack_init returns.

Here are two examples of how to allocate the space for an obstack and initialize it. First, an obstack that is a static variable:

  static struct obstack myobstack;

…

  obstack_init (&myobstack);

Second, an obstack that is itself dynamically allocated:

  struct obstack *myobstack_ptr
    = (struct obstack *) xmalloc (sizeof (struct obstack));
  obstack_init (myobstack_ptr);

3.4.3 Allocation in an Obstack

The most direct way to allocate an object in an obstack is with obstack_ alloc, which is invoked almost like malloc.

Function void * obstack alloc (struct obstack *obstack-ptr,

int size) This allocates an uninitialized block of size bytes in an obstack and returns its address. Here obstack-ptr specifies which obstack to allocate the block in; it is the address of the struct obstack object which represents the obstack. Each obstack function or macro requires you to specify an obstack-ptr as the first argument. This function calls the obstack’s obstack_chunk_alloc function if it needs to allocate a new chunk of memory; it returns a null pointer if obstack_chunk_alloc returns one. In that case, it has not changed the amount of memory allocated in the obstack. If you supply an obstack_chunk_alloc function that calls exit (see Section 22.3 [Program Termination], page 462) or longjmp (see Chapter 20 [Non-Local Exits], page 389) when out of memory, then obstack_alloc will never return a null pointer.

For example, here is a function that allocates a copy of a string str in a specific obstack, which is in the variable string_obstack:

  struct obstack string_obstack;
  char *
  copystring (char *string)
  {
    size_t len = strlen (string) + 1;
    char *s = (char *) obstack_alloc (&string_obstack, len);
    memcpy (s, string, len);
    return s;
  }

To allocate a block with specified contents, use the function obstack_ copy, declared like this:

Function void * obstack copy (struct obstack *obstack-ptr,

void *address, int size) This allocates a block and initializes it by copying size bytes of data starting at address. It can return a null pointer under the same con-

  ditions as obstack_alloc.

Function void * obstack copy0 (struct obstack *obstack-ptr,

void *address, int size) Like obstack_copy, but appends an extra byte containing a null character. This extra byte is not counted in the argument size.

The obstack_copy0 function is convenient for copying a sequence of characters into an obstack as a null-terminated string. Here is an example of its use:

     char *
     obstack_savestring (char *addr, int size)
     {
       return obstack_copy0 (&myobstack, addr, size);
     }

Contrast this with the previous example of savestring using malloc (see Section 3.3.1 [Basic Storage Allocation], page 32).

3.4.4 Freeing Objects in an Obstack

To free an object allocated in an obstack, use the function obstack_free. Since the obstack is a stack of objects, freeing one object automatically frees all other objects allocated more recently in the same obstack.

Function void obstack free (struct obstack *obstack-ptr, void

*object) If object is a null pointer, everything allocated in the obstack is freed. Otherwise, object must be the address of an object allocated in the obstack. Then object is freed, along with everything allocated in obstack since object.

Note that if object is a null pointer, the result is an uninitialized obstack. To free all storage in an obstack but leave it valid for further allocation, call obstack_free with the address of the first object allocated on the obstack:

     obstack_free (obstack_ptr, first_object_allocated_ptr);

Recall that the objects in an obstack are grouped into chunks. When all the objects in a chunk become free, the obstack library automatically frees the chunk (see Section 3.4.2 [Preparing for Using Obstacks], page 44). Then other obstacks, or non-obstack allocation, can reuse the space of the chunk.

3.4.5 Obstack Functions and Macros

The interfaces for using obstacks may be defined either as functions or as macros, depending on the compiler. The obstack facility works with all C compilers, including both ISO C and traditional C, but there are precautions you must take if you plan to use compilers other than GNU C.

If you are using an old-fashioned non-ISO C compiler, all the obstack “functions” are actually defined only as macros. You can call these macros like functions, but you cannot use them in any other way (for example, you cannot take their address).

Calling the macros requires a special precaution: namely, the first operand (the obstack pointer) may not contain any side effects, because it may be computed more than once. For example, if you write this:

     obstack_alloc (get_obstack (), 4);

you will find that get_obstack may be called several times. If you use *obstack_list_ptr++ as the obstack pointer argument, you will get very strange results since the incrementation may occur several times.

In ISO C, each function has both a macro definition and a function definition. The function definition is used if you take the address of the function without calling it. An ordinary call uses the macro definition by default, but you can request the function definition instead by writing the function name in parentheses, as shown here:

     char *x;
     void *(*funcp) ();

/* Use the macro. */

     x = (char *) obstack_alloc (obptr, size);

/* Call the function. */

     x = (char *) (obstack_alloc) (obptr, size);

/* Take the address of the function. */

     funcp = obstack_alloc;

This is the same situation that exists in ISO C for the standard library functions. See Section 1.3.2 [Macro Definitions of Functions], page 5.

Warning: When you do use the macros, you must observe the precaution of avoiding side effects in the first operand, even in ISO C.

If you use the GNU C compiler, this precaution is not necessary, because various language extensions in GNU C permit defining the macros so as to compute each argument only once.

3.4.6 Growing Objects

Because storage in obstack chunks is used sequentially, it is possible to build up an object step by step, adding one or more bytes at a time to the end of the object. With this technique, you do not need to know how much data you will put in the object until you come to the end of it. We call this the technique of growing objects. The special functions for adding data to the growing object are described in this section.

You don’t need to do anything special when you start to grow an object. Using one of the functions to add data to the object automatically starts it. However, it is necessary to say explicitly when the object is finished. This is done with the function obstack_finish.

The actual address of the object thus built up is not known until the object is finished. Until then, it always remains possible that you will add so much data that the object must be copied into a new chunk.

While the obstack is in use for a growing object, you cannot use it for ordinary allocation of another object. If you try to do so, the space already added to the growing object will become part of the other object.

Function void obstack blank (struct obstack *obstack-ptr,

int size) The most basic function for adding to a growing object is obstack_ blank, which adds space without initializing it.

Function void obstack grow (struct obstack *obstack-ptr,

         void *data, int size)

To add a block of initialized space, use obstack_grow, which is the growing-object analogue of obstack_copy. It adds size bytes of data to the growing object, copying the contents from data.

Function void obstack grow0 (struct obstack *obstack-ptr,

         void *data, int size)

This is the growing-object analogue of obstack_copy0. It adds size bytes copied from data, followed by an additional null character.

Function void obstack 1grow (struct obstack *obstack-ptr,

         char c)

To add one character at a time, use the function obstack_1grow. It adds a single byte containing c to the growing object.

Function void obstack ptr grow (struct obstack

*obstack-ptr, void *data) Adding the value of a pointer one can use the function obstack_ptr_ grow. It adds sizeof (void *) bytes containing the value of data.

Function void obstack int grow (struct obstack

*obstack-ptr, int data) A single value of type int can be added by using the obstack_int_ grow function. It adds sizeof (int) bytes to the growing object and initializes them with the value of data.

Function void * obstack finish (struct obstack *obstack-ptr)

When you are finished growing the object, use the function obstack_ finish to close it off and return its final address. Once you have finished the object, the obstack is available for ordinary allocation or for growing another object. This function can return a null pointer under the same conditions as obstack_alloc (see Section 3.4.3 [Allocation in an Obstack], page 45).

When you build an object by growing it, you will probably need to know afterward how long it became. You need not keep track of this as you grow the object, because you can find out the length from the obstack just before finishing the object with the function obstack_object_size, declared as follows:

Function int obstack object size (struct obstack

*obstack-ptr) This function returns the current size of the growing object, in bytes. Remember to call this function before finishing the object. After it is finished, obstack_object_size will return zero.

If you have started growing an object and wish to cancel it, you should finish it and then free it, like this:

     obstack_free (obstack_ptr, obstack_finish (obstack_ptr));

This has no effect if no object was growing.

You can use obstack_blank with a negative size argument to make the current object smaller. Just don’t try to shrink it beyond zero length— there’s no telling what will happen if you do that.

3.4.7 Extra Fast Growing Objects

The usual functions for growing objects incur overhead for checking whether there is room for the new growth in the current chunk. If you are frequently constructing objects in small steps of growth, this overhead can be significant.

You can reduce the overhead by using special “fast growth” functions that grow the object without checking. In order to have a robust program, you must do the checking yourself. If you do this checking in the simplest way each time you are about to add data to the object, you have not saved anything, because that is what the ordinary growth functions do. But if you can arrange to check less often, or check more efficiently, then you make the program faster.

The function obstack_room returns the amount of room available in the current chunk. It is declared as follows:

Function int obstack room (struct obstack *obstack-ptr)

This returns the number of bytes that can be added safely to the current growing object (or to an object about to be started) in obstack obstack using the fast growth functions.

While you know there is room, you can use these fast growth functions for adding data to a growing object:

Function void obstack 1grow fast (struct obstack

*obstack-ptr, char c) The function obstack_1grow_fast adds one byte containing the character c to the growing object in obstack obstack-ptr.

Function void obstack ptr grow fast (struct obstack

*obstack-ptr, void *data)

The function obstack_ptr_grow_fast adds sizeof (void *) bytes

containing the value of data to the growing object in obstack obstackptr.

Function void obstack int grow fast (struct obstack

*obstack-ptr, int data)

The function obstack_int_grow_fast adds sizeof (int) bytes con-

taining the value of data to the growing object in obstack obstack-ptr.

Function void obstack blank fast (struct obstack

*obstack-ptr, int size) The function obstack_blank_fast adds size bytes to the growing object in obstack obstack-ptr without initializing them.

When you check for space using obstack_room and there is not enough room for what you want to add, the fast growth functions are not safe. In this case, simply use the corresponding ordinary growth function instead. Very soon this will copy the object to a new chunk; then there will be lots of room available again.

So, each time you use an ordinary growth function, check afterward for sufficient space using obstack_room. Once the object is copied to a new chunk, there will be plenty of space again, so the program will start using the fast growth functions again.

Here is an example:

  void
  add_string (struct obstack *obstack, const char *ptr, int len)
  {
    while (len > 0)
      {
        int room = obstack_room (obstack);
        if (room == 0)
          {

/* Not enough room. Add one character slowly,

which may copy to a new chunk and make room. */

            obstack_1grow (obstack, *ptr++);
            len--;
          }
        else
          {
            if (room > len)
              room = len;

/* Add fast as much as we have room for. */

            len -= room;
            while (room-- > 0)
              obstack_1grow_fast (obstack, *ptr++);
          }
      }
  }

3.4.8 Status of an Obstack

Here are functions that provide information on the current status of allocation in an obstack. You can use them to learn about an object while still growing it.

Function void * obstack base (struct obstack *obstack-ptr)

This function returns the tentative address of the beginning of the currently growing object in obstack-ptr. If you finish the object immediately, it will have that address. If you make it larger first, it may outgrow the current chunk—then its address will change! If no object is growing, this value says where the next object you allocate will start (once again assuming it fits in the current chunk).

Function void * obstack next free (struct obstack

*obstack-ptr) This function returns the address of the first free byte in the current chunk of obstack obstack-ptr. This is the end of the currently growing object. If no object is growing, obstack_next_free returns the same

value as obstack_base.

Function int obstack object size (struct obstack

*obstack-ptr) This function returns the size in bytes of the currently growing object. This is equivalent to

        obstack_next_free (obstack-ptr) - obstack_base (obstack-ptr)

3.4.9 Alignment of Data in Obstacks

Each obstack has an alignment boundary; each object allocated in the obstack automatically starts on an address that is a multiple of the specified boundary. By default, this boundary is 4 bytes.

To access an obstack’s alignment boundary, use the macro obstack_ alignment_mask, whose function prototype looks like this:

Macro int obstack alignment mask (struct obstack

*obstack-ptr) The value is a bit mask; a bit that is 1 indicates that the corresponding bit in the address of an object should be 0. The mask value should be one less than a power of 2; the effect is that all object addresses are multiples of that power of 2. The default value of the mask is 3, so that addresses are multiples of 4. A mask value of 0 means an object can start on any multiple of 1 (that is, no alignment is required). The expansion of the macro obstack_alignment_mask is an lvalue, so you can alter the mask by assignment. For example, this statement:

        obstack_alignment_mask (obstack_ptr) = 0;

has the effect of turning off alignment processing in the specified obstack.

Note that a change in alignment mask does not take effect until after the next time an object is allocated or finished in the obstack. If you are not growing an object, you can make the new alignment mask take effect immediately by calling obstack_finish. This will finish a zero-length object and then do proper alignment for the next object.

3.4.10 Obstack Chunks

Obstacks work by allocating space for themselves in large chunks, and then parceling out space in the chunks to satisfy your requests. Chunks are normally 4096 bytes long unless you specify a different chunk size. The chunk size includes 8 bytes of overhead that are not actually used for storing objects. Regardless of the specified size, longer chunks will be allocated when necessary for long objects.

The obstack library allocates chunks by calling the function obstack_ chunk_alloc, which you must define. When a chunk is no longer needed because you have freed all the objects in it, the obstack library frees the chunk by calling obstack_chunk_free, which you must also define.

These two must be defined (as macros) or declared (as functions) in each source file that uses obstack_init (see Section 3.4.1 [Creating Obstacks], page 43). Most often they are defined as macros like this:

     #define obstack_chunk_alloc xmalloc
     #define obstack_chunk_free free

Note that these are simple macros (no arguments). Macro definitions with arguments will not work! It is necessary that obstack_chunk_alloc or obstack_chunk_free, alone, expand into a function name if it is not itself a function name.

If you allocate chunks with malloc, the chunk size should be a power of 2. The default chunk size, 4096, was chosen because it is long enough to satisfy many typical requests on the obstack yet short enough not to waste too much memory in the portion of the last chunk not yet used.

Macro int obstack chunk size (struct obstack *obstack-ptr)

This returns the chunk size of the given obstack.

Since this macro expands to an lvalue, you can specify a new chunk size by assigning it a new value. Doing so does not affect the chunks already allocated, but will change the size of chunks allocated for that particular obstack in the future. It is unlikely to be useful to make the chunk size smaller, but making it larger might improve efficiency if you are allocating many objects whose size is comparable to the chunk size. Here is how to do so cleanly:

     if (obstack_chunk_size (obstack_ptr) < new-chunk-size)
       obstack_chunk_size (obstack_ptr) = new-chunk-size;

3.4.11 Summary of Obstack Functions

Here is a summary of all the functions associated with obstacks. Each takes the address of an obstack (struct obstack *) as its first argument.

void obstack_init (struct obstack *obstack-ptr)

Initialize use of an obstack. See Section 3.4.1 [Creating Obstacks], page 43.

void *obstack_alloc (struct obstack *obstack-ptr, int size)

Allocate an object of size uninitialized bytes. See Section 3.4.3 [Allocation in an Obstack], page 45.

void *obstack_copy (struct obstack *obstack-ptr, void *address, int

size)

Allocate an object of size bytes, with contents copied from address. See Section 3.4.3 [Allocation in an Obstack], page 45.

void *obstack_copy0 (struct obstack *obstack-ptr, void *address,
int size)

Allocate an object of size+1 bytes, with size of them copied from address, followed by a null character at the end. See Section 3.4.3 [Allocation in an Obstack], page 45.

void obstack_free (struct obstack *obstack-ptr, void *object)

Free object (and everything allocated in the specified obstack more recently than object). See Section 3.4.4 [Freeing Objects in an Obstack], page 46.

void obstack_blank (struct obstack *obstack-ptr, int size)

Add size uninitialized bytes to a growing object. See Section 3.4.6 [Growing Objects], page 47.

void obstack_grow (struct obstack *obstack-ptr, void *address, int

size)

Add size bytes, copied from address, to a growing object. See Section 3.4.6 [Growing Objects], page 47.

void obstack_grow0 (struct obstack *obstack-ptr, void *address, int

size)

Add size bytes, copied from address, to a growing object, and then add another byte containing a null character. See Section 3.4.6 [Growing Objects], page 47.

void obstack_1grow (struct obstack *obstack-ptr, char data-char)

Add one byte containing data-char to a growing object. See Section 3.4.6 [Growing Objects], page 47.

void *obstack_finish (struct obstack *obstack-ptr)

Finalize the object that is growing and return its permanent address. See Section 3.4.6 [Growing Objects], page 47.

int obstack_object_size (struct obstack *obstack-ptr)

Get the current size of the currently growing object. See Section 3.4.6 [Growing Objects], page 47.

void obstack_blank_fast (struct obstack *obstack-ptr, int size)

Add size uninitialized bytes to a growing object without checking that there is enough room. See Section 3.4.7 [Extra Fast Growing Objects], page 49.

void obstack_1grow_fast (struct obstack *obstack-ptr, char

data-char)

Add one byte containing data-char to a growing object without checking that there is enough room. See Section 3.4.7 [Extra Fast Growing Objects], page 49.

int obstack_room (struct obstack *obstack-ptr)

Get the amount of room now available for growing the current object. See Section 3.4.7 [Extra Fast Growing Objects], page 49.

int obstack_alignment_mask (struct obstack *obstack-ptr)

The mask used for aligning the beginning of an object. This is an lvalue. See Section 3.4.9 [Alignment of Data in Obstacks], page 52.

int obstack_chunk_size (struct obstack *obstack-ptr)

The size for allocating chunks. This is an lvalue. See Section 3.4.10 [Obstack Chunks], page 52.

void *obstack_base (struct obstack *obstack-ptr)

Tentative starting address of the currently growing object. See Section 3.4.8 [Status of an Obstack], page 51.

void *obstack_next_free (struct obstack *obstack-ptr)

Address just after the end of the currently growing object. See Section 3.4.8 [Status of an Obstack], page 51.

3.5 Automatic Storage with Variable Size

The function alloca supports a kind of half-dynamic allocation in which blocks are allocated dynamically but freed automatically.

Allocating a block with alloca is an explicit action; you can allocate as many blocks as you wish, and compute the size at run time. But all the blocks are freed when you exit the function that alloca was called from, just as if they were automatic variables declared in that function. There is no way to free the space explicitly.

The prototype for alloca is in ‘stdlib.h’. This function is a BSD extension.

Function void * alloca (size_t size); The return value of alloca is the address of a block of size bytes of storage, allocated in the stack frame of the calling function.

Do not use alloca inside the arguments of a function call—you will get unpredictable results, because the stack space for the alloca would appear on the stack in the middle of the space for the function arguments. An example of what to avoid is foo (x, alloca (4), y).

3.5.1 alloca Example

As an example of use of alloca, here is a function that opens a file name made from concatenating two argument strings, and returns a file descriptor or minus one signifying failure:

int
open2 (char *str1, char *str2, int flags, int mode)
{
  char *name = (char *) alloca (strlen (str1) + strlen (str2) + 1);
  stpcpy (stpcpy (name, str1), str2);
  return open (name, flags, mode);
}

Here is how you would get the same results with malloc and free:

int
open2 (char *str1, char *str2, int flags, int mode)
{
  char *name = (char *) malloc (strlen (str1) + strlen (str2) + 1);
  int desc;
  if (name == 0)
    fatal ("virtual memory exceeded");
  stpcpy (stpcpy (name, str1), str2);
  desc = open (name, flags, mode);
  free (name);
  return desc;
}

As you can see, it is simpler with alloca. But alloca has other, more important advantages, and some disadvantages.

3.5.2 Advantages of alloca

Here are the reasons why alloca may be preferable to malloc:

• Using alloca wastes very little space and is very fast. (It is open-coded by the GNU C compiler.)

• Since alloca does not have separate pools for different sizes of block, space used for any size block can be reused for any other size. alloca does not cause storage fragmentation.

• Nonlocal exits done with longjmp (see Chapter 20 [Non-Local Exits], page 389) automatically free the space allocated with alloca when they exit through the function that called alloca. This is the most important reason to use alloca. To illustrate this, suppose you have a function open_or_report_error which returns a descriptor, like open, if it succeeds, but does not return to its caller if it fails. If the file cannot be opened, it prints an error message and jumps out to the command level of your program using longjmp. Let’s change open2 (see Section 3.5.1 [“codealloca Example], page 55) to use this subroutine:

     int
     open2 (char *str1, char *str2, int flags, int mode)

     {
       char *name = (char *) alloca (strlen (str1) + strlen (str2) + 1);
       stpcpy (stpcpy (name, str1), str2);
       return open_or_report_error (name, flags, mode);
     }

Because of the way alloca works, the storage it allocates is freed even when an error occurs, with no special effort required. By contrast, the previous definition of open2 (which uses malloc and free) would develop a storage leak if it were changed in this way. Even if you are willing to make more changes to fix it, there is no easy way to do so.

3.5.3 Disadvantages of alloca

These are the disadvantages of alloca in comparison with malloc:

• If you try to allocate more storage than the machine can provide, you don’t get a clean error message. Instead you get a fatal signal like the one you would get from an infinite recursion; probably a segmentation violation (see Section 21.2.1 [Program Error Signals], page 396).

• Some non-GNU systems fail to support alloca, so it is less portable. However, a slower emulation of alloca written in C is available for use on systems with this deficiency.

3.5.4 GNU C Variable-Size Arrays

In GNU C, you can replace most uses of alloca with an array of variable size. Here is how open2 would look then:

int open2 (char *str1, char *str2, int flags, int mode)
{
  char name[strlen (str1) + strlen (str2) + 1];
  stpcpy (stpcpy (name, str1), str2);
  return open (name, flags, mode);
}

But alloca is not always equivalent to a variable-sized array, for several reasons:

• A variable size array’s space is freed at the end of the scope of the name of the array. The space allocated with alloca remains until the end of the function.

• It is possible to use alloca within a loop, allocating an additional block on each iteration. This is impossible with variable-sized arrays. Note: If you mix use of alloca and variable-sized arrays within one

function, exiting a scope in which a variable-sized array was declared frees all blocks allocated with alloca during the execution of that scope.

3.6 Relocating Allocator

Any system of dynamic memory allocation has overhead: the amount of space it uses is more than the amount the program asks for. The relocating memory allocator achieves very low overhead by moving blocks in memory as necessary, on its own initiative.

3.6.1 Concepts of Relocating Allocation

When you allocate a block with malloc, the address of the block never changes unless you use realloc to change its size. Thus, you can safely store the address in various places, temporarily or permanently, as you like. This is not safe when you use the relocating memory allocator, because any and all relocatable blocks can move whenever you allocate memory in any fashion. Even calling malloc or realloc can move the relocatable blocks.

For each relocatable block, you must make a handle—a pointer object in memory, designated to store the address of that block. The relocating allocator knows where each block’s handle is, and updates the address stored there whenever it moves the block, so that the handle always points to the block. Each time you access the contents of the block, you should fetch its address anew from the handle.

To call any of the relocating allocator functions from a signal handler is almost certainly incorrect, because the signal could happen at any time and relocate all the blocks. The only way to make this safe is to block the signal around any access to the contents of any relocatable block—not a convenient mode of operation. See Section 21.4.6 [Signal Handling and Nonreentrant Functions], page 421.

3.6.2 Allocating and Freeing Relocatable Blocks

In the descriptions below, handleptr designates the address of the handle. All the functions are declared in ‘malloc.h’; all are GNU extensions.

Function void * r alloc (void **handleptr, size_t size)

This function allocates a relocatable block of size size. It stores the block’s address in *handleptr and returns a non-null pointer to indicate success. If r_alloc can’t get the space needed, it stores a null pointer in *handleptr, and returns a null pointer.

Function void r alloc free (void **handleptr)

This function is the way to free a relocatable block. It frees the block that *handleptr points to, and stores a null pointer in *handleptr to show it doesn’t point to an allocated block any more.

Function void * r re alloc (void **handleptr, size_t size)

The function r_re_alloc adjusts the size of the block that *handleptr points to, making it size bytes long. It stores the address of the resized block in *handleptr and returns a non-null pointer to indicate success. If enough memory is not available, this function returns a null pointer and does not modify *handleptr.

60 The GNU C Library

4 Character Handling

Programs that work with characters and strings often need to classify a character—is it alphabetic, is it a digit, is it whitespace, and so on— and perform case conversion operations on characters. The functions in the header file ‘ctype.h’ are provided for this purpose.

Since the choice of locale and character set can alter the classifications of particular character codes, all of these functions are affected by the current locale. (More precisely, they are affected by the locale currently selected for character classification—the LC_CTYPE category; see Section 19.3 [Categories of Activities that Locales Affect], page 380.)

4.1 Classification of Characters

This section explains the library functions for classifying characters. For example, isalpha is the function to test for an alphabetic character. It takes one argument, the character to test, and returns a nonzero integer if the character is alphabetic, and zero otherwise. You would use it like this:

     if (isalpha (c))
       printf ("The character `%c' is alphabetic.\n", c);

Each of the functions in this section tests for membership in a particular class of characters; each has a name starting with ‘is’. Each of them takes one argument, which is a character to test, and returns an int which is treated as a boolean value. The character argument is passed as an int, and it may be the constant value EOF instead of a real character.

The attributes of any given character can vary between locales. See Chapter 19 [Locales and Internationalization], page 379, for more information on locales.

These functions are declared in the header file ‘ctype.h’.

Function int islower (int c)

Returns true if c is a lower-case letter.

Function int isupper (int c)

Returns true if c is an upper-case letter.

Function int isalpha (int c)

Returns true if c is an alphabetic character (a letter). If islower or isupper is true of a character, then isalpha is also true. In some locales, there may be additional characters for which isalpha is true–letters which are neither upper case nor lower case. But in the standard "C" locale, there are no such additional characters.

Function int isdigit (int c)

Returns true if c is a decimal digit (‘0’ through ‘9’).

Function int isalnum (int c)

Returns true if c is an alphanumeric character (a letter or number); in other words, if either isalpha or isdigit is true of a character, then isalnum is also true.

Function int isxdigit (int c)

Returns true if c is a hexadecimal digit. Hexadecimal digits include the normal decimal digits ‘0’ through ‘9’ and the letters ‘A’ through ‘F’ and ‘a’ through ‘f’.

Function int ispunct (int c)

Returns true if c is a punctuation character. This means any printing character that is not alphanumeric or a space character.

Function int isspace (int c)

Returns true if c is a whitespace character. In the standard "C" locale, isspace returns true for only the standard whitespace characters:

' ' space

'\f' formfeed

'\n' newline

'\r' carriage return

'\t' horizontal tab

'\v' vertical tab

Function int isblank (int c)

Returns true if c is a blank character; that is, a space or a tab. This function is a GNU extension.

Function int isgraph (int c)

Returns true if c is a graphic character; that is, a character that has a glyph associated with it. The whitespace characters are not considered graphic.

Function int isprint (int c)

Returns true if c is a printing character. Printing characters include all the graphic characters, plus the space (‘ ’) character.

Function int iscntrl (int c)

Returns true if c is a control character (that is, a character that is not a printing character).

Function int isascii (int c)

Returns true if c is a 7-bit unsigned char value that fits into the US/UK ASCII character set. This function is a BSD extension and is also an SVID extension.

4.2 Case Conversion

This section explains the library functions for performing conversions such as case mappings on characters. For example, toupper converts any character to upper case if possible. If the character can’t be converted, toupper returns it unchanged.

These functions take one argument of type int, which is the character to convert, and return the converted character as an int. If the conversion is not applicable to the argument given, the argument is returned unchanged.

Compatibility Note: In pre-ISO C dialects, instead of returning the argument unchanged, these functions may fail when the argument is not suitable for the conversion. Thus for portability, you may need to write islower(c)

? toupper(c) : c rather than just toupper(c).

These functions are declared in the header file ‘ctype.h’.

Function int tolower (int c)

If c is an upper-case letter, tolower returns the corresponding lowercase letter. If c is not an upper-case letter, c is returned unchanged.

Function int toupper (int c)

If c is a lower-case letter, tolower returns the corresponding uppercase letter. Otherwise c is returned unchanged.

Function int toascii (int c)

This function converts c to a 7-bit unsigned char value that fits into the US/UK ASCII character set, by clearing the high-order bits. This function is a BSD extension and is also an SVID extension.

Function int tolower (int c)

This is identical to tolower, and is provided for compatibility with the SVID. See Section 1.2.4 [SVID (The System V Interface Description)], page 3.

Function int toupper (int c)

This is identical to toupper, and is provided for compatibility with the SVID.

64 The GNU C Library

5 String and Array Utilities

Operations on strings (or arrays of characters) are an important part of many programs. The GNU C library provides an extensive set of string utility functions, including functions for copying, concatenating, comparing, and searching strings. Many of these functions can also operate on arbitrary regions of storage; for example, the memcpy function can be used to copy the contents of any kind of array.

It’s fairly common for beginning C programmers to “reinvent the wheel” by duplicating this functionality in their own code, but it pays to become familiar with the library functions and to make use of them, since this offers benefits in maintenance, efficiency, and portability.

For instance, you could easily compare one string to another in two lines of C code, but if you use the built-in strcmp function, you’re less likely to make a mistake. And, since these library functions are typically highly optimized, your program may run faster too.

5.1 Representation of Strings

This section is a quick summary of string concepts for beginning C programmers. It describes how character strings are represented in C and some common pitfalls. If you are already familiar with this material, you can skip this section.

A string is an array of char objects. But string-valued variables are usually declared to be pointers of type char *. Such variables do not include space for the text of a string; that has to be stored somewhere else—in an array variable, a string constant, or dynamically allocated memory (see Chapter 3 [Memory Allocation], page 31). It’s up to you to store the address of the chosen memory space into the pointer variable. Alternatively you can store a null pointer in the pointer variable. The null pointer does not point anywhere, so attempting to reference the string it points to gets an error.

By convention, a null character, '\0', marks the end of a string. For example, in testing to see whether the char * variable p points to a null character marking the end of a string, you can write !*p or *p == '\0'.

A null character is quite different conceptually from a null pointer, although both are represented by the integer 0.

String literals appear in C program source as strings of characters between double-quote characters (‘"’). In ISO C, string literals can also be formed by string concatenation: "a" "b" is the same as "ab". Modification of string literals is not allowed by the GNU C compiler, because literals are placed in read-only storage.

Character arrays that are declared const cannot be modified either. It’s generally good style to declare non-modifiable string pointers to be of type const char *, since this often allows the C compiler to detect accidental modifications as well as providing some amount of documentation about what your program intends to do with the string.

The amount of memory allocated for the character array may extend past the null character that normally marks the end of the string. In this document, the term allocation size is always used to refer to the total amount of memory allocated for the string, while the term length refers to the number of characters up to (but not including) the terminating null character.

A notorious source of program bugs is trying to put more characters in a string than fit in its allocated size. When writing code that extends strings or moves characters into a pre-allocated array, you should be very careful to keep track of the length of the text and make explicit checks for overflowing the array. Many of the library functions do not do this for you! Remember also that you need to allocate an extra byte to hold the null character that marks the end of the string.

5.2 String and Array Conventions

This chapter describes both functions that work on arbitrary arrays or blocks of memory, and functions that are specific to null-terminated arrays of characters.

Functions that operate on arbitrary blocks of memory have names beginning with ‘mem’ (such as memcpy) and invariably take an argument which specifies the size (in bytes) of the block of memory to operate on. The array arguments and return values for these functions have type void *, and as a matter of style, the elements of these arrays are referred to as “bytes”. You can pass any kind of pointer to these functions, and the sizeof operator is useful in computing the value for the size argument.

In contrast, functions that operate specifically on strings have names beginning with ‘str’ (such as strcpy) and look for a null character to terminate the string instead of requiring an explicit size argument to be passed. (Some of these functions accept a specified maximum length, but they also check for premature termination with a null character.) The array arguments and return values for these functions have type char *, and the array elements are referred to as “characters”.

In many cases, there are both ‘mem’ and ‘str’ versions of a function. The one that is more appropriate to use depends on the exact situation. When your program is manipulating arbitrary arrays or blocks of storage, then you should always use the ‘mem’ functions. On the other hand, when you are manipulating null-terminated strings it is usually more convenient to use the ‘str’ functions, unless you already know the length of the string in advance.

5.3 String Length

You can get the length of a string using the strlen function. This function is declared in the header file ‘string.h’.

Function size_t strlen (const char *s)

The strlen function returns the length of the null-terminated string s. (In other words, it returns the offset of the terminating null character within the array.) For example,

        strlen ("hello, world")
            ⇒ 12

When applied to a character array, the strlen function returns the length of the string stored there, not its allocation size. You can get the allocation size of the character array that holds a string using the sizeof operator:

        char string[32] = "hello, world";
        sizeof (string)
            ⇒ 32
        strlen (string)
            ⇒ 12

5.4 Copying and Concatenation

You can use the functions described in this section to copy the contents of strings and arrays, or to append the contents of one string to another. These functions are declared in the header file ‘string.h’.

A helpful way to remember the ordering of the arguments to the functions in this section is that it corresponds to an assignment expression, with the destination array specified to the left of the source array. All of these functions return the address of the destination array.

Most of these functions do not work properly if the source and destination arrays overlap. For example, if the beginning of the destination array overlaps the end of the source array, the original contents of that part of the source array may get overwritten before it is copied. Even worse, in the case of the string functions, the null character marking the end of the string may be lost, and the copy function might get stuck in a loop trashing all the memory allocated to your program.

All functions that have problems copying between overlapping arrays are explicitly identified in this manual. In addition to functions in this section, there are a few others like sprintf (see Section 7.10.7 [Formatted Output Functions], page 113) and scanf (see Section 7.12.8 [Formatted Input Functions], page 134).

Function void * memcpy (void *to, const void *from, size_t

size) The memcpy function copies size bytes from the object beginning at from into the object beginning at to. The behavior of this function is undefined if the two arrays to and from overlap; use memmove instead if overlapping is possible. The value returned by memcpy is the value of to. Here is an example of how you might use memcpy to copy the contents of an array:

     struct foo *oldarray, *newarray;
     int arraysize;

…

     memcpy (new, old, arraysize * sizeof (struct foo));
Function void * memmove (void *to, const void *from,
      size_t size)

memmove copies the size bytes at from into the size bytes at to, even if those two blocks of space overlap. In the case of overlap, memmove is careful to copy the original values of the bytes in the block at from, including those bytes which also belong to the block at to.

Function void * memccpy (void *to, const void *from, int c,
      size_t size)

This function copies no more than size bytes from from to to, stopping if a byte matching c is found. The return value is a pointer into to one byte past where c was copied, or a null pointer if no byte matching c appeared in the first size bytes of from.

Function void * memset (void *block, int c, size_t size)

This function copies the value of c (converted to an unsigned char) into each of the first size bytes of the object beginning at block. It returns the value of block.

Function char * strcpy (char *to, const char *from)

This copies characters from the string from (up to and including the terminating null character) into the string to. Like memcpy, this function has undefined results if the strings overlap. The return value is the value of to.

Function char * strncpy (char *to, const char *from, size_t

size) This function is similar to strcpy but always copies exactly size characters into to. If the length of from is more than size, then strncpy copies just the first size characters. Note that in this case there is no null terminator written into to.

If the length of from is less than size, then strncpy copies all of from, followed by enough null characters to add up to size characters in all. This behavior is rarely useful, but it is specified by the ISO C standard. The behavior of strncpy is undefined if the strings overlap. Using strncpy as opposed to strcpy is a way to avoid bugs relating to writing past the end of the allocated space for to. However, it can also make your program much slower in one common case: copying a string which is probably small into a potentially large buffer. In this case, size may be large, and when it is, strncpy will waste a considerable amount of time copying null characters.

Function char * strdup (const char *s)

This function copies the null-terminated string s into a newly allocated string. The string is allocated using malloc; see Section 3.3 [Unconstrained Allocation], page 32. If malloc cannot allocate space for the new string, strdup returns a null pointer. Otherwise it returns a pointer to the new string.

Function char * strndup (const char *s, size_t size)

This function is similar to strdup but always copies at most size characters into the newly allocated string. If the length of s is more than size, then strndup copies just the first size characters and adds a closing null terminator. Otherwise all characters are copied and the string is terminated. This function is different to strncpy in that it always terminates the destination string.

Function char * stpcpy (char *to, const char *from)

This function is like strcpy, except that it returns a pointer to the end of the string to (that is, the address of the terminating null character) rather than the beginning. For example, this program uses stpcpy to concatenate ‘foo’ and ‘bar’ to produce ‘foobar’, which it then prints.

     #include <string.h>
     #include <stdio.h>
     int
     main (void)
     {
       char buffer[10];
       char *to = buffer;
       to = stpcpy (to, "foo");
       to = stpcpy (to, "bar");
       puts (buffer);
       return 0;

     }

This function is not part of the ISO or POSIX standards, and is not customary on Unix systems, but we did not invent it either. Perhaps it comes from MS-DOG. Its behavior is undefined if the strings overlap.

Function char * stpncpy (char *to, const char *from, size_t

size) This function is similar to stpcpy but copies always exactly size characters into to. If the length of from is more then size, then stpncpy copies just the first size characters and returns a pointer to the character directly following the one which was copied last. Note that in this case there is no null terminator written into to. If the length of from is less than size, then stpncpy copies all of from, followed by enough null characters to add up to size characters in all. This behaviour is rarely useful, but it is implemented to be useful in contexts where this behaviour of the strncpy is used. stpncpy returns a pointer to the first written null character. This function is not part of ISO or POSIX but was found useful while developing GNU C Library itself. Its behaviour is undefined if the strings overlap.

Function char * strdupa (const char *s)

This function is similar to strdup but allocates the new string using alloca instead of malloc see Section 3.5 [Automatic Storage with Variable Size], page 55. This means of course the returned string has the same limitations as any block of memory allocated using alloca. For obvious reasons strdupa is implemented only as a macro. I.e., you cannot get the address of this function. Despite this limitations it is a useful function. The following code shows a situation where using malloc would be a lot more expensive.

     #include <paths.h>
     #include <string.h>
     #include <stdio.h>
     const char path[] = _PATH_STDPATH;
     int
     main (void)
     {
       char *wr_path = strdupa (path);
       char *cp = strtok (wr_path, ":");

       while (cp != NULL)
         {
           puts (cp);
           cp = strtok (NULL, ":");
         }
       return 0;
     }

Please note that calling strtok using path directly is illegal.

This function is only available if GNU CC is used.

Function char * strndupa (const char *s, size_t size)

This function is similar to strndup but like strdupa it allocates the new string using alloca see Section 3.5 [Automatic Storage with Variable Size], page 55. The same advantages and limitations of strdupa are valid for strndupa, too.

This function is implemented only as a macro which means one cannot get the address of it.

strndupa is only available if GNU CC is used.

Function char * strcat (char *to, const char *from)

The strcat function is similar to strcpy, except that the characters from from are concatenated or appended to the end of to, instead of overwriting it. That is, the first character from from overwrites the null character marking the end of to.

An equivalent definition for strcat would be:

     char *
     strcat (char *to, const char *from)
     {
       strcpy (to + strlen (to), from);
       return to;
     }

This function has undefined results if the strings overlap.

Function char * strncat (char *to, const char *from, size_t

size) This function is like strcat except that not more than size characters from from are appended to the end of to. A single null character is also always appended to to, so the total allocated size of to must be at least size + 1 bytes longer than its initial length.

The strncat function could be implemented like this:

        char *
        strncat (char *to, const char *from, size_t size)
        {
          strncpy (to + strlen (to), from, size);
          return to;
        }

The behavior of strncat is undefined if the strings overlap.

Here is an example showing the use of strncpy and strncat. Notice how, in the call to strncat, the size parameter is computed to avoid overflowing the character array buffer.

     #include <string.h>
     #include <stdio.h>
     #define SIZE 10
     static char buffer[SIZE];
     main ()
     {
       strncpy (buffer, "hello", SIZE);
       puts (buffer);
       strncat (buffer, ", world", SIZE - strlen (buffer) - 1);
       puts (buffer);
     }

The output produced by this program looks like:

     hello
     hello, wo
Function void * bcopy (void *from, const void *to, size_t

size) This is a partially obsolete alternative for memmove, derived from BSD. Note that it is not quite equivalent to memmove, because the arguments are not in the same order.

Function void * bzero (void *block, size_t size)

This is a partially obsolete alternative for memset, derived from BSD. Note that it is not as general as memset, because the only value it can store is zero.

5.5 String/Array Comparison

You can use the functions in this section to perform comparisons on the contents of strings and arrays. As well as checking for equality, these functions can also be used as the ordering functions for sorting operations. See Chapter 15 [Searching and Sorting], page 317, for an example of this.

Unlike most comparison operations in C, the string comparison functions return a nonzero value if the strings are not equivalent rather than if they are. The sign of the value indicates the relative ordering of the first characters in the strings that are not equivalent: a negative value indicates that the first string is “less” than the second, while a positive value indicates that the first string is “greater”.

The most common use of these functions is to check only for equality. This is canonically done with an expression like ‘! strcmp (s1, s2)’.

All of these functions are declared in the header file ‘string.h’.

Function int memcmp (const void *a1, const void *a2,
         size_t size)

The function memcmp compares the size bytes of memory beginning at a1 against the size bytes of memory beginning at a2. The value returned has the same sign as the difference between the first differing pair of bytes (interpreted as unsigned char objects, then promoted to

  int).

If the contents of the two blocks are equal, memcmp returns 0.

On arbitrary arrays, the memcmp function is mostly useful for testing equality. It usually isn’t meaningful to do byte-wise ordering comparisons on arrays of things other than bytes. For example, a byte-wise comparison on the bytes that make up floating-point numbers isn’t likely to tell you anything about the relationship between the values of the floating-point numbers.

You should also be careful about using memcmp to compare objects that can contain “holes”, such as the padding inserted into structure objects to enforce alignment requirements, extra space at the end of unions, and extra characters at the ends of strings whose length is less than their allocated size. The contents of these “holes” are indeterminate and may cause strange behavior when performing byte-wise comparisons. For more predictable results, perform an explicit component-wise comparison.

For example, given a structure type definition like:

     struct foo
       {
         unsigned char tag;
         union
           {
             double f;
             long i;
             char *p;
           } value;

    };

you are better off writing a specialized comparison function to compare struct foo objects instead of comparing them with memcmp.

Function int strcmp (const char *s1, const char *s2)

The strcmp function compares the string s1 against s2, returning a value that has the same sign as the difference between the first differing pair of characters (interpreted as unsigned char objects, then promoted to int). If the two strings are equal, strcmp returns 0. A consequence of the ordering used by strcmp is that if s1 is an initial substring of s2, then s1 is considered to be “less than” s2.

Function int strcasecmp (const char *s1, const char *s2)

This function is like strcmp, except that differences in case are ignored. strcasecmp is derived from BSD.

Function int strncasecmp (const char *s1, const char *s2,
      size_t n)

This function is like strncmp, except that differences in case are ignored. strncasecmp is a GNU extension.

Function int strncmp (const char *s1, const char *s2, size_t

size) This function is the similar to strcmp, except that no more than size characters are compared. In other words, if the two strings are the same in their first size characters, the return value is zero.

Here are some examples showing the use of strcmp and strncmp. These examples assume the use of the ASCII character set. (If some other character set—say, EBCDIC—is used instead, then the glyphs are associated with different numeric codes, and the return values and ordering may differ.)

  strcmp ("hello", "hello")

⇒ 0 /* These two strings are the same. */

  strcmp ("hello", "Hello")

⇒ 32 /* Comparisons are case-sensitive. */

  strcmp ("hello", "world")

⇒ -15 /* The character 'h' comes before 'w'. */

  strcmp ("hello", "hello, world")

⇒ -44 /* Comparing a null character against a comma. */

  strncmp ("hello", "hello, world", 5)

⇒ 0 /* The initial 5 characters are the same. */

  strncmp ("hello, world", "hello, stupid world!!!", 5)

⇒ 0 /* The initial 5 characters are the same. */

Function int bcmp (const void *a1, const void *a2, size_t

size) This is an obsolete alias for memcmp, derived from BSD.

5.6 Collation Functions

In some locales, the conventions for lexicographic ordering differ from the strict numeric ordering of character codes. For example, in Spanish most glyphs with diacritical marks such as accents are not considered distinct letters for the purposes of collation. On the other hand, the two-character sequence ‘ll’ is treated as a single letter that is collated immediately after ‘l’.

You can use the functions strcoll and strxfrm (declared in the header file ‘string.h’) to compare strings using a collation ordering appropriate for the current locale. The locale used by these functions in particular can be specified by setting the locale for the LC_COLLATE category; see Chapter 19 [Locales and Internationalization], page 379.

In the standard C locale, the collation sequence for strcoll is the same as that for strcmp.

Effectively, the way these functions work is by applying a mapping to transform the characters in a string to a byte sequence that represents the string’s position in the collating sequence of the current locale. Comparing two such byte sequences in a simple fashion is equivalent to comparing the strings with the locale’s collating sequence.

The function strcoll performs this translation implicitly, in order to do one comparison. By contrast, strxfrm performs the mapping explicitly. If you are making multiple comparisons using the same string or set of strings, it is likely to be more efficient to use strxfrm to transform all the strings just once, and subsequently compare the transformed strings with strcmp.

Function int strcoll (const char *s1, const char *s2)

The strcoll function is similar to strcmp but uses the collating sequence of the current locale for collation (the LC_COLLATE locale).

Here is an example of sorting an array of strings, using strcoll to compare them. The actual sort algorithm is not written here; it comes from qsort (see Section 15.3 [Array Sort Function], page 318). The job of the code shown here is to say how to compare the strings while sorting them. (Later on in this section, we will show a way to do this more efficiently using

strxfrm.)

/* This is the comparison function used with qsort. */

     int
     compare_elements (char **p1, char **p2)

  {
    return strcoll (*p1, *p2);
  }

/* This is the entry point—the function to sort strings using the locale’s collating sequence. */

  void
  sort_strings (char **array, int nstrings)
  {

/* Sort temp_array by comparing the strings. */

    qsort (array, sizeof (char *),
           nstrings, compare_elements);
  }
Function size_t strxfrm (char *to, const char *from, size_t

size) The function strxfrm transforms string using the collation transformation determined by the locale currently selected for collation, and stores the transformed string in the array to. Up to size characters (including a terminating null character) are stored. The behavior is undefined if the strings to and from overlap; see Section 5.4 [Copying and Concatenation], page 67. The return value is the length of the entire transformed string. This value is not affected by the value of size, but if it is greater or equal than size, it means that the transformed string did not entirely fit in the array to. In this case, only as much of the string as actually fits was stored. To get the whole transformed string, call strxfrm again with a bigger output array. The transformed string may be longer than the original string, and it may also be shorter. If size is zero, no characters are stored in to. In this case, strxfrm simply returns the number of characters that would be the length of the transformed string. This is useful for determining what size string to allocate. It does not matter what to is if size is zero; to may even be a null pointer.

Here is an example of how you can use strxfrm when you plan to do many comparisons. It does the same thing as the previous example, but much faster, because it has to transform each string only once, no matter how many times it is compared with other strings. Even the time needed to allocate and free storage is much less than the time we save, when there are many strings.

  struct sorter { char *input; char *transformed; };

Chapter 5: String and Array Utilities 77

/* This is the comparison function used with qsort to sort an array of struct sorter. */

int
compare_elements (struct sorter *p1, struct sorter *p2)
{
return strcmp (p1->transformed, p2->transformed);
}

/* This is the entry point—the function to sort strings using the locale’s collating sequence. */

void
sort_strings_fast (char **array, int nstrings)
{
struct sorter temp_array[nstrings];
int i;

/* Set up temp_array. Each element contains

one input string and its transformed string. */

for (i = 0; i < nstrings; i++)
{
size_t length = strlen (array[i]) * 2;
char *transformed;
size_t transformed_lenght;
temp_array[i].input = array[i];

/* First try a buffer perhaps big enough. */

transformed = (char *) xmalloc (length);
/* Transform array[i]. */
transformed_length = strxfrm (transformed, array[i], length);

/* If the buffer was not large enough, resize it

and try again. */

if (transformed_length >= length)
  {

/* Allocate the needed space. +1 for terminating

NUL character. */

    transformed = (char *) xrealloc (transformed,
                                     transformed_length + 1);

/* The return value is not interesting because we know

how long the transformed string is. */

            (void) strxfrm (transformed, array[i], transformed_length + 1);
          }
        temp_array[i].transformed = transformed;
      }

/* Sort temp_array by comparing transformed strings. */

    qsort (temp_array, sizeof (struct sorter),
           nstrings, compare_elements);

/* Put the elements back in the permanent array

in their sorted order. */

    for (i = 0; i < nstrings; i++)
      array[i] = temp_array[i].input;

/* Free the strings we allocated. */

    for (i = 0; i < nstrings; i++)
      free (temp_array[i].transformed);
  }

Compatibility Note: The string collation functions are a new feature of ISO C. Older C dialects have no equivalent feature.

5.7 Search Functions

This section describes library functions which perform various kinds of searching operations on strings and arrays. These functions are declared in the header file ‘string.h’.

Function void * memchr (const void *block, int c, size_t

size) This function finds the first occurrence of the byte c (converted to an unsigned char) in the initial size bytes of the object beginning at block. The return value is a pointer to the located byte, or a null pointer if no match was found.

Function char * strchr (const char *string, int c)

The strchr function finds the first occurrence of the character c (converted to a char) in the null-terminated string beginning at string. The return value is a pointer to the located character, or a null pointer if no match was found. For example,

     strchr ("hello, world", 'l')

         ⇒ "llo, world"
     strchr ("hello, world", '?')
         ⇒ NULL

The terminating null character is considered to be part of the string, so you can use this function get a pointer to the end of a string by specifying a null character as the value of the c argument.

Function char * index (const char *string, int c)

index is another name for strchr; they are exactly the same.

Function char * strrchr (const char *string, int c)

The function strrchr is like strchr, except that it searches backwards from the end of the string string (instead of forwards from the front). For example,

     strrchr ("hello, world", 'l')
         ⇒ "ld"

Function char * rindex (const char *string, int c)

rindex is another name for strrchr; they are exactly the same.

Function char * strstr (const char *haystack, const char

*needle) This is like strchr, except that it searches haystack for a substring needle rather than just a single character. It returns a pointer into the string haystack that is the first character of the substring, or a null pointer if no match was found. If needle is an empty string, the function returns haystack. For example,

     strstr ("hello, world", "l")
         ⇒ "llo, world"
     strstr ("hello, world", "wo")
         ⇒ "world"

Function void * memmem (const void *haystack, size_t haystack-len,

      const void *needle, size_t needle-len)

This is like strstr, but needle and haystack are byte arrays rather than null-terminated strings. needle-len is the length of needle and haystack-len is the length of haystack. This function is a GNU extension.

Function size_t strspn (const char *string, const char

*skipset) The strspn (“string span”) function returns the length of the initial substring of string that consists entirely of characters that are members of the set specified by the string skipset. The order of the characters in skipset is not important. For example,

     strspn ("hello, world", "abcdefghijklmnopqrstuvwxyz")
         ⇒ 5
Function size_t strcspn (const char *string, const char

*stopset) The strcspn (“string complement span”) function returns the length of the initial substring of string that consists entirely of characters that are not members of the set specified by the string stopset. (In other words, it returns the offset of the first character in string that is a member of the set stopset.) For example,

     strcspn ("hello, world", " \t\n,.;!?")
         ⇒ 5
Function char * strpbrk (const char *string, const char

*stopset) The strpbrk (“string pointer break”) function is related to strcspn, except that it returns a pointer to the first character in string that is a member of the set stopset instead of the length of the initial substring. It returns a null pointer if no such character from stopset is found. For example,

     strpbrk ("hello, world", " \t\n,.;!?")
         ⇒ ", world"

5.8 Finding Tokens in a String

It’s fairly common for programs to have a need to do some simple kinds of lexical analysis and parsing, such as splitting a command string up into tokens. You can do this with the strtok function, declared in the header

file ‘string.h’.

Function char * strtok (char *newstring, const char

*delimiters) A string can be split into tokens by making a series of calls to the function strtok. The string to be split up is passed as the newstring argument on the first call only. The strtok function uses this to set up some internal state information. Subsequent calls to get additional tokens from the same string are indicated by passing a null pointer as the newstring argument. Calling strtok with another non-null newstring argument reinitializes the state information. It is guaranteed that no other library function ever calls strtok behind your back (which would mess up this internal state information). The delimiters argument is a string that specifies a set of delimiters that may surround the token being extracted. All the initial characters that are members of this set are discarded. The first character that is not a member of this set of delimiters marks the beginning of the next token. The end of the token is found by looking for the next character that is a member of the delimiter set. This character in the original string newstring is overwritten by a null character, and the pointer to the beginning of the token in newstring is returned. On the next call to strtok, the searching begins at the next character beyond the one that marked the end of the previous token. Note that the set of delimiters delimiters do not have to be the same on every call in a series of calls to strtok. If the end of the string newstring is reached, or if the remainder of string consists only of delimiter characters, strtok returns a null pointer.

Warning: Since strtok alters the string it is parsing, you always copy the string to a temporary buffer before parsing it with strtok. If you allow strtok to modify a string that came from another part of your program, you are asking for trouble; that string may be part of a data structure that could be used for other purposes during the parsing, when alteration by strtok makes the data structure temporarily inaccurate.

The string that you are operating on might even be a constant. Then when strtok tries to modify it, your program will get a fatal signal for writing in read-only memory. See Section 21.2.1 [Program Error Signals], page 396.

This is a special case of a general principle: if a part of a program does not have as its purpose the modification of a certain data structure, then it is error-prone to modify the data structure temporarily.

The function strtok is not reentrant. See Section 21.4.6 [Signal Handling and Nonreentrant Functions], page 421, for a discussion of where and why reentrancy is important.

Here is a simple example showing the use of strtok.

  #include <string.h>
  #include <stddef.h>

…

  char string[] = "words separated by spaces -- and, punctuation!";
  const char delimiters[] = " .,;:!-";
  char *token;

…

  token = strtok (string, delimiters);  /* token => "words" */
  token = strtok (NULL, delimiters);    /* token => "separated" */
  token = strtok (NULL, delimiters);    /* token => "by" */
  token = strtok (NULL, delimiters);    /* token => "spaces" */
  token = strtok (NULL, delimiters);    /* token => "and" */
  token = strtok (NULL, delimiters);    /* token => "punctuation" */
  token = strtok (NULL, delimiters);    /* token => NULL */

The GNU C library contains two more functions for tokenizing a string which overcome the limitation of non-reentrancy.

Function char * strtok r (char *newstring, const char

*delimiters, char **save ptr) Just like strtok this function splits the string into several tokens which can be accessed be successive calls to strtok_r. The difference is that the information about the next token is not set up in some internal state information. Instead the caller has to provide another argument save ptr which is a pointer to a string pointer. Calling strtok_r with a null pointer for newstring and leaving save ptr between the calls unchanged does the job without limiting reentrancy. This function was proposed for POSIX.1b and can be found on many systems which support multi-threading.

Function char * strsep (char **string ptr, const char

*delimiter) A second reentrant approach is to avoid the additional first argument. The initialization of the moving pointer has to be done by the user. Successive calls of strsep move the pointer along the tokens separated by delimiter, returning the address of the next token and updating string ptr to point to the beginning of the next token. This function was introduced in 4.3BSD and therefore is widely available.

Here is how the above example looks like when strsep is used.

  #include <string.h>
  #include <stddef.h>

…

  char string[] = "words separated by spaces -- and, punctuation!";
  const char delimiters[] = " .,;:!-";
  char *running;
  char *token;

Chapter 5: String and Array Utilities 83

…

running = string;
token = strsep (&running, delimiters);    /* token => "words" */
token = strsep (&running, delimiters);    /* token => "separated" */
token = strsep (&running, delimiters);    /* token => "by" */
token = strsep (&running, delimiters);    /* token => "spaces" */
token = strsep (&running, delimiters);    /* token => "and" */
token = strsep (&running, delimiters);    /* token => "punctuation" */
token = strsep (&running, delimiters);    /* token => NULL */

84 The GNU C Library

6 Input/Output Overview

Most programs need to do either input (reading data) or output (writing data), or most frequently both, in order to do anything useful. The GNU C library provides such a large selection of input and output functions that the hardest part is often deciding which function is most appropriate!

This chapter introduces concepts and terminology relating to input and output. Other chapters relating to the GNU I/O facilities are:

• Chapter 7 [Input/Output on Streams], page 93, which covers the highlevel functions that operate on streams, including formatted input and output.

• Chapter 8 [Low-Level Input/Output], page 151, which covers the basic I/O and control functions on file descriptors.

• Chapter 9 [File System Interface], page 181, which covers functions for operating on directories and for manipulating file attributes such as access modes and ownership.

• Chapter 10 [Pipes and FIFOs], page 215, which includes information on the basic interprocess communication facilities.

• Chapter 11 [Sockets], page 221, which covers a more complicated interprocess communication facility with support for networking.

• Chapter 12 [Low-Level Terminal Interface], page 271, which covers functions for changing how input and output to terminal or other serial devices are processed.

6.1 Input/Output Concepts

Before you can read or write the contents of a file, you must establish a connection or communications channel to the file. This process is called opening the file. You can open a file for reading, writing, or both.

The connection to an open file is represented either as a stream or as a file descriptor. You pass this as an argument to the functions that do the actual read or write operations, to tell them which file to operate on. Certain functions expect streams, and others are designed to operate on file descriptors.

When you have finished reading to or writing from the file, you can terminate the connection by closing the file. Once you have closed a stream or file descriptor, you cannot do any more input or output operations on it.

6.1.1 Streams and File Descriptors

When you want to do input or output to a file, you have a choice of two basic mechanisms for representing the connection between your program and the file: file descriptors and streams. File descriptors are represented as objects of type int, while streams are represented as FILE * objects.

File descriptors provide a primitive, low-level interface to input and output operations. Both file descriptors and streams can represent a connection to a device (such as a terminal), or a pipe or socket for communicating with another process, as well as a normal file. But, if you want to do control operations that are specific to a particular kind of device, you must use a file descriptor; there are no facilities to use streams in this way. You must also use file descriptors if your program needs to do input or output in special modes, such as nonblocking (or polled) input (see Section 8.10 [File Status Flags], page 170).

Streams provide a higher-level interface, layered on top of the primitive file descriptor facilities. The stream interface treats all kinds of files pretty much alike—the sole exception being the three styles of buffering that you can choose (see Section 7.17 [Stream Buffering], page 140).

The main advantage of using the stream interface is that the set of functions for performing actual input and output operations (as opposed to control operations) on streams is much richer and more powerful than the corresponding facilities for file descriptors. The file descriptor interface provides only simple functions for transferring blocks of characters, but the stream interface also provides powerful formatted input and output functions (printf and scanf) as well as functions for character- and line-oriented input and output.

Since streams are implemented in terms of file descriptors, you can extract the file descriptor from a stream and perform low-level operations directly on the file descriptor. You can also initially open a connection as a file descriptor and then make a stream associated with that file descriptor.

In general, you should stick with using streams rather than file descriptors, unless there is some specific operation you want to do that can only be done on a file descriptor. If you are a beginning programmer and aren’t sure what functions to use, we suggest that you concentrate on the formatted input functions (see Section 7.12 [Formatted Input], page 126) and formatted output functions (see Section 7.10 [Formatted Output], page 104).

If you are concerned about portability of your programs to systems other than GNU, you should also be aware that file descriptors are not as portable as streams. You can expect any system running ISO C to support streams, but non-GNU systems may not support file descriptors at all, or may only implement a subset of the GNU functions that operate on file descriptors. Most of the file descriptor functions in the GNU library are included in the POSIX.1 standard, however.

6.1.2 File Position

One of the attributes of an open file is its file position that keeps track of where in the file the next character is to be read or written. In the GNU system, and all POSIX.1 systems, the file position is simply an integer representing the number of bytes from the beginning of the file.

The file position is normally set to the beginning of the file when it is opened, and each time a character is read or written, the file position is incremented. In other words, access to the file is normally sequential.

Ordinary files permit read or write operations at any position within the file. Some other kinds of files may also permit this. Files which do permit this are sometimes referred to as random-access files. You can change the file position using the fseek function on a stream (see Section 7.15 [File Positioning], page 137) or the lseek function on a file descriptor (see Section 8.2 [Input and Output Primitives], page 153). If you try to change the file position on a file that doesn’t support random access, you get the

ESPIPE error.

Streams and descriptors that are opened for append access are treated specially for output: output to such files is always appended sequentially to the end of the file, regardless of the file position. However, the file position is still used to control where in the file reading is done.

If you think about it, you’ll realize that several programs can read a given file at the same time. In order for each program to be able to read the file at its own pace, each program must have its own file pointer, which is not affected by anything the other programs do.

In fact, each opening of a file creates a separate file position. Thus, if you open a file twice even in the same program, you get two streams or descriptors with independent file positions.

By contrast, if you open a descriptor and then duplicate it to get another descriptor, these two descriptors share the same file position: changing the file position of one descriptor will affect the other.

6.2 File Names

In order to open a connection to a file, or to perform other operations such as deleting a file, you need some way to refer to the file. Nearly all files have names that are strings—even files which are actually devices such as tape drives or terminals. These strings are called file names. You specify the file name to say which file you want to open or operate on.

This section describes the conventions for file names and how the operating system works with them.

6.2.1 Directories

In order to understand the syntax of file names, you need to understand how the file system is organized into a hierarchy of directories.

A directory is a file that contains information to associate other files with names; these associations are called links or directory entries. Sometimes, people speak of “files in a directory”, but in reality, a directory only contains pointers to files, not the files themselves.

The name of a file contained in a directory entry is called a file name component. In general, a file name consists of a sequence of one or more such components, separated by the slash character (‘/’). A file name which is just one component names a file with respect to its directory. A file name with multiple components names a directory, and then a file in that directory, and so on.

Some other documents, such as the POSIX standard, use the term pathname for what we call a file name, and either filename or pathname component for what this manual calls a file name component. We don’t use this terminology because a “path” is something completely different (a list of directories to search), and we think that “pathname” used for something else will confuse users. We always use “file name” and “file name component” (or sometimes just “component”, where the context is obvious) in GNU documentation. Some macros use the POSIX terminology in their names, such as PATH_MAX. These macros are defined by the POSIX standard, so we cannot change their names.

You can find more detailed information about operations on directories in Chapter 9 [File System Interface], page 181.

6.2.2 File Name Resolution

A file name consists of file name components separated by slash (‘/’) characters. On the systems that the GNU C library supports, multiple successive ‘/’ characters are equivalent to a single ‘/’ character.

The process of determining what file a file name refers to is called file name resolution. This is performed by examining the components that make up a file name in left-to-right order, and locating each successive component in the directory named by the previous component. Of course, each of the files that are referenced as directories must actually exist, be directories instead of regular files, and have the appropriate permissions to be accessible by the process; otherwise the file name resolution fails.

If a file name begins with a ‘/’, the first component in the file name is located in the root directory of the process (usually all processes on the system have the same root directory). Such a file name is called an absolute file name.

Otherwise, the first component in the file name is located in the current working directory (see Section 9.1 [Working Directory], page 181). This kind of file name is called a relative file name.

The file name components ‘.’ (“dot”) and ‘..’ (“dot-dot”) have special meanings. Every directory has entries for these file name components. The file name component ‘.’ refers to the directory itself, while the file name component ‘..’ refers to its parent directory (the directory that contains the link for the directory in question). As a special case, ‘..’ in the root directory refers to the root directory itself, since it has no parent; thus ‘/..’ is the same as ‘/’.

Here are some examples of file names: ‘/a’ The file named ‘a’, in the root directory. ‘/a/b’ The file named ‘b’, in the directory named ‘a’ in the root direc-

tory.

‘a’ The file named ‘a’, in the current working directory. ‘/a/./b’ This is the same as ‘/a/b’. ‘./a’ The file named ‘a’, in the current working directory. ‘../a’ The file named ‘a’, in the parent directory of the current working

directory. A file name that names a directory may optionally end in a ‘/’. You can

specify a file name of ‘/’ to refer to the root directory, but the empty string is not a meaningful file name. If you want to refer to the current working directory, use a file name of ‘.’ or ‘./’.

Unlike some other operating systems, the GNU system doesn’t have any built-in support for file types (or extensions) or file versions as part of its file name syntax. Many programs and utilities use conventions for file names— for example, files containing C source code usually have names suffixed with ‘.c’—but there is nothing in the file system itself that enforces this kind of convention.

6.2.3 File Name Errors

Functions that accept file name arguments usually detect these errno error conditions relating to the file name syntax or trouble finding the named file. These errors are referred to throughout this manual as the usual file name errors. EACCES The process does not have search permission for a directory com-

ponent of the file name.

ENAMETOOLONG

This error is used when either the the total length of a file name is greater than PATH_MAX, or when an individual file name component has a length greater than NAME_MAX. See Section 28.6 [Limits on File System Capacity], page 545. In the GNU system, there is no imposed limit on overall file name length, but some file systems may place limits on the length of a component.

ENOENT This error is reported when a file referenced as a directory component in the file name doesn’t exist, or when a component is a symbolic link whose target file does not exist. See Section 9.4 [Symbolic Links], page 190.

ENOTDIR A file that is referenced as a directory component in the file name exists, but it isn’t a directory.

ELOOP Too many symbolic links were resolved while trying to look up the file name. The system has an arbitrary limit on the number of symbolic links that may be resolved in looking up a single file name, as a primitive way to detect loops. See Section 9.4 [Symbolic Links], page 190.

6.2.4 Portability of File Names

The rules for the syntax of file names discussed in Section 6.2 [File Names], page 87, are the rules normally used by the GNU system and by other POSIX systems. However, other operating systems may use other conventions.

There are two reasons why it can be important for you to be aware of file name portability issues:

• If your program makes assumptions about file name syntax, or contains embedded literal file name strings, it is more difficult to get it to run under other operating systems that use different syntax conventions.

• Even if you are not concerned about running your program on machines that run other operating systems, it may still be possible to access files that use different naming conventions. For example, you may be able to access file systems on another computer running a different operating system over a network, or read and write disks in formats used by other operating systems. The ISO C standard says very little about file name syntax, only that

file names are strings. In addition to varying restrictions on the length of file names and what characters can validly appear in a file name, different operating systems use different conventions and syntax for concepts such as structured directories and file types or extensions. Some concepts such as file versions might be supported in some operating systems and not by others.

The POSIX.1 standard allows implementations to put additional restrictions on file name syntax, concerning what characters are permitted in file names and on the length of file name and file name component strings. However, in the GNU system, you do not need to worry about these restrictions; any character except the null character is permitted in a file name string, and there are no limits on the length of file name strings.

92 The GNU C Library

7 Input/Output on Streams

This chapter describes the functions for creating streams and performing input and output operations on them. As discussed in Chapter 6 [Input/Output Overview], page 85, a stream is a fairly abstract, high-level concept representing a communications channel to a file, device, or process.

7.1 Streams

For historical reasons, the type of the C data structure that represents a stream is called FILE rather than “stream”. Since most of the library functions deal with objects of type FILE *, sometimes the term file pointer is also used to mean “stream”. This leads to unfortunate confusion over terminology in many books on C. This manual, however, is careful to use the terms “file” and “stream” only in the technical sense.

The FILE type is declared in the header file ‘stdio.h’.

Data Type FILE

This is the data type used to represent stream objects. A FILE object holds all of the internal state information about the connection to the associated file, including such things as the file position indicator and buffering information. Each stream also has error and end-of-file status indicators that can be tested with the ferror and feof functions; see Section 7.13 [End-Of-File and Errors], page 135.

FILE objects are allocated and managed internally by the input/output library functions. Don’t try to create your own objects of type FILE; let the library do it. Your programs should deal only with pointers to these objects (that is, FILE * values) rather than the objects themselves.

7.2 Standard Streams

When the main function of your program is invoked, it already has three predefined streams open and available for use. These represent the “standard” input and output channels that have been established for the process.

These streams are declared in the header file ‘stdio.h’.

Variable FILE * stdin

The standard input stream, which is the normal source of input for the program.

Variable FILE * stdout

The standard output stream, which is used for normal output from the program.

Variable FILE * stderr

The standard error stream, which is used for error messages and diagnostics issued by the program.

In the GNU system, you can specify what files or processes correspond to these streams using the pipe and redirection facilities provided by the shell. (The primitives shells use to implement these facilities are described in Chapter 9 [File System Interface], page 181.) Most other operating systems provide similar mechanisms, but the details of how to use them can vary.

In the GNU C library, stdin, stdout, and stderr are normal variables which you can set just like any others. For example, to redirect the standard output to a file, you could do:

     fclose (stdout);
     stdout = fopen ("standard-output-file", "w");

Note however, that in other systems stdin, stdout, and stderr are macros that you cannot assign to in the normal way. But you can use freopen to get the effect of closing one and reopening it. See Section 7.3 [Opening Streams], page 94.

7.3 Opening Streams

Opening a file with the fopen function creates a new stream and establishes a connection between the stream and a file. This may involve creating a new file.

Everything described in this section is declared in the header file

‘stdio.h’.
Function FILE * fopen (const char *filename, const char

*opentype) The fopen function opens a stream for I/O to the file filename, and returns a pointer to the stream. The opentype argument is a string that controls how the file is opened and specifies attributes of the resulting stream. It must begin with one of the following sequences of characters:

‘r’ Open an existing file for reading only.

‘w’ Open the file for writing only. If the file already exists, it is truncated to zero length. Otherwise a new file is created.

‘a’ Open a file for append access; that is, writing at the end of file only. If the file already exists, its initial contents are unchanged and output to the stream is appended to the end of the file. Otherwise, a new, empty file is created.

‘r+’ Open an existing file for both reading and writing. The initial contents of the file are unchanged and the initial file position is at the beginning of the file.

‘w+’ Open a file for both reading and writing. If the file already exists, it is truncated to zero length. Otherwise, a new file is created.

‘a+’ Open or create file for both reading and appending. If the file exists, its initial contents are unchanged. Otherwise, a new file is created. The initial file position for reading is at the beginning of the file, but output is always appended to the end of the file.

As you can see, ‘+’ requests a stream that can do both input and output. The ISO standard says that when using such a stream, you must call fflush (see Section 7.17 [Stream Buffering], page 140) or a file positioning function such as fseek (see Section 7.15 [File Positioning], page 137) when switching from reading to writing or vice versa. Otherwise, internal buffers might not be emptied properly. The GNU C library does not have this limitation; you can do arbitrary reading and writing operations on a stream in whatever order. Additional characters may appear after these to specify flags for the call. Always put the mode (‘r’, ‘w+’, etc.) first; that is the only part you are guaranteed will be understood by all systems. The GNU C library defines one additional character for use in opentype: the character ‘x’ insists on creating a new file—if a file filename already exists, fopen fails rather than opening it. If you use ‘x’ you can are guaranteed that you will not clobber an existing file. This is equivalent to the O_EXCL option to the open function (see Section 8.1 [Opening and Closing Files], page 151). The character ‘b’ in opentype has a standard meaning; it requests a binary stream rather than a text stream. But this makes no difference in POSIX systems (including the GNU system). If both ‘+’ and ‘b’ are specified, they can appear in either order. See Section 7.14 [Text and Binary Streams], page 136. Any other characters in opentype are simply ignored. They may be meaningful in other systems. If the open fails, fopen returns a null pointer.

You can have multiple streams (or file descriptors) pointing to the same file open at the same time. If you do only input, this works straightforwardly, but you must be careful if any output streams are included. See Section 8.5 [Dangers of Mixing Streams and Descriptors], page 160. This is equally true whether the streams are in one program (not usual) or in several programs (which can easily happen). It may be advantageous to use the file locking facilities to avoid simultaneous access. See Section 8.11 [File Locks], page 175.

Macro int FOPEN MAX

The value of this macro is an integer constant expression that represents the minimum number of streams that the implementation guarantees can be open simultaneously. You might be able to open more than this many streams, but that is not guaranteed. The value of this constant is at least eight, which includes the three standard streams stdin, stdout, and stderr. In POSIX.1 systems this value is determined by the OPEN_MAX parameter; see Section 28.1 [General Capacity Limits], page 537. In BSD and GNU, it is controlled by the RLIMIT_ NOFILE resource limit; see Section 17.6 [Limiting Resource Usage], page 362.

Function FILE * freopen (const char *filename, const char

*opentype, FILE *stream) This function is like a combination of fclose and fopen. It first closes the stream referred to by stream, ignoring any errors that are detected in the process. (Because errors are ignored, you should not use freopen on an output stream if you have actually done any output using the stream.) Then the file named by filename is opened with mode opentype as for fopen, and associated with the same stream object stream. If the operation fails, a null pointer is returned; otherwise, freopen returns stream. freopen has traditionally been used to connect a standard stream such as stdin with a file of your own choice. This is useful in programs in which use of a standard stream for certain purposes is hard-coded. In the GNU C library, you can simply close the standard streams and open new ones with fopen. But other systems lack this ability, so using freopen is more portable.

7.4 Closing Streams

When a stream is closed with fclose, the connection between the stream and the file is cancelled. After you have closed a stream, you cannot perform any additional operations on it.

Function int fclose (FILE *stream)

This function causes stream to be closed and the connection to the corresponding file to be broken. Any buffered output is written and any buffered input is discarded. The fclose function returns a value of 0 if the file was closed successfully, and EOF if an error was detected.

It is important to check for errors when you call fclose to close an output stream, because real, everyday errors can be detected at this time. For example, when fclose writes the remaining buffered output, it might get an error because the disk is full. Even if you know the buffer is empty, errors can still occur when closing a file if you are using NFS. The function fclose is declared in ‘stdio.h’.

To close all streams currently available the GNU C Library provides another function.

Function int fcloseall (void)

This function causes all open streams of the process to be closed and the connection to corresponding files to be broken. All buffered data is written and any buffered inputis discarded. The fcloseall function returns a value of 0 if all the files were closed successfully, and EOF if an error was detected. This function should be used in only in special situation, e.g., when an error occurred and the program must be aborted. Normally each single stream should be closed separately so that problems with one stream can be identifier. It is also problematic since the standard streams (see Section 7.2 [Standard Streams], page 93) will also be closed. The function fcloseall is declared in ‘stdio.h’.

If the main function to your program returns, or if you call the exit function (see Section 22.3.1 [Normal Termination], page 462), all open streams are automatically closed properly. If your program terminates in any other manner, such as by calling the abort function (see Section 22.3.4 [Aborting a Program], page 465) or from a fatal signal (see Chapter 21 [Signal Handling], page 393), open streams might not be closed properly. Buffered output might not be flushed and files may be incomplete. For more information on buffering of streams, see Section 7.17 [Stream Buffering], page 140.

7.5 Simple Output by Characters or Lines

This section describes functions for performing character- and lineoriented output.

These functions are declared in the header file ‘stdio.h’.

Function int fputc (int c, FILE *stream)

The fputc function converts the character c to type unsigned char, and writes it to the stream stream. EOF is returned if a write error occurs; otherwise the character c is returned.

Function int putc (int c, FILE *stream)

This is just like fputc, except that most systems implement it as a macro, making it faster. One consequence is that it may evaluate the stream argument more than once, which is an exception to the general rule for macros. putc is usually the best function to use for writing a single character.

Function int putchar (int c)

The putchar function is equivalent to putc with stdout as the value of the stream argument.

Function int fputs (const char *s, FILE *stream)

The function fputs writes the string s to the stream stream. The terminating null character is not written. This function does not add a newline character, either. It outputs only the characters in the string. This function returns EOF if a write error occurs, and otherwise a nonnegative value. For example:

     fputs ("Are ", stdout);
     fputs ("you ", stdout);
     fputs ("hungry?\n", stdout);

outputs the text ‘Are you hungry?’ followed by a newline.

Function int puts (const char *s)

The puts function writes the string s to the stream stdout followed by a newline. The terminating null character of the string is not written. (Note that fputs does not write a newline as this function does.) puts is the most convenient function for printing simple messages. For example:

     puts ("This is a message.");

Function int putw (int w, FILE *stream)

This function writes the word w (that is, an int) to stream. It is provided for compatibility with SVID, but we recommend you use fwrite instead (see Section 7.9 [Block Input/Output], page 104).

7.6 Character Input

This section describes functions for performing character-oriented input. These functions are declared in the header file ‘stdio.h’.

These functions return an int value that is either a character of input, or the special value EOF (usually -1). It is important to store the result of these functions in a variable of type int instead of char, even when you plan to use it only as a character. Storing EOF in a char variable truncates its value to the size of a character, so that it is no longer distinguishable from the valid character ‘(char) -1’. So always use an int for the result of getc and friends, and check for EOF after the call; once you’ve verified that the result is not EOF, you can be sure that it will fit in a ‘char’ variable without loss of information.

Function int fgetc (FILE *stream)

This function reads the next character as an unsigned char from the stream stream and returns its value, converted to an int. If an endof-file condition or read error occurs, EOF is returned instead.

Function int getc (FILE *stream)

This is just like fgetc, except that it is permissible (and typical) for it to be implemented as a macro that evaluates the stream argument more than once. getc is often highly optimized, so it is usually the best function to use to read a single character.

Function int getchar (void)

The getchar function is equivalent to getc with stdin as the value of the stream argument.

Here is an example of a function that does input using fgetc. It would work just as well using getc instead, or using getchar () instead

of fgetc (stdin).
     int
     y_or_n_p (const char *question)
     {
       fputs (question, stdout);
       while (1)
         {
           int c, answer;

/* Write a space to separate answer from question. */

           fputc (' ', stdout);

/* Read the first character of the line.

This should be the answer character, but might not be. */

           c = tolower (fgetc (stdin));
           answer = c;

/* Discard rest of input line. */

           while (c != '\n' && c != EOF)
             c = fgetc (stdin);

/* Obey the answer if it was valid. */

           if (answer == 'y')
             return 1;
           if (answer == 'n')
             return 0;

/* Answer was invalid: ask for valid answer. */

        fputs ("Please answer y or n:", stdout);
      }
  }

Function int getw (FILE *stream)

This function reads a word (that is, an int) from stream. It’s provided for compatibility with SVID. We recommend you use fread instead (see Section 7.9 [Block Input/Output], page 104). Unlike getc, any int value could be a valid result. getw returns EOF when it encounters end-of-file or an error, but there is no way to distinguish this from an input word with value -1.

7.7 Line-Oriented Input

Since many programs interpret input on the basis of lines, it’s convenient to have functions to read a line of text from a stream.

Standard C has functions to do this, but they aren’t very safe: null characters and even (for gets) long lines can confuse them. So the GNU library provides the nonstandard getline function that makes it easy to read lines reliably.

Another GNU extension, getdelim, generalizes getline. It reads a delimited record, defined as everything through the next occurrence of a specified delimiter character.

All these functions are declared in ‘stdio.h’.

Function ssize_t getline (char **lineptr, size_t *n, FILE

*stream) This function reads an entire line from stream, storing the text (including the newline and a terminating null character) in a buffer and storing the buffer address in *lineptr. Before calling getline, you should place in *lineptr the address of a buffer *n bytes long, allocated with malloc. If this buffer is long enough to hold the line, getline stores the line in this buffer. Otherwise, getline makes the buffer bigger using realloc, storing the new buffer address back in *lineptr and the increased size back in *n. See Section 3.3 [Unconstrained Allocation], page 32. If you set *lineptr to a null pointer, and *n to zero, before the call, then getline allocates the initial buffer for you by calling malloc. In either case, when getline returns, *lineptr is a char * which points to the text of the line. When getline is successful, it returns the number of characters read (including the newline, but not including the terminating null). This value enables you to distinguish null characters that are part of the line from the null character inserted as a terminator.

This function is a GNU extension, but it is the recommended way to read lines from a stream. The alternative standard functions are unreliable. If an error occurs or end of file is reached, getline returns -1.

Function ssize_t getdelim (char **lineptr, size_t *n, int

delimiter, FILE *stream) This function is like getline except that the character which tells it to stop reading is not necessarily newline. The argument delimiter specifies the delimiter character; getdelim keeps reading until it sees that character (or end of file). The text is stored in lineptr, including the delimiter character and a terminating null. Like getline, getdelim makes lineptr bigger if it isn’t big enough. getline is in fact implemented in terms of getdelim, just like this:

     ssize_t
     getline (char **lineptr, size_t *n, FILE *stream)
     {
       return getdelim (lineptr, n, '\n', stream);
     }

Function char * fgets (char *s, int count, FILE *stream)

The fgets function reads characters from the stream stream up to and including a newline character and stores them in the string s, adding a null character to mark the end of the string. You must supply count characters worth of space in s, but the number of characters read is at most count − 1. The extra character space is used to hold the null character at the end of the string. If the system is already at end of file when you call fgets, then the contents of the array s are unchanged and a null pointer is returned. A null pointer is also returned if a read error occurs. Otherwise, the return value is the pointer s. Warning: If the input data has a null character, you can’t tell. So don’t use fgets unless you know the data cannot contain a null. Don’t use it to read files edited by the user because, if the user inserts a null character, you should either handle it properly or print a clear error message. We recommend using getline instead of fgets.

Deprecated function char * gets (char *s)

The function gets reads characters from the stream stdin up to the next newline character, and stores them in the string s. The newline character is discarded (note that this differs from the behavior of fgets, which copies the newline character into the string). If gets encounters a read error or end-of-file, it returns a null pointer; otherwise it returns s.

Warning: The gets function is very dangerous because it provides no protection against overflowing the string s. The GNU library includes it for compatibility only. You should always use fgets or getline instead. To remind you of this, the linker (if using GNU ld) will issue a warning whenever you use gets.

7.8 Unreading

In parser programs it is often useful to examine the next character in the input stream without removing it from the stream. This is called “peeking ahead” at the input because your program gets a glimpse of the input it will read next.

Using stream I/O, you can peek ahead at input by first reading it and then unreading it (also called pushing it back on the stream). Unreading a character makes it available to be input again from the stream, by the next call to fgetc or other input function on that stream.

7.8.1 What Unreading Means

Here is a pictorial explanation of unreading. Suppose you have a stream reading a file that contains just six characters, the letters ‘foobar’. Suppose you have read three characters so far. The situation looks like this:

     f  o  o  b  a  r
              ^

so the next input character will be ‘b’.

If instead of reading ‘b’ you unread the letter ‘o’, you get a situation like this:

     f  o  o  b  a  r
              |
           o--
           ^

so that the next input characters will be ‘o’ and ‘b’.

If you unread ‘9’ instead of ‘o’, you get this situation:

     f  o  o  b  a  r
              |
           9--
           ^

so that the next input characters will be ‘9’ and ‘b’.

7.8.2 Using ungetc To Do Unreading

The function to unread a character is called ungetc, because it reverses the action of getc.

Function int ungetc (int c, FILE *stream)

The ungetc function pushes back the character c onto the input stream stream. So the next input from stream will read c before anything else. If c is EOF, ungetc does nothing and just returns EOF. This lets you call ungetc with the return value of getc without needing to check for an error from getc. The character that you push back doesn’t have to be the same as the last character that was actually read from the stream. In fact, it isn’t necessary to actually read any characters from the stream before unreading them with ungetc! But that is a strange way to write a program; usually ungetc is used only to unread a character that was just read from the same stream. The GNU C library only supports one character of pushback—in other words, it does not work to call ungetc twice without doing input in between. Other systems might let you push back multiple characters; then reading from the stream retrieves the characters in the reverse order that they were pushed. Pushing back characters doesn’t alter the file; only the internal buffering for the stream is affected. If a file positioning function (such as fseek or rewind; see Section 7.15 [File Positioning], page 137) is called, any pending pushed-back characters are discarded. Unreading a character on a stream that is at end of file clears the end-of-file indicator for the stream, because it makes the character of input available. After you read that character, trying to read again will encounter end of file.

Here is an example showing the use of getc and ungetc to skip over whitespace characters. When this function reaches a non-whitespace character, it unreads that character to be seen again on the next read operation on the stream.

  #include <stdio.h>
  #include <ctype.h>
  void
  skip_whitespace (FILE *stream)
  {
    int c;
    do

/* No need to check for EOF because it is not

         isspace, and ungetc ignores EOF. */
      c = getc (stream);
    while (isspace (c));
    ungetc (c, stream);
  }

7.9 Block Input/Output

This section describes how to do input and output operations on blocks of data. You can use these functions to read and write binary data, as well as to read and write text in fixed-size blocks instead of by characters or lines.

Binary files are typically used to read and write blocks of data in the same format as is used to represent the data in a running program. In other words, arbitrary blocks of memory—not just character or string objects— can be written to a binary file, and meaningfully read in again by the same program.

Storing data in binary form is often considerably more efficient than using the formatted I/O functions. Also, for floating-point numbers, the binary form avoids possible loss of precision in the conversion process. On the other hand, binary files can’t be examined or modified easily using many standard file utilities (such as text editors), and are not portable between different implementations of the language, or different kinds of computers.

These functions are declared in ‘stdio.h’.

Function size_t fread (void *data, size_t size, size_t count,

FILE *stream) This function reads up to count objects of size size into the array data, from the stream stream. It returns the number of objects actually read, which might be less than count if a read error occurs or the end of the file is reached. This function returns a value of zero (and doesn’t read anything) if either size or count is zero. If fread encounters end of file in the middle of an object, it returns the number of complete objects read, and discards the partial object. Therefore, the stream remains at the actual end of the file.

Function size_t fwrite (const void *data, size_t size, size_t

count, FILE *stream) This function writes up to count objects of size size from the array data, to the stream stream. The return value is normally count, if the call succeeds. Any other value indicates some sort of error, such as running out of space.

7.10 Formatted Output

The functions described in this section (printf and related functions) provide a convenient way to perform formatted output. You call printf with a format string or template string that specifies how to format the values of the remaining arguments.

Unless your program is a filter that specifically performs line- or character-oriented processing, using printf or one of the other related functions described in this section is usually the easiest and most concise way to perform output. These functions are especially useful for printing error messages, tables of data, and the like.

7.10.1 Formatted Output Basics

The printf function can be used to print any number of arguments. The template string argument you supply in a call provides information not only about the number of additional arguments, but also about their types and what style should be used for printing them.

Ordinary characters in the template string are simply written to the output stream as-is, while conversion specifications introduced by a ‘%’ character in the template cause subsequent arguments to be formatted and written to the output stream. For example,

     int pct = 37;
     char filename[] = "foo.txt";
     printf ("Processing of `%s' is %d%% finished.\nPlease be patient.\n",
             filename, pct);

produces output like

     Processing of `foo.txt' is 37% finished.
     Please be patient.

This example shows the use of the ‘%d’ conversion to specify that an int argument should be printed in decimal notation, the ‘%s’ conversion to specify printing of a string argument, and the ‘%%’ conversion to print a literal ‘%’ character.

There are also conversions for printing an integer argument as an unsigned value in octal, decimal, or hexadecimal radix (‘%o’, ‘%u’, or ‘%x’, respectively); or as a character value (‘%c’).

Floating-point numbers can be printed in normal, fixed-point notation using the ‘%f’ conversion or in exponential notation using the ‘%e’ conversion. The ‘%g’ conversion uses either ‘%e’ or ‘%f’ format, depending on what is more appropriate for the magnitude of the particular number.

You can control formatting more precisely by writing modifiers between the ‘%’ and the character that indicates which conversion to apply. These slightly alter the ordinary behavior of the conversion. For example, most conversion specifications permit you to specify a minimum field width and a flag indicating whether you want the result left- or right-justified within the field.

The specific flags and modifiers that are permitted and their interpretation vary depending on the particular conversion. They’re all described in more detail in the following sections. Don’t worry if this all seems excessively complicated at first; you can almost always get reasonable free-format output without using any of the modifiers at all. The modifiers are mostly used to make the output look “prettier” in tables.

7.10.2 Output Conversion Syntax

This section provides details about the precise syntax of conversion specifications that can appear in a printf template string.

Characters in the template string that are not part of a conversion specification are printed as-is to the output stream. Multibyte character sequences (see Chapter 18 [Extended Characters], page 367) are permitted in a template string.

The conversion specifications in a printf template string have the general form:

% flags width [ . precision ] type conversion For example, in the conversion specifier ‘%-10.8ld’, the ‘-’ is a flag, ‘10’

specifies the field width, the precision is ‘8’, the letter ‘l’ is a type modifier, and ‘d’ specifies the conversion style. (This particular type specifier says to print a long int argument in decimal notation, with a minimum of 8 digits left-justified in a field at least 10 characters wide.)

In more detail, output conversion specifications consist of an initial ‘%’ character followed in sequence by:

• Zero or more flag characters that modify the normal behavior of the conversion specification.

• An optional decimal integer specifying the minimum field width. If the normal conversion produces fewer characters than this, the field is padded with spaces to the specified width. This is a minimum value; if the normal conversion produces more characters than this, the field is not truncated. Normally, the output is right-justified within the field. You can also specify a field width of ‘*’. This means that the next argument in the argument list (before the actual value to be printed) is used as the field width. The value must be an int. If the value is negative, this means to set the ‘-’ flag (see below) and to use the absolute value as the field width.

• An optional precision to specify the number of digits to be written for the numeric conversions. If the precision is specified, it consists of a period (‘.’) followed optionally by a decimal integer (which defaults to zero if omitted). You can also specify a precision of ‘*’. This means that the next argument in the argument list (before the actual value to be printed) is used as the precision. The value must be an int, and is ignored if it is negative. If you specify ‘*’ for both the field width and precision, the field width argument precedes the precision argument. Other C library versions may not recognize this syntax.

• An optional type modifier character, which is used to specify the data type of the corresponding argument if it differs from the default type. (For example, the integer conversions assume a type of int, but you can specify ‘h’, ‘l’, or ‘L’ for other integer types.)

• A character that specifies the conversion to be applied.

The exact options that are permitted and how they are interpreted vary between the different conversion specifiers. See the descriptions of the individual conversions for information about the particular options that they use.

With the ‘-Wformat’ option, the GNU C compiler checks calls to printf and related functions. It examines the format string and verifies that the correct number and types of arguments are supplied. There is also a GNU C syntax to tell the compiler that a function you write uses a printf-style format string. See section “Declaring Attributes of Functions” in Using GNU CC, for more information.

7.10.3 Table of Output Conversions

Here is a table summarizing what all the different conversions do:

‘%d’, ‘%i’ Print an integer as a signed decimal number. See Section 7.10.4 [Integer Conversions], page 108, for details. ‘%d’ and ‘%i’ are synonymous for output, but are different when used with scanf for input (see Section 7.12.3 [Table of Input Conversions], page 129).

‘%o’ Print an integer as an unsigned octal number. See Section 7.10.4 [Integer Conversions], page 108, for details.

‘%u’ Print an integer as an unsigned decimal number. See Section 7.10.4 [Integer Conversions], page 108, for details.

‘%x’, ‘%X’ Print an integer as an unsigned hexadecimal number. ‘%x’ uses lower-case letters and ‘%X’ uses upper-case. See Section 7.10.4 [Integer Conversions], page 108, for details.

‘%f’ Print a floating-point number in normal (fixed-point) notation. See Section 7.10.5 [Floating-Point Conversions], page 110, for details.

‘%e’, ‘%E’ Print a floating-point number in exponential notation. ‘%e’ uses lower-case letters and ‘%E’ uses upper-case. See Section 7.10.5 [Floating-Point Conversions], page 110, for details.

‘%g’, ‘%G’ Print a floating-point number in either normal or exponential notation, whichever is more appropriate for its magnitude. ‘%g’ uses lower-case letters and ‘%G’ uses upper-case. See Section 7.10.5 [Floating-Point Conversions], page 110, for details.

‘%c’ Print a single character. See Section 7.10.6 [Other Output Conversions], page 112.

‘%s’ Print a string. See Section 7.10.6 [Other Output Conversions], page 112.

‘%p’ Print the value of a pointer. See Section 7.10.6 [Other Output Conversions], page 112.

‘%n’ Get the number of characters printed so far. See Section 7.10.6 [Other Output Conversions], page 112. Note that this conversion specification never produces any output.

‘%m’ Print the string corresponding to the value of errno. (This is a GNU extension.) See Section 7.10.6 [Other Output Conversions], page 112.

‘%%’ Print a literal ‘%’ character. See Section 7.10.6 [Other Output Conversions], page 112.

If the syntax of a conversion specification is invalid, unpredictable things will happen, so don’t do this. If there aren’t enough function arguments provided to supply values for all the conversion specifications in the template string, or if the arguments are not of the correct types, the results are unpredictable. If you supply more arguments than conversion specifications, the extra argument values are simply ignored; this is sometimes useful.

7.10.4 Integer Conversions

This section describes the options for the ‘%d’, ‘%i’, ‘%o’, ‘%u’, ‘%x’, and ‘%X’ conversion specifications. These conversions print integers in various formats.

The ‘%d’ and ‘%i’ conversion specifications both print an int argument as a signed decimal number; while ‘%o’, ‘%u’, and ‘%x’ print the argument as an unsigned octal, decimal, or hexadecimal number (respectively). The ‘%X’ conversion specification is just like ‘%x’ except that it uses the characters ‘ABCDEF’ as digits instead of ‘abcdef’.

The following flags are meaningful:

‘-’ Left-justify the result in the field (instead of the normal rightjustification).

‘+’ For the signed ‘%d’ and ‘%i’ conversions, print a plus sign if the value is positive.

‘ ’ For the signed ‘%d’ and ‘%i’ conversions, if the result doesn’t start with a plus or minus sign, prefix it with a space character instead. Since the ‘+’ flag ensures that the result includes a sign, this flag is ignored if you supply both of them.

‘#’ For the ‘%o’ conversion, this forces the leading digit to be ‘0’, as if by increasing the precision. For ‘%x’ or ‘%X’, this prefixes a leading ‘0x’ or ‘0X’ (respectively) to the result. This doesn’t do anything useful for the ‘%d’, ‘%i’, or ‘%u’ conversions. Using this flag produces output which can be parsed by the strtoul function (see Section 14.7.1 [Parsing of Integers], page 310) and scanf with the ‘%i’ conversion (see Section 7.12.4 [Numeric Input Conversions], page 130).

‘'’ Separate the digits into groups as specified by the locale specified for the LC_NUMERIC category; see Section 19.6.1 [Generic Numeric Formatting Parameters], page 384. This flag is a GNU extension.

‘0’ Pad the field with zeros instead of spaces. The zeros are placed after any indication of sign or base. This flag is ignored if the ‘-’ flag is also specified, or if a precision is specified.

If a precision is supplied, it specifies the minimum number of digits to appear; leading zeros are produced if necessary. If you don’t specify a precision, the number is printed with as many digits as it needs. If you convert a value of zero with an explicit precision of zero, then no characters at all are produced.

Without a type modifier, the corresponding argument is treated as an int (for the signed conversions ‘%i’ and ‘%d’) or unsigned int (for the unsigned conversions ‘%o’, ‘%u’, ‘%x’, and ‘%X’). Recall that since printf and friends are variadic, any char and short arguments are automatically converted to int by the default argument promotions. For arguments of other integer types, you can use these modifiers: ‘h’ Specifies that the argument is a short int or unsigned short

int, as appropriate. A short argument is converted to an int or unsigned int by the default argument promotions anyway, but the ‘h’ modifier says to convert it back to a short again.

‘l’ Specifies that the argument is a long int or unsigned long int, as appropriate. Two ‘l’ characters is like the ‘L’ modifier, below.

‘L’

‘ll’

‘q’ Specifies that the argument is a long long int. (This type is an extension supported by the GNU C compiler. On systems that don’t support extra-long integers, this is the same as long

int.)

The ‘q’ modifier is another name for the same thing, which comes from 4.4 BSD; a long long int is sometimes called a “quad”

int.

‘Z’ Specifies that the argument is a size_t. This is a GNU extension.

Here is an example. Using the template string:

     "|%5d|%-5d|%+5d|%+-5d|% 5d|%05d|%5.0d|%5.2d|%d|\n"

to print numbers using the different options for the ‘%d’ conversion gives results like:

     |    0|0    |   +0|+0   |    0|00000|     |   00|0|
     |    1|1    |   +1|+1   |    1|00001|    1|   01|1|
     |   -1|-1   |   -1|-1   |   -1|-0001|   -1|  -01|-1|
     |100000|100000|+100000| 100000|100000|100000|100000|100000|

In particular, notice what happens in the last case where the number is too large to fit in the minimum field width specified.

Here are some more examples showing how unsigned integers print under various format options, using the template string:

     "|%5u|%5o|%5x|%5X|%#5o|%#5x|%#5X|%#10.8x|\n"
     |    0|    0|    0|    0|    0|  0x0|  0X0|0x00000000|
     |    1|    1|    1|    1|   01|  0x1|  0X1|0x00000001|
     |100000|303240|186a0|186A0|0303240|0x186a0|0X186A0|0x000186a0|

7.10.5 Floating-Point Conversions

This section discusses the conversion specifications for floating-point numbers: the ‘%f’, ‘%e’, ‘%E’, ‘%g’, and ‘%G’ conversions.

The ‘%f’ conversion prints its argument in fixed-point notation, producing output of the form [-]ddd.ddd, where the number of digits following the decimal point is controlled by the precision you specify.

The ‘%e’ conversion prints its argument in exponential notation, producing output of the form [-]d.ddde[+|-]dd. Again, the number of digits following the decimal point is controlled by the precision. The exponent always contains at least two digits. The ‘%E’ conversion is similar but the exponent is marked with the letter ‘E’ instead of ‘e’.

The ‘%g’ and ‘%G’ conversions print the argument in the style of ‘%e’ or ‘%E’ (respectively) if the exponent would be less than -4 or greater than or equal to the precision; otherwise they use the ‘%f’ style. Trailing zeros are removed from the fractional portion of the result and a decimal-point character appears only if it is followed by a digit.

The following flags can be used to modify the behavior:

‘-’ Left-justify the result in the field. Normally the result is rightjustified.

‘+’ Always include a plus or minus sign in the result.

‘ ’ If the result doesn’t start with a plus or minus sign, prefix it with a space instead. Since the ‘+’ flag ensures that the result includes a sign, this flag is ignored if you supply both of them.

‘#’ Specifies that the result should always include a decimal point, even if no digits follow it. For the ‘%g’ and ‘%G’ conversions, this also forces trailing zeros after the decimal point to be left in place where they would otherwise be removed.

‘'’ Separate the digits of the integer part of the result into groups as specified by the locale specified for the LC_NUMERIC category; see Section 19.6.1 [Generic Numeric Formatting Parameters], page 384. This flag is a GNU extension.

‘0’ Pad the field with zeros instead of spaces; the zeros are placed after any sign. This flag is ignored if the ‘-’ flag is also specified.

The precision specifies how many digits follow the decimal-point character for the ‘%f’, ‘%e’, and ‘%E’ conversions. For these conversions, the default precision is 6. If the precision is explicitly 0, this suppresses the decimal point character entirely. For the ‘%g’ and ‘%G’ conversions, the precision specifies how many significant digits to print. Significant digits are the first digit before the decimal point, and all the digits after it. If the precision 0 or not specified for ‘%g’ or ‘%G’, it is treated like a value of 1. If the value being printed cannot be expressed accurately in the specified number of digits, the value is rounded to the nearest number that fits.

Without a type modifier, the floating-point conversions use an argument of type double. (By the default argument promotions, any float arguments are automatically converted to double.) The following type modifier is supported: ‘L’ An uppercase ‘L’ specifies that the argument is a long double.

Here are some examples showing how numbers print using the various floating-point conversions. All of the numbers were printed using this template string:

     "|%12.4f|%12.4e|%12.4g|\n"

Here is the output:

     |      0.0000|  0.0000e+00|           0|
     |      1.0000|  1.0000e+00|           1|
     |     -1.0000| -1.0000e+00|          -1|
     |    100.0000|  1.0000e+02|         100|
     |   1000.0000|  1.0000e+03|        1000|
     |  10000.0000|  1.0000e+04|       1e+04|
     |  12345.0000|  1.2345e+04|   1.234e+04|
     | 100000.0000|  1.0000e+05|       1e+05|
     | 123456.0000|  1.2346e+05|   1.234e+05|

Notice how the ‘%g’ conversion drops trailing zeros.

7.10.6 Other Output Conversions

This section describes miscellaneous conversions for printf.

The ‘%c’ conversion prints a single character. The int argument is first converted to an unsigned char. The ‘-’ flag can be used to specify leftjustification in the field, but no other flags are defined, and no precision or type modifier can be given. For example:

     printf ("%c%c%c%c%c", 'h', 'e', 'l', 'l', 'o');

prints ‘hello’.

The ‘%s’ conversion prints a string. The corresponding argument must be of type char * (or const char *). A precision can be specified to indicate the maximum number of characters to write; otherwise characters in the string up to but not including the terminating null character are written to the output stream. The ‘-’ flag can be used to specify left-justification in the field, but no other flags or type modifiers are defined for this conversion. For example:

     printf ("%3s%-6s", "no", "where");

prints ‘ nowhere ’.

If you accidentally pass a null pointer as the argument for a ‘%s’ conversion, the GNU library prints it as ‘(null)’. We think this is more useful than crashing. But it’s not good practice to pass a null argument intentionally.

The ‘%m’ conversion prints the string corresponding to the error code in errno. See Section 2.3 [Error Messages], page 27. Thus:

     fprintf (stderr, "can't open `%s': %m\n", filename);

is equivalent to:

     fprintf (stderr, "can't open `%s': %s\n", filename, strerror (errno));

The ‘%m’ conversion is a GNU C library extension.

The ‘%p’ conversion prints a pointer value. The corresponding argument must be of type void *. In practice, you can use any type of pointer.

In the GNU system, non-null pointers are printed as unsigned integers, as if a ‘%#x’ conversion were used. Null pointers print as ‘(nil)’. (Pointers might print differently in other systems.)

For example:

     printf ("%p", "testing");

prints ‘0x’ followed by a hexadecimal number—the address of the string constant "testing". It does not print the word ‘testing’.

You can supply the ‘-’ flag with the ‘%p’ conversion to specify leftjustification, but no other flags, precision, or type modifiers are defined.

The ‘%n’ conversion is unlike any of the other output conversions. It uses an argument which must be a pointer to an int, but instead of printing anything it stores the number of characters printed so far by this call at that location. The ‘h’ and ‘l’ type modifiers are permitted to specify that the argument is of type short int * or long int * instead of int *, but no flags, field width, or precision are permitted.

For example,

  int nchar;
  printf ("%d %s%n\n", 3, "bears", &nchar);

prints:

  3 bears

and sets nchar to 7, because ‘3 bears’ is seven characters.

The ‘%%’ conversion prints a literal ‘%’ character. This conversion doesn’t use an argument, and no flags, field width, precision, or type modifiers are permitted.

7.10.7 Formatted Output Functions

This section describes how to call printf and related functions. Prototypes for these functions are in the header file ‘stdio.h’. Because these functions take a variable number of arguments, you must declare prototypes for them before using them. Of course, the easiest way to make sure you have all the right prototypes is to just include ‘stdio.h’.

Function int printf (const char *template, …)

The printf function prints the optional arguments under the control of the template string template to the stream stdout. It returns the number of characters printed, or a negative value if there was an output error.

Function int fprintf (FILE *stream, const char *template, …)

This function is just like printf, except that the output is written to the stream stream instead of stdout.

Function int sprintf (char *s, const char *template, …)

This is like printf, except that the output is stored in the character array s instead of written to a stream. A null character is written to mark the end of the string. The sprintf function returns the number of characters stored in the array s, not including the terminating null character. The behavior of this function is undefined if copying takes place between objects that overlap—for example, if s is also given as an argument to be printed under control of the ‘%s’ conversion. See Section 5.4 [Copying and Concatenation], page 67. Warning: The sprintf function can be dangerous because it can potentially output more characters than can fit in the allocation size of the string s. Remember that the field width given in a conversion specification is only a minimum value. To avoid this problem, you can use snprintf or asprintf, described below.

Function int snprintf (char *s, size_t size, const char

*template, …) The snprintf function is similar to sprintf, except that the size argument specifies the maximum number of characters to produce. The trailing null character is counted towards this limit, so you should allocate at least size characters for the string s. The return value is the number of characters stored, not including the terminating null. If this value equals size - 1, then there was not enough space in s for all the output. You should try again with a bigger output string. Here is an example of doing this:

/* Construct a message describing the value of a variable whose name is name and whose value is value. */

     char *
     make_message (char *name, char *value)
     {

/* Guess we need no more than 100 chars of space. */

       int size = 100;
       char *buffer = (char *) xmalloc (size);
       while (1)
         {

/* Try to print in the allocated space. */

           int nchars = snprintf (buffer, size,
                                  "value of %s is %s",
                                  name, value);

/* If that worked, return the string. */

           if (nchars < size)
             return buffer;

/* Else try again with twice as much space. */

           size *= 2;
           buffer = (char *) xrealloc (size, buffer);
         }
     }

In practice, it is often easier just to use asprintf, below.

7.10.8 Dynamically Allocating Formatted Output

The functions in this section do formatted output and place the results in dynamically allocated memory.

Function int asprintf (char **ptr, const char *template, …)

This function is similar to sprintf, except that it dynamically allocates a string (as with malloc; see Section 3.3 [Unconstrained Allocation], page 32) to hold the output, instead of putting the output in a buffer you allocate in advance. The ptr argument should be the address of a char * object, and asprintf stores a pointer to the newly allocated string at that location. Here is how to use asprintf to get the same result as the snprintf example, but more easily:

/* Construct a message describing the value of a variable whose name is name and whose value is value. */

        char *
        make_message (char *name, char *value)
        {
          char *result;
          asprintf (&result, "value of %s is %s", name, value);
          return result;
        }

Function int obstack printf (struct obstack *obstack, const

char *template, …) This function is similar to asprintf, except that it uses the obstack obstack to allocate the space. See Section 3.4 [Obstacks], page 43. The characters are written onto the end of the current object. To get at them, you must finish the object with obstack_finish (see Section 3.4.6 [Growing Objects], page 47).

7.10.9 Variable Arguments Output Functions

The functions vprintf and friends are provided so that you can define your own variadic printf-like functions that make use of the same internals as the built-in formatted output functions.

The most natural way to define such functions would be to use a language construct to say, “Call printf and pass this template plus all of my arguments after the first five.” But there is no way to do this in C, and it would be hard to provide a way, since at the C language level there is no way to tell how many arguments your function received.

Since that method is impossible, we provide alternative functions, the vprintf series, which lets you pass a va_list to describe “all of my arguments after the first five.”

When it is sufficient to define a macro rather than a real function, the GNU C compiler provides a way to do this much more easily with macros. For example:

     #define myprintf(a, b, c, d, e, rest...) printf (mytemplate , ## rest...)

See section “Macros with Variable Numbers of Arguments” in Using GNU CC, for details. But this is limited to macros, and does not apply to real functions at all.

Before calling vprintf or the other functions listed in this section, you must call va_start (see Section A.2 [Variadic Functions], page 555) to initialize a pointer to the variable arguments. Then you can call va_arg to fetch the arguments that you want to handle yourself. This advances the pointer past those arguments.

Once your va_list pointer is pointing at the argument of your choice, you are ready to call vprintf. That argument and all subsequent arguments that were passed to your function are used by vprintf along with the template that you specified separately.

In some other systems, the va_list pointer may become invalid after the call to vprintf, so you must not use va_arg after you call vprintf. Instead, you should call va_end to retire the pointer from service. However, you can safely call va_start on another pointer variable and begin fetching the arguments again through that pointer. Calling vprintf does not destroy the argument list of your function, merely the particular pointer that you passed to it.

GNU C does not have such restrictions. You can safely continue to fetch arguments from a va_list pointer after passing it to vprintf, and va_end is a no-op. (Note, however, that subsequent va_arg calls will fetch the same arguments which vprintf previously used.)

Prototypes for these functions are declared in ‘stdio.h’.

Function int vprintf (const char *template, va_list ap)

This function is similar to printf except that, instead of taking a variable number of arguments directly, it takes an argument list pointer ap.

Function int vfprintf (FILE *stream, const char *template,

         va_list ap)

This is the equivalent of fprintf with the variable argument list specified directly as for vprintf.

Function int vsprintf (char *s, const char *template, va_list

ap) This is the equivalent of sprintf with the variable argument list specified directly as for vprintf.

Function int vsnprintf (char *s, size_t size, const char

*template, va_list ap) This is the equivalent of snprintf with the variable argument list specified directly as for vprintf.

Function int vasprintf (char **ptr, const char *template,

         va_list ap)

The vasprintf function is the equivalent of asprintf with the variable argument list specified directly as for vprintf.

Function int obstack vprintf (struct obstack *obstack, const

         char *template, va_list ap)
  The obstack_vprintf function is the equivalent of obstack_printf

with the variable argument list specified directly as for vprintf.

Here’s an example showing how you might use vfprintf. This is a function that prints error messages to the stream stderr, along with a prefix indicating the name of the program (see Section 2.3 [Error Messages], page 27,

for a description of program_invocation_short_name).
     #include <stdio.h>
     #include <stdarg.h>
     void
     eprintf (const char *template, ...)
     {
       va_list ap;
       extern char *program_invocation_short_name;
       fprintf (stderr, "%s: ", program_invocation_short_name);
       va_start (ap, count);
       vfprintf (stderr, template, ap);
       va_end (ap);
     }

You could call eprintf like this:

     eprintf ("file `%s' does not exist\n", filename);

In GNU C, there is a special construct you can use to let the compiler know that a function uses a printf-style format string. Then it can check the number and types of arguments in each call to the function, and warn you when they do not match the format string. For example, take this declaration of eprintf:

     void eprintf (const char *template, ...)
             __attribute__((format (printf, 1, 2)));

This tells the compiler that eprintf uses a format string like printf (as opposed to scanf; see Section 7.12 [Formatted Input], page 126); the format string appears as the first argument; and the arguments to satisfy the format begin with the second. See section “Declaring Attributes of Functions” in Using GNU CC, for more information.

7.10.10 Parsing a Template String

You can use the function parse_printf_format to obtain information about the number and types of arguments that are expected by a given template string. This function permits interpreters that provide interfaces to printf to avoid passing along invalid arguments from the user’s program, which could cause a crash.

All the symbols described in this section are declared in the header file

‘printf.h’.

Function size_t parse printf format (const char *template,

      size_t n, int *argtypes)

This function returns information about the number and types of arguments expected by the printf template string template. The information is stored in the array argtypes; each element of this array describes one argument. This information is encoded using the various ‘PA_’ macros, listed below. The n argument specifies the number of elements in the array argtypes. This is the most elements that parse_printf_format will try to write. parse_printf_format returns the total number of arguments required by template. If this number is greater than n, then the information returned describes only the first n arguments. If you want information about more than that many arguments, allocate a bigger array and

call parse_printf_format again.

The argument types are encoded as a combination of a basic type and modifier flag bits.

Macro int PA FLAG MASK

This macro is a bitmask for the type modifier flag bits. You can write the expression (argtypes[i] & PA_FLAG_MASK) to extract just the flag bits for an argument, or (argtypes[i] & ~PA_FLAG_MASK) to extract just the basic type code.

Here are symbolic constants that represent the basic types; they stand for integer values.

PA_INT This specifies that the base type is int.

PA_CHAR This specifies that the base type is int, cast to char.

PA_STRING

This specifies that the base type is char *, a null-terminated string.

PA_POINTER

This specifies that the base type is void *, an arbitrary pointer.

PA_FLOAT This specifies that the base type is float.

PA_DOUBLE

This specifies that the base type is double.

PA_LAST You can define additional base types for your own programs as offsets from PA_LAST. For example, if you have data types ‘foo’ and ‘bar’ with their own specialized printf conversions, you could define encodings for these types as:

     #define PA_FOO  PA_LAST
     #define PA_BAR  (PA_LAST + 1)

Here are the flag bits that modify a basic type. They are combined with the code for the basic type using inclusive-or.

PA_FLAG_PTR

If this bit is set, it indicates that the encoded type is a pointer to the base type, rather than an immediate value. For example,

‘PA_INT|PA_FLAG_PTR’ represents the type ‘int *’.
PA_FLAG_SHORT

If this bit is set, it indicates that the base type is modified with short. (This corresponds to the ‘h’ type modifier.)

PA_FLAG_LONG

If this bit is set, it indicates that the base type is modified with long. (This corresponds to the ‘l’ type modifier.)

PA_FLAG_LONG_LONG

If this bit is set, it indicates that the base type is modified with long long. (This corresponds to the ‘L’ type modifier.)

PA_FLAG_LONG_DOUBLE

This is a synonym for PA_FLAG_LONG_LONG, used by convention with a base type of PA_DOUBLE to indicate a type of long double.

7.10.11 Example of Parsing a Template String

Here is an example of decoding argument types for a format string. We assume this is part of an interpreter which contains arguments of type NUMBER, CHAR, STRING and STRUCTURE (and perhaps others which are not valid here).

/* Test whether the nargs specified objects in the vector args are valid for the format string format: if so, return 1. If not, return 0 after printing an error message. */

int
validate_args (char *format, int nargs, OBJECT *args)
{

120 The GNU C Library

int *argtypes;
int nwanted;

/* Get the information about the arguments.

Each conversion specification must be at least two characters long, so there cannot be more specifications than half the length of the string. */

argtypes = (int *) alloca (strlen (format) / 2 * sizeof (int));
nwanted = parse_printf_format (string, nelts, argtypes);

/* Check the number of arguments. */

if (nwanted > nargs)
  {
    error ("too few arguments (at least %d required)", nwanted);
    return 0;
  }

/* Check the C type wanted for each argument

and see if the object given is suitable. */

for (i = 0; i < nwanted; i++)
  {
    int wanted;
    if (argtypes[i] & PA_FLAG_PTR)
      wanted = STRUCTURE;
    else
      switch (argtypes[i] & ~PA_FLAG_MASK)
        {
        case PA_INT:
        case PA_FLOAT:
        case PA_DOUBLE:
          wanted = NUMBER;
          break;
        case PA_CHAR:
          wanted = CHAR;
          break;
        case PA_STRING:
          wanted = STRING;
          break;
        case PA_POINTER:
          wanted = STRUCTURE;
          break;
        }

           if (TYPE (args[i]) != wanted)
             {
               error ("type mismatch for arg number %d", i);
               return 0;
             }
         }
       return 1;
     }

7.11 Customizing printf

The GNU C library lets you define your own custom conversion specifiers for printf template strings, to teach printf clever ways to print the important data structures of your program.

The way you do this is by registering the conversion with the function register_printf_function; see Section 7.11.1 [Registering New Conversions], page 121. One of the arguments you pass to this function is a pointer to a handler function that produces the actual output; see Section 7.11.3 [Defining the Output Handler], page 123, for information on how to write this function.

You can also install a function that just returns information about the number and type of arguments expected by the conversion specifier. See Section 7.10.10 [Parsing a Template String], page 118, for information about this.

The facilities of this section are declared in the header file ‘printf.h’. Portability Note: The ability to extend the syntax of printf template

strings is a GNU extension. ISO standard C has nothing similar.

7.11.1 Registering New Conversions

The function to register a new output conversion is register_printf_

function, declared in ‘printf.h’.

Function int register printf function (int spec,

         printf_function handler-function,
         printf_arginfo_function arginfo-function)

This function defines the conversion specifier character spec. Thus, if spec is 'z', it defines the conversion ‘%z’. You can redefine the builtin conversions like ‘%s’, but flag characters like ‘#’ and type modifiers like ‘l’ can never be used as conversions; calling register_printf_ function for those characters has no effect. The handler-function is the function called by printf and friends when this conversion appears in a template string. See Section 7.11.3 [Defining the Output Handler], page 123, for information about how to define a function to pass as this argument. If you specify a null pointer, any existing handler function for spec is removed. The arginfo-function is the function called by parse_printf_format when this conversion appears in a template string. See Section 7.10.10 [Parsing a Template String], page 118, for information about this. Attention: In the GNU C library version before 2.0 the arginfofunction function did not need to be installed unless the user uses the parse_printf_format function. This changed. Now a call to any of the printf functions will call this function when this format specifier appears in the format string. The return value is 0 on success, and -1 on failure (which occurs if spec is out of range). You can redefine the standard output conversions, but this is probably not a good idea because of the potential for confusion. Library routines written by other people could break if you do this.

7.11.2 Conversion Specifier Options

If you define a meaning for ‘%A’, what if the template contains ‘%+23A’ or ‘%-#A’? To implement a sensible meaning for these, the handler when called needs to be able to get the options specified in the template.

Both the handler-function and arginfo-function arguments to register_ printf_function accept an argument that points to a struct printf_ info, which contains information about the options appearing in an instance of the conversion specifier. This data type is declared in the header

file ‘printf.h’.

Type struct printf info

This structure is used to pass information about the options appearing in an instance of a conversion specifier in a printf template string to the handler and arginfo functions for that specifier. It contains the following members: int prec This is the precision specified. The value is -1 if no pre-

cision was specified. If the precision was given as ‘*’, the printf_info structure passed to the handler function contains the actual value retrieved from the argument list. But the structure passed to the arginfo function contains a value of INT_MIN, since the actual value is not known.

int width This is the minimum field width specified. The value is 0 if no width was specified. If the field width was given as ‘*’, the printf_info structure passed to the handler function contains the actual value retrieved from the argument list. But the structure passed to the arginfo function contains a value of INT_MIN, since the actual value is not known.

wchar_t spec

This is the conversion specifier character specified. It’s stored in the structure so that you can register the same handler function for multiple characters, but still have a way to tell them apart when the handler function is called.

unsigned int is_long_double

This is a boolean that is true if the ‘L’, ‘ll’, or ‘q’ type modifier was specified. For integer conversions, this indicates long long int, as opposed to long double for floating point conversions.

unsigned int is_short

This is a boolean that is true if the ‘h’ type modifier was specified.

unsigned int is_long

This is a boolean that is true if the ‘l’ type modifier was specified.

unsigned int alt

This is a boolean that is true if the ‘#’ flag was specified.

unsigned int space

This is a boolean that is true if the ‘ ’ flag was specified.

unsigned int left

This is a boolean that is true if the ‘-’ flag was specified.

unsigned int showsign

This is a boolean that is true if the ‘+’ flag was specified.

unsigned int group

This is a boolean that is true if the ‘'’ flag was specified.

unsigned int extra

This flag has a special meaning depending on the context. It could be used freely by the user-defined handlers but when called from the printf function this variable always contains the value 0.

wchar_t pad

This is the character to use for padding the output to the minimum field width. The value is '0' if the ‘0’ flag was specified, and ' ' otherwise.

7.11.3 Defining the Output Handler

Now let’s look at how to define the handler and arginfo functions which

are passed as arguments to register_printf_function.

Compatibility Note: The interface change in the GNU libc version 2.0. Previously the third argument was of type va_list *.

You should define your handler functions with a prototype like:

     int function (FILE *stream, const struct printf_info *info,
                         const void *const *args)

The stream argument passed to the handler function is the stream to which it should write output.

The info argument is a pointer to a structure that contains information about the various options that were included with the conversion in the template string. You should not modify this structure inside your handler function. See Section 7.11.2 [Conversion Specifier Options], page 122, for a description of this data structure.

The args is a vector of pointers to the arguments data. The number of arguments were determined by calling the argument information function provided by the user.

Your handler function should return a value just like printf does: it should return the number of characters it has written, or a negative value to indicate an error.

Data Type printf function

This is the data type that a handler function should have.

If you are going to use parse_printf_format in your application, you must also define a function to pass as the arginfo-function argument for each new conversion you install with register_printf_function.

You have to define these functions with a prototype like:

     int function (const struct printf_info *info,
                         size_t n, int *argtypes)

The return value from the function should be the number of arguments the conversion expects. The function should also fill in no more than n elements of the argtypes array with information about the types of each of these arguments. This information is encoded using the various ‘PA_’ macros. (You will notice that this is the same calling convention parse_

printf_format itself uses.)

Data Type printf arginfo function

This type is used to describe functions that return information about the number and type of arguments used by a conversion specifier.

7.11.4 printf Extension Example

Here is an example showing how to define a printf handler function. This program defines a data structure called a Widget and defines the ‘%W’ conversion to print information about Widget * arguments, including the Chapter 7: Input/Output on Streams 125

pointer value and the name stored in the data structure. The ‘%W’ conversion supports the minimum field width and left-justification options, but ignores everything else.

#include <stdio.h>
#include <printf.h>
#include <stdarg.h>
typedef struct
{
  char *name;
} Widget;
int
print_widget (FILE *stream, const struct printf_info *info, va_list *app)
{
Widget *w;
char *buffer;
int len;

/* Format the output into a string. */

w = va_arg (*app, Widget *);
len = asprintf (&buffer, "<Widget %p: %s>", w, w->name);
if (len == -1)
  return -1;

/* Pad to the minimum field width and print to the stream. */

len = fprintf (stream, "%*s",
               (info->left ? - info->width : info->width),
               buffer);

/* Clean up and return. */

free (buffer);
return len;
}

int
print_widget_arginfo (const struct printf_info *info, size_t n,
                    int *argtypes)
{

/* We always take exactly one argument and this is a pointer to the structure.. */

if (n > 0)
  argtypes[0] = PA_POINTER;
return 1;

     }

     int
     main (void)
     {

/* Make a widget to print. */

       Widget mywidget;
       mywidget.name = "mywidget";

/* Register the print function for widgets. */

       register_printf_function ('W', print_widget, print_widget_arginfo);

/* Now print the widget. */

       printf ("|%W|\n", &mywidget);
       printf ("|%35W|\n", &mywidget);
       printf ("|%-35W|\n", &mywidget);
       return 0;
     }

The output produced by this program looks like:

     |<Widget 0xffeffb7c: mywidget>|
     |      <Widget 0xffeffb7c: mywidget>|
     |<Widget 0xffeffb7c: mywidget>      |

7.12 Formatted Input

The functions described in this section (scanf and related functions) provide facilities for formatted input analogous to the formatted output facilities. These functions provide a mechanism for reading arbitrary values under the control of a format string or template string.

7.12.1 Formatted Input Basics

Calls to scanf are superficially similar to calls to printf in that arbitrary arguments are read under the control of a template string. While the syntax of the conversion specifications in the template is very similar to that for printf, the interpretation of the template is oriented more towards free-format input and simple pattern matching, rather than fixed-field formatting. For example, most scanf conversions skip over any amount of “white space” (including spaces, tabs, and newlines) in the input file, and there is no concept of precision for the numeric input conversions as there is for the corresponding output conversions. Ordinarily, non-whitespace characters in the template are expected to match characters in the input stream exactly, but a matching failure is distinct from an input error on the stream.

Another area of difference between scanf and printf is that you must remember to supply pointers rather than immediate values as the optional arguments to scanf; the values that are read are stored in the objects that the pointers point to. Even experienced programmers tend to forget this occasionally, so if your program is getting strange errors that seem to be related to scanf, you might want to double-check this.

When a matching failure occurs, scanf returns immediately, leaving the first non-matching character as the next character to be read from the stream. The normal return value from scanf is the number of values that were assigned, so you can use this to determine if a matching error happened before all the expected values were read.

The scanf function is typically used for things like reading in the contents of tables. For example, here is a function that uses scanf to initialize an array of double:

     void
     readarray (double *array, int n)
     {
       int i;
       for (i=0; i<n; i++)
         if (scanf (" %lf", &(array[i])) != 1)
           invalid_input_error ();
     }

The formatted input functions are not used as frequently as the formatted output functions. Partly, this is because it takes some care to use them properly. Another reason is that it is difficult to recover from a matching error.

If you are trying to read input that doesn’t match a single, fixed pattern, you may be better off using a tool such as Flex to generate a lexical scanner, or Bison to generate a parser, rather than using scanf. For more information about these tools, see section “” in Flex: The Lexical Scanner Generator, and section “” in The Bison Reference Manual.

7.12.2 Input Conversion Syntax

A scanf template string is a string that contains ordinary multibyte characters interspersed with conversion specifications that start with ‘%’.

Any whitespace character (as defined by the isspace function; see Section 4.1 [Classification of Characters], page 61) in the template causes any number of whitespace characters in the input stream to be read and discarded. The whitespace characters that are matched need not be exactly the same whitespace characters that appear in the template string. For example, write ‘ , ’ in the template to recognize a comma with optional whitespace before and after.

Other characters in the template string that are not part of conversion specifications must match characters in the input stream exactly; if this is not the case, a matching failure occurs.

The conversion specifications in a scanf template string have the general form:

% flags width type conversion

In more detail, an input conversion specification consists of an initial ‘%’ character followed in sequence by:

• An optional flag character ‘*’, which says to ignore the text read for this specification. When scanf finds a conversion specification that uses this flag, it reads input as directed by the rest of the conversion specification, but it discards this input, does not use a pointer argument, and does not increment the count of successful assignments.

• An optional flag character ‘a’ (valid with string conversions only) which requests allocation of a buffer long enough to store the string in. (This is a GNU extension.) See Section 7.12.6 [Dynamically Allocating String Conversions], page 133.

• An optional decimal integer that specifies the maximum field width. Reading of characters from the input stream stops either when this maximum is reached or when a non-matching character is found, whichever happens first. Most conversions discard initial whitespace characters (those that don’t are explicitly documented), and these discarded characters don’t count towards the maximum field width. String input conversions store a null character to mark the end of the input; the maximum field width does not include this terminator.

• An optional type modifier character. For example, you can specify a type modifier of ‘l’ with integer conversions such as ‘%d’ to specify that the argument is a pointer to a long int rather than a pointer to an

int.

• A character that specifies the conversion to be applied.

The exact options that are permitted and how they are interpreted vary between the different conversion specifiers. See the descriptions of the individual conversions for information about the particular options that they allow.

With the ‘-Wformat’ option, the GNU C compiler checks calls to scanf and related functions. It examines the format string and verifies that the correct number and types of arguments are supplied. There is also a GNU C syntax to tell the compiler that a function you write uses a scanf-style format string. See section “Declaring Attributes of Functions” in Using GNU CC, for more information.

7.12.3 Table of Input Conversions

Here is a table that summarizes the various conversion specifications: ‘%d’ Matches an optionally signed integer written in decimal. See

Section 7.12.4 [Numeric Input Conversions], page 130.

‘%i’ Matches an optionally signed integer in any of the formats that the C language defines for specifying an integer constant. See Section 7.12.4 [Numeric Input Conversions], page 130.

‘%o’ Matches an unsigned integer written in octal radix. See Section 7.12.4 [Numeric Input Conversions], page 130.

‘%u’ Matches an unsigned integer written in decimal radix. See Section 7.12.4 [Numeric Input Conversions], page 130.

‘%x’, ‘%X’ Matches an unsigned integer written in hexadecimal radix. See Section 7.12.4 [Numeric Input Conversions], page 130.

‘%e’, ‘%f’, ‘%g’, ‘%E’, ‘%G’ Matches an optionally signed floating-point number. See Section 7.12.4 [Numeric Input Conversions], page 130.

‘%s’ Matches a string containing only non-whitespace characters. See Section 7.12.5 [String Input Conversions], page 131.

‘%[’ Matches a string of characters that belong to a specified set. See Section 7.12.5 [String Input Conversions], page 131.

‘%c’ Matches a string of one or more characters; the number of characters read is controlled by the maximum field width given for the conversion. See Section 7.12.5 [String Input Conversions], page 131.

‘%p’ Matches a pointer value in the same implementation-defined format used by the ‘%p’ output conversion for printf. See Section 7.12.7 [Other Input Conversions], page 133.

‘%n’ This conversion doesn’t read any characters; it records the number of characters read so far by this call. See Section 7.12.7 [Other Input Conversions], page 133.

‘%%’ This matches a literal ‘%’ character in the input stream. No corresponding argument is used. See Section 7.12.7 [Other Input Conversions], page 133.

If the syntax of a conversion specification is invalid, the behavior is undefined. If there aren’t enough function arguments provided to supply addresses for all the conversion specifications in the template strings that perform assignments, or if the arguments are not of the correct types, the behavior is also undefined. On the other hand, extra arguments are simply ignored.

7.12.4 Numeric Input Conversions

This section describes the scanf conversions for reading numeric values. The ‘%d’ conversion matches an optionally signed integer in decimal radix.

The syntax that is recognized is the same as that for the strtol function (see Section 14.7.1 [Parsing of Integers], page 310) with the value 10 for the base argument.

The ‘%i’ conversion matches an optionally signed integer in any of the formats that the C language defines for specifying an integer constant. The syntax that is recognized is the same as that for the strtol function (see Section 14.7.1 [Parsing of Integers], page 310) with the value 0 for the base argument. (You can print integers in this syntax with printf by using the ‘#’ flag character with the ‘%x’, ‘%o’, or ‘%d’ conversion. See Section 7.10.4 [Integer Conversions], page 108.)

For example, any of the strings ‘10’, ‘0xa’, or ‘012’ could be read in as integers under the ‘%i’ conversion. Each of these specifies a number with decimal value 10.

The ‘%o’, ‘%u’, and ‘%x’ conversions match unsigned integers in octal, decimal, and hexadecimal radices, respectively. The syntax that is recognized is the same as that for the strtoul function (see Section 14.7.1 [Parsing of Integers], page 310) with the appropriate value (8, 10, or 16) for the base argument.

The ‘%X’ conversion is identical to the ‘%x’ conversion. They both permit either uppercase or lowercase letters to be used as digits.

The default type of the corresponding argument for the %d and %i conversions is int *, and unsigned int * for the other integer conversions. You can use the following type modifiers to specify other sizes of integer:

‘h’ Specifies that the argument is a short int * or unsigned short

          int *.

‘l’ Specifies that the argument is a long int * or unsigned long int *. Two ‘l’ characters is like the ‘L’ modifier, below.

‘ll’

‘L’ ‘q’ Specifies that the argument is a long long int * or unsigned

long long int *. (The long long type is an extension supported by the GNU C compiler. For systems that don’t provide extra-long integers, this is the same as long int.) The ‘q’ modifier is another name for the same thing, which comes from 4.4 BSD; a long long int is sometimes called a “quad”

          int.

All of the ‘%e’, ‘%f’, ‘%g’, ‘%E’, and ‘%G’ input conversions are interchangeable. They all match an optionally signed floating point number, in the same syntax as for the strtod function (see Section 14.7.2 [Parsing of Floats], page 314).

For the floating-point input conversions, the default argument type is float *. (This is different from the corresponding output conversions, where the default type is double; remember that float arguments to printf are converted to double by the default argument promotions, but float * arguments are not promoted to double *.) You can specify other sizes of float using these type modifiers:

‘l’ Specifies that the argument is of type double *.

‘L’ Specifies that the argument is of type long double *.

For all the above number parsing formats there is an additional optional flag ‘'’. When this flag is given the scanf function expects the number represented in the input string to be formatted according to the grouping rules of the currently selected locale (see Section 19.6.1 [Generic Numeric Formatting Parameters], page 384).

If the "C" or "POSIX" locale is selected there is no difference. But for a locale which specifies values for the appropriate fields in the locale the input must have the correct form in the input. Otherwise the longest prefix with a correct form is processed.

7.12.5 String Input Conversions

This section describes the scanf input conversions for reading string and character values: ‘%s’, ‘%[’, and ‘%c’.

You have two options for how to receive the input from these conversions:

• Provide a buffer to store it in. This is the default. You should provide an argument of type char *. Warning: To make a robust program, you must make sure that the input (plus its terminating null) cannot possibly exceed the size of the buffer you provide. In general, the only way to do this is to specify a maximum field width one less than the buffer size. If you provide the buffer, always specify a maximum field width to prevent overflow.

• Ask scanf to allocate a big enough buffer, by specifying the ‘a’ flag character. This is a GNU extension. You should provide an argument of type char ** for the buffer address to be stored in. See Section 7.12.6 [Dynamically Allocating String Conversions], page 133. The ‘%c’ conversion is the simplest: it matches a fixed number of char-

acters, always. The maximum field with says how many characters to read; if you don’t specify the maximum, the default is 1. This conversion doesn’t append a null character to the end of the text it reads. It also does not skip over initial whitespace characters. It reads precisely the next n characters, and fails if it cannot get that many. Since there is always a maximum field width with ‘%c’ (whether specified, or 1 by default), you can always prevent overflow by making the buffer long enough.

The ‘%s’ conversion matches a string of non-whitespace characters. It skips and discards initial whitespace, but stops when it encounters more whitespace after having read something. It stores a null character at the end of the text that it reads.

For example, reading the input:

      hello, world

with the conversion ‘%10c’ produces " hello, wo", but reading the same input with the conversion ‘%10s’ produces "hello,".

Warning: If you do not specify a field width for ‘%s’, then the number of characters read is limited only by where the next whitespace character appears. This almost certainly means that invalid input can make your program crash—which is a bug.

To read in characters that belong to an arbitrary set of your choice, use the ‘%[’ conversion. You specify the set between the ‘[’ character and a following ‘]’ character, using the same syntax used in regular expressions. As special cases:

• A literal ‘]’ character can be specified as the first character of the set.

• An embedded ‘-’ character (that is, one that is not the first or last character of the set) is used to specify a range of characters.

• If a caret character ‘^’ immediately follows the initial ‘[’, then the set of allowed input characters is the everything except the characters listed.

The ‘%[’ conversion does not skip over initial whitespace characters. Here are some examples of ‘%[’ conversions and what they mean:

‘%25[1234567890]’

Matches a string of up to 25 digits.

‘%25[][]’ Matches a string of up to 25 square brackets.

‘%25[^\f\n\r\t\v]’

Matches a string up to 25 characters long that doesn’t contain any of the standard whitespace characters. This is slightly different from ‘%s’, because if the input begins with a whitespace character, ‘%[’ reports a matching failure while ‘%s’ simply discards the initial whitespace.

‘%25[a-z]’

Matches up to 25 lowercase characters.

One more reminder: the ‘%s’ and ‘%[’ conversions are dangerous if you don’t specify a maximum width or use the ‘a’ flag, because input too long would overflow whatever buffer you have provided for it. No matter how long your buffer is, a user could supply input that is longer. A well-written program reports invalid input with a comprehensible error message, not with a crash.

7.12.6 Dynamically Allocating String Conversions

A GNU extension to formatted input lets you safely read a string with no maximum size. Using this feature, you don’t supply a buffer; instead, scanf allocates a buffer big enough to hold the data and gives you its address. To use this feature, write ‘a’ as a flag character, as in ‘%as’ or ‘%a[0-9a-z]’.

The pointer argument you supply for where to store the input should have type char **. The scanf function allocates a buffer and stores its address in the word that the argument points to. You should free the buffer with free when you no longer need it.

Here is an example of using the ‘a’ flag with the ‘%[…]’ conversion specification to read a “variable assignment” of the form ‘variable = value’.

     {
       char *variable, *value;
       if (2 > scanf ("%a[a-zA-Z0-9] = %a[^\n]\n",
                      &variable, &value))
         {
           invalid_input_error ();
           return 0;
         }

…

     }

7.12.7 Other Input Conversions

This section describes the miscellaneous input conversions.

The ‘%p’ conversion is used to read a pointer value. It recognizes the same syntax as is used by the ‘%p’ output conversion for printf (see Section 7.10.6 [Other Output Conversions], page 112); that is, a hexadecimal number just as the ‘%x’ conversion accepts. The corresponding argument should be of type void **; that is, the address of a place to store a pointer.

The resulting pointer value is not guaranteed to be valid if it was not originally written during the same program execution that reads it in.

The ‘%n’ conversion produces the number of characters read so far by this call. The corresponding argument should be of type int *. This conversion works in the same way as the ‘%n’ conversion for printf; see Section 7.10.6 [Other Output Conversions], page 112, for an example.

The ‘%n’ conversion is the only mechanism for determining the success of literal matches or conversions with suppressed assignments. If the ‘%n’ follows the locus of a matching failure, then no value is stored for it since scanf returns before processing the ‘%n’. If you store -1 in that argument slot before calling scanf, the presence of -1 after scanf indicates an error occurred before the ‘%n’ was reached.

Finally, the ‘%%’ conversion matches a literal ‘%’ character in the input stream, without using an argument. This conversion does not permit any flags, field width, or type modifier to be specified.

7.12.8 Formatted Input Functions

Here are the descriptions of the functions for performing formatted input. Prototypes for these functions are in the header file ‘stdio.h’.

Function int scanf (const char *template, …)

The scanf function reads formatted input from the stream stdin under the control of the template string template. The optional arguments are pointers to the places which receive the resulting values. The return value is normally the number of successful assignments. If an end-of-file condition is detected before any matches are performed (including matches against whitespace and literal characters in the template), then EOF is returned.

Function int fscanf (FILE *stream, const char *template, …)

This function is just like scanf, except that the input is read from the stream stream instead of stdin.

Function int sscanf (const char *s, const char *template, …)

This is like scanf, except that the characters are taken from the nullterminated string s instead of from a stream. Reaching the end of the string is treated as an end-of-file condition. The behavior of this function is undefined if copying takes place between objects that overlap—for example, if s is also given as an argument to receive a string read under control of the ‘%s’ conversion.

7.12.9 Variable Arguments Input Functions

The functions vscanf and friends are provided so that you can define your own variadic scanf-like functions that make use of the same internals as the built-in formatted output functions. These functions are analogous to the vprintf series of output functions. See Section 7.10.9 [Variable Arguments Output Functions], page 115, for important information on how to use them.

Portability Note: The functions listed in this section are GNU extensions.

Function int vscanf (const char *template, va_list ap)

This function is similar to scanf except that, instead of taking a variable number of arguments directly, it takes an argument list pointer ap of type va_list (see Section A.2 [Variadic Functions], page 555).

Function int vfscanf (FILE *stream, const char *template,

      va_list ap)

This is the equivalent of fscanf with the variable argument list specified directly as for vscanf.

Function int vsscanf (const char *s, const char *template,
      va_list ap)

This is the equivalent of sscanf with the variable argument list specified directly as for vscanf.

In GNU C, there is a special construct you can use to let the compiler know that a function uses a scanf-style format string. Then it can check the number and types of arguments in each call to the function, and warn you when they do not match the format string. See section “Declaring Attributes of Functions” in Using GNU CC, for details.

7.13 End-Of-File and Errors

Many of the functions described in this chapter return the value of the macro EOF to indicate unsuccessful completion of the operation. Since EOF is used to report both end of file and random errors, it’s often better to use the feof function to check explicitly for end of file and ferror to check for errors. These functions check indicators that are part of the internal state of the stream object, indicators set if the appropriate condition was detected by a previous I/O operation on that stream.

These symbols are declared in the header file ‘stdio.h’.

Macro int EOF

This macro is an integer value that is returned by a number of functions to indicate an end-of-file condition, or some other error situation. With the GNU library, EOF is -1. In other libraries, its value may be some other negative number.

Function void clearerr (FILE *stream)

This function clears the end-of-file and error indicators for the stream stream. The file positioning functions (see Section 7.15 [File Positioning], page 137) also clear the end-of-file indicator for the stream.

Function int feof (FILE *stream)

The feof function returns nonzero if and only if the end-of-file indicator for the stream stream is set.

Function int ferror (FILE *stream)

The ferror function returns nonzero if and only if the error indicator for the stream stream is set, indicating that an error has occurred on a previous operation on the stream.

In addition to setting the error indicator associated with the stream, the functions that operate on streams also set errno in the same way as the corresponding low-level functions that operate on file descriptors. For example, all of the functions that perform output to a stream—such as fputc, printf, and fflush—are implemented in terms of write, and all of the errno error conditions defined for write are meaningful for these functions. For more information about the descriptor-level I/O functions, see Chapter 8 [Low-Level Input/Output], page 151.

7.14 Text and Binary Streams

The GNU system and other POSIX-compatible operating systems organize all files as uniform sequences of characters. However, some other systems make a distinction between files containing text and files containing binary data, and the input and output facilities of ISO C provide for this distinction. This section tells you how to write programs portable to such systems.

When you open a stream, you can specify either a text stream or a binary stream. You indicate that you want a binary stream by specifying the ‘b’ modifier in the opentype argument to fopen; see Section 7.3 [Opening Streams], page 94. Without this option, fopen opens the file as a text stream.

Text and binary streams differ in several ways:

• The data read from a text stream is divided into lines which are terminated by newline ('\n') characters, while a binary stream is simply a long series of characters. A text stream might on some systems fail to handle lines more than 254 characters long (including the terminating newline character).

• On some systems, text files can contain only printing characters, horizontal tab characters, and newlines, and so text streams may not support other characters. However, binary streams can handle any character value.

• Space characters that are written immediately preceding a newline character in a text stream may disappear when the file is read in again.

• More generally, there need not be a one-to-one mapping between characters that are read from or written to a text stream, and the characters in the actual file.

Since a binary stream is always more capable and more predictable than a text stream, you might wonder what purpose text streams serve. Why not simply always use binary streams? The answer is that on these operating systems, text and binary streams use different file formats, and the only way to read or write “an ordinary file of text” that can work with other text-oriented programs is through a text stream.

In the GNU library, and on all POSIX systems, there is no difference between text streams and binary streams. When you open a stream, you get the same kind of stream regardless of whether you ask for binary. This stream can handle any file content, and has none of the restrictions that text streams sometimes have.

7.15 File Positioning

The file position of a stream describes where in the file the stream is currently reading or writing. I/O on the stream advances the file position through the file. In the GNU system, the file position is represented as an integer, which counts the number of bytes from the beginning of the file. See Section 6.1.2 [File Position], page 87.

During I/O to an ordinary disk file, you can change the file position whenever you wish, so as to read or write any portion of the file. Some other kinds of files may also permit this. Files which support changing the file position are sometimes referred to as random-access files.

You can use the functions in this section to examine or modify the file position indicator associated with a stream. The symbols listed below are declared in the header file ‘stdio.h’.

Function long int ftell (FILE *stream)

This function returns the current file position of the stream stream.

This function can fail if the stream doesn’t support file positioning, or if the file position can’t be represented in a long int, and possibly for other reasons as well. If a failure occurs, a value of -1 is returned.

Function int fseek (FILE *stream, long int offset, int whence)

The fseek function is used to change the file position of the stream stream. The value of whence must be one of the constants SEEK_SET, SEEK_CUR, or SEEK_END, to indicate whether the offset is relative to the beginning of the file, the current file position, or the end of the file, respectively. This function returns a value of zero if the operation was successful, and a nonzero value to indicate failure. A successful call also clears the end-of-file indicator of stream and discards any characters that were “pushed back” by the use of ungetc. fseek either flushes any buffered output before setting the file position or else remembers it so it will be written later in its proper place in the file.

Portability Note: In non-POSIX systems, ftell and fseek might work reliably only on binary streams. See Section 7.14 [Text and Binary Streams], page 136.

The following symbolic constants are defined for use as the whence argument to fseek. They are also used with the lseek function (see Section 8.2 [Input and Output Primitives], page 153) and to specify offsets for file locks (see Section 8.7 [Control Operations on Files], page 165).

Macro int SEEK SET

This is an integer constant which, when used as the whence argument to the fseek function, specifies that the offset provided is relative to the beginning of the file.

Macro int SEEK CUR

This is an integer constant which, when used as the whence argument to the fseek function, specifies that the offset provided is relative to the current file position.

Macro int SEEK END

This is an integer constant which, when used as the whence argument to the fseek function, specifies that the offset provided is relative to the end of the file.

Function void rewind (FILE *stream)

The rewind function positions the stream stream at the begining of the file. It is equivalent to calling fseek on the stream with an offset argument of 0L and a whence argument of SEEK_SET, except that the return value is discarded and the error indicator for the stream is reset.

These three aliases for the ‘SEEK_…’ constants exist for the sake of compatibility with older BSD systems. They are defined in two different header

files: ‘fcntl.h’ and ‘sys/file.h’.
L_SET  An alias for SEEK_SET.
L_INCR  An alias for SEEK_CUR.
L_XTND  An alias for SEEK_END.

7.16 Portable File-Position Functions

On the GNU system, the file position is truly a character count. You can specify any character count value as an argument to fseek and get reliable results for any random access file. However, some ISO C systems do not represent file positions in this way.

On some systems where text streams truly differ from binary streams, it is impossible to represent the file position of a text stream as a count of characters from the beginning of the file. For example, the file position on some systems must encode both a record offset within the file, and a character offset within the record.

As a consequence, if you want your programs to be portable to these systems, you must observe certain rules:

• The value returned from ftell on a text stream has no predictable relationship to the number of characters you have read so far. The only thing you can rely on is that you can use it subsequently as the offset argument to fseek to move back to the same file position.

• In a call to fseek on a text stream, either the offset must either be zero; or whence must be SEEK_SET and the offset must be the result of an earlier call to ftell on the same stream.

• The value of the file position indicator of a text stream is undefined while there are characters that have been pushed back with ungetc that haven’t been read or discarded. See Section 7.8 [Unreading], page 102. But even if you observe these rules, you may still have trouble for long

files, because ftell and fseek use a long int value to represent the file position. This type may not have room to encode all the file positions in a large file.

So if you do want to support systems with peculiar encodings for the file positions, it is better to use the functions fgetpos and fsetpos instead. These functions represent the file position using the data type fpos_t, whose internal representation varies from system to system.

These symbols are declared in the header file ‘stdio.h’.

Data Type fpos t

This is the type of an object that can encode information about the file position of a stream, for use by the functions fgetpos and fsetpos. In the GNU system, fpos_t is equivalent to off_t or long int. In other systems, it might have a different internal representation.

Function int fgetpos (FILE *stream, fpos_t *position)

This function stores the value of the file position indicator for the stream stream in the fpos_t object pointed to by position. If successful, fgetpos returns zero; otherwise it returns a nonzero value and stores an implementation-defined positive value in errno.

Function int fsetpos (FILE *stream, const fpos_t position)

This function sets the file position indicator for the stream stream to the position position, which must have been set by a previous call to fgetpos on the same stream. If successful, fsetpos clears the end-of-file indicator on the stream, discards any characters that were “pushed back” by the use of ungetc, and returns a value of zero. Otherwise, fsetpos returns a nonzero value and stores an implementation-defined positive value in errno.

7.17 Stream Buffering

Characters that are written to a stream are normally accumulated and transmitted asynchronously to the file in a block, instead of appearing as soon as they are output by the application program. Similarly, streams often retrieve input from the host environment in blocks rather than on a character-by-character basis. This is called buffering.

If you are writing programs that do interactive input and output using streams, you need to understand how buffering works when you design the user interface to your program. Otherwise, you might find that output (such as progress or prompt messages) doesn’t appear when you intended it to, or other unexpected behavior.

This section deals only with controlling when characters are transmitted between the stream and the file or device, and not with how things like echoing, flow control, and the like are handled on specific classes of devices. For information on common control operations on terminal devices, see Chapter 12 [Low-Level Terminal Interface], page 271.

You can bypass the stream buffering facilities altogether by using the low-level input and output functions that operate on file descriptors instead. See Chapter 8 [Low-Level Input/Output], page 151.

7.17.1 Buffering Concepts

There are three different kinds of buffering strategies:

• Characters written to or read from an unbuffered stream are transmitted individually to or from the file as soon as possible.

• Characters written to a line buffered stream are transmitted to the file in blocks when a newline character is encountered.

• Characters written to or read from a fully buffered stream are transmitted to or from the file in blocks of arbitrary size.

Newly opened streams are normally fully buffered, with one exception: a stream connected to an interactive device such as a terminal is initially line buffered. See Section 7.17.3 [Controlling Which Kind of Buffering], page 141, for information on how to select a different kind of buffering. Usually the automatic selection gives you the most convenient kind of buffering for the file or device you open.

The use of line buffering for interactive devices implies that output messages ending in a newline will appear immediately—which is usually what you want. Output that doesn’t end in a newline might or might not show up immediately, so if you want them to appear immediately, you should flush buffered output explicitly with fflush, as described in Section 7.17.2 [Flushing Buffers], page 141.

7.17.2 Flushing Buffers

Flushing output on a buffered stream means transmitting all accumulated characters to the file. There are many circumstances when buffered output on a stream is flushed automatically:

• When you try to do output and the output buffer is full.

• When the stream is closed. See Section 7.4 [Closing Streams], page 96.

• When the program terminates by calling exit. See Section 22.3.1 [Normal Termination], page 462.

• When a newline is written, if the stream is line buffered.

• Whenever an input operation on any stream actually reads data from its file.

If you want to flush the buffered output at another time, call fflush, which is declared in the header file ‘stdio.h’.

Function int fflush (FILE *stream)

This function causes any buffered output on stream to be delivered to the file. If stream is a null pointer, then fflush causes buffered output on all open output streams to be flushed. This function returns EOF if a write error occurs, or zero otherwise.

Compatibility Note: Some brain-damaged operating systems have been known to be so thoroughly fixated on line-oriented input and output that flushing a line buffered stream causes a newline to be written! Fortunately, this “feature” seems to be becoming less common. You do not need to worry about this in the GNU system.

7.17.3 Controlling Which Kind of Buffering

After opening a stream (but before any other operations have been performed on it), you can explicitly specify what kind of buffering you want it to have using the setvbuf function.

The facilities listed in this section are declared in the header file ‘stdio.h’.

Function int setvbuf (FILE *stream, char *buf, int mode,

      size_t size)

This function is used to specify that the stream stream should have the buffering mode mode, which can be either _IOFBF (for full buffering), _IOLBF (for line buffering), or _IONBF (for unbuffered input/output). If you specify a null pointer as the buf argument, then setvbuf allocates a buffer itself using malloc. This buffer will be freed when you close the stream. Otherwise, buf should be a character array that can hold at least size characters. You should not free the space for this array as long as the stream remains open and this array remains its buffer. You should usually either allocate it statically, or malloc (see Section 3.3 [Unconstrained Allocation], page 32) the buffer. Using an automatic array is not a good idea unless you close the file before exiting the block that declares the array. While the array remains a stream buffer, the stream I/O functions will use the buffer for their internal purposes. You shouldn’t try to access the values in the array directly while the stream is using it for buffering. The setvbuf function returns zero on success, or a nonzero value if the value of mode is not valid or if the request could not be honored.

Macro int IOFBF

The value of this macro is an integer constant expression that can be used as the mode argument to the setvbuf function to specify that the stream should be fully buffered.

Macro int IOLBF

The value of this macro is an integer constant expression that can be used as the mode argument to the setvbuf function to specify that the stream should be line buffered.

Macro int IONBF

The value of this macro is an integer constant expression that can be used as the mode argument to the setvbuf function to specify that the stream should be unbuffered.

Macro int BUFSIZ

The value of this macro is an integer constant expression that is good to use for the size argument to setvbuf. This value is guaranteed to be at least 256. The value of BUFSIZ is chosen on each system so as to make stream I/O efficient. So it is a good idea to use BUFSIZ as the size for the buffer when you call setvbuf.

Actually, you can get an even better value to use for the buffer size by means of the fstat system call: it is found in the st_blksize field of the file attributes. See Section 9.8.1 [What the File Attribute Values Mean], page 195. Sometimes people also use BUFSIZ as the allocation size of buffers used for related purposes, such as strings used to receive a line of input with fgets (see Section 7.6 [Character Input], page 98). There is no particular reason to use BUFSIZ for this instead of any other integer, except that it might lead to doing I/O in chunks of an efficient size.

Function void setbuf (FILE *stream, char *buf )

If buf is a null pointer, the effect of this function is equivalent to calling setvbuf with a mode argument of _IONBF. Otherwise, it is equivalent to calling setvbuf with buf, and a mode of _IOFBF and a size argument of BUFSIZ. The setbuf function is provided for compatibility with old code; use setvbuf in all new programs.

Function void setbuffer (FILE *stream, char *buf, size_t size)

If buf is a null pointer, this function makes stream unbuffered. Otherwise, it makes stream fully buffered using buf as the buffer. The size argument specifies the length of buf. This function is provided for compatibility with old BSD code. Use setvbuf instead.

Function void setlinebuf (FILE *stream)

This function makes stream be line buffered, and allocates the buffer for you. This function is provided for compatibility with old BSD code. Use setvbuf instead.

7.18 Other Kinds of Streams

The GNU library provides ways for you to define additional kinds of streams that do not necessarily correspond to an open file.

One such type of stream takes input from or writes output to a string. These kinds of streams are used internally to implement the sprintf and sscanf functions. You can also create such a stream explicitly, using the functions described in Section 7.18.1 [String Streams], page 144.

More generally, you can define streams that do input/output to arbitrary objects using functions supplied by your program. This protocol is discussed in Section 7.18.3 [Programming Your Own Custom Streams], page 147.

Portability Note: The facilities described in this section are specific to GNU. Other systems or C implementations might or might not provide equivalent functionality.

7.18.1 String Streams

The fmemopen and open_memstream functions allow you to do I/O to a string or memory buffer. These facilities are declared in ‘stdio.h’.

Function FILE * fmemopen (void *buf, size_t size, const

char *opentype) This function opens a stream that allows the access specified by the opentype argument, that reads from or writes to the buffer specified by the argument buf. This array must be at least size bytes long. If you specify a null pointer as the buf argument, fmemopen dynamically allocates (as with malloc; see Section 3.3 [Unconstrained Allocation], page 32) an array size bytes long. This is really only useful if you are going to write things to the buffer and then read them back in again, because you have no way of actually getting a pointer to the buffer (for this, try open_memstream, below). The buffer is freed when the stream is open. The argument opentype is the same as in fopen (See Section 7.3 [Opening Streams], page 94). If the opentype specifies append mode, then the initial file position is set to the first null character in the buffer. Otherwise the initial file position is at the beginning of the buffer. When a stream open for writing is flushed or closed, a null character (zero byte) is written at the end of the buffer if it fits. You should add an extra byte to the size argument to account for this. Attempts to write more than size bytes to the buffer result in an error. For a stream open for reading, null characters (zero bytes) in the buffer do not count as “end of file”. Read operations indicate end of file only when the file position advances past size bytes. So, if you want to read characters from a null-terminated string, you should supply the length of the string as the size argument.

Here is an example of using fmemopen to create a stream for reading from a string:

  #include <stdio.h>
  static char buffer[] = "foobar";
  int
  main (void)
  {
    int ch;
    FILE *stream;
    stream = fmemopen (buffer, strlen (buffer), "r");
    while ((ch = fgetc (stream)) != EOF)

      printf ("Got %c\n", ch);
    fclose (stream);
    return 0;
  }

This program produces the following output:

  Got f
  Got o
  Got o
  Got b
  Got a
  Got r

Function FILE * open memstream (char **ptr, size_t

*sizeloc) This function opens a stream for writing to a buffer. The buffer is allocated dynamically (as with malloc; see Section 3.3 [Unconstrained Allocation], page 32) and grown as necessary. When the stream is closed with fclose or flushed with fflush, the locations ptr and sizeloc are updated to contain the pointer to the buffer and its size. The values thus stored remain valid only as long as no further output on the stream takes place. If you do more output, you must flush the stream again to store new values before you use them again. A null character is written at the end of the buffer. This null character is not included in the size value stored at sizeloc. You can move the stream’s file position with fseek (see Section 7.15 [File Positioning], page 137). Moving the file position past the end of the data already written fills the intervening space with zeroes.

Here is an example of using open_memstream:

  #include <stdio.h>
  int
  main (void)
  {
    char *bp;
    size_t size;
    FILE *stream;
    stream = open_memstream (&bp, &size);
    fprintf (stream, "hello");
    fflush (stream);
    printf ("buf = `%s', size = %d\n", bp, size);

    fprintf (stream, ", world");
    fclose (stream);
    printf ("buf = `%s', size = %d\n", bp, size);
    return 0;
  }

This program produces the following output:

  buf = `hello', size = 5
  buf = `hello, world', size = 12

7.18.2 Obstack Streams

You can open an output stream that puts it data in an obstack. See Section 3.4 [Obstacks], page 43.

Function FILE * open obstack stream (struct obstack

*obstack) This function opens a stream for writing data into the obstack obstack. This starts an object in the obstack and makes it grow as data is written (see Section 3.4.6 [Growing Objects], page 47). Calling fflush on this stream updates the current size of the object to match the amount of data that has been written. After a call to fflush, you can examine the object temporarily. You can move the file position of an obstack stream with fseek (see Section 7.15 [File Positioning], page 137). Moving the file position past the end of the data written fills the intervening space with zeros. To make the object permanent, update the obstack with fflush, and then use obstack_finish to finalize the object and get its address. The following write to the stream starts a new object in the obstack, and later writes add to that object until you do another fflush and

obstack_finish.

But how do you find out how long the object is? You can get the length in bytes by calling obstack_object_size (see Section 3.4.8 [Status of an Obstack], page 51), or you can null-terminate the object like this:

     obstack_1grow (obstack, 0);

Whichever one you do, you must do it before calling obstack_finish. (You can do both if you wish.)

Here is a sample function that uses open_obstack_stream:

  char *
  make_message_string (const char *a, int b)
  {
    FILE *stream = open_obstack_stream (&message_obstack);
    output_task (stream);

       fprintf (stream, ": ");
       fprintf (stream, a, b);
       fprintf (stream, "\n");
       fclose (stream);
       obstack_1grow (&message_obstack, 0);
       return obstack_finish (&message_obstack);
     }

7.18.3 Programming Your Own Custom Streams

This section describes how you can make a stream that gets input from an arbitrary data source or writes output to an arbitrary data sink programmed by you. We call these custom streams.

7.18.3.1 Custom Streams and Cookies

Inside every custom stream is a special object called the cookie. This is an object supplied by you which records where to fetch or store the data read or written. It is up to you to define a data type to use for the cookie. The stream functions in the library never refer directly to its contents, and they don’t even know what the type is; they record its address with type

void *.

To implement a custom stream, you must specify how to fetch or store the data in the specified place. You do this by defining hook functions to read, write, change “file position”, and close the stream. All four of these functions will be passed the stream’s cookie so they can tell where to fetch or store the data. The library functions don’t know what’s inside the cookie, but your functions will know.

When you create a custom stream, you must specify the cookie pointer, and also the four hook functions stored in a structure of type cookie_io_

functions_t.

These facilities are declared in ‘stdio.h’.

Data Type cookie io functions t

This is a structure type that holds the functions that define the communications protocol between the stream and its cookie. It has the following members:

  cookie_read_function_t *read

This is the function that reads data from the cookie. If the value is a null pointer instead of a function, then read operations on ths stream always return EOF.

cookie_write_function_t *write

This is the function that writes data to the cookie. If the value is a null pointer instead of a function, then data written to the stream is discarded.

cookie_seek_function_t *seek

This is the function that performs the equivalent of file positioning on the cookie. If the value is a null pointer instead of a function, calls to fseek on this stream can only seek to locations within the buffer; any attempt to seek outside the buffer will return an ESPIPE error.

cookie_close_function_t *close

This function performs any appropriate cleanup on the cookie when closing the stream. If the value is a null pointer instead of a function, nothing special is done to close the cookie when the stream is closed.

Function FILE * fopencookie (void *cookie, const char

*opentype, cookie_io_functions_t io-functions)

This function actually creates the stream for communicating with the cookie using the functions in the io-functions argument. The opentype argument is interpreted as for fopen; see Section 7.3 [Opening Streams], page 94. (But note that the “truncate on open” option is ignored.) The new stream is fully buffered. The fopencookie function returns the newly created stream, or a null pointer in case of an error.

7.18.3.2 Custom Stream Hook Functions

Here are more details on how you should define the four hook functions that a custom stream needs.

You should define the function to read data from the cookie as:

ssize_t reader (void *cookie, void *buffer, size_t size)

This is very similar to the read function; see Section 8.2 [Input and Output Primitives], page 153. Your function should transfer up to size bytes into the buffer, and return the number of bytes read, or zero to indicate endof-file. You can return a value of -1 to indicate an error.

You should define the function to write data to the cookie as:

ssize_t writer (void *cookie, const void *buffer, size_t size)

This is very similar to the write function; see Section 8.2 [Input and Output Primitives], page 153. Your function should transfer up to size bytes from the buffer, and return the number of bytes written. You can return a value of -1 to indicate an error.

You should define the function to perform seek operations on the cookie as:

int seeker (void *cookie, fpos_t *position, int whence)

For this function, the position and whence arguments are interpreted as for fgetpos; see Section 7.16 [Portable File-Position Functions], page 138. In the GNU library, fpos_t is equivalent to off_t or long int, and simply represents the number of bytes from the beginning of the file.

After doing the seek operation, your function should store the resulting file position relative to the beginning of the file in position. Your function should return a value of 0 on success and -1 to indicate an error.

You should define the function to do cleanup operations on the cookie appropriate for closing the stream as:

int cleaner (void *cookie)

Your function should return -1 to indicate an error, and 0 otherwise.

Data Type cookie read function

This is the data type that the read function for a custom stream should have. If you declare the function as shown above, this is the type it will have.

Data Type cookie write function

The data type of the write function for a custom stream.

Data Type cookie seek function

The data type of the seek function for a custom stream.

Data Type cookie close function

The data type of the close function for a custom stream.

150 The GNU C Library

8 Low-Level Input/Output

This chapter describes functions for performing low-level input/output operations on file descriptors. These functions include the primitives for the higher-level I/O functions described in Chapter 7 [Input/Output on Streams], page 93, as well as functions for performing low-level control operations for which there are no equivalents on streams.

Stream-level I/O is more flexible and usually more convenient; therefore, programmers generally use the descriptor-level functions only when necessary. These are some of the usual reasons:

• For reading binary files in large chunks.

• For reading an entire file into core before parsing it.

• To perform operations other than data transfer, which can only be done with a descriptor. (You can use fileno to get the descriptor corresponding to a stream.)

• To pass descriptors to a child process. (The child can create its own stream to use a descriptor that it inherits, but cannot inherit a stream directly.)

8.1 Opening and Closing Files

This section describes the primitives for opening and closing files using file descriptors. The open and creat functions are declared in the header file ‘fcntl.h’, while close is declared in ‘unistd.h’.

Function int open (const char *filename, int flags[, mode_t

mode]) The open function creates and returns a new file descriptor for the file named by filename. Initially, the file position indicator for the file is at the beginning of the file. The argument mode is used only when a file is created, but it doesn’t hurt to supply the argument in any case. The flags argument controls how the file is to be opened. This is a bit mask; you create the value by the bitwise OR of the appropriate parameters (using the ‘|’ operator in C). See Section 8.10 [File Status Flags], page 170, for the parameters available. The normal return value from open is a non-negative integer file descriptor. In the case of an error, a value of -1 is returned instead. In addition to the usual file name errors (see Section 6.2.3 [File Name Errors], page 89), the following errno error conditions are defined for this function:

EACCES The file exists but is not readable/writable as requested by the flags argument, the file does not exist and the directory is unwritable so it cannot be created.

EEXIST Both O_CREAT and O_EXCL are set, and the named file already exists.

EINTR The open operation was interrupted by a signal. See Section 21.5 [Primitives Interrupted by Signals], page 425.

EISDIR The flags argument specified write access, and the file is a directory.

EMFILE The process has too many files open. The maximum number of file descriptors is controlled by the RLIMIT_NOFILE resource limit; see Section 17.6 [Limiting Resource Usage], page 362.

ENFILE The entire system, or perhaps the file system which contains the directory, cannot support any additional open files at the moment. (This problem cannot happen on the GNU system.)

ENOENT The named file does not exist, and O_CREAT is not specified.

ENOSPC The directory or file system that would contain the new file cannot be extended, because there is no disk space left.

ENXIO O_NONBLOCK and O_WRONLY are both set in the flags argument, the file named by filename is a FIFO (see Chapter 10 [Pipes and FIFOs], page 215), and no process has the file open for reading.

EROFS The file resides on a read-only file system and any of O_WRONLY, O_RDWR, and O_TRUNC are set in the flags argument, or O_CREAT is set and the file does not already exist.

The open function is the underlying primitive for the fopen and freopen functions, that create streams.

Obsolete function int creat (const char *filename, mode_t

mode) This function is obsolete. The call:

creat (filename, mode)

is equivalent to:

open (filename, O_WRONLY | O_CREAT | O_TRUNC, mode)

Function int close (int filedes)

The function close closes the file descriptor filedes. Closing a file has the following consequences:

• The file descriptor is deallocated.

• Any record locks owned by the process on the file are unlocked.

• When all file descriptors associated with a pipe or FIFO have been closed, any unread data is discarded.

The normal return value from close is 0; a value of -1 is returned in case of failure. The following errno error conditions are defined for this function:

EBADF The filedes argument is not a valid file descriptor.

EINTR The close call was interrupted by a signal. See Section 21.5 [Primitives Interrupted by Signals], page 425. Here is an example of how to handle EINTR properly:

                 TEMP_FAILURE_RETRY (close (desc));
ENOSPC
EIO

EDQUOT When the file is accessed by NFS, these errors from write can sometimes not be detected until close. See Section 8.2 [Input and Output Primitives], page 153, for details on their meaning.

To close a stream, call fclose (see Section 7.4 [Closing Streams], page 96) instead of trying to close its underlying file descriptor with close. This flushes any buffered output and updates the stream object to indicate that it is closed.

8.2 Input and Output Primitives

This section describes the functions for performing primitive input and output operations on file descriptors: read, write, and lseek. These functions are declared in the header file ‘unistd.h’.

Data Type ssize t

This data type is used to represent the sizes of blocks that can be read or written in a single operation. It is similar to size_t, but must be a signed type.

Function ssize_t read (int filedes, void *buffer, size_t size)

The read function reads up to size bytes from the file with descriptor filedes, storing the results in the buffer. (This is not necessarily a character string and there is no terminating null character added.) The return value is the number of bytes actually read. This might be less than size; for example, if there aren’t that many bytes left in the file or if there aren’t that many bytes immediately available. The exact behavior depends on what kind of file it is. Note that reading less than size bytes is not an error.

A value of zero indicates end-of-file (except if the value of the size argument is also zero). This is not considered an error. If you keep calling read while at end-of-file, it will keep returning zero and doing nothing else. If read returns at least one character, there is no way you can tell whether end-of-file was reached. But if you did reach the end, the next read will return zero. In case of an error, read returns -1. The following errno error conditions are defined for this function:

EAGAIN Normally, when no input is immediately available, read waits for some input. But if the O_NONBLOCK flag is set for the file (see Section 8.10 [File Status Flags], page 170), read returns immediately without reading any data, and reports this error. Compatibility Note: Most versions of BSD Unix use a different error code for this: EWOULDBLOCK. In the GNU library, EWOULDBLOCK is an alias for EAGAIN, so it doesn’t matter which name you use. On some systems, reading a large amount of data from a character special file can also fail with EAGAIN if the kernel cannot find enough physical memory to lock down the user’s pages. This is limited to devices that transfer with direct memory access into the user’s memory, which means it does not include terminals, since they always use separate buffers inside the kernel. This problem never happens in the GNU system. Any condition that could result in EAGAIN can instead result in a successful read which returns fewer bytes than requested. Calling read again immediately would result

in EAGAIN.

EBADF The filedes argument is not a valid file descriptor, or is not open for reading.

EINTR read was interrupted by a signal while it was waiting for input. See Section 21.5 [Primitives Interrupted by Signals], page 425. A signal will not necessary cause read to return EINTR; it may instead result in a successful read which returns fewer bytes than requested.

EIO For many devices, and for disk files, this error code indicates a hardware error. EIO also occurs when a background process tries to read from the controlling terminal, and the normal action of stopping the process by sending it a SIGTTIN signal isn’t working. This might happen if signal is being blocked or ignored, or because the process group is orphaned. See Chapter 24 [Job Control], page 479, for more information about job control, and Chapter 21 [Signal Handling], page 393, for information about signals.

The read function is the underlying primitive for all of the functions that read from streams, such as fgetc.

Function ssize_t write (int filedes, const void *buffer, size_t

size) The write function writes up to size bytes from buffer to the file with descriptor filedes. The data in buffer is not necessarily a character string and a null character is output like any other character. The return value is the number of bytes actually written. This may be size, but can always be smaller. Your program should always call write in a loop, iterating until all the data is written. Once write returns, the data is enqueued to be written and can be read back right away, but it is not necessarily written out to permanent storage immediately. You can use fsync when you need to be sure your data has been permanently stored before continuing. (It is more efficient for the system to batch up consecutive writes and do them all at once when convenient. Normally they will always be written to disk within a minute or less.) You can use the O_FSYNC open mode to make write always store the data to disk before returning; see Section 8.10.3 [I/O Operating Modes], page 173. In the case of an error, write returns -1. The following errno error conditions are defined for this function:

EAGAIN Normally, write blocks until the write operation is complete. But if the O_NONBLOCK flag is set for the file (see Section 8.7 [Control Operations on Files], page 165), it returns immediately without writing any data, and reports this error. An example of a situation that might cause the process to block on output is writing to a terminal device that supports flow control, where output has been suspended by receipt of a STOP character. Compatibility Note: Most versions of BSD Unix use a different error code for this: EWOULDBLOCK. In the GNU library, EWOULDBLOCK is an alias for EAGAIN, so it doesn’t matter which name you use. On some systems, writing a large amount of data from a character special file can also fail with EAGAIN if the kernel cannot find enough physical memory to lock down the user’s pages. This is limited to devices that transfer with direct memory access into the user’s memory, which means it does not include terminals, since they always use separate buffers inside the kernel. This problem does not arise in the GNU system.

EBADF The filedes argument is not a valid file descriptor, or is not open for writing.

EFBIG The size of the file would become larger than the implementation can support.

EINTR The write operation was interrupted by a signal while it was blocked waiting for completion. A signal will not necessary cause write to return EINTR; it may instead result in a successful write which writes fewer bytes than requested. See Section 21.5 [Primitives Interrupted by Signals], page 425.

EIO For many devices, and for disk files, this error code indicates a hardware error.

ENOSPC The device containing the file is full.

EPIPE This error is returned when you try to write to a pipe or FIFO that isn’t open for reading by any process. When this happens, a SIGPIPE signal is also sent to the process; see Chapter 21 [Signal Handling], page 393.

Unless you have arranged to prevent EINTR failures, you should check errno after each failing call to write, and if the error was EINTR, you should simply repeat the call. See Section 21.5 [Primitives Interrupted by Signals], page 425. The easy way to do this is with the macro TEMP_

FAILURE_RETRY, as follows:
nbytes = TEMP_FAILURE_RETRY (write (desc, buffer, count));

The write function is the underlying primitive for all of the functions that write to streams, such as fputc.

8.3 Setting the File Position of a Descriptor

Just as you can set the file position of a stream with fseek, you can set the file position of a descriptor with lseek. This specifies the position in the file for the next read or write operation. See Section 7.15 [File Positioning], page 137, for more information on the file position and what it means.

To read the current file position value from a descriptor, use lseek (desc,

0, SEEK_CUR).

Function off_t lseek (int filedes, off_t offset, int whence)

The lseek function is used to change the file position of the file with descriptor filedes. The whence argument specifies how the offset should be interpreted in the same way as for the fseek function, and must be one of the

symbolic constants SEEK_SET, SEEK_CUR, or SEEK_END.

SEEK_SET Specifies that whence is a count of characters from the beginning of the file.

SEEK_CUR Specifies that whence is a count of characters from the current file position. This count may be positive or negative.

SEEK_END Specifies that whence is a count of characters from the end of the file. A negative count specifies a position within the current extent of the file; a positive count specifies a position past the current end. If you set the position past the current end, and actually write data, you will extend the file with zeros up to that position.

The return value from lseek is normally the resulting file position, measured in bytes from the beginning of the file. You can use this feature together with SEEK_CUR to read the current file position. If you want to append to the file, setting the file position to the current end of file with SEEK_END is not sufficient. Another process may write more data after you seek but before you write, extending the file so the position you write onto clobbers their data. Instead, use the O_APPEND operating mode; see Section 8.10.3 [I/O Operating Modes], page 173. You can set the file position past the current end of the file. This does not by itself make the file longer; lseek never changes the file. But subsequent output at that position will extend the file. Characters between the previous end of file and the new position are filled with zeros. Extending the file in this way can create a “hole”: the blocks of zeros are not actually allocated on disk, so the file takes up less space than it appears so; it is then called a “sparse file”. If the file position cannot be changed, or the operation is in some way invalid, lseek returns a value of -1. The following errno error conditions are defined for this function:

EBADF The filedes is not a valid file descriptor.

EINVAL The whence argument value is not valid, or the resulting file offset is not valid. A file offset is invalid.

ESPIPE The filedes corresponds to an object that cannot be positioned, such as a pipe, FIFO or terminal device. (POSIX.1 specifies this error only for pipes and FIFOs, but in the GNU system, you always get ESPIPE if the object is not seekable.)

The lseek function is the underlying primitive for the fseek, ftell and rewind functions, which operate on streams instead of file descriptors.

You can have multiple descriptors for the same file if you open the file more than once, or if you duplicate a descriptor with dup. Descriptors that come from separate calls to open have independent file positions; using lseek on one descriptor has no effect on the other. For example,

     {
       int d1, d2;
       char buf[4];
       d1 = open ("foo", O_RDONLY);
       d2 = open ("foo", O_RDONLY);
       lseek (d1, 1024, SEEK_SET);
       read (d2, buf, 4);
     }

will read the first four characters of the file ‘foo’. (The error-checking code necessary for a real program has been omitted here for brevity.)

By contrast, descriptors made by duplication share a common file position with the original descriptor that was duplicated. Anything which alters the file position of one of the duplicates, including reading or writing data, affects all of them alike. Thus, for example,

     {
       int d1, d2, d3;
       char buf1[4], buf2[4];
       d1 = open ("foo", O_RDONLY);
       d2 = dup (d1);
       d3 = dup (d2);
       lseek (d3, 1024, SEEK_SET);
       read (d1, buf1, 4);
       read (d2, buf2, 4);
     }

will read four characters starting with the 1024’th character of ‘foo’, and then four more characters starting with the 1028’th character.

Data Type off t

This is an arithmetic data type used to represent file sizes. In the GNU system, this is equivalent to fpos_t or long int.

These aliases for the ‘SEEK_…’ constants exist for the sake of compatibility with older BSD systems. They are defined in two different header files:

‘fcntl.h’ and ‘sys/file.h’.

L_SET  An alias for SEEK_SET.
L_INCR  An alias for SEEK_CUR.
L_XTND  An alias for SEEK_END.

8.4 Descriptors and Streams

Given an open file descriptor, you can create a stream for it with the fdopen function. You can get the underlying file descriptor for an existing stream with the fileno function. These functions are declared in the header

file ‘stdio.h’.

Function FILE * fdopen (int filedes, const char *opentype)

The fdopen function returns a new stream for the file descriptor filedes. The opentype argument is interpreted in the same way as for the fopen function (see Section 7.3 [Opening Streams], page 94), except that the ‘b’ option is not permitted; this is because GNU makes no distinction between text and binary files. Also, "w" and "w+" do not cause truncation of the file; these have affect only when opening a file, and in this case the file has already been opened. You must make sure that the opentype argument matches the actual mode of the open file descriptor. The return value is the new stream. If the stream cannot be created (for example, if the modes for the file indicated by the file descriptor do not permit the access specified by the opentype argument), a null pointer is returned instead. In some other systems, fdopen may fail to detect that the modes for file descriptor do not permit the access specified by opentype. The GNU C library always checks for this.

For an example showing the use of the fdopen function, see Section 10.1 [Creating a Pipe], page 215.

Function int fileno (FILE *stream)

This function returns the file descriptor associated with the stream stream. If an error is detected (for example, if the stream is not valid) or if stream does not do I/O to a file, fileno returns -1.

There are also symbolic constants defined in ‘unistd.h’ for the file descriptors belonging to the standard streams stdin, stdout, and stderr; see Section 7.2 [Standard Streams], page 93.

STDIN_FILENO

This macro has value 0, which is the file descriptor for standard input.

STDOUT_FILENO

This macro has value 1, which is the file descriptor for standard output.

STDERR_FILENO

This macro has value 2, which is the file descriptor for standard error output.

8.5 Dangers of Mixing Streams and Descriptors

You can have multiple file descriptors and streams (let’s call both streams and descriptors “channels” for short) connected to the same file, but you must take care to avoid confusion between channels. There are two cases to consider: linked channels that share a single file position value, and independent channels that have their own file positions.

It’s best to use just one channel in your program for actual data transfer to any given file, except when all the access is for input. For example, if you open a pipe (something you can only do at the file descriptor level), either do all I/O with the descriptor, or construct a stream from the descriptor with fdopen and then do all I/O with the stream.

8.5.1 Linked Channels

Channels that come from a single opening share the same file position; we call them linked channels. Linked channels result when you make a stream from a descriptor using fdopen, when you get a descriptor from a stream with fileno, when you copy a descriptor with dup or dup2, and when descriptors are inherited during fork. For files that don’t support random access, such as terminals and pipes, all channels are effectively linked. On random-access files, all append-type output streams are effectively linked to each other.

If you have been using a stream for I/O, and you want to do I/O using another channel (either a stream or a descriptor) that is linked to it, you must first clean up the stream that you have been using. See Section 8.5.3 [Cleaning Streams], page 161.

Terminating a process, or executing a new program in the process, destroys all the streams in the process. If descriptors linked to these streams persist in other processes, their file positions become undefined as a result. To prevent this, you must clean up the streams before destroying them.

8.5.2 Independent Channels

When you open channels (streams or descriptors) separately on a seekable file, each channel has its own file position. These are called independent channels.

The system handles each channel independently. Most of the time, this is quite predictable and natural (especially for input): each channel can read or write sequentially at its own place in the file. However, if some of the channels are streams, you must take these precautions:

• You should clean an output stream after use, before doing anything else that might read or write from the same part of the file.

• You should clean an input stream before reading data that may have been modified using an independent channel. Otherwise, you might read obsolete data that had been in the stream’s buffer.

If you do output to one channel at the end of the file, this will certainly leave the other independent channels positioned somewhere before the new end. You cannot reliably set their file positions to the new end of file before writing, because the file can always be extended by another process between when you set the file position and when you write the data. Instead, use an append-type descriptor or stream; they always output at the current end of the file. In order to make the end-of-file position accurate, you must clean the output channel you were using, if it is a stream.

It’s impossible for two channels to have separate file pointers for a file that doesn’t support random access. Thus, channels for reading or writing such files are always linked, never independent. Append-type channels are also always linked. For these channels, follow the rules for linked channels; see Section 8.5.1 [Linked Channels], page 160.

8.5.3 Cleaning Streams

On the GNU system, you can clean up any stream with fclean:

Function int fclean (FILE *stream)

Clean up the stream stream so that its buffer is empty. If stream is doing output, force it out. If stream is doing input, give the data in the buffer back to the system, arranging to reread it.

On other systems, you can use fflush to clean a stream in most cases. You can skip the fclean or fflush if you know the stream is already

clean. A stream is clean whenever its buffer is empty. For example, an unbuffered stream is always clean. An input stream that is at end-of-file is clean. A line-buffered stream is clean when the last character output was a newline.

There is one case in which cleaning a stream is impossible on most systems. This is when the stream is doing input from a file that is not randomaccess. Such streams typically read ahead, and when the file is not random access, there is no way to give back the excess data already read. When an input stream reads from a random-access file, fflush does clean the stream, but leaves the file pointer at an unpredictable place; you must set the file pointer before doing any further I/O. On the GNU system, using fclean avoids both of these problems.

Closing an output-only stream also does fflush, so this is a valid way of cleaning an output stream. On the GNU system, closing an input stream

does fclean.

You need not clean a stream before using its descriptor for control operations such as setting terminal modes; these operations don’t affect the file position and are not affected by it. You can use any descriptor for these operations, and all channels are affected simultaneously. However, text already “output” to a stream but still buffered by the stream will be subject to the new terminal modes when subsequently flushed. To make sure “past” output is covered by the terminal settings that were in effect at the time, flush the output streams for that terminal before setting the modes. See Section 12.4 [Terminal Modes], page 273.

8.6 Waiting for Input or Output

Sometimes a program needs to accept input on multiple input channels whenever input arrives. For example, some workstations may have devices such as a digitizing tablet, function button box, or dial box that are connected via normal asynchronous serial interfaces; good user interface style requires responding immediately to input on any device. Another example is a program that acts as a server to several other processes via pipes or sockets.

You cannot normally use read for this purpose, because this blocks the program until input is available on one particular file descriptor; input on other channels won’t wake it up. You could set nonblocking mode and poll each file descriptor in turn, but this is very inefficient.

A better solution is to use the select function. This blocks the program until input or output is ready on a specified set of file descriptors, or until a timer expires, whichever comes first. This facility is declared in the header

file ‘sys/types.h’.

In the case of a server socket (see Section 11.8.2 [Listening for Connections], page 248), we say that “input” is available when there are pending connections that could be accepted (see Section 11.8.3 [Accepting Connections], page 248). accept for server sockets blocks and interacts with select just as read does for normal input.

The file descriptor sets for the select function are specified as fd_set objects. Here is the description of the data type and some macros for manipulating these objects.

Data Type fd set

The fd_set data type represents file descriptor sets for the select function. It is actually a bit array.

Macro int FD SETSIZE

The value of this macro is the maximum number of file descriptors that a fd_set object can hold information about. On systems with a fixed maximum number, FD_SETSIZE is at least that number. On some systems, including GNU, there is no absolute limit on the number of descriptors open, but this macro still has a constant value which controls the number of bits in an fd_set; if you get a file descriptor with a value as high as FD_SETSIZE, you cannot put that descriptor into an fd_set.

Macro void FD ZERO (fd_set *set)

This macro initializes the file descriptor set set to be the empty set.

Macro void FD SET (int filedes, fd_set *set)

This macro adds filedes to the file descriptor set set.

Macro void FD CLR (int filedes, fd_set *set)

This macro removes filedes from the file descriptor set set.

Macro int FD ISSET (int filedes, fd_set *set)

This macro returns a nonzero value (true) if filedes is a member of the the file descriptor set set, and zero (false) otherwise.

Next, here is the description of the select function itself.

Function int select (int nfds, fd_set *read-fds, fd_set

*write-fds, fd_set *except-fds, struct timeval *timeout) The select function blocks the calling process until there is activity on any of the specified sets of file descriptors, or until the timeout period has expired. The file descriptors specified by the read-fds argument are checked to see if they are ready for reading; the write-fds file descriptors are checked to see if they are ready for writing; and the except-fds file descriptors are checked for exceptional conditions. You can pass a null pointer for any of these arguments if you are not interested in checking for that kind of condition. A file descriptor is considered ready for reading if it is at end of file. A server socket is considered ready for reading if there is a pending connection which can be accepted with accept; see Section 11.8.3 [Accepting Connections], page 248. A client socket is ready for writing when its connection is fully established; see Section 11.8.1 [Making a Connection], page 246. “Exceptional conditions” does not mean errors—errors are reported immediately when an erroneous system call is executed, and do not constitute a state of the descriptor. Rather, they include conditions such as the presence of an urgent message on a socket. (See Chapter 11 [Sockets], page 221, for information on urgent messages.)

The select function checks only the first nfds file descriptors. The usual thing is to pass FD_SETSIZE as the value of this argument. The timeout specifies the maximum time to wait. If you pass a null pointer for this argument, it means to block indefinitely until one of the file descriptors is ready. Otherwise, you should provide the time in struct timeval format; see Section 17.2.2 [High-Resolution Calendar], page 342. Specify zero as the time (a struct timeval containing all zeros) if you want to find out which descriptors are ready without waiting if none are ready. The normal return value from select is the total number of ready file descriptors in all of the sets. Each of the argument sets is overwritten with information about the descriptors that are ready for the corresponding operation. Thus, to see if a particular descriptor desc has input, use FD_ISSET (desc, read-fds) after select returns. If select returns because the timeout period expires, it returns a value of zero. Any signal will cause select to return immediately. So if your program uses signals, you can’t rely on select to keep waiting for the full time specified. If you want to be sure of waiting for a particular amount of time, you must check for EINTR and repeat the select with a newly calculated timeout based on the current time. See the example below. See also Section 21.5 [Primitives Interrupted by Signals], page 425. If an error occurs, select returns -1 and does not modify the argument file descriptor sets. The following errno error conditions are defined for this function:

EBADF One of the file descriptor sets specified an invalid file descriptor.

EINTR The operation was interrupted by a signal. See Section 21.5 [Primitives Interrupted by Signals], page 425.

EINVAL The timeout argument is invalid; one of the components is negative or too large.

Portability Note: The select function is a BSD Unix feature. Here is an example showing how you can use select to establish a timeout

period for reading from a file descriptor. The input_timeout function blocks the calling process until input is available on the file descriptor, or until the timeout period expires.

  #include <stdio.h>
  #include <unistd.h>
  #include <sys/types.h>
  #include <sys/time.h>

     int
     input_timeout (int filedes, unsigned int seconds)
     {
       fd_set set;
       struct timeval timeout;

/* Initialize the file descriptor set. */

       FD_ZERO (&set);
       FD_SET (filedes, &set);

/* Initialize the timeout data structure. */

       timeout.tv_sec = seconds;
       timeout.tv_usec = 0;

/* select returns 0 if timeout, 1 if input available, -1 if error. */

       return TEMP_FAILURE_RETRY (select (FD_SETSIZE,
                                          &set, NULL, NULL,
                                          &timeout));
     }
     int
     main (void)
     {
       fprintf (stderr, "select returned %d.\n",
                input_timeout (STDIN_FILENO, 5));
       return 0;
     }

There is another example showing the use of select to multiplex input from multiple sockets in Section 11.8.7 [Byte Stream Connection Server Example], page 254.

8.7 Control Operations on Files

This section describes how you can perform various other operations on file descriptors, such as inquiring about or setting flags describing the status of the file descriptor, manipulating record locks, and the like. All of these operations are performed by the function fcntl.

The second argument to the fcntl function is a command that specifies which operation to perform. The function and macros that name various flags that are used with it are declared in the header file ‘fcntl.h’. Many of these flags are also used by the open function; see Section 8.1 [Opening and Closing Files], page 151.

Function int fcntl (int filedes, int command, …)

The fcntl function performs the operation specified by command on the file descriptor filedes. Some commands require additional arguments to be supplied. These additional arguments and the return value and error conditions are given in the detailed descriptions of the individual commands. Briefly, here is a list of what the various commands are. F_DUPFD Duplicate the file descriptor (return another file descriptor

pointing to the same open file). See Section 8.8 [Duplicating Descriptors], page 166.

F_GETFD Get flags associated with the file descriptor. See Section 8.9 [File Descriptor Flags], page 168.

F_SETFD Set flags associated with the file descriptor. See Section 8.9 [File Descriptor Flags], page 168.

F_GETFL Get flags associated with the open file. See Section 8.10 [File Status Flags], page 170.

F_SETFL Set flags associated with the open file. See Section 8.10 [File Status Flags], page 170.

F_GETLK Get a file lock. See Section 8.11 [File Locks], page 175. F_SETLK Set or clear a file lock. See Section 8.11 [File Locks],

page 175.

F_SETLKW Like F_SETLK, but wait for completion. See Section 8.11 [File Locks], page 175.

F_GETOWN Get process or process group ID to receive SIGIO signals. See Section 8.12 [Interrupt-Driven Input], page 179.

F_SETOWN Set process or process group ID to receive SIGIO signals. See Section 8.12 [Interrupt-Driven Input], page 179.

8.8 Duplicating Descriptors

You can duplicate a file descriptor, or allocate another file descriptor that refers to the same open file as the original. Duplicate descriptors share one file position and one set of file status flags (see Section 8.10 [File Status Flags], page 170), but each has its own set of file descriptor flags (see Section 8.9 [File Descriptor Flags], page 168).

The major use of duplicating a file descriptor is to implement redirection of input or output: that is, to change the file or pipe that a particular file descriptor corresponds to.

You can perform this operation using the fcntl function with the F_ DUPFD command, but there are also convenient functions dup and dup2 for duplicating descriptors.

The fcntl function and flags are declared in ‘fcntl.h’, while prototypes for dup and dup2 are in the header file ‘unistd.h’.

Function int dup (int old)

This function copies descriptor old to the first available descriptor number (the first number not currently open). It is equivalent to

fcntl (old, F_DUPFD, 0).

Function int dup2 (int old, int new)

This function copies the descriptor old to descriptor number new. If old is an invalid descriptor, then dup2 does nothing; it does not close new. Otherwise, the new duplicate of old replaces any previous meaning of descriptor new, as if new were closed first. If old and new are different numbers, and old is a valid descriptor number, then dup2 is equivalent to:

     close (new);
     fcntl (old, F_DUPFD, new)

However, dup2 does this atomically; there is no instant in the middle of calling dup2 at which new is closed and not yet a duplicate of old.

Macro int F DUPFD

This macro is used as the command argument to fcntl, to copy the file descriptor given as the first argument. The form of the call in this case is:

     fcntl (old, F_DUPFD, next-filedes)

The next-filedes argument is of type int and specifies that the file descriptor returned should be the next available one greater than or equal to this value. The return value from fcntl with this command is normally the value of the new file descriptor. A return value of -1 indicates an error. The following errno error conditions are defined for this command:

EBADF The old argument is invalid.

EINVAL The next-filedes argument is invalid.

EMFILE There are no more file descriptors available—your program is already using the maximum. In BSD and GNU, the maximum is controlled by a resource limit that can be changed; see Section 17.6 [Limiting Resource Usage], page 362, for more information about the RLIMIT_NOFILE limit.

ENFILE is not a possible error code for dup2 because dup2 does not create a new opening of a file; duplicate descriptors do not count toward the limit which ENFILE indicates. EMFILE is possible because it refers to the limit on distinct descriptor numbers in use in one process.

Here is an example showing how to use dup2 to do redirection. Typically, redirection of the standard streams (like stdin) is done by a shell or shell-like program before calling one of the exec functions (see Section 23.5 [Executing a File], page 470) to execute a new program in a child process. When the new program is executed, it creates and initializes the standard streams to point to the corresponding file descriptors, before its main function is invoked.

So, to redirect standard input to a file, the shell could do something like:

     pid = fork ();
     if (pid == 0)
       {
         char *filename;
         char *program;
         int file;

…

         file = TEMP_FAILURE_RETRY (open (filename, O_RDONLY));
         dup2 (file, STDIN_FILENO);
         TEMP_FAILURE_RETRY (close (file));
         execv (program, NULL);
       }

There is also a more detailed example showing how to implement redirection in the context of a pipeline of processes in Section 24.6.3 [Launching Jobs], page 486.

8.9 File Descriptor Flags

File descriptor flags are miscellaneous attributes of a file descriptor. These flags are associated with particular file descriptors, so that if you have created duplicate file descriptors from a single opening of a file, each descriptor has its own set of flags.

Currently there is just one file descriptor flag: FD_CLOEXEC, which causes the descriptor to be closed if you use any of the exec… functions (see Section 23.5 [Executing a File], page 470).

The symbols in this section are defined in the header file ‘fcntl.h’.

Macro int F GETFD

This macro is used as the command argument to fcntl, to specify that it should return the file descriptor flags associated with the filedes argument. The normal return value from fcntl with this command is a nonnegative number which can be interpreted as the bitwise OR of the individual flags (except that currently there is only one flag to use). In case of an error, fcntl returns -1. The following errno error conditions are defined for this command:

EBADF The filedes argument is invalid.

Macro int F SETFD

This macro is used as the command argument to fcntl, to specify that it should set the file descriptor flags associated with the filedes argument. This requires a third int argument to specify the new flags, so the form of the call is:

     fcntl (filedes, F_SETFD, new-flags)

The normal return value from fcntl with this command is an unspecified value other than -1, which indicates an error. The flags and error conditions are the same as for the F_GETFD command.

The following macro is defined for use as a file descriptor flag with the fcntl function. The value is an integer constant usable as a bit mask value.

Macro int FD CLOEXEC

This flag specifies that the file descriptor should be closed when an exec function is invoked; see Section 23.5 [Executing a File], page 470. When a file descriptor is allocated (as with open or dup), this bit is initially cleared on the new file descriptor, meaning that descriptor will survive into the new program after exec.

If you want to modify the file descriptor flags, you should get the current flags with F_GETFD and modify the value. Don’t assume that the flags listed here are the only ones that are implemented; your program may be run years from now and more flags may exist then. For example, here is a function to set or clear the flag FD_CLOEXEC without altering any other flags:

/* Set the FD_CLOEXEC flag of desc if value is nonzero, or clear the flag if value is 0. Return 0 on success, or -1 on error with errno set. */

  int
  set_cloexec_flag (int desc, int value)
  {
    int oldflags = fcntl (desc, F_GETFD, 0);

/* If reading the flags failed, return error indication now.

    if (oldflags < 0)
      return oldflags;

/* Set just the flag we want to set. */

    if (value != 0)
      oldflags |= FD_CLOEXEC;
    else
      oldflags &= ~FD_CLOEXEC;

/* Store modified flag word in the descriptor. */

    return fcntl (desc, F_SETFD, oldflags);
  }

8.10 File Status Flags

File status flags are used to specify attributes of the opening of a file. Unlike the file descriptor flags discussed in Section 8.9 [File Descriptor Flags], page 168, the file status flags are shared by duplicated file descriptors resulting from a single opening of the file. The file status flags are specified with the flags argument to open; see Section 8.1 [Opening and Closing Files], page 151.

File status flags fall into three categories, which are described in the following sections.

• Section 8.10.1 [File Access Modes], page 170, specify what type of access is allowed to the file: reading, writing, or both. They are set by open and are returned by fcntl, but cannot be changed.

• Section 8.10.2 [Open-time Flags], page 171, control details of what open will do. These flags are not preserved after the open call.

• Section 8.10.3 [I/O Operating Modes], page 173, affect how operations such as read and write are done. They are set by open, and can be fetched or changed with fcntl. The symbols in this section are defined in the header file ‘fcntl.h’.

8.10.1 File Access Modes

The file access modes allow a file descriptor to be used for reading, writing, or both. (In the GNU system, they can also allow none of these, and allow execution of the file as a program.) The access modes are chosen when the file is opened, and never change.

Macro int O RDONLY

Open the file for read access.

Macro int O WRONLY

Open the file for write access.

Macro int O RDWR

Open the file for both reading and writing.

In the GNU system (and not in other systems), O_RDONLY and O_WRONLY are independent bits that can be bitwise-ORed together, and it is valid for either bit to be set or clear. This means that O_RDWR is the same as O_ RDONLY|O_WRONLY. A file access mode of zero is permissible; it allows no operations that do input or output to the file, but does allow other operations such as fchmod. On the GNU system, since “read-only” or “write-only” is a misnomer, ‘fcntl.h’ defines additional names for the file access modes. These names are preferred when writing GNU-specific code. But most programs will want to be portable to other POSIX.1 systems and should use the POSIX.1 names above instead.

Macro int O READ

Open the file for reading. Same as O_RDWR; only defined on GNU.

Macro int O WRITE

Open the file for reading. Same as O_WRONLY; only defined on GNU.

Macro int O EXEC

Open the file for executing. Only defined on GNU.

To determine the file access mode with fcntl, you must extract the access mode bits from the retrieved file status flags. In the GNU system, you can just test the O_READ and O_WRITE bits in the flags word. But in other POSIX.1 systems, reading and writing access modes are not stored as distinct bit flags. The portable way to extract the file access mode bits is

with O_ACCMODE.

Macro int O ACCMODE

This macro stands for a mask that can be bitwise-ANDed with the file status flag value to produce a value representing the file access mode. The mode will be O_RDONLY, O_WRONLY, or O_RDWR. (In the GNU system it could also be zero, and it never includes the O_EXEC bit.)

8.10.2 Open-time Flags

The open-time flags specify options affecting how open will behave. These options are not preserved once the file is open. The exception to this is O_ NONBLOCK, which is also an I/O operating mode and so it is saved. See Section 8.1 [Opening and Closing Files], page 151, for how to call open.

There are two sorts of options specified by open-time flags.

• File name translation flags affect how open looks up the file name to locate the file, and whether the file can be created.

• Open-time action flags specify extra operations that open will perform on the file once it is open.

Here are the file name translation flags.

Macro int O CREAT

If set, the file will be created if it doesn’t already exist.

Macro int O EXCL

If both O_CREAT and O_EXCL are set, then open fails if the specified file already exists. This is guaranteed to never clobber an existing file.

Macro int O NONBLOCK

This prevents open from blocking for a “long time” to open the file. This is only meaningful for some kinds of files, usually devices such as serial ports; when it is not meaningful, it is harmless and ignored. Often opening a port to a modem blocks until the modem reports carrier detection; if O_NONBLOCK is specified, open will return immediately without a carrier. Note that the O_NONBLOCK flag is overloaded as both an I/O operating mode and a file name translation flag. This means that specifying O_ NONBLOCK in open also sets nonblocking I/O mode; see Section 8.10.3 [I/O Operating Modes], page 173. To open the file without blocking but do normal I/O that blocks, you must call open with O_NONBLOCK set and then call fcntl to turn the bit off.

Macro int O NOCTTY

If the named file is a terminal device, don’t make it the controlling terminal for the process. See Chapter 24 [Job Control], page 479, for information about what it means to be the controlling terminal. In the GNU system and 4.4 BSD, opening a file never makes it the controlling terminal and O_NOCTTY is zero. However, other systems may use a nonzero value for O_NOCTTY and set the controlling terminal when you open a file that is a terminal device; so to be portable, use O_NOCTTY when it is important to avoid this.

The following three file name translation flags exist only in the GNU system.

Macro int O IGNORE CTTY

Do not recognize the named file as the controlling terminal, even if it refers to the process’s existing controlling terminal device. Operations on the new file descriptor will never induce job control signals. See Chapter 24 [Job Control], page 479.

Macro int O NOLINK

If the named file is a symbolic link, open the link itself instead of the file it refers to. (fstat on the new file descriptor will return the information returned by lstat on the link’s name.)

Macro int O NOTRANS

If the named file is specially translated, do not invoke the translator. Open the bare file the translator itself sees.

The open-time action flags tell open to do additional operations which are not really related to opening the file. The reason to do them as part of open instead of in separate calls is that open can do them atomically.

Macro int O TRUNC

Truncate the file to zero length. This option is only useful for regular files, not special files such as directories or FIFOs. POSIX.1 requires that you open the file for writing to use O_TRUNC. In BSD and GNU you must have permission to write the file to truncate it, but you need not open for write access. This is the only open-time action flag specified by POSIX.1. There is no good reason for truncation to be done by open, instead of by calling ftruncate afterwards. The O_TRUNC flag existed in Unix before ftruncate was invented, and is retained for backward compatibility.

Macro int O SHLOCK

Acquire a shared lock on the file, as with flock. See Section 8.11 [File Locks], page 175. If O_CREAT is specified, the locking is done atomically when creating the file. You are guaranteed that no other process will get the lock on the new file first.

Macro int O EXLOCK

Acquire an exclusive lock on the file, as with flock. See Section 8.11 [File Locks], page 175. This is atomic like O_SHLOCK.

8.10.3 I/O Operating Modes

The operating modes affect how input and output operations using a file descriptor work. These flags are set by open and can be fetched and changed

with fcntl.

Macro int O APPEND

The bit that enables append mode for the file. If set, then all write operations write the data at the end of the file, extending it, regardless of the current file position. This is the only reliable way to append to a file. In append mode, you are guaranteed that the data you write will always go to the current end of the file, regardless of other processes writing to the file. Conversely, if you simply set the file position to the end of file and write, then another process can extend the file after you set the file position but before you write, resulting in your data appearing someplace before the real end of file.

Macro int O NONBLOCK

The bit that enables nonblocking mode for the file. If this bit is set, read requests on the file can return immediately with a failure status if there is no input immediately available, instead of blocking. Likewise, write requests can also return immediately with a failure status if the output can’t be written immediately. Note that the O_NONBLOCK flag is overloaded as both an I/O operating mode and a file name translation flag; see Section 8.10.2 [Open-time Flags], page 171.

Macro int O NDELAY

This is an obsolete name for O_NONBLOCK, provided for compatibility with BSD. It is not defined by the POSIX.1 standard.

The remaining operating modes are BSD and GNU extensions. They exist only on some systems. On other systems, these macros are not defined.

Macro int O ASYNC

The bit that enables asynchronous input mode. If set, then SIGIO signals will be generated when input is available. See Section 8.12 [Interrupt-Driven Input], page 179. Asynchronous input mode is a BSD feature.

Macro int O FSYNC

The bit that enables synchronous writing for the file. If set, each write call will make sure the data is reliably stored on disk before returning. Synchronous writing is a BSD feature.

Macro int O SYNC

This is another name for O_FSYNC. They have the same value.

Macro int O NOATIME

If this bit is set, read will not update the access time of the file. See Section 9.8.9 [File Times], page 208. This is used by programs that do backups, so that backing a file up does not count as reading it. Only the owner of the file or the superuser may use this bit. This is a GNU extension.

8.10.4 Getting and Setting File Status Flags

The fcntl function can fetch or change file status flags.

Macro int F GETFL

This macro is used as the command argument to fcntl, to read the file status flags for the open file with descriptor filedes. The normal return value from fcntl with this command is a nonnegative number which can be interpreted as the bitwise OR of the individual flags. Since the file access modes are not single-bit values, you can mask off other bits in the returned flags with O_ACCMODE to compare them. In case of an error, fcntl returns -1. The following errno error conditions are defined for this command:

EBADF The filedes argument is invalid.

Macro int F SETFL

This macro is used as the command argument to fcntl, to set the file status flags for the open file corresponding to the filedes argument. This command requires a third int argument to specify the new flags, so the call looks like this:

     fcntl (filedes, F_SETFL, new-flags)

You can’t change the access mode for the file in this way; that is, whether the file descriptor was opened for reading or writing. The normal return value from fcntl with this command is an unspecified value other than -1, which indicates an error. The error conditions are the same as for the F_GETFL command.

If you want to modify the file status flags, you should get the current flags with F_GETFL and modify the value. Don’t assume that the flags listed here are the only ones that are implemented; your program may be run years from now and more flags may exist then. For example, here is a function to set or clear the flag O_NONBLOCK without altering any other flags:

/* Set the O_NONBLOCK flag of desc if value is nonzero, or clear the flag if value is 0. Return 0 on success, or -1 on error with errno set. */

  int
  set_nonblock_flag (int desc, int value)
  {
    int oldflags = fcntl (desc, F_GETFL, 0);

/* If reading the flags failed, return error indication now. */

    if (oldflags == -1)
      return -1;

/* Set just the flag we want to set. */

    if (value != 0)
      oldflags |= O_NONBLOCK;
    else
      oldflags &= ~O_NONBLOCK;

/* Store modified flag word in the descriptor. */

    return fcntl (desc, F_SETFL, oldflags);
  }

8.11 File Locks

The remaining fcntl commands are used to support record locking, which permits multiple cooperating programs to prevent each other from simultaneously accessing parts of a file in error-prone ways.

An exclusive or write lock gives a process exclusive access for writing to the specified part of the file. While a write lock is in place, no other process can lock that part of the file.

A shared or read lock prohibits any other process from requesting a write lock on the specified part of the file. However, other processes can request read locks.

The read and write functions do not actually check to see whether there are any locks in place. If you want to implement a locking protocol for a file shared by multiple processes, your application must do explicit fcntl calls to request and clear locks at the appropriate points.

Locks are associated with processes. A process can only have one kind of lock set for each byte of a given file. When any file descriptor for that file is closed by the process, all of the locks that process holds on that file are released, even if the locks were made using other descriptors that remain open. Likewise, locks are released when a process exits, and are not inherited by child processes created using fork (see Section 23.4 [Creating a Process], page 469).

When making a lock, use a struct flock to specify what kind of lock and where. This data type and the associated macros for the fcntl function are declared in the header file ‘fcntl.h’.

Data Type struct flock

This structure is used with the fcntl function to describe a file lock. It has these members:

  short int l_type

Specifies the type of the lock; one of F_RDLCK, F_WRLCK,

             or F_UNLCK.
  short int l_whence

This corresponds to the whence argument to fseek or lseek, and specifies what the offset is relative to. Its value

             can be one of SEEK_SET, SEEK_CUR, or SEEK_END.
  off_t l_start

This specifies the offset of the start of the region to which the lock applies, and is given in bytes relative to the point specified by l_whence member.

  off_t l_len

This specifies the length of the region to be locked. A value of 0 is treated specially; it means the region extends to the end of the file.

  pid_t l_pid

This field is the process ID (see Section 23.2 [Process Creation Concepts], page 468) of the process holding the lock.

It is filled in by calling fcntl with the F_GETLK command, but is ignored when making a lock.

Macro int F GETLK

This macro is used as the command argument to fcntl, to specify that it should get information about a lock. This command requires a third argument of type struct flock * to be passed to fcntl, so that the form of the call is:

     fcntl (filedes, F_GETLK, lockp)

If there is a lock already in place that would block the lock described by the lockp argument, information about that lock overwrites *lockp. Existing locks are not reported if they are compatible with making a new lock as specified. Thus, you should specify a lock type of F_WRLCK if you want to find out about both read and write locks, or F_RDLCK if you want to find out about write locks only. There might be more than one lock affecting the region specified by the lockp argument, but fcntl only returns information about one of them. The l_whence member of the lockp structure is set to SEEK_SET and the l_start and l_len fields set to identify the locked region. If no lock applies, the only change to the lockp structure is to update the l_type to a value of F_UNLCK. The normal return value from fcntl with this command is an unspecified value other than -1, which is reserved to indicate an error. The following errno error conditions are defined for this command:

EBADF The filedes argument is invalid.

EINVAL Either the lockp argument doesn’t specify valid lock information, or the file associated with filedes doesn’t support locks.

Macro int F SETLK

This macro is used as the command argument to fcntl, to specify that it should set or clear a lock. This command requires a third argument of type struct flock * to be passed to fcntl, so that the form of the call is:

     fcntl (filedes, F_SETLK, lockp)

If the process already has a lock on any part of the region, the old lock on that part is replaced with the new lock. You can remove a lock by specifying a lock type of F_UNLCK. If the lock cannot be set, fcntl returns immediately with a value of -1. This function does not block waiting for other processes to release locks. If fcntl succeeds, it return a value other than -1. The following errno error conditions are defined for this function:

EAGAIN

EACCES The lock cannot be set because it is blocked by an existing lock on the file. Some systems use EAGAIN in this case, and other systems use EACCES; your program should treat them alike, after F_SETLK. (The GNU system always uses

EAGAIN.)

EBADF Either: the filedes argument is invalid; you requested a read lock but the filedes is not open for read access; or, you requested a write lock but the filedes is not open for write access.

EINVAL Either the lockp argument doesn’t specify valid lock information, or the file associated with filedes doesn’t support locks.

ENOLCK The system has run out of file lock resources; there are already too many file locks in place. Well-designed file systems never report this error, because they have no limitation on the number of locks. However, you must still take account of the possibility of this error, as it could result from network access to a file system on another machine.

Macro int F SETLKW

This macro is used as the command argument to fcntl, to specify that it should set or clear a lock. It is just like the F_SETLK command, but causes the process to block (or wait) until the request can be specified. This command requires a third argument of type struct flock *, as for the F_SETLK command. The fcntl return values and errors are the same as for the F_SETLK command, but these additional errno error conditions are defined for this command:

EINTR The function was interrupted by a signal while it was waiting. See Section 21.5 [Primitives Interrupted by Signals], page 425.

EDEADLK The specified region is being locked by another process. But that process is waiting to lock a region which the current process has locked, so waiting for the lock would result in deadlock. The system does not guarantee that it will detect all such conditions, but it lets you know if it notices one.

The following macros are defined for use as values for the l_type member of the flock structure. The values are integer constants.

F_RDLCK This macro is used to specify a read (or shared) lock.

F_WRLCK This macro is used to specify a write (or exclusive) lock.

F_UNLCK This macro is used to specify that the region is unlocked.

As an example of a situation where file locking is useful, consider a program that can be run simultaneously by several different users, that logs status information to a common file. One example of such a program might be a game that uses a file to keep track of high scores. Another example might be a program that records usage or accounting information for billing purposes.

Having multiple copies of the program simultaneously writing to the file could cause the contents of the file to become mixed up. But you can prevent this kind of problem by setting a write lock on the file before actually writing to the file.

If the program also needs to read the file and wants to make sure that the contents of the file are in a consistent state, then it can also use a read lock. While the read lock is set, no other process can lock that part of the file for writing.

Remember that file locks are only a voluntary protocol for controlling access to a file. There is still potential for access to the file by programs that don’t use the lock protocol.

8.12 Interrupt-Driven Input

If you set the O_ASYNC status flag on a file descriptor (see Section 8.10 [File Status Flags], page 170), a SIGIO signal is sent whenever input or output becomes possible on that file descriptor. The process or process group to receive the signal can be selected by using the F_SETOWN command to the fcntl function. If the file descriptor is a socket, this also selects the recipient of SIGURG signals that are delivered when out-of-band data arrives on that socket; see Section 11.8.8 [Out-of-Band Data], page 257. (SIGURG is sent in any situation where select would report the socket as having an “exceptional condition”. See Section 8.6 [Waiting for Input or Output], page 162.)

If the file descriptor corresponds to a terminal device, then SIGIO signals are sent to the foreground process group of the terminal. See Chapter 24 [Job Control], page 479.

The symbols in this section are defined in the header file ‘fcntl.h’.

Macro int F GETOWN

This macro is used as the command argument to fcntl, to specify that it should get information about the process or process group to which SIGIO signals are sent. (For a terminal, this is actually the foreground process group ID, which you can get using tcgetpgrp; see Section 24.7.3 [Functions for Controlling Terminal Access], page 499.) The return value is interpreted as a process ID; if negative, its absolute value is the process group ID. The following errno error condition is defined for this command:

EBADF The filedes argument is invalid.

Macro int F SETOWN

This macro is used as the command argument to fcntl, to specify that it should set the process or process group to which SIGIO signals are sent. This command requires a third argument of type pid_t to be passed to fcntl, so that the form of the call is:

     fcntl (filedes, F_SETOWN, pid)

The pid argument should be a process ID. You can also pass a negative number whose absolute value is a process group ID. The return value from fcntl with this command is -1 in case of error and some other value if successful. The following errno error conditions are defined for this command:

EBADF The filedes argument is invalid.

ESRCH There is no process or process group corresponding to pid.

9 File System Interface

This chapter describes the GNU C library’s functions for manipulating files. Unlike the input and output functions described in Chapter 7 [Input/Output on Streams], page 93 and Chapter 8 [Low-Level Input/Output], page 151, these functions are concerned with operating on the files themselves, rather than on their contents.

Among the facilities described in this chapter are functions for examining or modifying directories, functions for renaming and deleting files, and functions for examining and setting file attributes such as access permissions and modification times.

9.1 Working Directory

Each process has associated with it a directory, called its current working directory or simply working directory, that is used in the resolution of relative file names (see Section 6.2.2 [File Name Resolution], page 88).

When you log in and begin a new session, your working directory is initially set to the home directory associated with your login account in the system user database. You can find any user’s home directory using the getpwuid or getpwnam functions; see Section 26.12 [User Database], page 521.

Users can change the working directory using shell commands like cd. The functions described in this section are the primitives used by those commands and by other programs for examining and changing the working directory.

Prototypes for these functions are declared in the header file ‘unistd.h’.

Function char * getcwd (char *buffer, size_t size)

The getcwd function returns an absolute file name representing the current working directory, storing it in the character array buffer that you provide. The size argument is how you tell the system the allocation size of buffer. The GNU library version of this function also permits you to specify a null pointer for the buffer argument. Then getcwd allocates a buffer automatically, as with malloc (see Section 3.3 [Unconstrained Allocation], page 32). If the size is greater than zero, then the buffer is that large; otherwise, the buffer is as large as necessary to hold the result. The return value is buffer on success and a null pointer on failure. The following errno error conditions are defined for this function:

EINVAL The size argument is zero and buffer is not a null pointer.

ERANGE The size argument is less than the length of the working directory name. You need to allocate a bigger array and try again.

EACCES Permission to read or search a component of the file name was denied.

Here is an example showing how you could implement the behavior of GNU’s getcwd (NULL, 0) using only the standard behavior of getcwd:

  char *
  gnu_getcwd ()
  {
    int size = 100;
    char *buffer = (char *) xmalloc (size);
    while (1)
      {
        char *value = getcwd (buffer, size);
        if (value != 0)
          return buffer;
        size *= 2;
        free (buffer);
        buffer = (char *) xmalloc (size);
      }
  }

See Section 3.3.2 [Examples of “codemalloc], page 33, for information about xmalloc, which is not a library function but is a customary name used in most GNU software.

Function char * getwd (char *buffer)

This is similar to getcwd, but has no way to specify the size of the buffer. The GNU library provides getwd only for backwards compatibility with BSD. The buffer argument should be a pointer to an array at least PATH_ MAX bytes long (see Section 28.6 [Limits on File System Capacity], page 545). In the GNU system there is no limit to the size of a file name, so this is not necessarily enough space to contain the directory name. That is why this function is deprecated.

Function int chdir (const char *filename)

This function is used to set the process’s working directory to filename. The normal, successful return value from chdir is 0. A value of -1 is returned to indicate an error. The errno error conditions defined for this function are the usual file name syntax errors (see Section 6.2.3 [File Name Errors], page 89), plus ENOTDIR if the file filename is not a directory.

9.2 Accessing Directories

The facilities described in this section let you read the contents of a directory file. This is useful if you want your program to list all the files in a directory, perhaps as part of a menu.

The opendir function opens a directory stream whose elements are directory entries. You use the readdir function on the directory stream to retrieve these entries, represented as struct dirent objects. The name of the file for each entry is stored in the d_name member of this structure. There are obvious parallels here to the stream facilities for ordinary files, described in Chapter 7 [Input/Output on Streams], page 93.

9.2.1 Format of a Directory Entry

This section describes what you find in a single directory entry, as you might obtain it from a directory stream. All the symbols are declared in the header file ‘dirent.h’.

Data Type struct dirent

This is a structure type used to return information about directory entries. It contains the following fields:

char d_name[]

This is the null-terminated file name component. This is the only field you can count on in all POSIX systems.

ino_t d_fileno

This is the file serial number. For BSD compatibility, you can also refer to this member as d_ino. In the GNU system and most POSIX systems, for most files this the same as the st_ino member that stat will return for the file. See Section 9.8 [File Attributes], page 195.

unsigned char d_namlen

This is the length of the file name, not including the terminating null character. Its type is unsigned char because that is the integer type of the appropriate size

unsigned char d_type

This is the type of the file, possibly unknown. The following constants are defined for its value:

DT_UNKNOWN

The type is unknown. On some systems this is the only value returned.

DT_REG A regular file.

DT_DIR A directory.

DT_FIFO A named pipe, or FIFO. See Section 10.3 [FIFO Special Files], page 219.

DT_SOCK A local-domain socket.

DT_CHR A character device.

DT_BLK A block device.

This member is a BSD extension. Each value except DT UNKNOWN corresponds to the file type bits in the st_mode member of struct statbuf. These two macros convert between d_type values and st_mode values:

Function int IFTODT (mode_t mode)

This returns the d_type value corresponding to mode.

Function mode_t DTTOIF (int dirtype)

This returns the st_mode value corresponding to dirtype.

This structure may contain additional members in the future. When a file has multiple names, each name has its own directory entry. The only way you can tell that the directory entries belong to a single file is that they have the same value for the d_fileno field. File attributes such as size, modification times, and the like are part of the file itself, not any particular directory entry. See Section 9.8 [File Attributes], page 195.

9.2.2 Opening a Directory Stream

This section describes how to open a directory stream. All the symbols are declared in the header file ‘dirent.h’.

Data Type DIR

The DIR data type represents a directory stream.

You shouldn’t ever allocate objects of the struct dirent or DIR data types, since the directory access functions do that for you. Instead, you refer to these objects using the pointers returned by the following functions.

Function DIR * opendir (const char *dirname)

The opendir function opens and returns a directory stream for reading the directory whose file name is dirname. The stream has type DIR *. If unsuccessful, opendir returns a null pointer. In addition to the usual file name errors (see Section 6.2.3 [File Name Errors], page 89), the following errno error conditions are defined for this function:

EACCES Read permission is denied for the directory named by

          dirname.

EMFILE The process has too many files open.

ENFILE The entire system, or perhaps the file system which contains the directory, cannot support any additional open files at the moment. (This problem cannot happen on the GNU system.)

The DIR type is typically implemented using a file descriptor, and the opendir function in terms of the open function. See Chapter 8 [Low-Level Input/Output], page 151. Directory streams and the underlying file descriptors are closed on exec (see Section 23.5 [Executing a File], page 470).

9.2.3 Reading and Closing a Directory Stream

This section describes how to read directory entries from a directory stream, and how to close the stream when you are done with it. All the symbols are declared in the header file ‘dirent.h’.

Function struct dirent * readdir (DIR *dirstream)

This function reads the next entry from the directory. It normally returns a pointer to a structure containing information about the file. This structure is statically allocated and can be rewritten by a subsequent call. Portability Note: On some systems, readdir may not return entries for ‘.’ and ‘..’, even though these are always valid file names in any directory. See Section 6.2.2 [File Name Resolution], page 88. If there are no more entries in the directory or an error is detected, readdir returns a null pointer. The following errno error conditions are defined for this function:

EBADF The dirstream argument is not valid.

readdir is not thread safe. Multiple threads using readdir on the same dirstream may overwrite the return value. Use readdir_r when this is critical.

Function int readdir r (DIR *dirstream, struct dirent *entry,

      struct dirent **result)

This function is the reentrant version of readdir. Like readdir it returns the next entry from the directory. But to prevent conflicts for simultaneously running threads the result is not stored in some internal memory. Instead the argument entry has to point to a place where the result is stored.

The return value is 0 in case the next entry was read successfully. In this case a pointer to the result is returned in *result. It is not required that *result is the same as entry. If something goes wrong while executing readdir_r the function returns -1. The errno variable is set like described for readdir. Portability Note: On some systems, readdir_r may not return a terminated string as the file name even if no d_reclen element is available in struct dirent and the file name as the maximal allowed size. Modern systems all have the d_reclen field and on old systems multi threading is not critical. In any case, there is no such problem with the readdir function so that even on systems without d_reclen field one could use multiple threads by using external locking.

Function int closedir (DIR *dirstream)

This function closes the directory stream dirstream. It returns 0 on success and -1 on failure. The following errno error conditions are defined for this function:

EBADF The dirstream argument is not valid.

9.2.4 Simple Program to List a Directory

Here’s a simple program that prints the names of the files in the current working directory:

  #include <stddef.h>
  #include <stdio.h>
  #include <sys/types.h>
  #include <dirent.h>
  int
  main (void)
  {
    DIR *dp;
    struct dirent *ep;
    dp = opendir ("./");
    if (dp != NULL)
      {
        while (ep = readdir (dp))
          puts (ep->d_name);
        (void) closedir (dp);
      }
    else
      puts ("Couldn't open the directory.");
    return 0;

  }

The order in which files appear in a directory tends to be fairly random. A more useful program would sort the entries (perhaps by alphabetizing them) before printing them; see Section 9.2.6 [Scanning the Content of a Directory], page 187 and Section 15.3 [Array Sort Function], page 318.

9.2.5 Random Access in a Directory Stream

This section describes how to reread parts of a directory that you have already read from an open directory stream. All the symbols are declared in the header file ‘dirent.h’.

Function void rewinddir (DIR *dirstream)

The rewinddir function is used to reinitialize the directory stream dirstream, so that if you call readdir it returns information about the first entry in the directory again. This function also notices if files have been added or removed to the directory since it was opened with opendir. (Entries for these files might or might not be returned by readdir if they were added or removed since you last called opendir

or rewinddir.)

Function off_t telldir (DIR *dirstream)

The telldir function returns the file position of the directory stream dirstream. You can use this value with seekdir to restore the directory stream to that position.

Function void seekdir (DIR *dirstream, off_t pos)

The seekdir function sets the file position of the directory stream dirstream to pos. The value pos must be the result of a previous call to telldir on this particular stream; closing and reopening the directory can invalidate values returned by telldir.

9.2.6 Scanning the Content of a Directory

A higher-level interface to the directory handling functions is the scandir function. With its help one can select a subset of the entries in a directory, possibly sort them and get as the result a list of names.

Function int scandir (const char *dir, struct dirent
      ***namelist, int (*selector) (struct dirent *), int
      (*cmp) (const void *, const void *))

The scandir function scans the contents of the directory selected by dir. The result in namelist is an array of pointers to structure of type struct dirent which describe all selected directory entries and which is allocated using malloc. Instead of always getting all directory entries returned, the user supplied function selector can be used to decide which entries are in the result. Only the entries for which selector returns a nonzero value are selected. Finally the entries in the namelist are sorted using the user supplied function cmp. The arguments of the cmp function are of type struct dirent **. I.e., one cannot directly use the strcmp or strcoll function; see the function alphasort below. The return value of the function gives the number of entries placed in namelist. If it is -1 an error occurred and the global variable errno contains more information on the error.

As said above the fourth argument to the scandir function must be a pointer to a sorting function. For the convenience of the programmer the GNU C library contains an implementation of a function which is very helpful for this purpose.

Function int alphasort (const void *a, const void *b)

The alphasort function behaves like the strcmp function (see Section 5.5 [String/Array Comparison], page 72). The difference is that the arguments are not string pointers but instead they are of type

struct dirent **.

Return value of is less than, equal to, or greater than zero depending on the order of the two entries a and b.

9.2.7 Simple Program to List a Directory, Mark II

Here is a revised version of the directory lister found above (see Section 9.2.4 [Simple Program to List a Directory], page 186). Using the scandir function we can avoid using the functions which directly work with the directory contents. After the call the found entries are available for direct used.

  #include <stdio.h>
  #include <dirent.h>
  static int
  one (struct dirent *unused)
  {
    return 1;
  }
  int
  main (void)
  {
    struct dirent **eps;
    int n;

       n = scandir ("./", &eps, one, alphasort);
       if (n >= 0)
         {
           int cnt;
           for (cnt = 0; cnt < n; ++cnt)
             puts (eps[cnt]->d_name);
         }
       else
         perror ("Couldn't open the directory");
       return 0;
     }

Please note the simple selector function for this example. Since we want to see all directory entries we always return 1.

9.3 Hard Links

In POSIX systems, one file can have many names at the same time. All of the names are equally real, and no one of them is preferred to the others.

To add a name to a file, use the link function. (The new name is also called a hard link to the file.) Creating a new link to a file does not copy the contents of the file; it simply makes a new name by which the file can be known, in addition to the file’s existing name or names.

One file can have names in several directories, so the the organization of the file system is not a strict hierarchy or tree.

In most implementations, it is not possible to have hard links to the same file in multiple file systems. link reports an error if you try to make a hard link to the file from another file system when this cannot be done.

The prototype for the link function is declared in the header file

‘unistd.h’.

Function int link (const char *oldname, const char *newname)

The link function makes a new link to the existing file named by oldname, under the new name newname. This function returns a value of 0 if it is successful and -1 on failure. In addition to the usual file name errors (see Section 6.2.3 [File Name Errors], page 89) for both oldname and newname, the following errno error conditions are defined for this function:

EACCES You are not allowed to write the directory in which the new link is to be written.

EEXIST There is already a file named newname. If you want to replace this link with a new link, you must remove the old link explicitly first.

EMLINK There are already too many links to the file named by oldname. (The maximum number of links to a file is LINK_MAX; see Section 28.6 [Limits on File System Capacity], page 545.)

ENOENT The file named by oldname doesn’t exist. You can’t make a link to a file that doesn’t exist.

ENOSPC The directory or file system that would contain the new link is full and cannot be extended.

EPERM In the GNU system and some others, you cannot make links to directories. Many systems allow only privileged users to do so. This error is used to report the problem.

EROFS The directory containing the new link can’t be modified because it’s on a read-only file system.

EXDEV The directory specified in newname is on a different file system than the existing file.

EIO A hardware error occurred while trying to read or write the to filesystem.

9.4 Symbolic Links

The GNU system supports soft links or symbolic links. This is a kind of “file” that is essentially a pointer to another file name. Unlike hard links, symbolic links can be made to directories or across file systems with no restrictions. You can also make a symbolic link to a name which is not the name of any file. (Opening this link will fail until a file by that name is created.) Likewise, if the symbolic link points to an existing file which is later deleted, the symbolic link continues to point to the same file name even though the name no longer names any file.

The reason symbolic links work the way they do is that special things happen when you try to open the link. The open function realizes you have specified the name of a link, reads the file name contained in the link, and opens that file name instead. The stat function likewise operates on the file that the symbolic link points to, instead of on the link itself.

By contrast, other operations such as deleting or renaming the file operate on the link itself. The functions readlink and lstat also refrain from following symbolic links, because their purpose is to obtain information about the link. So does link, the function that makes a hard link—it makes a hard link to the symbolic link, which one rarely wants.

Prototypes for the functions listed in this section are in ‘unistd.h’.

Function int symlink (const char *oldname, const char

*newname) The symlink function makes a symbolic link to oldname named newname. The normal return value from symlink is 0. A return value of -1 indicates an error. In addition to the usual file name syntax errors (see Section 6.2.3 [File Name Errors], page 89), the following errno error conditions are defined for this function:

EEXIST There is already an existing file named newname.

EROFS The file newname would exist on a read-only file system.

ENOSPC The directory or file system cannot be extended to make the new link.

EIO A hardware error occurred while reading or writing data on the disk.

Function int readlink (const char *filename, char *buffer,

      size_t size)

The readlink function gets the value of the symbolic link filename. The file name that the link points to is copied into buffer. This file name string is not null-terminated; readlink normally returns the number of characters copied. The size argument specifies the maximum number of characters to copy, usually the allocation size of buffer. If the return value equals size, you cannot tell whether or not there was room to return the entire name. So make a bigger buffer and call readlink again. Here is an example:

     char *
     readlink_malloc (char *filename)
     {
       int size = 100;
       while (1)
         {
           char *buffer = (char *) xmalloc (size);
           int nchars = readlink (filename, buffer, size);
           if (nchars < size)
             return buffer;
           free (buffer);
           size *= 2;
         }
     }

A value of -1 is returned in case of error. In addition to the usual file name errors (see Section 6.2.3 [File Name Errors], page 89), the following errno error conditions are defined for this function:

EINVAL The named file is not a symbolic link.

EIO A hardware error occurred while reading or writing data on the disk.

9.5 Deleting Files

You can delete a file with the functions unlink or remove. Deletion actually deletes a file name. If this is the file’s only name, then

the file is deleted as well. If the file has other names as well (see Section 9.3 [Hard Links], page 189), it remains accessible under its other names.

Function int unlink (const char *filename)

The unlink function deletes the file name filename. If this is a file’s sole name, the file itself is also deleted. (Actually, if any process has the file open when this happens, deletion is postponed until all processes have closed the file.) The function unlink is declared in the header file ‘unistd.h’. This function returns 0 on successful completion, and -1 on error. In addition to the usual file name errors (see Section 6.2.3 [File Name Errors], page 89), the following errno error conditions are defined for this function:

EACCES Write permission is denied for the directory from which the file is to be removed, or the directory has the sticky bit set and you do not own the file.

EBUSY This error indicates that the file is being used by the system in such a way that it can’t be unlinked. For example, you might see this error if the file name specifies the root directory or a mount point for a file system.

ENOENT The file name to be deleted doesn’t exist.

EPERM On some systems, unlink cannot be used to delete the name of a directory, or can only be used this way by a privileged user. To avoid such problems, use rmdir to delete directories. (In the GNU system unlink can never delete the name of a directory.)

EROFS The directory in which the file name is to be deleted is on a read-only file system, and can’t be modified.

Function int rmdir (const char *filename)

The rmdir function deletes a directory. The directory must be empty before it can be removed; in other words, it can only contain entries for ‘.’ and ‘..’. In most other respects, rmdir behaves like unlink. There are two additional errno error conditions defined for rmdir:

ENOTEMPTY

EEXIST The directory to be deleted is not empty.

These two error codes are synonymous; some systems use one, and some use the other. The GNU system always uses ENOTEMPTY. The prototype for this function is declared in the header file

‘unistd.h’.

Function int remove (const char *filename)

This is the ISO C function to remove a file. It works like unlink for files and like rmdir for directories. remove is declared in ‘stdio.h’.

9.6 Renaming Files

The rename function is used to change a file’s name.

Function int rename (const char *oldname, const char

*newname) The rename function renames the file name oldname with newname. The file formerly accessible under the name oldname is afterward accessible as newname instead. (If the file had any other names aside from oldname, it continues to have those names.) The directory containing the name newname must be on the same file system as the file (as indicated by the name oldname). One special case for rename is when oldname and newname are two names for the same file. The consistent way to handle this case is to delete oldname. However, POSIX requires that in this case rename do nothing and report success—which is inconsistent. We don’t know what your operating system will do. If the oldname is not a directory, then any existing file named newname is removed during the renaming operation. However, if newname is the name of a directory, rename fails in this case. If the oldname is a directory, then either newname must not exist or it must name a directory that is empty. In the latter case, the existing directory named newname is deleted first. The name newname must not specify a subdirectory of the directory oldname which is being renamed.

One useful feature of rename is that the meaning of the name newname changes “atomically” from any previously existing file by that name to its new meaning (the file that was called oldname). There is no instant at which newname is nonexistent “in between” the old meaning and the new meaning. If there is a system crash during the operation, it is possible for both names to still exist; but newname will always be intact if it exists at all. If rename fails, it returns -1. In addition to the usual file name errors (see Section 6.2.3 [File Name Errors], page 89), the following errno error conditions are defined for this function:

EACCES One of the directories containing newname or oldname refuses write permission; or newname and oldname are directories and write permission is refused for one of them.

EBUSY A directory named by oldname or newname is being used by the system in a way that prevents the renaming from working. This includes directories that are mount points for filesystems, and directories that are the current working directories of processes.

ENOTEMPTY

EEXIST The directory newname isn’t empty. The GNU system always returns ENOTEMPTY for this, but some other systems return EEXIST.

EINVAL The oldname is a directory that contains newname.

EISDIR The newname names a directory, but the oldname doesn’t.

EMLINK The parent directory of newname would have too many links.

ENOENT The file named by oldname doesn’t exist.

ENOSPC The directory that would contain newname has no room for another entry, and there is no space left in the file system to expand it.

EROFS The operation would involve writing to a directory on a read-only file system.

EXDEV The two file names newname and oldnames are on different file systems.

9.7 Creating Directories

Directories are created with the mkdir function. (There is also a shell command mkdir which does the same thing.)

Function int mkdir (const char *filename, mode_t mode)

The mkdir function creates a new, empty directory whose name is filename. The argument mode specifies the file permissions for the new directory file. See Section 9.8.5 [The Mode Bits for Access Permission], page 202, for more information about this. A return value of 0 indicates successful completion, and -1 indicates failure. In addition to the usual file name syntax errors (see Section 6.2.3 [File Name Errors], page 89), the following errno error conditions are defined for this function:

EACCES Write permission is denied for the parent directory in which the new directory is to be added.

EEXIST A file named filename already exists.

EMLINK The parent directory has too many links.

Well-designed file systems never report this error, because they permit more links than your disk could possibly hold. However, you must still take account of the possibility of this error, as it could result from network access to a file system on another machine.

ENOSPC The file system doesn’t have enough room to create the new directory.

EROFS The parent directory of the directory being created is on a read-only file system, and cannot be modified.

To use this function, your program should include the header file

‘sys/stat.h’.

9.8 File Attributes

When you issue an ‘ls -l’ shell command on a file, it gives you information about the size of the file, who owns it, when it was last modified, and the like. This kind of information is called the file attributes; it is associated with the file itself and not a particular one of its names.

This section contains information about how you can inquire about and modify these attributes of files.

9.8.1 What the File Attribute Values Mean

When you read the attributes of a file, they come back in a structure called struct stat. This section describes the names of the attributes, their data types, and what they mean. For the functions to read the attributes of a file, see Section 9.8.2 [Reading the Attributes of a File], page 198.

The header file ‘sys/stat.h’ declares all the symbols defined in this section.

Data Type struct stat

The stat structure type is used to return information about the attributes of a file. It contains at least the following members:

mode_t st_mode

Specifies the mode of the file. This includes file type information (see Section 9.8.3 [Testing the Type of a File], page 199) and the file permission bits (see Section 9.8.5 [The Mode Bits for Access Permission], page 202).

ino_t st_ino

The file serial number, which distinguishes this file from all other files on the same device.

dev_t st_dev

Identifies the device containing the file. The st_ino and st_dev, taken together, uniquely identify the file. The st_dev value is not necessarily consistent across reboots or system crashes, however.

nlink_t st_nlink

The number of hard links to the file. This count keeps track of how many directories have entries for this file. If the count is ever decremented to zero, then the file itself is discarded as soon as no process still holds it open. Symbolic links are not counted in the total.

uid_t st_uid

The user ID of the file’s owner. See Section 9.8.4 [File Owner], page 200.

gid_t st_gid

The group ID of the file. See Section 9.8.4 [File Owner], page 200.

off_t st_size

This specifies the size of a regular file in bytes. For files that are really devices and the like, this field isn’t usually meaningful. For symbolic links, this specifies the length of the file name the link refers to.

time_t st_atime

This is the last access time for the file. See Section 9.8.9 [File Times], page 208.

unsigned long int st_atime_usec

This is the fractional part of the last access time for the file. See Section 9.8.9 [File Times], page 208.

time_t st_mtime

This is the time of the last modification to the contents of the file. See Section 9.8.9 [File Times], page 208.

unsigned long int st_mtime_usec

This is the fractional part of the time of last modification to the contents of the file. See Section 9.8.9 [File Times], page 208.

time_t st_ctime

This is the time of the last modification to the attributes of the file. See Section 9.8.9 [File Times], page 208.

unsigned long int st_ctime_usec

This is the fractional part of the time of last modification to the attributes of the file. See Section 9.8.9 [File Times], page 208.

unsigned int st_blocks

This is the amount of disk space that the file occupies, measured in units of 512-byte blocks. The number of disk blocks is not strictly proportional to the size of the file, for two reasons: the file system may use some blocks for internal record keeping; and the file may be sparse—it may have “holes” which contain zeros but do not actually take up space on the disk. You can tell (approximately) whether a file is sparse by comparing this value with st_size, like this:

     (st.st_blocks * 512 < st.st_size)

This test is not perfect because a file that is just slightly sparse might not be detected as sparse at all. For practical applications, this is not a problem.

unsigned int st_blksize

The optimal block size for reading of writing this file, in bytes. You might use this size for allocating the buffer space for reading of writing the file. (This is unrelated to

st_blocks.)

Some of the file attributes have special data type names which exist specifically for those attributes. (They are all aliases for well-known integer types that you know and love.) These typedef names are defined in the header file ‘sys/types.h’ as well as in ‘sys/stat.h’. Here is a list of them.

Data Type mode t

This is an integer data type used to represent file modes. In the GNU system, this is equivalent to unsigned int.

Data Type ino t

This is an arithmetic data type used to represent file serial numbers. (In Unix jargon, these are sometimes called inode numbers.) In the GNU system, this type is equivalent to unsigned long int.

Data Type dev t

This is an arithmetic data type used to represent file device numbers. In the GNU system, this is equivalent to int.

Data Type nlink t

This is an arithmetic data type used to represent file link counts. In the GNU system, this is equivalent to unsigned short int.

9.8.2 Reading the Attributes of a File

To examine the attributes of files, use the functions stat, fstat and lstat. They return the attribute information in a struct stat object. All three functions are declared in the header file ‘sys/stat.h’.

Function int stat (const char *filename, struct stat *buf )

The stat function returns information about the attributes of the file named by filename in the structure pointed at by buf. If filename is the name of a symbolic link, the attributes you get describe the file that the link points to. If the link points to a nonexistent file name, then stat fails, reporting a nonexistent file. The return value is 0 if the operation is successful, and -1 on failure. In addition to the usual file name errors (see Section 6.2.3 [File Name Errors], page 89, the following errno error conditions are defined for this function:

ENOENT The file named by filename doesn’t exist.

Function int fstat (int filedes, struct stat *buf )

The fstat function is like stat, except that it takes an open file descriptor as an argument instead of a file name. See Chapter 8 [Low-Level Input/Output], page 151. Like stat, fstat returns 0 on success and -1 on failure. The following errno error conditions are defined for fstat:

EBADF The filedes argument is not a valid file descriptor.

Function int lstat (const char *filename, struct stat *buf )

The lstat function is like stat, except that it does not follow symbolic links. If filename is the name of a symbolic link, lstat returns information about the link itself; otherwise, lstat works like stat. See Section 9.4 [Symbolic Links], page 190.

9.8.3 Testing the Type of a File

The file mode, stored in the st_mode field of the file attributes, contains two kinds of information: the file type code, and the access permission bits. This section discusses only the type code, which you can use to tell whether the file is a directory, whether it is a socket, and so on. For information about the access permission, Section 9.8.5 [The Mode Bits for Access Permission], page 202.

There are two predefined ways you can access the file type portion of the file mode. First of all, for each type of file, there is a predicate macro which examines a file mode value and returns true or false—is the file of that type, or not. Secondly, you can mask out the rest of the file mode to get just a file type code. You can compare this against various constants for the supported file types.

All of the symbols listed in this section are defined in the header file

‘sys/stat.h’.

The following predicate macros test the type of a file, given the value m which is the st_mode field returned by stat on that file:

Macro int S ISDIR (mode_t m)

This macro returns nonzero if the file is a directory.

Macro int S ISCHR (mode_t m)

This macro returns nonzero if the file is a character special file (a device like a terminal).

Macro int S ISBLK (mode_t m)

This macro returns nonzero if the file is a block special file (a device like a disk).

Macro int S ISREG (mode_t m)

This macro returns nonzero if the file is a regular file.

Macro int S ISFIFO (mode_t m)

This macro returns nonzero if the file is a FIFO special file, or a pipe. See Chapter 10 [Pipes and FIFOs], page 215.

Macro int S ISLNK (mode_t m)

This macro returns nonzero if the file is a symbolic link. See Section 9.4 [Symbolic Links], page 190.

Macro int S ISSOCK (mode_t m)

This macro returns nonzero if the file is a socket. See Chapter 11 [Sockets], page 221.

An alterate non-POSIX method of testing the file type is supported for compatibility with BSD. The mode can be bitwise ANDed with S_IFMT to extract the file type code, and compared to the appropriate type code constant. For example,

     S_ISCHR (mode)

is equivalent to:

     ((mode & S_IFMT) == S_IFCHR)

Macro int S IFMT

This is a bit mask used to extract the file type code portion of a mode value.

These are the symbolic names for the different file type codes:

S_IFDIR This macro represents the value of the file type code for a directory file.

S_IFCHR This macro represents the value of the file type code for a character-oriented device file.

S_IFBLK This macro represents the value of the file type code for a blockoriented device file.

S_IFREG This macro represents the value of the file type code for a regular file.

S_IFLNK This macro represents the value of the file type code for a symbolic link.

S_IFSOCK This macro represents the value of the file type code for a socket.

S_IFIFO This macro represents the value of the file type code for a FIFO or pipe.

9.8.4 File Owner

Every file has an owner which is one of the registered user names defined on the system. Each file also has a group, which is one of the defined groups. The file owner can often be useful for showing you who edited the file (especially when you edit with GNU Emacs), but its main purpose is for access control.

The file owner and group play a role in determining access because the file has one set of access permission bits for the user that is the owner, another set that apply to users who belong to the file’s group, and a third set of bits that apply to everyone else. See Section 9.8.6 [How Your Access to a File is Decided], page 204, for the details of how access is decided based on this data.

When a file is created, its owner is set from the effective user ID of the process that creates it (see Section 26.2 [The Persona of a Process], page 511). The file’s group ID may be set from either effective group ID of the process, or the group ID of the directory that contains the file, depending on the system where the file is stored. When you access a remote file system, it behaves according to its own rule, not according to the system your program is running on. Thus, your program must be prepared to encounter either kind of behavior, no matter what kind of system you run it on.

You can change the owner and/or group owner of an existing file using the chown function. This is the primitive for the chown and chgrp shell commands.

The prototype for this function is declared in ‘unistd.h’.

Function int chown (const char *filename, uid_t owner, gid_t

group) The chown function changes the owner of the file filename to owner, and its group owner to group. Changing the owner of the file on certain systems clears the set-user-ID and set-group-ID bits of the file’s permissions. (This is because those bits may not be appropriate for the new owner.) The other file permission bits are not changed. The return value is 0 on success and -1 on failure. In addition to the usual file name errors (see Section 6.2.3 [File Name Errors], page 89), the following errno error conditions are defined for this function: EPERM This process lacks permission to make the requested

change.

Only privileged users or the file’s owner can change the file’s group. On most file systems, only privileged users can change the file owner; some file systems allow you to change the owner if you are currently the owner. When you access a remote file system, the behavior you encounter is determined by the system that actually holds the file, not by the system your program is running on. See Section 28.7 [Optional Features in File Support], page 546, for information about the _POSIX_CHOWN_

          RESTRICTED macro.

EROFS The file is on a read-only file system.

Function int fchown (int filedes, int owner, int group)

This is like chown, except that it changes the owner of the file with open file descriptor filedes. The return value from fchown is 0 on success and -1 on failure. The following errno error codes are defined for this function: EBADF The filedes argument is not a valid file descriptor. EINVAL The filedes argument corresponds to a pipe or socket, not

an ordinary file.

EPERM This process lacks permission to make the requested change. For details, see chmod, above.

EROFS The file resides on a read-only file system.

9.8.5 The Mode Bits for Access Permission

The file mode, stored in the st_mode field of the file attributes, contains two kinds of information: the file type code, and the access permission bits. This section discusses only the access permission bits, which control who can read or write the file. See Section 9.8.3 [Testing the Type of a File], page 199, for information about the file type code.

All of the symbols listed in this section are defined in the header file

‘sys/stat.h’.

These symbolic constants are defined for the file mode bits that control access permission for the file:

S_IRUSR

S_IREAD Read permission bit for the owner of the file. On many systems, this bit is 0400. S_IREAD is an obsolete synonym provided for BSD compatibility.

S_IWUSR

S_IWRITE Write permission bit for the owner of the file. Usually 0200. S_IWRITE is an obsolete synonym provided for BSD compatibility.

S_IXUSR

S_IEXEC Execute (for ordinary files) or search (for directories) permission bit for the owner of the file. Usually 0100. S_IEXEC is an obsolete synonym provided for BSD compatibility.

S_IRWXU This is equivalent to ‘(S_IRUSR | S_IWUSR | S_IXUSR)’.

S_IRGRP Read permission bit for the group owner of the file. Usually 040.

S_IWGRP Write permission bit for the group owner of the file. Usually 020.

S_IXGRP Execute or search permission bit for the group owner of the file. Usually 010.

S_IRWXG This is equivalent to ‘(S_IRGRP | S_IWGRP | S_IXGRP)’.

S_IROTH Read permission bit for other users. Usually 04.

S_IWOTH Write permission bit for other users. Usually 02.

S_IXOTH Execute or search permission bit for other users. Usually 01.

S_IRWXO This is equivalent to ‘(S_IROTH | S_IWOTH | S_IXOTH)’.

S_ISUID This is the set-user-ID on execute bit, usually 04000. See Section 26.4 [How an Application Can Change Persona], page 513.

S_ISGID This is the set-group-ID on execute bit, usually 02000. See Section 26.4 [How an Application Can Change Persona], page 513.

S_ISVTX This is the sticky bit, usually 01000.

On a directory, it gives permission to delete a file in the directory only if you own that file. Ordinarily, a user either can delete all the files in the directory or cannot delete any of them (based on whether the user has write permission for the directory). The same restriction applies—you must both have write permission for the directory and own the file you want to delete. The one exception is that the owner of the directory can delete any file in the directory, no matter who owns it (provided the owner has given himself write permission for the directory). This is commonly used for the ‘/tmp’ directory, where anyone may create files, but not delete files created by other users.

Originally the sticky bit on an executable file modified the swapping policies of the system. Normally, when a program terminated, its pages in core were immediately freed and reused. If the sticky bit was set on the executable file, the system kept the pages in core for a while as if the program were still running. This was advantageous for a program likely to be run many times in succession. This usage is obsolete in modern systems. When a program terminates, its pages always remain in core as long as there is no shortage of memory in the system. When the program is next run, its pages will still be in core if no shortage arose since the last run.

On some modern systems where the sticky bit has no useful meaning for an executable file, you cannot set the bit at all for a non-directory. If you try, chmod fails with EFTYPE; see Section 9.8.7 [Assigning File Permissions], page 204.

Some systems (particularly SunOS) have yet another use for the sticky bit. If the sticky bit is set on a file that is not executable, it means the opposite: never cache the pages of this file at all. The main use of this is for the files on an NFS server machine which are used as the swap area of diskless client machines. The idea is that the pages of the file will be cached in the client’s memory, so it is a waste of the server’s memory to cache them a second time. In this use the sticky bit also says that the filesystem may fail to record the file’s modification time onto disk reliably (the idea being that noone cares for a swap file).

The actual bit values of the symbols are listed in the table above so you can decode file mode values when debugging your programs. These bit values are correct for most systems, but they are not guaranteed.

Warning: Writing explicit numbers for file permissions is bad practice. It is not only nonportable, it also requires everyone who reads your program to remember what the bits mean. To make your program clean, use the symbolic names.

9.8.6 How Your Access to a File is Decided

Recall that the operating system normally decides access permission for a file based on the effective user and group IDs of the process, and its supplementary group IDs, together with the file’s owner, group and permission bits. These concepts are discussed in detail in Section 26.2 [The Persona of a Process], page 511.

If the effective user ID of the process matches the owner user ID of the file, then permissions for read, write, and execute/search are controlled by the corresponding “user” (or “owner”) bits. Likewise, if any of the effective group ID or supplementary group IDs of the process matches the group owner ID of the file, then permissions are controlled by the “group” bits. Otherwise, permissions are controlled by the “other” bits.

Privileged users, like ‘root’, can access any file, regardless of its file permission bits. As a special case, for a file to be executable even for a privileged user, at least one of its execute bits must be set.

9.8.7 Assigning File Permissions

The primitive functions for creating files (for example, open or mkdir) take a mode argument, which specifies the file permissions for the newly created file. But the specified mode is modified by the process’s file creation mask, or umask, before it is used.

The bits that are set in the file creation mask identify permissions that are always to be disabled for newly created files. For example, if you set all the “other” access bits in the mask, then newly created files are not accessible at all to processes in the “other” category, even if the mode argument specified to the creation function would permit such access. In other words, the file creation mask is the complement of the ordinary access permissions you want to grant.

Programs that create files typically specify a mode argument that includes all the permissions that make sense for the particular file. For an ordinary file, this is typically read and write permission for all classes of users. These permissions are then restricted as specified by the individual user’s own file creation mask.

To change the permission of an existing file given its name, call chmod. This function ignores the file creation mask; it uses exactly the specified permission bits.

In normal use, the file creation mask is initialized in the user’s login shell (using the umask shell command), and inherited by all subprocesses. Application programs normally don’t need to worry about the file creation mask. It will do automatically what it is supposed to do.

When your program should create a file and bypass the umask for its access permissions, the easiest way to do this is to use fchmod after opening the file, rather than changing the umask.

In fact, changing the umask is usually done only by shells. They use the umask function.

The functions in this section are declared in ‘sys/stat.h’.

Function mode_t umask (mode_t mask)

The umask function sets the file creation mask of the current process to mask, and returns the previous value of the file creation mask. Here is an example showing how to read the mask with umask without changing it permanently:

     mode_t
     read_umask (void)
     {
       mask = umask (0);
       umask (mask);
     }

However, it is better to use getumask if you just want to read the mask value, because that is reentrant (at least if you use the GNU operating system).

Function mode_t getumask (void)

Return the current value of the file creation mask for the current process. This function is a GNU extension.

Function int chmod (const char *filename, mode_t mode)

The chmod function sets the access permission bits for the file named by filename to mode. If the filename names a symbolic link, chmod changes the permission of the file pointed to by the link, not those of the link itself. This function returns 0 if successful and -1 if not. In addition to the usual file name errors (see Section 6.2.3 [File Name Errors], page 89), the following errno error conditions are defined for this function: ENOENT The named file doesn’t exist. EPERM This process does not have permission to change the access

permission of this file. Only the file’s owner (as judged by the effective user ID of the process) or a privileged user can change them.

EROFS The file resides on a read-only file system.

EFTYPE mode has the S_ISVTX bit (the “sticky bit”) set, and the named file is not a directory. Some systems do not allow setting the sticky bit on non-directory files, and some do (and only some of those assign a useful meaning to the bit for non-directory files). You only get EFTYPE on systems where the sticky bit has no useful meaning for non-directory files, so it is always safe to just clear the bit in mode and call chmod again. See Section 9.8.5 [The Mode Bits for Access Permission], page 202, for full details on the sticky bit.

Function int fchmod (int filedes, int mode)

This is like chmod, except that it changes the permissions of the file currently open via descriptor filedes. The return value from fchmod is 0 on success and -1 on failure. The following errno error codes are defined for this function:

EBADF The filedes argument is not a valid file descriptor.

EINVAL The filedes argument corresponds to a pipe or socket, or something else that doesn’t really have access permissions.

EPERM This process does not have permission to change the access permission of this file. Only the file’s owner (as judged by the effective user ID of the process) or a privileged user can change them.

EROFS The file resides on a read-only file system.

9.8.8 Testing Permission to Access a File

When a program runs as a privileged user, this permits it to access files off-limits to ordinary users—for example, to modify ‘/etc/passwd’. Programs designed to be run by ordinary users but access such files use the setuid bit feature so that they always run with root as the effective user ID.

Such a program may also access files specified by the user, files which conceptually are being accessed explicitly by the user. Since the program runs as root, it has permission to access whatever file the user specifies— but usually the desired behavior is to permit only those files which the user could ordinarily access.

The program therefore must explicitly check whether the user would have the necessary access to a file, before it reads or writes the file.

To do this, use the function access, which checks for access permission based on the process’s real user ID rather than the effective user ID. (The setuid feature does not alter the real user ID, so it reflects the user who actually ran the program.)

There is another way you could check this access, which is easy to describe, but very hard to use. This is to examine the file mode bits and mimic the system’s own access computation. This method is undesirable because many systems have additional access control features; your program cannot portably mimic them, and you would not want to try to keep track of the diverse features that different systems have. Using access is simple and automatically does whatever is appropriate for the system you are using.

access is only only appropriate to use in setuid programs. A non-setuid program will always use the effective ID rather than the real ID.

The symbols in this section are declared in ‘unistd.h’.

Function int access (const char *filename, int how)

The access function checks to see whether the file named by filename can be accessed in the way specified by the how argument. The how argument either can be the bitwise OR of the flags R_OK, W_OK, X_OK, or the existence test F_OK. This function uses the real user and group ID’s of the calling process, rather than the effective ID’s, to check for access permission. As a result, if you use the function from a setuid or setgid program (see Section 26.4 [How an Application Can Change Persona], page 513), it gives information relative to the user who actually ran the program. The return value is 0 if the access is permitted, and -1 otherwise. (In other words, treated as a predicate function, access returns true if the requested access is denied.) In addition to the usual file name errors (see Section 6.2.3 [File Name Errors], page 89), the following errno error conditions are defined for this function:

EACCES The access specified by how is denied.

ENOENT The file doesn’t exist.

EROFS Write permission was requested for a file on a read-only file system.

These macros are defined in the header file ‘unistd.h’ for use as the how argument to the access function. The values are integer constants.

Macro int R OK

Argument that means, test for read permission.

Macro int W OK

Argument that means, test for write permission.

Macro int X OK

Argument that means, test for execute/search permission.

Macro int F OK

Argument that means, test for existence of the file.

9.8.9 File Times

Each file has three timestamps associated with it: its access time, its modification time, and its attribute modification time. These correspond to the st_atime, st_mtime, and st_ctime members of the stat structure; see Section 9.8 [File Attributes], page 195.

All of these times are represented in calendar time format, as time_t objects. This data type is defined in ‘time.h’. For more information about representation and manipulation of time values, see Section 17.2 [Calendar Time], page 341.

Reading from a file updates its access time attribute, and writing updates its modification time. When a file is created, all three timestamps for that file are set to the current time. In addition, the attribute change time and modification time fields of the directory that contains the new entry are updated.

Adding a new name for a file with the link function updates the attribute change time field of the file being linked, and both the attribute change time and modification time fields of the directory containing the new name. These same fields are affected if a file name is deleted with unlink, remove, or rmdir. Renaming a file with rename affects only the attribute change time and modification time fields of the two parent directories involved, and not the times for the file being renamed.

Changing attributes of a file (for example, with chmod) updates its attribute change time field.

You can also change some of the timestamps of a file explicitly using the utime function—all except the attribute change time. You need to include the header file ‘utime.h’ to use this facility.

Data Type struct utimbuf

The utimbuf structure is used with the utime function to specify new access and modification times for a file. It contains the following members:

  time_t actime

This is the access time for the file.

  time_t modtime

This is the modification time for the file.

Function int utime (const char *filename, const struct
      utimbuf *times)

This function is used to modify the file times associated with the file named filename. If times is a null pointer, then the access and modification times of the file are set to the current time. Otherwise, they are set to the values from the actime and modtime members (respectively) of the utimbuf structure pointed at by times. The attribute modification time for the file is set to the current time in either case (since changing the timestamps is itself a modification of the file attributes). The utime function returns 0 if successful and -1 on failure. In addition to the usual file name errors (see Section 6.2.3 [File Name Errors], page 89), the following errno error conditions are defined for this function:

EACCES There is a permission problem in the case where a null pointer was passed as the times argument. In order to update the timestamp on the file, you must either be the owner of the file, have write permission on the file, or be a privileged user.

ENOENT The file doesn’t exist.

EPERM If the times argument is not a null pointer, you must either be the owner of the file or be a privileged user. This error is used to report the problem.

EROFS The file lives on a read-only file system.

Each of the three time stamps has a corresponding microsecond part, which extends its resolution. These fields are called st_atime_usec, st_ mtime_usec, and st_ctime_usec; each has a value between 0 and 999,999, which indicates the time in microseconds. They correspond to the tv_usec field of a timeval structure; see Section 17.2.2 [High-Resolution Calendar], page 342.

The utimes function is like utime, but also lets you specify the fractional part of the file times. The prototype for this function is in the header file

‘sys/time.h’.
Function int utimes (const char *filename, struct timeval

tvp[2]) This function sets the file access and modification times for the file named by filename. The new file access time is specified by tvp[0], and the new modification time by tvp[1]. This function comes from BSD.

The return values and error conditions are the same as for the utime function.

9.9 Making Special Files

The mknod function is the primitive for making special files, such as files that correspond to devices. The GNU library includes this function for compatibility with BSD.

The prototype for mknod is declared in ‘sys/stat.h’.

Function int mknod (const char *filename, int mode, int

dev) The mknod function makes a special file with name filename. The mode specifies the mode of the file, and may include the various special file bits, such as S_IFCHR (for a character special file) or S_IFBLK (for a block special file). See Section 9.8.3 [Testing the Type of a File], page 199. The dev argument specifies which device the special file refers to. Its exact interpretation depends on the kind of special file being created. The return value is 0 on success and -1 on error. In addition to the usual file name errors (see Section 6.2.3 [File Name Errors], page 89), the following errno error conditions are defined for this function:

EPERM The calling process is not privileged. Only the superuser can create special files.

ENOSPC The directory or file system that would contain the new file is full and cannot be extended.

EROFS The directory containing the new file can’t be modified because it’s on a read-only file system.

EEXIST There is already a file named filename. If you want to replace this file, you must remove the old file explicitly first.

9.10 Temporary Files

If you need to use a temporary file in your program, you can use the tmpfile function to open it. Or you can use the tmpnam (better: tmpnam_r) function make a name for a temporary file and then open it in the usual way

with fopen.

The tempnam function is like tmpnam but lets you choose what directory temporary files will go in, and something about what their file names will look like. Important for multi threaded programs is that tempnam is reentrant while tmpnam is not since it returns a pointer to a static buffer.

These facilities are declared in the header file ‘stdio.h’.

Function FILE * tmpfile (void)

This function creates a temporary binary file for update mode, as if by calling fopen with mode "wb+". The file is deleted automatically when it is closed or when the program terminates. (On some other ISO C systems the file may fail to be deleted if the program terminates abnormally). This function is reentrant.

Function char * tmpnam (char *result)

This function constructs and returns a file name that is a valid file name and that does not name any existing file. If the result argument is a null pointer, the return value is a pointer to an internal static string, which might be modified by subsequent calls and therefore makes this function non-reentrant. Otherwise, the result argument should be a pointer to an array of at least L_tmpnam characters, and the result is written into that array. It is possible for tmpnam to fail if you call it too many times without removing previously created files. This is because the fixed length of a temporary file name gives room for only a finite number of different names. If tmpnam fails, it returns a null pointer.

Function char * tmpnam r (char *result)

This function is nearly identical to the tmpnam function. But it does not allow result to be a null pointer. In the later case a null pointer is returned. This function is reentrant because the non-reentrant situation of tmpnam cannot happen here.

Macro int L tmpnam

The value of this macro is an integer constant expression that represents the minimum allocation size of a string large enough to hold the file name generated by the tmpnam function.

Macro int TMP MAX

The macro TMP_MAX is a lower bound for how many temporary names you can create with tmpnam. You can rely on being able to call tmpnam at least this many times before it might fail saying you have made too many temporary file names. With the GNU library, you can create a very large number of temporary file names—if you actually create the files, you will probably run out of disk space before you run out of names. Some other systems have a fixed, small limit on the number of temporary files. The limit is never less than 25.

Function char * tempnam (const char *dir, const char

*prefix) This function generates a unique temporary filename. If prefix is not a null pointer, up to five characters of this string are used as a prefix for the file name. The return value is a string newly allocated with malloc; you should release its storage with free when it is no longer needed. Because the string is dynamically allocated this function is reentrant. The directory prefix for the temporary file name is determined by testing each of the following, in sequence. The directory must exist and be writable.

• The environment variable TMPDIR, if it is defined. For security reasons this only happens if the program is not SUID or SGID enabled.

• The dir argument, if it is not a null pointer.

• The value of the P_tmpdir macro.

• The directory ‘/tmp’.

This function is defined for SVID compatibility.

SVID Macro char * P tmpdir

This macro is the name of the default directory for temporary files.

Older Unix systems did not have the functions just described. Instead they used mktemp and mkstemp. Both of these functions work by modifying a file name template string you pass. The last six characters of this string must be ‘XXXXXX’. These six ‘X’s are replaced with six characters which make the whole string a unique file name. Usually the template string is something like ‘/tmp/prefixXXXXXX’, and each program uses a unique prefix.

Note: Because mktemp and mkstemp modify the template string, you must not pass string constants to them. String constants are normally in read-only storage, so your program would crash when mktemp or mkstemp tried to modify the string.

Function char * mktemp (char *template)

The mktemp function generates a unique file name by modifying template as described above. If successful, it returns template as modified. If mktemp cannot find a unique file name, it makes template an empty string and returns that. If template does not end with ‘XXXXXX’, mktemp returns a null pointer.

Function int mkstemp (char *template)

The mkstemp function generates a unique file name just as mktemp does, but it also opens the file for you with open (see Section 8.1 [Opening and Closing Files], page 151). If successful, it modifies template in place and returns a file descriptor open on that file for reading and writing. If mkstemp cannot create a uniquely-named file, it makes template an empty string and returns -1. If template does not end with ‘XXXXXX’, mkstemp returns -1 and does not modify template.

Unlike mktemp, mkstemp is actually guaranteed to create a unique file that cannot possibly clash with any other program trying to create a temporary file. This is because it works by calling open with the O_EXCL flag bit, which says you want to always create a new file, and get an error if the file already exists.

214 The GNU C Library

10 Pipes and FIFOs

A pipe is a mechanism for interprocess communication; data written to the pipe by one process can be read by another process. The data is handled in a first-in, first-out (FIFO) order. The pipe has no name; it is created for one use and both ends must be inherited from the single process which created the pipe.

A FIFO special file is similar to a pipe, but instead of being an anonymous, temporary connection, a FIFO has a name or names like any other file. Processes open the FIFO by name in order to communicate through it.

A pipe or FIFO has to be open at both ends simultaneously. If you read from a pipe or FIFO file that doesn’t have any processes writing to it (perhaps because they have all closed the file, or exited), the read returns end-of-file. Writing to a pipe or FIFO that doesn’t have a reading process is treated as an error condition; it generates a SIGPIPE signal, and fails with error code EPIPE if the signal is handled or blocked.

Neither pipes nor FIFO special files allow file positioning. Both reading and writing operations happen sequentially; reading from the beginning of the file and writing at the end.

10.1 Creating a Pipe

The primitive for creating a pipe is the pipe function. This creates both the reading and writing ends of the pipe. It is not very useful for a single process to use a pipe to talk to itself. In typical use, a process creates a pipe just before it forks one or more child processes (see Section 23.4 [Creating a Process], page 469). The pipe is then used for communication either between the parent or child processes, or between two sibling processes.

The pipe function is declared in the header file ‘unistd.h’.

Function int pipe (int filedes[2])

The pipe function creates a pipe and puts the file descriptors for the reading and writing ends of the pipe (respectively) into filedes[0] and filedes[1]. An easy way to remember that the input end comes first is that file descriptor 0 is standard input, and file descriptor 1 is standard output. If successful, pipe returns a value of 0. On failure, -1 is returned. The following errno error conditions are defined for this function:

EMFILE The process has too many files open.

ENFILE There are too many open files in the entire system. See Section 2.2 [Error Codes], page 16, for more information about ENFILE. This error never occurs in the GNU system.

Here is an example of a simple program that creates a pipe. This program uses the fork function (see Section 23.4 [Creating a Process], page 469) to create a child process. The parent process writes data to the pipe, which is read by the child process.

#include <sys/types.h>
#include <unistd.h>
#include <stdio.h>
#include <stdlib.h>

/* Read characters from the pipe and echo them to stdout. */

void
read_from_pipe (int file)
{
  FILE *stream;
  int c;
  stream = fdopen (file, "r");
  while ((c = fgetc (stream)) != EOF)
    putchar (c);
  fclose (stream);
}

/* Write some random text to the pipe. */

void
write_to_pipe (int file)
{
  FILE *stream;
  stream = fdopen (file, "w");
  fprintf (stream, "hello, world!\n");
  fprintf (stream, "goodbye, world!\n");
  fclose (stream);
}
int
main (void)
{
  pid_t pid;
  int mypipe[2];

/* Create the pipe. */

       if (pipe (mypipe))
         {
           fprintf (stderr, "Pipe failed.\n");
           return EXIT_FAILURE;
         }

/* Create the child process. */

       pid = fork ();
       if (pid == (pid_t) 0)
         {

/* This is the child process. */

           read_from_pipe (mypipe[0]);
           return EXIT_SUCCESS;
         }
       else if (pid < (pid_t) 0)
         {

/* The fork failed. */

           fprintf (stderr, "Fork failed.\n");
           return EXIT_FAILURE;
         }
       else
         {

/* This is the parent process. */

           write_to_pipe (mypipe[1]);
           return EXIT_SUCCESS;
         }
     }

10.2 Pipe to a Subprocess

A common use of pipes is to send data to or receive data from a program being run as subprocess. One way of doing this is by using a combination of pipe (to create the pipe), fork (to create the subprocess), dup2 (to force the subprocess to use the pipe as its standard input or output channel), and exec (to execute the new program). Or, you can use popen and pclose.

The advantage of using popen and pclose is that the interface is much simpler and easier to use. But it doesn’t offer as much flexibility as using the low-level functions directly.

Function FILE * popen (const char *command, const char

*mode) The popen function is closely related to the system function; see Section 23.1 [Running a Command], page 467. It executes the shell command command as a subprocess. However, instead of waiting for the command to complete, it creates a pipe to the subprocess and returns a stream that corresponds to that pipe.

If you specify a mode argument of "r", you can read from the stream to retrieve data from the standard output channel of the subprocess. The subprocess inherits its standard input channel from the parent process.

Similarly, if you specify a mode argument of "w", you can write to the stream to send data to the standard input channel of the subprocess. The subprocess inherits its standard output channel from the parent process.

In the event of an error, popen returns a null pointer. This might happen if the pipe or stream cannot be created, if the subprocess cannot be forked, or if the program cannot be executed.

Function int pclose (FILE *stream)

The pclose function is used to close a stream created by popen. It waits for the child process to terminate and returns its status value, as for the system function.

Here is an example showing how to use popen and pclose to filter output through another program, in this case the paging program more.

  #include <stdio.h>
  #include <stdlib.h>
  void
  write_data (FILE * stream)
  {
    int i;
    for (i = 0; i < 100; i++)
      fprintf (stream, "%d\n", i);
    if (ferror (stream))
      {
        fprintf (stderr, "Output to stream failed.\n");
        exit (EXIT_FAILURE);
      }
  }

  int
  main (void)
  {
    FILE *output;
    output = popen ("more", "w");
    if (!output)
      {
        fprintf (stderr, "Could not run more.\n");
        return EXIT_FAILURE;
      }
    write_data (output);
    pclose (output);
    return EXIT_SUCCESS;
  }

10.3 FIFO Special Files

A FIFO special file is similar to a pipe, except that it is created in a different way. Instead of being an anonymous communications channel, a FIFO special file is entered into the file system by calling mkfifo.

Once you have created a FIFO special file in this way, any process can open it for reading or writing, in the same way as an ordinary file. However, it has to be open at both ends simultaneously before you can proceed to do any input or output operations on it. Opening a FIFO for reading normally blocks until some other process opens the same FIFO for writing, and vice versa.

The mkfifo function is declared in the header file ‘sys/stat.h’.

Function int mkfifo (const char *filename, mode_t mode)

The mkfifo function makes a FIFO special file with name filename. The mode argument is used to set the file’s permissions; see Section 9.8.7 [Assigning File Permissions], page 204. The normal, successful return value from mkfifo is 0. In the case of an error, -1 is returned. In addition to the usual file name errors (see Section 6.2.3 [File Name Errors], page 89), the following errno error conditions are defined for this function:

EEXIST The named file already exists.

ENOSPC The directory or file system cannot be extended.

EROFS The directory that would contain the file resides on a readonly file system.

10.4 Atomicity of Pipe I/O

Reading or writing pipe data is atomic if the size of data written is not greater than PIPE_BUF. This means that the data transfer seems to be an instantaneous unit, in that nothing else in the system can observe a state in which it is partially complete. Atomic I/O may not begin right away (it may need to wait for buffer space or for data), but once it does begin, it finishes immediately.

Reading or writing a larger amount of data may not be atomic; for example, output data from other processes sharing the descriptor may be interspersed. Also, once PIPE_BUF characters have been written, further writes will block until some characters are read.

See Section 28.6 [Limits on File System Capacity], page 545, for information about the PIPE_BUF parameter.

11 Sockets

This chapter describes the GNU facilities for interprocess communication using sockets.

A socket is a generalized interprocess communication channel. Like a pipe, a socket is represented as a file descriptor. But, unlike pipes, sockets support communication between unrelated processes, and even between processes running on different machines that communicate over a network. Sockets are the primary means of communicating with other machines; telnet, rlogin, ftp, talk, and the other familiar network programs use sockets.

Not all operating systems support sockets. In the GNU library, the header file ‘sys/socket.h’ exists regardless of the operating system, and the socket functions always exist, but if the system does not really support sockets, these functions always fail.

Incomplete: We do not currently document the facilities for broadcast messages or for configuring Internet interfaces.

11.1 Socket Concepts

When you create a socket, you must specify the style of communication you want to use and the type of protocol that should implement it. The communication style of a socket defines the user-level semantics of sending and receiving data on the socket. Choosing a communication style specifies the answers to questions such as these:

• What are the units of data transmission? Some communication styles regard the data as a sequence of bytes, with no larger structure; others group the bytes into records (which are known in this context as packets).

• Can data be lost during normal operation? Some communication styles guarantee that all the data sent arrives in the order it was sent (barring system or network crashes); other styles occasionally lose data as a normal part of operation, and may sometimes deliver packets more than once or in the wrong order. Designing a program to use unreliable communication styles usually involves taking precautions to detect lost or misordered packets and to retransmit data as needed.

• Is communication entirely with one partner? Some communication styles are like a telephone call—you make a connection with one remote socket, and then exchange data freely. Other styles are like mailing letters—you specify a destination address for each message you send.

You must also choose a namespace for naming the socket. A socket name (“address”) is meaningful only in the context of a particular namespace.

In fact, even the data type to use for a socket name may depend on the namespace. Namespaces are also called “domains”, but we avoid that word as it can be confused with other usage of the same term. Each namespace has a symbolic name that starts with ‘PF_’. A corresponding symbolic name starting with ‘AF_’ designates the address format for that namespace.

Finally you must choose the protocol to carry out the communication. The protocol determines what low-level mechanism is used to transmit and receive data. Each protocol is valid for a particular namespace and communication style; a namespace is sometimes called a protocol family because of this, which is why the namespace names start with ‘PF_’.

The rules of a protocol apply to the data passing between two programs, perhaps on different computers; most of these rules are handled by the operating system, and you need not know about them. What you do need to know about protocols is this:

• In order to have communication between two sockets, they must specify the same protocol.

• Each protocol is meaningful with particular style/namespace combinations and cannot be used with inappropriate combinations. For example, the TCP protocol fits only the byte stream style of communication and the Internet namespace.

• For each combination of style and namespace, there is a default protocol which you can request by specifying 0 as the protocol number. And that’s what you should normally do—use the default. Throughout the following description at various places variables/parameters

to denote sizes are required. And here the trouble starts. In the first implementations the type of these variables was simply int. This type was on almost all machines of this time 32 bits wide and so a de-factor standard required 32 bit variables. This is important since references to variables of this type are passed to the kernel.

But now the POSIX people came and unified the interface with their words "all size values are of type size_t". But on 64 bit machines size_t is 64 bits wide and so variable references are not anymore possible.

A solution provides the Unix98 specification which finally introduces a type socklen_t. This type is used in all of the cases in previously changed to use size_t. The only requirement of this type is that it is an unsigned type of at least 32 bits. Therefore, implementations which require references to 32 bit variables be passed can be as happy as implementations which right from the start of 64 bit values.

11.2 Communication Styles

The GNU library includes support for several different kinds of sockets, each with different characteristics. This section describes the supported socket types. The symbolic constants listed here are defined in

‘sys/socket.h’.

Macro int SOCK STREAM

The SOCK_STREAM style is like a pipe (see Chapter 10 [Pipes and FI-FOs], page 215); it operates over a connection with a particular remote socket, and transmits data reliably as a stream of bytes. Use of this style is covered in detail in Section 11.8 [Using Sockets with Connections], page 246.

Macro int SOCK DGRAM

The SOCK_DGRAM style is used for sending individually-addressed packets, unreliably. It is the diametrical opposite of SOCK_STREAM. Each time you write data to a socket of this kind, that data becomes one packet. Since SOCK_DGRAM sockets do not have connections, you must specify the recipient address with each packet. The only guarantee that the system makes about your requests to transmit data is that it will try its best to deliver each packet you send. It may succeed with the sixth packet after failing with the fourth and fifth packets; the seventh packet may arrive before the sixth, and may arrive a second time after the sixth. The typical use for SOCK_DGRAM is in situations where it is acceptable to simply resend a packet if no response is seen in a reasonable amount of time. See Section 11.9 [Datagram Socket Operations], page 260, for detailed information about how to use datagram sockets.

Macro int SOCK RAW

This style provides access to low-level network protocols and interfaces. Ordinary user programs usually have no need to use this style.

11.3 Socket Addresses

The name of a socket is normally called an address. The functions and symbols for dealing with socket addresses were named inconsistently, sometimes using the term “name” and sometimes using “address”. You can regard these terms as synonymous where sockets are concerned.

A socket newly created with the socket function has no address. Other processes can find it for communication only if you give it an address. We call this binding the address to the socket, and the way to do it is with the bind function.

You need be concerned with the address of a socket if other processes are to find it and start communicating with it. You can specify an address for other sockets, but this is usually pointless; the first time you send data from a socket, or use it to initiate a connection, the system assigns an address automatically if you have not specified one.

Occasionally a client needs to specify an address because the server discriminates based on addresses; for example, the rsh and rlogin protocols look at the client’s socket address and don’t bypass password checking unless it is less than IPPORT_RESERVED (see Section 11.5.3 [Internet Ports], page 237).

The details of socket addresses vary depending on what namespace you are using. See Section 11.4 [The File Namespace], page 227, or Section 11.5 [The Internet Namespace], page 229, for specific information.

Regardless of the namespace, you use the same functions bind and getsockname to set and examine a socket’s address. These functions use a phony data type, struct sockaddr *, to accept the address. In practice, the address lives in a structure of some other data type appropriate to the address format you are using, but you cast its address to struct sockaddr * when you pass it to bind.

11.3.1 Address Formats

The functions bind and getsockname use the generic data type struct sockaddr * to represent a pointer to a socket address. You can’t use this data type effectively to interpret an address or construct one; for that, you must use the proper data type for the socket’s namespace.

Thus, the usual practice is to construct an address in the proper namespace-specific type, then cast a pointer to struct sockaddr * when

you call bind or getsockname.

The one piece of information that you can get from the struct sockaddr data type is the address format designator which tells you which data type to use to understand the address fully.

The symbols in this section are defined in the header file ‘sys/socket.h’.

Date Type struct sockaddr

The struct sockaddr type itself has the following members:

  short int sa_family

This is the code for the address format of this address. It identifies the format of the data which follows.

  char sa_data[14]

This is the actual socket address data, which is formatdependent. Its length also depends on the format, and may well be more than 14. The length 14 of sa_data is essentially arbitrary.

Each address format has a symbolic name which starts with ‘AF_’. Each of them corresponds to a ‘PF_’ symbol which designates the corresponding namespace. Here is a list of address format names:

AF_FILE This designates the address format that goes with the file namespace. (PF_FILE is the name of that namespace.) See Section 11.4.2 [Details of File Namespace], page 227, for information about this address format.

AF_UNIX This is a synonym for AF_FILE, for compatibility. (PF_UNIX is likewise a synonym for PF_FILE.)

AF_INET This designates the address format that goes with the Internet namespace. (PF_INET is the name of that namespace.) See Section 11.5.1 [Internet Socket Address Formats], page 230.

AF_INET6 This is similar to AF_INET, but refers to the IPv6 protocol. (PF_ INET6 is the name of the corresponding namespace.)

AF_UNSPEC

This designates no particular address format. It is used only in rare cases, such as to clear out the default destination address of a “connected” datagram socket. See Section 11.9.1 [Sending Datagrams], page 260. The corresponding namespace designator symbol PF_UNSPEC exists for completeness, but there is no reason to use it in a program.

‘sys/socket.h’ defines symbols starting with ‘AF_’ for many different kinds of networks, all or most of which are not actually implemented. We will document those that really work, as we receive information about how to use them.

11.3.2 Setting the Address of a Socket

Use the bind function to assign an address to a socket. The prototype for bind is in the header file ‘sys/socket.h’. For examples of use, see Section 11.4 [The File Namespace], page 227, or see Section 11.5.7 [Internet Socket Example], page 242.

Function int bind (int socket, struct sockaddr *addr,
socklen_t length)

The bind function assigns an address to the socket socket. The addr and length arguments specify the address; the detailed format of the address depends on the namespace. The first part of the address is always the format designator, which specifies a namespace, and says that the address is in the format for that namespace. The return value is 0 on success and -1 on failure. The following errno error conditions are defined for this function:

EBADF The socket argument is not a valid file descriptor.

ENOTSOCK The descriptor socket is not a socket.

EADDRNOTAVAIL

The specified address is not available on this machine.

EADDRINUSE

Some other socket is already using the specified address.

EINVAL The socket socket already has an address.

EACCES You do not have permission to access the requested address. (In the Internet domain, only the super-user is allowed to specify a port number in the range 0 through IPPORT_RESERVED minus one; see Section 11.5.3 [Internet Ports], page 237.)

Additional conditions may be possible depending on the particular namespace of the socket.

11.3.3 Reading the Address of a Socket

Use the function getsockname to examine the address of an Internet socket. The prototype for this function is in the header file ‘sys/socket.h’.

Function int getsockname (int socket, struct sockaddr

*addr, socklen_t *length-ptr) The getsockname function returns information about the address of the socket socket in the locations specified by the addr and length-ptr arguments. Note that the length-ptr is a pointer; you should initialize it to be the allocation size of addr, and on return it contains the actual size of the address data. The format of the address data depends on the socket namespace. The length of the information is usually fixed for a given namespace, so normally you can know exactly how much space is needed and can provide that much. The usual practice is to allocate a place for the value using the proper data type for the socket’s namespace, then cast

its address to struct sockaddr * to pass it to getsockname.

The return value is 0 on success and -1 on error. The following errno error conditions are defined for this function:

EBADF The socket argument is not a valid file descriptor.

ENOTSOCK The descriptor socket is not a socket.

ENOBUFS There are not enough internal buffers available for the operation.

You can’t read the address of a socket in the file namespace. This is consistent with the rest of the system; in general, there’s no way to find a file’s name from a descriptor for that file.

11.4 The File Namespace

This section describes the details of the file namespace, whose symbolic name (required when you create a socket) is PF_FILE.

11.4.1 File Namespace Concepts

In the file namespace, socket addresses are file names. You can specify any file name you want as the address of the socket, but you must have write permission on the directory containing it. In order to connect to a socket, you must have read permission for it. It’s common to put these files in the ‘/tmp’ directory.

One peculiarity of the file namespace is that the name is only used when opening the connection; once that is over with, the address is not meaningful and may not exist.

Another peculiarity is that you cannot connect to such a socket from another machine–not even if the other machine shares the file system which contains the name of the socket. You can see the socket in a directory listing, but connecting to it never succeeds. Some programs take advantage of this, such as by asking the client to send its own process ID, and using the process IDs to distinguish between clients. However, we recommend you not use this method in protocols you design, as we might someday permit connections from other machines that mount the same file systems. Instead, send each new client an identifying number if you want it to have one.

After you close a socket in the file namespace, you should delete the file name from the file system. Use unlink or remove to do this; see Section 9.5 [Deleting Files], page 192.

The file namespace supports just one protocol for any communication style; it is protocol number 0.

11.4.2 Details of File Namespace

To create a socket in the file namespace, use the constant PF_FILE as the namespace argument to socket or socketpair. This constant is defined in

‘sys/socket.h’.

Macro int PF FILE

This designates the file namespace, in which socket addresses are file names, and its associated family of protocols.

Macro int PF UNIX

This is a synonym for PF_FILE, for compatibility’s sake.

The structure for specifying socket names in the file namespace is defined in the header file ‘sys/un.h’:

Data Type struct sockaddr un

This structure is used to specify file namespace socket addresses. It has the following members:

short int sun_family

This identifies the address family or format of the socket address. You should store the value AF_FILE to designate the file namespace. See Section 11.3 [Socket Addresses], page 223.

char sun_path[108]

This is the file name to use. Incomplete: Why is 108 a magic number? RMS suggests making this a zero-length array and tweaking the example following to use alloca to allocate an appropriate amount of storage based on the length of the filename.

You should compute the length parameter for a socket address in the file namespace as the sum of the size of the sun_family component and the string length (not the allocation size!) of the file name string.

11.4.3 Example of File-Namespace Sockets

Here is an example showing how to create and name a socket in the file namespace.

#include <stddef.h>
#include <stdio.h>
#include <errno.h>
#include <stdlib.h>
#include <sys/socket.h>
#include <sys/un.h>
int
make_named_socket (const char *filename)
{
struct sockaddr_un name;
int sock;
size_t size;

/* Create the socket. */

sock = socket (PF_UNIX, SOCK_DGRAM, 0);
if (sock < 0)
{
perror ("socket");
exit (EXIT_FAILURE);

         }

/* Bind a name to the socket. */

       name.sun_family = AF_FILE;
       strcpy (name.sun_path, filename);

/* The size of the address is

the offset of the start of the filename, plus its length, plus one for the terminating null byte. */

       size = (offsetof (struct sockaddr_un, sun_path)
               + strlen (name.sun_path) + 1);
       if (bind (sock, (struct sockaddr *) &name, size) < 0)
         {
           perror ("bind");
           exit (EXIT_FAILURE);
         }
       return sock;
     }

11.5 The Internet Namespace

This section describes the details the protocols and socket naming conventions used in the Internet namespace.

To create a socket in the Internet namespace, use the symbolic name PF_INET of this namespace as the namespace argument to socket or socketpair. This macro is defined in ‘sys/socket.h’.

Macro int PF INET

This designates the Internet namespace and associated family of protocols.

A socket address for the Internet namespace includes the following components:

• The address of the machine you want to connect to. Internet addresses can be specified in several ways; these are discussed in Section 11.5.1 [Internet Socket Address Formats], page 230, Section 11.5.2 [Host Addresses], page 231, and Section 11.5.2.4 [Host Names], page 235.

• A port number for that machine. See Section 11.5.3 [Internet Ports], page 237.

You must ensure that the address and port number are represented in a canonical format called network byte order. See Section 11.5.5 [Byte Order Conversion], page 239, for information about this.

11.5.1 Internet Socket Address Formats

In the Internet namespace, for both IPv4 (AF_INET) and IPv6 (AF_ INET6), a socket address consists of a host address and a port on that host. In addition, the protocol you choose serves effectively as a part of the address because local port numbers are meaningful only within a particular protocol.

The data types for representing socket addresses in the Internet namespace are defined in the header file ‘netinet/in.h’.

Data Type struct sockaddr in

This is the data type used to represent socket addresses in the Internet namespace. It has the following members:

short int sin_family

This identifies the address family or format of the socket address. You should store the value of AF_INET in this member. See Section 11.3 [Socket Addresses], page 223.

struct in_addr sin_addr

This is the Internet address of the host machine. See Section 11.5.2 [Host Addresses], page 231, and Section 11.5.2.4 [Host Names], page 235, for how to get a value to store here.

unsigned short int sin_port

This is the port number. See Section 11.5.3 [Internet Ports], page 237.

When you call bind or getsockname, you should specify sizeof (struct sockaddr_in) as the length parameter if you are using an Internet namespace socket address.

Data Type struct sockaddr in6

This is the data type used to represent socket addresses in the IPv6 namespace. It has the following members:

short int sin6_family

This identifies the address family or format of the socket address. You should store the value of AF_INET6 in this member. See Section 11.3 [Socket Addresses], page 223.

struct in6_addr sin6_addr

This is the IPv6 address of the host machine. See Section 11.5.2 [Host Addresses], page 231, and Section 11.5.2.4 [Host Names], page 235, for how to get a value to store here.

  uint32_t sin6_flowinfo

This is a currently unimplemented field.

  uint16_t sin6_port

This is the port number. See Section 11.5.3 [Internet Ports], page 237.

11.5.2 Host Addresses

Each computer on the Internet has one or more Internet addresses, numbers which identify that computer among all those on the Internet. Users typically write IPv4 numeric host addresses as sequences of four numbers, separated by periods, as in ‘128.52.46.32’, and IPv6 numeric host addresses as sequences of up to eight numbers seperated by colons, as in

‘5f03:1200:836f:c100::1’.

Each computer also has one or more host names, which are strings of words separated by periods, as in ‘churchy.gnu.ai.mit.edu’.

Programs that let the user specify a host typically accept both numeric addresses and host names. But the program needs a numeric address to open a connection; to use a host name, you must convert it to the numeric address it stands for.

11.5.2.1 Internet Host Addresses

An Internet host address is a number containing four bytes of data. These are divided into two parts, a network number and a local network address number within that network. The network number consists of the first one, two or three bytes; the rest of the bytes are the local address.

Network numbers are registered with the Network Information Center (NIC), and are divided into three classes—A, B, and C. The local network address numbers of individual machines are registered with the administrator of the particular network.

Class A networks have single-byte numbers in the range 0 to 127. There are only a small number of Class A networks, but they can each support a very large number of hosts. Medium-sized Class B networks have twobyte network numbers, with the first byte in the range 128 to 191. Class C networks are the smallest; they have three-byte network numbers, with the first byte in the range 192-255. Thus, the first 1, 2, or 3 bytes of an Internet address specifies a network. The remaining bytes of the Internet address specify the address within that network.

The Class A network 0 is reserved for broadcast to all networks. In addition, the host number 0 within each network is reserved for broadcast to all hosts in that network.

The Class A network 127 is reserved for loopback; you can always use the Internet address ‘127.0.0.1’ to refer to the host machine.

Since a single machine can be a member of multiple networks, it can have multiple Internet host addresses. However, there is never supposed to be more than one machine with the same host address.

There are four forms of the standard numbers-and-dots notation for Internet addresses:

a.b.c.d This specifies all four bytes of the address individually.

a.b.c The last part of the address, c, is interpreted as a 2-byte quantity. This is useful for specifying host addresses in a Class B network with network address number a.b.

a.b The last part of the address, c, is interpreted as a 3-byte quantity. This is useful for specifying host addresses in a Class A network with network address number a.

a If only one part is given, this corresponds directly to the host address number.

Within each part of the address, the usual C conventions for specifying the radix apply. In other words, a leading ‘0x’ or ‘0X’ implies hexadecimal radix; a leading ‘0’ implies octal; and otherwise decimal radix is assumed.

11.5.2.2 Host Address Data Type

Internet host addresses are represented in some contexts as integers (type unsigned long int). In other contexts, the integer is packaged inside a structure of type struct in_addr. It would be better if the usage were made consistent, but it is not hard to extract the integer from the structure or put the integer into a structure.

The following basic definitions for Internet addresses appear in the header

file ‘netinet/in.h’:

Data Type struct in addr

This data type is used in certain contexts to contain an Internet host address. It has just one field, named s_addr, which records the host address number as an unsigned long int.

Macro unsigned int INADDR LOOPBACK

You can use this constant to stand for “the address of this machine,” instead of finding its actual address. It is the Internet address ‘127.0.0.1’, which is usually called ‘localhost’. This special constant saves you the trouble of looking up the address of your own machine. Also, the system usually implements INADDR_LOOPBACK specially, avoiding any network traffic for the case of one machine talking to itself.

Macro unsigned int INADDR ANY

You can use this constant to stand for “any incoming address,” when binding to an address. See Section 11.3.2 [Setting the Address of a Socket], page 225. This is the usual address to give in the sin_addr member of struct sockaddr_in when you want to accept Internet connections.

Macro unsigned int INADDR BROADCAST

This constant is the address you use to send a broadcast message.

Macro unsigned int INADDR NONE

This constant is returned by some functions to indicate an error.

Data Type struct in6 addr

This data type is used to store an IPv6 address. It stores 128 bits of data, which can be accessed (via a union) in a variety of ways.

Constant struct in6_addr in6addr loopback.

This constant is the IPv6 address ‘::1’, the loopback address. See above for a description of what this means. The macro IN6ADDR_ LOOPBACK_INIT is provided to allow you to initialise your own variables to this value.

Constant struct in6_addr in6addr any

This constant is the IPv6 address ‘::’, the unspecified address. See above for a description of what this means. The macro IN6ADDR_ANY_ INIT is provided to allow you to initialise your own variables to this value.

11.5.2.3 Host Address Functions

These additional functions for manipulating Internet addresses are declared in ‘arpa/inet.h’. They represent Internet addresses in network byte order; they represent network numbers and local-address-within-network numbers in host byte order. See Section 11.5.5 [Byte Order Conversion], page 239, for an explanation of network and host byte order.

Function int inet aton (const char *name, struct in_addr

*addr) This function converts the Internet host address name from the standard numbers-and-dots notation into binary data and stores it in the struct in_addr that addr points to. inet_aton returns nonzero if the address is valid, zero if not.

Function unsigned long int inet addr (const char *name)

This function converts the Internet host address name from the standard numbers-and-dots notation into binary data. If the input is not valid, inet_addr returns INADDR_NONE. This is an obsolete interface to inet_aton, described immediately above; it is obsolete because INADDR_NONE is a valid address (255.255.255.255), and inet_aton provides a cleaner way to indicate error return.

Function unsigned long int inet network (const char

*name) This function extracts the network number from the address name, given in the standard numbers-and-dots notation. If the input is not

valid, inet_network returns -1.

Function char * inet ntoa (struct in_addr addr)

This function converts the Internet host address addr to a string in the standard numbers-and-dots notation. The return value is a pointer into a statically-allocated buffer. Subsequent calls will overwrite the same buffer, so you should copy the string if you need to save it. In multi-threaded programs each thread has an own staticallyallocated buffer. But still subsequent calls of inet_ntoa in the same thread will overwrite the result of the last call.

Function struct in_addr inet makeaddr (int net, int local)

This function makes an Internet host address by combining the network number net with the local-address-within-network number local.

Function int inet lnaof (struct in_addr addr)

This function returns the local-address-within-network part of the Internet host address addr.

Function int inet netof (struct in_addr addr)

This function returns the network number part of the Internet host address addr.

Function int inet pton (int af, const char *cp, void *buf )

This function converts an Internet address (either IPv4 or IPv6) from presentation (textual) to network (binary) format. af should be either AF_INET or AF_INET6, as appropriate for the type of address being converted. cp is a pointer to the input string, and buf is a pointer to a buffer for the result. It is the caller’s responsibility to make sure the buffer is large enough.

Function char * inet ntop (int af, const void *cp, char *buf,
      size_t len)

This function converts an Internet address (either IPv4 or IPv6) from network (binary) to presentation (textual) form. af should be either AF_INET or AF_INET6, as appropriate. cp is a pointer to the address to be converted. buf should be a pointer to a buffer to hold the result, and len is the length of this buffer. The return value from the function will be this buffer address.

11.5.2.4 Host Names

Besides the standard numbers-and-dots notation for Internet addresses, you can also refer to a host by a symbolic name. The advantage of a symbolic name is that it is usually easier to remember. For example, the machine with Internet address ‘128.52.46.32’ is also known as

‘churchy.gnu.ai.mit.edu’; and other machines in the ‘gnu.ai.mit.edu’

domain can refer to it simply as ‘churchy’.

Internally, the system uses a database to keep track of the mapping between host names and host numbers. This database is usually either the file ‘/etc/hosts’ or an equivalent provided by a name server. The functions and other symbols for accessing this database are declared in ‘netdb.h’. They are BSD features, defined unconditionally if you include ‘netdb.h’.

Data Type struct hostent

This data type is used to represent an entry in the hosts database. It has the following members:

char *h_name

This is the “official” name of the host.

char **h_aliases

These are alternative names for the host, represented as a null-terminated vector of strings.

int h_addrtype

This is the host address type; in practice, its value is always either AF_INET or AF_INET6, with the latter being used for IPv6 hosts. In principle other kinds of addresses could be represented in the data base as well as Internet addresses; if this were done, you might find a value in this field other than AF_INET or AF_INET6. See Section 11.3 [Socket Addresses], page 223.

int h_length

This is the length, in bytes, of each address.

char **h_addr_list

This is the vector of addresses for the host. (Recall that the host might be connected to multiple networks and have different addresses on each one.) The vector is terminated by a null pointer.

  char *h_addr

This is a synonym for h_addr_list[0]; in other words, it is the first host address.

As far as the host database is concerned, each address is just a block of memory h_length bytes long. But in other contexts there is an implicit assumption that you can convert this to a struct in_addr or an unsigned long int. Host addresses in a struct hostent structure are always given in network byte order; see Section 11.5.5 [Byte Order Conversion], page 239.

   You can use gethostbyname, gethostbyname2 or gethostbyaddr to

search the hosts database for information about a particular host. The information is returned in a statically-allocated structure; you must copy the information if you need to save it across calls. You can also use getaddrinfo and getnameinfo to obtain this information.

Function struct hostent * gethostbyname (const char

*name) The gethostbyname function returns information about the host named name. If the lookup fails, it returns a null pointer.

Function struct hostent * gethostbyname2 (const char

*name, int af ) The gethostbyname2 function is like gethostbyname, but allows the caller to specify the desired address family (e.g. AF_INET or AF_INET6) for the result.

Function struct hostent * gethostbyaddr (const char

*addr, int length, int format) The gethostbyaddr function returns information about the host with Internet address addr. The length argument is the size (in bytes) of the address at addr. format specifies the address format; for an Internet address, specify a value of AF_INET. If the lookup fails, gethostbyaddr returns a null pointer.

If the name lookup by gethostbyname or gethostbyaddr fails, you can find out the reason by looking at the value of the variable h_errno. (It would be cleaner design for these functions to set errno, but use of h_errno is compatible with other systems.) Before using h_errno, you must declare it like this:

     extern int h_errno;

Here are the error codes that you may find in h_errno:

HOST_NOT_FOUND

No such host is known in the data base.

TRY_AGAIN

This condition happens when the name server could not be contacted. If you try again later, you may succeed then.

NO_RECOVERY

A non-recoverable error occurred.

NO_ADDRESS

The host database contains an entry for the name, but it doesn’t have an associated Internet address.

You can also scan the entire hosts database one entry at a time using sethostent, gethostent, and endhostent. Be careful in using these functions, because they are not reentrant.

Function void sethostent (int stayopen)

This function opens the hosts database to begin scanning it. You can then call gethostent to read the entries. If the stayopen argument is nonzero, this sets a flag so that subsequent calls to gethostbyname or gethostbyaddr will not close the database (as they usually would). This makes for more efficiency if you call those functions several times, by avoiding reopening the database for each call.

Function struct hostent * gethostent ()

This function returns the next entry in the hosts database. It returns a null pointer if there are no more entries.

Function void endhostent ()

This function closes the hosts database.

11.5.3 Internet Ports

A socket address in the Internet namespace consists of a machine’s Internet address plus a port number which distinguishes the sockets on a given machine (for a given protocol). Port numbers range from 0 to 65,535.

Port numbers less than IPPORT_RESERVED are reserved for standard servers, such as finger and telnet. There is a database that keeps track of these, and you can use the getservbyname function to map a service name onto a port number; see Section 11.5.4 [The Services Database], page 238.

If you write a server that is not one of the standard ones defined in the database, you must choose a port number for it. Use a number greater than IPPORT_USERRESERVED; such numbers are reserved for servers and won’t ever be generated automatically by the system. Avoiding conflicts with servers being run by other users is up to you.

When you use a socket without specifying its address, the system generates a port number for it. This number is between IPPORT_RESERVED and

IPPORT_USERRESERVED.

On the Internet, it is actually legitimate to have two different sockets with the same port number, as long as they never both try to communicate with the same socket address (host address plus port number). You shouldn’t duplicate a port number except in special circumstances where a higherlevel protocol requires it. Normally, the system won’t let you do it; bind normally insists on distinct port numbers. To reuse a port number, you must set the socket option SO_REUSEADDR. See Section 11.11.2 [Socket-Level Options], page 267.

These macros are defined in the header file ‘netinet/in.h’.

Macro int IPPORT RESERVED

Port numbers less than IPPORT_RESERVED are reserved for superuser use.

Macro int IPPORT USERRESERVED

Port numbers greater than or equal to IPPORT_USERRESERVED are reserved for explicit use; they will never be allocated automatically.

11.5.4 The Services Database

The database that keeps track of “well-known” services is usually either the file ‘/etc/services’ or an equivalent from a name server. You can use these utilities, declared in ‘netdb.h’, to access the services database.

Data Type struct servent

This data type holds information about entries from the services database. It has the following members:

  char *s_name

This is the “official” name of the service.

  char **s_aliases

These are alternate names for the service, represented as an array of strings. A null pointer terminates the array.

  int s_port

This is the port number for the service. Port numbers are given in network byte order; see Section 11.5.5 [Byte Order Conversion], page 239.

  char *s_proto

This is the name of the protocol to use with this service. See Section 11.5.6 [Protocols Database], page 240.

To get information about a particular service, use the getservbyname or getservbyport functions. The information is returned in a staticallyallocated structure; you must copy the information if you need to save it across calls.

Function struct servent * getservbyname (const char
      *name, const char *proto)

The getservbyname function returns information about the service named name using protocol proto. If it can’t find such a service, it returns a null pointer. This function is useful for servers as well as for clients; servers use it to determine which port they should listen on (see Section 11.8.2 [Listening for Connections], page 248).

Function struct servent * getservbyport (int port, const

char *proto) The getservbyport function returns information about the service at port port using protocol proto. If it can’t find such a service, it returns a null pointer.

You can also scan the services database using setservent, getservent, and endservent. Be careful in using these functions, because they are not reentrant.

Function void setservent (int stayopen)

This function opens the services database to begin scanning it. If the stayopen argument is nonzero, this sets a flag so that subsequent calls to getservbyname or getservbyport will not close the database (as they usually would). This makes for more efficiency if you call those functions several times, by avoiding reopening the database for each call.

Function struct servent * getservent (void)

This function returns the next entry in the services database. If there are no more entries, it returns a null pointer.

Function void endservent (void)

This function closes the services database.

11.5.5 Byte Order Conversion

Different kinds of computers use different conventions for the ordering of bytes within a word. Some computers put the most significant byte within a word first (this is called “big-endian” order), and others put it last (“littleendian” order).

So that machines with different byte order conventions can communicate, the Internet protocols specify a canonical byte order convention for data transmitted over the network. This is known as the network byte order.

When establishing an Internet socket connection, you must make sure that the data in the sin_port and sin_addr members of the sockaddr_in structure are represented in the network byte order. If you are encoding integer data in the messages sent through the socket, you should convert this to network byte order too. If you don’t do this, your program may fail when running on or talking to other kinds of machines.

   If you use getservbyname and gethostbyname or inet_addr to get the

port number and host address, the values are already in the network byte order, and you can copy them directly into the sockaddr_in structure.

Otherwise, you have to convert the values explicitly. Use htons and ntohs to convert values for the sin_port member. Use htonl and ntohl to convert values for the sin_addr member. (Remember, struct in_addr is equivalent to unsigned long int.) These functions are declared in ‘netinet/in.h’.

Function unsigned short int htons (unsigned short int

hostshort) This function converts the short integer hostshort from host byte order to network byte order.

Function unsigned short int ntohs (unsigned short int

netshort) This function converts the short integer netshort from network byte order to host byte order.

Function unsigned long int htonl (unsigned long int

hostlong) This function converts the long integer hostlong from host byte order to network byte order.

Function unsigned long int ntohl (unsigned long int netlong)

This function converts the long integer netlong from network byte order to host byte order.

11.5.6 Protocols Database

The communications protocol used with a socket controls low-level details of how data is exchanged. For example, the protocol implements things like checksums to detect errors in transmissions, and routing instructions for messages. Normal user programs have little reason to mess with these details directly.

The default communications protocol for the Internet namespace depends on the communication style. For stream communication, the default is TCP (“transmission control protocol”). For datagram communication, the default is UDP (“user datagram protocol”). For reliable datagram communication, the default is RDP (“reliable datagram protocol”). You should nearly always use the default.

Internet protocols are generally specified by a name instead of a number. The network protocols that a host knows about are stored in a database.

This is usually either derived from the file ‘/etc/protocols’, or it may be an equivalent provided by a name server. You look up the protocol number associated with a named protocol in the database using the getprotobyname function.

Here are detailed descriptions of the utilities for accessing the protocols database. These are declared in ‘netdb.h’.

Data Type struct protoent

This data type is used to represent entries in the network protocols database. It has the following members:

char *p_name

This is the official name of the protocol.

char **p_aliases

These are alternate names for the protocol, specified as an array of strings. The last element of the array is a null pointer.

int p_proto

This is the protocol number (in host byte order); use this member as the protocol argument to socket.

You can use getprotobyname and getprotobynumber to search the pro-

tocols database for a specific protocol. The information is returned in a statically-allocated structure; you must copy the information if you need to save it across calls.

Function struct protoent * getprotobyname (const char

*name) The getprotobyname function returns information about the network protocol named name. If there is no such protocol, it returns a null pointer.

Function struct protoent * getprotobynumber (int

protocol) The getprotobynumber function returns information about the network protocol with number protocol. If there is no such protocol, it returns a null pointer.

You can also scan the whole protocols database one protocol at a time

by using setprotoent, getprotoent, and endprotoent. Be careful in using

these functions, because they are not reentrant.

Function void setprotoent (int stayopen)

This function opens the protocols database to begin scanning it. If the stayopen argument is nonzero, this sets a flag so that subse-

quent calls to getprotobyname or getprotobynumber will not close

the database (as they usually would). This makes for more efficiency if you call those functions several times, by avoiding reopening the database for each call.

Function struct protoent * getprotoent (void)

This function returns the next entry in the protocols database. It returns a null pointer if there are no more entries.

Function void endprotoent (void)

This function closes the protocols database.

11.5.7 Internet Socket Example

Here is an example showing how to create and name a socket in the Internet namespace. The newly created socket exists on the machine that the program is running on. Rather than finding and using the machine’s Internet address, this example specifies INADDR_ANY as the host address; the system replaces that with the machine’s actual address.

  #include <stdio.h>
  #include <stdlib.h>
  #include <sys/socket.h>
  #include <netinet/in.h>
  int
  make_socket (unsigned short int port)
  {
    int sock;
    struct sockaddr_in name;

/* Create the socket. */

    sock = socket (PF_INET, SOCK_STREAM, 0);
    if (sock < 0)
      {
        perror ("socket");
        exit (EXIT_FAILURE);
      }

/* Give the socket a name. */

    name.sin_family = AF_INET;
    name.sin_port = htons (port);
    name.sin_addr.s_addr = htonl (INADDR_ANY);
    if (bind (sock, (struct sockaddr *) &name, sizeof (name)) < 0)
      {
        perror ("bind");
        exit (EXIT_FAILURE);

         }
       return sock;
     }

Here is another example, showing how you can fill in a sockaddr_in structure, given a host name string and a port number:

     #include <stdio.h>
     #include <stdlib.h>
     #include <sys/socket.h>
     #include <netinet/in.h>
     #include <netdb.h>
     void
     init_sockaddr (struct sockaddr_in *name,
                    const char *hostname,
                    unsigned short int port)
     {
       struct hostent *hostinfo;
       name->sin_family = AF_INET;
       name->sin_port = htons (port);
       hostinfo = gethostbyname (hostname);
       if (hostinfo == NULL)
         {
           fprintf (stderr, "Unknown host %s.\n", hostname);
           exit (EXIT_FAILURE);
         }
       name->sin_addr = *(struct in_addr *) hostinfo->h_addr;
     }

11.6 Other Namespaces

Certain other namespaces and associated protocol families are supported but not documented yet because they are not often used. PF_NS refers to the Xerox Network Software protocols. PF_ISO stands for Open Systems Interconnect. PF_CCITT refers to protocols from CCITT. ‘socket.h’ defines these symbols and others naming protocols not actually implemented.

PF_IMPLINK is used for communicating between hosts and Internet Message Processors. For information on this, and on PF_ROUTE, an occasionallyused local area routing protocol, see the GNU Hurd Manual (to appear in the future).

11.7 Opening and Closing Sockets

This section describes the actual library functions for opening and closing sockets. The same functions work for all namespaces and connection styles.

11.7.1 Creating a Socket

The primitive for creating a socket is the socket function, declared in

‘sys/socket.h’.

Function int socket (int namespace, int style, int protocol)

This function creates a socket and specifies communication style style, which should be one of the socket styles listed in Section 11.2 [Communication Styles], page 222. The namespace argument specifies the namespace; it must be PF_FILE (see Section 11.4 [The File Namespace], page 227) or PF_INET (see Section 11.5 [The Internet Namespace], page 229). protocol designates the specific protocol (see Section 11.1 [Socket Concepts], page 221); zero is usually right for protocol. The return value from socket is the file descriptor for the new socket, or -1 in case of error. The following errno error conditions are defined for this function:

EPROTONOSUPPORT

The protocol or style is not supported by the namespace specified.

EMFILE The process already has too many file descriptors open.

ENFILE The system already has too many file descriptors open.

EACCESS The process does not have privilege to create a socket of the specified style or protocol.

ENOBUFS The system ran out of internal buffer space. The file descriptor returned by the socket function supports both read and write operations. But, like pipes, sockets do not support file positioning operations.

For examples of how to call the socket function, see Section 11.4 [The File Namespace], page 227, or Section 11.5.7 [Internet Socket Example], page 242.

11.7.2 Closing a Socket

When you are finished using a socket, you can simply close its file descriptor with close; see Section 8.1 [Opening and Closing Files], page 151. If there is still data waiting to be transmitted over the connection, normally close tries to complete this transmission. You can control this behavior using the SO_LINGER socket option to specify a timeout period; see Section 11.11 [Socket Options], page 266.

You can also shut down only reception or only transmission on a connection by calling shutdown, which is declared in ‘sys/socket.h’.

Function int shutdown (int socket, int how)

The shutdown function shuts down the connection of socket socket. The argument how specifies what action to perform:

0 Stop receiving data for this socket. If further data arrives, reject it.

1 Stop trying to transmit data from this socket. Discard any data waiting to be sent. Stop looking for acknowledgement of data already sent; don’t retransmit it if it is lost.

2 Stop both reception and transmission. The return value is 0 on success and -1 on failure. The following errno error conditions are defined for this function:

EBADF socket is not a valid file descriptor.

ENOTSOCK socket is not a socket.

ENOTCONN socket is not connected.

11.7.3 Socket Pairs

A socket pair consists of a pair of connected (but unnamed) sockets. It is very similar to a pipe and is used in much the same way. Socket pairs are created with the socketpair function, declared in ‘sys/socket.h’. A socket pair is much like a pipe; the main difference is that the socket pair is bidirectional, whereas the pipe has one input-only end and one output-only end (see Chapter 10 [Pipes and FIFOs], page 215).

Function int socketpair (int namespace, int style, int

protocol, int filedes[2]) This function creates a socket pair, returning the file descriptors in filedes[0] and filedes[1]. The socket pair is a full-duplex communications channel, so that both reading and writing may be performed at either end. The namespace, style, and protocol arguments are interpreted as for the socket function. style should be one of the communication styles listed in Section 11.2 [Communication Styles], page 222. The namespace argument specifies the namespace, which must be AF_FILE (see Section 11.4 [The File Namespace], page 227); protocol specifies the communications protocol, but zero is the only meaningful value.

If style specifies a connectionless communication style, then the two sockets you get are not connected, strictly speaking, but each of them knows the other as the default destination address, so they can send packets to each other. The socketpair function returns 0 on success and -1 on failure. The following errno error conditions are defined for this function:

EMFILE The process has too many file descriptors open.

EAFNOSUPPORT

The specified namespace is not supported.

EPROTONOSUPPORT

The specified protocol is not supported.

EOPNOTSUPP

The specified protocol does not support the creation of socket pairs.

11.8 Using Sockets with Connections

The most common communication styles involve making a connection to a particular other socket, and then exchanging data with that socket over and over. Making a connection is asymmetric; one side (the client) acts to request a connection, while the other side (the server) makes a socket and waits for the connection request.

• Section 11.8.1 [Making a Connection], page 246, describes what the client program must do to initiate a connection with a server.

• Section 11.8.2 [Listening for Connections], page 248, and Section 11.8.3 [Accepting Connections], page 248, describe what the server program must do to wait for and act upon connection requests from clients.

• Section 11.8.5 [Transferring Data], page 250, describes how data is transferred through the connected socket.

11.8.1 Making a Connection

In making a connection, the client makes a connection while the server waits for and accepts the connection. Here we discuss what the client program must do, using the connect function, which is declared in

‘sys/socket.h’.

Function int connect (int socket, struct sockaddr *addr,

      socklen_t length)

The connect function initiates a connection from the socket with file descriptor socket to the socket whose address is specified by the addr and length arguments. (This socket is typically on another machine, and it must be already set up as a server.) See Section 11.3 [Socket Addresses], page 223, for information about how these arguments are interpreted. Normally, connect waits until the server responds to the request before it returns. You can set nonblocking mode on the socket socket to make connect return immediately without waiting for the response. See Section 8.10 [File Status Flags], page 170, for information about nonblocking mode. The normal return value from connect is 0. If an error occurs, connect returns -1. The following errno error conditions are defined for this function:

EBADF The socket socket is not a valid file descriptor.

ENOTSOCK File descriptor socket is not a socket.

EADDRNOTAVAIL

The specified address is not available on the remote machine.

EAFNOSUPPORT

The namespace of the addr is not supported by this socket.

EISCONN The socket socket is already connected.

ETIMEDOUT

The attempt to establish the connection timed out.

ECONNREFUSED

The server has actively refused to establish the connection.

ENETUNREACH

The network of the given addr isn’t reachable from this host.

EADDRINUSE

The socket address of the given addr is already in use.

EINPROGRESS

The socket socket is non-blocking and the connection could not be established immediately. You can determine when the connection is completely established with select; see Section 8.6 [Waiting for Input or Output], page 162. Another connect call on the same socket, before the connection is completely established, will fail with

          EALREADY.

EALREADY The socket socket is non-blocking and already has a pending connection in progress (see EINPROGRESS above).

11.8.2 Listening for Connections

Now let us consider what the server process must do to accept connections on a socket. First it must use the listen function to enable connection requests on the socket, and then accept each incoming connection with a call to accept (see Section 11.8.3 [Accepting Connections], page 248). Once connection requests are enabled on a server socket, the select function reports when the socket has a connection ready to be accepted (see Section 8.6 [Waiting for Input or Output], page 162).

The listen function is not allowed for sockets using connectionless communication styles.

You can write a network server that does not even start running until a connection to it is requested. See Section 11.10.1 [“codeinetd Servers], page 265.

In the Internet namespace, there are no special protection mechanisms for controlling access to connect to a port; any process on any machine can make a connection to your server. If you want to restrict access to your server, make it examine the addresses associated with connection requests or implement some other handshaking or identification protocol.

In the File namespace, the ordinary file protection bits control who has access to connect to the socket.

Function int listen (int socket, unsigned int n)

The listen function enables the socket socket to accept connections, thus making it a server socket. The argument n specifies the length of the queue for pending connections. When the queue fills, new clients attempting to connect fail with ECONNREFUSED until the server calls accept to accept a connection from the queue. The listen function returns 0 on success and -1 on failure. The following errno error conditions are defined for this function:

EBADF The argument socket is not a valid file descriptor.

ENOTSOCK The argument socket is not a socket.

  EOPNOTSUPP

The socket socket does not support this operation.

11.8.3 Accepting Connections

When a server receives a connection request, it can complete the connection by accepting the request. Use the function accept to do this.

A socket that has been established as a server can accept connection requests from multiple clients. The server’s original socket does not become part of the connection; instead, accept makes a new socket which participates in the connection. accept returns the descriptor for this socket. The server’s original socket remains available for listening for further connection requests.

The number of pending connection requests on a server socket is finite. If connection requests arrive from clients faster than the server can act upon them, the queue can fill up and additional requests are refused with a ECONNREFUSED error. You can specify the maximum length of this queue as an argument to the listen function, although the system may also impose its own internal limit on the length of this queue.

Function int accept (int socket, struct sockaddr *addr,

socklen_t *length-ptr) This function is used to accept a connection request on the server socket socket. The accept function waits if there are no connections pending, unless the socket socket has nonblocking mode set. (You can use select to wait for a pending connection, with a nonblocking socket.) See Section 8.10 [File Status Flags], page 170, for information about nonblocking mode. The addr and length-ptr arguments are used to return information about the name of the client socket that initiated the connection. See Section 11.3 [Socket Addresses], page 223, for information about the format of the information. Accepting a connection does not make socket part of the connection. Instead, it creates a new socket which becomes connected. The normal return value of accept is the file descriptor for the new socket. After accept, the original socket socket remains open and unconnected, and continues listening until you close it. You can accept further connections with socket by calling accept again. If an error occurs, accept returns -1. The following errno error conditions are defined for this function:

EBADF The socket argument is not a valid file descriptor.

ENOTSOCK The descriptor socket argument is not a socket.

EOPNOTSUPP

The descriptor socket does not support this operation.

EWOULDBLOCK

socket has nonblocking mode set, and there are no pending connections immediately available.

The accept function is not allowed for sockets using connectionless communication styles.

11.8.4 Who is Connected to Me?

Function int getpeername (int socket, struct sockaddr

*addr, size_t *length-ptr) The getpeername function returns the address of the socket that socket is connected to; it stores the address in the memory space specified by addr and length-ptr. It stores the length of the address in *length-ptr. See Section 11.3 [Socket Addresses], page 223, for information about the format of the address. In some operating systems, getpeername works only for sockets in the Internet domain. The return value is 0 on success and -1 on error. The following errno error conditions are defined for this function:

EBADF The argument socket is not a valid file descriptor.

ENOTSOCK The descriptor socket is not a socket.

ENOTCONN The socket socket is not connected.

ENOBUFS There are not enough internal buffers available.

11.8.5 Transferring Data

Once a socket has been connected to a peer, you can use the ordinary read and write operations (see Section 8.2 [Input and Output Primitives], page 153) to transfer data. A socket is a two-way communications channel, so read and write operations can be performed at either end.

There are also some I/O modes that are specific to socket operations. In order to specify these modes, you must use the recv and send functions instead of the more generic read and write functions. The recv and send functions take an additional argument which you can use to specify various flags to control the special I/O modes. For example, you can specify the MSG_OOB flag to read or write out-of-band data, the MSG_PEEK flag to peek at input, or the MSG_DONTROUTE flag to control inclusion of routing information on output.

11.8.5.1 Sending Data

The send function is declared in the header file ‘sys/socket.h’. If your flags argument is zero, you can just as well use write instead of send; see Section 8.2 [Input and Output Primitives], page 153. If the socket was connected but the connection has broken, you get a SIGPIPE signal for any use of send or write (see Section 21.2.7 [Miscellaneous Signals], page 404).

Function int send (int socket, void *buffer, size_t size, int

flags) The send function is like write, but with the additional flags flags. The possible values of flags are described in Section 11.8.5.3 [Socket Data Options], page 252. This function returns the number of bytes transmitted, or -1 on failure. If the socket is nonblocking, then send (like write) can return after sending just part of the data. See Section 8.10 [File Status Flags], page 170, for information about nonblocking mode. Note, however, that a successful return value merely indicates that the message has been sent without error, not necessarily that it has been received without error. The following errno error conditions are defined for this function: EBADF The socket argument is not a valid file descriptor. EINTR The operation was interrupted by a signal before any data

was sent. See Section 21.5 [Primitives Interrupted by Signals], page 425.

ENOTSOCK The descriptor socket is not a socket. EMSGSIZE The socket type requires that the message be sent atomi-

cally, but the message is too large for this to be possible.

EWOULDBLOCK

Nonblocking mode has been set on the socket, and the write operation would block. (Normally send blocks until the operation can be completed.)

ENOBUFS There is not enough internal buffer space available. ENOTCONN You never connected this socket. EPIPE This socket was connected but the connection is now bro-

ken. In this case, send generates a SIGPIPE signal first; if that signal is ignored or blocked, or if its handler returns, then send fails with EPIPE.

11.8.5.2 Receiving Data

The recv function is declared in the header file ‘sys/socket.h’. If your flags argument is zero, you can just as well use read instead of recv; see Section 8.2 [Input and Output Primitives], page 153.

Function int recv (int socket, void *buffer, size_t size, int

flags) The recv function is like read, but with the additional flags flags. The possible values of flags are described In Section 11.8.5.3 [Socket Data Options], page 252.

If nonblocking mode is set for socket, and no data is available to be read, recv fails immediately rather than waiting. See Section 8.10 [File Status Flags], page 170, for information about nonblocking mode. This function returns the number of bytes received, or -1 on failure. The following errno error conditions are defined for this function:

EBADF The socket argument is not a valid file descriptor.

ENOTSOCK The descriptor socket is not a socket.

EWOULDBLOCK

Nonblocking mode has been set on the socket, and the read operation would block. (Normally, recv blocks until there is input available to be read.)

EINTR The operation was interrupted by a signal before any data was read. See Section 21.5 [Primitives Interrupted by Signals], page 425.

ENOTCONN You never connected this socket.

11.8.5.3 Socket Data Options

The flags argument to send and recv is a bit mask. You can bitwise-OR the values of the following macros together to obtain a value for this argument. All are defined in the header file ‘sys/socket.h’.

Macro int MSG OOB

Send or receive out-of-band data. See Section 11.8.8 [Out-of-Band Data], page 257.

Macro int MSG PEEK

Look at the data but don’t remove it from the input queue. This is only meaningful with input functions such as recv, not with send.

Macro int MSG DONTROUTE

Don’t include routing information in the message. This is only meaningful with output operations, and is usually only of interest for diagnostic or routing programs. We don’t try to explain it here.

11.8.6 Byte Stream Socket Example

Here is an example client program that makes a connection for a byte stream socket in the Internet namespace. It doesn’t do anything particularly interesting once it has connected to the server; it just sends a text string to the server and exits.

  #include <stdio.h>
  #include <errno.h>

Chapter 11: Sockets 253

#include <stdlib.h>
#include <unistd.h>
#include <sys/types.h>
#include <sys/socket.h>
#include <netinet/in.h>
#include <netdb.h>
#define PORT            5555
#define MESSAGE         "Yow!!! Are we having fun yet?!?"
#define SERVERHOST      "churchy.gnu.ai.mit.edu"
void
write_to_server (int filedes)
{
  int nbytes;
  nbytes = write (filedes, MESSAGE, strlen (MESSAGE) + 1);
  if (nbytes < 0)
    {
      perror ("write");
      exit (EXIT_FAILURE);
    }
}

int
main (void)
{
  extern void init_sockaddr (struct sockaddr_in *name,
                             const char *hostname,
                             unsigned short int port);
  int sock;
  struct sockaddr_in servername;

/* Create the socket. */

  sock = socket (PF_INET, SOCK_STREAM, 0);
  if (sock < 0)
    {
      perror ("socket (client)");
      exit (EXIT_FAILURE);
    }

/* Connect to the server. */

  init_sockaddr (&servername, SERVERHOST, PORT);

       if (0 > connect (sock,
                        (struct sockaddr *) &servername,
                        sizeof (servername)))
         {
           perror ("connect (client)");
           exit (EXIT_FAILURE);
         }

/* Send data to the server. */

       write_to_server (sock);
       close (sock);
       exit (EXIT_SUCCESS);
     }

11.8.7 Byte Stream Connection Server Example

The server end is much more complicated. Since we want to allow multiple clients to be connected to the server at the same time, it would be incorrect to wait for input from a single client by simply calling read or recv. Instead, the right thing to do is to use select (see Section 8.6 [Waiting for Input or Output], page 162) to wait for input on all of the open sockets. This also allows the server to deal with additional connection requests.

This particular server doesn’t do anything interesting once it has gotten a message from a client. It does close the socket for that client when it detects an end-of-file condition (resulting from the client shutting down its end of the connection).

This program uses make_socket and init_sockaddr to set up the socket address; see Section 11.5.7 [Internet Socket Example], page 242.

     #include <stdio.h>
     #include <errno.h>
     #include <stdlib.h>
     #include <unistd.h>
     #include <sys/types.h>
     #include <sys/socket.h>
     #include <netinet/in.h>
     #include <netdb.h>
     #define PORT    5555
     #define MAXMSG  512
     int
     read_from_client (int filedes)
     {
       char buffer[MAXMSG];

int nbytes;
nbytes = read (filedes, buffer, MAXMSG);
if (nbytes < 0)
  {

/* Read error. */

    perror ("read");
    exit (EXIT_FAILURE);
  }
else if (nbytes == 0)

/* End-of-file. */

  return -1;
else
  {

/* Data read. */

    fprintf (stderr, "Server: got message: `%s'\n", buffer);
    return 0;
  }
}
int
main (void)
{
extern int make_socket (unsigned short int port);
int sock;
fd_set active_fd_set, read_fd_set;
int i;
struct sockaddr_in clientname;
size_t size;

/* Create the socket and set it up to accept connections. */

sock = make_socket (PORT);
if (listen (sock, 1) < 0)
  {
    perror ("listen");
    exit (EXIT_FAILURE);
  }

/* Initialize the set of active sockets. */

FD_ZERO (&active_fd_set);
FD_SET (sock, &active_fd_set);

/* Block until input arrives on one or more active sockets. */

read_fd_set = active_fd_set;
if (select (FD_SETSIZE, &read_fd_set, NULL, NULL, NULL) < 0)
{
perror ("select");
exit (EXIT_FAILURE);
}

/* Service all the sockets with input pending. */

for (i = 0; i < FD_SETSIZE; ++i)
if (FD_ISSET (i, &read_fd_set))
{
if (i == sock)
{

/* Connection request on original socket. */

int new;
size = sizeof (clientname);
new = accept (sock,
              (struct sockaddr *) &clientname,
              &size);
if (new < 0)
  {
    perror ("accept");
    exit (EXIT_FAILURE);
  }
fprintf (stderr,
         "Server: connect from host %s, port %hd.\n",
         inet_ntoa (clientname.sin_addr),
         ntohs (clientname.sin_port));
FD_SET (new, &active_fd_set);
}
else
{

/* Data arriving on an already-connected socket. */

if (read_from_client (i) < 0)
  {
    close (i);
    FD_CLR (i, &active_fd_set);
  }
}
}
}
}

11.8.8 Out-of-Band Data

Streams with connections permit out-of-band data that is delivered with higher priority than ordinary data. Typically the reason for sending out-of-band data is to send notice of an exceptional condition. The way to send outof-band data is using send, specifying the flag MSG_OOB (see Section 11.8.5.1 [Sending Data], page 250).

Out-of-band data is received with higher priority because the receiving process need not read it in sequence; to read the next available out-of-band data, use recv with the MSG_OOB flag (see Section 11.8.5.2 [Receiving Data], page 251). Ordinary read operations do not read out-of-band data; they read only the ordinary data.

When a socket finds that out-of-band data is on its way, it sends a SIGURG signal to the owner process or process group of the socket. You can specify the owner using the F_SETOWN command to the fcntl function; see Section 8.12 [Interrupt-Driven Input], page 179. You must also establish a handler for this signal, as described in Chapter 21 [Signal Handling], page 393, in order to take appropriate action such as reading the out-of-band data.

Alternatively, you can test for pending out-of-band data, or wait until there is out-of-band data, using the select function; it can wait for an exceptional condition on the socket. See Section 8.6 [Waiting for Input or Output], page 162, for more information about select.

Notification of out-of-band data (whether with SIGURG or with select) indicates that out-of-band data is on the way; the data may not actually arrive until later. If you try to read the out-of-band data before it arrives, recv fails with an EWOULDBLOCK error.

Sending out-of-band data automatically places a “mark” in the stream of ordinary data, showing where in the sequence the out-of-band data “would have been”. This is useful when the meaning of out-of-band data is “cancel everything sent so far”. Here is how you can test, in the receiving process, whether any ordinary data was sent before the mark:

     success = ioctl (socket, SIOCATMARK, &result);

Here’s a function to discard any ordinary data preceding the out-of-band mark:

     int
     discard_until_mark (int socket)
     {
       while (1)
         {

/* This is not an arbitrary limit; any size will do. */

           char buffer[1024];
           int result, success;

/* If we have reached the mark, return. */

      success = ioctl (socket, SIOCATMARK, &result);
      if (success < 0)
        perror ("ioctl");
      if (result)
        return;

/* Otherwise, read a bunch of ordinary data and discard it. This is guaranteed not to read past the mark if it starts before the mark. */

      success = read (socket, buffer, sizeof buffer);
      if (success < 0)
        perror ("read");
    }
}

If you don’t want to discard the ordinary data preceding the mark, you may need to read some of it anyway, to make room in internal system buffers for the out-of-band data. If you try to read out-of-band data and get an EWOULDBLOCK error, try reading some ordinary data (saving it so that you can use it when you want it) and see if that makes room. Here is an example:

struct buffer
{
  char *buffer;
  int size;
  struct buffer *next;
};

/* Read the out-of-band data from SOCKET and return it as a ‘struct buffer’, which records the address of the data and its size.

It may be necessary to read some ordinary data

in order to make room for the out-of-band data.

If so, the ordinary data is saved as a chain of buffers found in the ‘next’ field of the value. */

struct buffer *
read_oob (int socket)
{
  struct buffer *tail = 0;
  struct buffer *list = 0;

/* This is an arbitrary limit.

Does anyone know how to do this without a limit? */

char *buffer = (char *) xmalloc (1024);
struct buffer *link;
int success;
int result;

/* Try again to read the out-of-band data. */

success = recv (socket, buffer, sizeof buffer, MSG_OOB);
if (success >= 0)
  {

/* We got it, so return it. */

    struct buffer *link
      = (struct buffer *) xmalloc (sizeof (struct buffer));
    link->buffer = buffer;
    link->size = success;
    link->next = list;
    return link;
  }

/* If we fail, see if we are at the mark. */

success = ioctl (socket, SIOCATMARK, &result);
if (success < 0)
  perror ("ioctl");
if (result)
  {

/* At the mark; skipping past more ordinary data cannot help.

So just wait a while. */

    sleep (1);
    continue;
  }

/* Otherwise, read a bunch of ordinary data and save it. This is guaranteed not to read past the mark if it starts before the mark. */

success = read (socket, buffer, sizeof buffer);
if (success < 0)
  perror ("read");

/* Save this data in the buffer list. */

{
  struct buffer *link
    = (struct buffer *) xmalloc (sizeof (struct buffer));
  link->buffer = buffer;

             link->size = success;

/* Add the new link to the end of the list. */

             if (tail)
               tail->next = link;
             else
               list = link;
             tail = link;
           }
         }
     }

11.9 Datagram Socket Operations

This section describes how to use communication styles that don’t use connections (styles SOCK_DGRAM and SOCK_RDM). Using these styles, you group data into packets and each packet is an independent communication. You specify the destination for each packet individually.

Datagram packets are like letters: you send each one independently, with its own destination address, and they may arrive in the wrong order or not at all.

The listen and accept functions are not allowed for sockets using connectionless communication styles.

11.9.1 Sending Datagrams

The normal way of sending data on a datagram socket is by using the sendto function, declared in ‘sys/socket.h’.

You can call connect on a datagram socket, but this only specifies a default destination for further data transmission on the socket. When a socket has a default destination, then you can use send (see Section 11.8.5.1 [Sending Data], page 250) or even write (see Section 8.2 [Input and Output Primitives], page 153) to send a packet there. You can cancel the default destination by calling connect using an address format of AF_UNSPEC in the addr argument. See Section 11.8.1 [Making a Connection], page 246, for more information about the connect function.

Function int sendto (int socket, void *buffer. size_t size, int

         flags, struct sockaddr *addr, socklen_t length)

The sendto function transmits the data in the buffer through the socket socket to the destination address specified by the addr and length arguments. The size argument specifies the number of bytes to be transmitted.

The flags are interpreted the same way as for send; see Section 11.8.5.3 [Socket Data Options], page 252. The return value and error conditions are also the same as for send, but you cannot rely on the system to detect errors and report them; the most common error is that the packet is lost or there is no one at the specified address to receive it, and the operating system on your machine usually does not know this. It is also possible for one call to sendto to report an error due to a problem related to a previous call.

11.9.2 Receiving Datagrams

The recvfrom function reads a packet from a datagram socket and also tells you where it was sent from. This function is declared in ‘sys/socket.h’.

Function int recvfrom (int socket, void *buffer, size_t size,

      int flags, struct sockaddr *addr, socklen_t *length-ptr)

The recvfrom function reads one packet from the socket socket into the buffer buffer. The size argument specifies the maximum number of bytes to be read. If the packet is longer than size bytes, then you get the first size bytes of the packet, and the rest of the packet is lost. There’s no way to read the rest of the packet. Thus, when you use a packet protocol, you must always know how long a packet to expect. The addr and length-ptr arguments are used to return the address where the packet came from. See Section 11.3 [Socket Addresses], page 223. For a socket in the file domain, the address information won’t be meaningful, since you can’t read the address of such a socket (see Section 11.4 [The File Namespace], page 227). You can specify a null pointer as the addr argument if you are not interested in this information. The flags are interpreted the same way as for recv (see Section 11.8.5.3 [Socket Data Options], page 252). The return value and error conditions are also the same as for recv.

You can use plain recv (see Section 11.8.5.2 [Receiving Data], page 251) instead of recvfrom if you know don’t need to find out who sent the packet (either because you know where it should come from or because you treat all possible senders alike). Even read can be used if you don’t want to specify flags (see Section 8.2 [Input and Output Primitives], page 153).

11.9.3 Datagram Socket Example

Here is a set of example programs that send messages over a datagram stream in the file namespace. Both the client and server programs use the make_named_socket function that was presented in Section 11.4 [The File Namespace], page 227, to create and name their sockets.

First, here is the server program. It sits in a loop waiting for messages to arrive, bouncing each message back to the sender. Obviously, this isn’t a particularly useful program, but it does show the general ideas involved.

#include <stdio.h>
#include <errno.h>
#include <stdlib.h>
#include <sys/socket.h>
#include <sys/un.h>
#define SERVER  "/tmp/serversocket"
#define MAXMSG  512
int
main (void)
{
  int sock;
  char message[MAXMSG];
  struct sockaddr_un name;
  size_t size;
  int nbytes;

/* Make the socket, then loop endlessly. */

  sock = make_named_socket (SERVER);
  while (1)
    {

/* Wait for a datagram. */

      size = sizeof (name);
      nbytes = recvfrom (sock, message, MAXMSG, 0,
                         (struct sockaddr *) & name, &size);
      if (nbytes < 0)
        {
          perror ("recfrom (server)");
          exit (EXIT_FAILURE);
        }

/* Give a diagnostic message. */

      fprintf (stderr, "Server: got message: %s\n", message);

/* Bounce the message back to the sender. */

      nbytes = sendto (sock, message, nbytes, 0,
                       (struct sockaddr *) & name, size);

      if (nbytes < 0)
        {
          perror ("sendto (server)");
          exit (EXIT_FAILURE);
        }
    }
}

11.9.4 Example of Reading Datagrams

Here is the client program corresponding to the server above. It sends a datagram to the server and then waits for a reply. Notice that

the socket for the client (as well as for the server) in this example has to be given a name. This is so that the server can direct a message back to the client. Since the socket has no associated connection state, the only way the server can do this is by referencing the name of the client.

#include <stdio.h>
#include <errno.h>
#include <unistd.h>
#include <stdlib.h>
#include <sys/socket.h>
#include <sys/un.h>
#define SERVER  "/tmp/serversocket"
#define CLIENT  "/tmp/mysocket"
#define MAXMSG  512
#define MESSAGE "Yow!!! Are we having fun yet?!?"
int
main (void)
{
  extern int make_named_socket (const char *name);
  int sock;
  char message[MAXMSG];
  struct sockaddr_un name;
  size_t size;
  int nbytes;

/* Make the socket. */

  sock = make_named_socket (CLIENT);

/* Initialize the server socket address. */

  name.sun_family = AF_UNIX;
  strcpy (name.sun_path, SERVER);

       size = strlen (name.sun_path) + sizeof (name.sun_family);

/* Send the datagram. */

       nbytes = sendto (sock, MESSAGE, strlen (MESSAGE) + 1, 0,
                        (struct sockaddr *) & name, size);
       if (nbytes < 0)
         {
           perror ("sendto (client)");
           exit (EXIT_FAILURE);
         }

/* Wait for a reply. */

       nbytes = recvfrom (sock, message, MAXMSG, 0, NULL, 0);
       if (nbytes < 0)
         {
           perror ("recfrom (client)");
           exit (EXIT_FAILURE);
         }

/* Print a diagnostic message. */

       fprintf (stderr, "Client: got message: %s\n", message);

/* Clean up. */

       remove (CLIENT);
       close (sock);
     }

Keep in mind that datagram socket communications are unreliable. In this example, the client program waits indefinitely if the message never reaches the server or if the server’s response never comes back. It’s up to the user running the program to kill it and restart it, if desired. A more automatic solution could be to use select (see Section 8.6 [Waiting for Input or Output], page 162) to establish a timeout period for the reply, and in case of timeout either resend the message or shut down the socket and exit.

11.10 The inetd Daemon

We’ve explained above how to write a server program that does its own listening. Such a server must already be running in order for anyone to connect to it.

Another way to provide service for an Internet port is to let the daemon program inetd do the listening. inetd is a program that runs all the time and waits (using select) for messages on a specified set of ports. When it receives a message, it accepts the connection (if the socket style calls for connections) and then forks a child process to run the corresponding server program. You specify the ports and their programs in the file

‘/etc/inetd.conf’.

11.10.1 inetd Servers

Writing a server program to be run by inetd is very simple. Each time someone requests a connection to the appropriate port, a new server process starts. The connection already exists at this time; the socket is available as the standard input descriptor and as the standard output descriptor (descriptors 0 and 1) in the server process. So the server program can begin reading and writing data right away. Often the program needs only the ordinary I/O facilities; in fact, a general-purpose filter program that knows nothing about sockets can work as a byte stream server run by inetd.

You can also use inetd for servers that use connectionless communication styles. For these servers, inetd does not try to accept a connection, since no connection is possible. It just starts the server program, which can read the incoming datagram packet from descriptor 0. The server program can handle one request and then exit, or you can choose to write it to keep reading more requests until no more arrive, and then exit. You must specify which of these two techniques the server uses, when you configure inetd.

11.10.2 Configuring inetd

The file ‘/etc/inetd.conf’ tells inetd which ports to listen to and what server programs to run for them. Normally each entry in the file is one line, but you can split it onto multiple lines provided all but the first line of the entry start with whitespace. Lines that start with ‘#’ are comments.

Here are two standard entries in ‘/etc/inetd.conf’:

     ftp stream tcp nowait root /libexec/ftpd ftpd
     talk dgram udp wait root /libexec/talkd talkd

An entry has this format:

service style protocol wait username program arguments

The service field says which service this program provides. It should be the name of a service defined in ‘/etc/services’. inetd uses service to decide which port to listen on for this entry.

The fields style and protocol specify the communication style and the protocol to use for the listening socket. The style should be the name of a communication style, converted to lower case and with ‘SOCK_’ deleted—for example, ‘stream’ or ‘dgram’. protocol should be one of the protocols listed in ‘/etc/protocols’. The typical protocol names are ‘tcp’ for byte stream connections and ‘udp’ for unreliable datagrams.

The wait field should be either ‘wait’ or ‘nowait’. Use ‘wait’ if style is a connectionless style and the server, once started, handles multiple requests, as many as come in. Use ‘nowait’ if inetd should start a new process for each message or request that comes in. If style uses connections, then wait must be ‘nowait’.

user is the user name that the server should run as. inetd runs as root, so it can set the user ID of its children arbitrarily. It’s best to avoid using ‘root’ for user if you can; but some servers, such as Telnet and FTP, read a username and password themselves. These servers need to be root initially so they can log in as commanded by the data coming over the network.

program together with arguments specifies the command to run to start the server. program should be an absolute file name specifying the executable file to run. arguments consists of any number of whitespace-separated words, which become the command-line arguments of program. The first word in arguments is argument zero, which should by convention be the program name itself (sans directories).

If you edit ‘/etc/inetd.conf’, you can tell inetd to reread the file and obey its new contents by sending the inetd process the SIGHUP signal. You’ll have to use ps to determine the process ID of the inetd process, as it is not fixed.

11.11 Socket Options

This section describes how to read or set various options that modify the behavior of sockets and their underlying communications protocols.

When you are manipulating a socket option, you must specify which level the option pertains to. This describes whether the option applies to the socket interface, or to a lower-level communications protocol interface.

11.11.1 Socket Option Functions

Here are the functions for examining and modifying socket options. They are declared in ‘sys/socket.h’.

Function int getsockopt (int socket, int level, int optname,

void *optval, socklen_t *optlen-ptr) The getsockopt function gets information about the value of option optname at level level for socket socket. The option value is stored in a buffer that optval points to. Before the call, you should supply in *optlen-ptr the size of this buffer; on return, it contains the number of bytes of information actually stored in the buffer. Most options interpret the optval buffer as a single int value. The actual return value of getsockopt is 0 on success and -1 on failure. The following errno error conditions are defined:

EBADF The socket argument is not a valid file descriptor.

ENOTSOCK The descriptor socket is not a socket.

ENOPROTOOPT

The optname doesn’t make sense for the given level.

Function int setsockopt (int socket, int level, int optname,

void *optval, socklen_t optlen)

This function is used to set the socket option optname at level level for socket socket. The value of the option is passed in the buffer optval, which has size optlen. The return value and error codes for setsockopt are the same as for

getsockopt.

11.11.2 Socket-Level Options

Constant int SOL SOCKET

Use this constant as the level argument to getsockopt or setsockopt to manipulate the socket-level options described in this section.

Here is a table of socket-level option names; all are defined in the header

file ‘sys/socket.h’.
SO_DEBUG

This option toggles recording of debugging information in the underlying protocol modules. The value has type int; a nonzero value means “yes”.

SO_REUSEADDR

This option controls whether bind (see Section 11.3.2 [Setting the Address of a Socket], page 225) should permit reuse of local addresses for this socket. If you enable this option, you can actually have two sockets with the same Internet port number; but the system won’t allow you to use the two identically-named sockets in a way that would confuse the Internet. The reason for this option is that some higher-level Internet protocols, including FTP, require you to keep reusing the same socket number. The value has type int; a nonzero value means “yes”.

SO_KEEPALIVE

This option controls whether the underlying protocol should periodically transmit messages on a connected socket. If the peer fails to respond to these messages, the connection is considered broken. The value has type int; a nonzero value means “yes”.

SO_DONTROUTE

This option controls whether outgoing messages bypass the normal message routing facilities. If set, messages are sent directly to the network interface instead. The value has type int; a nonzero value means “yes”.

SO_LINGER

This option specifies what should happen when the socket of a type that promises reliable delivery still has untransmitted messages when it is closed; see Section 11.7.2 [Closing a Socket], page 244. The value has type struct linger.

Data Type struct linger

This structure type has the following members:

  int l_onoff

This field is interpreted as a boolean. If nonzero, close blocks until the data is transmitted or the timeout period has expired.

  int l_linger

This specifies the timeout period, in seconds.

SO_BROADCAST

This option controls whether datagrams may be broadcast from the socket. The value has type int; a nonzero value means “yes”.

SO_OOBINLINE

If this option is set, out-of-band data received on the socket is placed in the normal input queue. This permits it to be read using read or recv without specifying the MSG_OOB flag. See Section 11.8.8 [Out-of-Band Data], page 257. The value has type int; a nonzero value means “yes”.

SO_SNDBUF

This option gets or sets the size of the output buffer. The value is a size_t, which is the size in bytes.

SO_RCVBUF

This option gets or sets the size of the input buffer. The value is a size_t, which is the size in bytes.

SO_STYLE

SO_TYPE This option can be used with getsockopt only. It is used to get the socket’s communication style. SO_TYPE is the historical name, and SO_STYLE is the preferred name in GNU. The value has type int and its value designates a communication style; see Section 11.2 [Communication Styles], page 222.

SO_ERROR

This option can be used with getsockopt only. It is used to reset the error status of the socket. The value is an int, which represents the previous error status.

11.12 Networks Database

Many systems come with a database that records a list of networks known to the system developer. This is usually kept either in the file ‘/etc/networks’ or in an equivalent from a name server. This data base is useful for routing programs such as route, but it is not useful for programs that simply communicate over the network. We provide functions to access this data base, which are declared in ‘netdb.h’.

Data Type struct netent

This data type is used to represent information about entries in the networks database. It has the following members:

char *n_name

This is the “official” name of the network.

char **n_aliases

These are alternative names for the network, represented as a vector of strings. A null pointer terminates the array.

int n_addrtype

This is the type of the network number; this is always equal to AF_INET for Internet networks.

unsigned long int n_net

This is the network number. Network numbers are returned in host byte order; see Section 11.5.5 [Byte Order Conversion], page 239.

Use the getnetbyname or getnetbyaddr functions to search the networks database for information about a specific network. The information is returned in a statically-allocated structure; you must copy the information if you need to save it.

Function struct netent * getnetbyname (const char *name)

The getnetbyname function returns information about the network named name. It returns a null pointer if there is no such network.

Function struct netent * getnetbyaddr (long net, int type)

The getnetbyaddr function returns information about the network of type type with number net. You should specify a value of AF_INET for the type argument for Internet networks. getnetbyaddr returns a null pointer if there is no such network.

You can also scan the networks database using setnetent, getnetent, and endnetent. Be careful in using these functions, because they are not reentrant.

Function void setnetent (int stayopen)

This function opens and rewinds the networks database. If the stayopen argument is nonzero, this sets a flag so that subsequent calls to getnetbyname or getnetbyaddr will not close the database (as they usually would). This makes for more efficiency if you call those functions several times, by avoiding reopening the database for each call.

Function struct netent * getnetent (void)

This function returns the next entry in the networks database. It returns a null pointer if there are no more entries.

Function void endnetent (void)

This function closes the networks database.

12 Low-Level Terminal Interface

This chapter describes functions that are specific to terminal devices. You can use these functions to do things like turn off input echoing; set serial line characteristics such as line speed and flow control; and change which characters are used for end-of-file, command-line editing, sending signals, and similar control functions.

Most of the functions in this chapter operate on file descriptors. See Chapter 8 [Low-Level Input/Output], page 151, for more information about what a file descriptor is and how to open a file descriptor for a terminal device.

12.1 Identifying Terminals

The functions described in this chapter only work on files that correspond to terminal devices. You can find out whether a file descriptor is associated with a terminal by using the isatty function.

Prototypes for both isatty and ttyname are declared in the header file

‘unistd.h’.

Function int isatty (int filedes)

This function returns 1 if filedes is a file descriptor associated with an open terminal device, and 0 otherwise.

If a file descriptor is associated with a terminal, you can get its associated file name using the ttyname function. See also the ctermid function, described in Section 24.7.1 [Identifying the Controlling Terminal], page 497.

Function char * ttyname (int filedes)

If the file descriptor filedes is associated with a terminal device, the ttyname function returns a pointer to a statically-allocated, nullterminated string containing the file name of the terminal file. The value is a null pointer if the file descriptor isn’t associated with a terminal, or the file name cannot be determined.

12.2 I/O Queues

Many of the remaining functions in this section refer to the input and output queues of a terminal device. These queues implement a form of buffering within the kernel independent of the buffering implemented by I/O streams (see Chapter 7 [Input/Output on Streams], page 93).

The terminal input queue is also sometimes referred to as its typeahead buffer. It holds the characters that have been received from the terminal but not yet read by any process.

The size of the terminal’s input queue is described by the MAX_INPUT and _POSIX_MAX_INPUT parameters; see Section 28.6 [Limits on File System Capacity], page 545. You are guaranteed a queue size of at least MAX_INPUT, but the queue might be larger, and might even dynamically change size. If input flow control is enabled by setting the IXOFF input mode bit (see Section 12.4.4 [Input Modes], page 277), the terminal driver transmits STOP and START characters to the terminal when necessary to prevent the queue from overflowing. Otherwise, input may be lost if it comes in too fast from the terminal. In canonical mode, all input stays in the queue until a newline character is received, so the terminal input queue can fill up when you type a very long line. See Section 12.3 [Two Styles of Input: Canonical or Not], page 272.

The terminal output queue is like the input queue, but for output; it contains characters that have been written by processes, but not yet transmitted to the terminal. If output flow control is enabled by setting the IXON input mode bit (see Section 12.4.4 [Input Modes], page 277), the terminal driver obeys STOP and STOP characters sent by the terminal to stop and restart transmission of output.

Clearing the terminal input queue means discarding any characters that have been received but not yet read. Similarly, clearing the terminal output queue means discarding any characters that have been written but not yet transmitted.

12.3 Two Styles of Input: Canonical or Not

POSIX systems support two basic modes of input: canonical and noncanonical.

In canonical input processing mode, terminal input is processed in lines terminated by newline ('\n'), EOF, or EOL characters. No input can be read until an entire line has been typed by the user, and the read function (see Section 8.2 [Input and Output Primitives], page 153) returns at most a single line of input, no matter how many bytes are requested.

In canonical input mode, the operating system provides input editing facilities: some characters are interpreted specially to perform editing operations within the current line of text, such as ERASE and KILL. See Section 12.4.9.1 [Characters for Input Editing], page 287.

The constants _POSIX_MAX_CANON and MAX_CANON parameterize the maximum number of bytes which may appear in a single line of canonical input. See Section 28.6 [Limits on File System Capacity], page 545. You are guaranteed a maximum line length of at least MAX_CANON bytes, but the maximum might be larger, and might even dynamically change size.

In noncanonical input processing mode, characters are not grouped into lines, and ERASE and KILL processing is not performed. The granularity with which bytes are read in noncanonical input mode is controlled by the MIN and TIME settings. See Section 12.4.10 [Noncanonical Input], page 291.

Most programs use canonical input mode, because this gives the user a way to edit input line by line. The usual reason to use noncanonical mode is when the program accepts single-character commands or provides its own editing facilities.

The choice of canonical or noncanonical input is controlled by the ICANON flag in the c_lflag member of struct termios. See Section 12.4.7 [Local Modes], page 282.

12.4 Terminal Modes

This section describes the various terminal attributes that control how input and output are done. The functions, data structures, and symbolic constants are all declared in the header file ‘termios.h’.

12.4.1 Terminal Mode Data Types

The entire collection of attributes of a terminal is stored in a structure of type struct termios. This structure is used with the functions tcgetattr and tcsetattr to read and set the attributes.

Data Type struct termios

Structure that records all the I/O attributes of a terminal. The structure includes at least the following members:

  tcflag_t c_iflag

A bit mask specifying flags for input modes; see Section 12.4.4 [Input Modes], page 277.

  tcflag_t c_oflag

A bit mask specifying flags for output modes; see Section 12.4.5 [Output Modes], page 279.

  tcflag_t c_cflag

A bit mask specifying flags for control modes; see Section 12.4.6 [Control Modes], page 280.

  tcflag_t c_lflag

A bit mask specifying flags for local modes; see Section 12.4.7 [Local Modes], page 282.

  cc_t c_cc[NCCS]

An array specifying which characters are associated with various control functions; see Section 12.4.9 [Special Characters], page 286.

The struct termios structure also contains members which encode input and output transmission speeds, but the representation is not specified. See Section 12.4.8 [Line Speed], page 284, for how to examine and store the speed values.

The following sections describe the details of the members of the struct termios structure.

Data Type tcflag t

This is an unsigned integer type used to represent the various bit masks for terminal flags.

Data Type cc t

This is an unsigned integer type used to represent characters associated with various terminal control functions.

Macro int NCCS

The value of this macro is the number of elements in the c_cc array.

12.4.2 Terminal Mode Functions

Function int tcgetattr (int filedes, struct termios *termios-p)

This function is used to examine the attributes of the terminal device with file descriptor filedes. The attributes are returned in the structure that termios-p points to. If successful, tcgetattr returns 0. A return value of -1 indicates an error. The following errno error conditions are defined for this function:

EBADF The filedes argument is not a valid file descriptor.

ENOTTY The filedes is not associated with a terminal.

Function int tcsetattr (int filedes, int when, const struct

termios *termios-p) This function sets the attributes of the terminal device with file descriptor filedes. The new attributes are taken from the structure that termios-p points to. The when argument specifies how to deal with input and output already queued. It can be one of the following values:

TCSANOW Make the change immediately.

TCSADRAIN

Make the change after waiting until all queued output has been written. You should usually use this option when changing parameters that affect output.

  TCSAFLUSH

This is like TCSADRAIN, but also discards any queued input.

TCSASOFT This is a flag bit that you can add to any of the above alternatives. Its meaning is to inhibit alteration of the state of the terminal hardware. It is a BSD extension; it is only supported on BSD systems and the GNU system. Using TCSASOFT is exactly the same as setting the CIGNORE bit in the c_cflag member of the structure termios-p points to. See Section 12.4.6 [Control Modes], page 280, for a description of CIGNORE.

If this function is called from a background process on its controlling terminal, normally all processes in the process group are sent a SIGTTOU signal, in the same way as if the process were trying to write to the terminal. The exception is if the calling process itself is ignoring or blocking SIGTTOU signals, in which case the operation is performed and no signal is sent. See Chapter 24 [Job Control], page 479. If successful, tcsetattr returns 0. A return value of -1 indicates an error. The following errno error conditions are defined for this function:

EBADF The filedes argument is not a valid file descriptor.

ENOTTY The filedes is not associated with a terminal.

EINVAL Either the value of the when argument is not valid, or there is something wrong with the data in the termios-p argument.

Although tcgetattr and tcsetattr specify the terminal device with a file descriptor, the attributes are those of the terminal device itself and not of the file descriptor. This means that the effects of changing terminal attributes are persistent; if another process opens the terminal file later on, it will see the changed attributes even though it doesn’t have anything to do with the open file descriptor you originally specified in changing the attributes.

Similarly, if a single process has multiple or duplicated file descriptors for the same terminal device, changing the terminal attributes affects input and output to all of these file descriptors. This means, for example, that you can’t open one file descriptor or stream to read from a terminal in the normal line-buffered, echoed mode; and simultaneously have another file descriptor for the same terminal that you use to read from it in single-character, nonechoed mode. Instead, you have to explicitly switch the terminal back and forth between the two modes.

12.4.3 Setting Terminal Modes Properly

When you set terminal modes, you should call tcgetattr first to get the current modes of the particular terminal device, modify only those modes that you are really interested in, and store the result with tcsetattr.

It’s a bad idea to simply initialize a struct termios structure to a chosen set of attributes and pass it directly to tcsetattr. Your program may be run years from now, on systems that support members not documented in this manual. The way to avoid setting these members to unreasonable values is to avoid changing them.

What’s more, different terminal devices may require different mode settings in order to function properly. So you should avoid blindly copying attributes from one terminal device to another.

When a member contains a collection of independent flags, as the c_ iflag, c_oflag and c_cflag members do, even setting the entire member is a bad idea, because particular operating systems have their own flags. Instead, you should start with the current value of the member and alter only the flags whose values matter in your program, leaving any other flags unchanged.

Here is an example of how to set one flag (ISTRIP) in the struct termios structure while properly preserving all the other data in the structure:

     int
     set_istrip (int desc, int value)
     {
       struct termios settings;
       int result;
       result = tcgetattr (desc, &settings);
       if (result < 0)
         {
           perror ("error in tcgetattr");
           return 0;
         }
       settings.c_iflag &= ~ISTRIP;
       if (value)
         settings.c_iflag |= ISTRIP;
       result = tcsetattr (desc, TCSANOW, &settings);
       if (result < 0)
         {
           perror ("error in tcgetattr");
           return;
        }
       return 1;
     }

12.4.4 Input Modes

This section describes the terminal attribute flags that control fairly lowlevel aspects of input processing: handling of parity errors, break signals, flow control, and ^{⟨}\underline{\overline{RET}}^{⟩} and ^{⟨}\underline{\overline{LFD}}^{⟩} characters.

All of these flags are bits in the c_iflag member of the struct termios structure. The member is an integer, and you change flags using the operators &, | and ^. Don’t try to specify the entire value for c_iflag—instead, change only specific flags and leave the rest untouched (see Section 12.4.3 [Setting Terminal Modes Properly], page 276).

Macro tcflag_t INPCK

If this bit is set, input parity checking is enabled. If it is not set, no checking at all is done for parity errors on input; the characters are simply passed through to the application. Parity checking on input processing is independent of whether parity detection and generation on the underlying terminal hardware is enabled; see Section 12.4.6 [Control Modes], page 280. For example, you could clear the INPCK input mode flag and set the PARENB control mode flag to ignore parity errors on input, but still generate parity on output. If this bit is set, what happens when a parity error is detected depends on whether the IGNPAR or PARMRK bits are set. If neither of these bits are set, a byte with a parity error is passed to the application as a '\0' character.

Macro tcflag_t IGNPAR

If this bit is set, any byte with a framing or parity error is ignored. This is only useful if INPCK is also set.

Macro tcflag_t PARMRK

If this bit is set, input bytes with parity or framing errors are marked when passed to the program. This bit is meaningful only when INPCK is set and IGNPAR is not set. The way erroneous bytes are marked is with two preceding bytes, 377 and 0. Thus, the program actually reads three bytes for one erroneous byte received from the terminal. If a valid byte has the value 0377, and ISTRIP (see below) is not set, the program might confuse it with the prefix that marks a parity error. So a valid byte 0377 is passed to the program as two bytes, 03770377, in this case.

Macro tcflag_t ISTRIP

If this bit is set, valid input bytes are stripped to seven bits; otherwise, all eight bits are available for programs to read.

Macro tcflag_t IGNBRK

If this bit is set, break conditions are ignored. A break condition is defined in the context of asynchronous serial data transmission as a series of zero-value bits longer than a single byte.

Macro tcflag_t BRKINT

If this bit is set and IGNBRK is not set, a break condition clears the terminal input and output queues and raises a SIGINT signal for the foreground process group associated with the terminal. If neither BRKINT nor IGNBRK are set, a break condition is passed to the application as a single '\0' character if PARMRK is not set, or otherwise as a three-character sequence '\377', '\0', '\0'.

Macro tcflag_t IGNCR

If this bit is set, carriage return characters ('\r') are discarded on input. Discarding carriage return may be useful on terminals that send both carriage return and linefeed when you type the ^{⟨}\underline{\overline{RET}}^{⟩} key.

Macro tcflag_t ICRNL

If this bit is set and IGNCR is not set, carriage return characters ('\r') received as input are passed to the application as newline characters

('\n').

Macro tcflag_t INLCR

If this bit is set, newline characters ('\n') received as input are passed to the application as carriage return characters ('\r').

Macro tcflag_t IXOFF

If this bit is set, start/stop control on input is enabled. In other words, the computer sends STOP and START characters as necessary to prevent input from coming in faster than programs are reading it. The idea is that the actual terminal hardware that is generating the input data responds to a STOP character by suspending transmission, and to a START character by resuming transmission. See Section 12.4.9.3 [Special Characters for Flow Control], page 290.

Macro tcflag_t IXON

If this bit is set, start/stop control on output is enabled. In other words, if the computer receives a STOP character, it suspends output until a START character is received. In this case, the STOP and START characters are never passed to the application program. If this bit is not set, then START and STOP can be read as ordinary characters. See Section 12.4.9.3 [Special Characters for Flow Control], page 290.

Macro tcflag_t IXANY

If this bit is set, any input character restarts output when output has been suspended with the STOP character. Otherwise, only the START character restarts output. This is a BSD extension; it exists only on BSD systems and the GNU system.

Macro tcflag_t IMAXBEL

If this bit is set, then filling up the terminal input buffer sends a BEL character (code 007) to the terminal to ring the bell. This is a BSD extension.

12.4.5 Output Modes

This section describes the terminal flags and fields that control how output characters are translated and padded for display. All of these are contained in the c_oflag member of the struct termios structure.

The c_oflag member itself is an integer, and you change the flags and fields using the operators &, |, and ^. Don’t try to specify the entire value for c_oflag—instead, change only specific flags and leave the rest untouched (see Section 12.4.3 [Setting Terminal Modes Properly], page 276).

Macro tcflag_t OPOST

If this bit is set, output data is processed in some unspecified way so that it is displayed appropriately on the terminal device. This typically includes mapping newline characters ('\n') onto carriage return and linefeed pairs. If this bit isn’t set, the characters are transmitted as-is.

The following three bits are BSD features, and they exist only BSD systems and the GNU system. They are effective only if OPOST is set.

Macro tcflag_t ONLCR

If this bit is set, convert the newline character on output into a pair of characters, carriage return followed by linefeed.

Macro tcflag_t OXTABS

If this bit is set, convert tab characters on output into the appropriate number of spaces to emulate a tab stop every eight columns.

Macro tcflag_t ONOEOT

If this bit is set, discard C-d characters (code 004) on output. These characters cause many dial-up terminals to disconnect.

12.4.6 Control Modes

This section describes the terminal flags and fields that control parameters usually associated with asynchronous serial data transmission. These flags may not make sense for other kinds of terminal ports (such as a network connection pseudo-terminal). All of these are contained in the c_cflag member of the struct termios structure.

The c_cflag member itself is an integer, and you change the flags and fields using the operators &, |, and ^. Don’t try to specify the entire value for c_cflag—instead, change only specific flags and leave the rest untouched (see Section 12.4.3 [Setting Terminal Modes Properly], page 276).

Macro tcflag_t CLOCAL

If this bit is set, it indicates that the terminal is connected “locally” and that the modem status lines (such as carrier detect) should be ignored. On many systems if this bit is not set and you call open without the O_NONBLOCK flag set, open blocks until a modem connection is established. If this bit is not set and a modem disconnect is detected, a SIGHUP signal is sent to the controlling process group for the terminal (if it has one). Normally, this causes the process to exit; see Chapter 21 [Signal Handling], page 393. Reading from the terminal after a disconnect causes an end-of-file condition, and writing causes an EIO error to be returned. The terminal device must be closed and reopened to clear the condition.

Macro tcflag_t HUPCL

If this bit is set, a modem disconnect is generated when all processes that have the terminal device open have either closed the file or exited.

Macro tcflag_t CREAD

If this bit is set, input can be read from the terminal. Otherwise, input is discarded when it arrives.

Macro tcflag_t CSTOPB

If this bit is set, two stop bits are used. Otherwise, only one stop bit is used.

Macro tcflag_t PARENB

If this bit is set, generation and detection of a parity bit are enabled. See Section 12.4.4 [Input Modes], page 277, for information on how input parity errors are handled. If this bit is not set, no parity bit is added to output characters, and input characters are not checked for correct parity.

Macro tcflag_t PARODD

This bit is only useful if PARENB is set. If PARODD is set, odd parity is used, otherwise even parity is used.

The control mode flags also includes a field for the number of bits per character. You can use the CSIZE macro as a mask to extract the value, like

this: settings.c_cflag & CSIZE.

Macro tcflag_t CSIZE

This is a mask for the number of bits per character.

Macro tcflag_t CS5

This specifies five bits per byte.

Macro tcflag_t CS6

This specifies six bits per byte.

Macro tcflag_t CS7

This specifies seven bits per byte.

Macro tcflag_t CS8

This specifies eight bits per byte.

The following four bits are BSD extensions; this exist only on BSD systems and the GNU system.

Macro tcflag_t CCTS OFLOW

If this bit is set, enable flow control of output based on the CTS wire (RS232 protocol).

Macro tcflag_t CRTS IFLOW

If this bit is set, enable flow control of input based on the RTS wire (RS232 protocol).

Macro tcflag_t MDMBUF

If this bit is set, enable carrier-based flow control of output.

Macro tcflag_t CIGNORE

If this bit is set, it says to ignore the control modes and line speed values entirely. This is only meaningful in a call to tcsetattr. The c_cflag member and the line speed values returned by

cfgetispeed and cfgetospeed will be unaffected by the call. CIGNORE

is useful if you want to set all the software modes in the other members, but leave the hardware details in c_cflag unchanged. (This is

how the TCSASOFT flag to tcsettattr works.)

This bit is never set in the structure filled in by tcgetattr.

12.4.7 Local Modes

This section describes the flags for the c_lflag member of the struct termios structure. These flags generally control higher-level aspects of input processing than the input modes flags described in Section 12.4.4 [Input Modes], page 277, such as echoing, signals, and the choice of canonical or noncanonical input.

The c_lflag member itself is an integer, and you change the flags and fields using the operators &, |, and ^. Don’t try to specify the entire value for c_lflag—instead, change only specific flags and leave the rest untouched (see Section 12.4.3 [Setting Terminal Modes Properly], page 276).

Macro tcflag_t ICANON

This bit, if set, enables canonical input processing mode. Otherwise, input is processed in noncanonical mode. See Section 12.3 [Two Styles of Input: Canonical or Not], page 272.

Macro tcflag_t ECHO

If this bit is set, echoing of input characters back to the terminal is enabled.

Macro tcflag_t ECHOE

If this bit is set, echoing indicates erasure of input with the ERASE character by erasing the last character in the current line from the screen. Otherwise, the character erased is re-echoed to show what has happened (suitable for a printing terminal). This bit only controls the display behavior; the ICANON bit by itself controls actual recognition of the ERASE character and erasure of input, without which ECHOE is simply irrelevant.

Macro tcflag_t ECHOPRT

This bit is like ECHOE, enables display of the ERASE character in a way that is geared to a hardcopy terminal. When you type the ERASE character, a ‘\’ character is printed followed by the first character erased. Typing the ERASE character again just prints the next character erased. Then, the next time you type a normal character, a ‘/’ character is printed before the character echoes. This is a BSD extension, and exists only in BSD systems and the GNU system.

Macro tcflag_t ECHOK

This bit enables special display of the KILL character by moving to a new line after echoing the KILL character normally. The behavior of ECHOKE (below) is nicer to look at. If this bit is not set, the KILL character echoes just as it would if it were not the KILL character. Then it is up to the user to remember that the KILL character has erased the preceding input; there is no indication of this on the screen. This bit only controls the display behavior; the ICANON bit by itself controls actual recognition of the KILL character and erasure of input, without which ECHOK is simply irrelevant.

Macro tcflag_t ECHOKE

This bit is similar to ECHOK. It enables special display of the KILL character by erasing on the screen the entire line that has been killed. This is a BSD extension, and exists only in BSD systems and the GNU system.

Macro tcflag_t ECHONL

If this bit is set and the ICANON bit is also set, then the newline ('\n') character is echoed even if the ECHO bit is not set.

Macro tcflag_t ECHOCTL

If this bit is set and the ECHO bit is also set, echo control characters with ‘^’ followed by the corresponding text character. Thus, control-A echoes as ‘^A’. This is usually the preferred mode for interactive input, because echoing a control character back to the terminal could have some undesired effect on the terminal. This is a BSD extension, and exists only in BSD systems and the GNU system.

Macro tcflag_t ISIG

This bit controls whether the INTR, QUIT, and SUSP characters are recognized. The functions associated with these characters are performed if and only if this bit is set. Being in canonical or noncanonical input mode has no affect on the interpretation of these characters. You should use caution when disabling recognition of these characters. Programs that cannot be interrupted interactively are very userunfriendly. If you clear this bit, your program should provide some alternate interface that allows the user to interactively send the signals associated with these characters, or to escape from the program.

See Section 12.4.9.2 [Characters that Cause Signals], page 288.

Macro tcflag_t IEXTEN

POSIX.1 gives IEXTEN implementation-defined meaning, so you cannot rely on this interpretation on all systems. On BSD systems and the GNU system, it enables the LNEXT and DISCARD characters. See Section 12.4.9.4 [Other Special Characters], page 290.

Macro tcflag_t NOFLSH

Normally, the INTR, QUIT, and SUSP characters cause input and output queues for the terminal to be cleared. If this bit is set, the queues are not cleared.

Macro tcflag_t TOSTOP

If this bit is set and the system supports job control, then SIGTTOU signals are generated by background processes that attempt to write to the terminal. See Section 24.4 [Access to the Controlling Terminal], page 481.

The following bits are BSD extensions; they exist only in BSD systems and the GNU system.

Macro tcflag_t ALTWERASE

This bit determines how far the WERASE character should erase. The WERASE character erases back to the beginning of a word; the question is, where do words begin? If this bit is clear, then the beginning of a word is a nonwhitespace character following a whitespace character. If the bit is set, then the beginning of a word is an alphanumeric character or underscore following a character which is none of those. See Section 12.4.9.1 [Characters for Input Editing], page 287, for more information about the WERASE character.

Macro tcflag_t FLUSHO

This is the bit that toggles when the user types the DISCARD character. While this bit is set, all output is discarded. See Section 12.4.9.4 [Other Special Characters], page 290.

Macro tcflag_t NOKERNINFO

Setting this bit disables handling of the STATUS character. See Section 12.4.9.4 [Other Special Characters], page 290.

Macro tcflag_t PENDIN

If this bit is set, it indicates that there is a line of input that needs to be reprinted. Typing the REPRINT character sets this bit; the bit remains set until reprinting is finished. See Section 12.4.9.1 [Characters for Input Editing], page 287.

12.4.8 Line Speed

The terminal line speed tells the computer how fast to read and write data on the terminal.

If the terminal is connected to a real serial line, the terminal speed you specify actually controls the line—if it doesn’t match the terminal’s own idea of the speed, communication does not work. Real serial ports accept only certain standard speeds. Also, particular hardware may not support even all the standard speeds. Specifying a speed of zero hangs up a dialup connection and turns off modem control signals.

If the terminal is not a real serial line (for example, if it is a network connection), then the line speed won’t really affect data transmission speed, but some programs will use it to determine the amount of padding needed. It’s best to specify a line speed value that matches the actual speed of the actual terminal, but you can safely experiment with different values to vary the amount of padding.

There are actually two line speeds for each terminal, one for input and one for output. You can set them independently, but most often terminals use the same speed for both directions.

The speed values are stored in the struct termios structure, but don’t try to access them in the struct termios structure directly. Instead, you should use the following functions to read and store them:

Function speed_t cfgetospeed (const struct termios

*termios-p) This function returns the output line speed stored in the structure *termios-p.

Function speed_t cfgetispeed (const struct termios

*termios-p) This function returns the input line speed stored in the structure *termios-p.

Function int cfsetospeed (struct termios *termios-p, speed_t

speed) This function stores speed in *termios-p as the output speed. The normal return value is 0; a value of -1 indicates an error. If speed is not a speed, cfsetospeed returns -1.

Function int cfsetispeed (struct termios *termios-p, speed_t

speed) This function stores speed in *termios-p as the input speed. The normal return value is 0; a value of -1 indicates an error. If speed is not a speed, cfsetospeed returns -1.

Function int cfsetspeed (struct termios *termios-p, speed_t

speed) This function stores speed in *termios-p as both the input and output speeds. The normal return value is 0; a value of -1 indicates an error. If speed is not a speed, cfsetspeed returns -1. This function is an extension in 4.4 BSD.

Data Type speed t

The speed_t type is an unsigned integer data type used to represent line speeds.

The functions cfsetospeed and cfsetispeed report errors only for speed values that the system simply cannot handle. If you specify a speed value that is basically acceptable, then those functions will succeed. But they do not check that a particular hardware device can actually support the specified speeds—in fact, they don’t know which device you plan to set the speed for. If you use tcsetattr to set the speed of a particular device to a value that it cannot handle, tcsetattr returns -1.

Portability note: In the GNU library, the functions above accept speeds measured in bits per second as input, and return speed values measured in bits per second. Other libraries require speeds to be indicated by special codes. For POSIX.1 portability, you must use one of the following symbols to represent the speed; their precise numeric values are system-dependent, but each name has a fixed meaning: B110 stands for 110 bps, B300 for 300 bps, and so on. There is no portable way to represent any speed but these, but these are the only speeds that typical serial lines can support.

     B0  B50  B75  B110  B134  B150  B200
     B300  B600  B1200  B1800  B2400  B4800
     B9600  B19200  B38400

BSD defines two additional speed symbols as aliases: EXTA is an alias for B19200 and EXTB is an alias for B38400. These aliases are obsolete.

12.4.9 Special Characters

In canonical input, the terminal driver recognizes a number of special characters which perform various control functions. These include the ERASE character (usually ^{⟨}\underline{\overline{DEL}}^{⟩}) for editing input, and other editing characters. The INTR character (normally C-c) for sending a SIGINT signal, and other signal-raising characters, may be available in either canonical or noncanonical input mode. All these characters are described in this section.

The particular characters used are specified in the c_cc member of the struct termios structure. This member is an array; each element specifies the character for a particular role. Each element has a symbolic constant that stands for the index of that element—for example, INTR is the index of the element that specifies the INTR character, so storing '=' in termios.c_ cc[INTR] specifies ‘=’ as the INTR character.

On some systems, you can disable a particular special character function by specifying the value _POSIX_VDISABLE for that role. This value is unequal to any possible character code. See Section 28.7 [Optional Features in File Support], page 546, for more information about how to tell whether the operating system you are using supports _POSIX_VDISABLE.

12.4.9.1 Characters for Input Editing

These special characters are active only in canonical input mode. See Section 12.3 [Two Styles of Input: Canonical or Not], page 272.

Macro int VEOF

This is the subscript for the EOF character in the special control character array. termios.c_cc[VEOF] holds the character itself. The EOF character is recognized only in canonical input mode. It acts as a line terminator in the same way as a newline character, but if the EOF character is typed at the beginning of a line it causes read to return a byte count of zero, indicating end-of-file. The EOF character itself is discarded. Usually, the EOF character is C-d.

Macro int VEOL

This is the subscript for the EOL character in the special control character array. termios.c_cc[VEOL] holds the character itself. The EOL character is recognized only in canonical input mode. It acts as a line terminator, just like a newline character. The EOL character is not discarded; it is read as the last character in the input line. You don’t need to use the EOL character to make ^{⟨}\underline{\overline{RET}}^{⟩} end a line. Just set the ICRNL flag. In fact, this is the default state of affairs.

Macro int VEOL2

This is the subscript for the EOL2 character in the special control character array. termios.c_cc[VEOL2] holds the character itself. The EOL2 character works just like the EOL character (see above), but it can be a different character. Thus, you can specify two characters to terminate an input line, by setting EOL to one of them and EOL2 to the other. The EOL2 character is a BSD extension; it exists only on BSD systems and the GNU system.

Macro int VERASE

This is the subscript for the ERASE character in the special control character array. termios.c_cc[VERASE] holds the character itself. The ERASE character is recognized only in canonical input mode. When the user types the erase character, the previous character typed is discarded. (If the terminal generates multibyte character sequences, this may cause more than one byte of input to be discarded.) This cannot be used to erase past the beginning of the current line of text. The ERASE character itself is discarded. Usually, the ERASE character is ^{⟨}\underline{\overline{DEL}}^{⟩}.

Macro int VWERASE

This is the subscript for the WERASE character in the special control character array. termios.c_cc[VWERASE] holds the character itself. The WERASE character is recognized only in canonical mode. It erases an entire word of prior input, and any whitespace after it; whitespace characters before the word are not erased. The definition of a “word” depends on the setting of the ALTWERASE mode; see Section 12.4.7 [Local Modes], page 282. If the ALTWERASE mode is not set, a word is defined as a sequence of any characters except space or tab. If the ALTWERASE mode is set, a word is defined as a sequence of characters containing only letters, numbers, and underscores, optionally followed by one character that is not a letter, number, or underscore. The WERASE character is usually C-w. This is a BSD extension.

Macro int VKILL

This is the subscript for the KILL character in the special control character array. termios.c_cc[VKILL] holds the character itself. The KILL character is recognized only in canonical input mode. When the user types the kill character, the entire contents of the current line of input are discarded. The kill character itself is discarded too. The KILL character is usually C-u.

Macro int VREPRINT

This is the subscript for the REPRINT character in the special control character array. termios.c_cc[VREPRINT] holds the character itself. The REPRINT character is recognized only in canonical mode. It reprints the current input line. If some asynchronous output has come while you are typing, this lets you see the line you are typing clearly again. The REPRINT character is usually C-r. This is a BSD extension.

12.4.9.2 Characters that Cause Signals

These special characters may be active in either canonical or noncanonical input mode, but only when the ISIG flag is set (see Section 12.4.7 [Local Modes], page 282).

Macro int VINTR

This is the subscript for the INTR character in the special control character array. termios.c_cc[VINTR] holds the character itself.

The INTR (interrupt) character raises a SIGINT signal for all processes in the foreground job associated with the terminal. The INTR character itself is then discarded. See Chapter 21 [Signal Handling], page 393, for more information about signals. Typically, the INTR character is C-c.

Macro int VQUIT

This is the subscript for the QUIT character in the special control character array. termios.c_cc[VQUIT] holds the character itself. The QUIT character raises a SIGQUIT signal for all processes in the foreground job associated with the terminal. The QUIT character itself is then discarded. See Chapter 21 [Signal Handling], page 393, for more information about signals. Typically, the QUIT character is C-\.

Macro int VSUSP

This is the subscript for the SUSP character in the special control character array. termios.c_cc[VSUSP] holds the character itself. The SUSP (suspend) character is recognized only if the implementation supports job control (see Chapter 24 [Job Control], page 479). It causes a SIGTSTP signal to be sent to all processes in the foreground job associated with the terminal. The SUSP character itself is then discarded. See Chapter 21 [Signal Handling], page 393, for more information about signals. Typically, the SUSP character is C-z.

Few applications disable the normal interpretation of the SUSP character. If your program does this, it should provide some other mechanism for the user to stop the job. When the user invokes this mechanism, the program should send a SIGTSTP signal to the process group of the process, not just to the process itself. See Section 21.6.2 [Signaling Another Process], page 428.

Macro int VDSUSP

This is the subscript for the DSUSP character in the special control character array. termios.c_cc[VDSUSP] holds the character itself. The DSUSP (suspend) character is recognized only if the implementation supports job control (see Chapter 24 [Job Control], page 479). It sends a SIGTSTP signal, like the SUSP character, but not right away— only when the program tries to read it as input. Not all systems with job control support DSUSP; only BSD-compatible systems (including the GNU system). See Chapter 21 [Signal Handling], page 393, for more information about signals. Typically, the DSUSP character is C-y.

12.4.9.3 Special Characters for Flow Control

These special characters may be active in either canonical or noncanonical input mode, but their use is controlled by the flags IXON and IXOFF (see Section 12.4.4 [Input Modes], page 277).

Macro int VSTART

This is the subscript for the START character in the special control character array. termios.c_cc[VSTART] holds the character itself. The START character is used to support the IXON and IXOFF input modes. If IXON is set, receiving a START character resumes suspended output; the START character itself is discarded. If IXANY is set, receiving any character at all resumes suspended output; the resuming character is not discarded unless it is the START character. IXOFF is set, the system may also transmit START characters to the terminal. The usual value for the START character is C-q. You may not be able to change this value—the hardware may insist on using C-q regardless of what you specify.

Macro int VSTOP

This is the subscript for the STOP character in the special control character array. termios.c_cc[VSTOP] holds the character itself. The STOP character is used to support the IXON and IXOFF input modes. If IXON is set, receiving a STOP character causes output to be suspended; the STOP character itself is discarded. If IXOFF is set, the system may also transmit STOP characters to the terminal, to prevent the input queue from overflowing. The usual value for the STOP character is C-s. You may not be able to change this value—the hardware may insist on using C-s regardless of what you specify.

12.4.9.4 Other Special Characters

These special characters exist only in BSD systems and the GNU system.

Macro int VLNEXT

This is the subscript for the LNEXT character in the special control character array. termios.c_cc[VLNEXT] holds the character itself. The LNEXT character is recognized only when IEXTEN is set, but in both canonical and noncanonical mode. It disables any special significance of the next character the user types. Even if the character would normally perform some editing function or generate a signal, it is read as a plain character. This is the analogue of the C-q command in Emacs. “LNEXT” stands for “literal next.” The LNEXT character is usually C-v.

Macro int VDISCARD

This is the subscript for the DISCARD character in the special control character array. termios.c_cc[VDISCARD] holds the character itself. The DISCARD character is recognized only when IEXTEN is set, but in both canonical and noncanonical mode. Its effect is to toggle the discard-output flag. When this flag is set, all program output is discarded. Setting the flag also discards all output currently in the output buffer. Typing any other character resets the flag.

Macro int VSTATUS

This is the subscript for the STATUS character in the special control character array. termios.c_cc[VSTATUS] holds the character itself. The STATUS character’s effect is to print out a status message about how the current process is running. The STATUS character is recognized only in canonical mode, and only if NOKERNINFO is not set.

12.4.10 Noncanonical Input

In noncanonical input mode, the special editing characters such as ERASE and KILL are ignored. The system facilities for the user to edit input are disabled in noncanonical mode, so that all input characters (unless they are special for signal or flow-control purposes) are passed to the application program exactly as typed. It is up to the application program to give the user ways to edit the input, if appropriate.

Noncanonical mode offers special parameters called MIN and TIME for controlling whether and how long to wait for input to be available. You can even use them to avoid ever waiting—to return immediately with whatever input is available, or with no input.

The MIN and TIME are stored in elements of the c_cc array, which is a member of the struct termios structure. Each element of this array has a particular role, and each element has a symbolic constant that stands for the index of that element. VMIN and VMAX are the names for the indices in the array of the MIN and TIME slots.

Macro int VMIN

This is the subscript for the MIN slot in the c_cc array. Thus, termios.c_cc[VMIN] is the value itself. The MIN slot is only meaningful in noncanonical input mode; it specifies the minimum number of bytes that must be available in the input queue in order for read to return.

Macro int VTIME

This is the subscript for the TIME slot in the c_cc array. Thus, termios.c_cc[VTIME] is the value itself.

The TIME slot is only meaningful in noncanonical input mode; it specifies how long to wait for input before returning, in units of 0.1 seconds.

The MIN and TIME values interact to determine the criterion for when read should return; their precise meanings depend on which of them are nonzero. There are four possible cases:

• Both TIME and MIN are nonzero.

In this case, TIME specifies how long to wait after each input character to see if more input arrives. After the first character received, read keeps waiting until either MIN bytes have arrived in all, or TIME elapses with no further input. read always blocks until the first character arrives, even if TIME elapses first. read can return more than MIN characters if more than MIN happen to be in the queue.

• Both MIN and TIME are zero.

In this case, read always returns immediately with as many characters as are available in the queue, up to the number requested. If no input is immediately available, read returns a value of zero.

• MIN is zero but TIME has a nonzero value. In this case, read waits for time TIME for input to become available; the availability of a single byte is enough to satisfy the read request and cause read to return. When it returns, it returns as many characters as are available, up to the number requested. If no input is available before the timer expires, read returns a value of zero.

• TIME is zero but MIN has a nonzero value. In this case, read waits until at least MIN bytes are available in the queue. At that time, read returns as many characters as are available, up to the number requested. read can return more than MIN characters if more than MIN happen to be in the queue. What happens if MIN is 50 and you ask to read just 10 bytes? Normally,

read waits until there are 50 bytes in the buffer (or, more generally, the wait condition described above is satisfied), and then reads 10 of them, leaving the other 40 buffered in the operating system for a subsequent call to read.

Portability note: On some systems, the MIN and TIME slots are actually the same as the EOF and EOL slots. This causes no serious problem because the MIN and TIME slots are used only in noncanonical input and the EOF and EOL slots are used only in canonical input, but it isn’t very clean. The GNU library allocates separate slots for these uses.

Function int cfmakeraw (struct termios *termios-p)

This function provides an easy way to set up *termios-p for what has traditionally been called “raw mode” in BSD. This uses noncanonical input, and turns off most processing to give an unmodified channel to the terminal. It does exactly this:

       termios-p->c_iflag &= ~(IGNBRK|BRKINT|PARMRK|ISTRIP
                                     |INLCR|IGNCR|ICRNL|IXON);
       termios-p->c_oflag &= ~OPOST;
       termios-p->c_lflag &= ~(ECHO|ECHONL|ICANON|ISIG|IEXTEN);
       termios-p->c_cflag &= ~(CSIZE|PARENB);
       termios-p->c_cflag |= CS8;

12.5 Line Control Functions

These functions perform miscellaneous control actions on terminal devices. As regards terminal access, they are treated like doing output: if any of these functions is used by a background process on its controlling terminal, normally all processes in the process group are sent a SIGTTOU signal. The exception is if the calling process itself is ignoring or blocking SIGTTOU signals, in which case the operation is performed and no signal is sent. See Chapter 24 [Job Control], page 479.

Function int tcsendbreak (int filedes, int duration)

This function generates a break condition by transmitting a stream of zero bits on the terminal associated with the file descriptor filedes. The duration of the break is controlled by the duration argument. If zero, the duration is between 0.25 and 0.5 seconds. The meaning of a nonzero value depends on the operating system. This function does nothing if the terminal is not an asynchronous serial data port. The return value is normally zero. In the event of an error, a value of -1 is returned. The following errno error conditions are defined for this function:

EBADF The filedes is not a valid file descriptor.

ENOTTY The filedes is not associated with a terminal device.

Function int tcdrain (int filedes)

The tcdrain function waits until all queued output to the terminal filedes has been transmitted. The return value is normally zero. In the event of an error, a value of -1 is returned. The following errno error conditions are defined for this function:

EBADF The filedes is not a valid file descriptor.

ENOTTY The filedes is not associated with a terminal device.

EINTR The operation was interrupted by delivery of a signal. See Section 21.5 [Primitives Interrupted by Signals], page 425.

Function int tcflush (int filedes, int queue)

The tcflush function is used to clear the input and/or output queues associated with the terminal file filedes. The queue argument specifies which queue(s) to clear, and can be one of the following values:

TCIFLUSH

Clear any input data received, but not yet read.

TCOFLUSH

Clear any output data written, but not yet transmitted.

TCIOFLUSH

Clear both queued input and output.

The return value is normally zero. In the event of an error, a value of -1 is returned. The following errno error conditions are defined for this function: EBADF The filedes is not a valid file descriptor. ENOTTY The filedes is not associated with a terminal device. EINVAL A bad value was supplied as the queue argument. It is unfortunate that this function is named tcflush, because the term “flush” is normally used for quite another operation—waiting until all output is transmitted—and using it for discarding input or output would be confusing. Unfortunately, the name tcflush comes from POSIX and we cannot change it.

Function int tcflow (int filedes, int action)

The tcflow function is used to perform operations relating to XON/XOFF flow control on the terminal file specified by filedes. The action argument specifies what operation to perform, and can be one of the following values:

TCOOFF Suspend transmission of output. TCOON Restart transmission of output. TCIOFF Transmit a STOP character. TCION Transmit a START character. For more information about the STOP and START characters, see Section 12.4.9 [Special Characters], page 286. The return value is normally zero. In the event of an error, a value of -1 is returned. The following errno error conditions are defined for this function:

Chapter 12: Low-Level Terminal Interface 295

EBADF The filedes is not a valid file descriptor.

ENOTTY The filedes is not associated with a terminal device.

EINVAL A bad value was supplied as the action argument.

12.6 Noncanonical Mode Example

Here is an example program that shows how you can set up a terminal device to read single characters in noncanonical input mode, without echo.

#include <unistd.h>
#include <stdio.h>
#include <stdlib.h>
#include <termios.h>

/* Use this variable to remember original terminal attributes. */

struct termios saved_attributes;
void
reset_input_mode (void)
{
  tcsetattr (STDIN_FILENO, TCSANOW, &saved_attributes);
}
void
set_input_mode (void)
{
  struct termios tattr;
  char *name;

/* Make sure stdin is a terminal. */

  if (!isatty (STDIN_FILENO))
    {
      fprintf (stderr, "Not a terminal.\n");
      exit (EXIT_FAILURE);
    }

/* Save the terminal attributes so we can restore them later. */

  tcgetattr (STDIN_FILENO, &saved_attributes);
  atexit (reset_input_mode);

/* Set the funny terminal modes. */

       tcgetattr (STDIN_FILENO, &tattr);
       tattr.c_lflag &= ~(ICANON|ECHO); /* Clear ICANON and ECHO. */
       tattr.c_cc[VMIN] = 1;
       tattr.c_cc[VTIME] = 0;
       tcsetattr (STDIN_FILENO, TCSAFLUSH, &tattr);
     }
     int
     main (void)
     {
       char c;
       set_input_mode ();
       while (1)
         {
           read (STDIN_FILENO, &c, 1);
           if (c == '\004')          /* C-d */
             break;
           else
             putchar (c);
         }
       return EXIT_SUCCESS;
     }

This program is careful to restore the original terminal modes before exiting or terminating with a signal. It uses the atexit function (see Section 22.3.3 [Cleanups on Exit], page 464) to make sure this is done by exit.

The shell is supposed to take care of resetting the terminal modes when a process is stopped or continued; see Chapter 24 [Job Control], page 479. But some existing shells do not actually do this, so you may wish to establish handlers for job control signals that reset terminal modes. The above example does so.

13 Mathematics

This chapter contains information about functions for performing mathematical computations, such as trigonometric functions. Most of these functions have prototypes declared in the header file ‘math.h’.

All of the functions that operate on floating-point numbers accept arguments and return results of type double. In the future, there may be additional functions that operate on float and long double values. For example, cosf and cosl would be versions of the cos function that operate on float and long double arguments, respectively. In the meantime, you should avoid using these names yourself. See Section 1.3.3 [Reserved Names], page 6.

13.1 Domain and Range Errors

Many of the functions listed in this chapter are defined mathematically over a domain that is only a subset of real numbers. For example, the acos function is defined over the domain between -1 and 1. If you pass an argument to one of these functions that is outside the domain over which it is defined, the function sets errno to EDOM to indicate a domain error. On machines that support IEEE 754 floating point, functions reporting error EDOM also return a NaN.

Some of these functions are defined mathematically to result in a complex value over parts of their domains. The most familiar example of this is taking the square root of a negative number. The functions in this chapter take only real arguments and return only real values; therefore, if the value ought to be nonreal, this is treated as a domain error.

A related problem is that the mathematical result of a function may not be representable as a floating point number. If magnitude of the correct result is too large to be represented, the function sets errno to ERANGE to indicate a range error, and returns a particular very large value (named by the macro HUGE_VAL) or its negation (- HUGE_VAL).

If the magnitude of the result is too small, a value of zero is returned instead. In this case, errno might or might not be set to ERANGE.

The only completely reliable way to check for domain and range errors is to set errno to 0 before you call the mathematical function and test errno afterward. As a consequence of this use of errno, use of the mathematical functions is not reentrant if you check for errors.

None of the mathematical functions ever generates signals as a result of domain or range errors. In particular, this means that you won’t see SIGFPE signals generated within these functions. (See Chapter 21 [Signal Handling], page 393, for more information about signals.)

Macro double HUGE VAL

An expression representing a particular very large number. On machines that use IEEE 754/IEEE 854 floating point format, the value is “infinity”. On other machines, it’s typically the largest positive number that can be represented. The value of this macro is used as the return value from various mathematical double returning functions in overflow situations.

Macro float HUGE VALf

This macro is similar to the HUGE_VAL macro except that it is used by functions returning float values. This macro is a GNU extension.

Macro long double HUGE VALl

This macro is similar to the HUGE_VAL macro except that it is used by functions returning long double values. The value is only different from HUGE_VAL if the architecture really supports long double values. This macro is a GNU extension.

For more information about floating-point representations and limits, see Section A.5.3.2 [Floating Point Parameters], page 566. In particular, the macro DBL_MAX might be more appropriate than HUGE_VAL for many uses other than testing for an error in a mathematical function.

13.2 Trigonometric Functions

These are the familiar sin, cos, and tan functions. The arguments to all of these functions are in units of radians; recall that pi radians equals 180 degrees.

The math library doesn’t define a symbolic constant for pi, but you can define your own if you need one:

  #define PI 3.14159265358979323846264338327

You can also compute the value of pi with the expression acos (-1.0).

Function double sin (double x)

This function returns the sine of x, where x is given in radians. The return value is in the range -1 to 1.

Function double cos (double x)

This function returns the cosine of x, where x is given in radians. The return value is in the range -1 to 1.

Function double tan (double x)

This function returns the tangent of x, where x is given in radians. The following errno error conditions are defined for this function:

ERANGE Mathematically, the tangent function has singularities at odd multiples of pi/2. If the argument x is too close to one of these singularities, tan sets errno to ERANGE and returns either positive or negative HUGE_VAL.

13.3 Inverse Trigonometric Functions

These are the usual arc sine, arc cosine and arc tangent functions, which are the inverses of the sine, cosine and tangent functions, respectively.

Function double asin (double x)

This function computes the arc sine of x—that is, the value whose sine is x. The value is in units of radians. Mathematically, there are infinitely many such values; the one actually returned is the one between -pi/2 and pi/2 (inclusive). asin fails, and sets errno to EDOM, if x is out of range. The arc sine function is defined mathematically only over the domain -1 to 1.

Function double acos (double x)

This function computes the arc cosine of x—that is, the value whose cosine is x. The value is in units of radians. Mathematically, there are infinitely many such values; the one actually returned is the one between 0 and pi (inclusive). acos fails, and sets errno to EDOM, if x is out of range. The arc cosine function is defined mathematically only over the domain -1 to 1.

Function double atan (double x)

This function computes the arc tangent of x—that is, the value whose tangent is x. The value is in units of radians. Mathematically, there are infinitely many such values; the one actually returned is the one between -pi/2 and pi/2 (inclusive).

Function double atan2 (double y, double x)

This is the two argument arc tangent function. It is similar to computing the arc tangent of y/x, except that the signs of both arguments are used to determine the quadrant of the result, and x is permitted to be zero. The return value is given in radians and is in the range -pi to pi, inclusive. If x and y are coordinates of a point in the plane, atan2 returns the signed angle between the line from the origin to that point and the x-axis. Thus, atan2 is useful for converting Cartesian coordinates to polar coordinates. (To compute the radial coordinate, use hypot; see Section 13.4 [Exponentiation and Logarithms], page 300.) The function atan2 sets errno to EDOM if both x and y are zero; the return value is not defined in this case.

13.4 Exponentiation and Logarithms

Function double exp (double x)

The exp function returns the value of e (the base of natural logarithms) raised to power x. The function fails, and sets errno to ERANGE, if the magnitude of the result is too large to be representable.

Function double log (double x)

This function returns the natural logarithm of x. exp (log (x)) equals x, exactly in mathematics and approximately in C. The following errno error conditions are defined for this function:

EDOM The argument x is negative. The log function is defined mathematically to return a real result only on positive arguments.

ERANGE The argument is zero. The log of zero is not defined.

Function double log10 (double x)

This function returns the base-10 logarithm of x. Except for the different base, it is similar to the log function. In fact, log10 (x) equals

log (x) / log (10).

Function double pow (double base, double power)

This is a general exponentiation function, returning base raised to power. The following errno error conditions are defined for this function:

EDOM The argument base is negative and power is not an integral value. Mathematically, the result would be a complex number in this case.

ERANGE An underflow or overflow condition was detected in the result.

Function double sqrt (double x)

This function returns the nonnegative square root of x. The sqrt function fails, and sets errno to EDOM, if x is negative. Mathematically, the square root would be a complex number.

Function double cbrt (double x)

This function returns the cube root of x. This function cannot fail; every representable real value has a representable real cube root.

Function double hypot (double x, double y)

The hypot function returns sqrt (x*x + y*y). (This is the length of the hypotenuse of a right triangle with sides of length x and y, or the distance of the point (x, y) from the origin.) See also the function cabs in Section 14.3 [Absolute Value], page 306.

Function double expm1 (double x)

This function returns a value equivalent to exp (x) - 1. It is computed in a way that is accurate even if the value of x is near zero—a case where exp (x) - 1 would be inaccurate due to subtraction of two numbers that are nearly equal.

Function double log1p (double x)

This function returns a value equivalent to log (1 + x). It is computed in a way that is accurate even if the value of x is near zero.

13.5 Hyperbolic Functions

The functions in this section are related to the exponential functions; see Section 13.4 [Exponentiation and Logarithms], page 300.

Function double sinh (double x)

The sinh function returns the hyperbolic sine of x, defined mathematically as exp (x) - exp (-x) / 2. The function fails, and sets errno to ERANGE, if the value of x is too large; that is, if overflow occurs.

Function double cosh (double x)

The cosh function returns the hyperbolic cosine of x, defined mathematically as exp (x) + exp (-x) / 2. The function fails, and sets errno to ERANGE, if the value of x is too large; that is, if overflow occurs.

Function double tanh (double x)

This function returns the hyperbolic tangent of x, whose mathematical definition is sinh (x) / cosh (x).

Function double asinh (double x)

This function returns the inverse hyperbolic sine of x—the value whose hyperbolic sine is x.

Function double acosh (double x)

This function returns the inverse hyperbolic cosine of x—the value whose hyperbolic cosine is x. If x is less than 1, acosh returns HUGE_

VAL.

Function double atanh (double x)

This function returns the inverse hyperbolic tangent of x—the value whose hyperbolic tangent is x. If the absolute value of x is greater than or equal to 1, atanh returns HUGE_VAL.

13.6 Pseudo-Random Numbers

This section describes the GNU facilities for generating a series of pseudorandom numbers. The numbers generated are not truly random; typically, they form a sequence that repeats periodically, with a period so large that you can ignore it for ordinary purposes. The random number generator works by remembering at all times a seed value which it uses to compute the next random number and also to compute a new seed.

Although the generated numbers look unpredictable within one run of a program, the sequence of numbers is exactly the same from one run to the next. This is because the initial seed is always the same. This is convenient when you are debugging a program, but it is unhelpful if you want the program to behave unpredictably. If you want truly random numbers, not just pseudo-random, specify a seed based on the current time.

You can get repeatable sequences of numbers on a particular machine type by specifying the same initial seed value for the random number generator. There is no standard meaning for a particular seed value; the same seed, used in different C libraries or on different CPU types, will give you different random numbers.

The GNU library supports the standard ISO C random number functions plus another set derived from BSD. We recommend you use the standard

ones, rand and srand.

13.6.1 ISO C Random Number Functions

This section describes the random number functions that are part of the ISO C standard.

To use these facilities, you should include the header file ‘stdlib.h’ in your program.

Macro int RAND MAX

The value of this macro is an integer constant expression that represents the maximum possible value returned by the rand function. In the GNU library, it is 037777777, which is the largest signed integer representable in 32 bits. In other libraries, it may be as low as 32767.

Function int rand ()

The rand function returns the next pseudo-random number in the series. The value is in the range from 0 to RAND_MAX.

Function void srand (unsigned int seed)

This function establishes seed as the seed for a new series of pseudorandom numbers. If you call rand before a seed has been established with srand, it uses the value 1 as a default seed. To produce truly random numbers (not just pseudo-random), do srand

(time (0)).

13.6.2 BSD Random Number Functions

This section describes a set of random number generation functions that are derived from BSD. There is no advantage to using these functions with the GNU C library; we support them for BSD compatibility only.

The prototypes for these functions are in ‘stdlib.h’.

Function long int random ()

This function returns the next pseudo-random number in the sequence. The range of values returned is from 0 to RAND_MAX.

Function void srandom (unsigned int seed)

The srandom function sets the seed for the current random number state based on the integer seed. If you supply a seed value of 1, this will cause random to reproduce the default set of random numbers. To produce truly random numbers (not just pseudo-random), do

srandom (time (0)).

Function void * initstate (unsigned int seed, void *state,

      size_t size)

The initstate function is used to initialize the random number generator state. The argument state is an array of size bytes, used to hold the state information. The size must be at least 8 bytes, and optimal sizes are 8, 16, 32, 64, 128, and 256. The bigger the state array, the better. The return value is the previous value of the state information array. You can use this value later as an argument to setstate to restore that state.

Function void * setstate (void *state)

The setstate function restores the random number state information state. The argument must have been the result of a previous call to initstate or setstate. The return value is the previous value of the state information array. You can use thise value later as an argument to setstate to restore that state.

304 The GNU C Library

14 Low-Level Arithmetic Functions

This chapter contains information about functions for doing basic arithmetic operations, such as splitting a float into its integer and fractional parts. These functions are declared in the header file ‘math.h’.

14.1 “Not a Number” Values

The IEEE floating point format used by most modern computers supports values that are “not a number”. These values are called NaNs. “Not a number” values result from certain operations which have no meaningful numeric result, such as zero divided by zero or infinity divided by infinity.

One noteworthy property of NaNs is that they are not equal to themselves. Thus, x == x can be 0 if the value of x is a NaN. You can use this to test whether a value is a NaN or not: if it is not equal to itself, then it is a NaN. But the recommended way to test for a NaN is with the isnan function (see Section 14.2 [Predicates on Floats], page 305).

Almost any arithmetic operation in which one argument is a NaN returns a NaN.

Macro double NAN

An expression representing a value which is “not a number”. This macro is a GNU extension, available only on machines that support “not a number” values—that is to say, on all machines that support IEEE floating point. You can use ‘#ifdef NAN’ to test whether the machine supports NaNs. (Of course, you must arrange for GNU extensions to be visible, such as by defining _GNU_SOURCE, and then you must include ‘math.h’.)

14.2 Predicates on Floats

This section describes some miscellaneous test functions on doubles. Prototypes for these functions appear in ‘math.h’. These are BSD functions, and thus are available if you define _BSD_SOURCE or _GNU_SOURCE.

Function int isinf (double x)

This function returns -1 if x represents negative infinity, 1 if x represents positive infinity, and 0 otherwise.

Function int isnan (double x)

This function returns a nonzero value if x is a “not a number” value, and zero otherwise. (You can just as well use x != x to get the same result).

Function int finite (double x)

This function returns a nonzero value if x is finite or a “not a number” value, and zero otherwise.

Function double infnan (int error)

This function is provided for compatibility with BSD. The other mathematical functions use infnan to decide what to return on occasion of an error. Its argument is an error code, EDOM or ERANGE; infnan returns a suitable value to indicate this with. -ERANGE is also acceptable as an argument, and corresponds to -HUGE_VAL as a value. In the BSD library, on certain machines, infnan raises a fatal signal in all cases. The GNU library does not do likewise, because that does not fit the ISO C specification.

Portability Note: The functions listed in this section are BSD extensions.

14.3 Absolute Value

These functions are provided for obtaining the absolute value (or magnitude) of a number. The absolute value of a real number x is x is x is positive, −x if x is negative. For a complex number z, whose real part is x and whose imaginary part is y, the absolute value is sqrt (x*x + y*y).

Prototypes for abs and labs are in ‘stdlib.h’; fabs and cabs are declared in ‘math.h’.

Function int abs (int number)

This function returns the absolute value of number. Most computers use a two’s complement integer representation, in which the absolute value of INT_MIN (the smallest possible int) cannot be represented; thus, abs (INT_MIN) is not defined.

Function long int labs (long int number)

This is similar to abs, except that both the argument and result are of type long int rather than int.

Function double fabs (double number)

This function returns the absolute value of the floating-point number number.

Function double cabs (struct { double real, imag; } z)

The cabs function returns the absolute value of the complex number z, whose real part is z.real and whose imaginary part is z.imag. (See also the function hypot in Section 13.4 [Exponentiation and Logarithms], page 300.) The value is:

     sqrt (z.real*z.real + z.imag*z.imag)

14.4 Normalization Functions

The functions described in this section are primarily provided as a way to efficiently perform certain low-level manipulations on floating point numbers that are represented internally using a binary radix; see Section A.5.3.1 [Floating Point Representation Concepts], page 565. These functions are required to have equivalent behavior even if the representation does not use a radix of 2, but of course they are unlikely to be particularly efficient in those cases.

All these functions are declared in ‘math.h’.

Function double frexp (double value, int *exponent)

The frexp function is used to split the number value into a normalized fraction and an exponent. If the argument value is not zero, the return value is value times a power of two, and is always in the range 1/2 (inclusive) to 1 (exclusive). The corresponding exponent is stored in *exponent; the return value multiplied by 2 raised to this exponent equals the original number value. For example, frexp (12.8, &exponent) returns 0.8 and stores 4 in

exponent.

If value is zero, then the return value is zero and zero is stored in *exponent.

Function double ldexp (double value, int exponent)

This function returns the result of multiplying the floating-point number value by 2 raised to the power exponent. (It can be used to reassemble floating-point numbers that were taken apart by frexp.) For example, ldexp (0.8, 4) returns 12.8.

The following functions which come from BSD provide facilities equivalent to those of ldexp and frexp:

Function double scalb (double value, int exponent)

The scalb function is the BSD name for ldexp.

Function double logb (double x)

This BSD function returns the integer part of the base-2 logarithm of x, an integer value represented in type double. This is the highest integer power of 2 contained in x. The sign of x is ignored. For

example, logb (3.5) is 1.0 and logb (4.0) is 2.0.

When 2 raised to this power is divided into x, it gives a quotient between 1 (inclusive) and 2 (exclusive). If x is zero, the value is minus infinity (if the machine supports such a value), or else a very small number. If x is infinity, the value is infinity.

The value returned by logb is one less than the value that frexp would store into *exponent.

Function double copysign (double value, double sign)

The copysign function returns a value whose absolute value is the same as that of value, and whose sign matches that of sign. This is a BSD function.

14.5 Rounding and Remainder Functions

The functions listed here perform operations such as rounding, truncation, and remainder in division of floating point numbers. Some of these functions convert floating point numbers to integer values. They are all declared in ‘math.h’.

You can also convert floating-point numbers to integers simply by casting them to int. This discards the fractional part, effectively rounding towards zero. However, this only works if the result can actually be represented as an int—for very large numbers, this is impossible. The functions listed here return the result as a double instead to get around this problem.

Function double ceil (double x)

The ceil function rounds x upwards to the nearest integer, returning that value as a double. Thus, ceil (1.5) is 2.0.

Function double floor (double x)

The ceil function rounds x downwards to the nearest integer, returning that value as a double. Thus, floor (1.5) is 1.0 and floor

(-1.5) is -2.0.

Function double rint (double x)

This function rounds x to an integer value according to the current rounding mode. See Section A.5.3.2 [Floating Point Parameters], page 566, for information about the various rounding modes. The default rounding mode is to round to the nearest integer; some machines support other modes, but round-to-nearest is always used unless you explicit select another.

Function double modf (double value, double *integer-part)

This function breaks the argument value into an integer part and a fractional part (between -1 and 1, exclusive). Their sum equals value. Each of the parts has the same sign as value, so the rounding of the integer part is towards zero. modf stores the integer part in *integer-part, and returns the fractional part. For example, modf (2.5, &intpart) returns 0.5 and stores 2.0

into intpart.

Function double fmod (double numerator, double denominator)

This function computes the remainder from the division of numerator by denominator. Specifically, the return value is numerator - n * denominator, where n is the quotient of numerator divided by denominator, rounded towards zero to an integer. Thus,

fmod (6.5, 2.3) returns 1.9, which is 6.5 minus 4.6.

The result has the same sign as the numerator and has magnitude less than the magnitude of the denominator. If denominator is zero, fmod fails and sets errno to EDOM.

Function double drem (double numerator, double denominator)

The function drem is like fmod except that it rounds the internal quotient n to the nearest integer instead of towards zero to an integer. For example, drem (6.5, 2.3) returns -0.4, which is 6.5 minus 6.9. The absolute value of the result is less than or equal to half the absolute value of the denominator. The difference between fmod (numerator, denominator) and drem (numerator, denominator) is always either denominator, minus denominator, or zero. If denominator is zero, drem fails and sets errno to EDOM.

14.6 Integer Division

This section describes functions for performing integer division. These functions are redundant in the GNU C library, since in GNU C the ‘/’ operator always rounds towards zero. But in other C implementations, ‘/’ may round differently with negative arguments. div and ldiv are useful because they specify how to round the quotient: towards zero. The remainder has the same sign as the numerator.

These functions are specified to return a result r such that the value r.quot*denominator + r.rem equals numerator.

To use these facilities, you should include the header file ‘stdlib.h’ in your program.

Data Type div t

This is a structure type used to hold the result returned by the div function. It has the following members: int quot The quotient from the division. int rem The remainder from the division.

Function div_t div (int numerator, int denominator)

This function div computes the quotient and remainder from the division of numerator by denominator, returning the result in a structure of type div_t.

If the result cannot be represented (as in a division by zero), the behavior is undefined. Here is an example, albeit not a very useful one.

     div_t result;
     result = div (20, -6);
Now result.quot is -3 and result.rem is 2.

Data Type ldiv t

This is a structure type used to hold the result returned by the ldiv function. It has the following members:

long int quot

The quotient from the division.

long int rem

The remainder from the division.

(This is identical to div_t except that the components are of type long int rather than int.)

Function ldiv_t ldiv (long int numerator, long int

denominator) The ldiv function is similar to div, except that the arguments are of type long int and the result is returned as a structure of type ldiv.

14.7 Parsing of Numbers

This section describes functions for “reading” integer and floating-point numbers from a string. It may be more convenient in some cases to use sscanf or one of the related functions; see Section 7.12 [Formatted Input], page 126. But often you can make a program more robust by finding the tokens in the string by hand, then converting the numbers one by one.

14.7.1 Parsing of Integers

These functions are declared in ‘stdlib.h’.

Function long int strtol (const char *string, char **tailptr,

int base) The strtol (“string-to-long”) function converts the initial part of string to a signed integer, which is returned as a value of type long

int.

This function attempts to decompose string as follows:

• A (possibly empty) sequence of whitespace characters. Which characters are whitespace is determined by the isspace function (see Section 4.1 [Classification of Characters], page 61). These are discarded.

• An optional plus or minus sign (‘+’ or ‘-’).

• A nonempty sequence of digits in the radix specified by base. If base is zero, decimal radix is assumed unless the series of digits begins with ‘0’ (specifying octal radix), or ‘0x’ or ‘0X’ (specifying hexadecimal radix); in other words, the same syntax used for integer constants in C. Otherwise base must have a value between 2 and 35. If base is 16, the digits may optionally be preceded by ‘0x’ or ‘0X’. If base has no legal value the value returned is 0l and the global variable

    errno is set to EINVAL.

• Any remaining characters in the string. If tailptr is not a null pointer, strtol stores a pointer to this tail in *tailptr.

If the string is empty, contains only whitespace, or does not contain an initial substring that has the expected syntax for an integer in the specified base, no conversion is performed. In this case, strtol returns a value of zero and the value stored in *tailptr is the value of string. In a locale other than the standard "C" locale, this function may recognize additional implementation-dependent syntax. If the string has valid syntax for an integer but the value is not representable because of overflow, strtol returns either LONG_MAX or LONG_ MIN (see Section A.5.2 [Range of an Integer Type], page 563), as appropriate for the sign of the value. It also sets errno to ERANGE to indicate there was overflow. Because the value 0l is a correct result for strtol the user who is interested in handling errors should set the global variable errno to 0 before calling this function, so that the program can later test whether an error occurred. There is an example at the end of this section.

Function unsigned long int strtoul (const char *string, char

**tailptr, int base) The strtoul (“string-to-unsigned-long”) function is like strtol except it deals with unsigned numbers, and returns its value with type unsigned long int. No ‘+’ or ‘-’ sign may appear before the number, but the syntax is otherwise the same as described above for strtol. The value returned in case of overflow is ULONG_MAX (see Section A.5.2 [Range of an Integer Type], page 563). Like strtol this function sets errno and returns the value 0ul in case the value for base is not in the legal range. For strtoul this can happen in another situation. In case the number to be converted is negative strtoul also sets errno to EINVAL and returns 0ul.

Function long long int strtoq (const char *string, char

**tailptr, int base) The strtoq (“string-to-quad-word”) function is like strtol except that is deals with extra long numbers and it returns its value with

type long long int.

If the string has valid syntax for an integer but the value is not representable because of overflow, strtoq returns either LONG_LONG_MAX or LONG_LONG_MIN (see Section A.5.2 [Range of an Integer Type], page 563), as appropriate for the sign of the value. It also sets errno to ERANGE to indicate there was overflow.

Function long long int strtoll (const char *string, char

**tailptr, int base) strtoll is only an commonly used other name for the strtoq function. Everything said for strtoq applies to strtoll as well.

Function unsigned long long int strtouq (const char

*string, char **tailptr, int base) The strtouq (“string-to-unsigned-quad-word”) function is like strtoul except that is deals with extra long numbers and it returns its value with type unsigned long long int. The value returned in case of overflow is ULONG_LONG_MAX (see Section A.5.2 [Range of an Integer Type], page 563).

Function unsigned long long int strtoull (const char

*string, char **tailptr, int base) strtoull is only an commonly used other name for the strtouq function. Everything said for strtouq applies to strtoull as well.

Function long int atol (const char *string)

This function is similar to the strtol function with a base argument of 10, except that it need not detect overflow errors. The atol function is provided mostly for compatibility with existing code; using strtol is more robust.

Function int atoi (const char *string)

This function is like atol, except that it returns an int value rather than long int. The atoi function is also considered obsolete; use strtol instead.

The POSIX locales contain some information about how to format numbers (see Section 19.6.1 [Generic Numeric Formatting Parameters], page 384). This mainly deals with representing numbers for better readability for humans. The functions present so far in this section cannot handle numbers in this form.

If this functionality is needed in a program one can use the functions from the scanf family which know about the flag ‘'’ for parsing numeric input (see Section 7.12.4 [Numeric Input Conversions], page 130). Sometimes it is more desirable to have finer control.

In these situation one could use the function __strtoXXX_internal. XXX here stands for any of the above forms. All numeric conversion functions (including the functions to process floating-point numbers) have such a counterpart. The difference to the normal form is the extra argument at the end of the parameter list. If this value has an non-zero value the handling of number grouping is enabled. The advantage of using these functions is that the tailptr parameters allow to determine which part of the input is processed. The scanf functions don’t provide this information. The drawback of using these functions is that they are not portable. They only exist in the GNU C library.

Here is a function which parses a string as a sequence of integers and returns the sum of them:

     int
     sum_ints_from_string (char *string)
     {
       int sum = 0;
       while (1) {
         char *tail;
         int next;

/* Skip whitespace by hand, to detect the end. */

         while (isspace (*string)) string++;
         if (*string == 0)
           break;

/* There is more nonwhitespace, */ /* so it ought to be another number. */

         errno = 0;

/* Parse it. */

         next = strtol (string, &tail, 0);

/* Add it in, if not overflow. */

         if (errno)
           printf ("Overflow\n");
         else
           sum += next;

/* Advance past it. */

         string = tail;
       }
       return sum;
     }

14.7.2 Parsing of Floats

These functions are declared in ‘stdlib.h’.

Function double strtod (const char *string, char **tailptr)

The strtod (“string-to-double”) function converts the initial part of string to a floating-point number, which is returned as a value of type

double.

This function attempts to decompose string as follows:

• A (possibly empty) sequence of whitespace characters. Which characters are whitespace is determined by the isspace function (see Section 4.1 [Classification of Characters], page 61). These are discarded.

• An optional plus or minus sign (‘+’ or ‘-’).

• A nonempty sequence of digits optionally containing a decimalpoint character—normally ‘.’, but it depends on the locale (see Section 19.6 [Numeric Formatting], page 383).

• An optional exponent part, consisting of a character ‘e’ or ‘E’, an optional sign, and a sequence of digits.

• Any remaining characters in the string. If tailptr is not a null pointer, a pointer to this tail of the string is stored in *tailptr.

If the string is empty, contains only whitespace, or does not contain an initial substring that has the expected syntax for a floating-point number, no conversion is performed. In this case, strtod returns a value of zero and the value returned in *tailptr is the value of string. In a locale other than the standard "C" or "POSIX" locales, this function may recognize additional locale-dependent syntax. If the string has valid syntax for a floating-point number but the value is not representable because of overflow, strtod returns either positive or negative HUGE_VAL (see Chapter 13 [Mathematics], page 297), depending on the sign of the value. Similarly, if the value is not representable because of underflow, strtod returns zero. It also sets errno to ERANGE if there was overflow or underflow. Since the value zero which is returned in the error case is also a valid result the user should set the global variable errno to zero before calling this function. So one can test for failures after the call since all failures set errno to a non-zero value.

Function float strtof (const char *string, char **tailptr)

This function is similar to the strtod function but it returns a float value instead of a double value. If the precision of a float value is sufficient this function should be used since it is much faster than strtod on some architectures. The reasons are obvious: IEEE 754 defines float to have a mantissa of 23 bits while double has 53 bits and every additional bit of precision can require additional computation. If the string has valid syntax for a floating-point number but the value is not representable because of overflow, strtof returns either positive or negative HUGE_VALf (see Chapter 13 [Mathematics], page 297), depending on the sign of the value. This function is a GNU extension.

Function long double strtold (const char *string, char

**tailptr) This function is similar to the strtod function but it returns a long double value instead of a double value. It should be used when high precision is needed. On systems which define a long double type (i.e., on which it is not the same as double) running this function might take significantly more time since more bits of precision are required. If the string has valid syntax for a floating-point number but the value is not representable because of overflow, strtold returns either positive or negative HUGE_VALl (see Chapter 13 [Mathematics], page 297), depending on the sign of the value. This function is a GNU extension.

As for the integer parsing functions there are additional functions which will handle numbers represented using the grouping scheme of the current locale (see Section 14.7.1 [Parsing of Integers], page 310).

Function double atof (const char *string)

This function is similar to the strtod function, except that it need not detect overflow and underflow errors. The atof function is provided mostly for compatibility with existing code; using strtod is more robust.

316 The GNU C Library

15 Searching and Sorting

This chapter describes functions for searching and sorting arrays of arbitrary objects. You pass the appropriate comparison function to be applied as an argument, along with the size of the objects in the array and the total number of elements.

15.1 Defining the Comparison Function

In order to use the sorted array library functions, you have to describe how to compare the elements of the array.

To do this, you supply a comparison function to compare two elements of the array. The library will call this function, passing as arguments pointers to two array elements to be compared. Your comparison function should return a value the way strcmp (see Section 5.5 [String/Array Comparison], page 72) does: negative if the first argument is “less” than the second, zero if they are “equal”, and positive if the first argument is “greater”.

Here is an example of a comparison function which works with an array of numbers of type double:

     int
     compare_doubles (const double *a, const double *b)
     {
       return (int) (*a - *b);
     }

The header file ‘stdlib.h’ defines a name for the data type of comparison functions. This type is a GNU extension.

     int comparison_fn_t (const void *, const void *);

15.2 Array Search Function

To search a sorted array for an element matching the key, use the bsearch function. The prototype for this function is in the header file ‘stdlib.h’.

Function void * bsearch (const void *key, const void *array,
         size_t count, size_t size, comparison_fn_t compare)

The bsearch function searches the sorted array array for an object that is equivalent to key. The array contains count elements, each of which is of size size bytes. The compare function is used to perform the comparison. This function is called with two pointer arguments and should return an integer less than, equal to, or greater than zero corresponding to whether its first argument is considered less than, equal to, or greater than its second argument. The elements of the array must already be sorted in ascending order according to this comparison function. The return value is a pointer to the matching array element, or a null pointer if no match is found. If the array contains more than one element that matches, the one that is returned is unspecified. This function derives its name from the fact that it is implemented using the binary search algorithm.

15.3 Array Sort Function

To sort an array using an arbitrary comparison function, use the qsort function. The prototype for this function is in ‘stdlib.h’.

Function void qsort (void *array, size_t count, size_t size,

      comparison_fn_t compare)

The qsort function sorts the array array. The array contains count elements, each of which is of size size. The compare function is used to perform the comparison on the array elements. This function is called with two pointer arguments and should return an integer less than, equal to, or greater than zero corresponding to whether its first argument is considered less than, equal to, or greater than its second argument. Warning: If two objects compare as equal, their order after sorting is unpredictable. That is to say, the sorting is not stable. This can make a difference when the comparison considers only part of the elements. Two elements with the same sort key may differ in other respects. If you want the effect of a stable sort, you can get this result by writing the comparison function so that, lacking other reason distinguish between two elements, it compares them by their addresses. Note that doing this may make the sorting algorithm less efficient, so do it only if necessary. Here is a simple example of sorting an array of doubles in numerical order, using the comparison function defined above (see Section 15.1 [Defining the Comparison Function], page 317):

     {
       double *array;
       int size;

…

       qsort (array, size, sizeof (double), compare_doubles);
     }

The qsort function derives its name from the fact that it was originally implemented using the “quick sort” algorithm.

Chapter 15: Searching and Sorting 319

15.4 Searching and Sorting Example

Here is an example showing the use of qsort and bsearch with an array of structures. The objects in the array are sorted by comparing their name fields with the strcmp function. Then, we can look up individual objects based on their names.

#include <stdlib.h>
#include <stdio.h>
#include <string.h>

/* Define an array of critters to sort. */

struct critter
{
const char *name;
const char *species;
};
struct critter muppets[] =
{
{"Kermit", "frog"},
{"Piggy", "pig"},
{"Gonzo", "whatever"},
{"Fozzie", "bear"},
{"Sam", "eagle"},
{"Robin", "frog"},
{"Animal", "animal"},
{"Camilla", "chicken"},
{"Sweetums", "monster"},
{"Dr. Strangepork", "pig"},
{"Link Hogthrob", "pig"},
{"Zoot", "human"},
{"Dr. Bunsen Honeydew", "human"},
{"Beaker", "human"},
{"Swedish Chef", "human"}
};
int count = sizeof (muppets) / sizeof (struct critter);

/* This is the comparison function used for sorting and searching. */

int

critter_cmp (const struct critter *c1, const struct critter *c2)
{
  return strcmp (c1->name, c2->name);
}

/* Print information about a critter. */

void
print_critter (const struct critter *c)
{
  printf ("%s, the %s\n", c->name, c->species);
}

/* Do the lookup into the sorted array. */

void
find_critter (const char *name)
{
  struct critter target, *result;
  target.name = name;
  result = bsearch (&target, muppets, count, sizeof (struct critter),
                    critter_cmp);
  if (result)
    print_critter (result);
  else
    printf ("Couldn't find %s.\n", name);
}

/* Main program. */

int
main (void)
{
  int i;
  for (i = 0; i < count; i++)
    print_critter (&muppets[i]);
  printf ("\n");
  qsort (muppets, count, sizeof (struct critter), critter_cmp);
  for (i = 0; i < count; i++)
    print_critter (&muppets[i]);

Chapter 15: Searching and Sorting 321

  printf ("\n");
  find_critter ("Kermit");
  find_critter ("Gonzo");
  find_critter ("Janice");
  return 0;
}

The output from this program looks like:

Kermit, the frog
Piggy, the pig
Gonzo, the whatever
Fozzie, the bear
Sam, the eagle
Robin, the frog
Animal, the animal
Camilla, the chicken
Sweetums, the monster
Dr. Strangepork, the pig
Link Hogthrob, the pig
Zoot, the human
Dr. Bunsen Honeydew, the human
Beaker, the human
Swedish Chef, the human
Animal, the animal
Beaker, the human
Camilla, the chicken
Dr. Bunsen Honeydew, the human
Dr. Strangepork, the pig
Fozzie, the bear
Gonzo, the whatever
Kermit, the frog
Link Hogthrob, the pig
Piggy, the pig
Robin, the frog
Sam, the eagle
Swedish Chef, the human
Sweetums, the monster
Zoot, the human
Kermit, the frog
Gonzo, the whatever
Couldn't find Janice.

322 The GNU C Library

16 Pattern Matching

The GNU C Library provides pattern matching facilities for two kinds of patterns: regular expressions and file-name wildcards. The library also provides a facility for expanding variable and command references and parsing text into words in the way the shell does.

16.1 Wildcard Matching

This section describes how to match a wildcard pattern against a particular string. The result is a yes or no answer: does the string fit the pattern or not. The symbols described here are all declared in ‘fnmatch.h’.

Function int fnmatch (const char *pattern, const char

*string, int flags) This function tests whether the string string matches the pattern pattern. It returns 0 if they do match; otherwise, it returns the nonzero value FNM_NOMATCH. The arguments pattern and string are both strings. The argument flags is a combination of flag bits that alter the details of matching. See below for a list of the defined flags. In the GNU C Library, fnmatch cannot experience an “error”—it always returns an answer for whether the match succeeds. However, other implementations of fnmatch might sometimes report “errors”. They would do so by returning nonzero values that are not equal to

FNM_NOMATCH.

These are the available flags for the flags argument:

FNM_FILE_NAME

Treat the ‘/’ character specially, for matching file names. If this flag is set, wildcard constructs in pattern cannot match ‘/’ in string. Thus, the only way to match ‘/’ is with an explicit ‘/’ in pattern.

FNM_PATHNAME

This is an alias for FNM_FILE_NAME; it comes from POSIX.2. We don’t recommend this name because we don’t use the term “pathname” for file names.

FNM_PERIOD

Treat the ‘.’ character specially if it appears at the beginning of string. If this flag is set, wildcard constructs in pattern cannot match ‘.’ as the first character of string. If you set both FNM_PERIOD and FNM_FILE_NAME, then the special treatment applies to ‘.’ following ‘/’ as well as to ‘.’ at the beginning of string. (The shell uses the FNM_PERIOD and FNM_ FILE_NAME flags together for matching file names.)

FNM_NOESCAPE

Don’t treat the ‘\’ character specially in patterns. Normally, ‘\’ quotes the following character, turning off its special meaning (if any) so that it matches only itself. When quoting is enabled, the pattern ‘\?’ matches only the string ‘?’, because the question mark in the pattern acts like an ordinary character. If you use FNM_NOESCAPE, then ‘\’ is an ordinary character.

FNM_LEADING_DIR

Ignore a trailing sequence of characters starting with a ‘/’ in string; that is to say, test whether string starts with a directory name that pattern matches. If this flag is set, either ‘foo*’ or ‘foobar’ as a pattern would match the string ‘foobar/frobozz’.

FNM_CASEFOLD

Ignore case in comparing string to pattern.

16.2 Globbing

The archetypal use of wildcards is for matching against the files in a directory, and making a list of all the matches. This is called globbing.

You could do this using fnmatch, by reading the directory entries one by one and testing each one with fnmatch. But that would be slow (and complex, since you would have to handle subdirectories by hand).

The library provides a function glob to make this particular use of wildcards convenient. glob and the other symbols in this section are declared in

‘glob.h’.

16.2.1 Calling glob

The result of globbing is a vector of file names (strings). To return this vector, glob uses a special data type, glob_t, which is a structure. You pass glob the address of the structure, and it fills in the structure’s fields to tell you about the results.

Data Type glob t

This data type holds a pointer to a word vector. More precisely, it records both the address of the word vector and its size.

gl_pathc The number of elements in the vector.

gl_pathv The address of the vector. This field has type char **. gl_offs The offset of the first real element of the vector, from its

nominal address in the gl_pathv field. Unlike the other fields, this is always an input to glob, rather than an output from it. If you use a nonzero offset, then that many elements at the beginning of the vector are left empty. (The glob function fills them with null pointers.) The gl_offs field is meaningful only if you use the GLOB_ DOOFFS flag. Otherwise, the offset is always zero regardless of what is in this field, and the first real element comes at the beginning of the vector.

Function int glob (const char *pattern, int flags, int (*errfunc) (const char *filename, int error-code), glob_t *vector-ptr)

The function glob does globbing using the pattern pattern in the current directory. It puts the result in a newly allocated vector, and stores the size and address of this vector into *vector-ptr. The argument flags is a combination of bit flags; see Section 16.2.2 [Flags for Globbing], page 326, for details of the flags. The result of globbing is a sequence of file names. The function glob allocates a string for each resulting word, then allocates a vector of type char ** to store the addresses of these strings. The last element of the vector is a null pointer. This vector is called the word vector. To return this vector, glob stores both its address and its length (number of elements, not counting the terminating null pointer) into *vector-ptr. Normally, glob sorts the file names alphabetically before returning them. You can turn this off with the flag GLOB_NOSORT if you want to get the information as fast as possible. Usually it’s a good idea to let glob sort them—if you process the files in alphabetical order, the users will have a feel for the rate of progress that your application is making. If glob succeeds, it returns 0. Otherwise, it returns one of these error codes:

GLOB_ABORTED

There was an error opening a directory, and you used the flag GLOB_ERR or your specified errfunc returned a nonzero value. See below for an explanation of the GLOB_ERR flag and errfunc.

GLOB_NOMATCH

The pattern didn’t match any existing files. If you use the GLOB_NOCHECK flag, then you never get this error code, because that flag tells glob to pretend that the pattern matched at least one file.

GLOB_NOSPACE

It was impossible to allocate memory to hold the result.

In the event of an error, glob stores information in *vector-ptr about all the matches it has found so far.

16.2.2 Flags for Globbing

This section describes the flags that you can specify in the flags argument to glob. Choose the flags you want, and combine them with the C bitwise OR operator |.

GLOB_APPEND

Append the words from this expansion to the vector of words produced by previous calls to glob. This way you can effectively expand several words as if they were concatenated with spaces between them. In order for appending to work, you must not modify the contents of the word vector structure between calls to glob. And, if you set GLOB_DOOFFS in the first call to glob, you must also set it when you append to the results. Note that the pointer stored in gl_pathv may no longer be valid after you call glob the second time, because glob might have relocated the vector. So always fetch gl_pathv from the glob_ t structure after each glob call; never save the pointer across calls.

GLOB_DOOFFS

Leave blank slots at the beginning of the vector of words. The gl_offs field says how many slots to leave. The blank slots contain null pointers.

GLOB_ERR Give up right away and report an error if there is any difficulty reading the directories that must be read in order to expand pattern fully. Such difficulties might include a directory in which you don’t have the requisite access. Normally, glob tries its best to keep on going despite any errors, reading whatever directories it can. You can exercise even more control than this by specifying an error-handler function errfunc when you call glob. If errfunc is not a null pointer, then glob doesn’t give up right away when it can’t read a directory; instead, it calls errfunc with two arguments, like this:

(*errfunc) (filename, error-code)

The argument filename is the name of the directory that glob couldn’t open or couldn’t read, and error-code is the errno value that was reported to glob. If the error handler function returns nonzero, then glob gives up right away. Otherwise, it continues.

GLOB_MARK

If the pattern matches the name of a directory, append ‘/’ to the directory’s name when returning it.

GLOB_NOCHECK

If the pattern doesn’t match any file names, return the pattern itself as if it were a file name that had been matched. (Normally, when the pattern doesn’t match anything, glob returns that there were no matches.)

GLOB_NOSORT

Don’t sort the file names; return them in no particular order. (In practice, the order will depend on the order of the entries in the directory.) The only reason not to sort is to save time.

GLOB_NOESCAPE

Don’t treat the ‘\’ character specially in patterns. Normally, ‘\’ quotes the following character, turning off its special meaning (if any) so that it matches only itself. When quoting is enabled, the pattern ‘\?’ matches only the string ‘?’, because the question mark in the pattern acts like an ordinary character. If you use GLOB_NOESCAPE, then ‘\’ is an ordinary character. glob does its work by calling the function fnmatch repeatedly. It handles the flag GLOB_NOESCAPE by turning on the FNM_

NOESCAPE flag in calls to fnmatch.

16.3 Regular Expression Matching

The GNU C library supports two interfaces for matching regular expressions. One is the standard POSIX.2 interface, and the other is what the GNU system has had for many years.

Both interfaces are declared in the header file ‘regex.h’. If you define _POSIX_C_SOURCE, then only the POSIX.2 functions, structures, and constants are declared.

16.3.1 POSIX Regular Expression Compilation

Before you can actually match a regular expression, you must compile it. This is not true compilation—it produces a special data structure, not machine instructions. But it is like ordinary compilation in that its purpose is to enable you to “execute” the pattern fast. (See Section 16.3.3 [Matching a Compiled POSIX Regular Expression], page 330, for how to use the compiled regular expression for matching.)

There is a special data type for compiled regular expressions:

Data Type regex t

This type of object holds a compiled regular expression. It is actually a structure. It has just one field that your programs should look at:

re_nsub This field holds the number of parenthetical subexpressions in the regular expression that was compiled.

There are several other fields, but we don’t describe them here, because only the functions in the library should use them.

After you create a regex_t object, you can compile a regular expression into it by calling regcomp.

Function int regcomp (regex_t *compiled, const char

*pattern, int cflags) The function regcomp “compiles” a regular expression into a data structure that you can use with regexec to match against a string. The compiled regular expression format is designed for efficient matching. regcomp stores it into *compiled. It’s up to you to allocate an object of type regex_t and pass its address

to regcomp.

The argument cflags lets you specify various options that control the syntax and semantics of regular expressions. See Section 16.3.2 [Flags for POSIX Regular Expressions], page 329. If you use the flag REG_NOSUB, then regcomp omits from the compiled regular expression the information necessary to record how subexpressions actually match. In this case, you might as well pass 0 for the matchptr and nmatch arguments when you call regexec. If you don’t use REG_NOSUB, then the compiled regular expression does have the capacity to record how subexpressions match. Also, regcomp tells you how many subexpressions pattern has, by storing the number in compiled->re_nsub. You can use that value to decide how long an array to allocate to hold information about subexpression matches. regcomp returns 0 if it succeeds in compiling the regular expression; otherwise, it returns a nonzero error code (see the table below). You can use regerror to produce an error message string describing the reason for a nonzero value; see Section 16.3.6 [POSIX Regexp Matching Cleanup], page 332.

Here are the possible nonzero values that regcomp can return:

REG_BADBR

There was an invalid ‘\{…\}’ construct in the regular expression. A valid ‘\{…\}’ construct must contain either a single number, or two numbers in increasing order separated by a comma.

REG_BADPAT

There was a syntax error in the regular expression.

REG_BADRPT

A repetition operator such as ‘?’ or ‘*’ appeared in a bad position (with no preceding subexpression to act on).

REG_ECOLLATE

The regular expression referred to an invalid collating element (one not defined in the current locale for string collation). See Section 19.3 [Categories of Activities that Locales Affect], page 380.

REG_ECTYPE

The regular expression referred to an invalid character class name.

REG_EESCAPE

The regular expression ended with ‘\’.

REG_ESUBREG

There was an invalid number in the ‘\digit’ construct.

REG_EBRACK

There were unbalanced square brackets in the regular expression.

REG_EPAREN

An extended regular expression had unbalanced parentheses, or a basic regular expression had unbalanced ‘\(’ and ‘\)’.

REG_EBRACE

The regular expression had unbalanced ‘\{’ and ‘\}’.

REG_ERANGE

One of the endpoints in a range expression was invalid.

REG_ESPACE

regcomp ran out of memory.

16.3.2 Flags for POSIX Regular Expressions

These are the bit flags that you can use in the cflags operand when compiling a regular expression with regcomp.

REG_EXTENDED

Treat the pattern as an extended regular expression, rather than as a basic regular expression.

REG_ICASE

Ignore case when matching letters.

REG_NOSUB

Don’t bother storing the contents of the matches-ptr array.

REG_NEWLINE

Treat a newline in string as dividing string into multiple lines, so that ‘$’ can match before the newline and ‘^’ can match after. Also, don’t permit ‘.’ to match a newline, and don’t permit ‘[^…]’ to match a newline. Otherwise, newline acts like any other ordinary character.

16.3.3 Matching a Compiled POSIX Regular Expression

Once you have compiled a regular expression, as described in Section 16.3.1 [POSIX Regular Expression Compilation], page 327, you can match it against strings using regexec. A match anywhere inside the string counts as success, unless the regular expression contains anchor characters (‘^’ or ‘$’).

Function int regexec (regex_t *compiled, char *string, size_t

nmatch, regmatch_t matchptr [], int eflags) This function tries to match the compiled regular expression *compiled against string. regexec returns 0 if the regular expression matches; otherwise, it returns a nonzero value. See the table below for what nonzero values mean. You can use regerror to produce an error message string describing the reason for a nonzero value; see Section 16.3.6 [POSIX Regexp Matching Cleanup], page 332. The argument eflags is a word of bit flags that enable various options. If you want to get information about what part of string actually matched the regular expression or its subexpressions, use the arguments matchptr and nmatch. Otherwise, pass 0 for nmatch, and NULL for matchptr. See Section 16.3.4 [Match Results with Subexpressions], page 331.

You must match the regular expression with the same set of current locales that were in effect when you compiled the regular expression.

The function regexec accepts the following flags in the eflags argument:

REG_NOTBOL

Do not regard the beginning of the specified string as the beginning of a line; more generally, don’t make any assumptions about what text might precede it.

REG_NOTEOL

Do not regard the end of the specified string as the end of a line; more generally, don’t make any assumptions about what text might follow it.

Here are the possible nonzero values that regexec can return:

REG_NOMATCH

The pattern didn’t match the string. This isn’t really an error.

REG_ESPACE

regexec ran out of memory.

16.3.4 Match Results with Subexpressions

When regexec matches parenthetical subexpressions of pattern, it records which parts of string they match. It returns that information by storing the offsets into an array whose elements are structures of type regmatch_t. The first element of the array (index 0) records the part of the string that matched the entire regular expression. Each other element of the array records the beginning and end of the part that matched a single parenthetical subexpression.

Data Type regmatch t

This is the data type of the matcharray array that you pass to regexec. It contains two structure fields, as follows:

rm_so The offset in string of the beginning of a substring. Add this value to string to get the address of that part.

rm_eo The offset in string of the end of the substring.

Data Type regoff t

regoff_t is an alias for another signed integer type. The fields of

  regmatch_t have type regoff_t.

The regmatch_t elements correspond to subexpressions positionally; the first element (index 1) records where the first subexpression matched, the second element records the second subexpression, and so on. The order of the subexpressions is the order in which they begin.

When you call regexec, you specify how long the matchptr array is, with the nmatch argument. This tells regexec how many elements to store. If the actual regular expression has more than nmatch subexpressions, then you won’t get offset information about the rest of them. But this doesn’t alter whether the pattern matches a particular string or not.

If you don’t want regexec to return any information about where the subexpressions matched, you can either supply 0 for nmatch, or use the flag REG_NOSUB when you compile the pattern with regcomp.

16.3.5 Complications in Subexpression Matching

Sometimes a subexpression matches a substring of no characters. This happens when ‘f\(o*\)’ matches the string ‘fum’. (It really matches just the ‘f’.) In this case, both of the offsets identify the point in the string where the null substring was found. In this example, the offsets are both 1.

Sometimes the entire regular expression can match without using some of its subexpressions at all—for example, when ‘ba\(na\)*’ matches the string ‘ba’, the parenthetical subexpression is not used. When this happens, regexec stores -1 in both fields of the element for that subexpression.

Sometimes matching the entire regular expression can match a particular subexpression more than once—for example, when ‘ba\(na\)*’ matches the string ‘bananana’, the parenthetical subexpression matches three times. When this happens, regexec usually stores the offsets of the last part of the string that matched the subexpression. In the case of ‘bananana’, these offsets are 6 and 8.

But the last match is not always the one that is chosen. It’s more accurate to say that the last opportunity to match is the one that takes precedence. What this means is that when one subexpression appears within another, then the results reported for the inner subexpression reflect whatever happened on the last match of the outer subexpression. For an example, consider ‘\(ba\(na\)*s \)*’ matching the string ‘bananas bas ’. The last time the inner expression actually matches is near the end of the first word. But it is considered again in the second word, and fails to match there. regexec reports nonuse of the “na” subexpression.

Another place where this rule applies is when the regular expression

‘\(ba\(na\)*s \|nefer\(ti\)* \)*’ matches ‘bananas nefertiti’. The

“na” subexpression does match in the first word, but it doesn’t match in the second word because the other alternative is used there. Once again, the second repetition of the outer subexpression overrides the first, and within that second repetition, the “na” subexpression is not used. So regexec reports nonuse of the “na” subexpression.

16.3.6 POSIX Regexp Matching Cleanup

When you are finished using a compiled regular expression, you can free the storage it uses by calling regfree.

Function void regfree (regex_t *compiled)

Calling regfree frees all the storage that *compiled points to. This includes various internal fields of the regex_t structure that aren’t documented in this manual. regfree does not free the object *compiled itself.

You should always free the space in a regex_t structure with regfree before using the structure to compile another regular expression.

When regcomp or regexec reports an error, you can use the function regerror to turn it into an error message string.

Function size_t regerror (int errcode, regex_t *compiled,

      char *buffer, size_t length)

This function produces an error message string for the error code errcode, and stores the string in length bytes of memory starting at buffer. For the compiled argument, supply the same compiled regular expression structure that regcomp or regexec was working with when it got the error. Alternatively, you can supply NULL for compiled; you will still get a meaningful error message, but it might not be as detailed. If the error message can’t fit in length bytes (including a terminating null character), then regerror truncates it. The string that regerror stores is always null-terminated even if it has been truncated. The return value of regerror is the minimum length needed to store the entire error message. If this is less than length, then the error message was not truncated, and you can use it. Otherwise, you should call regerror again with a larger buffer. Here is a function which uses regerror, but always dynamically allocates a buffer for the error message:

     char *get_regerror (int errcode, regex_t *compiled)
     {
       size_t length = regerror (errcode, compiled, NULL, 0);
       char *buffer = xmalloc (length);
       (void) regerror (errcode, compiled, buffer, length);
       return buffer;
     }

16.4 Shell-Style Word Expansion

Word expansion means the process of splitting a string into words and substituting for variables, commands, and wildcards just as the shell does.

For example, when you write ‘ls -l foo.c’, this string is split into three separate words—‘ls’, ‘-l’ and ‘foo.c’. This is the most basic function of word expansion.

When you write ‘ls *.c’, this can become many words, because the word ‘*.c’ can be replaced with any number of file names. This is called wildcard expansion, and it is also a part of word expansion.

When you use ‘echo $PATH’ to print your path, you are taking advantage of variable substitution, which is also part of word expansion.

Ordinary programs can perform word expansion just like the shell by calling the library function wordexp.

16.4.1 The Stages of Word Expansion

When word expansion is applied to a sequence of words, it performs the following transformations in the order shown here:

1. Tilde expansion: Replacement of ‘~foo’ with the name of the home directory of ‘foo’.

2. Next, three different transformations are applied in the same step, from

left to right:

• Variable substitution: Environment variables are substituted for references such as ‘$foo’.

• Command substitution: Constructs such as ‘`cat foo`’ and the equivalent ‘$(cat foo)’ are replaced with the output from the inner command.

• Arithmetic expansion: Constructs such as ‘$(($x-1))’ are replaced

with the result of the arithmetic computation.

3. Field splitting: subdivision of the text into words.

4. Wildcard expansion: The replacement of a construct such as ‘*.c’ with a list of ‘.c’ file names. Wildcard expansion applies to an entire word at a time, and replaces that word with 0 or more file names that are themselves words.

5. Quote removal: The deletion of string-quotes, now that they have done their job by inhibiting the above transformations when appropriate.

For the details of these transformations, and how to write the constructs that use them, see The BASH Manual (to appear).

16.4.2 Calling wordexp

All the functions, constants and data types for word expansion are declared in the header file ‘wordexp.h’.

Word expansion produces a vector of words (strings). To return this vector, wordexp uses a special data type, wordexp_t, which is a structure. You pass wordexp the address of the structure, and it fills in the structure’s fields to tell you about the results.

Data Type wordexp t

This data type holds a pointer to a word vector. More precisely, it records both the address of the word vector and its size. we_wordc The number of elements in the vector. we_wordv The address of the vector. This field has type char **. we_offs The offset of the first real element of the vector, from its

nominal address in the we_wordv field. Unlike the other fields, this is always an input to wordexp, rather than an output from it. If you use a nonzero offset, then that many elements at the beginning of the vector are left empty. (The wordexp function fills them with null pointers.) The we_offs field is meaningful only if you use the WRDE_ DOOFFS flag. Otherwise, the offset is always zero regardless of what is in this field, and the first real element comes at the beginning of the vector.

Function int wordexp (const char *words, wordexp_t

*word-vector-ptr, int flags) Perform word expansion on the string words, putting the result in a newly allocated vector, and store the size and address of this vector into *word-vector-ptr. The argument flags is a combination of bit flags; see Section 16.4.3 [Flags for Word Expansion], page 336, for details of the flags. You shouldn’t use any of the characters ‘|&;<>’ in the string words unless they are quoted; likewise for newline. If you use these characters unquoted, you will get the WRDE_BADCHAR error code. Don’t use parentheses or braces unless they are quoted or part of a word expansion construct. If you use quotation characters ‘'"`’, they should come in pairs that balance. The results of word expansion are a sequence of words. The function wordexp allocates a string for each resulting word, then allocates a vector of type char ** to store the addresses of these strings. The last element of the vector is a null pointer. This vector is called the word vector. To return this vector, wordexp stores both its address and its length (number of elements, not counting the terminating null pointer) into *word-vector-ptr. If wordexp succeeds, it returns 0. Otherwise, it returns one of these error codes:

WRDE_BADCHAR

The input string words contains an unquoted invalid character such as ‘|’.

WRDE_BADVAL

The input string refers to an undefined shell variable, and you used the flag WRDE_UNDEF to forbid such references.

WRDE_CMDSUB

The input string uses command substitution, and you used the flag WRDE_NOCMD to forbid command substitution.

WRDE_NOSPACE

It was impossible to allocate memory to hold the result. In this case, wordexp can store part of the results—as much as it could allocate room for.

WRDE_SYNTAX

There was a syntax error in the input string. For example, an unmatched quoting character is a syntax error.

Function void wordfree (wordexp_t *word-vector-ptr)

Free the storage used for the word-strings and vector that *wordvector-ptr points to. This does not free the structure *word-vector-ptr itself—only the other data it points to.

16.4.3 Flags for Word Expansion

This section describes the flags that you can specify in the flags argument to wordexp. Choose the flags you want, and combine them with the C operator |.

WRDE_APPEND

Append the words from this expansion to the vector of words produced by previous calls to wordexp. This way you can effectively expand several words as if they were concatenated with spaces between them. In order for appending to work, you must not modify the contents of the word vector structure between calls to wordexp. And, if you set WRDE_DOOFFS in the first call to wordexp, you must also set it when you append to the results.

WRDE_DOOFFS

Leave blank slots at the beginning of the vector of words. The we_offs field says how many slots to leave. The blank slots contain null pointers.

WRDE_NOCMD

Don’t do command substitution; if the input requests command substitution, report an error.

WRDE_REUSE

Reuse a word vector made by a previous call to wordexp. Instead of allocating a new vector of words, this call to wordexp will use the vector that already exists (making it larger if necessary). Note that the vector may move, so it is not safe to save an old pointer and use it again after calling wordexp. You must fetch we_pathv anew after each call.

WRDE_SHOWERR

Do show any error messages printed by commands run by command substitution. More precisely, allow these commands to inherit the standard error output stream of the current process. By default, wordexp gives these commands a standard error stream that discards all output.

WRDE_UNDEF

If the input refers to a shell variable that is not defined, report an error.

16.4.4 wordexp Example

Here is an example of using wordexp to expand several strings and use the results to run a shell command. It also shows the use of WRDE_APPEND to concatenate the expansions and of wordfree to free the space allocated

by wordexp.
int
expand_and_execute (const char *program, const char *options)
{
wordexp_t result;
pid_t pid
int status, i;

/* Expand the string for the program to run. */

switch (wordexp (program, &result, 0))
{
case 0:  /* Successful.  */
break;
case WRDE_NOSPACE:
/* If the error was WRDE_NOSPACE,

then perhaps part of the result was allocated. */

wordfree (&result);

default: /* Some other error. */

return -1;
}

/* Expand the strings specified for the arguments. */

for (i = 0; args[i]; i++)
  {
    if (wordexp (options, &result, WRDE_APPEND))
      {
        wordfree (&result);
        return -1;
      }
  }
pid = fork ();
if (pid == 0)
  {

/* This is the child process. Execute the command. */

    execv (result.we_wordv[0], result.we_wordv);
    exit (EXIT_FAILURE);
  }
else if (pid < 0)

/* The fork failed. Report failure. */

  status = -1;
else

/* This is the parent process. Wait for the child to complete. */

  if (waitpid (pid, &status, 0) != pid)
    status = -1;
wordfree (&result);
return status;
}

In practice, since wordexp is executed by running a subshell, it would be faster to do this by concatenating the strings with spaces between them and running that as a shell command using ‘sh -c’.

17 Date and Time

This chapter describes functions for manipulating dates and times, including functions for determining what the current time is and conversion between different time representations.

The time functions fall into three main categories:

• Functions for measuring elapsed CPU time are discussed in Section 17.1 [Processor Time], page 339.

• Functions for measuring absolute clock or calendar time are discussed in Section 17.2 [Calendar Time], page 341.

• Functions for setting alarms and timers are discussed in Section 17.3 [Setting an Alarm], page 356.

17.1 Processor Time

If you’re trying to optimize your program or measure its efficiency, it’s very useful to be able to know how much processor time or CPU time it has used at any given point. Processor time is different from actual wall clock time because it doesn’t include any time spent waiting for I/O or when some other process is running. Processor time is represented by the data type clock_t, and is given as a number of clock ticks relative to an arbitrary base time marking the beginning of a single program invocation.

17.1.1 Basic CPU Time Inquiry

To get the elapsed CPU time used by a process, you can use the clock function. This facility is declared in the header file ‘time.h’.

In typical usage, you call the clock function at the beginning and end of the interval you want to time, subtract the values, and then divide by CLOCKS_PER_SEC (the number of clock ticks per second), like this:

     #include <time.h>
     clock_t start, end;
     double elapsed;
     start = clock();

… /* Do the work. */

     end = clock();
     elapsed = ((double) (end - start)) / CLOCKS_PER_SEC;

Different computers and operating systems vary wildly in how they keep track of processor time. It’s common for the internal processor clock to have a resolution somewhere between hundredths and millionths of a second.

In the GNU system, clock_t is equivalent to long int and CLOCKS_ PER_SEC is an integer value. But in other systems, both clock_t and the type of the macro CLOCKS_PER_SEC can be either integer or floating-point types. Casting processor time values to double, as in the example above, makes sure that operations such as arithmetic and printing work properly and consistently no matter what the underlying representation is.

Macro int CLOCKS PER SEC

The value of this macro is the number of clock ticks per second measured by the clock function.

Macro int CLK TCK

This is an obsolete name for CLOCKS_PER_SEC.

Data Type clock t

This is the type of the value returned by the clock function. Values of type clock_t are in units of clock ticks.

Function clock_t clock (void)

This function returns the elapsed processor time. The base time is arbitrary but doesn’t change within a single process. If the processor time is not available or cannot be represented, clock returns the value

(clock_t)(-1).

17.1.2 Detailed Elapsed CPU Time Inquiry

The times function returns more detailed information about elapsed processor time in a struct tms object. You should include the header file ‘sys/times.h’ to use this facility.

Data Type struct tms

The tms structure is used to return information about process times. It contains at least the following members:

clock_t tms_utime

This is the CPU time used in executing the instructions of the calling process.

clock_t tms_stime

This is the CPU time used by the system on behalf of the calling process.

clock_t tms_cutime

This is the sum of the tms_utime values and the tms_ cutime values of all terminated child processes of the calling process, whose status has been reported to the parent process by wait or waitpid; see Section 23.6 [Process Completion], page 473. In other words, it represents the total CPU time used in executing the instructions of all the terminated child processes of the calling process, excluding child processes which have not yet been reported

          by wait or waitpid.
clock_t tms_cstime

This is similar to tms_cutime, but represents the total CPU time used by the system on behalf of all the terminated child processes of the calling process.

All of the times are given in clock ticks. These are absolute values; in a newly created process, they are all zero. See Section 23.4 [Creating a Process], page 469.

Function clock_t times (struct tms *buffer)

The times function stores the processor time information for the calling process in buffer. The return value is the same as the value of clock(): the elapsed real time relative to an arbitrary base. The base is a constant within a particular process, and typically represents the time since system start-up. A value of (clock_t)(-1) is returned to indicate failure.

Portability Note: The clock function described in Section 17.1.1 [Basic CPU Time Inquiry], page 339, is specified by the ISO C standard. The times function is a feature of POSIX.1. In the GNU system, the value returned by the clock function is equivalent to the sum of the tms_utime and tms_stime fields returned by times.

17.2 Calendar Time

This section describes facilities for keeping track of dates and times according to the Gregorian calendar.

There are three representations for date and time information:

• Calendar time (the time_t data type) is a compact representation, typically giving the number of seconds elapsed since some implementationspecific base time.

• There is also a high-resolution time representation (the struct timeval data type) that includes fractions of a second. Use this time representation instead of ordinary calendar time when you need greater precision.

• Local time or broken-down time (the struct tm data type) represents the date and time as a set of components specifying the year, month, and so on, for a specific time zone. This time representation is usually used in conjunction with formatting date and time values.

17.2.1 Simple Calendar Time

This section describes the time_t data type for representing calendar time, and the functions which operate on calendar time objects. These facilities are declared in the header file ‘time.h’.

Data Type time t

This is the data type used to represent calendar time. When interpreted as an absolute time value, it represents the number of seconds elapsed since 00:00:00 on January 1, 1970, Coordinated Universal Time. (This date is sometimes referred to as the epoch.) POSIX requires that this count ignore leap seconds, but on some hosts this count includes leap seconds if you set TZ to certain values (see Section 17.2.5 [Specifying the Time Zone with “codeTZ], page 352). In the GNU C library, time_t is equivalent to long int. In other systems, time_t might be either an integer or floating-point type.

Function double difftime (time_t time1, time_t time0)

The difftime function returns the number of seconds elapsed between time time1 and time time0, as a value of type double. The difference ignores leap seconds unless leap second support is enabled. In the GNU system, you can simply subtract time_t values. But on other systems, the time_t data type might use some other encoding where subtraction doesn’t work directly.

Function time_t time (time_t *result)

The time function returns the current time as a value of type time_t. If the argument result is not a null pointer, the time value is also stored in *result. If the calendar time is not available, the value

(time_t)(-1) is returned.

17.2.2 High-Resolution Calendar

The time_t data type used to represent calendar times has a resolution of only one second. Some applications need more precision.

So, the GNU C library also contains functions which are capable of representing calendar times to a higher resolution than one second. The functions and the associated data types described in this section are declared in

‘sys/time.h’.

Data Type struct timeval

The struct timeval structure represents a calendar time. It has the following members:

long int tv_sec

This represents the number of seconds since the epoch. It is equivalent to a normal time_t value.

long int tv_usec

This is the fractional second value, represented as the number of microseconds. Some times struct timeval values are used for time intervals. Then the tv_sec member is the number of seconds in the interval, and tv_usec is the number of additional microseconds.

Data Type struct timezone

The struct timezone structure is used to hold minimal information about the local time zone. It has the following members:

int tz_minuteswest

This is the number of minutes west of UTC.

int tz_dsttime

If nonzero, daylight saving time applies during some part of the year.

The struct timezone type is obsolete and should never be used. Instead, use the facilities described in Section 17.2.6 [Functions and Variables for Time Zones], page 354.

It is often necessary to subtract two values of type struct timeval. Here is the best way to do this. It works even on some peculiar operating systems where the tv_sec member has an unsigned type.

/* Subtract the ‘struct timeval’ values X and Y, storing the result in RESULT. Return 1 if the difference is negative, otherwise 0. */

int
timeval_subtract (result, x, y)
struct timeval *result, *x, *y;
{

/* Perform the carry for the later subtraction by updating y. */

if (x->tv_usec < y->tv_usec) {
int nsec = (y->tv_usec - x->tv_usec) / 1000000 + 1;
y->tv_usec -= 1000000 * nsec;
y->tv_sec += nsec;
}
if (x->tv_usec - y->tv_usec > 1000000) {
int nsec = (y->tv_usec - x->tv_usec) / 1000000;
y->tv_usec += 1000000 * nsec;
y->tv_sec -= nsec;
}

/* Compute the time remaining to wait.

tv_usec is certainly positive. */

    result->tv_sec = x->tv_sec - y->tv_sec;
    result->tv_usec = x->tv_usec - y->tv_usec;

/* Return 1 if result is negative. */

    return x->tv_sec < y->tv_sec;
  }
Function int gettimeofday (struct timeval *tp, struct
      timezone *tzp)

The gettimeofday function returns the current date and time in the struct timeval structure indicated by tp. Information about the time zone is returned in the structure pointed at tzp. If the tzp argument is a null pointer, time zone information is ignored. The return value is 0 on success and -1 on failure. The following errno error condition is defined for this function:

ENOSYS The operating system does not support getting time zone information, and tzp is not a null pointer. The GNU operating system does not support using struct timezone to represent time zone information; that is an obsolete feature of 4.3 BSD. Instead, use the facilities described in Section 17.2.6 [Functions and Variables for Time Zones], page 354.

Function int settimeofday (const struct timeval *tp, const
      struct timezone *tzp)

The settimeofday function sets the current date and time according to the arguments. As for gettimeofday, time zone information is ignored if tzp is a null pointer. You must be a privileged user in order to use settimeofday. The return value is 0 on success and -1 on failure. The following errno error conditions are defined for this function:

EPERM This process cannot set the time because it is not privileged.

ENOSYS The operating system does not support setting time zone information, and tzp is not a null pointer.

Function int adjtime (const struct timeval *delta, struct

timeval *olddelta) This function speeds up or slows down the system clock in order to make gradual adjustments in the current time. This ensures that the time reported by the system clock is always monotonically increasing, which might not happen if you simply set the current time.

The delta argument specifies a relative adjustment to be made to the current time. If negative, the system clock is slowed down for a while until it has lost this much time. If positive, the system clock is speeded up for a while. If the olddelta argument is not a null pointer, the adjtime function returns information about any previous time adjustment that has not yet completed. This function is typically used to synchronize the clocks of computers in a local network. You must be a privileged user to use it. The return value is 0 on success and -1 on failure. The following errno error condition is defined for this function:

EPERM You do not have privilege to set the time.

Portability Note: The gettimeofday, settimeofday, and adjtime func-

tions are derived from BSD.

17.2.3 Broken-down Time

Calendar time is represented as a number of seconds. This is convenient for calculation, but has no resemblance to the way people normally represent dates and times. By contrast, broken-down time is a binary representation separated into year, month, day, and so on. Broken down time values are not useful for calculations, but they are useful for printing human readable time.

A broken-down time value is always relative to a choice of local time zone, and it also indicates which time zone was used.

The symbols in this section are declared in the header file ‘time.h’.

Data Type struct tm

This is the data type used to represent a broken-down time. The structure contains at least the following members, which can appear in any order:

int tm_sec

This is the number of seconds after the minute, normally in the range 0 through 59. (The actual upper limit is 60, to allow for leap seconds if leap second support is available.)

int tm_min

This is the number of minutes after the hour, in the range 0 through 59.

int tm_hour

This is the number of hours past midnight, in the range 0 through 23.

int tm_mday

This is the day of the month, in the range 1 through 31.

int tm_mon

This is the number of months since January, in the range 0 through 11.

int tm_year

This is the number of years since 1900.

int tm_wday

This is the number of days since Sunday, in the range 0 through 6.

int tm_yday

This is the number of days since January 1, in the range 0 through 365.

int tm_isdst

This is a flag that indicates whether Daylight Saving Time is (or was, or will be) in effect at the time described. The value is positive if Daylight Saving Time is in effect, zero if it is not, and negative if the information is not available.

long int tm_gmtoff

This field describes the time zone that was used to compute this broken-down time value, including any adjustment for daylight saving; it is the number of seconds that you must add to UTC to get local time. You can also think of this as the number of seconds east of UTC. For example, for U.S. Eastern Standard Time, the value is - 5*60*60. The tm_gmtoff field is derived from BSD and is a GNU library extension; it is not visible in a strict ISO C environment.

const char *tm_zone

This field is the name for the time zone that was used to compute this broken-down time value. Like tm_gmtoff, this field is a BSD and GNU extension, and is not visible in a strict ISO C environment.

Function struct tm * localtime (const time_t *time)

The localtime function converts the calendar time pointed to by time to broken-down time representation, expressed relative to the user’s specified time zone. The return value is a pointer to a static broken-down time structure, which might be overwritten by subsequent calls to ctime, gmtime, or localtime. (But no other library function overwrites the contents of this object.)

Calling localtime has one other effect: it sets the variable tzname with information about the current time zone. See Section 17.2.6 [Functions and Variables for Time Zones], page 354.

Function struct tm * gmtime (const time_t *time)

This function is similar to localtime, except that the broken-down time is expressed as Coordinated Universal Time (UTC)—that is, as Greenwich Mean Time (GMT)—rather than relative to the local time zone. Recall that calendar times are always expressed in coordinated universal time.

Function time_t mktime (struct tm *brokentime)

The mktime function is used to convert a broken-down time structure to a calendar time representation. It also “normalizes” the contents of the broken-down time structure, by filling in the day of week and day of year based on the other date and time components. The mktime function ignores the specified contents of the tm_wday and tm_yday members of the broken-down time structure. It uses the values of the other components to compute the calendar time; it’s permissible for these components to have unnormalized values outside of their normal ranges. The last thing that mktime does is adjust the components of the brokentime structure (including the tm_wday and

tm_yday).

If the specified broken-down time cannot be represented as a calendar time, mktime returns a value of (time_t)(-1) and does not modify the contents of brokentime. Calling mktime also sets the variable tzname with information about the current time zone. See Section 17.2.6 [Functions and Variables for Time Zones], page 354.

17.2.4 Formatting Date and Time

The functions described in this section format time values as strings. These functions are declared in the header file ‘time.h’.

Function char * asctime (const struct tm *brokentime)

The asctime function converts the broken-down time value that brokentime points to into a string in a standard format:

     "Tue May 21 13:46:22 1991\n"

The abbreviations for the days of week are: ‘Sun’, ‘Mon’, ‘Tue’, ‘Wed’, ‘Thu’, ‘Fri’, and ‘Sat’. The abbreviations for the months are: ‘Jan’, ‘Feb’, ‘Mar’, ‘Apr’, ‘May’, ‘Jun’, ‘Jul’, ‘Aug’, ‘Sep’, ‘Oct’, ‘Nov’, and ‘Dec’.

The return value points to a statically allocated string, which might be overwritten by subsequent calls to asctime or ctime. (But no other library function overwrites the contents of this string.)

Function char * ctime (const time_t *time)

The ctime function is similar to asctime, except that the time value is specified as a time_t calendar time value rather than in broken-down local time format. It is equivalent to

     asctime (localtime (time))

ctime sets the variable tzname, because localtime does so. See Section 17.2.6 [Functions and Variables for Time Zones], page 354.

Function size_t strftime (char *s, size_t size, const char

*template, const struct tm *brokentime) This function is similar to the sprintf function (see Section 7.12 [Formatted Input], page 126), but the conversion specifications that can appear in the format template template are specialized for printing components of the date and time brokentime according to the locale currently specified for time conversion (see Chapter 19 [Locales and Internationalization], page 379). Ordinary characters appearing in the template are copied to the output string s; this can include multibyte character sequences. Conversion specifiers are introduced by a ‘%’ character, followed by an optional flag which can be one of the following. These flags, which are GNU extensions, affect only the output of numbers: _The number is padded with spaces.

- The number is not padded at all. 0 The number is padded with zeros even if the format spec-

ifies padding with spaces.

^The output uses uppercase characters, but only if this is possible (see Section 4.2 [Case Conversion], page 63).

The default action is to pad the number with zeros to keep it a constant width. Numbers that do not have a range indicated below are never padded, since there is no natural width for them. Following the flag an optional specification of the width is possible. This is specified in decimal notation. If the natural size of the output is of the field has less than the specified number of characters, the result is written right adjusted and space padded to the given size. An optional modifier can follow the optional flag and width specification. The modifiers, which are POSIX.2 extensions, are: E Use the locale’s alternate representation for date and time.

This modifier applies to the %c, %C, %x, %X, %y and %Y format specifiers. In a Japanese locale, for example, %Ex might yield a date format based on the Japanese Emperors’ reigns.

O Use the locale’s alternate numeric symbols for numbers. This modifier applies only to numeric format specifiers.

If the format supports the modifier but no alternate representation is available, it is ignored. The conversion specifier ends with a format specifier taken from the following list. The whole ‘%’ sequence is replaced in the output string as follows:

%a The abbreviated weekday name according to the current locale.

%A The full weekday name according to the current locale.

%b The abbreviated month name according to the current locale.

%B The full month name according to the current locale.

%c The preferred date and time representation for the current locale.

%C The century of the year. This is equivalent to the greatest integer not greater than the year divided by 100. This format is a POSIX.2 extension.

%d The day of the month as a decimal number (range 01 through 31).

%D The date using the format %m/%d/%y.

This format is a POSIX.2 extension.

%e The day of the month like with %d, but padded with blank (range 1 through 31). This format is a POSIX.2 extension.

%g The year corresponding to the ISO week number, but without the century (range 00 through 99). This has the same format and value as %y, except that if the ISO week number (see %V) belongs to the previous or next year, that year is used instead. This format is a GNU extension.

%G The year corresponding to the ISO week number. This has the same format and value as %Y, except that if the ISO week number (see %V) belongs to the previous or next year, that year is used instead. This format is a GNU extension.

%h The abbreviated month name according to the current locale. The action is the same as for %b. This format is a POSIX.2 extension.

%H The hour as a decimal number, using a 24-hour clock (range 00 through 23).

%I The hour as a decimal number, using a 12-hour clock (range 01 through 12).

%j The day of the year as a decimal number (range 001 through 366).

%k The hour as a decimal number, using a 24-hour clock like %H, but padded with blank (range 0 through 23). This format is a GNU extension.

%l The hour as a decimal number, using a 12-hour clock like %I, but padded with blank (range 1 through 12). This format is a GNU extension.

%m The month as a decimal number (range 01 through 12).

%M The minute as a decimal number (range 00 through 59).

%n A single ‘\n’ (newline) character.

This format is a POSIX.2 extension.

%p Either ‘AM’ or ‘PM’, according to the given time value; or the corresponding strings for the current locale. Noon is treated as ‘PM’ and midnight as ‘AM’.

%P Either ‘am’ or ‘pm’, according to the given time value; or the corresponding strings for the current locale, printed in lowercase characters. Noon is treated as ‘pm’ and midnight as ‘am’. This format is a GNU extension.

%r The complete time using the AM/PM format of the current locale. This format is a POSIX.2 extension.

%R The hour and minute in decimal numbers using the format

%H:%M.

This format is a GNU extension.

%s The number of seconds since the epoch, i.e., since 1970-01-01 00:00:00 UTC. Leap seconds are not counted unless leap second support is available. This format is a GNU extension.

%S The second as a decimal number (range 00 through 60).

%t A single ‘\t’ (tabulator) character.

This format is a POSIX.2 extension.

%T The time using decimal numbers using the format

%H:%M:%S.

This format is a POSIX.2 extension.

%u The day of the week as a decimal number (range 1 through 7), Monday being 1. This format is a POSIX.2 extension.

%U The week number of the current year as a decimal number (range 00 through 53), starting with the first Sunday as the first day of the first week. Days preceding the first Sunday in the year are considered to be in week 00.

%V The ISO 8601:1988 week number as a decimal number (range 01 through 53). ISO weeks start with Monday and end with Sunday. Week 01 of a year is the first week which has the majority of its days in that year; this is equivalent to the week containing the year’s first Thursday, and it is also equivalent to the week containing January 4. Week 01 of a year can contain days from the previous year. The week before week 01 of a year is the last week (52 or 53) of the previous year even if it contains days from the new year. This format is a POSIX.2 extension.

%w The day of the week as a decimal number (range 0 through 6), Sunday being 0.

%W The week number of the current year as a decimal number (range 00 through 53), starting with the first Monday as the first day of the first week. All days preceding the first Monday in the year are considered to be in week 00.

%x The preferred date representation for the current locale, but without the time.

%X The preferred time representation for the current locale, but with no date.

%y The year without a century as a decimal number (range 00 through 99). This is equivalent to the year modulo 100.

%Y The year as a decimal number, using the Gregorian calendar. Years before the year 1 are numbered 0, -1, and so on.

%z RFC 822/ISO 8601:1988 style numeric time zone (e.g., - 0600 or +0100), or nothing if no time zone is determinable. This format is a GNU extension.

%Z The time zone abbreviation (empty if the time zone can’t be determined).

%% A literal ‘%’ character.

The size parameter can be used to specify the maximum number of characters to be stored in the array s, including the terminating null character. If the formatted time requires more than size characters, the excess characters are discarded. The return value from strftime is the number of characters placed in the array s, not including the terminating null character. If the value equals size, it means that the array s was too small; you should repeat the call, providing a bigger array. If s is a null pointer, strftime does not actually write anything, but instead returns the number of characters it would have written. According to POSIX.1 every call to strftime implies a call to tzset. So the contents of the environment variable TZ is examined before any output is produced. For an example of strftime, see Section 17.2.7 [Time Functions Example], page 355.

17.2.5 Specifying the Time Zone with TZ

In POSIX systems, a user can specify the time zone by means of the TZ environment variable. For information about how to set environment variables, see Section 22.2 [Environment Variables], page 458. The functions for accessing the time zone are declared in ‘time.h’.

You should not normally need to set TZ. If the system is configured properly, the default time zone will be correct. You might set TZ if you are using a computer over the network from a different time zone, and would like times reported to you in the time zone that local for you, rather than what is local for the computer.

In POSIX.1 systems the value of the TZ variable can be of one of three formats. With the GNU C library, the most common format is the last one, which can specify a selection from a large database of time zone information for many regions of the world. The first two formats are used to describe the time zone information directly, which is both more cumbersome and less precise. But the POSIX.1 standard only specifies the details of the first two formats, so it is good to be familiar with them in case you come across a POSIX.1 system that doesn’t support a time zone information database.

The first format is used when there is no Daylight Saving Time (or summer time) in the local time zone:

std offset

The std string specifies the name of the time zone. It must be three or more characters long and must not contain a leading colon or embedded digits, commas, or plus or minus signs. There is no space character separating the time zone name from the offset, so these restrictions are necessary to parse the specification correctly.

The offset specifies the time value one must add to the local time to get a Coordinated Universal Time value. It has syntax like [+|-]hh[:mm[:ss]]. This is positive if the local time zone is west of the Prime Meridian and negative if it is east. The hour must be between 0 and 23, and the minute and seconds between 0 and 59.

For example, here is how we would specify Eastern Standard Time, but without any daylight saving time alternative:

     EST+5

The second format is used when there is Daylight Saving Time:

std offset dst [offset],start[/time],end[/time]

The initial std and offset specify the standard time zone, as described above. The dst string and offset specify the name and offset for the corresponding daylight saving time time zone; if the offset is omitted, it defaults to one hour ahead of standard time.

The remainder of the specification describes when daylight saving time is in effect. The start field is when daylight saving time goes into effect and the end field is when the change is made back to standard time. The following formats are recognized for these fields:

Jn This specifies the Julian day, with n between 1 and 365. February 29 is never counted, even in leap years.

n This specifies the Julian day, with n between 0 and 365. February 29 is counted in leap years.

Mm.w.d This specifies day d of week w of month m. The day d must be between 0 (Sunday) and 6. The week w must be between 1 and 5; week 1 is the first week in which day d occurs, and week 5 specifies the last d day in the month. The month m should be between 1 and 12.

The time fields specify when, in the local time currently in effect, the change to the other time occurs. If omitted, the default is 02:00:00.

For example, here is how one would specify the Eastern time zone in the United States, including the appropriate daylight saving time and its dates of applicability. The normal offset from UTC is 5 hours; since this is west of the prime meridian, the sign is positive. Summer time begins on the first Sunday in April at 2:00am, and ends on the last Sunday in October at 2:00am.

     EST+5EDT,M4.1.0/2,M10.5.0/2

The schedule of daylight saving time in any particular jurisdiction has changed over the years. To be strictly correct, the conversion of dates and times in the past should be based on the schedule that was in effect then. However, this format has no facilities to let you specify how the schedule has changed from year to year. The most you can do is specify one particular schedule—usually the present day schedule—and this is used to convert any date, no matter when. For precise time zone specifications, it is best to use the time zone information database (see below).

The third format looks like this:

:characters

Each operating system interprets this format differently; in the GNU C library, characters is the name of a file which describes the time zone.

If the TZ environment variable does not have a value, the operation chooses a time zone by default. In the GNU C library, the default time zone is like the specification ‘TZ=:/etc/localtime’ (or ‘TZ=:/usr/local/etc/localtime’, depending on how GNU C library was configured; see Section C.1 [How to Install the GNU C Library], page 667). Other C libraries use their own rule for choosing the default time zone, so there is little we can say about them.

If characters begins with a slash, it is an absolute file name; otherwise the library looks for the file ‘/share/lib/zoneinfo/characters’. The ‘zoneinfo’ directory contains data files describing local time zones in many different parts of the world. The names represent major cities, with subdirectories for geographical areas; for example, ‘America/New_York’, ‘Europe/London’, ‘Asia/Hong_Kong’. These data files are installed by the system administrator, who also sets ‘/etc/localtime’ to point to the data file for the local time zone. The GNU C library comes with a large database of time zone information for most regions of the world, which is maintained by a community of volunteers and put in the public domain.

17.2.6 Functions and Variables for Time Zones

Variable char * tzname [2]

The array tzname contains two strings, which are the standard names of the pair of time zones (standard and daylight saving) that the user has selected. tzname[0] is the name of the standard time zone (for example, "EST"), and tzname[1] is the name for the time zone when daylight saving time is in use (for example, "EDT"). These correspond to the std and dst strings (respectively) from the TZ environment variable. If daylight saving time is never used, tzname[1] is the empty string. The tzname array is initialized from the TZ environment variable when-

ever tzset, ctime, strftime, mktime, or localtime is called. If mul-

tiple abbreviations have been used (e.g. "EWT" and "EDT" for U.S. Eastern War Time and Eastern Daylight Time), the array contains the most recent abbreviation. The tzname array is required for POSIX.1 compatibility, but in GNU programs it is better to use the tm_zone member of the broken-down time structure, since tm_zone reports the correct abbreviation even when it is not the latest one.

Function void tzset (void)

The tzset function initializes the tzname variable from the value of the TZ environment variable. It is not usually necessary for your program to call this function, because it is called automatically when you use the other time conversion functions that depend on the time zone.

The following variables are defined for compatibility with System V Unix. Like tzname, these variables are set by calling tzset or the other time conversion functions.

Variable long int timezone

This contains the difference between UTC and the latest local standard time, in seconds west of UTC. For example, in the U.S. Eastern time zone, the value is 5*60*60. Unlike the tm_gmtoff member of the broken-down time structure, this value is not adjusted for daylight saving, and its sign is reversed. In GNU programs it is better to use tm_gmtoff, since it contains the correct offset even when it is not the latest one.

Variable int daylight

This variable has a nonzero value if daylight savings time rules apply. A nonzero value does not necessarily mean that daylight savings time is now in effect; it means only that daylight savings time is sometimes in effect.

17.2.7 Time Functions Example

Here is an example program showing the use of some of the local time and calendar time functions.

  #include <time.h>
  #include <stdio.h>
  #define SIZE 256

int
main (void)
{
  char buffer[SIZE];
  time_t curtime;
  struct tm *loctime;

/* Get the current time. */

  curtime = time (NULL);

/* Convert it to local time representation. */

  loctime = localtime (&curtime);

/* Print out the date and time in the standard format. */

  fputs (asctime (loctime), stdout);

/* Print it out in a nice format. */

  strftime (buffer, SIZE, "Today is %A, %B %d.\n", loctime);
  fputs (buffer, stdout);
  strftime (buffer, SIZE, "The time is %I:%M %p.\n", loctime);
  fputs (buffer, stdout);
  return 0;
}

It produces output like this:

Wed Jul 31 13:02:36 1991
Today is Wednesday, July 31.
The time is 01:02 PM.

17.3 Setting an Alarm

The alarm and setitimer functions provide a mechanism for a process to interrupt itself at some future time. They do this by setting a timer; when the timer expires, the process receives a signal.

Each process has three independent interval timers available:

• A real-time timer that counts clock time. This timer sends a SIGALRM signal to the process when it expires.

• A virtual timer that counts CPU time used by the process. This timer sends a SIGVTALRM signal to the process when it expires.

• A profiling timer that counts both CPU time used by the process, and CPU time spent in system calls on behalf of the process. This timer sends a SIGPROF signal to the process when it expires.

This timer is useful for profiling in interpreters. The interval timer mechanism does not have the fine granularity necessary for profiling native code. You can only have one timer of each kind set at any given time. If you

set a timer that has not yet expired, that timer is simply reset to the new value.

You should establish a handler for the appropriate alarm signal using signal or sigaction before issuing a call to setitimer or alarm. Otherwise, an unusual chain of events could cause the timer to expire before your program establishes the handler, and in that case it would be terminated, since that is the default action for the alarm signals. See Chapter 21 [Signal Handling], page 393.

The setitimer function is the primary means for setting an alarm. This facility is declared in the header file ‘sys/time.h’. The alarm function, declared in ‘unistd.h’, provides a somewhat simpler interface for setting the real-time timer.

Data Type struct itimerval

This structure is used to specify when a timer should expire. It contains the following members:

struct timeval it_interval

This is the interval between successive timer interrupts. If zero, the alarm will only be sent once.

struct timeval it_value

This is the interval to the first timer interrupt. If zero, the alarm is disabled.

The struct timeval data type is described in Section 17.2.2 [High-Resolution Calendar], page 342.

Function int setitimer (int which, struct itimerval *new,

      struct itimerval *old)

The setitimer function sets the timer specified by which according to new. The which argument can have a value of ITIMER_REAL, ITIMER_

VIRTUAL, or ITIMER_PROF.

If old is not a null pointer, setitimer returns information about any previous unexpired timer of the same kind in the structure it points to. The return value is 0 on success and -1 on failure. The following errno error conditions are defined for this function: EINVAL The timer interval was too large.

Function int getitimer (int which, struct itimerval *old)

The getitimer function stores information about the timer specified by which in the structure pointed at by old.

The return value and error conditions are the same as for setitimer.

ITIMER_REAL

This constant can be used as the which argument to the setitimer and getitimer functions to specify the real-time timer.

ITIMER_VIRTUAL

This constant can be used as the which argument to the setitimer and getitimer functions to specify the virtual timer.

ITIMER_PROF

This constant can be used as the which argument to the setitimer and getitimer functions to specify the profiling timer.

Function unsigned int alarm (unsigned int seconds)

The alarm function sets the real-time timer to expire in seconds seconds. If you want to cancel any existing alarm, you can do this by calling alarm with a seconds argument of zero. The return value indicates how many seconds remain before the previous alarm would have been sent. If there is no previous alarm, alarm returns zero.

The alarm function could be defined in terms of setitimer like this:

     unsigned int
     alarm (unsigned int seconds)
     {
       struct itimerval old, new;
       new.it_interval.tv_usec = 0;
       new.it_interval.tv_sec = 0;
       new.it_value.tv_usec = 0;
       new.it_value.tv_sec = (long int) seconds;
       if (setitimer (ITIMER_REAL, &new, &old) < 0)
         return 0;
       else
         return old.it_value.tv_sec;
     }

There is an example showing the use of the alarm function in Section 21.4.1 [Signal Handlers that Return], page 414.

If you simply want your process to wait for a given number of seconds, you should use the sleep function. See Section 17.4 [Sleeping], page 359.

You shouldn’t count on the signal arriving precisely when the timer expires. In a multiprocessing environment there is typically some amount of delay involved.

Portability Note: The setitimer and getitimer functions are derived from BSD Unix, while the alarm function is specified by the POSIX.1 standard. setitimer is more powerful than alarm, but alarm is more widely used.

17.4 Sleeping

The function sleep gives a simple way to make the program wait for short periods of time. If your program doesn’t use signals (except to terminate), then you can expect sleep to wait reliably for the specified amount of time. Otherwise, sleep can return sooner if a signal arrives; if you want to wait for a given period regardless of signals, use select (see Section 8.6 [Waiting for Input or Output], page 162) and don’t specify any descriptors to wait for.

Function unsigned int sleep (unsigned int seconds)

The sleep function waits for seconds or until a signal is delivered, whichever happens first. If sleep function returns because the requested time has elapsed, it returns a value of zero. If it returns because of delivery of a signal, its return value is the remaining time in the sleep period. The sleep function is declared in ‘unistd.h’.

Resist the temptation to implement a sleep for a fixed amount of time by using the return value of sleep, when nonzero, to call sleep again. This will work with a certain amount of accuracy as long as signals arrive infrequently. But each signal can cause the eventual wakeup time to be off by an additional second or so. Suppose a few signals happen to arrive in rapid succession by bad luck—there is no limit on how much this could shorten or lengthen the wait.

Instead, compute the time at which the program should stop waiting, and keep trying to wait until that time. This won’t be off by more than a second. With just a little more work, you can use select and make the waiting period quite accurate. (Of course, heavy system load can cause unavoidable additional delays—unless the machine is dedicated to one application, there is no way you can avoid this.)

On some systems, sleep can do strange things if your program uses SIGALRM explicitly. Even if SIGALRM signals are being ignored or blocked when sleep is called, sleep might return prematurely on delivery of a SIGALRM signal. If you have established a handler for SIGALRM signals and a SIGALRM signal is delivered while the process is sleeping, the action taken might be just to cause sleep to return instead of invoking your handler. And, if sleep is interrupted by delivery of a signal whose handler requests an alarm or alters the handling of SIGALRM, this handler and sleep will interfere.

On the GNU system, it is safe to use sleep and SIGALRM in the same program, because sleep does not work by means of SIGALRM.

17.5 Resource Usage

The function getrusage and the data type struct rusage are used for examining the usage figures of a process. They are declared in

‘sys/resource.h’.

Function int getrusage (int processes, struct rusage *rusage)

This function reports the usage totals for processes specified by processes, storing the information in *rusage. In most systems, processes has only two valid values:

RUSAGE_SELF

Just the current process.

RUSAGE_CHILDREN

All child processes (direct and indirect) that have terminated already.

In the GNU system, you can also inquire about a particular child process by specifying its process ID. The return value of getrusage is zero for success, and -1 for failure. EINVAL The argument processes is not valid.

One way of getting usage figures for a particular child process is with the function wait4, which returns totals for a child when it terminates. See Section 23.8 [BSD Process Wait Functions], page 476.

Data Type struct rusage

This data type records a collection usage amounts for various sorts of resources. It has the following members, and possibly others:

struct timeval ru_utime

Time spent executing user instructions.

struct timeval ru_stime

Time spent in operating system code on behalf of processes.

long int ru_maxrss

The maximum resident set size used, in kilobytes. That is, the maximum number of kilobytes that processes used in real memory simultaneously.

long int ru_ixrss

An integral value expressed in kilobytes times ticks of execution, which indicates the amount of memory used by text that was shared with other processes.

long int ru_idrss

An integral value expressed the same way, which is the amount of unshared memory used in data.

long int ru_isrss

An integral value expressed the same way, which is the amount of unshared memory used in stack space.

long int ru_minflt

The number of page faults which were serviced without requiring any I/O.

long int ru_majflt

The number of page faults which were serviced by doing I/O.

long int ru_nswap

The number of times processes was swapped entirely out of main memory.

long int ru_inblock

The number of times the file system had to read from the disk on behalf of processes.

long int ru_oublock

The number of times the file system had to write to the disk on behalf of processes.

long int ru_msgsnd

Number of IPC messages sent.

long ru_msgrcv

Number of IPC messages received.

long int ru_nsignals

Number of signals received.

long int ru_nvcsw

The number of times processes voluntarily invoked a context switch (usually to wait for some service).

long int ru_nivcsw

The number of times an involuntary context switch took place (because the time slice expired, or another process of higher priority became runnable).

An additional historical function for examining usage figures, vtimes, is supported but not documented here. It is declared in ‘sys/vtimes.h’.

17.6 Limiting Resource Usage

You can specify limits for the resource usage of a process. When the process tries to exceed a limit, it may get a signal, or the system call by which it tried to do so may fail, depending on the limit. Each process initially inherits its limit values from its parent, but it can subsequently change them.

The symbols in this section are defined in ‘sys/resource.h’.

Function int getrlimit (int resource, struct rlimit *rlp)

Read the current value and the maximum value of resource resource and store them in *rlp. The return value is 0 on success and -1 on failure. The only possible errno error condition is EFAULT.

Function int setrlimit (int resource, struct rlimit *rlp)

Store the current value and the maximum value of resource resource in *rlp. The return value is 0 on success and -1 on failure. The following errno error condition is possible:

EPERM You tried to change the maximum permissible limit value, but you don’t have privileges to do so.

Data Type struct rlimit

This structure is used with getrlimit to receive limit values, and with setrlimit to specify limit values. It has two fields:

rlim_cur The current value of the limit in question. This is also called the “soft limit”.

rlim_max The maximum permissible value of the limit in question. You cannot set the current value of the limit to a larger number than this maximum. Only the super user can change the maximum permissible value. This is also called the “hard limit”.

In getrlimit, the structure is an output; it receives the current values. In setrlimit, it specifies the new values.

Here is a list of resources that you can specify a limit for. Those that are sizes are measured in bytes.

RLIMIT_CPU

The maximum amount of cpu time the process can use. If it runs for longer than this, it gets a signal: SIGXCPU. The value is measured in seconds. See Section 21.2.6 [Operation Error Signals], page 403.

RLIMIT_FSIZE

The maximum size of file the process can create. Trying to write a larger file causes a signal: SIGXFSZ. See Section 21.2.6 [Operation Error Signals], page 403.

RLIMIT_DATA

The maximum size of data memory for the process. If the process tries to allocate data memory beyond this amount, the allocation function fails.

RLIMIT_STACK

The maximum stack size for the process. If the process tries to extend its stack past this size, it gets a SIGSEGV signal. See Section 21.2.1 [Program Error Signals], page 396.

RLIMIT_CORE

The maximum size core file that this process can create. If the process terminates and would dump a core file larger than this maximum size, then no core file is created. So setting this limit to zero prevents core files from ever being created.

RLIMIT_RSS

The maximum amount of physical memory that this process should get. This parameter is a guide for the system’s scheduler and memory allocator; the system may give the process more memory when there is a surplus.

RLIMIT_MEMLOCK

The maximum amount of memory that can be locked into physical memory (so it will never be paged out).

RLIMIT_NPROC

The maximum number of processes that can be created with the same user ID. If you have reached the limit for your user ID, fork will fail with EAGAIN. See Section 23.4 [Creating a Process], page 469.

RLIMIT_NOFILE
RLIMIT_OFILE

The maximum number of files that the process can open. If it tries to open more files than this, it gets error code EMFILE. See Section 2.2 [Error Codes], page 16. Not all systems support this limit; GNU does, and 4.4 BSD does.

RLIM_NLIMITS

The number of different resource limits. Any valid resource operand must be less than RLIM_NLIMITS.

Constant int RLIM INFINITY This constant stands for a value of “infinity” when supplied as the limit value in setrlimit.

Two historical functions for setting resource limits, ulimit and vlimit, are not documented here. The latter is declared in ‘sys/vlimit.h’ and comes from BSD.

17.7 Process Priority

When several processes try to run, their respective priorities determine what share of the CPU each process gets. This section describes how you can read and set the priority of a process. All these functions and macros are declared in ‘sys/resource.h’.

The range of valid priority values depends on the operating system, but typically it runs from -20 to 20. A lower priority value means the process runs more often. These constants describe the range of priority values:

PRIO_MIN The smallest valid priority value.

PRIO_MAX The smallest valid priority value.

Function int getpriority (int class, int id)

Read the priority of a class of processes; class and id specify which ones (see below). If the processes specified do not all have the same priority, this returns the smallest value that any of them has. The return value is the priority value on success, and -1 on failure. The following errno error condition are possible for this function:

ESRCH The combination of class and id does not match any existing process.

EINVAL The value of class is not valid. When the return value is -1, it could indicate failure, or it could be the priority value. The only way to make certain is to set errno = 0 before calling getpriority, then use errno != 0 afterward as the criterion for failure.

Function int setpriority (int class, int id, int priority)

Set the priority of a class of processes to priority; class and id specify which ones (see below). The return value is 0 on success and -1 on failure. The following errno error condition are defined for this function:

ESRCH The combination of class and id does not match any existing process.

EINVAL The value of class is not valid.

EPERM You tried to set the priority of some other user’s process, and you don’t have privileges for that.

EACCES You tried to lower the priority of a process, and you don’t have privileges for that.

The arguments class and id together specify a set of processes you are interested in. These are the possible values for class:

PRIO_PROCESS

Read or set the priority of one process. The argument id is a process ID.

PRIO_PGRP

Read or set the priority of one process group. The argument id is a process group ID.

PRIO_USER

Read or set the priority of one user’s processes. The argument id is a user ID.

If the argument id is 0, it stands for the current process, current process group, or the current user, according to class.

Function int nice (int increment)

Increment the priority of the current process by increment. The return value is the same as for setpriority. Here is an equivalent definition for nice:

        int
        nice (int increment)
        {
          int old = getpriority (PRIO_PROCESS, 0);
          return setpriority (PRIO_PROCESS, 0, old + increment);
        }

366 The GNU C Library

18 Extended Characters

A number of languages use character sets that are larger than the range of values of type char. Japanese and Chinese are probably the most familiar examples.

The GNU C library includes support for two mechanisms for dealing with extended character sets: multibyte characters and wide characters. This chapter describes how to use these mechanisms, and the functions for converting between them.

The behavior of the functions in this chapter is affected by the current locale for character classification—the LC_CTYPE category; see Section 19.3 [Categories of Activities that Locales Affect], page 380. This choice of locale selects which multibyte code is used, and also controls the meanings and characteristics of wide character codes.

18.1 Introduction to Extended Characters

You can represent extended characters in either of two ways:

• As multibyte characters which can be embedded in an ordinary string, an array of char objects. Their advantage is that many programs and operating systems can handle occasional multibyte characters scattered among ordinary ASCII characters, without any change.

• As wide characters, which are like ordinary characters except that they occupy more bits. The wide character data type, wchar_t, has a range large enough to hold extended character codes as well as old-fashioned ASCII codes. An advantage of wide characters is that each character is a single data object, just like ordinary ASCII characters. There are a few disadvantages:

• Each existing program must be modified and recompiled to make it use wide characters.

• Files of wide characters cannot be read by programs that expect ordinary characters.

Typically, you use the multibyte character representation as part of the external program interface, such as reading or writing text to files. However, it’s usually easier to perform internal manipulations on strings containing extended characters on arrays of wchar_t objects, since the uniform representation makes most editing operations easier. If you do use multibyte characters for files and wide characters for internal operations, you need to convert between them when you read and write data.

If your system supports extended characters, then it supports them both as multibyte characters and as wide characters. The library includes functions you can use to convert between the two representations. These functions are described in this chapter.

18.2 Locales and Extended Characters

A computer system can support more than one multibyte character code, and more than one wide character code. The user controls the choice of codes through the current locale for character classification (see Chapter 19 [Locales and Internationalization], page 379). Each locale specifies a particular multibyte character code and a particular wide character code. The choice of locale influences the behavior of the conversion functions in the library.

Some locales support neither wide characters nor nontrivial multibyte characters. In these locales, the library conversion functions still work, even though what they do is basically trivial.

If you select a new locale for character classification, the internal shift state maintained by these functions can become confused, so it’s not a good idea to change the locale while you are in the middle of processing a string.

18.3 Multibyte Characters

In the ordinary ASCII code, a sequence of characters is a sequence of bytes, and each character is one byte. This is very simple, but allows for only 256 distinct characters.

In a multibyte character code, a sequence of characters is a sequence of bytes, but each character may occupy one or more consecutive bytes of the sequence.

There are many different ways of designing a multibyte character code; different systems use different codes. To specify a particular code means designating the basic byte sequences—those which represent a single character—and what characters they stand for. A code that a computer can actually use must have a finite number of these basic sequences, and typically none of them is more than a few characters long.

These sequences need not all have the same length. In fact, many of them are just one byte long. Because the basic ASCII characters in the range from 0 to 0177 are so important, they stand for themselves in all multibyte character codes. That is to say, a byte whose value is 0 through 0177 is always a character in itself. The characters which are more than one byte must always start with a byte in the range from 0200 through 0377.

The byte value 0 can be used to terminate a string, just as it is often used in a string of ASCII characters.

Specifying the basic byte sequences that represent single characters automatically gives meanings to many longer byte sequences, as more than one character. For example, if the two byte sequence 0205 049 stands for the Greek letter alpha, then 0205 049 065 must stand for an alpha followed by an ‘A’ (ASCII code 065), and 0205 049 0205 049 must stand for two alphas in a row.

If any byte sequence can have more than one meaning as a sequence of characters, then the multibyte code is ambiguous—and no good. The codes that systems actually use are all unambiguous.

In most codes, there are certain sequences of bytes that have no meaning as a character or characters. These are called invalid.

The simplest possible multibyte code is a trivial one:

The basic sequences consist of single bytes.

This particular code is equivalent to not using multibyte characters at all. It has no invalid sequences. But it can handle only 256 different characters.

Here is another possible code which can handle 9376 different characters: The basic sequences consist of

• single bytes with values in the range 0 through 0237.

• two-byte sequences, in which both of the bytes have values in the range from 0240 through 0377.

This code or a similar one is used on some systems to represent Japanese characters. The invalid sequences are those which consist of an odd number of consecutive bytes in the range from 0240 through 0377.

Here is another multibyte code which can handle more distinct extended characters—in fact, almost thirty million:

The basic sequences consist of • single bytes with values in the range 0 through 0177.

• sequences of up to four bytes in which the first byte is in the range from 0200 through 0237, and the remaining bytes are in the range from 0240 through 0377.

In this code, any sequence that starts with a byte in the range from 0240 through 0377 is invalid.

And here is another variant which has the advantage that removing the last byte or bytes from a valid character can never produce another valid character. (This property is convenient when you want to search strings for particular characters.)

The basic sequences consist of • single bytes with values in the range 0 through 0177.

• two-byte sequences in which the first byte is in the range from 0200 through 0207, and the second byte is in the range from

        0240 through 0377.

• three-byte sequences in which the first byte is in the range from 0210 through 0217, and the other bytes are in the range from

        0240 through 0377.

• four-byte sequences in which the first byte is in the range from 0220 through 0227, and the other bytes are in the range from

        0240 through 0377.

The list of invalid sequences for this code is long and not worth stating in full; examples of invalid sequences include 0240 and 0220 0300 065.

The number of possible multibyte codes is astronomical. But a given computer system will support at most a few different codes. (One of these codes may allow for thousands of different characters.) Another computer system may support a completely different code. The library facilities described in this chapter are helpful because they package up the knowledge of the details of a particular computer system’s multibyte code, so your programs need not know them.

You can use special standard macros to find out the maximum possible number of bytes in a character in the currently selected multibyte code with MB_CUR_MAX, and the maximum for any multibyte code supported on your computer with MB_LEN_MAX.

Macro int MB LEN MAX

This is the maximum length of a multibyte character for any supported locale. It is defined in ‘limits.h’.

Macro int MB CUR MAX

This macro expands into a (possibly non-constant) positive integer expression that is the maximum number of bytes in a multibyte character in the current locale. The value is never greater than MB_LEN_MAX.

  MB_CUR_MAX is defined in ‘stdlib.h’.

Normally, each basic sequence in a particular character code stands for one character, the same character regardless of context. Some multibyte character codes have a concept of shift state; certain codes, called shift sequences, change to a different shift state, and the meaning of some or all basic sequences varies according to the current shift state. In fact, the set of basic sequences might even be different depending on the current shift state. See Section 18.9 [Multibyte Codes Using Shift Sequences], page 377, for more information on handling this sort of code.

What happens if you try to pass a string containing multibyte characters to a function that doesn’t know about them? Normally, such a function treats a string as a sequence of bytes, and interprets certain byte values specially; all other byte values are “ordinary”. As long as a multibyte character doesn’t contain any of the special byte values, the function should pass it through as if it were several ordinary characters.

For example, let’s figure out what happens if you use multibyte characters in a file name. The functions such as open and unlink that operate on file names treat the name as a sequence of byte values, with ‘/’ as the only special value. Any other byte values are copied, or compared, in sequence, and all byte values are treated alike. Thus, you may think of the file name as a sequence of bytes or as a string containing multibyte characters; the same behavior makes sense equally either way, provided no multibyte character contains a ‘/’.

18.4 Wide Character Introduction

Wide characters are much simpler than multibyte characters. They are simply characters with more than eight bits, so that they have room for more than 256 distinct codes. The wide character data type, wchar_t, has a range large enough to hold extended character codes as well as old-fashioned ASCII codes.

An advantage of wide characters is that each character is a single data object, just like ordinary ASCII characters. Wide characters also have some disadvantages:

• A program must be modified and recompiled in order to use wide characters at all.

• Files of wide characters cannot be read by programs that expect ordinary characters.

Wide character values 0 through 0177 are always identical in meaning to the ASCII character codes. The wide character value zero is often used to terminate a string of wide characters, just as a single byte with value zero often terminates a string of ordinary characters.

Data Type wchar t

This is the “wide character” type, an integer type whose range is large enough to represent all distinct values in any extended character set in the supported locales. See Chapter 19 [Locales and Internationalization], page 379, for more information about locales. This type is defined in the header file ‘stddef.h’.

If your system supports extended characters, then each extended character has both a wide character code and a corresponding multibyte basic sequence.

In this chapter, the term code is used to refer to a single extended character object to emphasize the distinction from the char data type.

18.5 Conversion of Extended Strings

The mbstowcs function converts a string of multibyte characters to a wide character array. The wcstombs function does the reverse. These functions are declared in the header file ‘stdlib.h’.

In most programs, these functions are the only ones you need for conversion between wide strings and multibyte character strings. But they have limitations. If your data is not null-terminated or is not all in core at once, you probably need to use the low-level conversion functions to convert one character at a time. See Section 18.7 [Conversion of Extended Characters One by One], page 374.

Function size_t mbstowcs (wchar_t *wstring, const char

*string, size_t size) The mbstowcs (“multibyte string to wide character string”) function converts the null-terminated string of multibyte characters string to an array of wide character codes, storing not more than size wide characters into the array beginning at wstring. The terminating null character counts towards the size, so if size is less than the actual number of wide characters resulting from string, no terminating null character is stored. The conversion of characters from string begins in the initial shift state. If an invalid multibyte character sequence is found, this function returns a value of -1. Otherwise, it returns the number of wide characters stored in the array wstring. This number does not include the terminating null character, which is present if the number is less than size. Here is an example showing how to convert a string of multibyte characters, allocating enough space for the result.

     wchar_t *
     mbstowcs_alloc (const char *string)
     {
       size_t size = strlen (string) + 1;
       wchar_t *buf = xmalloc (size * sizeof (wchar_t));
       size = mbstowcs (buf, string, size);
       if (size == (size_t) -1)
         return NULL;
       buf = xrealloc (buf, (size + 1) * sizeof (wchar_t));
       return buf;
     }

Function size_t wcstombs (char *string, const wchar_t

wstring, size_t size) The wcstombs (“wide character string to multibyte string”) function converts the null-terminated wide character array wstring into a string containing multibyte characters, storing not more than size bytes starting at string, followed by a terminating null character if there is room. The conversion of characters begins in the initial shift state. The terminating null character counts towards the size, so if size is less than or equal to the number of bytes needed in wstring, no terminating null character is stored. If a code that does not correspond to a valid multibyte character is found, this function returns a value of -1. Otherwise, the return value is the number of bytes stored in the array string. This number does not include the terminating null character, which is present if the number is less than size.

18.6 Multibyte Character Length

This section describes how to scan a string containing multibyte characters, one character at a time. The difficulty in doing this is to know how many bytes each character contains. Your program can use mblen to find this out.

Function int mblen (const char *string, size_t size)

The mblen function with a non-null string argument returns the number of bytes that make up the multibyte character beginning at string, never examining more than size bytes. (The idea is to supply for size the number of bytes of data you have in hand.) The return value of mblen distinguishes three possibilities: the first size bytes at string start with valid multibyte character, they start with an invalid byte sequence or just part of a character, or string points to an empty string (a null character). For a valid multibyte character, mblen returns the number of bytes in that character (always at least 1, and never more than size). For an invalid byte sequence, mblen returns -1. For an empty string, it returns 0. If the multibyte character code uses shift characters, then mblen maintains and updates a shift state as it scans. If you call mblen with a null pointer for string, that initializes the shift state to its standard initial value. It also returns nonzero if the multibyte character code in use actually has a shift state. See Section 18.9 [Multibyte Codes Using Shift Sequences], page 377. The function mblen is declared in ‘stdlib.h’.

18.7 Conversion of Extended Characters One by One

You can convert multibyte characters one at a time to wide characters with the mbtowc function. The wctomb function does the reverse. These functions are declared in ‘stdlib.h’.

Function int mbtowc (wchar_t *result, const char *string,

      size_t size)

The mbtowc (“multibyte to wide character”) function when called with non-null string converts the first multibyte character beginning at string to its corresponding wide character code. It stores the result in *result. mbtowc never examines more than size bytes. (The idea is to supply for size the number of bytes of data you have in hand.) mbtowc with non-null string distinguishes three possibilities: the first size bytes at string start with valid multibyte character, they start with an invalid byte sequence or just part of a character, or string points to an empty string (a null character). For a valid multibyte character, mbtowc converts it to a wide character and stores that in *result, and returns the number of bytes in that character (always at least 1, and never more than size). For an invalid byte sequence, mbtowc returns -1. For an empty string, it returns 0, also storing 0 in *result. If the multibyte character code uses shift characters, then mbtowc maintains and updates a shift state as it scans. If you call mbtowc with a null pointer for string, that initializes the shift state to its standard initial value. It also returns nonzero if the multibyte character code in use actually has a shift state. See Section 18.9 [Multibyte Codes Using Shift Sequences], page 377.

Function int wctomb (char *string, wchar_t wchar)

The wctomb (“wide character to multibyte”) function converts the wide character code wchar to its corresponding multibyte character sequence, and stores the result in bytes starting at string. At most MB_CUR_MAX characters are stored. wctomb with non-null string distinguishes three possibilities for wchar: a valid wide character code (one that can be translated to a multibyte character), an invalid code, and 0. Given a valid code, wctomb converts it to a multibyte character, storing the bytes starting at string. Then it returns the number of bytes in that character (always at least 1, and never more than MB_CUR_MAX). If wchar is an invalid wide character code, wctomb returns -1. If wchar is 0, it returns 0, also storing 0 in *string.

If the multibyte character code uses shift characters, then wctomb maintains and updates a shift state as it scans. If you call wctomb with a null pointer for string, that initializes the shift state to its standard initial value. It also returns nonzero if the multibyte character code in use actually has a shift state. See Section 18.9 [Multibyte Codes Using Shift Sequences], page 377. Calling this function with a wchar argument of zero when string is not null has the side-effect of reinitializing the stored shift state as well as storing the multibyte character 0 and returning 0.

18.8 Character-by-Character Conversion Example

Here is an example that reads multibyte character text from descriptor input and writes the corresponding wide characters to descriptor output. We need to convert characters one by one for this example because mbstowcs is unable to continue past a null character, and cannot cope with an apparently invalid partial character by reading more input.

  int
  file_mbstowcs (int input, int output)
  {
    char buffer[BUFSIZ + MB_LEN_MAX];
    int filled = 0;
    int eof = 0;
    while (!eof)
      {
        int nread;
        int nwrite;
        char *inp = buffer;
        wchar_t outbuf[BUFSIZ];
        wchar_t *outp = outbuf;

/* Fill up the buffer from the input file. */

        nread = read (input, buffer + filled, BUFSIZ);
        if (nread < 0)
          {
            perror ("read");
            return 0;
          }

/* If we reach end of file, make a note to read no more. */

        if (nread == 0)
          eof = 1;

/* filled is now the number of bytes in buffer. */

filled += nread;

/* Convert those bytes to wide characters–as many as we can. */

while (1)
{
int thislen = mbtowc (outp, inp, filled);
/* Stop converting at invalid character;
   this can mean we have read just the first part
   of a valid character.  */
if (thislen == -1)
  break;

/* Treat null character like any other,

but also reset shift state. */

if (thislen == 0) {
  thislen = 1;
  mbtowc (NULL, NULL, 0);
}

/* Advance past this character. */

inp += thislen;
filled -= thislen;
outp++;
}

/* Write the wide characters we just made. */

nwrite = write (output, outbuf,
            (outp - outbuf) * sizeof (wchar_t));
if (nwrite < 0)
{
perror ("write");
return 0;
}

/* See if we have a real invalid character. */

if ((eof && filled > 0) || filled >= MB_CUR_MAX)
{
error ("invalid multibyte character");
return 0;
}

/* If any characters must be carried forward,

put them at the beginning of buffer. */

if (filled > 0)
memcpy (inp, buffer, filled);
}

         }
       return 1;
     }

18.9 Multibyte Codes Using Shift Sequences

In some multibyte character codes, the meaning of any particular byte sequence is not fixed; it depends on what other sequences have come earlier in the same string. Typically there are just a few sequences that can change the meaning of other sequences; these few are called shift sequences and we say that they set the shift state for other sequences that follow.

To illustrate shift state and shift sequences, suppose we decide that the sequence 0200 (just one byte) enters Japanese mode, in which pairs of bytes in the range from 0240 to 0377 are single characters, while 0201 enters Latin-1 mode, in which single bytes in the range from 0240 to 0377 are characters, and interpreted according to the ISO Latin-1 character set. This is a multibyte code which has two alternative shift states (“Japanese mode” and “Latin-1 mode”), and two shift sequences that specify particular shift states.

When the multibyte character code in use has shift states, then mblen, mbtowc and wctomb must maintain and update the current shift state as they scan the string. To make this work properly, you must follow these rules:

• Before starting to scan a string, call the function with a null pointer for the multibyte character address—for example, mblen (NULL, 0). This initializes the shift state to its standard initial value.

• Scan the string one character at a time, in order. Do not “back up” and rescan characters already scanned, and do not intersperse the processing of different strings.

Here is an example of using mblen following these rules:

     void
     scan_string (char *s)
     {
       int length = strlen (s);

/* Initialize shift state. */

       mblen (NULL, 0);
       while (1)
         {
           int thischar = mblen (s, length);

/* Deal with end of string and invalid characters. */

           if (thischar == 0)

             break;
           if (thischar == -1)
             {
               error ("invalid multibyte character");
               break;
             }

/* Advance past this character. */

           s += thischar;
           length -= thischar;
         }
     }

The functions mblen, mbtowc and wctomb are not reentrant when using a multibyte code that uses a shift state. However, no other library functions call these functions, so you don’t have to worry that the shift state will be changed mysteriously.

19 Locales and Internationalization

Different countries and cultures have varying conventions for how to communicate. These conventions range from very simple ones, such as the format for representing dates and times, to very complex ones, such as the language spoken.

Internationalization of software means programming it to be able to adapt to the user’s favorite conventions. In ISO C, internationalization works by means of locales. Each locale specifies a collection of conventions, one convention for each purpose. The user chooses a set of conventions by specifying a locale (via environment variables).

All programs inherit the chosen locale as part of their environment. Provided the programs are written to obey the choice of locale, they will follow the conventions preferred by the user.

19.1 What Effects a Locale Has

Each locale specifies conventions for several purposes, including the following:

• What multibyte character sequences are valid, and how they are interpreted (see Chapter 18 [Extended Characters], page 367).

• Classification of which characters in the local character set are considered alphabetic, and upper- and lower-case conversion conventions (see Chapter 4 [Character Handling], page 61).

• The collating sequence for the local language and character set (see Section 5.6 [Collation Functions], page 75).

• Formatting of numbers and currency amounts (see Section 19.6 [Numeric Formatting], page 383).

• Formatting of dates and times (see Section 17.2.4 [Formatting Date and Time], page 347).

• What language to use for output, including error messages. (The C library doesn’t yet help you implement this.)

• What language to use for user answers to yes-or-no questions.

• What language to use for more complex user input. (The C library doesn’t yet help you implement this.)

Some aspects of adapting to the specified locale are handled automatically by the library subroutines. For example, all your program needs to do in order to use the collating sequence of the chosen locale is to use strcoll or strxfrm to compare strings.

Other aspects of locales are beyond the comprehension of the library. For example, the library can’t automatically translate your program’s output messages into other languages. The only way you can support output in the user’s favorite language is to program this more or less by hand. (Eventually, we hope to provide facilities to make this easier.)

This chapter discusses the mechanism by which you can modify the current locale. The effects of the current locale on specific library functions are discussed in more detail in the descriptions of those functions.

19.2 Choosing a Locale

The simplest way for the user to choose a locale is to set the environment variable LANG. This specifies a single locale to use for all purposes. For example, a user could specify a hypothetical locale named ‘espana-castellano’ to use the standard conventions of most of Spain.

The set of locales supported depends on the operating system you are using, and so do their names. We can’t make any promises about what locales will exist, except for one standard locale called ‘C’ or ‘POSIX’.

A user also has the option of specifying different locales for different purposes—in effect, choosing a mixture of multiple locales.

For example, the user might specify the locale ‘espana-castellano’ for most purposes, but specify the locale ‘usa-english’ for currency formatting. This might make sense if the user is a Spanish-speaking American, working in Spanish, but representing monetary amounts in US dollars.

Note that both locales ‘espana-castellano’ and ‘usa-english’, like all locales, would include conventions for all of the purposes to which locales apply. However, the user can choose to use each locale for a particular subset of those purposes.

19.3 Categories of Activities that Locales Affect

The purposes that locales serve are grouped into categories, so that a user or a program can choose the locale for each category independently. Here is a table of categories; each name is both an environment variable that a user can set, and a macro name that you can use as an argument to setlocale.

LC_COLLATE

This category applies to collation of strings (functions strcoll and strxfrm); see Section 5.6 [Collation Functions], page 75.

LC_CTYPE This category applies to classification and conversion of characters, and to multibyte and wide characters; see Chapter 4 [Character Handling], page 61 and Chapter 18 [Extended Characters], page 367.

LC_MONETARY

This category applies to formatting monetary values; see Section 19.6 [Numeric Formatting], page 383.

LC_NUMERIC

This category applies to formatting numeric values that are not monetary; see Section 19.6 [Numeric Formatting], page 383.

LC_TIME This category applies to formatting date and time values; see Section 17.2.4 [Formatting Date and Time], page 347.

LC_MESSAGES

This category applies to selecting the language used in the user interface for message translation.

LC_ALL This is not an environment variable; it is only a macro that you can use with setlocale to set a single locale for all purposes.

LANG If this environment variable is defined, its value specifies the locale to use for all purposes except as overridden by the variables above.

19.4 How Programs Set the Locale

A C program inherits its locale environment variables when it starts up. This happens automatically. However, these variables do not automatically control the locale used by the library functions, because ISO C says that all programs start by default in the standard ‘C’ locale. To use the locales specified by the environment, you must call setlocale. Call it as follows:

     setlocale (LC_ALL, "");

to select a locale based on the appropriate environment variables.

You can also use setlocale to specify a particular locale, for general use or for a specific category.

The symbols in this section are defined in the header file ‘locale.h’.

Function char * setlocale (int category, const char *locale)

The function setlocale sets the current locale for category category to locale. If category is LC_ALL, this specifies the locale for all purposes. The other possible values of category specify an individual purpose (see Section 19.3 [Categories of Activities that Locales Affect], page 380). You can also use this function to find out the current locale by passing a null pointer as the locale argument. In this case, setlocale returns a string that is the name of the locale currently selected for category category.

The string returned by setlocale can be overwritten by subsequent calls, so you should make a copy of the string (see Section 5.4 [Copying and Concatenation], page 67) if you want to save it past any further calls to setlocale. (The standard library is guaranteed never to call

setlocale itself.)

You should not modify the string returned by setlocale. It might be the same string that was passed as an argument in a previous call to

setlocale.

When you read the current locale for category LC_ALL, the value encodes the entire combination of selected locales for all categories. In this case, the value is not just a single locale name. In fact, we don’t make any promises about what it looks like. But if you specify the same “locale name” with LC_ALL in a subsequent call to setlocale, it restores the same combination of locale selections. When the locale argument is not a null pointer, the string returned by setlocale reflects the newly modified locale. If you specify an empty string for locale, this means to read the appropriate environment variable and use its value to select the locale for category. If you specify an invalid locale name, setlocale returns a null pointer and leaves the current locale unchanged.

Here is an example showing how you might use setlocale to temporarily switch to a new locale.

  #include <stddef.h>
  #include <locale.h>
  #include <stdlib.h>
  #include <string.h>
  void
  with_other_locale (char *new_locale,
                     void (*subroutine) (int),
                     int argument)
  {
    char *old_locale, *saved_locale;

/* Get the name of the current locale. */

    old_locale = setlocale (LC_ALL, NULL);

/* Copy the name so it won’t be clobbered by setlocale. */

    saved_locale = strdup (old_locale);
    if (old_locale == NULL)
      fatal ("Out of memory");

/* Now change the locale and do some stuff with it. */

       setlocale (LC_ALL, new_locale);
       (*subroutine) (argument);

/* Restore the original locale. */

       setlocale (LC_ALL, saved_locale);
       free (saved_locale);
     }

Portability Note: Some ISO C systems may define additional locale categories. For portability, assume that any symbol beginning with ‘LC_’ might be defined in ‘locale.h’.

19.5 Standard Locales

The only locale names you can count on finding on all operating systems are these three standard ones:

"C" This is the standard C locale. The attributes and behavior it provides are specified in the ISO C standard. When your program starts up, it initially uses this locale by default.

"POSIX" This is the standard POSIX locale. Currently, it is an alias for the standard C locale.

"" The empty name says to select a locale based on environment variables. See Section 19.3 [Categories of Activities that Locales Affect], page 380.

Defining and installing named locales is normally a responsibility of the system administrator at your site (or the person who installed the GNU C library). Some systems may allow users to create locales, but we don’t discuss that here.

If your program needs to use something other than the ‘C’ locale, it will be more portable if you use whatever locale the user specifies with the environment, rather than trying to specify some non-standard locale explicitly by name. Remember, different machines might have different sets of locales installed.

19.6 Numeric Formatting

When you want to format a number or a currency amount using the conventions of the current locale, you can use the function localeconv to get the data on how to do it. The function localeconv is declared in the header file ‘locale.h’.

Function struct lconv * localeconv (void)

The localeconv function returns a pointer to a structure whose components contain information about how numeric and monetary values should be formatted in the current locale. You shouldn’t modify the structure or its contents. The structure might be overwritten by subsequent calls to localeconv, or by calls to setlocale, but no other function in the library overwrites this value.

Data Type struct lconv

This is the data type of the value returned by localeconv.

If a member of the structure struct lconv has type char, and the value is CHAR_MAX, it means that the current locale has no value for that parameter.

19.6.1 Generic Numeric Formatting Parameters

These are the standard members of struct lconv; there may be others.

char *decimal_point
char *mon_decimal_point

These are the decimal-point separators used in formatting nonmonetary and monetary quantities, respectively. In the ‘C’ locale, the value of decimal_point is ".", and the value of mon_

decimal_point is "".
char *thousands_sep
char *mon_thousands_sep

These are the separators used to delimit groups of digits to the left of the decimal point in formatting non-monetary and monetary quantities, respectively. In the ‘C’ locale, both members have a value of "" (the empty string).

char *grouping
char *mon_grouping

These are strings that specify how to group the digits to the left of the decimal point. grouping applies to non-monetary quantities and mon_grouping applies to monetary quantities.

Use either thousands_sep or mon_thousands_sep to separate

the digit groups. Each string is made up of decimal numbers separated by semicolons. Successive numbers (from left to right) give the sizes of successive groups (from right to left, starting at the decimal point). The last number in the string is used over and over for all the remaining groups.

If the last integer is -1, it means that there is no more grouping— or, put another way, any remaining digits form one large group without separators. For example, if grouping is "4;3;2", the correct grouping for

the number 123456787654321 is ‘12’, ‘34’, ‘56’, ‘78’, ‘765’,

‘4321’. This uses a group of 4 digits at the end, preceded by a group of 3 digits, preceded by groups of 2 digits (as many as needed). With a separator of ‘,’, the number would be printed

as ‘12,34,56,78,765,4321’.

A value of "3" indicates repeated groups of three digits, as normally used in the U.S. In the standard ‘C’ locale, both grouping and mon_grouping have a value of "". This value specifies no grouping at all.

char int_frac_digits
char frac_digits

These are small integers indicating how many fractional digits (to the right of the decimal point) should be displayed in a monetary value in international and local formats, respectively. (Most often, both members have the same value.) In the standard ‘C’ locale, both of these members have the value CHAR_MAX, meaning “unspecified”. The ISO standard doesn’t say what to do when you find this the value; we recommend printing no fractional digits. (This locale also specifies the empty string for mon_decimal_point, so printing any fractional digits would be confusing!)

19.6.2 Printing the Currency Symbol

These members of the struct lconv structure specify how to print the symbol to identify a monetary value—the international analog of ‘$’ for US dollars.

Each country has two standard currency symbols. The local currency symbol is used commonly within the country, while the international currency symbol is used internationally to refer to that country’s currency when it is necessary to indicate the country unambiguously.

For example, many countries use the dollar as their monetary unit, and when dealing with international currencies it’s important to specify that one is dealing with (say) Canadian dollars instead of U.S. dollars or Australian dollars. But when the context is known to be Canada, there is no need to make this explicit—dollar amounts are implicitly assumed to be in Canadian dollars.

char *currency_symbol

The local currency symbol for the selected locale.

In the standard ‘C’ locale, this member has a value of "" (the empty string), meaning “unspecified”. The ISO standard doesn’t say what to do when you find this value; we recommend you simply print the empty string as you would print any other string found in the appropriate member.

char *int_curr_symbol

The international currency symbol for the selected locale. The value of int_curr_symbol should normally consist of a three-letter abbreviation determined by the international standard ISO 4217 Codes for the Representation of Currency and Funds, followed by a one-character separator (often a space). In the standard ‘C’ locale, this member has a value of "" (the empty string), meaning “unspecified”. We recommend you simply print the empty string as you would print any other string found in the appropriate member.

char p_cs_precedes
char n_cs_precedes

These members are 1 if the currency_symbol string should precede the value of a monetary amount, or 0 if the string should follow the value. The p_cs_precedes member applies to positive amounts (or zero), and the n_cs_precedes member applies to negative amounts. In the standard ‘C’ locale, both of these members have a value of CHAR_MAX, meaning “unspecified”. The ISO standard doesn’t say what to do when you find this value, but we recommend printing the currency symbol before the amount. That’s right for most countries. In other words, treat all nonzero values alike in these members. The POSIX standard says that these two members apply to

the int_curr_symbol as well as the currency_symbol. The

ISO C standard seems to imply that they should apply only to

the currency_symbol—so the int_curr_symbol should always

precede the amount. We can only guess which of these (if either) matches the usual conventions for printing international currency symbols. Our guess is that they should always precede the amount. If we find out a reliable answer, we will put it here.

char p_sep_by_space
char n_sep_by_space

These members are 1 if a space should appear between the currency_symbol string and the amount, or 0 if no space should appear. The p_sep_by_space member applies to positive amounts (or zero), and the n_sep_by_space member applies to negative amounts. In the standard ‘C’ locale, both of these members have a value of CHAR_MAX, meaning “unspecified”. The ISO standard doesn’t say what you should do when you find this value; we suggest you treat it as one (print a space). In other words, treat all nonzero values alike in these members. These members apply only to currency_symbol. When you use int_curr_symbol, you never print an additional space, because int_curr_symbol itself contains the appropriate separator. The POSIX standard says that these two members apply to the

int_curr_symbol as well as the currency_symbol. But an ex-

ample in the ISO C standard clearly implies that they should ap-

ply only to the currency_symbol—that the int_curr_symbol

contains any appropriate separator, so you should never print an additional space. Based on what we know now, we recommend you ignore these members when printing international currency symbols, and print no extra space.

19.6.3 Printing the Sign of an Amount of Money

These members of the struct lconv structure specify how to print the sign (if any) in a monetary value.

char *positive_sign
char *negative_sign

These are strings used to indicate positive (or zero) and negative (respectively) monetary quantities. In the standard ‘C’ locale, both of these members have a value of "" (the empty string), meaning “unspecified”. The ISO standard doesn’t say what to do when you find this value; we recommend printing positive_sign as you find it, even if it is empty. For a negative value, print negative_sign as you find it unless both it and positive_sign are empty, in which case print ‘-’ instead. (Failing to indicate the sign at all seems rather unreasonable.)

char p_sign_posn
char n_sign_posn

These members have values that are small integers indicating how to position the sign for nonnegative and negative monetary quantities, respectively. (The string used by the sign is what was specified with positive_sign or negative_sign.) The possible values are as follows:

0 The currency symbol and quantity should be surrounded by parentheses.

1 Print the sign string before the quantity and currency symbol.

2 Print the sign string after the quantity and currency symbol.

3 Print the sign string right before the currency symbol.

4 Print the sign string right after the currency symbol.

CHAR_MAX “Unspecified”. Both members have this value in the standard ‘C’ locale.

The ISO standard doesn’t say what you should do when the value is CHAR_MAX. We recommend you print the sign after the currency symbol.

It is not clear whether you should let these members apply to the international currency format or not. POSIX says you should, but intuition plus the examples in the ISO C standard suggest you should not. We hope that someone who knows well the conventions for formatting monetary quantities will tell us what we should recommend.

20 Non-Local Exits

Sometimes when your program detects an unusual situation inside a deeply nested set of function calls, you would like to be able to immediately return to an outer level of control. This section describes how you can do such non-local exits using the setjmp and longjmp functions.

20.1 Introduction to Non-Local Exits

As an example of a situation where a non-local exit can be useful, suppose you have an interactive program that has a “main loop” that prompts for and executes commands. Suppose the “read” command reads input from a file, doing some lexical analysis and parsing of the input while processing it. If a low-level input error is detected, it would be useful to be able to return immediately to the “main loop” instead of having to make each of the lexical analysis, parsing, and processing phases all have to explicitly deal with error situations initially detected by nested calls.

(On the other hand, if each of these phases has to do a substantial amount of cleanup when it exits—such as closing files, deallocating buffers or other data structures, and the like—then it can be more appropriate to do a normal return and have each phase do its own cleanup, because a non-local exit would bypass the intervening phases and their associated cleanup code entirely. Alternatively, you could use a non-local exit but do the cleanup explicitly either before or after returning to the “main loop”.)

In some ways, a non-local exit is similar to using the ‘return’ statement to return from a function. But while ‘return’ abandons only a single function call, transferring control back to the point at which it was called, a non-local exit can potentially abandon many levels of nested function calls.

You identify return points for non-local exits calling the function setjmp. This function saves information about the execution environment in which the call to setjmp appears in an object of type jmp_buf. Execution of the program continues normally after the call to setjmp, but if a exit is later made to this return point by calling longjmp with the corresponding jmp_buf object, control is transferred back to the point where setjmp was called. The return value from setjmp is used to distinguish between an ordinary return and a return made by a call to longjmp, so calls to setjmp usually appear in an ‘if’ statement.

Here is how the example program described above might be set up:

     #include <setjmp.h>
     #include <stdlib.h>
     #include <stdio.h>
     jmp_buf main_loop;

     void
     abort_to_main_loop (int status)
     {
       longjmp (main_loop, status);
     }
     int
     main (void)
     {
       while (1)
         if (setjmp (main_loop))
           puts ("Back at main loop....");
         else
           do_command ();
     }

     void
     do_command (void)
     {
       char buffer[128];
       if (fgets (buffer, 128, stdin) == NULL)
         abort_to_main_loop (-1);
       else
         exit (EXIT_SUCCESS);
     }

The function abort_to_main_loop causes an immediate transfer of control back to the main loop of the program, no matter where it is called from.

The flow of control inside the main function may appear a little mysterious at first, but it is actually a common idiom with setjmp. A normal call to setjmp returns zero, so the “else” clause of the conditional is executed. If abort_to_main_loop is called somewhere within the execution of do_ command, then it actually appears as if the same call to setjmp in main were returning a second time with a value of -1.

So, the general pattern for using setjmp looks something like:

     if (setjmp (buffer))

/* Code to clean up after premature return. */

…

     else

/* Code to be executed normally after setting up the return point. */ …

20.2 Details of Non-Local Exits

Here are the details on the functions and data structures used for performing non-local exits. These facilities are declared in ‘setjmp.h’.

Data Type jmp buf

Objects of type jmp_buf hold the state information to be restored by a non-local exit. The contents of a jmp_buf identify a specific place to return to.

Macro int setjmp (jmp_buf state)

When called normally, setjmp stores information about the execution state of the program in state and returns zero. If longjmp is later used to perform a non-local exit to this state, setjmp returns a nonzero value.

Function void longjmp (jmp_buf state, int value)

This function restores current execution to the state saved in state, and continues execution from the call to setjmp that established that return point. Returning from setjmp by means of longjmp returns the value argument that was passed to longjmp, rather than 0. (But if value is given as 0, setjmp returns 1).

There are a lot of obscure but important restrictions on the use of setjmp and longjmp. Most of these restrictions are present because non-local exits require a fair amount of magic on the part of the C compiler and can interact with other parts of the language in strange ways.

The setjmp function is actually a macro without an actual function definition, so you shouldn’t try to ‘#undef’ it or take its address. In addition, calls to setjmp are safe in only the following contexts:

• As the test expression of a selection or iteration statement (such as ‘if’,

    ‘switch’, or ‘while’).

• As one operand of a equality or comparison operator that appears as the test expression of a selection or iteration statement. The other operand must be an integer constant expression.

• As the operand of a unary ‘!’ operator, that appears as the test expression of a selection or iteration statement.

• By itself as an expression statement.

Return points are valid only during the dynamic extent of the function that called setjmp to establish them. If you longjmp to a return point that was established in a function that has already returned, unpredictable and disastrous things are likely to happen.

You should use a nonzero value argument to longjmp. While longjmp refuses to pass back a zero argument as the return value from setjmp, this is intended as a safety net against accidental misuse and is not really good programming style.

When you perform a non-local exit, accessible objects generally retain whatever values they had at the time longjmp was called. The exception is that the values of automatic variables local to the function containing the setjmp call that have been changed since the call to setjmp are indeterminate, unless you have declared them volatile.

20.3 Non-Local Exits and Signals

In BSD Unix systems, setjmp and longjmp also save and restore the set of blocked signals; see Section 21.7 [Blocking Signals], page 431. However, the POSIX.1 standard requires setjmp and longjmp not to change the set of blocked signals, and provides an additional pair of functions (sigsetjmp and sigsetjmp) to get the BSD behavior.

The behavior of setjmp and longjmp in the GNU library is controlled by feature test macros; see Section 1.3.4 [Feature Test Macros], page 8. The default in the GNU system is the POSIX.1 behavior rather than the BSD behavior.

The facilities in this section are declared in the header file ‘setjmp.h’.

Data Type sigjmp buf

This is similar to jmp_buf, except that it can also store state information about the set of blocked signals.

Function int sigsetjmp (sigjmp_buf state, int savesigs)

This is similar to setjmp. If savesigs is nonzero, the set of blocked signals is saved in state and will be restored if a siglongjmp is later performed with this state.

Function void siglongjmp (sigjmp_buf state, int value)

This is similar to longjmp except for the type of its state argument. If the sigsetjmp call that set this state used a nonzero savesigs flag, siglongjmp also restores the set of blocked signals.

21 Signal Handling

A signal is a software interrupt delivered to a process. The operating system uses signals to report exceptional situations to an executing program. Some signals report errors such as references to invalid memory addresses; others report asynchronous events, such as disconnection of a phone line.

The GNU C library defines a variety of signal types, each for a particular kind of event. Some kinds of events make it inadvisable or impossible for the program to proceed as usual, and the corresponding signals normally abort the program. Other kinds of signals that report harmless events are ignored by default.

If you anticipate an event that causes signals, you can define a handler function and tell the operating system to run it when that particular type of signal arrives.

Finally, one process can send a signal to another process; this allows a parent process to abort a child, or two related processes to communicate and synchronize.

21.1 Basic Concepts of Signals

This section explains basic concepts of how signals are generated, what happens after a signal is delivered, and how programs can handle signals.

21.1.1 Some Kinds of Signals

A signal reports the occurrence of an exceptional event. These are some of the events that can cause (or generate, or raise) a signal:

• A program error such as dividing by zero or issuing an address outside the valid range.

• A user request to interrupt or terminate the program. Most environments are set up to let a user suspend the program by typing C-z, or terminate it with C-c. Whatever key sequence is used, the operating system sends the proper signal to interrupt the process.

• The termination of a child process.

• Expiration of a timer or alarm.

• A call to kill or raise by the same process.

• A call to kill from another process. Signals are a limited but useful form of interprocess communication.

• An attempt to perform an I/O operation that cannot be done. Examples are reading from a pipe that has no writer (see Chapter 10 [Pipes and FIFOs], page 215), and reading or writing to a terminal in certain situations (see Chapter 24 [Job Control], page 479).

Each of these kinds of events (excepting explicit calls to kill and raise) generates its own particular kind of signal. The various kinds of signals are listed and described in detail in Section 21.2 [Standard Signals], page 395.

21.1.2 Concepts of Signal Generation

In general, the events that generate signals fall into three major categories: errors, external events, and explicit requests.

An error means that a program has done something invalid and cannot continue execution. But not all kinds of errors generate signals—in fact, most do not. For example, opening a nonexistent file is an error, but it does not raise a signal; instead, open returns -1. In general, errors that are necessarily associated with certain library functions are reported by returning a value that indicates an error. The errors which raise signals are those which can happen anywhere in the program, not just in library calls. These include division by zero and invalid memory addresses.

An external event generally has to do with I/O or other processes. These include the arrival of input, the expiration of a timer, and the termination of a child process.

An explicit request means the use of a library function such as kill whose purpose is specifically to generate a signal.

Signals may be generated synchronously or asynchronously. A synchronous signal pertains to a specific action in the program, and is delivered (unless blocked) during that action. Most errors generate signals synchronously, and so do explicit requests by a process to generate a signal for that same process. On some machines, certain kinds of hardware errors (usually floating-point exceptions) are not reported completely synchronously, but may arrive a few instructions later.

Asynchronous signals are generated by events outside the control of the process that receives them. These signals arrive at unpredictable times during execution. External events generate signals asynchronously, and so do explicit requests that apply to some other process.

A given type of signal is either typically synchronous or typically asynchronous. For example, signals for errors are typically synchronous because errors generate signals synchronously. But any type of signal can be generated synchronously or asynchronously with an explicit request.

21.1.3 How Signals Are Delivered

When a signal is generated, it becomes pending. Normally it remains pending for just a short period of time and then is delivered to the process that was signaled. However, if that kind of signal is currently blocked, it may remain pending indefinitely—until signals of that kind are unblocked.

Once unblocked, it will be delivered immediately. See Section 21.7 [Blocking Signals], page 431.

When the signal is delivered, whether right away or after a long delay, the specified action for that signal is taken. For certain signals, such as SIGKILL and SIGSTOP, the action is fixed, but for most signals, the program has a choice: ignore the signal, specify a handler function, or accept the default action for that kind of signal. The program specifies its choice using functions such as signal or sigaction (see Section 21.3 [Specifying Signal Actions], page 406). We sometimes say that a handler catches the signal. While the handler is running, that particular signal is normally blocked.

If the specified action for a kind of signal is to ignore it, then any such signal which is generated is discarded immediately. This happens even if the signal is also blocked at the time. A signal discarded in this way will never be delivered, not even if the program subsequently specifies a different action for that kind of signal and then unblocks it.

If a signal arrives which the program has neither handled nor ignored, its default action takes place. Each kind of signal has its own default action, documented below (see Section 21.2 [Standard Signals], page 395). For most kinds of signals, the default action is to terminate the process. For certain kinds of signals that represent “harmless” events, the default action is to do nothing.

When a signal terminates a process, its parent process can determine the cause of termination by examining the termination status code reported by the wait or waitpid functions. (This is discussed in more detail in Section 23.6 [Process Completion], page 473.) The information it can get includes the fact that termination was due to a signal, and the kind of signal involved. If a program you run from a shell is terminated by a signal, the shell typically prints some kind of error message.

The signals that normally represent program errors have a special property: when one of these signals terminates the process, it also writes a core dump file which records the state of the process at the time of termination. You can examine the core dump with a debugger to investigate what caused the error.

If you raise a “program error” signal by explicit request, and this terminates the process, it makes a core dump file just as if the signal had been due directly to an error.

21.2 Standard Signals

This section lists the names for various standard kinds of signals and describes what kind of event they mean. Each signal name is a macro which stands for a positive integer—the signal number for that kind of signal. Your programs should never make assumptions about the numeric code for a particular kind of signal, but rather refer to them always by the names defined here. This is because the number for a given kind of signal can vary from system to system, but the meanings of the names are standardized and fairly uniform.

The signal names are defined in the header file ‘signal.h’.

Macro int NSIG

The value of this symbolic constant is the total number of signals defined. Since the signal numbers are allocated consecutively, NSIG is also one greater than the largest defined signal number.

21.2.1 Program Error Signals

The following signals are generated when a serious program error is detected by the operating system or the computer itself. In general, all of these signals are indications that your program is seriously broken in some way, and there’s usually no way to continue the computation which encountered the error.

Some programs handle program error signals in order to tidy up before terminating; for example, programs that turn off echoing of terminal input should handle program error signals in order to turn echoing back on. The handler should end by specifying the default action for the signal that happened and then reraising it; this will cause the program to terminate with that signal, as if it had not had a handler. (See Section 21.4.2 [Handlers That Terminate the Process], page 415.)

Termination is the sensible ultimate outcome from a program error in most programs. However, programming systems such as Lisp that can load compiled user programs might need to keep executing even if a user program incurs an error. These programs have handlers which use longjmp to return control to the command level.

The default action for all of these signals is to cause the process to terminate. If you block or ignore these signals or establish handlers for them that return normally, your program will probably break horribly when such signals happen, unless they are generated by raise or kill instead of a real error.

When one of these program error signals terminates a process, it also writes a core dump file which records the state of the process at the time of termination. The core dump file is named ‘core’ and is written in whichever directory is current in the process at the time. (On the GNU system, you can specify the file name for core dumps with the environment variable COREFILE.) The purpose of core dump files is so that you can examine them with a debugger to investigate what caused the error.

Macro int SIGFPE

The SIGFPE signal reports a fatal arithmetic error. Although the name is derived from “floating-point exception”, this signal actually covers all arithmetic errors, including division by zero and overflow. If a program stores integer data in a location which is then used in a floatingpoint operation, this often causes an “invalid operation” exception, because the processor cannot recognize the data as a floating-point number. Actual floating-point exceptions are a complicated subject because there are many types of exceptions with subtly different meanings, and the SIGFPE signal doesn’t distinguish between them. The IEEE Standard for Binary Floating-Point Arithmetic (ANSI/IEEE Std 754-1985 and ANSI/IEEE Std 854-1987) defines various floating-point exceptions and requires conforming computer systems to report their occurrences. However, this standard does not specify how the exceptions are reported, or what kinds of handling and control the operating system can offer to the programmer.

BSD systems provide the SIGFPE handler with an extra argument that distinguishes various causes of the exception. In order to access this argument, you must define the handler to accept two arguments, which means you must cast it to a one-argument function type in order to establish the handler. The GNU library does provide this extra argument, but the value is meaningful only on operating systems that provide the information (BSD systems and GNU systems).

FPE_INTOVF_TRAP

Integer overflow (impossible in a C program unless you enable overflow trapping in a hardware-specific fashion).

FPE_INTDIV_TRAP

Integer division by zero.

FPE_SUBRNG_TRAP

Subscript-range (something that C programs never check for).

FPE_FLTOVF_TRAP

Floating overflow trap.

FPE_FLTDIV_TRAP

Floating/decimal division by zero.

FPE_FLTUND_TRAP

Floating underflow trap. (Trapping on floating underflow is not normally enabled.)

FPE_DECOVF_TRAP

Decimal overflow trap. (Only a few machines have decimal arithmetic and C never uses it.)

Macro int SIGILL

The name of this signal is derived from “illegal instruction”; it usually means your program is trying to execute garbage or a privileged instruction. Since the C compiler generates only valid instructions, SIGILL typically indicates that the executable file is corrupted, or that you are trying to execute data. Some common ways of getting into the latter situation are by passing an invalid object where a pointer to a function was expected, or by writing past the end of an automatic array (or similar problems with pointers to automatic variables) and corrupting other data on the stack such as the return address of a stack frame. SIGILL can also be generated when the stack overflows, or when the system has trouble running the handler for a signal.

Macro int SIGSEGV

This signal is generated when a program tries to read or write outside the memory that is allocated for it, or to write memory that can only be read. (Actually, the signals only occur when the program goes far enough outside to be detected by the system’s memory protection mechanism.) The name is an abbreviation for “segmentation violation”. Common ways of getting a SIGSEGV condition include dereferencing a null or uninitialized pointer, or when you use a pointer to step through an array, but fail to check for the end of the array. It varies among systems whether dereferencing a null pointer generates SIGSEGV or

SIGBUS.

Macro int SIGBUS

This signal is generated when an invalid pointer is dereferenced. Like SIGSEGV, this signal is typically the result of dereferencing an uninitialized pointer. The difference between the two is that SIGSEGV indicates an invalid access to valid memory, while SIGBUS indicates an access to an invalid address. In particular, SIGBUS signals often result from dereferencing a misaligned pointer, such as referring to a four-word integer at an address not divisible by four. (Each kind of computer has its own requirements for address alignment.) The name of this signal is an abbreviation for “bus error”.

Macro int SIGABRT

This signal indicates an error detected by the program itself and reported by calling abort. See Section 22.3.4 [Aborting a Program], page 465.

Macro int SIGIOT

Generated by the PDP-11 “iot” instruction. On most machines, this is just another name for SIGABRT.

Macro int SIGTRAP

Generated by the machine’s breakpoint instruction, and possibly other trap instructions. This signal is used by debuggers. Your program will probably only see SIGTRAP if it is somehow executing bad instructions.

Macro int SIGEMT

Emulator trap; this results from certain unimplemented instructions which might be emulated in software, or the operating system’s failure to properly emulate them.

Macro int SIGSYS

Bad system call; that is to say, the instruction to trap to the operating system was executed, but the code number for the system call to perform was invalid.

21.2.2 Termination Signals

These signals are all used to tell a process to terminate, in one way or another. They have different names because they’re used for slightly different purposes, and programs might want to handle them differently.

The reason for handling these signals is usually so your program can tidy up as appropriate before actually terminating. For example, you might want to save state information, delete temporary files, or restore the previous terminal modes. Such a handler should end by specifying the default action for the signal that happened and then reraising it; this will cause the program to terminate with that signal, as if it had not had a handler. (See Section 21.4.2 [Handlers That Terminate the Process], page 415.)

The (obvious) default action for all of these signals is to cause the process to terminate.

Macro int SIGTERM

The SIGTERM signal is a generic signal used to cause program termination. Unlike SIGKILL, this signal can be blocked, handled, and ignored. It is the normal way to politely ask a program to terminate. The shell command kill generates SIGTERM by default.

Macro int SIGINT

The SIGINT (“program interrupt”) signal is sent when the user types the INTR character (normally C-c). See Section 12.4.9 [Special Characters], page 286, for information about terminal driver support for

C-c.

Macro int SIGQUIT

The SIGQUIT signal is similar to SIGINT, except that it’s controlled by a different key—the QUIT character, usually C-\—and produces a core dump when it terminates the process, just like a program error signal. You can think of this as a program error condition “detected” by the user.

See Section 21.2.1 [Program Error Signals], page 396, for information about core dumps. See Section 12.4.9 [Special Characters], page 286, for information about terminal driver support.

Certain kinds of cleanups are best omitted in handling SIGQUIT. For example, if the program creates temporary files, it should handle the other termination requests by deleting the temporary files. But it is better for SIGQUIT not to delete them, so that the user can examine them in conjunction with the core dump.

Macro int SIGKILL

The SIGKILL signal is used to cause immediate program termination. It cannot be handled or ignored, and is therefore always fatal. It is also not possible to block this signal.

This signal is usually generated only by explicit request. Since it cannot be handled, you should generate it only as a last resort, after first trying a less drastic method such as C-c or SIGTERM. If a process does not respond to any other termination signals, sending it a SIGKILL signal will almost always cause it to go away.

In fact, if SIGKILL fails to terminate a process, that by itself constitutes an operating system bug which you should report.

The system will generate SIGKILL for a process itself under some unusual conditions where the program cannot possible continue to run (even to run a signal handler).

Macro int SIGHUP

The SIGHUP (“hang-up”) signal is used to report that the user’s terminal is disconnected, perhaps because a network or telephone connection was broken. For more information about this, see Section 12.4.6 [Control Modes], page 280.

This signal is also used to report the termination of the controlling process on a terminal to jobs associated with that session; this termination effectively disconnects all processes in the session from the controlling terminal. For more information, see Section 22.3.5 [Termination Internals], page 465.

21.2.3 Alarm Signals

These signals are used to indicate the expiration of timers. See Section 17.3 [Setting an Alarm], page 356, for information about functions that cause these signals to be sent.

The default behavior for these signals is to cause program termination. This default is rarely useful, but no other default would be useful; most of the ways of using these signals would require handler functions in any case.

Macro int SIGALRM

This signal typically indicates expiration of a timer that measures real or clock time. It is used by the alarm function, for example.

Macro int SIGVTALRM

This signal typically indicates expiration of a timer that measures CPU time used by the current process. The name is an abbreviation for “virtual time alarm”.

Macro int SIGPROF

This signal is typically indicates expiration of a timer that measures both CPU time used by the current process, and CPU time expended on behalf of the process by the system. Such a timer is used to implement code profiling facilities, hence the name of this signal.

21.2.4 Asynchronous I/O Signals

The signals listed in this section are used in conjunction with asynchronous I/O facilities. You have to take explicit action by calling fcntl to enable a particular file descriptor to generate these signals (see Section 8.12 [Interrupt-Driven Input], page 179). The default action for these signals is to ignore them.

Macro int SIGIO

This signal is sent when a file descriptor is ready to perform input or output. On most operating systems, terminals and sockets are the only kinds of files that can generate SIGIO; other kinds, including ordinary files, never generate SIGIO even if you ask them to. In the GNU system SIGIO will always be generated properly if you successfully set asynchronous mode with fcntl.

Macro int SIGURG

This signal is sent when “urgent” or out-of-band data arrives on a socket. See Section 11.8.8 [Out-of-Band Data], page 257.

Macro int SIGPOLL

This is a System V signal name, more or less similar to SIGIO. It is defined only for compatibility.

21.2.5 Job Control Signals

These signals are used to support job control. If your system doesn’t support job control, then these macros are defined but the signals themselves can’t be raised or handled.

You should generally leave these signals alone unless you really understand how job control works. See Chapter 24 [Job Control], page 479.

Macro int SIGCHLD

This signal is sent to a parent process whenever one of its child processes terminates or stops. The default action for this signal is to ignore it. If you establish a handler for this signal while there are child processes that have terminated but not reported their status via wait or waitpid (see Section 23.6 [Process Completion], page 473), whether your new handler applies to those processes or not depends on the particular operating system.

Macro int SIGCLD

This is an obsolete name for SIGCHLD.

Macro int SIGCONT

You can send a SIGCONT signal to a process to make it continue. This signal is special—it always makes the process continue if it is stopped, before the signal is delivered. The default behavior is to do nothing else. You cannot block this signal. You can set a handler, but SIGCONT always makes the process continue regardless. Most programs have no reason to handle SIGCONT; they simply resume execution without realizing they were ever stopped. You can use a handler for SIGCONT to make a program do something special when it is stopped and continued—for example, to reprint a prompt when it is suspended while waiting for input.

Macro int SIGSTOP

The SIGSTOP signal stops the process. It cannot be handled, ignored, or blocked.

Macro int SIGTSTP

The SIGTSTP signal is an interactive stop signal. Unlike SIGSTOP, this signal can be handled and ignored. Your program should handle this signal if you have a special need to leave files or system tables in a secure state when a process is stopped.

For example, programs that turn off echoing should handle SIGTSTP so they can turn echoing back on before stopping. This signal is generated when the user types the SUSP character (normally C-z). For more information about terminal driver support, see Section 12.4.9 [Special Characters], page 286.

Macro int SIGTTIN

A process cannot read from the the user’s terminal while it is running as a background job. When any process in a background job tries to read from the terminal, all of the processes in the job are sent a SIGTTIN signal. The default action for this signal is to stop the process. For more information about how this interacts with the terminal driver, see Section 24.4 [Access to the Controlling Terminal], page 481.

Macro int SIGTTOU

This is similar to SIGTTIN, but is generated when a process in a background job attempts to write to the terminal or set its modes. Again, the default action is to stop the process. SIGTTOU is only generated for an attempt to write to the terminal if the TOSTOP output mode is set; see Section 12.4.5 [Output Modes], page 279.

While a process is stopped, no more signals can be delivered to it until it is continued, except SIGKILL signals and (obviously) SIGCONT signals. The signals are marked as pending, but not delivered until the process is continued. The SIGKILL signal always causes termination of the process and can’t be blocked, handled or ignored. You can ignore SIGCONT, but it always causes the process to be continued anyway if it is stopped. Sending a SIGCONT signal to a process causes any pending stop signals for that process to be discarded. Likewise, any pending SIGCONT signals for a process are discarded when it receives a stop signal.

When a process in an orphaned process group (see Section 24.5 [Orphaned Process Groups], page 481) receives a SIGTSTP, SIGTTIN, or SIGTTOU signal and does not handle it, the process does not stop. Stopping the process would probably not be very useful, since there is no shell program that will notice it stop and allow the user to continue it. What happens instead depends on the operating system you are using. Some systems may do nothing; others may deliver another signal instead, such as SIGKILL or SIGHUP. In the GNU system, the process dies with SIGKILL; this avoids the problem of many stopped, orphaned processes lying around the system.

21.2.6 Operation Error Signals

These signals are used to report various errors generated by an operation done by the program. They do not necessarily indicate a programming error in the program, but an error that prevents an operating system call from completing. The default action for all of them is to cause the process to terminate.

Macro int SIGPIPE

Broken pipe. If you use pipes or FIFOs, you have to design your application so that one process opens the pipe for reading before another starts writing. If the reading process never starts, or terminates unexpectedly, writing to the pipe or FIFO raises a SIGPIPE signal. If SIGPIPE is blocked, handled or ignored, the offending call fails with EPIPE instead. Pipes and FIFO special files are discussed in more detail in Chapter 10 [Pipes and FIFOs], page 215. Another cause of SIGPIPE is when you try to output to a socket that isn’t connected. See Section 11.8.5.1 [Sending Data], page 250.

Macro int SIGLOST

Resource lost. This signal is generated when you have an advisory lock on an NFS file, and the NFS server reboots and forgets about your lock. In the GNU system, SIGLOST is generated when any server program dies unexpectedly. It is usually fine to ignore the signal; whatever call was made to the server that died just returns an error.

Macro int SIGXCPU

CPU time limit exceeded. This signal is generated when the process exceeds its soft resource limit on CPU time. See Section 17.6 [Limiting Resource Usage], page 362.

Macro int SIGXFSZ

File size limit exceeded. This signal is generated when the process attempts to extend a file so it exceeds the process’s soft resource limit on file size. See Section 17.6 [Limiting Resource Usage], page 362.

21.2.7 Miscellaneous Signals

These signals are used for various other purposes. In general, they will not affect your program unless it explicitly uses them for something.

Macro int SIGUSR1 Macro int SIGUSR2

The SIGUSR1 and SIGUSR2 signals are set aside for you to use any way you want. They’re useful for simple interprocess communication, if you write a signal handler for them in the program that receives the signal.

There is an example showing the use of SIGUSR1 and SIGUSR2 in Section 21.6.2 [Signaling Another Process], page 428. The default action is to terminate the process.

Macro int SIGWINCH

Window size change. This is generated on some systems (including GNU) when the terminal driver’s record of the number of rows and columns on the screen is changed. The default action is to ignore it. If a program does full-screen display, it should handle SIGWINCH. When the signal arrives, it should fetch the new screen size and reformat its display accordingly.

Macro int SIGINFO

Information request. In 4.4 BSD and the GNU system, this signal is sent to all the processes in the foreground process group of the controlling terminal when the user types the STATUS character in canonical mode; see Section 12.4.9.2 [Characters that Cause Signals], page 288. If the process is the leader of the process group, the default action is to print some status information about the system and what the process is doing. Otherwise the default is to do nothing.

21.2.8 Signal Messages

We mentioned above that the shell prints a message describing the signal that terminated a child process. The clean way to print a message describing a signal is to use the functions strsignal and psignal. These functions use a signal number to specify which kind of signal to describe. The signal number may come from the termination status of a child process (see Section 23.6 [Process Completion], page 473) or it may come from a signal handler in the same process.

Function char * strsignal (int signum)

This function returns a pointer to a statically-allocated string containing a message describing the signal signum. You should not modify the contents of this string; and, since it can be rewritten on subsequent calls, you should save a copy of it if you need to reference it later. This function is a GNU extension, declared in the header file

‘string.h’.

Function void psignal (int signum, const char *message)

This function prints a message describing the signal signum to the standard error output stream stderr; see Section 7.2 [Standard Streams], page 93.

If you call psignal with a message that is either a null pointer or an empty string, psignal just prints the message corresponding to signum, adding a trailing newline. If you supply a non-null message argument, then psignal prefixes its output with this string. It adds a colon and a space character to separate the message from the string corresponding to signum. This function is a BSD feature, declared in the header file ‘signal.h’.

There is also an array sys_siglist which contains the messages for the various signal codes. This array exists on BSD systems, unlike strsignal.

21.3 Specifying Signal Actions

The simplest way to change the action for a signal is to use the signal function. You can specify a built-in action (such as to ignore the signal), or you can establish a handler.

The GNU library also implements the more versatile sigaction facility. This section describes both facilities and gives suggestions on which to use when.

21.3.1 Basic Signal Handling

The signal function provides a simple interface for establishing an action for a particular signal. The function and associated macros are declared in the header file ‘signal.h’.

Data Type sighandler t

This is the type of signal handler functions. Signal handlers take one integer argument specifying the signal number, and have return type void. So, you should define handler functions like this:

     void handler (int signum) { . . . }

The name sighandler_t for this data type is a GNU extension.

Function sighandler_t signal (int signum, sighandler_t

action) The signal function establishes action as the action for the signal signum. The first argument, signum, identifies the signal whose behavior you want to control, and should be a signal number. The proper way to specify a signal number is with one of the symbolic signal names described in Section 21.2 [Standard Signals], page 395—don’t use an explicit number, because the numerical code for a given kind of signal may vary from operating system to operating system. The second argument, action, specifies the action to use for the signal signum. This can be one of the following:

SIG_DFL SIG_DFL specifies the default action for the particular signal. The default actions for various kinds of signals are stated in Section 21.2 [Standard Signals], page 395.

SIG_IGN SIG_IGN specifies that the signal should be ignored. Your program generally should not ignore signals that represent serious events or that are normally used to request termination. You cannot ignore the SIGKILL or SIGSTOP signals at all. You can ignore program error signals like SIGSEGV, but ignoring the error won’t enable the program to continue executing meaningfully. Ignoring user requests such as SIGINT, SIGQUIT, and SIGTSTP is unfriendly. When you do not wish signals to be delivered during a certain part of the program, the thing to do is to block them, not ignore them. See Section 21.7 [Blocking Signals], page 431.

handler Supply the address of a handler function in your program, to specify running this handler as the way to deliver the signal. For more information about defining signal handler functions, see Section 21.4 [Defining Signal Handlers], page 413.

If you set the action for a signal to SIG_IGN, or if you set it to SIG_DFL and the default action is to ignore that signal, then any pending signals of that type are discarded (even if they are blocked). Discarding the pending signals means that they will never be delivered, not even if you subsequently specify another action and unblock this kind of signal. The signal function returns the action that was previously in effect for the specified signum. You can save this value and restore it later by calling signal again. If signal can’t honor the request, it returns SIG_ERR instead. The following errno error conditions are defined for this function:

EINVAL You specified an invalid signum; or you tried to ignore or provide a handler for SIGKILL or SIGSTOP.

Here is a simple example of setting up a handler to delete temporary files when certain fatal signals happen:

#include <signal.h>
void
termination_handler (int signum)
{
struct temp_file *p;

       for (p = temp_file_list; p; p = p->next)
         unlink (p->name);
     }
     int
     main (void)
     {

…

       if (signal (SIGINT, termination_handler) == SIG_IGN)
         signal (SIGINT, SIG_IGN);
       if (signal (SIGHUP, termination_handler) == SIG_IGN)
         signal (SIGHUP, SIG_IGN);
       if (signal (SIGTERM, termination_handler) == SIG_IGN)
         signal (SIGTERM, SIG_IGN);

…

     }

Note how if a given signal was previously set to be ignored, this code avoids altering that setting. This is because non-job-control shells often ignore certain signals when starting children, and it is important for the children to respect this.

We do not handle SIGQUIT or the program error signals in this example because these are designed to provide information for debugging (a core dump), and the temporary files may give useful information.

Function sighandler_t ssignal (int signum, sighandler_t

action) The ssignal function does the same thing as signal; it is provided only for compatibility with SVID.

Macro sighandler_t SIG ERR

The value of this macro is used as the return value from signal to indicate an error.

21.3.2 Advanced Signal Handling

The sigaction function has the same basic effect as signal: to specify how a signal should be handled by the process. However, sigaction offers more control, at the expense of more complexity. In particular, sigaction allows you to specify additional flags to control when the signal is generated and how the handler is invoked.

The sigaction function is declared in ‘signal.h’.

Data Type struct sigaction

Structures of type struct sigaction are used in the sigaction function to specify all the information about how to handle a particular signal. This structure contains at least the following members:

sighandler_t sa_handler

This is used in the same way as the action argument to the signal function. The value can be SIG_DFL, SIG_ IGN, or a function pointer. See Section 21.3.1 [Basic Signal Handling], page 406.

sigset_t sa_mask

This specifies a set of signals to be blocked while the handler runs. Blocking is explained in Section 21.7.5 [Blocking Signals for a Handler], page 436. Note that the signal that was delivered is automatically blocked by default before its handler is started; this is true regardless of the value in sa_mask. If you want that signal not to be blocked within its handler, you must write code in the handler to unblock it.

int sa_flags

This specifies various flags which can affect the behavior of the signal. These are described in more detail in Section 21.3.5 [Flags for “codesigaction], page 412.

Function int sigaction (int signum, const struct sigaction
      *action, struct sigaction *old-action)

The action argument is used to set up a new action for the signal signum, while the old-action argument is used to return information about the action previously associated with this symbol. (In other words, old-action has the same purpose as the signal function’s return value—you can check to see what the old action in effect for the signal was, and restore it later if you want.) Either action or old-action can be a null pointer. If old-action is a null pointer, this simply suppresses the return of information about the old action. If action is a null pointer, the action associated with the signal signum is unchanged; this allows you to inquire about how a signal is being handled without changing that handling. The return value from sigaction is zero if it succeeds, and -1 on failure. The following errno error conditions are defined for this function:

EINVAL The signum argument is not valid, or you are trying to trap or ignore SIGKILL or SIGSTOP.

21.3.3 Interaction of signal and sigaction

It’s possible to use both the signal and sigaction functions within a single program, but you have to be careful because they can interact in slightly strange ways.

The sigaction function specifies more information than the signal function, so the return value from signal cannot express the full range of sigaction possibilities. Therefore, if you use signal to save and later reestablish an action, it may not be able to reestablish properly a handler that was established with sigaction.

To avoid having problems as a result, always use sigaction to save and restore a handler if your program uses sigaction at all. Since sigaction is more general, it can properly save and reestablish any action, regardless of whether it was established originally with signal or sigaction.

On some systems if you establish an action with signal and then examine it with sigaction, the handler address that you get may not be the same as what you specified with signal. It may not even be suitable for use as an action argument with signal. But you can rely on using it as an argument to sigaction. This problem never happens on the GNU system.

So, you’re better off using one or the other of the mechanisms consistently within a single program.

Portability Note: The basic signal function is a feature of ISO C, while sigaction is part of the POSIX.1 standard. If you are concerned about portability to non-POSIX systems, then you should use the signal function instead.

21.3.4 sigaction Function Example

In Section 21.3.1 [Basic Signal Handling], page 406, we gave an example of establishing a simple handler for termination signals using signal. Here is an equivalent example using sigaction:

     #include <signal.h>
     void
     termination_handler (int signum)
     {
       struct temp_file *p;
       for (p = temp_file_list; p; p = p->next)
         unlink (p->name);
     }
     int
     main (void)

Chapter 21: Signal Handling 411

{

…

struct sigaction new_action, old_action;

/* Set up the structure to specify the new action. */

new_action.sa_handler = termination_handler;
sigemptyset (&new_action.sa_mask);
new_action.sa_flags = 0;
sigaction (SIGINT, NULL, &old_action);
if (old_action.sa_handler != SIG_IGN)
  sigaction (SIGINT, &new_action, NULL);
sigaction (SIGHUP, NULL, &old_action);
if (old_action.sa_handler != SIG_IGN)
  sigaction (SIGHUP, &new_action, NULL);
sigaction (SIGTERM, NULL, &old_action);
if (old_action.sa_handler != SIG_IGN)
  sigaction (SIGTERM, &new_action, NULL);

…

}

The program just loads the new_action structure with the desired parameters and passes it in the sigaction call. The usage of sigemptyset is described later; see Section 21.7 [Blocking Signals], page 431.

As in the example using signal, we avoid handling signals previously set to be ignored. Here we can avoid altering the signal handler even momentarily, by using the feature of sigaction that lets us examine the current action without specifying a new one.

Here is another example. It retrieves information about the current action for SIGINT without changing that action.

struct sigaction query_action;
if (sigaction (SIGINT, NULL, &query_action) < 0)

/* sigaction returns -1 in case of error. */

else if (query_action.sa_handler == SIG_DFL)

/* SIGINT is handled in the default, fatal manner. */

else if (query_action.sa_handler == SIG_IGN)
/* SIGINT is ignored. */
else

/* A programmer-defined signal handler is in effect. */

21.3.5 Flags for sigaction

The sa_flags member of the sigaction structure is a catch-all for special features. Most of the time, SA_RESTART is a good value to use for this field.

The value of sa_flags is interpreted as a bit mask. Thus, you should choose the flags you want to set, or those flags together, and store the result in the sa_flags member of your sigaction structure.

Each signal number has its own set of flags. Each call to sigaction affects one particular signal number, and the flags that you specify apply only to that particular signal.

In the GNU C library, establishing a handler with signal sets all the flags to zero except for SA_RESTART, whose value depends on the settings you have made with siginterrupt. See Section 21.5 [Primitives Interrupted by Signals], page 425, to see what this is about.

These macros are defined in the header file ‘signal.h’.

Macro int SA NOCLDSTOP

This flag is meaningful only for the SIGCHLD signal. When the flag is set, the system delivers the signal for a terminated child process but not for one that is stopped. By default, SIGCHLD is delivered for both terminated children and stopped children. Setting this flag for a signal other than SIGCHLD has no effect.

Macro int SA ONSTACK

If this flag is set for a particular signal number, the system uses the signal stack when delivering that kind of signal. See Section 21.9 [Using a Separate Signal Stack], page 442. If a signal with this flag arrives and you have not set a signal stack, the system terminates the program

with SIGILL.

Macro int SA RESTART

This flag controls what happens when a signal is delivered during certain primitives (such as open, read or write), and the signal handler returns normally. There are two alternatives: the library function can resume, or it can return failure with error code EINTR. The choice is controlled by the SA_RESTART flag for the particular kind of signal that was delivered. If the flag is set, returning from a handler resumes the library function. If the flag is clear, returning from a handler makes the function fail. See Section 21.5 [Primitives Interrupted by Signals], page 425.

21.3.6 Initial Signal Actions

When a new process is created (see Section 23.4 [Creating a Process], page 469), it inherits handling of signals from its parent process. However, when you load a new process image using the exec function (see Section 23.5 [Executing a File], page 470), any signals that you’ve defined your own handlers for revert to their SIG_DFL handling. (If you think about it a little, this makes sense; the handler functions from the old program are specific to that program, and aren’t even present in the address space of the new program image.) Of course, the new program can establish its own handlers.

When a program is run by a shell, the shell normally sets the initial actions for the child process to SIG_DFL or SIG_IGN, as appropriate. It’s a good idea to check to make sure that the shell has not set up an initial action of SIG_IGN before you establish your own signal handlers.

Here is an example of how to establish a handler for SIGHUP, but not if SIGHUP is currently ignored:

…

     struct sigaction temp;
     sigaction (SIGHUP, NULL, &temp);
     if (temp.sa_handler != SIG_IGN)
       {
         temp.sa_handler = handle_sighup;
         sigemptyset (&temp.sa_mask);
         sigaction (SIGHUP, &temp, NULL);
       }

21.4 Defining Signal Handlers

This section describes how to write a signal handler function that can be established with the signal or sigaction functions.

A signal handler is just a function that you compile together with the rest of the program. Instead of directly invoking the function, you use signal or sigaction to tell the operating system to call it when a signal arrives. This is known as establishing the handler. See Section 21.3 [Specifying Signal Actions], page 406.

There are two basic strategies you can use in signal handler functions:

• You can have the handler function note that the signal arrived by tweaking some global data structures, and then return normally.

• You can have the handler function terminate the program or transfer control to a point where it can recover from the situation that caused the signal.

You need to take special care in writing handler functions because they can be called asynchronously. That is, a handler might be called at any point in the program, unpredictably. If two signals arrive during a very short interval, one handler can run within another. This section describes what your handler should do, and what you should avoid.

21.4.1 Signal Handlers that Return

Handlers which return normally are usually used for signals such as SIGALRM and the I/O and interprocess communication signals. But a handler for SIGINT might also return normally after setting a flag that tells the program to exit at a convenient time.

It is not safe to return normally from the handler for a program error signal, because the behavior of the program when the handler function returns is not defined after a program error. See Section 21.2.1 [Program Error Signals], page 396.

Handlers that return normally must modify some global variable in order to have any effect. Typically, the variable is one that is examined periodically by the program during normal operation. Its data type should be sig_ atomic_t for reasons described in Section 21.4.7 [Atomic Data Access and Signal Handling], page 423.

Here is a simple example of such a program. It executes the body of the loop until it has noticed that a SIGALRM signal has arrived. This technique is useful because it allows the iteration in progress when the signal arrives to complete before the loop exits.

     #include <signal.h>
     #include <stdio.h>
     #include <stdlib.h>

/* This flag controls termination of the main loop. */

     volatile sig_atomic_t keep_going = 1;

/* The signal handler just clears the flag and re-enables itself. */

     void
     catch_alarm (int sig)
     {
       keep_going = 0;
       signal (sig, catch_alarm);
     }
     void
     do_stuff (void)
     {
       puts ("Doing stuff while waiting for alarm....");

}
int
main (void)
{

/* Establish a handler for SIGALRM signals. */

signal (SIGALRM, catch_alarm);

/* Set an alarm to go off in a little while. */

alarm (2);

/* Check the flag once in a while to see when to quit. */

while (keep_going)
  do_stuff ();
return EXIT_SUCCESS;
}

21.4.2 Handlers That Terminate the Process

Handler functions that terminate the program are typically used to cause orderly cleanup or recovery from program error signals and interactive interrupts.

The cleanest way for a handler to terminate the process is to raise the same signal that ran the handler in the first place. Here is how to do this:

volatile sig_atomic_t fatal_error_in_progress = 0;
void
fatal_error_signal (int sig)
{

/* Since this handler is established for more than one kind of signal, it might still get invoked recursively by delivery of some other kind of signal. Use a static variable to keep track of that. */

if (fatal_error_in_progress)
  raise (sig);
fatal_error_in_progress = 1;

/* Now do the clean up actions:

- reset terminal modes

- kill child processes

- remove lock files */ … /* Now reraise the signal. Since the signal is blocked, it will receive its default handling, which is to terminate the process. We could just call exit or abort, but reraising the signal sets the return status from the process correctly. */

       raise (sig);
     }

21.4.3 Nonlocal Control Transfer in Handlers

You can do a nonlocal transfer of control out of a signal handler using the setjmp and longjmp facilities (see Chapter 20 [Non-Local Exits], page 389).

When the handler does a nonlocal control transfer, the part of the program that was running will not continue. If this part of the program was in the middle of updating an important data structure, the data structure will remain inconsistent. Since the program does not terminate, the inconsistency is likely to be noticed later on.

There are two ways to avoid this problem. One is to block the signal for the parts of the program that update important data structures. Blocking the signal delays its delivery until it is unblocked, once the critical updating is finished. See Section 21.7 [Blocking Signals], page 431.

The other way to re-initialize the crucial data structures in the signal handler, or make their values consistent.

Here is a rather schematic example showing the reinitialization of one global variable.

     #include <signal.h>
     #include <setjmp.h>
     jmp_buf return_to_top_level;
     volatile sig_atomic_t waiting_for_input;
     void
     handle_sigint (int signum)
     {

/* We may have been waiting for input when the signal arrived, but we are no longer waiting once we transfer control. */

       waiting_for_input = 0;
       longjmp (return_to_top_level, 1);
     }

     int
     main (void)
     {

…

       signal (SIGINT, sigint_handler);

…

       while (1) {
         prepare_for_command ();
         if (setjmp (return_to_top_level) == 0)
           read_and_execute_command ();
       }
     }

/* Imagine this is a subroutine used by various commands. */

     char *
     read_data ()
     {
       if (input_from_terminal) {
         waiting_for_input = 1;

…

         waiting_for_input = 0;
       } else {

…

       }
     }

21.4.4 Signals Arriving While a Handler Runs

What happens if another signal arrives while your signal handler function is running?

When the handler for a particular signal is invoked, that signal is automatically blocked until the handler returns. That means that if two signals of the same kind arrive close together, the second one will be held until the first has been handled. (The handler can explicitly unblock the signal using sigprocmask, if you want to allow more signals of this type to arrive; see Section 21.7.3 [Process Signal Mask], page 433.)

However, your handler can still be interrupted by delivery of another kind of signal. To avoid this, you can use the sa_mask member of the action structure passed to sigaction to explicitly specify which signals should be blocked while the signal handler runs. These signals are in addition to the signal for which the handler was invoked, and any other signals that are normally blocked by the process. See Section 21.7.5 [Blocking Signals for a Handler], page 436.

When the handler returns, the set of blocked signals is restored to the value it had before the handler ran. So using sigprocmask inside the handler only affects what signals can arrive during the execution of the handler itself, not what signals can arrive once the handler returns.

Portability Note: Always use sigaction to establish a handler for a signal that you expect to receive asynchronously, if you want your program to work properly on System V Unix. On this system, the handling of a signal whose handler was established with signal automatically sets the signal’s action back to SIG_DFL, and the handler must re-establish itself each time it runs. This practice, while inconvenient, does work when signals cannot arrive in succession. However, if another signal can arrive right away, it may arrive before the handler can re-establish itself. Then the second signal would receive the default handling, which could terminate the process.

21.4.5 Signals Close Together Merge into One

If multiple signals of the same type are delivered to your process before your signal handler has a chance to be invoked at all, the handler may only be invoked once, as if only a single signal had arrived. In effect, the signals merge into one. This situation can arise when the signal is blocked, or in a multiprocessing environment where the system is busy running some other processes while the signals are delivered. This means, for example, that you cannot reliably use a signal handler to count signals. The only distinction you can reliably make is whether at least one signal has arrived since a given time in the past.

Here is an example of a handler for SIGCHLD that compensates for the fact that the number of signals recieved may not equal the number of child processes generate them. It assumes that the program keeps track of all the child processes with a chain of structures as follows:

     struct process
     {
       struct process *next;

/* The process ID of this child. */

       int pid;

/* The descriptor of the pipe or pseudo terminal on which output comes from this child. */

       int input_descriptor;

/* Nonzero if this process has stopped or terminated. */

       sig_atomic_t have_status;

/* The status of this child; 0 if running,

otherwise a status value from waitpid. */

       int status;
     };
     struct process *process_list;

This example also uses a flag to indicate whether signals have arrived since some time in the past—whenever the program last cleared it to zero.

/* Nonzero means some child’s status has changed

so look at process_list for the details. */

int process_status_change;

Here is the handler itself:

void
sigchld_handler (int signo)
{
int old_errno = errno;
while (1) {
register int pid;
int w;
struct process *p;

/* Keep asking for a status until we get a definitive result. */

do
  {
    errno = 0;
    pid = waitpid (WAIT_ANY, &w, WNOHANG | WUNTRACED);
  }
while (pid <= 0 && errno == EINTR);
if (pid <= 0) {

/* A real failure means there are no more

stopped or terminated child processes, so return. */

  errno = old_errno;
  return;
}

/* Find the process that signaled us, and record its status. */

for (p = process_list; p; p = p->next)
  if (p->pid == pid) {
    p->status = w;

/* Indicate that the status field

has data to look at. We do this only after storing it. */

    p->have_status = 1;

/* If process has terminated, stop waiting for its output. */

    if (WIFSIGNALED (w) || WIFEXITED (w))
      if (p->input_descriptor)
        FD_CLR (p->input_descriptor, &input_wait_mask);

/* The program should check this flag from time to time to see if there is any news in process_list. */

             ++process_status_change;
           }

/* Loop around to handle all the processes that have something to tell us. */

       }
     }

Here is the proper way to check the flag process_status_change:

     if (process_status_change) {
       struct process *p;
       process_status_change = 0;
       for (p = process_list; p; p = p->next)
         if (p->have_status) {

… Examine p->status …

         }
     }

It is vital to clear the flag before examining the list; otherwise, if a signal were delivered just before the clearing of the flag, and after the appropriate element of the process list had been checked, the status change would go unnoticed until the next signal arrived to set the flag again. You could, of course, avoid this problem by blocking the signal while scanning the list, but it is much more elegant to guarantee correctness by doing things in the right order.

The loop which checks process status avoids examining p->status until it sees that status has been validly stored. This is to make sure that the status cannot change in the middle of accessing it. Once p->have_status is set, it means that the child process is stopped or terminated, and in either case, it cannot stop or terminate again until the program has taken notice. See Section 21.4.7.3 [Atomic Usage Patterns], page 425, for more information about coping with interruptions during accessings of a variable.

Here is another way you can test whether the handler has run since the last time you checked. This technique uses a counter which is never changed outside the handler. Instead of clearing the count, the program remembers the previous value and sees whether it has changed since the previous check. The advantage of this method is that different parts of the program can check independently, each part checking whether there has been a signal since that part last checked.

     sig_atomic_t process_status_change;
     sig_atomic_t last_process_status_change;

…

     {
       sig_atomic_t prev = last_process_status_change;
       last_process_status_change = process_status_change;
       if (last_process_status_change != prev) {
         struct process *p;
         for (p = process_list; p; p = p->next)
           if (p->have_status) {

… Examine p->status …

           }
       }
     }

21.4.6 Signal Handling and Nonreentrant Functions

Handler functions usually don’t do very much. The best practice is to write a handler that does nothing but set an external variable that the program checks regularly, and leave all serious work to the program. This is best because the handler can be called at asynchronously, at unpredictable times—perhaps in the middle of a primitive function, or even between the beginning and the end of a C operator that requires multiple instructions. The data structures being manipulated might therefore be in an inconsistent state when the handler function is invoked. Even copying one int variable into another can take two instructions on most machines.

This means you have to be very careful about what you do in a signal handler.

• If your handler needs to access any global variables from your program, declare those variables volatile. This tells the compiler that the value of the variable might change asynchronously, and inhibits certain optimizations that would be invalidated by such modifications.

• If you call a function in the handler, make sure it is reentrant with respect to signals, or else make sure that the signal cannot interrupt a call to a related function.

A function can be non-reentrant if it uses memory that is not on the stack.

• If a function uses a static variable or a global variable, or a dynamicallyallocated object that it finds for itself, then it is non-reentrant and any two calls to the function can interfere. For example, suppose that the signal handler uses gethostbyname. This function returns its value in a static object, reusing the same object each time. If the signal happens to arrive during a call to gethostbyname, or even after one (while the program is still using the value), it will clobber the value that the program asked for. However, if the program does not use gethostbyname or any other function that returns information in the same object, or if it always blocks signals around each use, then you are safe. There are a large number of library functions that return values in a fixed object, always reusing the same object in this fashion, and all of them cause the same problem. The description of a function in this manual always mentions this behavior.

• If a function uses and modifies an object that you supply, then it is potentially non-reentrant; two calls can interfere if they use the same object. This case arises when you do I/O using streams. Suppose that the signal handler prints a message with fprintf. Suppose that the program was in the middle of an fprintf call using the same stream when the signal was delivered. Both the signal handler’s message and the program’s data could be corrupted, because both calls operate on the same data structure—the stream itself. However, if you know that the stream that the handler uses cannot possibly be used by the program at a time when signals can arrive, then you are safe. It is no problem if the program uses some other stream.

• On most systems, malloc and free are not reentrant, because they use a static data structure which records what memory blocks are free. As a result, no library functions that allocate or free memory are reentrant. This includes functions that allocate space to store a result. The best way to avoid the need to allocate memory in a handler is to allocate in advance space for signal handlers to use. The best way to avoid freeing memory in a handler is to flag or record the objects to be freed, and have the program check from time to time whether anything is waiting to be freed. But this must be done with care, because placing an object on a chain is not atomic, and if it is interrupted by another signal handler that does the same thing, you could “lose” one of the objects. The relocating allocation functions (see Section 3.6 [Relocating Allocator], page 58) are certainly not safe to use in a signal handler.

• Any function that modifies errno is non-reentrant, but you can correct for this: in the handler, save the original value of errno and restore it before returning normally. This prevents errors that occur within the signal handler from being confused with errors from system calls at the point the program is interrupted to run the handler. This technique is generally applicable; if you want to call in a handler a function that modifies a particular object in memory, you can make this safe by saving and restoring that object.

• Merely reading from a memory object is safe provided that you can deal with any of the values that might appear in the object at a time when the signal can be delivered. Keep in mind that assignment to some data types requires more than one instruction, which means that the handler could run “in the middle of” an assignment to the variable if its type is not atomic. See Section 21.4.7 [Atomic Data Access and Signal Handling], page 423.

• Merely writing into a memory object is safe as long as a sudden change in the value, at any time when the handler might run, will not disturb anything.

21.4.7 Atomic Data Access and Signal Handling

Whether the data in your application concerns atoms, or mere text, you have to be careful about the fact that access to a single datum is not necessarily atomic. This means that it can take more than one instruction to read or write a single object. In such cases, a signal handler might in the middle of reading or writing the object.

There are three ways you can cope with this problem. You can use data types that are always accessed atomically; you can carefully arrange that nothing untoward happens if an access is interrupted, or you can block all signals around any access that had better not be interrupted (see Section 21.7 [Blocking Signals], page 431).

21.4.7.1 Problems with Non-Atomic Access

Here is an example which shows what can happen if a signal handler runs in the middle of modifying a variable. (Interrupting the reading of a variable can also lead to paradoxical results, but here we only show writing.)

     #include <signal.h>
     #include <stdio.h>
     struct two_words { int a, b; } memory;
     void
     handler(int signum)
     {
        printf ("%d,%d\n", memory.a, memory.b);
        alarm (1);
     }

     int
     main (void)
     {
        static struct two_words zeros = { 0, 0 }, ones = { 1, 1 };
        signal (SIGALRM, handler);
        memory = zeros;
        alarm (1);
        while (1)
          {
            memory = zeros;
            memory = ones;
          }
     }

This program fills memory with zeros, ones, zeros, ones, alternating forever; meanwhile, once per second, the alarm signal handler prints the current contents. (Calling printf in the handler is safe in this program because it is certainly not being called outside the handler when the signal happens.)

Clearly, this program can print a pair of zeros or a pair of ones. But that’s not all it can do! On most machines, it takes several instructions to store a new value in memory, and the value is stored one word at a time. If the signal is delivered in between these instructions, the handler might find that memory.a is zero and memory.b is one (or vice versa).

On some machines it may be possible to store a new value in memory with just one instruction that cannot be interrupted. On these machines, the handler will always print two zeros or two ones.

21.4.7.2 Atomic Types

To avoid uncertainty about interrupting access to a variable, you can use a particular data type for which access is always atomic: sig_atomic_ t. Reading and writing this data type is guaranteed to happen in a single instruction, so there’s no way for a handler to run “in the middle” of an access.

The type sig_atomic_t is always an integer data type, but which one it is, and how many bits it contains, may vary from machine to machine.

Data Type sig atomic t

This is an integer data type. Objects of this type are always accessed atomically.

In practice, you can assume that int and other integer types no longer than int are atomic. You can also assume that pointer types are atomic; that is very convenient. Both of these are true on all of the machines that the GNU C library supports, and on all POSIX systems we know of.

21.4.7.3 Atomic Usage Patterns

Certain patterns of access avoid any problem even if an access is interrupted. For example, a flag which is set by the handler, and tested and cleared by the main program from time to time, is always safe even if access actually requires two instructions. To show that this is so, we must consider each access that could be interrupted, and show that there is no problem if it is interrupted.

An interrupt in the middle of testing the flag is safe because either it’s recognized to be nonzero, in which case the precise value doesn’t matter, or it will be seen to be nonzero the next time it’s tested.

An interrupt in the middle of clearing the flag is no problem because either the value ends up zero, which is what happens if a signal comes in just before the flag is cleared, or the value ends up nonzero, and subsequent events occur as if the signal had come in just after the flag was cleared. As long as the code handles both of these cases properly, it can also handle a signal in the middle of clearing the flag. (This is an example of the sort of reasoning you need to do to figure out whether non-atomic usage is safe.)

Sometimes you can insure uninterrupted access to one object by protecting its use with another object, perhaps one whose type guarantees atomicity. See Section 21.4.5 [Signals Close Together Merge into One], page 418, for an example.

21.5 Primitives Interrupted by Signals

A signal can arrive and be handled while an I/O primitive such as open or read is waiting for an I/O device. If the signal handler returns, the system faces the question: what should happen next?

POSIX specifies one approach: make the primitive fail right away. The error code for this kind of failure is EINTR. This is flexible, but usually inconvenient. Typically, POSIX applications that use signal handlers must check for EINTR after each library function that can return it, in order to try the call again. Often programmers forget to check, which is a common source of error.

The GNU library provides a convenient way to retry a call after a temporary failure, with the macro TEMP_FAILURE_RETRY:

Macro TEMP FAILURE RETRY (expression)

This macro evaluates expression once. If it fails and reports error code EINTR, TEMP_FAILURE_RETRY evaluates it again, and over and over until the result is not a temporary failure. The value returned by TEMP_FAILURE_RETRY is whatever value expression produced.

BSD avoids EINTR entirely and provides a more convenient approach: to restart the interrupted primitive, instead of making it fail. If you choose this approach, you need not be concerned with EINTR.

You can choose either approach with the GNU library. If you use sigaction to establish a signal handler, you can specify how that handler should behave. If you specify the SA_RESTART flag, return from that handler will resume a primitive; otherwise, return from that handler will cause EINTR. See Section 21.3.5 [Flags for “codesigaction], page 412.

Another way to specify the choice is with the siginterrupt function. See Section 21.10.1 [BSD Function to Establish a Handler], page 445.

When you don’t specify with sigaction or siginterrupt what a particular handler should do, it uses a default choice. The default choice in the GNU library depends on the feature test macros you have defined. If you define _BSD_SOURCE or _GNU_SOURCE before calling signal, the default is to resume primitives; otherwise, the default is to make them fail with EINTR. (The library contains alternate versions of the signal function, and the feature test macros determine which one you really call.) See Section 1.3.4 [Feature Test Macros], page 8.

The description of each primitive affected by this issue lists EINTR among the error codes it can return.

There is one situation where resumption never happens no matter which choice you make: when a data-transfer function such as read or write is interrupted by a signal after transferring part of the data. In this case, the function returns the number of bytes already transferred, indicating partial success.

This might at first appear to cause unreliable behavior on record-oriented devices (including datagram sockets; see Section 11.9 [Datagram Socket Operations], page 260), where splitting one read or write into two would read or write two records. Actually, there is no problem, because interruption after a partial transfer cannot happen on such devices; they always transfer an entire record in one burst, with no waiting once data transfer has started.

21.6 Generating Signals

Besides signals that are generated as a result of a hardware trap or interrupt, your program can explicitly send signals to itself or to another process.

21.6.1 Signaling Yourself

A process can send itself a signal with the raise function. This function is declared in ‘signal.h’.

Function int raise (int signum)

The raise function sends the signal signum to the calling process. It returns zero if successful and a nonzero value if it fails. About the only reason for failure would be if the value of signum is invalid.

Function int gsignal (int signum)

The gsignal function does the same thing as raise; it is provided only for compatibility with SVID.

One convenient use for raise is to reproduce the default behavior of a signal that you have trapped. For instance, suppose a user of your program types the SUSP character (usually C-z; see Section 12.4.9 [Special Characters], page 286) to send it an interactive stop stop signal (SIGTSTP), and you want to clean up some internal data buffers before stopping. You might set this up like this:

     #include <signal.h>

/* When a stop signal arrives, set the action back to the default and then resend the signal after doing cleanup actions. */

     void
     tstp_handler (int sig)
     {
       signal (SIGTSTP, SIG_DFL);

/* Do cleanup actions here. */ …

       raise (SIGTSTP);
     }

/* When the process is continued again, restore the signal handler. */

     void
     cont_handler (int sig)
     {
       signal (SIGCONT, cont_handler);
       signal (SIGTSTP, tstp_handler);
     }

/* Enable both handlers during program initialization. */

  int
  main (void)
  {
    signal (SIGCONT, cont_handler);
    signal (SIGTSTP, tstp_handler);

…

  }

Portability note: raise was invented by the ISO C committee. Older systems may not support it, so using kill may be more portable. See Section 21.6.2 [Signaling Another Process], page 428.

21.6.2 Signaling Another Process

The kill function can be used to send a signal to another process. In spite of its name, it can be used for a lot of things other than causing a process to terminate. Some examples of situations where you might want to send signals between processes are:

• A parent process starts a child to perform a task—perhaps having the child running an infinite loop—and then terminates the child when the task is no longer needed.

• A process executes as part of a group, and needs to terminate or notify the other processes in the group when an error or other event occurs.

• Two processes need to synchronize while working together.

This section assumes that you know a little bit about how processes work. For more information on this subject, see Chapter 23 [Processes], page 467.

The kill function is declared in ‘signal.h’.

Function int kill (pid_t pid, int signum)

The kill function sends the signal signum to the process or process group specified by pid. Besides the signals listed in Section 21.2 [Standard Signals], page 395, signum can also have a value of zero to check the validity of the pid. The pid specifies the process or process group to receive the signal:

pid > 0 The process whose identifier is pid.

pid == 0 All processes in the same process group as the sender.

pid < -1 The process group whose identifier is −pid.

pid == -1 If the process is privileged, send the signal to all processes except for some special system processes. Otherwise, send the signal to all processes with the same effective user ID.

A process can send a signal signum to itself with a call like kill (getpid(), signum). If kill is used by a process to send a signal to itself, and the signal is not blocked, then kill delivers at least one signal (which might be some other pending unblocked signal instead of the signal signum) to that process before it returns. The return value from kill is zero if the signal can be sent successfully. Otherwise, no signal is sent, and a value of -1 is returned. If pid specifies sending a signal to several processes, kill succeeds if it can send the signal to at least one of them. There’s no way you can tell which of the processes got the signal or whether all of them did. The following errno error conditions are defined for this function:

EINVAL The signum argument is an invalid or unsupported number.

EPERM You do not have the privilege to send a signal to the process or any of the processes in the process group named by pid.

ESCRH The pid argument does not refer to an existing process or group.

Function int killpg (int pgid, int signum)

This is similar to kill, but sends signal signum to the process group pgid. This function is provided for compatibility with BSD; using kill to do this is more portable.

As a simple example of kill, the call kill (getpid (), sig) has the same effect as raise (sig).

21.6.3 Permission for using kill

There are restrictions that prevent you from using kill to send signals to any random process. These are intended to prevent antisocial behavior such as arbitrarily killing off processes belonging to another user. In typical use, kill is used to pass signals between parent, child, and sibling processes, and in these situations you normally do have permission to send signals. The only common exception is when you run a setuid program in a child process; if the program changes its real UID as well as its effective UID, you may not have permission to send a signal. The su program does this.

Whether a process has permission to send a signal to another process is determined by the user IDs of the two processes. This concept is discussed in detail in Section 26.2 [The Persona of a Process], page 511.

Generally, for a process to be able to send a signal to another process, either the sending process must belong to a privileged user (like ‘root’), or the real or effective user ID of the sending process must match the real or effective user ID of the receiving process. If the receiving process has changed its effective user ID from the set-user-ID mode bit on its process image file, then the owner of the process image file is used in place of its current effective user ID. In some implementations, a parent process might be able to send signals to a child process even if the user ID’s don’t match, and other implementations might enforce other restrictions.

The SIGCONT signal is a special case. It can be sent if the sender is part of the same session as the receiver, regardless of user IDs.

21.6.4 Using kill for Communication

Here is a longer example showing how signals can be used for interprocess communication. This is what the SIGUSR1 and SIGUSR2 signals are provided for. Since these signals are fatal by default, the process that is supposed to receive them must trap them through signal or sigaction.

In this example, a parent process forks a child process and then waits for the child to complete its initialization. The child process tells the parent when it is ready by sending it a SIGUSR1 signal, using the kill function.

     #include <signal.h>
     #include <stdio.h>
     #include <sys/types.h>
     #include <unistd.h>

/* When a SIGUSR1 signal arrives, set this variable. */

     volatile sig_atomic_t usr_interrupt = 0;
     void
     synch_signal (int sig)
     {
       usr_interrupt = 1;
     }

/* The child process executes this function. */

     void
     child_function (void)
     {

/* Perform initialization. */

       printf ("I'm here!!!  My pid is %d.\n", (int) getpid ());

/* Let parent know you’re done. */

       kill (getppid (), SIGUSR1);

/* Continue with execution. */

       puts ("Bye, now....");
       exit (0);

     }
     int
     main (void)
     {
       struct sigaction usr_action;
       sigset_t block_mask;
       pid_t child_id;

/* Establish the signal handler. */

       sigfillset (&block_mask);
       usr_action.sa_handler = synch_signal;
       usr_action.sa_mask = block_mask;
       usr_action.sa_flags = 0;
       sigaction (SIGUSR1, &usr_action, NULL);

/* Create the child process. */

       child_id = fork ();
       if (child_id == 0)
         child_function ();          /* Does not return. */

/* Busy wait for the child to send a signal. */

       while (!usr_interrupt)
         ;

/* Now continue execution. */

       puts ("That's all, folks!");
       return 0;
     }

This example uses a busy wait, which is bad, because it wastes CPU cycles that other programs could otherwise use. It is better to ask the system to wait until the signal arrives. See the example in Section 21.8 [Waiting for a Signal], page 439.

21.7 Blocking Signals

Blocking a signal means telling the operating system to hold it and deliver it later. Generally, a program does not block signals indefinitely—it might as well ignore them by setting their actions to SIG_IGN. But it is useful to block signals briefly, to prevent them from interrupting sensitive operations. For instance:

• You can use the sigprocmask function to block signals while you modify global variables that are also modified by the handlers for these signals.

• You can set sa_mask in your sigaction call to block certain signals while a particular signal handler runs. This way, the signal handler can run without being interrupted itself by signals.

21.7.1 Why Blocking Signals is Useful

Temporary blocking of signals with sigprocmask gives you a way to prevent interrupts during critical parts of your code. If signals arrive in that part of the program, they are delivered later, after you unblock them.

One example where this is useful is for sharing data between a signal handler and the rest of the program. If the type of the data is not sig_ atomic_t (see Section 21.4.7 [Atomic Data Access and Signal Handling], page 423), then the signal handler could run when the rest of the program has only half finished reading or writing the data. This would lead to confusing consequences.

To make the program reliable, you can prevent the signal handler from running while the rest of the program is examining or modifying that data— by blocking the appropriate signal around the parts of the program that touch the data.

Blocking signals is also necessary when you want to perform a certain action only if a signal has not arrived. Suppose that the handler for the signal sets a flag of type sig_atomic_t; you would like to test the flag and perform the action if the flag is not set. This is unreliable. Suppose the signal is delivered immediately after you test the flag, but before the consequent action: then the program will perform the action even though the signal has arrived.

The only way to test reliably for whether a signal has yet arrived is to test while the signal is blocked.

21.7.2 Signal Sets

All of the signal blocking functions use a data structure called a signal set to specify what signals are affected. Thus, every activity involves two stages: creating the signal set, and then passing it as an argument to a library function.

These facilities are declared in the header file ‘signal.h’.

Data Type sigset t

The sigset_t data type is used to represent a signal set. Internally, it may be implemented as either an integer or structure type. For portability, use only the functions described in this section to initialize, change, and retrieve information from sigset_t objects— don’t try to manipulate them directly.

There are two ways to initialize a signal set. You can initially specify it to be empty with sigemptyset and then add specified signals individually. Or you can specify it to be full with sigfillset and then delete specified signals individually.

You must always initialize the signal set with one of these two functions before using it in any other way. Don’t try to set all the signals explicitly because the sigset_t object might include some other information (like a version field) that needs to be initialized as well. (In addition, it’s not wise to put into your program an assumption that the system has no signals aside from the ones you know about.)

Function int sigemptyset (sigset_t *set)

This function initializes the signal set set to exclude all of the defined signals. It always returns 0.

Function int sigfillset (sigset_t *set)

This function initializes the signal set set to include all of the defined signals. Again, the return value is 0.

Function int sigaddset (sigset_t *set, int signum)

This function adds the signal signum to the signal set set. All sigaddset does is modify set; it does not block or unblock any signals. The return value is 0 on success and -1 on failure. The following errno error condition is defined for this function:

EINVAL The signum argument doesn’t specify a valid signal.

Function int sigdelset (sigset_t *set, int signum)

This function removes the signal signum from the signal set set. All sigdelset does is modify set; it does not block or unblock any signals. The return value and error conditions are the same as for sigaddset.

Finally, there is a function to test what signals are in a signal set:

Function int sigismember (const sigset_t *set, int signum)

The sigismember function tests whether the signal signum is a member of the signal set set. It returns 1 if the signal is in the set, 0 if not, and -1 if there is an error. The following errno error condition is defined for this function:

EINVAL The signum argument doesn’t specify a valid signal.

21.7.3 Process Signal Mask

The collection of signals that are currently blocked is called the signal mask. Each process has its own signal mask. When you create a new process (see Section 23.4 [Creating a Process], page 469), it inherits its parent’s mask.

You can block or unblock signals with total flexibility by modifying the signal mask.

The prototype for the sigprocmask function is in ‘signal.h’.

Function int sigprocmask (int how, const sigset_t *set,

      sigset_t *oldset)

The sigprocmask function is used to examine or change the calling process’s signal mask. The how argument determines how the signal mask is changed, and must be one of the following values:

SIG_BLOCK

Block the signals in set—add them to the existing mask. In other words, the new mask is the union of the existing mask and set.

SIG_UNBLOCK

Unblock the signals in set—remove them from the existing mask.

SIG_SETMASK

Use set for the mask; ignore the previous value of the mask.

The last argument, oldset, is used to return information about the old process signal mask. If you just want to change the mask without looking at it, pass a null pointer as the oldset argument. Similarly, if you want to know what’s in the mask without changing it, pass a null pointer for set (in this case the how argument is not significant). The oldset argument is often used to remember the previous signal mask in order to restore it later. (Since the signal mask is inherited over fork and exec calls, you can’t predict what its contents are when your program starts running.) If invoking sigprocmask causes any pending signals to be unblocked, at least one of those signals is delivered to the process before sigprocmask returns. The order in which pending signals are delivered is not specified, but you can control the order explicitly by making multiple sigprocmask calls to unblock various signals one at a time. The sigprocmask function returns 0 if successful, and -1 to indicate an error. The following errno error conditions are defined for this function:

EINVAL The how argument is invalid.

You can’t block the SIGKILL and SIGSTOP signals, but if the signal set includes these, sigprocmask just ignores them instead of returning an error status.

Remember, too, that blocking program error signals such as SIGFPE leads to undesirable results for signals generated by an actual program error (as opposed to signals sent with raise or kill). This is because your program may be too broken to be able to continue executing to a point where the signal is unblocked again. See Section 21.2.1 [Program Error Signals], page 396.

21.7.4 Blocking to Test for Delivery of a Signal

Now for a simple example. Suppose you establish a handler for SIGALRM signals that sets a flag whenever a signal arrives, and your main program checks this flag from time to time and then resets it. You can prevent additional SIGALRM signals from arriving in the meantime by wrapping the critical part of the code with calls to sigprocmask, like this:

/* This variable is set by the SIGALRM signal handler. */

  volatile sig_atomic_t flag = 0;
  int
  main (void)
  {
    sigset_t block_alarm;

…

/* Initialize the signal mask. */

    sigemptyset (&block_alarm);
    sigaddset (&block_alarm, SIGALRM);
    while (1)
      {

/* Check if a signal has arrived; if so, reset the flag. */

        sigprocmask (SIG_BLOCK, &block_alarm, NULL);
        if (flag)
          {

actions-if-not-arrived

            flag = 0;
          }
        sigprocmask (SIG_UNBLOCK, &block_alarm, NULL);

…

      }
  }

21.7.5 Blocking Signals for a Handler

When a signal handler is invoked, you usually want it to be able to finish without being interrupted by another signal. From the moment the handler starts until the moment it finishes, you must block signals that might confuse it or corrupt its data.

When a handler function is invoked on a signal, that signal is automatically blocked (in addition to any other signals that are already in the process’s signal mask) during the time the handler is running. If you set up a handler for SIGTSTP, for instance, then the arrival of that signal forces further SIGTSTP signals to wait during the execution of the handler.

However, by default, other kinds of signals are not blocked; they can arrive during handler execution.

The reliable way to block other kinds of signals during the execution of the handler is to use the sa_mask member of the sigaction structure.

Here is an example:

     #include <signal.h>
     #include <stddef.h>
     void catch_stop ();
     void
     install_handler (void)
     {
       struct sigaction setup_action;
       sigset_t block_mask;
       sigemptyset (&block_mask);

/* Block other terminal-generated signals while handler runs. */

       sigaddset (&block_mask, SIGINT);
       sigaddset (&block_mask, SIGQUIT);
       setup_action.sa_handler = catch_stop;
       setup_action.sa_mask = block_mask;
       setup_action.sa_flags = 0;
       sigaction (SIGTSTP, &setup_action, NULL);
     }

This is more reliable than blocking the other signals explicitly in the code for the handler. If you block signals explicitly in the handler, you can’t avoid at least a short interval at the beginning of the handler where they are not yet blocked.

You cannot remove signals from the process’s current mask using this mechanism. However, you can make calls to sigprocmask within your handler to block or unblock signals as you wish.

In any case, when the handler returns, the system restores the mask that was in place before the handler was entered. If any signals that become unblocked by this restoration are pending, the process will receive those signals immediately, before returning to the code that was interrupted.

21.7.6 Checking for Pending Signals

You can find out which signals are pending at any time by calling sigpending. This function is declared in ‘signal.h’.

Function int sigpending (sigset_t *set)

The sigpending function stores information about pending signals in set. If there is a pending signal that is blocked from delivery, then that signal is a member of the returned set. (You can test whether a particular signal is a member of this set using sigismember; see Section 21.7.2 [Signal Sets], page 432.) The return value is 0 if successful, and -1 on failure.

Testing whether a signal is pending is not often useful. Testing when that signal is not blocked is almost certainly bad design.

Here is an example.

     #include <signal.h>
     #include <stddef.h>
     sigset_t base_mask, waiting_mask;
     sigemptyset (&base_mask);
     sigaddset (&base_mask, SIGINT);
     sigaddset (&base_mask, SIGTSTP);

/* Block user interrupts while doing other processing. */

     sigprocmask (SIG_SETMASK, &base_mask, NULL);

…

/* After a while, check to see whether any signals are pending. */

     sigpending (&waiting_mask);
     if (sigismember (&waiting_mask, SIGINT)) {

/* User has tried to kill the process. */

     }
     else if (sigismember (&waiting_mask, SIGTSTP)) {

/* User has tried to stop the process. */

     }

Remember that if there is a particular signal pending for your process, additional signals of that same type that arrive in the meantime might be discarded. For example, if a SIGINT signal is pending when another SIGINT signal arrives, your program will probably only see one of them when you unblock this signal.

Portability Note: The sigpending function is new in POSIX.1. Older systems have no equivalent facility.

21.7.7 Remembering a Signal to Act On Later

Instead of blocking a signal using the library facilities, you can get almost the same results by making the handler set a flag to be tested later, when you “unblock”. Here is an example:

/* If this flag is nonzero, don’t handle the signal right away. */

volatile sig_atomic_t signal_pending;

/* This is nonzero if a signal arrived and was not handled. */

volatile sig_atomic_t defer_signal;
void
handler (int signum)
{
  if (defer_signal)
    signal_pending = signum;
  else

… /* “Really” handle the signal. */

}

…

void
update_mumble (int frob)
{

/* Prevent signals from having immediate effect. */

  defer_signal++;

/* Now update mumble, without worrying about interruption. */

  mumble.a = 1;
  mumble.b = hack ();
  mumble.c = frob;

/* We have updated mumble. Handle any signal that came in. */

  defer_signal--;
  if (defer_signal == 0 && signal_pending != 0)
    raise (signal_pending);
}

Note how the particular signal that arrives is stored in signal_pending. That way, we can handle several types of inconvenient signals with the same mechanism.

We increment and decrement defer_signal so that nested critical sections will work properly; thus, if update_mumble were called with signal_ pending already nonzero, signals would be deferred not only within update_ mumble, but also within the caller. This is also why we do not check signal_

pending if defer_signal is still nonzero.

The incrementing and decrementing of defer_signal require more than one instruction; it is possible for a signal to happen in the middle. But that does not cause any problem. If the signal happens early enough to see the value from before the increment or decrement, that is equivalent to a signal which came before the beginning of the increment or decrement, which is a case that works properly.

It is absolutely vital to decrement defer_signal before testing signal_ pending, because this avoids a subtle bug. If we did these things in the other order, like this,

       if (defer_signal == 1 && signal_pending != 0)
         raise (signal_pending);
       defer_signal--;

then a signal arriving in between the if statement and the decrement would be effectively “lost” for an indefinite amount of time. The handler would merely set defer_signal, but the program having already tested this variable, it would not test the variable again.

Bugs like these are called timing errors. They are especially bad because they happen only rarely and are nearly impossible to reproduce. You can’t expect to find them with a debugger as you would find a reproducible bug. So it is worth being especially careful to avoid them.

(You would not be tempted to write the code in this order, given the use of defer_signal as a counter which must be tested along with signal_ pending. After all, testing for zero is cleaner than testing for one. But if you did not use defer_signal as a counter, and gave it values of zero and one only, then either order might seem equally simple. This is a further advantage of using a counter for defer_signal: it will reduce the chance you will write the code in the wrong order and create a subtle bug.)

21.8 Waiting for a Signal

If your program is driven by external events, or uses signals for synchronization, then when it has nothing to do it should probably wait until a signal arrives.

21.8.1 Using pause

The simple way to wait until a signal arrives is to call pause. Please read about its disadvantages, in the following section, before you use it.

Function int pause ()

The pause function suspends program execution until a signal arrives whose action is either to execute a handler function, or to terminate the process. If the signal causes a handler function to be executed, then pause returns. This is considered an unsuccessful return (since “successful” behavior would be to suspend the program forever), so the return value is -1. Even if you specify that other primitives should resume when a system handler returns (see Section 21.5 [Primitives Interrupted by Signals], page 425), this has no effect on pause; it always fails when a signal is handled. The following errno error conditions are defined for this function:

EINTR The function was interrupted by delivery of a signal.

If the signal causes program termination, pause doesn’t return (obviously). The pause function is declared in ‘unistd.h’.

21.8.2 Problems with pause

The simplicity of pause can conceal serious timing errors that can make a program hang mysteriously.

It is safe to use pause if the real work of your program is done by the signal handlers themselves, and the “main program” does nothing but call pause. Each time a signal is delivered, the handler will do the next batch of work that is to be done, and then return, so that the main loop of the program can call pause again.

You can’t safely use pause to wait until one more signal arrives, and then resume real work. Even if you arrange for the signal handler to cooperate by setting a flag, you still can’t use pause reliably. Here is an example of this problem:

/* usr_interrupt is set by the signal handler. */

  if (!usr_interrupt)
    pause ();

/* Do work once the signal arrives. */

…

This has a bug: the signal could arrive after the variable usr_interrupt is checked, but before the call to pause. If no further signals arrive, the process would never wake up again.

You can put an upper limit on the excess waiting by using sleep in a loop, instead of using pause. (See Section 17.4 [Sleeping], page 359, for more about sleep.) Here is what this looks like:

/* usr_interrupt is set by the signal handler.

  while (!usr_interrupt)
    sleep (1);

/* Do work once the signal arrives. */

…

For some purposes, that is good enough. But with a little more complexity, you can wait reliably until a particular signal handler is run, using

sigsuspend.

21.8.3 Using sigsuspend

The clean and reliable way to wait for a signal to arrive is to block it and then use sigsuspend. By using sigsuspend in a loop, you can wait for certain kinds of signals, while letting other kinds of signals be handled by their handlers.

Function int sigsuspend (const sigset_t *set)

This function replaces the process’s signal mask with set and then suspends the process until a signal is delivered whose action is either to terminate the process or invoke a signal handling function. In other words, the program is effectively suspended until one of the signals that is not a member of set arrives. If the process is woken up by deliver of a signal that invokes a handler function, and the handler function returns, then sigsuspend also returns. The mask remains set only as long as sigsuspend is waiting. The function sigsuspend always restores the previous signal mask when it returns. The return value and error conditions are the same as for pause.

With sigsuspend, you can replace the pause or sleep loop in the previous section with something completely reliable:

  sigset_t mask, oldmask;

…

/* Set up the mask of signals to temporarily block. */

  sigemptyset (&mask);
  sigaddset (&mask, SIGUSR1);

…

/* Wait for a signal to arrive. */

     sigprocmask (SIG_BLOCK, &mask, &oldmask);
     while (!usr_interrupt)
       sigsuspend (&oldmask);
     sigprocmask (SIG_UNBLOCK, &mask, NULL);

This last piece of code is a little tricky. The key point to remember here is that when sigsuspend returns, it resets the process’s signal mask to the original value, the value from before the call to sigsuspend—in this case, the SIGUSR1 signal is once again blocked. The second call to sigprocmask is necessary to explicitly unblock this signal.

One other point: you may be wondering why the while loop is necessary at all, since the program is apparently only waiting for one SIGUSR1 signal. The answer is that the mask passed to sigsuspend permits the process to be woken up by the delivery of other kinds of signals, as well—for example, job control signals. If the process is woken up by a signal that doesn’t set usr_interrupt, it just suspends itself again until the “right” kind of signal eventually arrives.

This technique takes a few more lines of preparation, but that is needed just once for each kind of wait criterion you want to use. The code that actually waits is just four lines.

21.9 Using a Separate Signal Stack

A signal stack is a special area of memory to be used as the execution stack during signal handlers. It should be fairly large, to avoid any danger that it will overflow in turn; the macro SIGSTKSZ is defined to a canonical size for signal stacks. You can use malloc to allocate the space for the stack. Then call sigaltstack or sigstack to tell the system to use that space for the signal stack.

You don’t need to write signal handlers differently in order to use a signal stack. Switching from one stack to the other happens automatically. (Some non-GNU debuggers on some machines may get confused if you examine a stack trace while a handler that uses the signal stack is running.)

There are two interfaces for telling the system to use a separate signal stack. sigstack is the older interface, which comes from 4.2 BSD. sigaltstack is the newer interface, and comes from 4.4 BSD. The sigaltstack interface has the advantage that it does not require your program to know which direction the stack grows, which depends on the specific machine and operating system.

Data Type struct sigaltstack

This structure describes a signal stack. It contains the following members:

  void *ss_sp

This points to the base of the signal stack.

size_t ss_size

This is the size (in bytes) of the signal stack which ‘ss_sp’ points to. You should set this to however much space you allocated for the stack. There are two macros defined in ‘signal.h’ that you should use in calculating this size:

SIGSTKSZ This is the canonical size for a signal stack. It is judged to be sufficient for normal uses.

MINSIGSTKSZ

This is the amount of signal stack space the operating system needs just to implement signal delivery. The size of a signal stack must be greater than this. For most cases, just using SIGSTKSZ for ss_ size is sufficient. But if you know how much stack space your program’s signal handlers will need, you may want to use a different size. In this case, you should allocate MINSIGSTKSZ additional bytes for the signal stack and increase ss_size accordingly.

int ss_flags

This field contains the bitwise or of these flags:

SA_DISABLE

This tells the system that it should not use the signal stack.

SA_ONSTACK

This is set by the system, and indicates that the signal stack is currently in use. If this bit is not set, then signals will be delivered on the normal user stack.

Function int sigaltstack (const struct sigaltstack *stack,
struct sigaltstack *oldstack)

The sigaltstack function specifies an alternate stack for use during signal handling. When a signal is received by the process and its action indicates that the signal stack is used, the system arranges a switch to the currently installed signal stack while the handler for that signal is executed. If oldstack is not a null pointer, information about the currently installed signal stack is returned in the location it points to. If stack is not a null pointer, then this is installed as the new stack for use by signal handlers.

The return value is 0 on success and -1 on failure. If sigaltstack fails, it sets errno to one of these values:

EINVAL You tried to disable a stack that was in fact currently in use.

ENOMEM The size of the alternate stack was too small. It must be greater than MINSIGSTKSZ.

Here is the older sigstack interface. You should use sigaltstack instead on systems that have it.

Data Type struct sigstack

This structure describes a signal stack. It contains the following members:

void *ss_sp

This is the stack pointer. If the stack grows downwards on your machine, this should point to the top of the area you allocated. If the stack grows upwards, it should point to the bottom.

int ss_onstack

This field is true if the process is currently using this stack.

Function int sigstack (const struct sigstack *stack, struct
      sigstack *oldstack)

The sigstack function specifies an alternate stack for use during signal handling. When a signal is received by the process and its action indicates that the signal stack is used, the system arranges a switch to the currently installed signal stack while the handler for that signal is executed. If oldstack is not a null pointer, information about the currently installed signal stack is returned in the location it points to. If stack is not a null pointer, then this is installed as the new stack for use by signal handlers. The return value is 0 on success and -1 on failure.

21.10 BSD Signal Handling

This section describes alternative signal handling functions derived from BSD Unix. These facilities were an advance, in their time; today, they are mostly obsolete, and supported mainly for compatibility with BSD Unix.

There are many similarities between the BSD and POSIX signal handling facilities, because the POSIX facilities were inspired by the BSD facilities.

Besides having different names for all the functions to avoid conflicts, the main differences between the two are:

• BSD Unix represents signal masks as an int bit mask, rather than as a

 sigset_t object.

• The BSD facilities use a different default for whether an interrupted primitive should fail or resume. The POSIX facilities make system calls fail unless you specify that they should resume. With the BSD facility, the default is to make system calls resume unless you say they should fail. See Section 21.5 [Primitives Interrupted by Signals], page 425.

The BSD facilities are declared in ‘signal.h’.

21.10.1 BSD Function to Establish a Handler

Data Type struct sigvec

This data type is the BSD equivalent of struct sigaction (see Section 21.3.2 [Advanced Signal Handling], page 408); it is used to specify signal actions to the sigvec function. It contains the following members:

sighandler_t sv_handler

This is the handler function.

int sv_mask

This is the mask of additional signals to be blocked while the handler function is being called.

int sv_flags

This is a bit mask used to specify various flags which affect the behavior of the signal. You can also refer to this field

          as sv_onstack.

These symbolic constants can be used to provide values for the sv_flags field of a sigvec structure. This field is a bit mask value, so you bitwise-OR the flags of interest to you together.

Macro int SV ONSTACK

If this bit is set in the sv_flags field of a sigvec structure, it means to use the signal stack when delivering the signal.

Macro int SV INTERRUPT

If this bit is set in the sv_flags field of a sigvec structure, it means that system calls interrupted by this kind of signal should not be restarted if the handler returns; instead, the system calls should return with a EINTR error status. See Section 21.5 [Primitives Interrupted by Signals], page 425.

Macro int SV RESETHAND

If this bit is set in the sv_flags field of a sigvec structure, it means to reset the action for the signal back to SIG_DFL when the signal is received.

Function int sigvec (int signum, const struct sigvec

*action,struct sigvec *old-action) This function is the equivalent of sigaction (see Section 21.3.2 [Advanced Signal Handling], page 408); it installs the action action for the signal signum, returning information about the previous action in effect for that signal in old-action.

Function int siginterrupt (int signum, int failflag)

This function specifies which approach to use when certain primitives are interrupted by handling signal signum. If failflag is false, signal signum restarts primitives. If failflag is true, handling signum causes these primitives to fail with error code EINTR. See Section 21.5 [Primitives Interrupted by Signals], page 425.

21.10.2 BSD Functions for Blocking Signals

Macro int sigmask (int signum)

This macro returns a signal mask that has the bit for signal signum set. You can bitwise-OR the results of several calls to sigmask together to specify more than one signal. For example,

     (sigmask (SIGTSTP) | sigmask (SIGSTOP)
      | sigmask (SIGTTIN) | sigmask (SIGTTOU))

specifies a mask that includes all the job-control stop signals.

Function int sigblock (int mask)

This function is equivalent to sigprocmask (see Section 21.7.3 [Process Signal Mask], page 433) with a how argument of SIG_BLOCK: it adds the signals specified by mask to the calling process’s set of blocked signals. The return value is the previous set of blocked signals.

Function int sigsetmask (int mask)

This function equivalent to sigprocmask (see Section 21.7.3 [Process Signal Mask], page 433) with a how argument of SIG_SETMASK: it sets the calling process’s signal mask to mask. The return value is the previous set of blocked signals.

Function int sigpause (int mask)

This function is the equivalent of sigsuspend (see Section 21.8 [Waiting for a Signal], page 439): it sets the calling process’s signal mask to mask, and waits for a signal to arrive. On return the previous set of blocked signals is restored.

22 Process Startup and Termination

Processes are the primitive units for allocation of system resources. Each process has its own address space and (usually) one thread of control. A process executes a program; you can have multiple processes executing the same program, but each process has its own copy of the program within its own address space and executes it independently of the other copies.

This chapter explains what your program should do to handle the startup of a process, to terminate its process, and to receive information (arguments and the environment) from the parent process.

22.1 Program Arguments

The system starts a C program by calling the function main. It is up to you to write a function named main—otherwise, you won’t even be able to link your program without errors.

In ISO C you can define main either to take no arguments, or to take two arguments that represent the command line arguments to the program, like this:

     int main (int argc, char *argv[])

The command line arguments are the whitespace-separated tokens given in the shell command used to invoke the program; thus, in ‘cat foo bar’, the arguments are ‘foo’ and ‘bar’. The only way a program can look at its command line arguments is via the arguments of main. If main doesn’t take arguments, then you cannot get at the command line.

The value of the argc argument is the number of command line arguments. The argv argument is a vector of C strings; its elements are the individual command line argument strings. The file name of the program being run is also included in the vector as the first element; the value of argc counts this element. A null pointer always follows the last element: argv[argc] is this null pointer.

For the command ‘cat foo bar’, argc is 3 and argv has three elements,

"cat", "foo" and "bar".

If the syntax for the command line arguments to your program is simple enough, you can simply pick the arguments off from argv by hand. But unless your program takes a fixed number of arguments, or all of the arguments are interpreted in the same way (as file names, for example), you are usually better off using getopt to do the parsing.

In Unix systems you can define main a third way, using three arguments:

     int main (int argc, char *argv[], char *envp)

The first two arguments are just the same. The third argument envp gives the process’s environment; it is the same as the value of environ. See Section 22.2 [Environment Variables], page 458. POSIX.1 does not allow this three-argument form, so to be portable it is best to write main to take two arguments, and use the value of environ.

22.1.1 Program Argument Syntax Conventions

POSIX recommends these conventions for command line arguments. getopt (see Section 22.1.2 [Parsing Program Options], page 449) makes it easy to implement them.

• Arguments are options if they begin with a hyphen delimiter (‘-’).

• Multiple options may follow a hyphen delimiter in a single token if the options do not take arguments. Thus, ‘-abc’ is equivalent to ‘-a -b -c’.

• Option names are single alphanumeric characters (as for isalnum; see Section 4.1 [Classification of Characters], page 61).

• Certain options require an argument. For example, the ‘-o’ command of the ld command requires an argument—an output file name.

• An option and its argument may or may not appear as separate tokens. (In other words, the whitespace separating them is optional.) Thus, ‘-o foo’ and ‘-ofoo’ are equivalent.

• Options typically precede other non-option arguments.

The implementation of getopt in the GNU C library normally makes it appear as if all the option arguments were specified before all the non-option arguments for the purposes of parsing, even if the user of your program intermixed option and non-option arguments. It does this by reordering the elements of the argv array. This behavior is nonstandard; if you want to suppress it, define the _POSIX_OPTION_ ORDER environment variable. See Section 22.2.2 [Standard Environment Variables], page 460.

• The argument ‘--’ terminates all options; any following arguments are treated as non-option arguments, even if they begin with a hyphen.

• A token consisting of a single hyphen character is interpreted as an ordinary non-option argument. By convention, it is used to specify input from or output to the standard input and output streams.

• Options may be supplied in any order, or appear multiple times. The interpretation is left up to the particular application program. GNU adds long options to these conventions. Long options consist of ‘--’

followed by a name made of alphanumeric characters and dashes. Option names are typically one to three words long, with hyphens to separate words. Users can abbreviate the option names as long as the abbreviations are unique.

To specify an argument for a long option, write ‘--name=value’. This syntax enables a long option to accept an argument that is itself optional.

Eventually, the GNU system will provide completion for long option names in the shell.

22.1.2 Parsing Program Options

Here are the details about how to call the getopt function. To use this facility, your program must include the header file ‘unistd.h’.

Variable int opterr

If the value of this variable is nonzero, then getopt prints an error message to the standard error stream if it encounters an unknown option character or an option with a missing required argument. This is the default behavior. If you set this variable to zero, getopt does not print any messages, but it still returns the character ? to indicate an error.

Variable int optopt

When getopt encounters an unknown option character or an option with a missing required argument, it stores that option character in this variable. You can use this for providing your own diagnostic messages.

Variable int optind

This variable is set by getopt to the index of the next element of the argv array to be processed. Once getopt has found all of the option arguments, you can use this variable to determine where the remaining non-option arguments begin. The initial value of this variable is 1.

Variable char * optarg

This variable is set by getopt to point at the value of the option argument, for those options that accept arguments.

Function int getopt (int argc, char **argv, const char

*options) The getopt function gets the next option argument from the argument list specified by the argv and argc arguments. Normally these values come directly from the arguments received by main. The options argument is a string that specifies the option characters that are valid for this program. An option character in this string can be followed by a colon (‘:’) to indicate that it takes a required argument. If the options argument string begins with a hyphen (‘-’), this is treated specially. It permits arguments that are not options to be returned as if they were associated with option character ‘\0’. The getopt function returns the option character for the next command line option. When no more option arguments are available, it returns -1. There may still be more non-option arguments; you must compare the external variable optind against the argc parameter to check this.

If the option has an argument, getopt returns the argument by storing it in the variable optarg. You don’t ordinarily need to copy the optarg string, since it is a pointer into the original argv array, not into a static area that might be overwritten.

If getopt finds an option character in argv that was not included in options, or a missing option argument, it returns ‘?’ and sets the external variable optopt to the actual option character. If the first character of options is a colon (‘:’), then getopt returns ‘:’ instead of ‘?’ to indicate a missing option argument. In addition, if the external variable opterr is nonzero (which is the default), getopt prints an error message.

22.1.3 Example of Parsing Arguments with getopt

Here is an example showing how getopt is typically used. The key points to notice are:

• Normally, getopt is called in a loop. When getopt returns -1, indicating no more options are present, the loop terminates.

• A switch statement is used to dispatch on the return value from getopt. In typical use, each case just sets a variable that is used later in the program.

• A second loop is used to process the remaining non-option arguments.

  #include <unistd.h>
  #include <stdio.h>
  int
  main (int argc, char **argv)
  {
    int aflag = 0;
    int bflag = 0;
    char *cvalue = NULL;
    int index;
    int c;
    opterr = 0;

Chapter 22: Process Startup and Termination 451

while ((c = getopt (argc, argv, "abc:")) != -1)
switch (c)
{
case 'a':
aflag = 1;
break;
case 'b':
bflag = 1;
break;
case 'c':
cvalue = optarg;
break;
case '?':
if (isprint (optopt))
  fprintf (stderr, "Unknown option `-%c'.\n", optopt);
else
  fprintf (stderr,
           "Unknown option character `\\x%x'.\n",
           optopt);
return 1;
default:
abort ();
}
printf ("aflag = %d, bflag = %d, cvalue = %s\n", aflag, bflag, cvalue);
for (index = optind; index < argc; index++)
printf ("Non-option argument %s\n", argv[index]);
return 0;
}

Here are some examples showing what this program prints with different combinations of arguments:

% testopt
aflag = 0, bflag = 0, cvalue = (null)
% testopt -a -b
aflag = 1, bflag = 1, cvalue = (null)
% testopt -ab
aflag = 1, bflag = 1, cvalue = (null)
% testopt -c foo
aflag = 0, bflag = 0, cvalue = foo
% testopt -cfoo

  aflag = 0, bflag = 0, cvalue = foo
  % testopt arg1
  aflag = 0, bflag = 0, cvalue = (null)
  Non-option argument arg1
  % testopt -a arg1
  aflag = 1, bflag = 0, cvalue = (null)
  Non-option argument arg1
  % testopt -c foo arg1
  aflag = 0, bflag = 0, cvalue = foo
  Non-option argument arg1
  % testopt -a -- -b
  aflag = 1, bflag = 0, cvalue = (null)
  Non-option argument -b
  % testopt -a -
  aflag = 1, bflag = 0, cvalue = (null)
  Non-option argument -

22.1.4 Parsing Long Options

To accept GNU-style long options as well as single-character options, use getopt_long instead of getopt. This function is declared in ‘getopt.h’, not ‘unistd.h’. You should make every program accept long options if it uses any options, for this takes little extra work and helps beginners remember how to use the program.

Data Type struct option

This structure describes a single long option name for the sake of getopt_long. The argument longopts must be an array of these structures, one for each long option. Terminate the array with an element containing all zeros. The struct option structure has these fields:

const char *name

This field is the name of the option. It is a string.

int has_arg

This field says whether the option takes an argument. It is an integer, and there are three legitimate val-

          ues: no_argument, required_argument and optional_
          argument.

int *flag

int val These fields control how to report or act on the option when it occurs. If flag is a null pointer, then the val is a value which identifies this option. Often these values are chosen to uniquely identify particular long options. If flag is not a null pointer, it should be the address of an int variable which is the flag for this option. The value in val is the value to store in the flag to indicate that the option was seen.

Function int getopt long (int argc, char **argv, const char

*shortopts, struct option *longopts, int *indexptr) Decode options from the vector argv (whose length is argc). The argument shortopts describes the short options to accept, just as it does in getopt. The argument longopts describes the long options to accept (see above). When getopt_long encounters a short option, it does the same thing that getopt would do: it returns the character code for the option, and stores the options argument (if it has one) in optarg. When getopt_long encounters a long option, it takes actions based on the flag and val fields of the definition of that option. If flag is a null pointer, then getopt_long returns the contents of val to indicate which option it found. You should arrange distinct values in the val field for options with different meanings, so you can decode these values after getopt_long returns. If the long option is equivalent to a short option, you can use the short option’s character code in val. If flag is not a null pointer, that means this option should just set a flag in the program. The flag is a variable of type int that you define. Put the address of the flag in the flag field. Put in the val field the value you would like this option to store in the flag. In this case, getopt_long returns 0. For any long option, getopt_long tells you the index in the array longopts of the options definition, by storing it into *indexptr. You can get the name of the option with longopts[*indexptr].name. So you can distinguish among long options either by the values in their val fields or by their indices. You can also distinguish in this way among long options that set flags. When a long option has an argument, getopt_long puts the argument value in the variable optarg before returning. When the option has no argument, the value in optarg is a null pointer. This is how you can tell whether an optional argument was supplied.

When getopt_long has no more options to handle, it returns -1, and leaves in the variable optind the index in argv of the next remaining argument.

22.1.5 Example of Parsing Long Options

#include <stdio.h>
#include <stdlib.h>
#include <getopt.h>
/* Flag set by ‘--verbose’. */
static int verbose_flag;
int
main (argc, argv)
int argc;
char **argv;
{
int c;
while (1)
{
static struct option long_options[] =
{

/* These options set a flag. */

{"verbose", 0, &verbose_flag, 1},
{"brief", 0, &verbose_flag, 0},

/* These options don’t set a flag. We distinguish them by their indices. */

{"add", 1, 0, 0},
{"append", 0, 0, 0},
{"delete", 1, 0, 0},
{"create", 0, 0, 0},
{"file", 1, 0, 0},
{0, 0, 0, 0}
};

/* getopt_long stores the option index here. */

int option_index = 0;
c = getopt_long (argc, argv, "abc:d:",
             long_options, &option_index);

/* Detect the end of the options. */

if (c == -1)
break;

switch (c)
{
case 0:

/* If this option set a flag, do nothing else now. */

if (long_options[option_index].flag != 0)
  break;
printf ("option %s", long_options[option_index].name);
if (optarg)
  printf (" with arg %s", optarg);
printf ("\n");
break;
case 'a':
puts ("option -a\n");
break;
case 'b':
puts ("option -b\n");
break;
case 'c':
printf ("option -c with value `%s'\n", optarg);
break;
case 'd':
printf ("option -d with value `%s'\n", optarg);
break;
case '?':

/* getopt_long already printed an error message. */

break;
default:
abort ();
}
}

/* Instead of reporting ‘--verbose’ and ‘--brief’ as they are encountered, we report the final status resulting from them. */

if (verbose_flag)
puts ("verbose flag is set");

/* Print any remaining command line arguments (not options). */

    if (optind < argc)
      {
        printf ("non-option ARGV-elements: ");
        while (optind < argc)
          printf ("%s ", argv[optind++]);
        putchar ('\n');
      }
    exit (0);
  }

22.1.6 Parsing of Suboptions

Having a single level of options is sometimes not enough. There might be too many options which have to be available or a set of options is closely related.

For this case some programs use suboptions. One of the most prominent programs is certainly mount(8). The -o option take one argument which itself is a comma separated list of options. To ease the programming of code like this the function getsubopt is available.

Function int getsubopt (char **optionp, const char* const

*tokens, char **valuep) The optionp parameter must be a pointer to a variable containing the address of the string to process. When the function returns the reference is updated to point to the next suboption or to the terminating ‘\0’ character if there is no more suboption available.

The tokens parameter references an array of strings containing the known suboptions. All strings must be ‘\0’ terminated and to mark the end a null pointer must be stored. When getsubopt finds a possible legal suboption it compares it with all strings available in the tokens array and returns the index in the string as the indicator.

In case the suboption has an associated value introduced by a ‘=’ character, a pointer to the value is returned in valuep. The string is ‘\0’ terminated. If no argument is available valuep is set to the null pointer. By doing this the caller can check whether a necessary value is given or whether no unexpected value is present.

In case the next suboption in the string is not mentioned in the tokens array the starting address of the suboption including a possible value is returned in valuep and the return value of the function is ‘-1’.

22.1.7 Parsing of Suboptions Example

The code which might appear in the mount(8) program is a perfect example of the use of getsubopt:

#include <stdio.h>
#include <stdlib.h>
int do_all;
const char *type;
int read_size;
int write_size;
int read_only;
enum
{
  RO_OPTION = 0,
  RW_OPTION,
  READ_SIZE_OPTION,
  WRITE_SIZE_OPTION
};
const char *mount_opts[] =
{
  [RO_OPTION] = "ro",
  [RW_OPTION] = "rw",
  [READ_SIZE_OPTION] = "rsize",
  [WRITE_SIZE_OPTION] = "wsize"
};
int
main (int argc, char *argv[])
{
  char *subopts, *value;
  int opt;
  while ((opt = getopt (argc, argv, "at:o:")) != -1)
    switch (opt)
      {
      case 'a':
        do_all = 1;
        break;
      case 't':
        type = optarg;
        break;

case 'o':
subopts = optarg;
while (*subopts != '\0')
switch (getsubopt (&subopts, mount_opts, &value))
{
case RO_OPTION:
read_only = 1;
break;
case RW_OPTION:
read_only = 0;
break;
case READ_SIZE_OPTION:
if (value == NULL)
  abort ();
read_size = atoi (value);
break;
case WRITE_SIZE_OPTION:
if (value == NULL)
  abort ();
write_size = atoi (value);
break;
default:

/* Unknown suboption. */

printf ("Unknown suboption `%s'\n", value);
break;
}
break;
default:
abort ();
}

/* Do the real work. */

return 0;
}

22.2 Environment Variables

When a program is executed, it receives information about the context in which it was invoked in two ways. The first mechanism uses the argv and argc arguments to its main function, and is discussed in Section 22.1 [Program Arguments], page 447. The second mechanism uses environment variables and is discussed in this section.

The argv mechanism is typically used to pass command-line arguments specific to the particular program being invoked. The environment, on the other hand, keeps track of information that is shared by many programs, changes infrequently, and that is less frequently used.

The environment variables discussed in this section are the same environment variables that you set using assignments and the export command in the shell. Programs executed from the shell inherit all of the environment variables from the shell.

Standard environment variables are used for information about the user’s home directory, terminal type, current locale, and so on; you can define additional variables for other purposes. The set of all environment variables that have values is collectively known as the environment.

Names of environment variables are case-sensitive and must not contain the character ‘=’. System-defined environment variables are invariably uppercase.

The values of environment variables can be anything that can be represented as a string. A value must not contain an embedded null character, since this is assumed to terminate the string.

22.2.1 Environment Access

The value of an environment variable can be accessed with the getenv function. This is declared in the header file ‘stdlib.h’.

Function char * getenv (const char *name)

This function returns a string that is the value of the environment variable name. You must not modify this string. In some non-Unix systems not using the GNU library, it might be overwritten by subsequent calls to getenv (but not by any other library function). If the environment variable name is not defined, the value is a null pointer.

Function int putenv (const char *string)

The putenv function adds or removes definitions from the environment. If the string is of the form ‘name=value’, the definition is added to the environment. Otherwise, the string is interpreted as the name of an environment variable, and any definition for this variable in the environment is removed. The GNU library provides this function for compatibility with SVID; it may not be available in other systems.

You can deal directly with the underlying representation of environment objects to add more variables to the environment (for example, to communicate with another program you are about to execute; see Section 23.5 [Executing a File], page 470).

Variable char ** environ

The environment is represented as an array of strings. Each string is of the format ‘name=value’. The order in which strings appear in the environment is not significant, but the same name must not appear more than once. The last element of the array is a null pointer. This variable is declared in the header file ‘unistd.h’. If you just want to get the value of an environment variable, use

getenv.

Unix systems, and the GNU system, pass the initial value of environ as the third argument to main. See Section 22.1 [Program Arguments], page 447.

22.2.2 Standard Environment Variables

These environment variables have standard meanings. This doesn’t mean that they are always present in the environment; but if these variables are present, they have these meanings. You shouldn’t try to use these environment variable names for some other purpose.

HOME

This is a string representing the user’s home directory, or initial default working directory. The user can set HOME to any value. If you need to make sure to obtain the proper home directory for a particular user, you should not use HOME; instead, look up the user’s name in the user database (see Section 26.12 [User Database], page 521). For most purposes, it is better to use HOME, precisely because this lets the user specify the value.

LOGNAME

This is the name that the user used to log in. Since the value in the environment can be tweaked arbitrarily, this is not a reliable way to identify the user who is running a process; a function like getlogin (see Section 26.11 [Identifying Who Logged In], page 520) is better for that purpose. For most purposes, it is better to use LOGNAME, precisely because this lets the user specify the value.

PATH

A path is a sequence of directory names which is used for searching for a file. The variable PATH holds a path used for searching for programs to be run. The execlp and execvp functions (see Section 23.5 [Executing a File], page 470) use this environment variable, as do many shells and other utilities which are implemented in terms of those functions. The syntax of a path is a sequence of directory names separated by colons. An empty string instead of a directory name stands for the current directory (see Section 9.1 [Working Directory], page 181). A typical value for this environment variable might be a string like:

     :/bin:/etc:/usr/bin:/usr/new/X11:/usr/new:/usr/local/bin

This means that if the user tries to execute a program named foo, the system will look for files named ‘foo’, ‘/bin/foo’, ‘/etc/foo’, and so on. The first of these files that exists is the one that is executed.

TERM

This specifies the kind of terminal that is receiving program output. Some programs can make use of this information to take advantage of special escape sequences or terminal modes supported by particular kinds of terminals. Many programs which use the termcap library (see section “Finding a Terminal Description” in The Termcap Library Manual) use the TERM environment variable, for example.

TZ

This specifies the time zone. See Section 17.2.5 [Specifying the Time Zone with “codeTZ], page 352, for information about the format of this string and how it is used.

LANG

This specifies the default locale to use for attribute categories where neither LC_ALL nor the specific environment variable for that category is set. See Chapter 19 [Locales and Internationalization], page 379, for more information about locales.

LC_COLLATE

This specifies what locale to use for string sorting.

LC_CTYPE

This specifies what locale to use for character sets and character classification.

LC_MONETARY

This specifies what locale to use for formatting monetary values.

LC_NUMERIC

This specifies what locale to use for formatting numbers.

LC_TIME

This specifies what locale to use for formatting date/time values.

_POSIX_OPTION_ORDER

If this environment variable is defined, it suppresses the usual reordering of command line arguments by getopt. See Section 22.1.1 [Program Argument Syntax Conventions], page 448.

22.3 Program Termination

The usual way for a program to terminate is simply for its main function to return. The exit status value returned from the main function is used to report information back to the process’s parent process or shell.

A program can also terminate normally by calling the exit function. In addition, programs can be terminated by signals; this is discussed in

more detail in Chapter 21 [Signal Handling], page 393. The abort function causes a signal that kills the program.

22.3.1 Normal Termination

A process terminates normally when the program calls exit. Returning from main is equivalent to calling exit, and the value that main returns is used as the argument to exit.

Function void exit (int status)

The exit function terminates the process with status status. This function does not return.

Normal termination causes the following actions:

1. Functions that were registered with the atexit or on_exit functions are called in the reverse order of their registration. This mechanism allows your application to specify its own “cleanup” actions to be performed at program termination. Typically, this is used to do things like saving program state information in a file, or unlocking locks in shared data bases.

2. All open streams are closed, writing out any buffered output data. See Section 7.4 [Closing Streams], page 96. In addition, temporary files opened with the tmpfile function are removed; see Section 9.10 [Temporary Files], page 210.

3. _exit is called, terminating the program. See Section 22.3.5 [Termination Internals], page 465.

22.3.2 Exit Status

When a program exits, it can return to the parent process a small amount of information about the cause of termination, using the exit status. This is a value between 0 and 255 that the exiting process passes as an argument

to exit.

Normally you should use the exit status to report very broad information about success or failure. You can’t provide a lot of detail about the reasons for the failure, and most parent processes would not want much detail anyway.

There are conventions for what sorts of status values certain programs should return. The most common convention is simply 0 for success and 1 for failure. Programs that perform comparison use a different convention: they use status 1 to indicate a mismatch, and status 2 to indicate an inability to compare. Your program should follow an existing convention if an existing convention makes sense for it.

A general convention reserves status values 128 and up for special purposes. In particular, the value 128 is used to indicate failure to execute another program in a subprocess. This convention is not universally obeyed, but it is a good idea to follow it in your programs.

Warning: Don’t try to use the number of errors as the exit status. This is actually not very useful; a parent process would generally not care how many errors occurred. Worse than that, it does not work, because the status value is truncated to eight bits. Thus, if the program tried to report 256 errors, the parent would receive a report of 0 errors—that is, success.

For the same reason, it does not work to use the value of errno as the exit status—these can exceed 255.

Portability note: Some non-POSIX systems use different conventions for exit status values. For greater portability, you can use the macros EXIT_ SUCCESS and EXIT_FAILURE for the conventional status value for success and failure, respectively. They are declared in the file ‘stdlib.h’.

Macro int EXIT SUCCESS

This macro can be used with the exit function to indicate successful program completion. On POSIX systems, the value of this macro is 0. On other systems, the value might be some other (possibly non-constant) integer expression.

Macro int EXIT FAILURE

This macro can be used with the exit function to indicate unsuccessful program completion in a general sense. On POSIX systems, the value of this macro is 1. On other systems, the value might be some other (possibly non-constant) integer expression. Other nonzero status values also indicate failures. Certain programs use different nonzero status values to indicate particular kinds of "nonsuccess". For example, diff uses status value 1 to mean that the files are different, and 2 or more to mean that there was difficulty in opening the files.

22.3.3 Cleanups on Exit

Your program can arrange to run its own cleanup functions if normal termination happens. If you are writing a library for use in various application programs, then it is unreliable to insist that all applications call the library’s cleanup functions explicitly before exiting. It is much more robust to make the cleanup invisible to the application, by setting up a cleanup function in the library itself using atexit or on_exit.

Function int atexit (void (*function) (void))

The atexit function registers the function function to be called at normal program termination. The function is called with no arguments. The return value from atexit is zero on success and nonzero if the function cannot be registered.

Function int on exit (void (*function)(int status, void *arg),

      void *arg)

This function is a somewhat more powerful variant of atexit. It accepts two arguments, a function function and an arbitrary pointer arg. At normal program termination, the function is called with two arguments: the status value passed to exit, and the arg. This function is included in the GNU C library only for compatibility for SunOS, and may not be supported by other implementations.

Here’s a trivial program that illustrates the use of exit and atexit:

  #include <stdio.h>
  #include <stdlib.h>
  void
  bye (void)
  {
    puts ("Goodbye, cruel world....");
  }
  int
  main (void)
  {
    atexit (bye);
    exit (EXIT_SUCCESS);
  }

When this program is executed, it just prints the message and exits.

22.3.4 Aborting a Program

You can abort your program using the abort function. The prototype for this function is in ‘stdlib.h’.

Function void abort (void)

The abort function causes abnormal program termination. This does not execute cleanup functions registered with atexit or on_exit. This function actually terminates the process by raising a SIGABRT signal, and your program can include a handler to intercept this signal; see Chapter 21 [Signal Handling], page 393.

Future Change Warning: Proposed Federal censorship regulations may prohibit us from giving you information about the possibility of calling this function. We would be required to say that this is not an acceptable way of terminating a program.

22.3.5 Termination Internals

The _exit function is the primitive used for process termination by exit. It is declared in the header file ‘unistd.h’.

Function void exit (int status)

The _exit function is the primitive for causing a process to terminate with status status. Calling this function does not execute cleanup functions registered with atexit or on_exit.

When a process terminates for any reason—either by an explicit termination call, or termination as a result of a signal—the following things happen:

• All open file descriptors in the process are closed. See Chapter 8 [Low-Level Input/Output], page 151. Note that streams are not flushed automatically when the process terminates; See Chapter 7 [Input/Output on Streams], page 93.

• The low-order 8 bits of the return status code are saved to be reported back to the parent process via wait or waitpid; see Section 23.6 [Process Completion], page 473.

• Any child processes of the process being terminated are assigned a new parent process. (On most systems, including GNU, this is the init process, with process ID 1.)

• A SIGCHLD signal is sent to the parent process.

• If the process is a session leader that has a controlling terminal, then a SIGHUP signal is sent to each process in the foreground job, and the controlling terminal is disassociated from that session. See Chapter 24 [Job Control], page 479.

• If termination of a process causes a process group to become orphaned, and any member of that process group is stopped, then a SIGHUP signal and a SIGCONT signal are sent to each process in the group. See Chapter 24 [Job Control], page 479.

23 Processes

Processes are the primitive units for allocation of system resources. Each process has its own address space and (usually) one thread of control. A process executes a program; you can have multiple processes executing the same program, but each process has its own copy of the program within its own address space and executes it independently of the other copies.

Processes are organized hierarchically. Each process has a parent process which explicitly arranged to create it. The processes created by a given parent are called its child processes. A child inherits many of its attributes from the parent process.

This chapter describes how a program can create, terminate, and control child processes. Actually, there are three distinct operations involved: creating a new child process, causing the new process to execute a program, and coordinating the completion of the child process with the original program.

The system function provides a simple, portable mechanism for running another program; it does all three steps automatically. If you need more control over the details of how this is done, you can use the primitive functions to do each step individually instead.

23.1 Running a Command

The easy way to run another program is to use the system function. This function does all the work of running a subprogram, but it doesn’t give you much control over the details: you have to wait until the subprogram terminates before you can do anything else.

Function int system (const char *command)

This function executes command as a shell command. In the GNU C library, it always uses the default shell sh to run the command. In particular, it searches the directories in PATH to find programs to execute. The return value is -1 if it wasn’t possible to create the shell process, and otherwise is the status of the shell process. See Section 23.6 [Process Completion], page 473, for details on how this status code can be interpreted. The system function is declared in the header file ‘stdlib.h’.

Portability Note: Some C implementations may not have any notion of a command processor that can execute other programs. You can determine whether a command processor exists by executing system (NULL); if the return value is nonzero, a command processor is available.

The popen and pclose functions (see Section 10.2 [Pipe to a Subprocess], page 217) are closely related to the system function. They allow the parent process to communicate with the standard input and output channels of the command being executed.

23.2 Process Creation Concepts

This section gives an overview of processes and of the steps involved in creating a process and making it run another program.

Each process is named by a process ID number. A unique process ID is allocated to each process when it is created. The lifetime of a process ends when its termination is reported to its parent process; at that time, all of the process resources, including its process ID, are freed.

Processes are created with the fork system call (so the operation of creating a new process is sometimes called forking a process). The child process created by fork is a copy of the original parent process, except that it has its own process ID.

After forking a child process, both the parent and child processes continue to execute normally. If you want your program to wait for a child process to finish executing before continuing, you must do this explicitly after the fork operation, by calling wait or waitpid (see Section 23.6 [Process Completion], page 473). These functions give you limited information about why the child terminated—for example, its exit status code.

A newly forked child process continues to execute the same program as its parent process, at the point where the fork call returns. You can use the return value from fork to tell whether the program is running in the parent process or the child.

Having several processes run the same program is only occasionally useful. But the child can execute another program using one of the exec functions; see Section 23.5 [Executing a File], page 470. The program that the process is executing is called its process image. Starting execution of a new program causes the process to forget all about its previous process image; when the new program exits, the process exits too, instead of returning to the previous process image.

23.3 Process Identification

The pid_t data type represents process IDs. You can get the process ID of a process by calling getpid. The function getppid returns the process ID of the parent of the current process (this is also known as the parent process ID). Your program should include the header files ‘unistd.h’ and ‘sys/types.h’ to use these functions.

Data Type pid t

The pid_t data type is a signed integer type which is capable of representing a process ID. In the GNU library, this is an int.

Function pid_t getpid (void)

The getpid function returns the process ID of the current process.

Function pid_t getppid (void)

The getppid function returns the process ID of the parent of the current process.

23.4 Creating a Process

The fork function is the primitive for creating a process. It is declared in the header file ‘unistd.h’.

Function pid_t fork (void)

The fork function creates a new process. If the operation is successful, there are then both parent and child processes and both see fork return, but with different values: it returns a value of 0 in the child process and returns the child’s process ID in the parent process. If process creation failed, fork returns a value of -1 in the parent process. The following errno error conditions are defined for fork:

EAGAIN There aren’t enough system resources to create another process, or the user already has too many processes running. This means exceeding the RLIMIT_NPROC resource limit, which can usually be increased; see Section 17.6 [Limiting Resource Usage], page 362.

ENOMEM The process requires more space than the system can supply.

The specific attributes of the child process that differ from the parent process are:

• The child process has its own unique process ID.

• The parent process ID of the child process is the process ID of its parent process.

• The child process gets its own copies of the parent process’s open file descriptors. Subsequently changing attributes of the file descriptors in the parent process won’t affect the file descriptors in the child, and vice versa. See Section 8.7 [Control Operations on Files], page 165. However, the file position associated with each descriptor is shared by both processes; see Section 6.1.2 [File Position], page 87.

• The elapsed processor times for the child process are set to zero; see Section 17.1 [Processor Time], page 339.

• The child doesn’t inherit file locks set by the parent process. See Section 8.7 [Control Operations on Files], page 165.

• The child doesn’t inherit alarms set by the parent process. See Section 17.3 [Setting an Alarm], page 356.

• The set of pending signals (see Section 21.1.3 [How Signals Are Delivered], page 394) for the child process is cleared. (The child process inherits its mask of blocked signals and signal actions from the parent process.)

Function pid_t vfork (void)

The vfork function is similar to fork but on systems it is more efficient; however, there are restrictions you must follow to use it safely. While fork makes a complete copy of the calling process’s address space and allows both the parent and child to execute independently, vfork does not make this copy. Instead, the child process created with vfork shares its parent’s address space until it calls exits or one of the exec functions. In the meantime, the parent process suspends execution. You must be very careful not to allow the child process created with vfork to modify any global data or even local variables shared with the parent. Furthermore, the child process cannot return from (or do a long jump out of) the function that called vfork! This would leave the parent process’s control information very confused. If in doubt, use fork instead. Some operating systems don’t really implement vfork. The GNU C library permits you to use vfork on all systems, but actually executes fork if vfork isn’t available. If you follow the proper precautions for using vfork, your program will still work even if the system uses fork instead.

23.5 Executing a File

This section describes the exec family of functions, for executing a file as a process image. You can use these functions to make a child process execute a new program after it has been forked.

The functions in this family differ in how you specify the arguments, but otherwise they all do the same thing. They are declared in the header file

‘unistd.h’.

Function int execv (const char *filename, char *const argv[])

The execv function executes the file named by filename as a new process image. The argv argument is an array of null-terminated strings that is used to provide a value for the argv argument to the main function of the program to be executed. The last element of this array must be a null pointer. By convention, the first element of this array is the file name of the program sans directory names. See Section 22.1 [Program Arguments], page 447, for full details on how programs can access these arguments. The environment for the new process image is taken from the environ variable of the current process image; see Section 22.2 [Environment Variables], page 458, for information about environments.

Function int execl (const char *filename, const char *arg0,

…) This is similar to execv, but the argv strings are specified individually instead of as an array. A null pointer must be passed as the last such argument.

Function int execve (const char *filename, char *const argv[],

      char *const env[])

This is similar to execv, but permits you to specify the environment for the new program explicitly as the env argument. This should be an array of strings in the same format as for the environ variable; see Section 22.2.1 [Environment Access], page 459.

Function int execle (const char *filename, const char *arg0,

      char *const env[], . . .)

This is similar to execl, but permits you to specify the environment for the new program explicitly. The environment argument is passed following the null pointer that marks the last argv argument, and should be an array of strings in the same format as for the environ variable.

Function int execvp (const char *filename, char *const argv[])

The execvp function is similar to execv, except that it searches the directories listed in the PATH environment variable (see Section 22.2.2 [Standard Environment Variables], page 460) to find the full file name of a file from filename if filename does not contain a slash. This function is useful for executing system utility programs, because it looks for them in the places that the user has chosen. Shells use it to run the commands that users type.

Function int execlp (const char *filename, const char *arg0,

…) This function is like execl, except that it performs the same file name searching as the execvp function.

The size of the argument list and environment list taken together must not be greater than ARG_MAX bytes. See Section 28.1 [General Capacity Limits], page 537. In the GNU system, the size (which compares against ARG_MAX) includes, for each string, the number of characters in the string, plus the size of a char *, plus one, rounded up to a multiple of the size of a char *. Other systems may have somewhat different rules for counting.

These functions normally don’t return, since execution of a new program causes the currently executing program to go away completely. A value of -1 is returned in the event of a failure. In addition to the usual file name errors (see Section 6.2.3 [File Name Errors], page 89), the following errno error conditions are defined for these functions: E2BIG The combined size of the new program’s argument list and en-

vironment list is larger than ARG_MAX bytes. The GNU system has no specific limit on the argument list size, so this error code cannot result, but you may get ENOMEM instead if the arguments are too big for available memory.

ENOEXEC The specified file can’t be executed because it isn’t in the right format.

ENOMEM Executing the specified file requires more storage than is available.

If execution of the new file succeeds, it updates the access time field of the file as if the file had been read. See Section 9.8.9 [File Times], page 208, for more details about access times of files.

The point at which the file is closed again is not specified, but is at some point before the process exits or before another process image is executed.

Executing a new process image completely changes the contents of memory, copying only the argument and environment strings to new locations. But many other attributes of the process are unchanged:

• The process ID and the parent process ID. See Section 23.2 [Process Creation Concepts], page 468.

• Session and process group membership. See Section 24.1 [Concepts of Job Control], page 479.

• Real user ID and group ID, and supplementary group IDs. See Section 26.2 [The Persona of a Process], page 511.

• Pending alarms. See Section 17.3 [Setting an Alarm], page 356.

• Current working directory and root directory. See Section 9.1 [Working Directory], page 181. In the GNU system, the root directory is not copied when executing a setuid program; instead the system default root directory is used for the new program.

• File mode creation mask. See Section 9.8.7 [Assigning File Permissions], page 204.

• Process signal mask; see Section 21.7.3 [Process Signal Mask], page 433.

• Pending signals; see Section 21.7 [Blocking Signals], page 431.

• Elapsed processor time associated with the process; see Section 17.1 [Processor Time], page 339.

If the set-user-ID and set-group-ID mode bits of the process image file are set, this affects the effective user ID and effective group ID (respectively) of the process. These concepts are discussed in detail in Section 26.2 [The Persona of a Process], page 511.

Signals that are set to be ignored in the existing process image are also set to be ignored in the new process image. All other signals are set to the default action in the new process image. For more information about signals, see Chapter 21 [Signal Handling], page 393.

File descriptors open in the existing process image remain open in the new process image, unless they have the FD_CLOEXEC (close-on-exec) flag set. The files that remain open inherit all attributes of the open file description from the existing process image, including file locks. File descriptors are discussed in Chapter 8 [Low-Level Input/Output], page 151.

Streams, by contrast, cannot survive through exec functions, because they are located in the memory of the process itself. The new process image has no streams except those it creates afresh. Each of the streams in the pre-exec process image has a descriptor inside it, and these descriptors do survive through exec (provided that they do not have FD_CLOEXEC set). The new process image can reconnect these to new streams using fdopen (see Section 8.4 [Descriptors and Streams], page 159).

23.6 Process Completion

The functions described in this section are used to wait for a child process to terminate or stop, and determine its status. These functions are declared in the header file ‘sys/wait.h’.

Function pid_t waitpid (pid_t pid, int *status-ptr, int

options) The waitpid function is used to request status information from a child process whose process ID is pid. Normally, the calling process is suspended until the child process makes status information available by terminating. Other values for the pid argument have special interpretations. A value of -1 or WAIT_ANY requests status information for any child process; a value of 0 or WAIT_MYPGRP requests information for any child process in the same process group as the calling process; and any other negative value − pgid requests information for any child process whose process group ID is pgid. If status information for a child process is available immediately, this function returns immediately without waiting. If more than one eligible child process has status information available, one of them is chosen randomly, and its status is returned immediately. To get the status from the other eligible child processes, you need to call waitpid again. The options argument is a bit mask. Its value should be the bitwise OR (that is, the ‘|’ operator) of zero or more of the WNOHANG and WUNTRACED flags. You can use the WNOHANG flag to indicate that the parent process shouldn’t wait; and the WUNTRACED flag to request status information from stopped processes as well as processes that have terminated. The status information from the child process is stored in the object that status-ptr points to, unless status-ptr is a null pointer. The return value is normally the process ID of the child process whose status is reported. If the WNOHANG option was specified and no child process is waiting to be noticed, the value is zero. A value of -1 is returned in case of error. The following errno error conditions are defined for this function:

EINTR The function was interrupted by delivery of a signal to the calling process. See Section 21.5 [Primitives Interrupted by Signals], page 425.

ECHILD There are no child processes to wait for, or the specified pid is not a child of the calling process.

EINVAL An invalid value was provided for the options argument.

These symbolic constants are defined as values for the pid argument to the waitpid function.

WAIT_ANY

This constant macro (whose value is -1) specifies that waitpid should return status information about any child process.

WAIT_MYPGRP

This constant (with value 0) specifies that waitpid should return status information about any child process in the same process group as the calling process.

These symbolic constants are defined as flags for the options argument to the waitpid function. You can bitwise-OR the flags together to obtain a value to use as the argument.

WNOHANG

This flag specifies that waitpid should return immediately instead of waiting, if there is no child process ready to be noticed.

WUNTRACED

This flag specifies that waitpid should report the status of any child processes that have been stopped as well as those that have terminated.

Function pid_t wait (int *status-ptr)

This is a simplified version of waitpid, and is used to wait until any one child process terminates. The call:

     wait (&status)

is exactly equivalent to:

     waitpid (-1, &status, 0)

Function pid_t wait4 (pid_t pid, int *status-ptr, int options,

      struct rusage *usage)

If usage is a null pointer, wait4 is equivalent to waitpid (pid, statusptr, options). If usage is not null, wait4 stores usage figures for the child process in *rusage (but only if the child has terminated, not if it has stopped). See Section 17.5 [Resource Usage], page 360. This function is a BSD extension.

Here’s an example of how to use waitpid to get the status from all child processes that have terminated, without ever waiting. This function is designed to be a handler for SIGCHLD, the signal that indicates that at least one child process has terminated.

  void
  sigchld_handler (int signum)
  {
    int pid;
    int status;
    while (1)
      {
        pid = waitpid (WAIT_ANY, &status, WNOHANG);
        if (pid < 0)
          {
            perror ("waitpid");
            break;
          }
        if (pid == 0)
          break;
        notice_termination (pid, status);
      }
  }

23.7 Process Completion Status

If the exit status value (see Section 22.3 [Program Termination], page 462) of the child process is zero, then the status value reported by waitpid or wait is also zero. You can test for other kinds of information encoded in the returned status value using the following macros. These macros are defined in the header file ‘sys/wait.h’.

Macro int WIFEXITED (int status)

This macro returns a nonzero value if the child process terminated normally with exit or _exit.

Macro int WEXITSTATUS (int status)

If WIFEXITED is true of status, this macro returns the low-order 8 bits of the exit status value from the child process. See Section 22.3.2 [Exit Status], page 463.

Macro int WIFSIGNALED (int status)

This macro returns a nonzero value if the child process terminated because it received a signal that was not handled. See Chapter 21 [Signal Handling], page 393.

Macro int WTERMSIG (int status)

If WIFSIGNALED is true of status, this macro returns the signal number of the signal that terminated the child process.

Macro int WCOREDUMP (int status)

This macro returns a nonzero value if the child process terminated and produced a core dump.

Macro int WIFSTOPPED (int status)

This macro returns a nonzero value if the child process is stopped.

Macro int WSTOPSIG (int status)

If WIFSTOPPED is true of status, this macro returns the signal number of the signal that caused the child process to stop.

23.8 BSD Process Wait Functions

The GNU library also provides these related facilities for compatibility with BSD Unix. BSD uses the union wait data type to represent status values rather than an int. The two representations are actually interchangeable; they describe the same bit patterns. The GNU C Library defines macros such as WEXITSTATUS so that they will work on either kind of object, and the wait function is defined to accept either type of pointer as its status-ptr argument.

These functions are declared in ‘sys/wait.h’.

Data Type union wait

This data type represents program termination status values. It has the following members:

int w_termsig

The value of this member is the same as the result of the

          WTERMSIG macro.
int w_coredump

The value of this member is the same as the result of the

          WCOREDUMP macro.
int w_retcode

The value of this member is the same as the result of the

          WEXITSTATUS macro.
int w_stopsig

The value of this member is the same as the result of the

          WSTOPSIG macro.

Instead of accessing these members directly, you should use the equivalent macros.

The wait3 function is the predecessor to wait4, which is more flexible. wait3 is now obsolete.

Function pid_t wait3 (union wait *status-ptr, int options,

      struct rusage *usage)

If usage is a null pointer, wait3 is equivalent to waitpid (-1, statusptr, options). If usage is not null, wait3 stores usage figures for the child process in *rusage (but only if the child has terminated, not if it has stopped). See Section 17.5 [Resource Usage], page 360.

23.9 Process Creation Example

Here is an example program showing how you might write a function similar to the built-in system. It executes its command argument using the equivalent of ‘sh -c command’.

  #include <stddef.h>
  #include <stdlib.h>
  #include <unistd.h>
  #include <sys/types.h>
  #include <sys/wait.h>

/* Execute the command using this shell program. */

  #define SHELL "/bin/sh"

     int
     my_system (const char *command)
     {
       int status;
       pid_t pid;
       pid = fork ();
       if (pid == 0)
         {

/* This is the child process. Execute the shell command. */

           execl (SHELL, SHELL, "-c", command, NULL);
           _exit (EXIT_FAILURE);
         }
       else if (pid < 0)

/* The fork failed. Report failure. */

         status = -1;
       else

/* This is the parent process. Wait for the child to complete. */

         if (waitpid (pid, &status, 0) != pid)
           status = -1;
       return status;
     }

There are a couple of things you should pay attention to in this example. Remember that the first argv argument supplied to the program repre-

sents the name of the program being executed. That is why, in the call to execl, SHELL is supplied once to name the program to execute and a second time to supply a value for argv[0].

The execl call in the child process doesn’t return if it is successful. If it fails, you must do something to make the child process terminate. Just returning a bad status code with return would leave two processes running the original program. Instead, the right behavior is for the child process to report failure to its parent process.

Call _exit to accomplish this. The reason for using _exit instead of exit is to avoid flushing fully buffered streams such as stdout. The buffers of these streams probably contain data that was copied from the parent process by the fork, data that will be output eventually by the parent process. Calling exit in the child would output the data twice. See Section 22.3.5 [Termination Internals], page 465.

24 Job Control

Job control refers to the protocol for allowing a user to move between multiple process groups (or jobs) within a single login session. The job control facilities are set up so that appropriate behavior for most programs happens automatically and they need not do anything special about job control. So you can probably ignore the material in this chapter unless you are writing a shell or login program.

You need to be familiar with concepts relating to process creation (see Section 23.2 [Process Creation Concepts], page 468) and signal handling (see Chapter 21 [Signal Handling], page 393) in order to understand this material presented in this chapter.

24.1 Concepts of Job Control

The fundamental purpose of an interactive shell is to read commands from the user’s terminal and create processes to execute the programs specified by those commands. It can do this using the fork (see Section 23.4 [Creating a Process], page 469) and exec (see Section 23.5 [Executing a File], page 470) functions.

A single command may run just one process—but often one command uses several processes. If you use the ‘|’ operator in a shell command, you explicitly request several programs in their own processes. But even if you run just one program, it can use multiple processes internally. For example, a single compilation command such as ‘cc -c foo.c’ typically uses four processes (though normally only two at any given time). If you run make, its job is to run other programs in separate processes.

The processes belonging to a single command are called a process group or job. This is so that you can operate on all of them at once. For example, typing C-c sends the signal SIGINT to terminate all the processes in the foreground process group.

A session is a larger group of processes. Normally all the processes that stem from a single login belong to the same session.

Every process belongs to a process group. When a process is created, it becomes a member of the same process group and session as its parent process. You can put it in another process group using the setpgid function, provided the process group belongs to the same session.

The only way to put a process in a different session is to make it the initial process of a new session, or a session leader, using the setsid function. This also puts the session leader into a new process group, and you can’t move it out of that process group again.

Usually, new sessions are created by the system login program, and the session leader is the process running the user’s login shell.

A shell that supports job control must arrange to control which job can use the terminal at any time. Otherwise there might be multiple jobs trying to read from the terminal at once, and confusion about which process should receive the input typed by the user. To prevent this, the shell must cooperate with the terminal driver using the protocol described in this chapter.

The shell can give unlimited access to the controlling terminal to only one process group at a time. This is called the foreground job on that controlling terminal. Other process groups managed by the shell that are executing without such access to the terminal are called background jobs.

If a background job needs to read from its controlling terminal, it is stopped by the terminal driver; if the TOSTOP mode is set, likewise for writing. The user can stop a foreground job by typing the SUSP character (see Section 12.4.9 [Special Characters], page 286) and a program can stop any job by sending it a SIGSTOP signal. It’s the responsibility of the shell to notice when jobs stop, to notify the user about them, and to provide mechanisms for allowing the user to interactively continue stopped jobs and switch jobs between foreground and background.

See Section 24.4 [Access to the Controlling Terminal], page 481, for more information about I/O to the controlling terminal,

24.2 Job Control is Optional

Not all operating systems support job control. The GNU system does support job control, but if you are using the GNU library on some other system, that system may not support job control itself.

You can use the _POSIX_JOB_CONTROL macro to test at compile-time whether the system supports job control. See Section 28.2 [Overall System Options], page 538.

If job control is not supported, then there can be only one process group per session, which behaves as if it were always in the foreground. The functions for creating additional process groups simply fail with the error code

ENOSYS.

The macros naming the various job control signals (see Section 21.2.5 [Job Control Signals], page 402) are defined even if job control is not supported. However, the system never generates these signals, and attempts to send a job control signal or examine or specify their actions report errors or do nothing.

24.3 Controlling Terminal of a Process

One of the attributes of a process is its controlling terminal. Child processes created with fork inherit the controlling terminal from their parent process. In this way, all the processes in a session inherit the controlling terminal from the session leader. A session leader that has control of a terminal is called the controlling process of that terminal.

You generally do not need to worry about the exact mechanism used to allocate a controlling terminal to a session, since it is done for you by the system when you log in.

An individual process disconnects from its controlling terminal when it calls setsid to become the leader of a new session. See Section 24.7.2 [Process Group Functions], page 498.

24.4 Access to the Controlling Terminal

Processes in the foreground job of a controlling terminal have unrestricted access to that terminal; background processes do not. This section describes in more detail what happens when a process in a background job tries to access its controlling terminal.

When a process in a background job tries to read from its controlling terminal, the process group is usually sent a SIGTTIN signal. This normally causes all of the processes in that group to stop (unless they handle the signal and don’t stop themselves). However, if the reading process is ignoring or blocking this signal, then read fails with an EIO error instead.

Similarly, when a process in a background job tries to write to its controlling terminal, the default behavior is to send a SIGTTOU signal to the process group. However, the behavior is modified by the TOSTOP bit of the local modes flags (see Section 12.4.7 [Local Modes], page 282). If this bit is not set (which is the default), then writing to the controlling terminal is always permitted without sending a signal. Writing is also permitted if the SIGTTOU signal is being ignored or blocked by the writing process.

Most other terminal operations that a program can do are treated as reading or as writing. (The description of each operation should say which.)

For more information about the primitive read and write functions, see Section 8.2 [Input and Output Primitives], page 153.

24.5 Orphaned Process Groups

When a controlling process terminates, its terminal becomes free and a new session can be established on it. (In fact, another user could log in on the terminal.) This could cause a problem if any processes from the old session are still trying to use that terminal.

To prevent problems, process groups that continue running even after the session leader has terminated are marked as orphaned process groups.

When a process group becomes an orphan, its processes are sent a SIGHUP signal. Ordinarily, this causes the processes to terminate. However, if a program ignores this signal or establishes a handler for it (see Chapter 21 [Signal Handling], page 393), it can continue running as in the orphan process group even after its controlling process terminates; but it still cannot access the terminal any more.

24.6 Implementing a Job Control Shell

This section describes what a shell must do to implement job control, by presenting an extensive sample program to illustrate the concepts involved.

• Section 24.6.1 [Data Structures for the Shell], page 482, introduces the example and presents its primary data structures.

• Section 24.6.2 [Initializing the Shell], page 484, discusses actions which the shell must perform to prepare for job control.

• Section 24.6.3 [Launching Jobs], page 486, includes information about how to create jobs to execute commands.

• Section 24.6.4 [Foreground and Background], page 490, discusses what the shell should do differently when launching a job in the foreground as opposed to a background job.

• Section 24.6.5 [Stopped and Terminated Jobs], page 492, discusses reporting of job status back to the shell.

• Section 24.6.6 [Continuing Stopped Jobs], page 496, tells you how to continue jobs that have been stopped.

• Section 24.6.7 [The Missing Pieces], page 496, discusses other parts of the shell.

24.6.1 Data Structures for the Shell

All of the program examples included in this chapter are part of a simple shell program. This section presents data structures and utility functions which are used throughout the example.

The sample shell deals mainly with two data structures. The job type contains information about a job, which is a set of subprocesses linked together with pipes. The process type holds information about a single subprocess. Here are the relevant data structure declarations:

Chapter 24: Job Control 483

/* A process is a single process. */

typedef struct process
{
struct process *next;       /* next process in pipeline */
char **argv;                /* for exec */
pid_t pid;                  /* process ID */

char completed; /* true if process has completed */ char stopped; /* true if process has stopped */ int status; /* reported status value */

} process;

/* A job is a pipeline of processes. */

typedef struct job
{
struct job *next;           /* next active job */

char *command; /* command line, used for messages */

process *first_process;     /* list of processes in this job */
pid_t pgid;                 /* process group ID */

char notified; /* true if user told about stopped job */

struct termios tmodes;      /* saved terminal modes */
int stdin, stdout, stderr;  /* standard i/o channels */
} job;

/* The active jobs are linked into a list. This is its head. */

job *first_job = NULL;

Here are some utility functions that are used for operating on job objects.

/* Find the active job with the indicated pgid. */

job *
find_job (pid_t pgid)
{
job *j;
for (j = first_job; j; j = j->next)
  if (j->pgid == pgid)
    return j;
return NULL;
}

/* Return true if all processes in the job have stopped or completed. */

     int
     job_is_stopped (job *j)
     {
       process *p;
       for (p = j->first_process; p; p = p->next)
         if (!p->completed && !p->stopped)
           return 0;
       return 1;
     }

/* Return true if all processes in the job have completed. */

     int
     job_is_completed (job *j)
     {
       process *p;
       for (p = j->first_process; p; p = p->next)
         if (!p->completed)
           return 0;
       return 1;
     }

24.6.2 Initializing the Shell

When a shell program that normally performs job control is started, it has to be careful in case it has been invoked from another shell that is already doing its own job control.

A subshell that runs interactively has to ensure that it has been placed in the foreground by its parent shell before it can enable job control itself. It does this by getting its initial process group ID with the getpgrp function, and comparing it to the process group ID of the current foreground job associated with its controlling terminal (which can be retrieved using the tcgetpgrp function).

If the subshell is not running as a foreground job, it must stop itself by sending a SIGTTIN signal to its own process group. It may not arbitrarily put itself into the foreground; it must wait for the user to tell the parent shell to do this. If the subshell is continued again, it should repeat the check and stop itself again if it is still not in the foreground.

Once the subshell has been placed into the foreground by its parent shell, it can enable its own job control. It does this by calling setpgid to put itself into its own process group, and then calling tcsetpgrp to place this process group into the foreground.

When a shell enables job control, it should set itself to ignore all the job control stop signals so that it doesn’t accidentally stop itself. You can do this by setting the action for all the stop signals to SIG_IGN.

A subshell that runs non-interactively cannot and should not support job control. It must leave all processes it creates in the same process group as the shell itself; this allows the non-interactive shell and its child processes to be treated as a single job by the parent shell. This is easy to do—just don’t use any of the job control primitives—but you must remember to make the shell do it.

Here is the initialization code for the sample shell that shows how to do all of this.

/* Keep track of attributes of the shell. */

#include <sys/types.h>
#include <termios.h>
#include <unistd.h>
pid_t shell_pgid;
struct termios shell_tmodes;
int shell_terminal;
int shell_is_interactive;

/* Make sure the shell is running interactively as the foreground job before proceeding. */

void
init_shell ()
{

/* See if we are running interactively. */

  shell_terminal = STDIN_FILENO;
  shell_is_interactive = isatty (shell_terminal);
  if (shell_is_interactive)
    {

/* Loop until we are in the foreground. */

      while (tcgetpgrp (shell_terminal) != (shell_pgid = getpgrp ()))
        kill (- shell_pgid, SIGTTIN);

/* Ignore interactive and job-control signals. */

      signal (SIGINT, SIG_IGN);
      signal (SIGQUIT, SIG_IGN);
      signal (SIGTSTP, SIG_IGN);

           signal (SIGTTIN, SIG_IGN);
           signal (SIGTTOU, SIG_IGN);
           signal (SIGCHLD, SIG_IGN);

/* Put ourselves in our own process group. */

           shell_pgid = getpid ();
           if (setpgid (shell_pgid, shell_pgid) < 0)
             {
               perror ("Couldn't put the shell in its own process group");
               exit (1);
             }

/* Grab control of the terminal. */

           tcsetpgrp (shell_terminal, shell_pgid);

/* Save default terminal attributes for shell. */

           tcgetattr (shell_terminal, &shell_tmodes);
         }
     }

24.6.3 Launching Jobs

Once the shell has taken responsibility for performing job control on its controlling terminal, it can launch jobs in response to commands typed by the user.

To create the processes in a process group, you use the same fork and exec functions described in Section 23.2 [Process Creation Concepts], page 468. Since there are multiple child processes involved, though, things are a little more complicated and you must be careful to do things in the right order. Otherwise, nasty race conditions can result.

You have two choices for how to structure the tree of parent-child relationships among the processes. You can either make all the processes in the process group be children of the shell process, or you can make one process in group be the ancestor of all the other processes in that group. The sample shell program presented in this chapter uses the first approach because it makes bookkeeping somewhat simpler.

As each process is forked, it should put itself in the new process group by calling setpgid; see Section 24.7.2 [Process Group Functions], page 498. The first process in the new group becomes its process group leader, and its process ID becomes the process group ID for the group.

The shell should also call setpgid to put each of its child processes into the new process group. This is because there is a potential timing problem: each child process must be put in the process group before it begins executing a new program, and the shell depends on having all the child processes in the group before it continues executing. If both the child processes and the shell call setpgid, this ensures that the right things happen no matter which process gets to it first.

If the job is being launched as a foreground job, the new process group also needs to be put into the foreground on the controlling terminal using tcsetpgrp. Again, this should be done by the shell as well as by each of its child processes, to avoid race conditions.

The next thing each child process should do is to reset its signal actions. During initialization, the shell process set itself to ignore job control sig-

nals; see Section 24.6.2 [Initializing the Shell], page 484. As a result, any child processes it creates also ignore these signals by inheritance. This is definitely undesirable, so each child process should explicitly set the actions for these signals back to SIG_DFL just after it is forked.

Since shells follow this convention, applications can assume that they inherit the correct handling of these signals from the parent process. But every application has a responsibility not to mess up the handling of stop signals. Applications that disable the normal interpretation of the SUSP character should provide some other mechanism for the user to stop the job. When the user invokes this mechanism, the program should send a SIGTSTP signal to the process group of the process, not just to the process itself. See Section 21.6.2 [Signaling Another Process], page 428.

Finally, each child process should call exec in the normal way. This is also the point at which redirection of the standard input and output channels should be handled. See Section 8.8 [Duplicating Descriptors], page 166, for an explanation of how to do this.

Here is the function from the sample shell program that is responsible for launching a program. The function is executed by each child process immediately after it has been forked by the shell, and never returns.

     void
     launch_process (process *p, pid_t pgid,
                     int infile, int outfile, int errfile,
                     int foreground)
     {
       pid_t pid;
       if (shell_is_interactive)
         {

/* Put the process into the process group and give the process group the terminal, if appropriate. This has to be done both by the shell and in the individual child processes because of potential race conditions. */

           pid = getpid ();
           if (pgid == 0) pgid = pid;

setpgid (pid, pgid);
if (foreground)
  tcsetpgrp (shell_terminal, pgid);

/* Set the handling for job control signals back to the default. */

signal (SIGINT, SIG_DFL);
signal (SIGQUIT, SIG_DFL);
signal (SIGTSTP, SIG_DFL);
signal (SIGTTIN, SIG_DFL);
signal (SIGTTOU, SIG_DFL);
signal (SIGCHLD, SIG_DFL);
}

/* Set the standard input/output channels of the new process. */

if (infile != STDIN_FILENO)
{
dup2 (infile, STDIN_FILENO);
close (infile);
}
if (outfile != STDOUT_FILENO)
{
dup2 (outfile, STDOUT_FILENO);
close (outfile);
}
if (errfile != STDERR_FILENO)
{
dup2 (errfile, STDERR_FILENO);
close (errfile);
}

/* Exec the new process. Make sure we exit. */

execvp (p->argv[0], p->argv);
perror ("execvp");
exit (1);
}

If the shell is not running interactively, this function does not do anything with process groups or signals. Remember that a shell not performing job control must keep all of its subprocesses in the same process group as the shell itself.

Next, here is the function that actually launches a complete job. After creating the child processes, this function calls some other functions to put the newly created job into the foreground or background; these are discussed in Section 24.6.4 [Foreground and Background], page 490.

void

launch_job (job *j, int foreground)
{
process *p;
pid_t pid;
int mypipe[2], infile, outfile;
infile = j->stdin;
for (p = j->first_process; p; p = p->next)
{

/* Set up pipes, if necessary. */

if (p->next)
  {
    if (pipe (mypipe) < 0)
      {
        perror ("pipe");
        exit (1);
      }
    outfile = mypipe[1];
  }
else
  outfile = j->stdout;

/* Fork the child processes. */

pid = fork ();
if (pid == 0)

/* This is the child process. */

  launch_process (p, j->pgid, infile,
                  outfile, j->stderr, foreground);
else if (pid < 0)
  {

/* The fork failed. */

    perror ("fork");
    exit (1);
  }
else
  {

/* This is the parent process. */

    p->pid = pid;
    if (shell_is_interactive)
      {
        if (!j->pgid)
          j->pgid = pid;
        setpgid (pid, j->pgid);
      }

             }

/* Clean up after pipes. */

           if (infile != j->stdin)
             close (infile);
           if (outfile != j->stdout)
             close (outfile);
           infile = mypipe[0];
         }
       format_job_info (j, "launched");
       if (!shell_is_interactive)
         wait_for_job (j);
       else if (foreground)
         put_job_in_foreground (j, 0);
       else
         put_job_in_background (j, 0);
     }

24.6.4 Foreground and Background

Now let’s consider what actions must be taken by the shell when it launches a job into the foreground, and how this differs from what must be done when a background job is launched.

When a foreground job is launched, the shell must first give it access to the controlling terminal by calling tcsetpgrp. Then, the shell should wait for processes in that process group to terminate or stop. This is discussed in more detail in Section 24.6.5 [Stopped and Terminated Jobs], page 492.

When all of the processes in the group have either completed or stopped, the shell should regain control of the terminal for its own process group by calling tcsetpgrp again. Since stop signals caused by I/O from a background process or a SUSP character typed by the user are sent to the process group, normally all the processes in the job stop together.

The foreground job may have left the terminal in a strange state, so the shell should restore its own saved terminal modes before continuing. In case the job is merely been stopped, the shell should first save the current terminal modes so that it can restore them later if the job is continued. The functions for dealing with terminal modes are tcgetattr and tcsetattr; these are described in Section 12.4 [Terminal Modes], page 273.

Here is the sample shell’s function for doing all of this.

/* Put job j in the foreground. If cont is nonzero, restore the saved terminal modes and send the process group a SIGCONT signal to wake it up before we block. */

void
put_job_in_foreground (job *j, int cont)
{

/* Put the job into the foreground. */

tcsetpgrp (shell_terminal, j->pgid);

/* Send the job a continue signal, if necessary. */

if (cont)
  {
    tcsetattr (shell_terminal, TCSADRAIN, &j->tmodes);
    if (kill (- j->pgid, SIGCONT) < 0)
      perror ("kill (SIGCONT)");
  }

/* Wait for it to report. */

wait_for_job (j);

/* Put the shell back in the foreground. */

tcsetpgrp (shell_terminal, shell_pgid);

/* Restore the shell’s terminal modes. */

tcgetattr (shell_terminal, &j->tmodes);
tcsetattr (shell_terminal, TCSADRAIN, &shell_tmodes);
}

If the process group is launched as a background job, the shell should remain in the foreground itself and continue to read commands from the terminal.

In the sample shell, there is not much that needs to be done to put a job into the background. Here is the function it uses:

/* Put a job in the background. If the cont argument is true, send the process group a SIGCONT signal to wake it up. */

void
put_job_in_background (job *j, int cont)
{

/* Send the job a continue signal, if necessary. */

if (cont)
  if (kill (-j->pgid, SIGCONT) < 0)
    perror ("kill (SIGCONT)");
}

24.6.5 Stopped and Terminated Jobs

When a foreground process is launched, the shell must block until all of the processes in that job have either terminated or stopped. It can do this by calling the waitpid function; see Section 23.6 [Process Completion], page 473. Use the WUNTRACED option so that status is reported for processes that stop as well as processes that terminate.

The shell must also check on the status of background jobs so that it can report terminated and stopped jobs to the user; this can be done by calling waitpid with the WNOHANG option. A good place to put a such a check for terminated and stopped jobs is just before prompting for a new command.

The shell can also receive asynchronous notification that there is status information available for a child process by establishing a handler for SIGCHLD signals. See Chapter 21 [Signal Handling], page 393.

In the sample shell program, the SIGCHLD signal is normally ignored. This is to avoid reentrancy problems involving the global data structures the shell manipulates. But at specific times when the shell is not using these data structures—such as when it is waiting for input on the terminal—it makes sense to enable a handler for SIGCHLD. The same function that is used to do the synchronous status checks (do_job_notification, in this case) can also be called from within this handler.

Here are the parts of the sample shell program that deal with checking the status of jobs and reporting the information to the user.

/* Store the status of the process pid that was returned by waitpid. Return 0 if all went well, nonzero otherwise. */

     int
     mark_process_status (pid_t pid, int status)
     {
       job *j;
       process *p;

if (pid > 0)
  {

/* Update the record for the process. */

    for (j = first_job; j; j = j->next)
      for (p = j->first_process; p; p = p->next)
        if (p->pid == pid)
          {
            p->status = status;
            if (WIFSTOPPED (status))
              p->stopped = 1;
            else
              {
                p->completed = 1;
                if (WIFSIGNALED (status))
                  fprintf (stderr, "%d: Terminated by signal %d.\n",
                           (int) pid, WTERMSIG (p->status));
              }
            return 0;
           }
    fprintf (stderr, "No child process %d.\n", pid);
    return -1;
  }
else if (pid == 0 || errno == ECHILD)

/* No processes ready to report. */

  return -1;
else {

/* Other weird errors. */

  perror ("waitpid");
  return -1;
}
}

/* Check for processes that have status information available, without blocking. */

void
update_status (void)
{
int status;
pid_t pid;
do
  pid = waitpid (WAIT_ANY, &status, WUNTRACED|WNOHANG);
while (!mark_process_status (pid, status));
}

/* Check for processes that have status information available, blocking until all processes in the given job have reported. */

void
wait_for_job (job *j)
{
  int status;
  pid_t pid;
  do
    pid = waitpid (WAIT_ANY, &status, WUNTRACED);
  while (!mark_process_status (pid, status)
         && !job_is_stopped (j)
         && !job_is_completed (j));
}

/* Format information about job status for the user to look at. */

void
format_job_info (job *j, const char *status)
{
  fprintf (stderr, "%ld (%s): %s\n", (long)j->pgid, status, j->command);
}

/* Notify the user about stopped or terminated jobs. Delete terminated jobs from the active job list. */

void
do_job_notification (void)
{
job *j, *jlast, *jnext;
process *p;

/* Update status information for child processes. */

update_status ();
jlast = NULL;
for (j = first_job; j; j = jnext)
{
jnext = j->next;

/* If all processes have completed, tell the user the job has

completed and delete it from the list of active jobs. */

if (job_is_completed (j)) {
  format_job_info (j, "completed");
  if (jlast)
    jlast->next = jnext;
  else
    first_job = jnext;
  free_job (j);
}

/* Notify the user about stopped jobs,

marking them so that we won’t do this more than once. */

else if (job_is_stopped (j) && !j->notified) {
  format_job_info (j, "stopped");
  j->notified = 1;
  jlast = j;
}

/* Don’t say anything about jobs that are still running. */

else
  jlast = j;
}
}

24.6.6 Continuing Stopped Jobs

The shell can continue a stopped job by sending a SIGCONT signal to its process group. If the job is being continued in the foreground, the shell should first invoke tcsetpgrp to give the job access to the terminal, and restore the saved terminal settings. After continuing a job in the foreground, the shell should wait for the job to stop or complete, as if the job had just been launched in the foreground.

The sample shell program handles both newly created and continued jobs with the same pair of functions, put_job_in_foreground and put_job_in_background. The definitions of these functions were given in Section 24.6.4 [Foreground and Background], page 490. When continuing a stopped job, a nonzero value is passed as the cont argument to ensure that the SIGCONT signal is sent and the terminal modes reset, as appropriate.

This leaves only a function for updating the shell’s internal bookkeeping about the job being continued:

/* Mark a stopped job J as being running again. */

     void
     mark_job_as_running (job *j)
     {
       Process *p;
       for (p = j->first_process; p; p = p->next)
         p->stopped = 0;
       j->notified = 0;
     }

/* Continue the job J. */

     void
     continue_job (job *j, int foreground)
     {
       mark_job_as_running (j);
       if (foreground)
         put_job_in_foreground (j, 1);
       else
         put_job_in_background (j, 1);
     }

24.6.7 The Missing Pieces

The code extracts for the sample shell included in this chapter are only a part of the entire shell program. In particular, nothing at all has been said about how job and program data structures are allocated and initialized.

Most real shells provide a complex user interface that has support for a command language; variables; abbreviations, substitutions, and pattern matching on file names; and the like. All of this is far too complicated to explain here! Instead, we have concentrated on showing how to implement the core process creation and job control functions that can be called from such a shell.

Here is a table summarizing the major entry points we have presented:

void init_shell (void)

Initialize the shell’s internal state. See Section 24.6.2 [Initializing the Shell], page 484.

void launch_job (job *j, int foreground)

Launch the job j as either a foreground or background job. See Section 24.6.3 [Launching Jobs], page 486.

void do_job_notification (void)

Check for and report any jobs that have terminated or stopped. Can be called synchronously or within a handler for SIGCHLD signals. See Section 24.6.5 [Stopped and Terminated Jobs], page 492.

void continue_job (job *j, int foreground)

Continue the job j. See Section 24.6.6 [Continuing Stopped Jobs], page 496.

Of course, a real shell would also want to provide other functions for managing jobs. For example, it would be useful to have commands to list all active jobs or to send a signal (such as SIGKILL) to a job.

24.7 Functions for Job Control

This section contains detailed descriptions of the functions relating to job control.

24.7.1 Identifying the Controlling Terminal

You can use the ctermid function to get a file name that you can use to open the controlling terminal. In the GNU library, it returns the same string all the time: "/dev/tty". That is a special “magic” file name that refers to the controlling terminal of the current process (if it has one). To find the name of the specific terminal device, use ttyname; see Section 12.1 [Identifying Terminals], page 271.

The function ctermid is declared in the header file ‘stdio.h’.

Function char * ctermid (char *string)

The ctermid function returns a string containing the file name of the controlling terminal for the current process. If string is not a null pointer, it should be an array that can hold at least L_ctermid characters; the string is returned in this array. Otherwise, a pointer to a string in a static area is returned, which might get overwritten on subsequent calls to this function. An empty string is returned if the file name cannot be determined for any reason. Even if a file name is returned, access to the file it represents is not guaranteed.

Macro int L ctermid

The value of this macro is an integer constant expression that represents the size of a string large enough to hold the file name returned

by ctermid.

See also the isatty and ttyname functions, in Section 12.1 [Identifying Terminals], page 271.

24.7.2 Process Group Functions

Here are descriptions of the functions for manipulating process groups. Your program should include the header files ‘sys/types.h’ and ‘unistd.h’ to use these functions.

Function pid_t setsid (void)

The setsid function creates a new session. The calling process becomes the session leader, and is put in a new process group whose process group ID is the same as the process ID of that process. There are initially no other processes in the new process group, and no other process groups in the new session. This function also makes the calling process have no controlling terminal. The setsid function returns the new process group ID of the calling process if successful. A return value of -1 indicates an error. The following errno error conditions are defined for this function:

EPERM The calling process is already a process group leader, or there is already another process group around that has the same process group ID.

The getpgrp function has two definitions: one derived from BSD Unix, and one from the POSIX.1 standard. The feature test macros you have selected (see Section 1.3.4 [Feature Test Macros], page 8) determine which definition you get. Specifically, you get the BSD version if you define _BSD_ SOURCE; otherwise, you get the POSIX version if you define _POSIX_SOURCE or _GNU_SOURCE. Programs written for old BSD systems will not include ‘unistd.h’, which defines getpgrp specially under _BSD_SOURCE. You must link such programs with the -lbsd-compat option to get the BSD definition.

POSIX.1 Function pid_t getpgrp (void)

The POSIX.1 definition of getpgrp returns the process group ID of the calling process.

BSD Function pid_t getpgrp (pid_t pid)

The BSD definition of getpgrp returns the process group ID of the process pid. You can supply a value of 0 for the pid argument to get information about the calling process.

Function int setpgid (pid_t pid, pid_t pgid)

The setpgid function puts the process pid into the process group pgid. As a special case, either pid or pgid can be zero to indicate the process ID of the calling process. This function fails on a system that does not support job control. See Section 24.2 [Job Control is Optional], page 480, for more information. If the operation is successful, setpgid returns zero. Otherwise it returns -1. The following errno error conditions are defined for this function:

EACCES The child process named by pid has executed an exec function since it was forked.

EINVAL The value of the pgid is not valid.

ENOSYS The system doesn’t support job control.

EPERM The process indicated by the pid argument is a session leader, or is not in the same session as the calling process, or the value of the pgid argument doesn’t match a process group ID in the same session as the calling process.

ESRCH The process indicated by the pid argument is not the calling process or a child of the calling process.

Function int setpgrp (pid_t pid, pid_t pgid)

This is the BSD Unix name for setpgid. Both functions do exactly the same thing.

24.7.3 Functions for Controlling Terminal Access

These are the functions for reading or setting the foreground process group of a terminal. You should include the header files ‘sys/types.h’ and ‘unistd.h’ in your application to use these functions.

Although these functions take a file descriptor argument to specify the terminal device, the foreground job is associated with the terminal file itself and not a particular open file descriptor.

Function pid_t tcgetpgrp (int filedes)

This function returns the process group ID of the foreground process group associated with the terminal open on descriptor filedes. If there is no foreground process group, the return value is a number greater than 1 that does not match the process group ID of any existing process group. This can happen if all of the processes in the job that was formerly the foreground job have terminated, and no other job has yet been moved into the foreground. In case of an error, a value of -1 is returned. The following errno error conditions are defined for this function:

EBADF The filedes argument is not a valid file descriptor.

ENOSYS The system doesn’t support job control.

ENOTTY The terminal file associated with the filedes argument isn’t the controlling terminal of the calling process.

Function int tcsetpgrp (int filedes, pid_t pgid)

This function is used to set a terminal’s foreground process group ID. The argument filedes is a descriptor which specifies the terminal; pgid specifies the process group. The calling process must be a member of the same session as pgid and must have the same controlling terminal. For terminal access purposes, this function is treated as output. If it is called from a background process on its controlling terminal, normally all processes in the process group are sent a SIGTTOU signal. The exception is if the calling process itself is ignoring or blocking SIGTTOU signals, in which case the operation is performed and no signal is sent. If successful, tcsetpgrp returns 0. A return value of -1 indicates an error. The following errno error conditions are defined for this function:

EBADF The filedes argument is not a valid file descriptor.

EINVAL The pgid argument is not valid.

ENOSYS The system doesn’t support job control.

ENOTTY The filedes isn’t the controlling terminal of the calling process.

EPERM The pgid isn’t a process group in the same session as the calling process.

25 System Databases and Name Service Switch

Various functions in the C Library need to be configured to work correctly in the local environment. Traditionally, this was done by using files (e.g., ‘/etc/passwd’), but other nameservices (like the Network Information Service (NIS) and the Domain Name Service (DNS)) became popular, and were hacked into the C library, usually with a fixed search order (see section “frobnicate” in The Jargon File).

The GNU C Library contains a cleaner solution of this problem. It is designed after a method used by Sun Microsystems in the C library of Solaris 2. GNU C Library follows their name and calls this scheme Name Service Switch (NSS).

Though the interface might be similar to Sun’s version there is no common code. We never saw any source code of Sun’s implementation and so the internal interface is incompatible. This also manifests in the file names we use as we will see later.

25.1 NSS Basics

The basic idea is to put the implementation of the different services offered to access the databases in separate modules. This has some advantages:

1. Contributors can add new services without adding them to GNU C Library.

2. The modules can be updated separately.

3. The C library image is smaller.

To fulfill the first goal above the ABI of the modules will be described below. For getting the implementation of a new service right it is important to understand how the functions in the modules get called. They are in no way designed to be used by the programmer directly. Instead the programmer should only use the documented and standardized functions to access the databases. The databases available in the NSS are

aliases Mail aliases

ethers Ethernet numbers,

group Groups of users, see Section 26.13 [Group Database], page 525.

hosts Host names and numbers, see Section 11.5.2.4 [Host Names], page 235.

netgroup Network wide list of host and users, see Section 26.14 [Netgroup Database], page 528. network Network names and numbers, see Section 11.12 [Networks

Database], page 269.

protocols

Network protocols, see Section 11.5.6 [Protocols Database], page 240.

passwd User passwords, see Section 26.12 [User Database], page 521. rpc Remote procedure call names and numbers, services Network services, see Section 11.5.4 [The Services Database],

page 238.

shadow Shadow user passwords, There will be some more added later (automount, bootparams, netmasks,

and publickey).

25.2 The NSS Configuration File

Somehow the NSS code must be told about the wishes of the user. For this reason there is the file ‘/etc/nsswitch.conf’. For each database this file contain a specification how the lookup process should work. The file could look like this:

    # /etc/nsswitch.conf
    #
    # Name Service Switch configuration file.
    #
    passwd:     db files nis
    shadow:     files
    group:      db files nis
    hosts:      files nisplus nis dns
    networks:   nisplus [NOTFOUND=return] files
    ethers:     nisplus [NOTFOUND=return] db files
    protocols:  nisplus [NOTFOUND=return] db files
    rpc:        nisplus [NOTFOUND=return] db files
    services:   nisplus [NOTFOUND=return] db files

The first column is the database as you can guess from the table above. The rest of the line specifies how the lookup process works. Please note that you specify the way it works for each database individually. This cannot be done with the old way of a monolithic implementation.

The configuration specification for each database can contain two different items:

• the service specification like files, db, or nis.

• the reaction on lookup result like [NOTFOUND=return].

25.2.1 Services in the NSS configuration File

The above example file mentions four different services: files, db, nis, and nisplus. This does not mean these services are available on all sites and it does also not mean these are all the services which will ever be available.

In fact, these names are simply strings which the NSS code uses to find the implicitly addressed functions. The internal interface will be described later. Visible to the user are the modules which implement an individual service.

Assume the service name shall be used for a lookup. The code for this service is implemented in a module called ‘libnss_name’. On a system supporting shared libraries this is in fact a shared library with the name (for example) ‘libnss_name.so.1’. The number at the end is the currently used version of the interface which will not change frequently. Normally the user should not have to be cognizant of these files since they should be placed in a directory where they are found automatically. Only the names of all available services are important.

25.2.2 Actions in the NSS configuration

The second item in the specification gives the user much finer control on the lookup process. Action items are placed between two service names and are written within brackets. The general form is

[ ( !? status = action )+ ] where

     status ⇒ success | notfound | unavail | tryagain
     action ⇒ return | continue

The case of the keywords is insignificant. The status values are the results of a call to a lookup function of a specific service. They mean

‘success’ No error occurred and the wanted entry is returned. The default action for this is return.

‘notfound’

The lookup process works ok but the needed value was not found. The default action is continue.

‘unavail’ The service is permanently unavailable. This can either mean the needed file is not available, or, for DNS, the server is not available or does not allow queries. The default action is

          continue.

‘tryagain’

The service is temporarily unavailable. This could mean a file is locked or a server currently cannot accept more connections. The default action is continue.

If we have a line like

     ethers: nisplus [NOTFOUND=return] db files

this is equivalent to

     ethers: nisplus [SUCCESS=return NOTFOUND=return UNAVAIL=continue
                      TRYAGAIN=continue]
             db      [SUCCESS=return NOTFOUND=continue UNAVAIL=continue
                      TRYAGAIN=continue]
             files

(except that it would have to be written on one line). The default value for the actions are normally what you want, and only need to be changed in exceptional cases.

If the optional ! is placed before the status this means the following action is used for all statii but status itself. I.e., ! is negation as in the C language (and others).

Before we explain the exception which makes this action item necessary one more remark: obviously it makes no sense to add another action item after the files service. Since there is no other service following the action always is return.

Now, why is this [NOTFOUND=return] action useful? To understand this we should know that the nisplus service is often complete; i.e., if an entry is not available in the NIS+ tables it is not available anywhere else. This is what is expressed by this action item: it is useless to examine further services since they will not give us a result.

The situation would be different if the NIS+ service is not available because the machine is booting. In this case the return value of the lookup function is not notfound but instead unavail. And as you can see in the complete form above: in this situation the db and files services are used. Neat, isn’t it? The system administrator need not pay special care for the time the system is not completely ready to work (while booting or shutdown or network problems).

25.2.3 Notes on the NSS Configuration File

Finally a few more hints. The NSS implementation is not completely helpless if ‘/etc/nsswitch.conf’ does not exist. For all supported databases there is a default value so it should normally be possible to get the system running even if the file is corrupted or missing.

For the hosts and network databases the default value is dns [!UNAVAIL=return] files. I.e., the system is prepared for the DNS service not to be available but if it is available the answer it returns is ultimative.

The passwd, group, and shadow databases are traditionally handled in a special way. The appropriate files in the ‘/etc’ directory are read but if an entry with a name starting with a + character is found NIS is used. This kind of lookup remains possible by using the special lookup service compat and the default value for the three databases above is compat [NOTFOUND=return]

files.

For all other databases the default value is nis [NOTFOUND=return] files. This solution give the best chance to be correct since NIS and file based lookup is used.

A second point is that the user should try to optimize the lookup process. The different service have different response times. A simple file look up on a local file could be fast, but if the file is long and the needed entry is near the end of the file this may take quite some time. In this case it might be better to use the db service which allows fast local access to large data sets.

Often the situation is that some global information like NIS must be used. So it is unavoidable to use service entries like nis etc. But one should avoid slow services like this if possible.

25.3 NSS Module Internals

Now it is time to described how the modules look like. The functions contained in a module are identified by their names. I.e., there is no jump table or the like. How this is done is of no interest here; those interested in this topic should read about Dynamic Linking.

25.3.1 The Naming Scheme of the NSS Modules

The name of each function consist of various parts:

nss service function service of course corresponds to the name of the module this function is

found in.^{1} The function part is derived from the interface function in the C library itself. If the user calls the function gethostbyname and the service used is files the function

            _nss_files_gethostbyname_r

in the module

            libnss_files.so.1

1 Now you might ask why to duplicate this information. The answer is that we want to keep the possibility to link directly with these shared objects. is used. You see, what is explained above in not the whole truth. In fact the NSS modules only contain reentrant versions of the lookup functions. I.e., if the user would call the gethostbyname_r function this also would end in the above function. For all user interface functions the C library maps this call to a call to the reentrant function. For reentrant functions this is trivial since the interface is (nearly) the same. For the non-reentrant version The library keeps internal buffers which are used to replace the user supplied buffer.

I.e., the reentrant functions can have counterparts. No service module is forced to have functions for all databases and all kinds to access them. If a function is not available it is simply treated as if the function would return unavail (see Section 25.2.2 [Actions in the NSS configuration], page 503).

The file name ‘libnss_files.so.1’ would be on a Solaris 2 system ‘nss_files.so.1’. This is the difference mentioned above. Sun’s NSS modules are usable as modules which get indirectly loaded only.

The NSS modules in the GNU C Library are prepared to be used as normal libraries itself. This is not true in the moment, though. But the different organization of the name space in the modules does not make it impossible like it is for Solaris. Now you can see why the modules are still libraries.^{2}

25.3.2 The Interface of the Function in NSS Modules

Now we know about the functions contained in the modules. It is now time to describe the types. When we mentioned the reentrant versions of the functions above, this means there are some additional arguments (compared with the standard, non-reentrant version). The prototypes for the nonreentrant and reentrant versions of our function above are:

     struct hostent *gethostbyname (const char *name)
     int gethostbyname_r (const char *name, struct hostent *result_buf,
                          char *buf, size_t buflen, struct hostent **result,
                          int *h_errnop)

The actual prototype of the function in the NSS modules in this case is

     enum nss_status _nss_files_gethostbyname_r (const char *name,
                                                 struct hostent *result_buf,
                                                 char *buf, size_t buflen,
                                                 int *h_errnop)

I.e., the interface function is in fact the reentrant function with the change of the return value and the omission of the result parameter. While the user-

2 There is a second explanation: we were too lazy to change the Makefiles to allow the generation of shared objects not starting with ‘lib’ but do not tell this anybody. level function returns a pointer to the result the reentrant function return

an enum nss_status value:
NSS_STATUS_TRYAGAIN

numeric value -2

NSS_STATUS_UNAVAIL

numeric value -1

NSS_STATUS_NOTFOUND

numeric value 0

NSS_STATUS_SUCCESS

numeric value 1

Now you see where the action items of the ‘/etc/nsswitch.conf’ file are used.

If you study the source code you will find there is a fifth value: NSS_ STATUS_RETURN. This is an internal use only value, used by a few functions in places where none of the above value can be used. If necessary the source code should be examined to learn about the details.

The above function has something special which is missing for almost all the other module functions. There is an argument h errnop. This points to a variable which will be filled with the error code in case the execution of the function fails for some reason. The reentrant function cannot use the global variable h errno; gethostbyname calls gethostbyname_r with the last argument set to &h_errno.

The getXXXbyYYY functions are the most important functions in the NSS modules. But there are others which implement the other ways to access system databases (say for the password database, there are setpwent, getpwent, and endpwent). These will be described in more detail later. Here we give a general way to determine the signature of the module function:

• the return value is int;

• the name is as explain in see Section 25.3.1 [The Naming Scheme of the NSS Modules], page 505;

• the first arguments are identical to the arguments of the non-reentrant function;

• the next three arguments are:

    STRUCT_TYPE *result_buf

pointer to buffer where the result is stored. STRUCT_TYPE is normally a struct which corresponds to the database.

    char *buffer

pointer to a buffer where the function can store additional adata for the result etc.

    size_t buflen

length of the buffer pointed to by buffer.

• possibly a last argument h errnop, for the host name and network name lookup functions.

This table is correct for all functions but the set…ent and end…ent functions.

25.4 Extending NSS

One of the advantages of NSS mentioned above is that it can be extended quite easily. There are two ways in which the extension can happen: adding another database or adding another service. The former is normally done only by the C library developers. It is here only important to remember that adding another database is independent from adding another service because a service need not support all databases or lookup functions.

A designer/implementor of a new service is therefore free to choose the databases s/he is interested in and leave the rest for later (or completely aside).

25.4.1 Adding another Service to NSS

The sources for a new service need not (and should not) be part of the GNU C Library itself. The developer retains complete control over the sources and its development. The links between the C library and the new service module consists solely of the interface functions.

Each module is designed following a specific interface specification. For now the version is 1 and this manifests in the version number of the shared library object of the NSS modules: they have the extension .1. If the interface ever changes in an incompatible way, this number will be increased— hopefully this will never be necessary. Modules using the old interface will still be usable.

Developers of a new service will have to make sure that their module is created using the correct interface number. This means the file itself must have the correct name and on ElF systems the soname (Shared Object Name) must also have this number. Building a module from a bunch of object files on an ELF system using GNU CC could be done like this:

     gcc -shared -o libnss_NAME.so.1 -Wl,-soname,libnss_NAME.so.1 OBJECTS

section “Link Options” in GNU CC, to learn more about this command line.

To use the new module the library must be able to find it. This can be achieved by using options for the dynamic linker so that it will search directory where the binary is placed. For an ELF system this could be done by adding the wanted directory to the value of LD_LIBRARY_PATH.

But this is not always possible since some program (those which run under IDs which do not belong to the user) ignore this variable. Therefore the stable version of the module should be placed into a directory which is searched by the dynamic linker. Normally this should be the directory ‘$prefix/lib’, where ‘$prefix’ corresponds to the value given to configure using the --prefix option. But be careful: this should only be done if it is clear the module does not cause any harm. System administrators should be careful.

25.4.2 Internals of the NSS Module Functions

Until now we only provided the syntactic interface for the functions in the NSS module. In fact there is not more much we can tell since the implementation obviously is different for each function. But a few general rules must be followed by all functions.

In fact there are four kinds of different functions which may appear in the interface. All derive from the traditional ones for system databases. db in the following table is normally an abbreviation for the database (e.g., it is pw for the password database).

enum nss_status _nss_database_setdbent (void)

This function prepares the service for following operations. For a simple file based lookup this means files could be opened, for other services this function simply is a noop. One special case for this function is that it takes an additional argument for some databases (i.e., the interface is int setdbent (int)). Section 11.5.2.4 [Host Names], page 235, which describes the sethostent function. The return value should be NSS STATUS SUCCESS or according to the table above in case of an error (see Section 25.3.2 [The Interface of the Function in NSS Modules], page 506).

enum nss_status _nss_database_enddbent (void)

This function simply closes all files which are still open or removes buffer caches. If there are no files or buffers to remove this is again a simple noop. There normally is no return value different to NSS STATUS SUCCESS.

enum nss_status _nss_database_getdbent_r (STRUCTURE *result,
char *buffer, size_t buflen)

Since this function will be called several times in a row to retrieve one entry after the other it must keep some kind of state. But this also means the functions are not really reentrant. They are reentrant only in that simultaneous calls to this function will not try to write the retrieved data in the same place (as it would be the case for the non-reentrant functions); instead, it writes to the structure pointed to by the result parameter. But the calls share a common state and in the case of a file access this means they return neighboring entries in the file. The buffer of length buflen pointed to by buffer can be used for storing some additional data for the result. It is not guaranteed that the same buffer will be passed for the next call of this function. Therefore one must not misuse this buffer to save some state information from one call to another. As explained above this function could also have an additional last argument. This depends on the database used; it happens

only for host and network.

The function shall return NSS_STATUS_SUCCESS as long as their are more entries. When the last entry was read it should return NSS_STATUS_NOTFOUND. When the buffer given as an argument is too small for the data to be returned NSS_STATUS_TRYAGAIN should be returned. When the service was not formerly initialized by a call to _nss_DATABASE_setdbent all return value allowed for this function can also be returned here.

enum nss_status _nss_DATABASE_getdbbyXX_r (PARAMS,
STRUCTURE *result, char *buffer, size_t buflen)

This function shall return the entry from the database which is addressed by the PARAMS. The type and number of these arguments vary. It must be individually determined by looking to the user-level interface functions. All arguments given to the non-reentrant version are here described by PARAMS. The result must be stored in the structure pointed to by result. If there is additional data to return (say strings, where the result structure only contains pointers) the function must use the buffer or length buflen. There must not be any references to non-constant global data. The implementation of this function should honour the stayopen flag set by the setDBent function whenever this makes sense. Again, this function takes an additional last argument for the host and network database. The return value should as always follow the rules given above (see Section 25.3.2 [The Interface of the Function in NSS Modules], page 506).

26 Users and Groups

Every user who can log in on the system is identified by a unique number called the user ID. Each process has an effective user ID which says which user’s access permissions it has.

Users are classified into groups for access control purposes. Each process has one or more group ID values which say which groups the process can use for access to files.

The effective user and group IDs of a process collectively form its persona. This determines which files the process can access. Normally, a process inherits its persona from the parent process, but under special circumstances a process can change its persona and thus change its access permissions.

Each file in the system also has a user ID and a group ID. Access control works by comparing the user and group IDs of the file with those of the running process.

The system keeps a database of all the registered users, and another database of all the defined groups. There are library functions you can use to examine these databases.

26.1 User and Group IDs

Each user account on a computer system is identified by a user name (or login name) and user ID. Normally, each user name has a unique user ID, but it is possible for several login names to have the same user ID. The user names and corresponding user IDs are stored in a data base which you can access as described in Section 26.12 [User Database], page 521.

Users are classified in groups. Each user name also belongs to one or more groups, and has one default group. Users who are members of the same group can share resources (such as files) that are not accessible to users who are not a member of that group. Each group has a group name and group ID. See Section 26.13 [Group Database], page 525, for how to find information about a group ID or group name.

26.2 The Persona of a Process

At any time, each process has a single user ID and a group ID which determine the privileges of the process. These are collectively called the persona of the process, because they determine “who it is” for purposes of access control. These IDs are also called the effective user ID and effective group ID of the process.

Your login shell starts out with a persona which consists of your user ID and your default group ID. In normal circumstances, all your other processes inherit these values.

A process also has a real user ID which identifies the user who created the process, and a real group ID which identifies that user’s default group. These values do not play a role in access control, so we do not consider them part of the persona. But they are also important.

Both the real and effective user ID can be changed during the lifetime of a process. See Section 26.3 [Why Change the Persona of a Process?], page 512.

In addition, a user can belong to multiple groups, so the persona includes supplementary group IDs that also contribute to access permission.

For details on how a process’s effective user IDs and group IDs affect its permission to access files, see Section 9.8.6 [How Your Access to a File is Decided], page 204.

The user ID of a process also controls permissions for sending signals using the kill function. See Section 21.6.2 [Signaling Another Process], page 428.

26.3 Why Change the Persona of a Process?

The most obvious situation where it is necessary for a process to change its user and/or group IDs is the login program. When login starts running, its user ID is root. Its job is to start a shell whose user and group IDs are those of the user who is logging in. (To accomplish this fully, login must set the real user and group IDs as well as its persona. But this is a special case.)

The more common case of changing persona is when an ordinary user program needs access to a resource that wouldn’t ordinarily be accessible to the user actually running it.

For example, you may have a file that is controlled by your program but that shouldn’t be read or modified directly by other users, either because it implements some kind of locking protocol, or because you want to preserve the integrity or privacy of the information it contains. This kind of restricted access can be implemented by having the program change its effective user or group ID to match that of the resource.

Thus, imagine a game program that saves scores in a file. The game program itself needs to be able to update this file no matter who is running it, but if users can write the file without going through the game, they can give themselves any scores they like. Some people consider this undesirable, or even reprehensible. It can be prevented by creating a new user ID and login name (say, games) to own the scores file, and make the file writable only by this user. Then, when the game program wants to update this file, it can change its effective user ID to be that for games. In effect, the program must adopt the persona of games so it can write the scores file.

26.4 How an Application Can Change Persona

The ability to change the persona of a process can be a source of unintentional privacy violations, or even intentional abuse. Because of the potential for problems, changing persona is restricted to special circumstances.

You can’t arbitrarily set your user ID or group ID to anything you want; only privileged processes can do that. Instead, the normal way for a program to change its persona is that it has been set up in advance to change to a particular user or group. This is the function of the setuid and setgid bits of a file’s access mode. See Section 9.8.5 [The Mode Bits for Access Permission], page 202.

When the setuid bit of an executable file is set, executing that file automatically changes the effective user ID to the user that owns the file. Likewise, executing a file whose setgid bit is set changes the effective group ID to the group of the file. See Section 23.5 [Executing a File], page 470. Creating a file that changes to a particular user or group ID thus requires full access to that user or group ID.

See Section 9.8 [File Attributes], page 195, for a more general discussion of file modes and accessibility.

A process can always change its effective user (or group) ID back to its real ID. Programs do this so as to turn off their special privileges when they are not needed, which makes for more robustness.

26.5 Reading the Persona of a Process

Here are detailed descriptions of the functions for reading the user and group IDs of a process, both real and effective. To use these facilities, you must include the header files ‘sys/types.h’ and ‘unistd.h’.

Data Type uid t

This is an integer data type used to represent user IDs. In the GNU library, this is an alias for unsigned int.

Data Type gid t

This is an integer data type used to represent group IDs. In the GNU library, this is an alias for unsigned int.

Function uid_t getuid (void)

The getuid function returns the real user ID of the process.

Function gid_t getgid (void)

The getgid function returns the real group ID of the process.

Function uid_t geteuid (void)

The geteuid function returns the effective user ID of the process.

Function gid_t getegid (void)

The getegid function returns the effective group ID of the process.

Function int getgroups (int count, gid_t *groups)

The getgroups function is used to inquire about the supplementary group IDs of the process. Up to count of these group IDs are stored in the array groups; the return value from the function is the number of group IDs actually stored. If count is smaller than the total number of supplementary group IDs, then getgroups returns a value of -1 and

errno is set to EINVAL.

If count is zero, then getgroups just returns the total number of supplementary group IDs. On systems that do not support supplementary groups, this will always be zero. Here’s how to use getgroups to read all the supplementary group IDs:

     gid_t *
     read_all_groups (void)
     {
       int ngroups = getgroups (0, NULL);
       gid_t *groups
         = (gid_t *) xmalloc (ngroups * sizeof (gid_t));
       int val = getgroups (ngroups, groups);
       if (val < 0)
         {
           free (groups);
           return NULL;
         }
       return groups;
     }

26.6 Setting the User ID

This section describes the functions for altering the user ID (real and/or effective) of a process. To use these facilities, you must include the header

files ‘sys/types.h’ and ‘unistd.h’.

Function int setuid (uid_t newuid)

This function sets both the real and effective user ID of the process to newuid, provided that the process has appropriate privileges. If the process is not privileged, then newuid must either be equal to the real user ID or the saved user ID (if the system supports the _ POSIX_SAVED_IDS feature). In this case, setuid sets only the effective user ID and not the real user ID.

The setuid function returns a value of 0 to indicate successful completion, and a value of -1 to indicate an error. The following errno error conditions are defined for this function:

EINVAL The value of the newuid argument is invalid.

EPERM The process does not have the appropriate privileges; you do not have permission to change to the specified ID.

Function int setreuid (uid_t ruid, uid_t euid)

This function sets the real user ID of the process to ruid and the effective user ID to euid. If ruid is -1, it means not to change the real user ID; likewise if euid is -1, it means not to change the effective user ID. The setreuid function exists for compatibility with 4.3 BSD Unix, which does not support saved IDs. You can use this function to swap the effective and real user IDs of the process. (Privileged processes are not limited to this particular usage.) If saved IDs are supported, you should use that feature instead of this function. See Section 26.8 [Enabling and Disabling Setuid Access], page 516. The return value is 0 on success and -1 on failure. The following errno error conditions are defined for this function:

EPERM The process does not have the appropriate privileges; you do not have permission to change to the specified ID.

26.7 Setting the Group IDs

This section describes the functions for altering the group IDs (real and effective) of a process. To use these facilities, you must include the header

files ‘sys/types.h’ and ‘unistd.h’.

Function int setgid (gid_t newgid)

This function sets both the real and effective group ID of the process to newgid, provided that the process has appropriate privileges. If the process is not privileged, then newgid must either be equal to the real group ID or the saved group ID. In this case, setgid sets only the effective group ID and not the real group ID. The return values and error conditions for setgid are the same as those for setuid.

Function int setregid (gid_t rgid, fid_t egid)

This function sets the real group ID of the process to rgid and the effective group ID to egid. If rgid is -1, it means not to change the real group ID; likewise if egid is -1, it means not to change the effective group ID.

The setregid function is provided for compatibility with 4.3 BSD Unix, which does not support saved IDs. You can use this function to swap the effective and real group IDs of the process. (Privileged processes are not limited to this usage.) If saved IDs are supported, you should use that feature instead of using this function. See Section 26.8 [Enabling and Disabling Setuid Access], page 516. The return values and error conditions for setregid are the same as those for setreuid.

The GNU system also lets privileged processes change their supplementary group IDs. To use setgroups or initgroups, your programs should include the header file ‘grp.h’.

Function int setgroups (size_t count, gid_t *groups)

This function sets the process’s supplementary group IDs. It can only be called from privileged processes. The count argument specifies the number of group IDs in the array groups. This function returns 0 if successful and -1 on error. The following errno error conditions are defined for this function: EPERM The calling process is not privileged.

Function int initgroups (const char *user, gid_t gid)

The initgroups function effectively calls setgroups to set the process’s supplementary group IDs to be the normal default for the user name user. The group ID gid is also included.

26.8 Enabling and Disabling Setuid Access

A typical setuid program does not need its special access all of the time. It’s a good idea to turn off this access when it isn’t needed, so it can’t possibly give unintended access.

If the system supports the saved user ID feature, you can accomplish this with setuid. When the game program starts, its real user ID is jdoe, its effective user ID is games, and its saved user ID is also games. The program should record both user ID values once at the beginning, like this:

  user_user_id = getuid ();
  game_user_id = geteuid ();

Then it can turn off game file access with

  setuid (user_user_id);

and turn it on with

  setuid (game_user_id);

Throughout this process, the real user ID remains jdoe and the saved user ID remains games, so the program can always set its effective user ID to either one.

On other systems that don’t support the saved user ID feature, you can turn setuid access on and off by using setreuid to swap the real and effective user IDs of the process, as follows:

     setreuid (geteuid (), getuid ());

This special case is always allowed—it cannot fail.

Why does this have the effect of toggling the setuid access? Suppose a game program has just started, and its real user ID is jdoe while its effective user ID is games. In this state, the game can write the scores file. If it swaps the two uids, the real becomes games and the effective becomes jdoe; now the program has only jdoe access. Another swap brings games back to the effective user ID and restores access to the scores file.

In order to handle both kinds of systems, test for the saved user ID feature with a preprocessor conditional, like this:

     #ifdef _POSIX_SAVED_IDS
       setuid (user_user_id);
     #else
       setreuid (geteuid (), getuid ());
     #endif

26.9 Setuid Program Example

Here’s an example showing how to set up a program that changes its effective user ID.

This is part of a game program called caber-toss that manipulates a file ‘scores’ that should be writable only by the game program itself. The program assumes that its executable file will be installed with the set-user-ID bit set and owned by the same user as the ‘scores’ file. Typically, a system administrator will set up an account like games for this purpose.

The executable file is given mode 4755, so that doing an ‘ls -l’ on it produces output like:

     -rwsr-xr-x   1 games    184422 Jul 30 15:17 caber-toss

The set-user-ID bit shows up in the file modes as the ‘s’.

The scores file is given mode 644, and doing an ‘ls -l’ on it shows:

     -rw-r--r--  1 games           0 Jul 31 15:33 scores

Here are the parts of the program that show how to set up the changed user ID. This program is conditionalized so that it makes use of the saved IDs feature if it is supported, and otherwise uses setreuid to swap the effective and real user IDs.

     #include <stdio.h>
     #include <sys/types.h>
     #include <unistd.h>
     #include <stdlib.h>

/* Save the effective and real UIDs. */

static uid_t euid, ruid;

/* Restore the effective UID to its original value. */

void
do_setuid (void)
{
  int status;
#ifdef _POSIX_SAVED_IDS
  status = setuid (euid);
#else
  status = setreuid (ruid, euid);
#endif
  if (status < 0) {
    fprintf (stderr, "Couldn't set uid.\n");
    exit (status);
    }
}

/* Set the effective UID to the real UID. */

void
undo_setuid (void)
{
  int status;
#ifdef _POSIX_SAVED_IDS
  status = setuid (ruid);
#else
  status = setreuid (euid, ruid);
#endif
  if (status < 0) {
    fprintf (stderr, "Couldn't set uid.\n");
    exit (status);
    }
}

/* Main program. */

Chapter 26: Users and Groups 519

int
main (void)
{

/* Save the real and effective user IDs. */

ruid = getuid ();
euid = geteuid ();
undo_setuid ();

/* Do the game and record the score. */ …

}

Notice how the first thing the main function does is to set the effective user ID back to the real user ID. This is so that any other file accesses that are performed while the user is playing the game use the real user ID for determining permissions. Only when the program needs to open the scores file does it switch back to the original effective user ID, like this:

/* Record the score. */

int
record_score (int score)
{
FILE *stream;
char *myname;

/* Open the scores file. */

do_setuid ();
stream = fopen (SCORES_FILE, "a");
undo_setuid ();

/* Write the score to the file. */

if (stream)
  {
    myname = cuserid (NULL);
    if (score < 0)
      fprintf (stream, "%10s: Couldn't lift the caber.\n", myname);
    else
      fprintf (stream, "%10s: %d feet.\n", myname, score);
    fclose (stream);
    return 0;
  }
else
  return -1;
}

26.10 Tips for Writing Setuid Programs

It is easy for setuid programs to give the user access that isn’t intended— in fact, if you want to avoid this, you need to be careful. Here are some guidelines for preventing unintended access and minimizing its consequences when it does occur:

• Don’t have setuid programs with privileged user IDs such as root unless it is absolutely necessary. If the resource is specific to your particular program, it’s better to define a new, nonprivileged user ID or group ID just to manage that resource.

• Be cautious about using the system and exec functions in combination with changing the effective user ID. Don’t let users of your program execute arbitrary programs under a changed user ID. Executing a shell is especially bad news. Less obviously, the execlp and execvp functions are a potential risk (since the program they execute depends on the user’s PATH environment variable). If you must exec another program under a changed ID, specify an absolute file name (see Section 6.2.2 [File Name Resolution], page 88) for the executable, and make sure that the protections on that executable and all containing directories are such that ordinary users cannot replace it with some other program.

• Only use the user ID controlling the resource in the part of the program that actually uses that resource. When you’re finished with it, restore the effective user ID back to the actual user’s user ID. See Section 26.8 [Enabling and Disabling Setuid Access], page 516.

• If the setuid part of your program needs to access other files besides the controlled resource, it should verify that the real user would ordinarily have permission to access those files. You can use the access function (see Section 9.8.6 [How Your Access to a File is Decided], page 204) to check this; it uses the real user and group IDs, rather than the effective IDs.

26.11 Identifying Who Logged In

You can use the functions listed in this section to determine the login name of the user who is running a process, and the name of the user who logged in the current session. See also the function getuid and friends (see Section 26.5 [Reading the Persona of a Process], page 513).

The getlogin function is declared in ‘unistd.h’, while cuserid and

L_cuserid are declared in ‘stdio.h’.

Function char * getlogin (void)

The getlogin function returns a pointer to a string containing the name of the user logged in on the controlling terminal of the process, or a null pointer if this information cannot be determined. The string is statically allocated and might be overwritten on subsequent calls to this function or to cuserid.

Function char * cuserid (char *string)

The cuserid function returns a pointer to a string containing a user name associated with the effective ID of the process. If string is not a null pointer, it should be an array that can hold at least L_cuserid characters; the string is returned in this array. Otherwise, a pointer to a string in a static area is returned. This string is statically allocated and might be overwritten on subsequent calls to this function or to

getlogin.

The use of this function is deprecated since it is marked to be withdrawn in XPG4.2 and it is already removed in POSIX.1.

Macro int L cuserid

An integer constant that indicates how long an array you might need to store a user name.

These functions let your program identify positively the user who is running or the user who logged in this session. (These can differ when setuid programs are involved; See Section 26.2 [The Persona of a Process], page 511.) The user cannot do anything to fool these functions.

For most purposes, it is more useful to use the environment variable LOGNAME to find out who the user is. This is more flexible precisely because the user can set LOGNAME arbitrarily. See Section 22.2.2 [Standard Environment Variables], page 460.

26.12 User Database

This section describes all about how to search and scan the database of registered users. The database itself is kept in the file ‘/etc/passwd’ on most systems, but on some systems a special network server gives access to it.

26.12.1 The Data Structure that Describes a User

The functions and data structures for accessing the system user database are declared in the header file ‘pwd.h’.

Data Type struct passwd

The passwd data structure is used to hold information about entries in the system user data base. It has at least the following members:

char *pw_name

The user’s login name.

char *pw_passwd.

The encrypted password string.

uid_t pw_uid

The user ID number.

gid_t pw_gid

The user’s default group ID number.

char *pw_gecos

A string typically containing the user’s real name, and possibly other information such as a phone number.

char *pw_dir

The user’s home directory, or initial working directory. This might be a null pointer, in which case the interpretation is system-dependent.

char *pw_shell

The user’s default shell, or the initial program run when the user logs in. This might be a null pointer, indicating that the system default should be used.

26.12.2 Looking Up One User

You can search the system user database for information about a specific user using getpwuid or getpwnam. These functions are declared in ‘pwd.h’.

Function struct passwd * getpwuid (uid_t uid)

This function returns a pointer to a statically-allocated structure containing information about the user whose user ID is uid. This structure may be overwritten on subsequent calls to getpwuid. A null pointer value indicates there is no user in the data base with user ID uid.

Function int getpwuid r (uid_t uid, struct passwd
      *result buf, char *buffer, size_t buflen, struct passwd

**result) This function is similar to getpwuid in that is returns information about the user whose user ID is uid. But the result is not placed in a static buffer. Instead the user supplied structure pointed to by result buf is filled with the information. The first buflen bytes of the additional buffer pointed to by buffer are used to contain additional information, normally strings which are pointed to by the elements of the result structure. If the return value is 0 the pointer returned in result points to the record which contains the wanted data (i.e., result contains the value result buf ). In case the return value is non null there is no user in the data base with user ID uid or the buffer buffer is too small to contain all the needed information. In the later case the global errno variable is set to ERANGE.

Function struct passwd * getpwnam (const char *name)

This function returns a pointer to a statically-allocated structure containing information about the user whose user name is name. This structure may be overwritten on subsequent calls to getpwnam. A null pointer value indicates there is no user named name.

Function int getpwnam r (const char *name, struct passwd
      *result buf, char *buffer, size_t buflen, struct passwd

**result) This function is similar to getpwnam in that is returns information about the user whose user name is name. But the result is not placed in a static buffer. Instead the user supplied structure pointed to by result buf is filled with the information. The first buflen bytes of the additional buffer pointed to by buffer are used to contain additional information, normally strings which are pointed to by the elements of the result structure. If the return value is 0 the pointer returned in result points to the record which contains the wanted data (i.e., result contains the value result buf ). In case the return value is non null there is no user in the data base with user name name or the buffer buffer is too small to contain all the needed information. In the later case the global errno variable is set to ERANGE.

26.12.3 Scanning the List of All Users

This section explains how a program can read the list of all users in the system, one user at a time. The functions described here are declared in

‘pwd.h’.

You can use the fgetpwent function to read user entries from a particular file.

Function struct passwd * fgetpwent (FILE *stream)

This function reads the next user entry from stream and returns a pointer to the entry. The structure is statically allocated and is rewritten on subsequent calls to fgetpwent. You must copy the contents of the structure if you wish to save the information. This stream must correspond to a file in the same format as the standard password database file. This function comes from System V.

Function int fgetpwent r (FILE *stream, struct passwd

      *result buf, char *buffer, size_t buflen, struct passwd

**result) This function is similar to fgetpwent in that it reads the next user entry from stream. But the result is returned in the structure pointed to by result buf. The first buflen bytes of the additional buffer pointed to by buffer are used to contain additional information, normally strings which are pointed to by the elements of the result structure. This stream must correspond to a file in the same format as the standard password database file. If the function returns null result points to the structure with the wanted data (normally this is in result buf ). If errors occurred the return value is non-null and result contains a null pointer.

The way to scan all the entries in the user database is with setpwent,

getpwent, and endpwent.

Function void setpwent (void)

This function initializes a stream which getpwent and getpwent_r use to read the user database.

Function struct passwd * getpwent (void)

The getpwent function reads the next entry from the stream initialized by setpwent. It returns a pointer to the entry. The structure is statically allocated and is rewritten on subsequent calls to getpwent. You must copy the contents of the structure if you wish to save the information. A null pointer is returned in case no further entry is available.

Function int getpwent r (struct passwd *result buf, char

      *buffer, int buflen, struct passwd **result)

This function is similar to getpwent in that it returns the next entry from the stream initialized by setpwent. But in contrast to the getpwent function this function is reentrant since the result is placed in the user supplied structure pointed to by result buf. Additional data, normally the strings pointed to by the elements of the result structure, are placed in the additional buffer or length buflen starting at buffer. If the function returns zero result points to the structure with the wanted data (normally this is in result buf ). If errors occurred the return value is non-zero and result contains a null pointer.

Function void endpwent (void)

This function closes the internal stream used by getpwent or

getpwent_r.

26.12.4 Writing a User Entry

Function int putpwent (const struct passwd *p, FILE

*stream) This function writes the user entry *p to the stream stream, in the format used for the standard user database file. The return value is zero on success and nonzero on failure. This function exists for compatibility with SVID. We recommend that you avoid using it, because it makes sense only on the assumption that the struct passwd structure has no members except the standard ones; on a system which merges the traditional Unix data base with other extended information about users, adding an entry using this function would inevitably leave out much of the important information. The function putpwent is declared in ‘pwd.h’.

26.13 Group Database

This section describes all about how to search and scan the database of registered groups. The database itself is kept in the file ‘/etc/group’ on most systems, but on some systems a special network service provides access to it.

26.13.1 The Data Structure for a Group

The functions and data structures for accessing the system group database are declared in the header file ‘grp.h’.

Data Type struct group

The group structure is used to hold information about an entry in the system group database. It has at least the following members:

char *gr_name

The name of the group.

gid_t gr_gid

The group ID of the group.

char **gr_mem

A vector of pointers to the names of users in the group. Each user name is a null-terminated string, and the vector itself is terminated by a null pointer.

26.13.2 Looking Up One Group

You can search the group database for information about a specific group using getgrgid or getgrnam. These functions are declared in ‘grp.h’.

Function struct group * getgrgid (gid_t gid)

This function returns a pointer to a statically-allocated structure containing information about the group whose group ID is gid. This structure may be overwritten by subsequent calls to getgrgid. A null pointer indicates there is no group with ID gid.

Function int getgrgid r (gid_t gid, struct group *result buf,

      char *buffer, size_t buflen, struct group **result)

This function is similar to getgrgid in that is returns information about the group whose group ID is gid. But the result is not placed in a static buffer. Instead the user supplied structure pointed to by result buf is filled with the information. The first buflen bytes of the additional buffer pointed to by buffer are used to contain additional information, normally strings which are pointed to by the elements of the result structure. If the return value is 0 the pointer returned in result points to the record which contains the wanted data (i.e., result contains the value result buf ). If the return value is non-zero there is no group in the data base with group ID gid or the buffer buffer is too small to contain all the needed information. In the later case the global errno variable is set to ERANGE.

Function struct group * getgrnam (const char *name)

This function returns a pointer to a statically-allocated structure containing information about the group whose group name is name. This structure may be overwritten by subsequent calls to getgrnam. A null pointer indicates there is no group named name.

Function int getgrnam r (const char *name, struct group
      *result buf, char *buffer, size_t buflen, struct group

**result) This function is similar to getgrnam in that is returns information about the group whose group name is name. But the result is not placed in a static buffer. Instead the user supplied structure pointed to by result buf is filled with the information. The first buflen bytes of the additional buffer pointed to by buffer are used to contain additional information, normally strings which are pointed to by the elements of the result structure. If the return value is 0 the pointer returned in result points to the record which contains the wanted data (i.e., result contains the value result buf ). If the return value is non-zero there is no group in the data base with group name name or the buffer buffer is too small to contain all the needed information. In the later case the global errno variable is set to ERANGE.

26.13.3 Scanning the List of All Groups

This section explains how a program can read the list of all groups in the system, one group at a time. The functions described here are declared in

‘grp.h’.

You can use the fgetgrent function to read group entries from a particular file.

Function struct group * fgetgrent (FILE *stream)

The fgetgrent function reads the next entry from stream. It returns a pointer to the entry. The structure is statically allocated and is rewritten on subsequent calls to fgetgrent. You must copy the contents of the structure if you wish to save the information. The stream must correspond to a file in the same format as the standard group database file.

Function int fgetgrent r (FILE *stream, struct group

      *result buf, char *buffer, size_t buflen, struct group

**result) This function is similar to fgetgrent in that it reads the next user entry from stream. But the result is returned in the structure pointed to by result buf. The first buflen bytes of the additional buffer pointed to by buffer are used to contain additional information, normally strings which are pointed to by the elements of the result structure. This stream must correspond to a file in the same format as the standard group database file. If the function returns zero result points to the structure with the wanted data (normally this is in result buf ). If errors occurred the return value is non-zero and result contains a null pointer.

The way to scan all the entries in the group database is with setgrent,

getgrent, and endgrent.

Function void setgrent (void)

This function initializes a stream for reading from the group data base. You use this stream by calling getgrent or getgrent_r.

Function struct group * getgrent (void)

The getgrent function reads the next entry from the stream initialized by setgrent. It returns a pointer to the entry. The structure is statically allocated and is rewritten on subsequent calls to getgrent. You must copy the contents of the structure if you wish to save the information.

Function int getgrent r (struct group *result buf, char

         *buffer, size_t buflen, struct group **result)

This function is similar to getgrent in that it returns the next entry from the stream initialized by setgrent. But in contrast to the getgrent function this function is reentrant since the result is placed in the user supplied structure pointed to by result buf. Additional data, normally the strings pointed to by the elements of the result structure, are placed in the additional buffer or length buflen starting at buffer. If the function returns zero result points to the structure with the wanted data (normally this is in result buf ). If errors occurred the return value is non-zero and result contains a null pointer.

Function void endgrent (void)

This function closes the internal stream used by getgrent or

  getgrent_r.

26.14 Netgroup Database

26.14.1 Netgroup Data

Sometimes it is useful group users according to other criterias like the ones used in the See Section 26.13 [Group Database], page 525. E.g., it is useful to associate a certain group of users with a certain machine. On the other hand grouping of host names is not supported so far.

In Sun Microsystems SunOS appeared a new kind of database, the netgroup database. It allows to group hosts, users, and domain freely, giving them individual names. More concrete: a netgroup is a list of triples consisting of a host name, a user name, and a domain name, where any of the entries can be a wildcard entry, matching all inputs. A last possibility is that names of other netgroups can also be given in the list specifying a netgroup. So one can construct arbitrary hierarchies without loops.

Sun’s implementation allows netgroups only for the nis or nisplus service see Section 25.2.1 [Services in the NSS configuration File], page 503. The implementation in the GNU C library has no such restriction. An entry in either of the input services must have the following form:

groupname ( groupname | (hostname,username,domainname) )+

Any of the fields in the triple can be empty which means anything matches. While describing the functions we will see that the opposite case is useful as well. I.e., there may be entries which will not match any input. For entries like a name consisting of the single character - shall be used.

26.14.2 Looking up one Netgroup

The lookup functions for netgroups are a bit different to all other system database handling functions. Since a single netgroup can contain many entries a two-step process is needed. First a single netgroup is selected and then one can iterate over all entries in this netgroup. These functions are declared in ‘netdb.h’.

Function int setnetgrent (const char *netgroup)

A call to this function initializes the internal state of the library to allow following calls of the getnetgrent iterate over all entries in the netgroup with name netgroup. When the call is successful (i.e., when a netgroup with this name exist) the return value is 1. When the return value is 0 no netgroup of this name is known or some other error occurred.

It is important to remember that there is only one single state for iterating the netgroups. Even if the programmer uses the getnetgrent_r function the result is not really reentrant since always only one single netgroup at a time can be processed. If the program needs to process more than one netgroup simultaneously she must protect this by using external locking. This problem was introduced in the original netgroups implementation in SunOS and since we must stay compatible it is not possible to change this.

Some other functions also use the netgroups state. Currently these are the innetgr function and parts of the implementation of the compat service part of the NSS implementation.

Function int getnetgrent (char **hostp, char **userp, char

**domainp) This function returns the next unprocessed entry of the currently selected netgroup. The string pointers, which addresses are passed in the arguments hostp, userp, and domainp, will contain after a successful call pointers to appropriate strings. If the string in the next entry is empty the pointer has the value NULL. The returned string pointers are only valid unless no of the netgroup related functions are called. The return value is 1 if the next entry was successfully read. A value of 0 means no further entries exist or internal errors occurred.

Function int getnetgrent r (char **hostp, char **userp, char

**domainp, char *buffer, int buflen) This function is similar to getnetgrent with only one exception: the strings the three string pointers hostp, userp, and domainp point to, are placed in the buffer of buflen bytes starting at buffer. This means the returned values are valid even after other netgroup related functions are called.

The return value is 1 if the next entry was successfully read and the buffer contains enough room to place the strings in it. 0 is returned in case no more entries are found, the buffer is too small, or internal errors occurred. This function is a GNU extension. The original implementation in the SunOS libc does not provide this function.

Function void endnetgrent (void)

This function free all buffers which were allocated to process the last selected netgroup. As a result all string pointers returned by calls to getnetgrent are invalid afterwards.

26.14.3 Testing for Netgroup Membership

It is often not necessary to scan the whole netgroup since often the only interesting question is whether a given entry is part of the selected netgroup.

Function int innetgr (const char *netgroup, const char *host,

      const char *user, const char *domain)

This function tests whether the triple specified by the parameters hostp, userp, and domainp is part of the netgroup netgroup. Using this function has the advantage that

1. no other netgroup function can use the global netgroup state since internal locking is used and

2. the function is implemented more efficiently than successive calls to the other set/get/endnetgrent functions.

Any of the pointers hostp, userp, and domainp can be NULL which means any value is excepted in this position. This is also true for the name - which should not match any other string otherwise. The return value is 1 if an entry matching the given triple is found in the netgroup. The return value is 0 if the netgroup itself is not found, the netgroup does not contain the triple or internal errors occurred.

26.15 User and Group Database Example

Here is an example program showing the use of the system database inquiry functions. The program prints some information about the user running the program.

  #include <grp.h>
  #include <pwd.h>
  #include <sys/types.h>
  #include <unistd.h>
  #include <stdlib.h>

Chapter 26: Users and Groups 531

int
main (void)
{
uid_t me;
struct passwd *my_passwd;
struct group *my_group;
char **members;

/* Get information about the user ID. */

me = getuid ();
my_passwd = getpwuid (me);
if (!my_passwd)
  {
    printf ("Couldn't find out about user %d.\n", (int) me);
    exit (EXIT_FAILURE);
  }

/* Print the information. */

printf ("I am %s.\n", my_passwd->pw_gecos);
printf ("My login name is %s.\n", my_passwd->pw_name);
printf ("My uid is %d.\n", (int) (my_passwd->pw_uid));
printf ("My home directory is %s.\n", my_passwd->pw_dir);
printf ("My default shell is %s.\n", my_passwd->pw_shell);

/* Get information about the default group ID. */

my_group = getgrgid (my_passwd->pw_gid);
if (!my_group)
  {
    printf ("Couldn't find out about group %d.\n",
            (int) my_passwd->pw_gid);
    exit (EXIT_FAILURE);
  }

/* Print the information. */

printf ("My default group is %s (%d).\n",
        my_group->gr_name, (int) (my_passwd->pw_gid));
printf ("The members of this group are:\n");
members = my_group->gr_mem;
while (*members)
  {
    printf ("  %s\n", *(members));
    members++;
  }

  return EXIT_SUCCESS;
}

Here is some output from this program:

I am Throckmorton Snurd.
My login name is snurd.
My uid is 31093.
My home directory is /home/fsg/snurd.
My default shell is /bin/sh.
My default group is guest (12).
The members of this group are:
  friedman
  tami

27 System Information

This chapter describes functions that return information about the particular machine that is in use—the type of hardware, the type of software, and the individual machine’s name.

27.1 Host Identification

This section explains how to identify the particular machine that your program is running on. The identification of a machine consists of its Internet host name and Internet address; see Section 11.5 [The Internet Namespace], page 229. The host name should always be a fully qualified domain name,

like ‘crispy-wheats-n-chicken.ai.mit.edu’, not a simple name like just
‘crispy-wheats-n-chicken’.

Prototypes for these functions appear in ‘unistd.h’. The shell commands hostname and hostid work by calling them.

Function int gethostname (char *name, size_t size)

This function returns the name of the host machine in the array name. The size argument specifies the size of this array, in bytes. The return value is 0 on success and -1 on failure. In the GNU C library, gethostname fails if size is not large enough; then you can try again with a larger array. The following errno error condition is defined for this function:

ENAMETOOLONG

The size argument is less than the size of the host name plus one.

On some systems, there is a symbol for the maximum possible host name length: MAXHOSTNAMELEN. It is defined in ‘sys/param.h’. But you can’t count on this to exist, so it is cleaner to handle failure and try again. gethostname stores the beginning of the host name in name even if the host name won’t entirely fit. For some purposes, a truncated host name is good enough. If it is, you can ignore the error code.

Function int sethostname (const char *name, size_t length)

The sethostname function sets the name of the host machine to name, a string with length length. Only privileged processes are allowed to do this. Usually it happens just once, at system boot time. The return value is 0 on success and -1 on failure. The following errno error condition is defined for this function:

EPERM This process cannot set the host name because it is not privileged.

Function long int gethostid (void)

This function returns the “host ID” of the machine the program is running on. By convention, this is usually the primary Internet address of that machine, converted to a long int. However, some systems it is a meaningless but unique number which is hard-coded for each machine.

Function int sethostid (long int id)

The sethostid function sets the “host ID” of the host machine to id. Only privileged processes are allowed to do this. Usually it happens just once, at system boot time. The return value is 0 on success and -1 on failure. The following errno error condition is defined for this function:

EPERM This process cannot set the host name because it is not privileged.

ENOSYS The operating system does not support setting the host ID. On some systems, the host ID is a meaningless but unique number hard-coded for each machine.

27.2 Hardware/Software Type Identification

You can use the uname function to find out some information about the type of computer your program is running on. This function and the associated data type are declared in the header file ‘sys/utsname.h’.

Data Type struct utsname

The utsname structure is used to hold information returned by the uname function. It has the following members:

char sysname[]

This is the name of the operating system in use.

char nodename[]

This is the network name of this particular computer. In the GNU library, the value is the same as that returned by gethostname; see Section 27.1 [Host Identification], page 533.

char release[]

This is the current release level of the operating system implementation.

char version[]

This is the current version level within the release of the operating system.

char machine[]

This is a description of the type of hardware that is in use. Some systems provide a mechanism to interrogate the kernel directly for this information. On systems without such a mechanism, the GNU C library fills in this field based on the configuration name that was specified when building and installing the library. GNU uses a three-part name to describe a system configuration; the three parts are cpu, manufacturer and systemtype, and they are separated with dashes. Any possible combination of three names is potentially meaningful, but most such combinations are meaningless in practice and even the meaningful ones are not necessarily supported by any particular GNU program. Since the value in machine is supposed to describe just the hardware, it consists of the first two parts of the configuration name: ‘cpu-manufacturer’. For example, it might be one of these:

    "sparc-sun", "i386-anything", "m68k-hp",
    "m68k-sony", "m68k-sun", "mips-dec"

Function int uname (struct utsname *info)

The uname function fills in the structure pointed to by info with information about the operating system and host machine. A non-negative value indicates that the data was successfully stored. -1 as the value indicates an error. The only error possible is EFAULT, which we normally don’t mention as it is always a possibility.

536 The GNU C Library

28 System Configuration Parameters

The functions and macros listed in this chapter give information about configuration parameters of the operating system—for example, capacity limits, presence of optional POSIX features, and the default path for executable files (see Section 28.12 [String-Valued Parameters], page 551).

28.1 General Capacity Limits

The POSIX.1 and POSIX.2 standards specify a number of parameters that describe capacity limitations of the system. These limits can be fixed constants for a given operating system, or they can vary from machine to machine. For example, some limit values may be configurable by the system administrator, either at run time or by rebuilding the kernel, and this should not require recompiling application programs.

Each of the following limit parameters has a macro that is defined in ‘limits.h’ only if the system has a fixed, uniform limit for the parameter in question. If the system allows different file systems or files to have different limits, then the macro is undefined; use sysconf to find out the limit that applies at a particular time on a particular machine. See Section 28.4 [Using “codesysconf], page 541.

Each of these parameters also has another macro, with a name starting with ‘_POSIX’, which gives the lowest value that the limit is allowed to have on any POSIX system. See Section 28.5 [Minimum Values for General Capacity Limits], page 544.

Macro int ARG MAX

If defined, the unvarying maximum combined length of the argv and environ arguments that can be passed to the exec functions.

Macro int CHILD MAX

If defined, the unvarying maximum number of processes that can exist with the same real user ID at any one time. In BSD and GNU, this is controlled by the RLIMIT_NPROC resource limit; see Section 17.6 [Limiting Resource Usage], page 362.

Macro int OPEN MAX

If defined, the unvarying maximum number of files that a single process can have open simultaneously. In BSD and GNU, this is controlled by the RLIMIT_NOFILE resource limit; see Section 17.6 [Limiting Resource Usage], page 362.

Macro int STREAM MAX

If defined, the unvarying maximum number of streams that a single process can have open simultaneously. See Section 7.3 [Opening Streams], page 94.

Macro int TZNAME MAX

If defined, the unvarying maximum length of a time zone name. See Section 17.2.6 [Functions and Variables for Time Zones], page 354.

These limit macros are always defined in ‘limits.h’.

Macro int NGROUPS MAX

The maximum number of supplementary group IDs that one process can have. The value of this macro is actually a lower bound for the maximum. That is, you can count on being able to have that many supplementary group IDs, but a particular machine might let you have even more. You can use sysconf to see whether a particular machine will let you have more (see Section 28.4 [Using “codesysconf], page 541).

Macro int SSIZE MAX

The largest value that can fit in an object of type ssize_t. Effectively, this is the limit on the number of bytes that can be read or written in a single operation. This macro is defined in all POSIX systems because this limit is never configurable.

Macro int RE DUP MAX

The largest number of repetitions you are guaranteed is allowed in the construct ‘\{min,max\}’ in a regular expression. The value of this macro is actually a lower bound for the maximum. That is, you can count on being able to have that many repetitions, but a particular machine might let you have even more. You can use sysconf to see whether a particular machine will let you have more (see Section 28.4 [Using “codesysconf], page 541). And even the value that sysconf tells you is just a lower bound—larger values might work. This macro is defined in all POSIX.2 systems, because POSIX.2 says it should always be defined even if there is no specific imposed limit.

28.2 Overall System Options

POSIX defines certain system-specific options that not all POSIX systems support. Since these options are provided in the kernel, not in the library, simply using the GNU C library does not guarantee any of these features is supported; it depends on the system you are using.

You can test for the availability of a given option using the macros in this section, together with the function sysconf. The macros are defined only if you include ‘unistd.h’.

For the following macros, if the macro is defined in ‘unistd.h’, then the option is supported. Otherwise, the option may or may not be supported; use sysconf to find out. See Section 28.4 [Using “codesysconf], page 541.

Macro int POSIX JOB CONTROL

If this symbol is defined, it indicates that the system supports job control. Otherwise, the implementation behaves as if all processes within a session belong to a single process group. See Chapter 24 [Job Control], page 479.

Macro int POSIX SAVED IDS

If this symbol is defined, it indicates that the system remembers the effective user and group IDs of a process before it executes an executable file with the set-user-ID or set-group-ID bits set, and that explicitly changing the effective user or group IDs back to these values is permitted. If this option is not defined, then if a nonprivileged process changes its effective user or group ID to the real user or group ID of the process, it can’t change it back again. See Section 26.8 [Enabling and Disabling Setuid Access], page 516.

For the following macros, if the macro is defined in ‘unistd.h’, then its value indicates whether the option is supported. A value of -1 means no, and any other value means yes. If the macro is not defined, then the option may or may not be supported; use sysconf to find out. See Section 28.4 [Using “codesysconf], page 541.

Macro int POSIX2 C DEV

If this symbol is defined, it indicates that the system has the POSIX.2 C compiler command, c89. The GNU C library always defines this as 1, on the assumption that you would not have installed it if you didn’t have a C compiler.

Macro int POSIX2 FORT DEV

If this symbol is defined, it indicates that the system has the POSIX.2 Fortran compiler command, fort77. The GNU C library never defines this, because we don’t know what the system has.

Macro int POSIX2 FORT RUN

If this symbol is defined, it indicates that the system has the POSIX.2 asa command to interpret Fortran carriage control. The GNU C library never defines this, because we don’t know what the system has.

Macro int POSIX2 LOCALEDEF

If this symbol is defined, it indicates that the system has the POSIX.2 localedef command. The GNU C library never defines this, because we don’t know what the system has.

Macro int POSIX2 SW DEV

If this symbol is defined, it indicates that the system has the POSIX.2 commands ar, make, and strip. The GNU C library always defines this as 1, on the assumption that you had to have ar and make to install the library, and it’s unlikely that strip would be absent when those are present.

28.3 Which Version of POSIX is Supported

Macro long int POSIX VERSION

This constant represents the version of the POSIX.1 standard to which the implementation conforms. For an implementation conforming to the 1990 POSIX.1 standard, the value is the integer 199009L. _POSIX_VERSION is always defined (in ‘unistd.h’) in any POSIX system. Usage Note: Don’t try to test whether the system supports POSIX by including ‘unistd.h’ and then checking whether _POSIX_VERSION is defined. On a non-POSIX system, this will probably fail because there is no ‘unistd.h’. We do not know of any way you can reliably test at compilation time whether your target system supports POSIX or whether ‘unistd.h’ exists. The GNU C compiler predefines the symbol __POSIX__if the target system is a POSIX system. Provided you do not use any other compilers on POSIX systems, testing defined (__POSIX__) will reliably detect such systems.

Macro long int POSIX2 C VERSION

This constant represents the version of the POSIX.2 standard which the library and system kernel support. We don’t know what value this will be for the first version of the POSIX.2 standard, because the value is based on the year and month in which the standard is officially adopted. The value of this symbol says nothing about the utilities installed on the system. Usage Note: You can use this macro to tell whether a POSIX.1 system library supports POSIX.2 as well. Any POSIX.1 system contains ‘unistd.h’, so include that file and then test defined (_POSIX2_C_

VERSION).

28.4 Using sysconf

When your system has configurable system limits, you can use the sysconf function to find out the value that applies to any particular machine. The function and the associated parameter constants are declared in the header file ‘unistd.h’.

28.4.1 Definition of sysconf

Function long int sysconf (int parameter)

This function is used to inquire about runtime system parameters. The parameter argument should be one of the ‘_SC_’ symbols listed below. The normal return value from sysconf is the value you requested. A value of -1 is returned both if the implementation does not impose a limit, and in case of an error. The following errno error conditions are defined for this function:

EINVAL The value of the parameter is invalid.

28.4.2 Constants for sysconf Parameters

Here are the symbolic constants for use as the parameter argument to sysconf. The values are all integer constants (more specifically, enumeration type values).

_SC_ARG_MAX

Inquire about the parameter corresponding to ARG_MAX.

_SC_CHILD_MAX

Inquire about the parameter corresponding to CHILD_MAX.

_SC_OPEN_MAX

Inquire about the parameter corresponding to OPEN_MAX.

_SC_STREAM_MAX

Inquire about the parameter corresponding to STREAM_MAX.

_SC_TZNAME_MAX

Inquire about the parameter corresponding to TZNAME_MAX.

_SC_NGROUPS_MAX

Inquire about the parameter corresponding to NGROUPS_MAX.

_SC_JOB_CONTROL

Inquire about the parameter corresponding to _POSIX_JOB_

          CONTROL.

_SC_SAVED_IDS

Inquire about the parameter corresponding to _POSIX_SAVED_

IDS.
_SC_VERSION

Inquire about the parameter corresponding to _POSIX_VERSION.

_SC_CLK_TCK

Inquire about the parameter corresponding to CLOCKS_PER_SEC; see Section 17.1.1 [Basic CPU Time Inquiry], page 339.

_SC_2_C_DEV

Inquire about whether the system has the POSIX.2 C compiler command, c89.

_SC_2_FORT_DEV

Inquire about whether the system has the POSIX.2 Fortran compiler command, fort77.

_SC_2_FORT_RUN

Inquire about whether the system has the POSIX.2 asa command to interpret Fortran carriage control.

_SC_2_LOCALEDEF

Inquire about whether the system has the POSIX.2 localedef command.

_SC_2_SW_DEV

Inquire about whether the system has the POSIX.2 commands

ar, make, and strip.
_SC_BC_BASE_MAX

Inquire about the maximum value of obase in the bc utility.

_SC_BC_DIM_MAX

Inquire about the maximum size of an array in the bc utility.

_SC_BC_SCALE_MAX

Inquire about the maximum value of scale in the bc utility.

_SC_BC_STRING_MAX

Inquire about the maximum size of a string constant in the bc utility.

_SC_COLL_WEIGHTS_MAX

Inquire about the maximum number of weights that can necessarily be used in defining the collating sequence for a locale.

_SC_EXPR_NEST_MAX

Inquire about the maximum number of expressions nested within parentheses when using the expr utility.

_SC_LINE_MAX

Inquire about the maximum size of a text line that the POSIX.2 text utilities can handle.

_SC_EQUIV_CLASS_MAX

Inquire about the maximum number of weights that can be assigned to an entry of the LC_COLLATE category ‘order’ keyword in a locale definition. The GNU C library does not presently support locale definitions.

_SC_VERSION

Inquire about the version number of POSIX.1 that the library and kernel support.

_SC_2_VERSION

Inquire about the version number of POSIX.2 that the system utilities support.

_SC_PAGESIZE

Inquire about the virtual memory page size of the machine. getpagesize returns the same value.

28.4.3 Examples of sysconf

We recommend that you first test for a macro definition for the parameter you are interested in, and call sysconf only if the macro is not defined. For example, here is how to test whether job control is supported:

int
have_job_control (void)
{
#ifdef _POSIX_JOB_CONTROL
return 1;
#else
int value = sysconf (_SC_JOB_CONTROL);
if (value < 0)

/* If the system is that badly wedged,

there’s no use trying to go on. */

fatal (strerror (errno));
return value;
#endif
}

Here is how to get the value of a numeric limit:

int
get_child_max ()
{
#ifdef CHILD_MAX

return CHILD_MAX;
#else
int value = sysconf (_SC_CHILD_MAX);
if (value < 0)
fatal (strerror (errno));
return value;
#endif
}

28.5 Minimum Values for General Capacity Limits

Here are the names for the POSIX minimum upper bounds for the system limit parameters. The significance of these values is that you can safely push to these limits without checking whether the particular system you are using can go that far.

_POSIX_ARG_MAX

The value of this macro is the most restrictive limit permitted by POSIX for the maximum combined length of the argv and environ arguments that can be passed to the exec functions. Its value is 4096.

_POSIX_CHILD_MAX

The value of this macro is the most restrictive limit permitted by POSIX for the maximum number of simultaneous processes per real user ID. Its value is 6.

_POSIX_NGROUPS_MAX

The value of this macro is the most restrictive limit permitted by POSIX for the maximum number of supplementary group IDs per process. Its value is 0.

_POSIX_OPEN_MAX

The value of this macro is the most restrictive limit permitted by POSIX for the maximum number of files that a single process can have open simultaneously. Its value is 16.

_POSIX_SSIZE_MAX

The value of this macro is the most restrictive limit permitted by POSIX for the maximum value that can be stored in an object of type ssize_t. Its value is 32767.

_POSIX_STREAM_MAX

The value of this macro is the most restrictive limit permitted by POSIX for the maximum number of streams that a single process can have open simultaneously. Its value is 8.

_POSIX_TZNAME_MAX

The value of this macro is the most restrictive limit permitted by POSIX for the maximum length of a time zone name. Its value is 3.

_POSIX2_RE_DUP_MAX

The value of this macro is the most restrictive limit permitted by POSIX for the numbers used in the ‘\{min,max\}’ construct in a regular expression. Its value is 255.

28.6 Limits on File System Capacity

The POSIX.1 standard specifies a number of parameters that describe the limitations of the file system. It’s possible for the system to have a fixed, uniform limit for a parameter, but this isn’t the usual case. On most systems, it’s possible for different file systems (and, for some parameters, even different files) to have different maximum limits. For example, this is very likely if you use NFS to mount some of the file systems from other machines.

Each of the following macros is defined in ‘limits.h’ only if the system has a fixed, uniform limit for the parameter in question. If the system allows different file systems or files to have different limits, then the macro is undefined; use pathconf or fpathconf to find out the limit that applies to a particular file. See Section 28.9 [Using “codepathconf], page 548.

Each parameter also has another macro, with a name starting with ‘_POSIX’, which gives the lowest value that the limit is allowed to have on any POSIX system. See Section 28.8 [Minimum Values for File System Limits], page 547.

Macro int LINK MAX

The uniform system limit (if any) for the number of names for a given file. See Section 9.3 [Hard Links], page 189.

Macro int MAX CANON

The uniform system limit (if any) for the amount of text in a line of input when input editing is enabled. See Section 12.3 [Two Styles of Input: Canonical or Not], page 272.

Macro int MAX INPUT

The uniform system limit (if any) for the total number of characters typed ahead as input. See Section 12.2 [I/O Queues], page 271.

Macro int NAME MAX

The uniform system limit (if any) for the length of a file name component.

Macro int PATH MAX

The uniform system limit (if any) for the length of an entire file name (that is, the argument given to system calls such as open).

Macro int PIPE BUF

The uniform system limit (if any) for the number of bytes that can be written atomically to a pipe. If multiple processes are writing to the same pipe simultaneously, output from different processes might be interleaved in chunks of this size. See Chapter 10 [Pipes and FIFOs], page 215.

These are alternative macro names for some of the same information.

Macro int MAXNAMLEN

This is the BSD name for NAME_MAX. It is defined in ‘dirent.h’.

Macro int FILENAME MAX

The value of this macro is an integer constant expression that represents the maximum length of a file name string. It is defined in

‘stdio.h’.

Unlike PATH_MAX, this macro is defined even if there is no actual limit imposed. In such a case, its value is typically a very large number. This is always the case on the GNU system. Usage Note: Don’t use FILENAME_MAX as the size of an array in which to store a file name! You can’t possibly make an array that big! Use dynamic allocation (see Chapter 3 [Memory Allocation], page 31) instead.

28.7 Optional Features in File Support

POSIX defines certain system-specific options in the system calls for operating on files. Some systems support these options and others do not. Since these options are provided in the kernel, not in the library, simply using the GNU C library does not guarantee any of these features is supported; it depends on the system you are using. They can also vary between file systems on a single machine.

This section describes the macros you can test to determine whether a particular option is supported on your machine. If a given macro is defined in ‘unistd.h’, then its value says whether the corresponding feature is supported. (A value of -1 indicates no; any other value indicates yes.) If the macro is undefined, it means particular files may or may not support the feature.

Since all the machines that support the GNU C library also support NFS, one can never make a general statement about whether all file systems

support the _POSIX_CHOWN_RESTRICTED and _POSIX_NO_TRUNC features. So

these names are never defined as macros in the GNU C library.

Macro int POSIX CHOWN RESTRICTED

If this option is in effect, the chown function is restricted so that the only changes permitted to nonprivileged processes is to change the group owner of a file to either be the effective group ID of the process, or one of its supplementary group IDs. See Section 9.8.4 [File Owner], page 200.

Macro int POSIX NO TRUNC

If this option is in effect, file name components longer than NAME_MAX generate an ENAMETOOLONG error. Otherwise, file name components that are too long are silently truncated.

Macro unsigned char POSIX VDISABLE

This option is only meaningful for files that are terminal devices. If it is enabled, then handling for special control characters can be disabled individually. See Section 12.4.9 [Special Characters], page 286.

If one of these macros is undefined, that means that the option might be in effect for some files and not for others. To inquire about a particular file, call pathconf or fpathconf. See Section 28.9 [Using “codepathconf], page 548.

28.8 Minimum Values for File System Limits

Here are the names for the POSIX minimum upper bounds for some of the above parameters. The significance of these values is that you can safely push to these limits without checking whether the particular system you are using can go that far.

_POSIX_LINK_MAX

The most restrictive limit permitted by POSIX for the maximum value of a file’s link count. The value of this constant is 8; thus, you can always make up to eight names for a file without running into a system limit.

_POSIX_MAX_CANON

The most restrictive limit permitted by POSIX for the maximum number of bytes in a canonical input line from a terminal device. The value of this constant is 255.

_POSIX_MAX_INPUT

The most restrictive limit permitted by POSIX for the maximum number of bytes in a terminal device input queue (or typeahead buffer). See Section 12.4.4 [Input Modes], page 277. The value of this constant is 255.

_POSIX_NAME_MAX

The most restrictive limit permitted by POSIX for the maximum number of bytes in a file name component. The value of this constant is 14.

_POSIX_PATH_MAX

The most restrictive limit permitted by POSIX for the maximum number of bytes in a file name. The value of this constant is 255.

_POSIX_PIPE_BUF

The most restrictive limit permitted by POSIX for the maximum number of bytes that can be written atomically to a pipe. The value of this constant is 512.

28.9 Using pathconf

When your machine allows different files to have different values for a file system parameter, you can use the functions in this section to find out the value that applies to any particular file.

These functions and the associated constants for the parameter argument are declared in the header file ‘unistd.h’.

Function long int pathconf (const char *filename, int

parameter) This function is used to inquire about the limits that apply to the file named filename. The parameter argument should be one of the ‘_PC_’ constants listed below. The normal return value from pathconf is the value you requested. A value of -1 is returned both if the implementation does not impose a limit, and in case of an error. In the former case, errno is not set, while in the latter case, errno is set to indicate the cause of the problem. So the only way to use this function robustly is to store 0 into errno just before calling it. Besides the usual file name errors (see Section 6.2.3 [File Name Errors], page 89), the following error condition is defined for this function:

EINVAL The value of parameter is invalid, or the implementation doesn’t support the parameter for the specific file.

Function long int fpathconf (int filedes, int parameter)

This is just like pathconf except that an open file descriptor is used to specify the file for which information is requested, instead of a file name. The following errno error conditions are defined for this function:

EBADF The filedes argument is not a valid file descriptor.

EINVAL The value of parameter is invalid, or the implementation doesn’t support the parameter for the specific file.

Here are the symbolic constants that you can use as the parameter argument to pathconf and fpathconf. The values are all integer constants.

_PC_LINK_MAX

Inquire about the value of LINK_MAX.

_PC_MAX_CANON

Inquire about the value of MAX_CANON.

_PC_MAX_INPUT

Inquire about the value of MAX_INPUT.

_PC_NAME_MAX

Inquire about the value of NAME_MAX.

_PC_PATH_MAX

Inquire about the value of PATH_MAX.

_PC_PIPE_BUF

Inquire about the value of PIPE_BUF.

_PC_CHOWN_RESTRICTED
          Inquire about the value of _POSIX_CHOWN_RESTRICTED.
_PC_NO_TRUNC

Inquire about the value of _POSIX_NO_TRUNC.

_PC_VDISABLE

Inquire about the value of _POSIX_VDISABLE.

28.10 Utility Program Capacity Limits

The POSIX.2 standard specifies certain system limits that you can access through sysconf that apply to utility behavior rather than the behavior of the library or the operating system.

The GNU C library defines macros for these limits, and sysconf returns values for them if you ask; but these values convey no meaningful information. They are simply the smallest values that POSIX.2 permits.

Macro int BC BASE MAX

The largest value of obase that the bc utility is guaranteed to support.

Macro int BC SCALE MAX

The largest value of scale that the bc utility is guaranteed to support.

Macro int BC DIM MAX

The largest number of elements in one array that the bc utility is guaranteed to support.

Macro int BC STRING MAX

The largest number of characters in one string constant that the bc utility is guaranteed to support.

Macro int BC DIM MAX

The largest number of elements in one array that the bc utility is guaranteed to support.

Macro int COLL WEIGHTS MAX

The largest number of weights that can necessarily be used in defining the collating sequence for a locale.

Macro int EXPR NEST MAX

The maximum number of expressions that can be nested within parenthesis by the expr utility.

Macro int LINE MAX

The largest text line that the text-oriented POSIX.2 utilities can support. (If you are using the GNU versions of these utilities, then there is no actual limit except that imposed by the available virtual memory, but there is no way that the library can tell you this.)

Macro int EQUIV CLASS MAX

The maximum number of weights that can be assigned to an entry of the LC_COLLATE category ‘order’ keyword in a locale definition. The GNU C library does not presently support locale definitions.

28.11 Minimum Values for Utility Limits

_POSIX2_BC_BASE_MAX

The most restrictive limit permitted by POSIX.2 for the maximum value of obase in the bc utility. Its value is 99.

_POSIX2_BC_DIM_MAX

The most restrictive limit permitted by POSIX.2 for the maximum size of an array in the bc utility. Its value is 2048.

_POSIX2_BC_SCALE_MAX

The most restrictive limit permitted by POSIX.2 for the maximum value of scale in the bc utility. Its value is 99.

_POSIX2_BC_STRING_MAX

The most restrictive limit permitted by POSIX.2 for the maximum size of a string constant in the bc utility. Its value is

_POSIX2_COLL_WEIGHTS_MAX

The most restrictive limit permitted by POSIX.2 for the maximum number of weights that can necessarily be used in defining the collating sequence for a locale. Its value is 2.

_POSIX2_EXPR_NEST_MAX

The most restrictive limit permitted by POSIX.2 for the maximum number of expressions nested within parenthesis when using the expr utility. Its value is 32.

_POSIX2_LINE_MAX

The most restrictive limit permitted by POSIX.2 for the maximum size of a text line that the text utilities can handle. Its value is 2048.

_POSIX2_EQUIV_CLASS_MAX

The most restrictive limit permitted by POSIX.2 for the maximum number of weights that can be assigned to an entry of the LC_COLLATE category ‘order’ keyword in a locale definition. Its value is 2. The GNU C library does not presently support locale definitions.

28.12 String-Valued Parameters

POSIX.2 defines a way to get string-valued parameters from the operating system with the function confstr:

Function size_t confstr (int parameter, char *buf, size_t len)

This function reads the value of a string-valued system parameter, storing the string into len bytes of memory space starting at buf. The parameter argument should be one of the ‘_CS_’ symbols listed below. The normal return value from confstr is the length of the string value that you asked for. If you supply a null pointer for buf, then confstr does not try to store the string; it just returns its length. A value of 0 indicates an error. If the string you asked for is too long for the buffer (that is, longer than len - 1), then confstr stores just that much (leaving room for the terminating null character). You can tell that this has happened because confstr returns a value greater than or equal to len. The following errno error conditions are defined for this function:

EINVAL The value of the parameter is invalid.

Currently there is just one parameter you can read with confstr:

_CS_PATH This parameter’s value is the recommended default path for searching for executable files. This is the path that a user has by default just after logging in.

The way to use confstr without any arbitrary limit on string size is to call it twice: first call it to get the length, allocate the buffer accordingly, and then call confstr again to fill the buffer, like this:

     char *
     get_default_path (void)
     {
       size_t len = confstr (_CS_PATH, NULL, 0);
       char *buffer = (char *) xmalloc (len);
       if (confstr (_CS_PATH, buf, len + 1) == 0)
         {
           free (buffer);
           return NULL;
         }
       return buffer;
     }

Appendix A C Language Facilities in the Library

Some of the facilities implemented by the C library really should be thought of as parts of the C language itself. These facilities ought to be documented in the C Language Manual, not in the library manual; but since we don’t have the language manual yet, and documentation for these features has been written, we are publishing it here.

A.1 Explicitly Checking Internal Consistency

When you’re writing a program, it’s often a good idea to put in checks at strategic places for “impossible” errors or violations of basic assumptions. These kinds of checks are helpful in debugging problems with the interfaces between different parts of the program, for example.

The assert macro, defined in the header file ‘assert.h’, provides a convenient way to abort the program while printing a message about where in the program the error was detected.

Once you think your program is debugged, you can disable the error checks performed by the assert macro by recompiling with the macro NDEBUG defined. This means you don’t actually have to change the program source code to disable these checks.

But disabling these consistency checks is undesirable unless they make the program significantly slower. All else being equal, more error checking is good no matter who is running the program. A wise user would rather have a program crash, visibly, than have it return nonsense without indicating anything might be wrong.

Macro void assert (int expression)

Verify the programmer’s belief that expression should be nonzero at this point in the program. If NDEBUG is not defined, assert tests the value of expression. If it is false (zero), assert aborts the program (see Section 22.3.4 [Aborting a Program], page 465) after printing a message of the form:

`file':linenum: function: Assertion `expression' failed.

on the standard error stream stderr (see Section 7.2 [Standard Streams], page 93). The filename and line number are taken from the C preprocessor macros __FILE__and __LINE__and specify where the call to assert was written. When using the GNU C compiler, the name of the function which calls assert is taken from the builtin variable __PRETTY_FUNCTION__; with older compilers, the function name and following colon are omitted.

If the preprocessor macro NDEBUG is defined before ‘assert.h’ is included, the assert macro is defined to do absolutely nothing. Warning: Even the argument expression expression is not evaluated if NDEBUG is in effect. So never use assert with arguments that involve side effects. For example, assert (++i > 0); is a bad idea, because i will not be incremented if NDEBUG is defined.

Sometimes the “impossible” condition you want to check for is an error return from an operating system function. Then it is useful to display not only where the program crashes, but also what error was returned. The assert_perror macro makes this easy.

Macro void assert perror (int errnum)

Similar to assert, but verifies that errnum is zero. If NDEBUG is defined, assert_perror tests the value of errnum. If it is nonzero, assert_perror aborts the program after a printing a message of the form:

`file':linenum: function: error text

on the standard error stream. The file name, line number, and function name are as for assert. The error text is the result of strerror (errnum). See Section 2.3 [Error Messages], page 27. Like assert, if NDEBUG is defined before ‘assert.h’ is included, the assert_perror macro does absolutely nothing. It does not evaluate the argument, so errnum should not have any side effects. It is best for errnum to be a just simple variable reference; often it will be errno. This macro is a GNU extension.

Usage note: The assert facility is designed for detecting internal inconsistency; it is not suitable for reporting invalid input or improper usage by the user of the program.

The information in the diagnostic messages printed by the assert macro is intended to help you, the programmer, track down the cause of a bug, but is not really useful for telling a user of your program why his or her input was invalid or why a command could not be carried out. So you can’t use assert or assert_perror to print the error messages for these eventualities.

What’s more, your program should not abort when given invalid input, as assert would do—it should exit with nonzero status (see Section 22.3.2 [Exit Status], page 463) after printing its error messages, or perhaps read another command or move on to the next input file.

See Section 2.3 [Error Messages], page 27, for information on printing error messages for problems that do not represent bugs in the program.

A.2 Variadic Functions

ISO C defines a syntax for declaring a function to take a variable number or type of arguments. (Such functions are referred to as varargs functions or variadic functions.) However, the language itself provides no mechanism for such functions to access their non-required arguments; instead, you use the variable arguments macros defined in ‘stdarg.h’.

This section describes how to declare variadic functions, how to write them, and how to call them properly.

Compatibility Note: Many older C dialects provide a similar, but incompatible, mechanism for defining functions with variable numbers of arguments, using ‘varargs.h’.

A.2.1 Why Variadic Functions are Used

Ordinary C functions take a fixed number of arguments. When you define a function, you specify the data type for each argument. Every call to the function should supply the expected number of arguments, with types that can be converted to the specified ones. Thus, if the function ‘foo’ is declared with int foo (int, char *); then you must call it with two arguments, a number (any kind will do) and a string pointer.

But some functions perform operations that can meaningfully accept an unlimited number of arguments.

In some cases a function can handle any number of values by operating on all of them as a block. For example, consider a function that allocates a one-dimensional array with malloc to hold a specified set of values. This operation makes sense for any number of values, as long as the length of the array corresponds to that number. Without facilities for variable arguments, you would have to define a separate function for each possible array size.

The library function printf (see Section 7.10 [Formatted Output], page 104) is an example of another class of function where variable arguments are useful. This function prints its arguments (which can vary in type as well as number) under the control of a format template string.

These are good reasons to define a variadic function which can handle as many arguments as the caller chooses to pass.

Some functions such as open take a fixed set of arguments, but occasionally ignore the last few. Strict adherence to ISO C requires these functions to be defined as variadic; in practice, however, the GNU C compiler and most other C compilers let you define such a function to take a fixed set of arguments—the most it can ever use—and then only declare the function as variadic (or not declare its arguments at all!).

A.2.2 How Variadic Functions are Defined and Used

Defining and using a variadic function involves three steps:

• Define the function as variadic, using an ellipsis (‘…’) in the argument list, and using special macros to access the variable arguments. See Section A.2.2.2 [Receiving the Argument Values], page 556.

• Declare the function as variadic, using a prototype with an ellipsis (‘…’), in all the files which call it. See Section A.2.2.1 [Syntax for Variable Arguments], page 556.

• Call the function by writing the fixed arguments followed by the additional variable arguments. See Section A.2.2.4 [Calling Variadic Functions], page 558.

A.2.2.1 Syntax for Variable Arguments

A function that accepts a variable number of arguments must be declared with a prototype that says so. You write the fixed arguments as usual, and then tack on ‘…’ to indicate the possibility of additional arguments. The syntax of ISO C requires at least one fixed argument before the ‘…’. For example,

     int
     func (const char *a, int b, . . .)
     {

…

     }

outlines a definition of a function func which returns an int and takes two required arguments, a const char * and an int. These are followed by any number of anonymous arguments.

Portability note: For some C compilers, the last required argument must not be declared register in the function definition. Furthermore, this argument’s type must be self-promoting: that is, the default promotions must not change its type. This rules out array and function types, as well as float, char (whether signed or not) and short int (whether signed or not). This is actually an ISO C requirement.

A.2.2.2 Receiving the Argument Values

Ordinary fixed arguments have individual names, and you can use these names to access their values. But optional arguments have no names— nothing but ‘…’. How can you access them?

The only way to access them is sequentially, in the order they were written, and you must use special macros from ‘stdarg.h’ in the following three step process:

1. You initialize an argument pointer variable of type va_list using va_ start. The argument pointer when initialized points to the first optional argument.

2. You access the optional arguments by successive calls to va_arg. The first call to va_arg gives you the first optional argument, the next call gives you the second, and so on. You can stop at any time if you wish to ignore any remaining optional arguments. It is perfectly all right for a function to access fewer arguments than were supplied in the call, but you will get garbage values if you try to access too many arguments.

3. You indicate that you are finished with the argument pointer variable by calling va_end. (In practice, with most C compilers, calling va_end does nothing and you do not really need to call it. This is always true in the GNU C compiler. But you might as well call va_end just in case your program is someday compiled with a peculiar compiler.)

See Section A.2.2.5 [Argument Access Macros], page 559, for the full

definitions of va_start, va_arg and va_end.

Steps 1 and 3 must be performed in the function that accepts the optional arguments. However, you can pass the va_list variable as an argument to another function and perform all or part of step 2 there.

You can perform the entire sequence of the three steps multiple times within a single function invocation. If you want to ignore the optional arguments, you can do these steps zero times.

You can have more than one argument pointer variable if you like. You can initialize each variable with va_start when you wish, and then you can fetch arguments with each argument pointer as you wish. Each argument pointer variable will sequence through the same set of argument values, but at its own pace.

Portability note: With some compilers, once you pass an argument pointer value to a subroutine, you must not keep using the same argument pointer value after that subroutine returns. For full portability, you should just pass it to va_end. This is actually an ISO C requirement, but most ANSI C compilers work happily regardless.

A.2.2.3 How Many Arguments Were Supplied

There is no general way for a function to determine the number and type of the optional arguments it was called with. So whoever designs the function typically designs a convention for the caller to tell it how many arguments it has, and what kind. It is up to you to define an appropriate calling convention for each variadic function, and write all calls accordingly.

One kind of calling convention is to pass the number of optional arguments as one of the fixed arguments. This convention works provided all of the optional arguments are of the same type.

A similar alternative is to have one of the required arguments be a bit mask, with a bit for each possible purpose for which an optional argument might be supplied. You would test the bits in a predefined sequence; if the bit is set, fetch the value of the next argument, otherwise use a default value.

A required argument can be used as a pattern to specify both the number and types of the optional arguments. The format string argument to printf is one example of this (see Section 7.10.7 [Formatted Output Functions], page 113).

Another possibility is to pass an “end marker” value as the last optional argument. For example, for a function that manipulates an arbitrary number of pointer arguments, a null pointer might indicate the end of the argument list. (This assumes that a null pointer isn’t otherwise meaningful to the function.) The execl function works in just this way; see Section 23.5 [Executing a File], page 470.

A.2.2.4 Calling Variadic Functions

You don’t have to write anything special when you call a variadic function. Just write the arguments (required arguments, followed by optional ones) inside parentheses, separated by commas, as usual. But you should prepare by declaring the function with a prototype, and you must know how the argument values are converted.

In principle, functions that are defined to be variadic must also be declared to be variadic using a function prototype whenever you call them. (See Section A.2.2.1 [Syntax for Variable Arguments], page 556, for how.) This is because some C compilers use a different calling convention to pass the same set of argument values to a function depending on whether that function takes variable arguments or fixed arguments.

In practice, the GNU C compiler always passes a given set of argument types in the same way regardless of whether they are optional or required. So, as long as the argument types are self-promoting, you can safely omit declaring them. Usually it is a good idea to declare the argument types for variadic functions, and indeed for all functions. But there are a few functions which it is extremely convenient not to have to declare as variadic—for example, open and printf.

Since the prototype doesn’t specify types for optional arguments, in a call to a variadic function the default argument promotions are performed on the optional argument values. This means the objects of type char or short int (whether signed or not) are promoted to either int or unsigned int, as appropriate; and that objects of type float are promoted to type double.

So, if the caller passes a char as an optional argument, it is promoted to an int, and the function should get it with va_arg (ap, int).

Conversion of the required arguments is controlled by the function prototype in the usual way: the argument expression is converted to the declared argument type as if it were being assigned to a variable of that type.

A.2.2.5 Argument Access Macros

Here are descriptions of the macros used to retrieve variable arguments. These macros are defined in the header file ‘stdarg.h’.

Data Type va list

The type va_list is used for argument pointer variables.

Macro void va start (va_list ap, last-required)

This macro initializes the argument pointer variable ap to point to the first of the optional arguments of the current function; last-required must be the last required argument to the function. See Section A.2.3.1 [Old-Style Variadic Functions], page 560, for an alternate definition of va_start found in the header file ‘varargs.h’.

Macro type va arg (va_list ap, type)

The va_arg macro returns the value of the next optional argument, and modifies the value of ap to point to the subsequent argument. Thus, successive uses of va_arg return successive optional arguments. The type of the value returned by va_arg is type as specified in the call. type must be a self-promoting type (not char or short int or float) that matches the type of the actual argument.

Macro void va end (va_list ap)

This ends the use of ap. After a va_end call, further va_arg calls with the same ap may not work. You should invoke va_end before returning from the function in which va_start was invoked with the same ap argument. In the GNU C library, va_end does nothing, and you need not ever use it except for reasons of portability.

A.2.3 Example of a Variadic Function

Here is a complete sample function that accepts a variable number of arguments. The first argument to the function is the count of remaining arguments, which are added up and the result returned. While trivial, this function is sufficient to illustrate how to use the variable arguments facility.

  #include <stdarg.h>
  #include <stdio.h>

     int
     add_em_up (int count,...)
     {
       va_list ap;
       int i, sum;

va_start (ap, count); /* Initialize the argument list. */

       sum = 0;
       for (i = 0; i < count; i++)
         sum += va_arg (ap, int);    /* Get the next argument value. */
       va_end (ap);                  /* Clean up. */
       return sum;
     }
     int
     main (void)
     {

/* This call prints 16. */

       printf ("%d\n", add_em_up (3, 5, 5, 6));

/* This call prints 55. */

       printf ("%d\n", add_em_up (10, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10));
       return 0;
     }

A.2.3.1 Old-Style Variadic Functions

Before ISO C, programmers used a slightly different facility for writing variadic functions. The GNU C compiler still supports it; currently, it is more portable than the ISO C facility, since support for ISO C is still not universal. The header file which defines the old-fashioned variadic facility is

called ‘varargs.h’.

Using ‘varargs.h’ is almost the same as using ‘stdarg.h’. There is no difference in how you call a variadic function; See Section A.2.2.4 [Calling Variadic Functions], page 558. The only difference is in how you define them. First of all, you must use old-style non-prototype syntax, like this:

     tree
     build (va_alist)
          va_dcl
     {

Secondly, you must give va_start just one argument, like this:

       va_list p;
       va_start (p);

These are the special macros used for defining old-style variadic functions:

Macro va alist

This macro stands for the argument name list required in a variadic function.

Macro va dcl

This macro declares the implicit argument or arguments for a variadic function.

Macro void va start (va_list ap)

This macro, as defined in ‘varargs.h’, initializes the argument pointer variable ap to point to the first argument of the current function.

The other argument macros, va_arg and va_end, are the same in ‘varargs.h’ as in ‘stdarg.h’; see Section A.2.2.5 [Argument Access Macros], page 559 for details.

It does not work to include both ‘varargs.h’ and ‘stdarg.h’ in the same compilation; they define va_start in conflicting ways.

A.3 Null Pointer Constant

The null pointer constant is guaranteed not to point to any real object. You can assign it to any pointer variable since it has type void *. The preferred way to write a null pointer constant is with NULL.

Macro void * NULL

This is a null pointer constant.

You can also use 0 or (void *)0 as a null pointer constant, but using NULL is cleaner because it makes the purpose of the constant more evident.

If you use the null pointer constant as a function argument, then for complete portability you should make sure that the function has a prototype declaration. Otherwise, if the target machine has two different pointer representations, the compiler won’t know which representation to use for that argument. You can avoid the problem by explicitly casting the constant to the proper pointer type, but we recommend instead adding a prototype for the function you are calling.

A.4 Important Data Types

The result of subtracting two pointers in C is always an integer, but the precise data type varies from C compiler to C compiler. Likewise, the data type of the result of sizeof also varies between compilers. ISO defines standard aliases for these two types, so you can refer to them in a portable fashion. They are defined in the header file ‘stddef.h’.

Data Type ptrdiff t

This is the signed integer type of the result of subtracting two pointers. For example, with the declaration char *p1, *p2;, the expression p2 - p1 is of type ptrdiff_t. This will probably be one of the standard signed integer types (short int, int or long int), but might be a nonstandard type that exists only for this purpose.

Data Type size t

This is an unsigned integer type used to represent the sizes of objects. The result of the sizeof operator is of this type, and functions such as malloc (see Section 3.3 [Unconstrained Allocation], page 32) and memcpy (see Section 5.4 [Copying and Concatenation], page 67) accept arguments of this type to specify object sizes. Usage Note: size_t is the preferred way to declare any arguments or variables that hold the size of an object.

In the GNU system size_t is equivalent to either unsigned int or unsigned long int. These types have identical properties on the GNU system, and for most purposes, you can use them interchangeably. However, they are distinct as data types, which makes a difference in certain contexts.

For example, when you specify the type of a function argument in a function prototype, it makes a difference which one you use. If the system header files declare malloc with an argument of type size_t and you declare malloc with an argument of type unsigned int, you will get a compilation error if size_t happens to be unsigned long int on your system. To avoid any possibility of error, when a function argument or value is supposed to have type size_t, never declare its type in any other way.

Compatibility Note: Implementations of C before the advent of ISO C generally used unsigned int for representing object sizes and int for pointer subtraction results. They did not necessarily define either size_t or ptrdiff_t. Unix systems did define size_t, in ‘sys/types.h’, but the definition was usually a signed type.

A.5 Data Type Measurements

Most of the time, if you choose the proper C data type for each object in your program, you need not be concerned with just how it is represented or how many bits it uses. When you do need such information, the C language itself does not provide a way to get it. The header files ‘limits.h’ and ‘float.h’ contain macros which give you this information in full detail.

A.5.1 Computing the Width of an Integer Data Type

The most common reason that a program needs to know how many bits are in an integer type is for using an array of long int as a bit vector. You can access the bit at index n with

     vector[n / LONGBITS] & (1 << (n % LONGBITS))

provided you define LONGBITS as the number of bits in a long int.

There is no operator in the C language that can give you the number of bits in an integer data type. But you can compute it from the macro CHAR_BIT, defined in the header file ‘limits.h’.

CHAR_BIT This is the number of bits in a char—eight, on most systems. The value has type int. You can compute the number of bits in any data type type like this:

                 sizeof (type) * CHAR_BIT

A.5.2 Range of an Integer Type

Suppose you need to store an integer value which can range from zero to one million. Which is the smallest type you can use? There is no general rule; it depends on the C compiler and target machine. You can use the ‘MIN’ and ‘MAX’ macros in ‘limits.h’ to determine which type will work.

Each signed integer type has a pair of macros which give the smallest and largest values that it can hold. Each unsigned integer type has one such macro, for the maximum value; the minimum value is, of course, zero.

The values of these macros are all integer constant expressions. The ‘MAX’ and ‘MIN’ macros for char and short int types have values of type int. The ‘MAX’ and ‘MIN’ macros for the other types have values of the same type described by the macro—thus, ULONG_MAX has type unsigned long int.

SCHAR_MIN

This is the minimum value that can be represented by a

          signed char.
SCHAR_MAX
UCHAR_MAX

These are the maximum values that can be represented by a

          signed char and unsigned char, respectively.
CHAR_MIN

This is the minimum value that can be represented by a char. It’s equal to SCHAR_MIN if char is signed, or zero otherwise.

CHAR_MAX

This is the maximum value that can be represented by a char. It’s equal to SCHAR_MAX if char is signed, or UCHAR_MAX otherwise.

SHRT_MIN

This is the minimum value that can be represented by a signed short int. On most machines that the GNU C library runs on, short integers are 16-bit quantities.

SHRT_MAX
USHRT_MAX

These are the maximum values that can be represented by a

signed short int and unsigned short int, respectively.
INT_MIN

This is the minimum value that can be represented by a signed int. On most machines that the GNU C system runs on, an int is a 32-bit quantity.

INT_MAX
UINT_MAX

These are the maximum values that can be represented by, respectively, the type signed int and the type unsigned int.

LONG_MIN

This is the minimum value that can be represented by a signed long int. On most machines that the GNU C system runs on, long integers are 32-bit quantities, the same size as

int.
LONG_MAX
ULONG_MAX

These are the maximum values that can be represented by a

signed long int and unsigned long int, respectively.
LONG_LONG_MIN

This is the minimum value that can be represented by a signed long long int. On most machines that the GNU C system runs on, long long integers are 64-bit quantities.

LONG_LONG_MAX
ULONG_LONG_MAX

These are the maximum values that can be represented by

a signed long long int and unsigned long long int, respec-

tively.

WCHAR_MAX

This is the maximum value that can be represented by a wchar_ t. See Section 18.4 [Wide Character Introduction], page 371.

The header file ‘limits.h’ also defines some additional constants that parameterize various operating system and file system limits. These constants are described in Chapter 28 [System Configuration Parameters], page 537.

A.5.3 Floating Type Macros

The specific representation of floating point numbers varies from machine to machine. Because floating point numbers are represented internally as approximate quantities, algorithms for manipulating floating point data often need to take account of the precise details of the machine’s floating point representation.

Some of the functions in the C library itself need this information; for example, the algorithms for printing and reading floating point numbers (see Chapter 7 [Input/Output on Streams], page 93) and for calculating trigonometric and irrational functions (see Chapter 13 [Mathematics], page 297) use it to avoid round-off error and loss of accuracy. User programs that implement numerical analysis techniques also often need this information in order to minimize or compute error bounds.

The header file ‘float.h’ describes the format used by your machine.

A.5.3.1 Floating Point Representation Concepts

This section introduces the terminology for describing floating point representations.

You are probably already familiar with most of these concepts in terms of scientific or exponential notation for floating point numbers. For example, the number 123456.0 could be expressed in exponential notation as 1.23456e+05, a shorthand notation indicating that the mantissa 1.23456 is multiplied by the base 10 raised to power 5.

More formally, the internal representation of a floating point number can be characterized in terms of the following parameters:

• The sign is either -1 or 1.

• The base or radix for exponentiation, an integer greater than 1. This is a constant for a particular representation.

• The exponent to which the base is raised. The upper and lower bounds of the exponent value are constants for a particular representation. Sometimes, in the actual bits representing the floating point number, the exponent is biased by adding a constant to it, to make it always be represented as an unsigned quantity. This is only important if you have some reason to pick apart the bit fields making up the floating point number by hand, which is something for which the GNU library provides no support. So this is ignored in the discussion that follows.

• The mantissa or significand, an unsigned integer which is a part of each floating point number.

• The precision of the mantissa. If the base of the representation is b, then the precision is the number of base-b digits in the mantissa. This is a constant for a particular representation. Many floating point representations have an implicit hidden bit in the mantissa. This is a bit which is present virtually in the mantissa, but not stored in memory because its value is always 1 in a normalized number. The precision figure (see above) includes any hidden bits. Again, the GNU library provides no facilities for dealing with such lowlevel aspects of the representation.

The mantissa of a floating point number actually represents an implicit fraction whose denominator is the base raised to the power of the precision. Since the largest representable mantissa is one less than this denominator, the value of the fraction is always strictly less than 1. The mathematical value of a floating point number is then the product of this fraction, the sign, and the base raised to the exponent.

We say that the floating point number is normalized if the fraction is at least 1/b, where b is the base. In other words, the mantissa would be too large to fit if it were multiplied by the base. Non-normalized numbers are sometimes called denormal; they contain less precision than the representation normally can hold.

If the number is not normalized, then you can subtract 1 from the exponent while multiplying the mantissa by the base, and get another floating point number with the same value. Normalization consists of doing this repeatedly until the number is normalized. Two distinct normalized floating point numbers cannot be equal in value.

(There is an exception to this rule: if the mantissa is zero, it is considered normalized. Another exception happens on certain machines where the exponent is as small as the representation can hold. Then it is impossible to subtract 1 from the exponent, so a number may be normalized even if its fraction is less than 1/b.)

A.5.3.2 Floating Point Parameters

These macro definitions can be accessed by including the header file ‘float.h’ in your program.

Macro names starting with ‘FLT_’ refer to the float type, while names beginning with ‘DBL_’ refer to the double type and names beginning with ‘LDBL_’ refer to the long double type. (Currently GCC does not support long double as a distinct data type, so the values for the ‘LDBL_’ constants are equal to the corresponding constants for the double type.)

Of these macros, only FLT_RADIX is guaranteed to be a constant expression. The other macros listed here cannot be reliably used in places that require constant expressions, such as ‘#if’ preprocessing directives or in the dimensions of static arrays.

Although the ISO C standard specifies minimum and maximum values for most of these parameters, the GNU C implementation uses whatever values describe the floating point representation of the target machine. So in principle GNU C actually satisfies the ISO C requirements only if the target machine is suitable. In practice, all the machines currently supported are suitable.

FLT_ROUNDS

This value characterizes the rounding mode for floating point addition. The following values indicate standard rounding modes: -1 The mode is indeterminable.

0 Rounding is towards zero.

1 Rounding is to the nearest number.

2 Rounding is towards positive infinity.

3 Rounding is towards negative infinity.

Any other value represents a machine-dependent nonstandard rounding mode. On most machines, the value is 1, in accordance with the IEEE standard for floating point. Here is a table showing how certain values round for each possible value of FLT_ROUNDS, if the other aspects of the representation match the IEEE single-precision standard.

                     0      1             2             3
      1.00000003    1.0    1.0           1.00000012    1.0
      1.00000007    1.0    1.00000012    1.00000012    1.0
     -1.00000003   -1.0   -1.0          -1.0          -1.00000012
     -1.00000007   -1.0   -1.00000012   -1.0          -1.00000012
FLT_RADIX

This is the value of the base, or radix, of exponent representation. This is guaranteed to be a constant expression, unlike the other macros described in this section. The value is 2 on all machines we know of except the IBM 360 and derivatives.

FLT_MANT_DIG

This is the number of base-FLT_RADIX digits in the floating point mantissa for the float data type. The following expression yields 1.0 (even though mathematically it should not) due to the limited number of mantissa digits:

     float radix = FLT_RADIX;
     1.0f + 1.0f / radix / radix / . . . / radix

where radix appears FLT_MANT_DIG times.

DBL_MANT_DIG
LDBL_MANT_DIG

This is the number of base-FLT_RADIX digits in the floating point mantissa for the data types double and long double, respectively.

FLT_DIG

This is the number of decimal digits of precision for the float data type. Technically, if p and b are the precision and base (respectively) for the representation, then the decimal precision q is the maximum number of decimal digits such that any floating point number with q base 10 digits can be rounded to a floating point number with p base b digits and back again, without change to the q decimal digits. The value of this macro is supposed to be at least 6, to satisfy ISO C.

DBL_DIG
LDBL_DIG

These are similar to FLT_DIG, but for the data types double and long double, respectively. The values of these macros are supposed to be at least 10.

FLT_MIN_EXP

This is the smallest possible exponent value for type float. More precisely, is the minimum negative integer such that the value FLT_RADIX raised to this power minus 1 can be represented as a normalized floating point number of type float.

DBL_MIN_EXP
LDBL_MIN_EXP

These are similar to FLT_MIN_EXP, but for the data types double and long double, respectively.

FLT_MIN_10_EXP

This is the minimum negative integer such that 10 raised to this power minus 1 can be represented as a normalized floating point number of type float. This is supposed to be -37 or even less.

DBL_MIN_10_EXP
LDBL_MIN_10_EXP

These are similar to FLT_MIN_10_EXP, but for the data types double and long double, respectively.

FLT_MAX_EXP

This is the largest possible exponent value for type float. More precisely, this is the maximum positive integer such that value FLT_RADIX raised to this power minus 1 can be represented as a floating point number of type float.

DBL_MAX_EXP
LDBL_MAX_EXP

These are similar to FLT_MAX_EXP, but for the data types double and long double, respectively.

FLT_MAX_10_EXP

This is the maximum positive integer such that 10 raised to this power minus 1 can be represented as a normalized floating point number of type float. This is supposed to be at least 37.

DBL_MAX_10_EXP
LDBL_MAX_10_EXP

These are similar to FLT_MAX_10_EXP, but for the data types double and long double, respectively.

FLT_MAX

The value of this macro is the maximum number representable in type float. It is supposed to be at least 1E+37. The value has type float. The smallest representable number is - FLT_MAX.

DBL_MAX
LDBL_MAX

These are similar to FLT_MAX, but for the data types double and long double, respectively. The type of the macro’s value is the same as the type it describes.

FLT_MIN

The value of this macro is the minimum normalized positive floating point number that is representable in type float. It is supposed to be no more than 1E-37.

DBL_MIN
LDBL_MIN

These are similar to FLT_MIN, but for the data types double and long double, respectively. The type of the macro’s value is the same as the type it describes.

FLT_EPSILON

This is the minimum positive floating point number of type

float such that 1.0 + FLT_EPSILON != 1.0 is true. It’s sup-

posed to be no greater than 1E-5.

DBL_EPSILON
LDBL_EPSILON

These are similar to FLT_EPSILON, but for the data types double and long double, respectively. The type of the macro’s value is the same as the type it describes. The values are not supposed to be greater than 1E-9.

A.5.3.3 IEEE Floating Point

Here is an example showing how the floating type measurements come out for the most common floating point representation, specified by the IEEE Standard for Binary Floating Point Arithmetic (ANSI/IEEE Std 754-1985). Nearly all computers designed since the 1980s use this format.

The IEEE single-precision float representation uses a base of 2. There is a sign bit, a mantissa with 23 bits plus one hidden bit (so the total precision is 24 base-2 digits), and an 8-bit exponent that can represent values in the range -125 to 128, inclusive.

So, for an implementation that uses this representation for the float data type, appropriate values for the corresponding parameters are:

     FLT_RADIX                             2
     FLT_MANT_DIG                         24
     FLT_DIG                               6
     FLT_MIN_EXP                        -125
     FLT_MIN_10_EXP                      -37
     FLT_MAX_EXP                         128
     FLT_MAX_10_EXP                      +38
     FLT_MIN                 1.17549435E-38F
     FLT_MAX                 3.40282347E+38F
     FLT_EPSILON             1.19209290E-07F

Here are the values for the double data type:

     DBL_MANT_DIG                         53
     DBL_DIG                              15
     DBL_MIN_EXP                       -1021
     DBL_MIN_10_EXP                     -307
     DBL_MAX_EXP                        1024
     DBL_MAX_10_EXP                      308
     DBL_MAX         1.7976931348623157E+308
     DBL_MIN         2.2250738585072014E-308
     DBL_EPSILON     2.2204460492503131E-016

A.5.4 Structure Field Offset Measurement

You can use offsetof to measure the location within a structure type of a particular structure member.

Macro size_t offsetof (type, member)

This expands to a integer constant expression that is the offset of the structure member named member in a the structure type type. For example, offsetof (struct s, elem) is the offset, in bytes, of the

member elem in a struct s.

This macro won’t work if member is a bit field; you get an error from the C compiler in that case.

572 The GNU C Library

Appendix B Summary of Library Facilities

This appendix is a complete list of the facilities declared within the header files supplied with the GNU C library. Each entry also lists the standard or other source from which each facility is derived, and tells you where in the manual you can find more information about how to use it.

void abort (void)

‘stdlib.h’ (ISO): Section 22.3.4 [Aborting a Program], page 465.

int abs (int number)

‘stdlib.h’ (ISO): Section 14.3 [Absolute Value], page 306.

int accept (int socket, struct sockaddr *addr, socklen_t

*length-ptr)

‘sys/socket.h’ (BSD): Section 11.8.3 [Accepting Connections], page 248.

int access (const char *filename, int how)

‘unistd.h’ (POSIX.1): Section 9.8.8 [Testing Permission to Access a File], page 206.

double acosh (double x)

‘math.h’ (BSD): Section 13.5 [Hyperbolic Functions], page 301.

double acos (double x)

‘math.h’ (ISO): Section 13.3 [Inverse Trigonometric Functions], page 299.

int adjtime (const struct timeval *delta, struct timeval *olddelta)

‘sys/time.h’ (BSD): Section 17.2.2 [High-Resolution Calendar], page 342.

AF_FILE

‘sys/socket.h’ (GNU): Section 11.3.1 [Address Formats], page 224.

AF_INET6

‘sys/socket.h’ (IPv6 Basic API): Section 11.3.1 [Address Formats], page 224.

AF_INET

‘sys/socket.h’ (BSD): Section 11.3.1 [Address Formats], page 224.

AF_UNIX

‘sys/socket.h’ (BSD): Section 11.3.1 [Address Formats], page 224.

AF_UNSPEC

‘sys/socket.h’ (BSD): Section 11.3.1 [Address Formats], page 224.

unsigned int alarm (unsigned int seconds)

‘unistd.h’ (POSIX.1): Section 17.3 [Setting an Alarm], page 356.

void * alloca (size_t size);

‘stdlib.h’ (GNU, BSD): Section 3.5 [Automatic Storage with Variable Size], page 55.

tcflag_t ALTWERASE

‘termios.h’ (BSD): Section 12.4.7 [Local Modes], page 282.

int ARG_MAX

‘limits.h’ (POSIX.1): Section 28.1 [General Capacity Limits], page 537.

char * asctime (const struct tm *brokentime)

‘time.h’ (ISO): Section 17.2.4 [Formatting Date and Time], page 347.

double asinh (double x)

‘math.h’ (BSD): Section 13.5 [Hyperbolic Functions], page 301.

double asin (double x)

‘math.h’ (ISO): Section 13.3 [Inverse Trigonometric Functions], page 299.

int asprintf (char **ptr, const char *template, . . .)

‘stdio.h’ (GNU): Section 7.10.8 [Dynamically Allocating Formatted Output], page 114.

void assert (int expression)

‘assert.h’ (ISO): Section A.1 [Explicitly Checking Internal Consistency], page 553.

void assert_perror (int errnum)

‘assert.h’ (GNU): Section A.1 [Explicitly Checking Internal Consistency], page 553.

double atan2 (double y, double x)

‘math.h’ (ISO): Section 13.3 [Inverse Trigonometric Functions], page 299.

double atanh (double x)

‘math.h’ (BSD): Section 13.5 [Hyperbolic Functions], page 301.

double atan (double x)

‘math.h’ (ISO): Section 13.3 [Inverse Trigonometric Functions], page 299.

int atexit (void (*function) (void))

‘stdlib.h’ (ISO): Section 22.3.3 [Cleanups on Exit], page 464.

double atof (const char *string)

‘stdlib.h’ (ISO): Section 14.7.2 [Parsing of Floats], page 314.

int atoi (const char *string)

‘stdlib.h’ (ISO): Section 14.7.1 [Parsing of Integers], page 310.

long int atol (const char *string)

‘stdlib.h’ (ISO): Section 14.7.1 [Parsing of Integers], page 310.

B0

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

B110

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

B1200

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

B134

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

B150

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

B1800

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

B19200

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

B200

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

B2400

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

B300

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

B38400

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

B4800

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

B50

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

B600

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

B75

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

B9600

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

int BC_BASE_MAX

‘limits.h’ (POSIX.2): Section 28.10 [Utility Program Capacity Limits], page 549.

int BC_DIM_MAX

‘limits.h’ (POSIX.2): Section 28.10 [Utility Program Capacity Limits], page 549.

int BC_DIM_MAX

‘limits.h’ (POSIX.2): Section 28.10 [Utility Program Capacity Limits], page 549.

int bcmp (const void *a1, const void *a2, size_t size)

‘string.h’ (BSD): Section 5.5 [String/Array Comparison], page 72.

void * bcopy (void *from, const void *to, size_t size)

‘string.h’ (BSD): Section 5.4 [Copying and Concatenation], page 67.

int BC_SCALE_MAX

‘limits.h’ (POSIX.2): Section 28.10 [Utility Program Capacity Limits], page 549.

int BC_STRING_MAX

‘limits.h’ (POSIX.2): Section 28.10 [Utility Program Capacity Limits], page 549.

int bind (int socket, struct sockaddr *addr, socklen_t length)

‘sys/socket.h’ (BSD): Section 11.3.2 [Setting the Address of a Socket], page 225.

tcflag_t BRKINT

‘termios.h’ (POSIX.1): Section 12.4.4 [Input Modes], page 277.

_BSD_SOURCE

(GNU): Section 1.3.4 [Feature Test Macros], page 8.

void * bsearch (const void *key, const void *array, size_t count,
size_t size, comparison_fn_t compare)

‘stdlib.h’ (ISO): Section 15.2 [Array Search Function], page 317.

int BUFSIZ

‘stdio.h’ (ISO): Section 7.17.3 [Controlling Which Kind of Buffering], page 141.

void * bzero (void *block, size_t size)

‘string.h’ (BSD): Section 5.4 [Copying and Concatenation], page 67.

double cabs (struct { double real, imag; } z)

‘math.h’ (BSD): Section 14.3 [Absolute Value], page 306.

void * calloc (size_t count, size_t eltsize)

‘malloc.h’, ‘stdlib.h’ (ISO): Section 3.3.5 [Allocating Cleared Space], page 36.

double cbrt (double x)

‘math.h’ (BSD): Section 13.4 [Exponentiation and Logarithms], page 300.

cc_t

‘termios.h’ (POSIX.1): Section 12.4.1 [Terminal Mode Data Types], page 273.

tcflag_t CCTS_OFLOW

‘termios.h’ (BSD): Section 12.4.6 [Control Modes], page 280.

double ceil (double x)

‘math.h’ (ISO): Section 14.5 [Rounding and Remainder Functions], page 308.

speed_t cfgetispeed (const struct termios *termios-p)

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

speed_t cfgetospeed (const struct termios *termios-p)

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

int cfmakeraw (struct termios *termios-p)

‘termios.h’ (BSD): Section 12.4.10 [Noncanonical Input], page 291.

void cfree (void *ptr)

‘stdlib.h’ (Sun): Section 3.3.3 [Freeing Memory Allocated with “codemalloc], page 34.

int cfsetispeed (struct termios *termios-p, speed_t speed)

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

int cfsetospeed (struct termios *termios-p, speed_t speed)

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

int cfsetspeed (struct termios *termios-p, speed_t speed)

‘termios.h’ (BSD): Section 12.4.8 [Line Speed], page 284.

CHAR_BIT

‘limits.h’ (ISO): Section A.5.1 [Computing the Width of an Integer Data Type], page 563.

CHAR_MAX

‘limits.h’ (ISO): Section A.5.2 [Range of an Integer Type], page 563.

CHAR_MIN

‘limits.h’ (ISO): Section A.5.2 [Range of an Integer Type], page 563.

int chdir (const char *filename)

‘unistd.h’ (POSIX.1): Section 9.1 [Working Directory], page 181.

int CHILD_MAX

‘limits.h’ (POSIX.1): Section 28.1 [General Capacity Limits], page 537.

int chmod (const char *filename, mode_t mode)

‘sys/stat.h’ (POSIX.1): Section 9.8.7 [Assigning File Permissions], page 204.

int chown (const char *filename, uid_t owner, gid_t group)

‘unistd.h’ (POSIX.1): Section 9.8.4 [File Owner], page 200.

tcflag_t CIGNORE

‘termios.h’ (BSD): Section 12.4.6 [Control Modes], page 280.

void clearerr (FILE *stream)

‘stdio.h’ (ISO): Section 7.13 [End-Of-File and Errors], page 135.

int CLK_TCK

‘time.h’ (POSIX.1): Section 17.1.1 [Basic CPU Time Inquiry], page 339.

tcflag_t CLOCAL

‘termios.h’ (POSIX.1): Section 12.4.6 [Control Modes], page 280.

clock_t clock (void)

‘time.h’ (ISO): Section 17.1.1 [Basic CPU Time Inquiry], page 339.

int CLOCKS_PER_SEC

‘time.h’ (ISO): Section 17.1.1 [Basic CPU Time Inquiry], page 339.

clock_t

‘time.h’ (ISO): Section 17.1.1 [Basic CPU Time Inquiry], page 339.

int closedir (DIR *dirstream)

‘dirent.h’ (POSIX.1): Section 9.2.3 [Reading and Closing a Directory Stream], page 185.

int close (int filedes)

‘unistd.h’ (POSIX.1): Section 8.1 [Opening and Closing Files], page 151.

int COLL_WEIGHTS_MAX

‘limits.h’ (POSIX.2): Section 28.10 [Utility Program Capacity Limits], page 549.

size_t confstr (int parameter, char *buf, size_t len)

‘unistd.h’ (POSIX.2): Section 28.12 [String-Valued Parameters], page 551.

int connect (int socket, struct sockaddr *addr, socklen_t length)

‘sys/socket.h’ (BSD): Section 11.8.1 [Making a Connection], page 246.

cookie_close_function

‘stdio.h’ (GNU): Section 7.18.3.2 [Custom Stream Hook Functions], page 148.

cookie_io_functions_t

‘stdio.h’ (GNU): Section 7.18.3.1 [Custom Streams and Cookies], page 147.

cookie_read_function

‘stdio.h’ (GNU): Section 7.18.3.2 [Custom Stream Hook Functions], page 148.

cookie_seek_function

‘stdio.h’ (GNU): Section 7.18.3.2 [Custom Stream Hook Functions], page 148.

cookie_write_function

‘stdio.h’ (GNU): Section 7.18.3.2 [Custom Stream Hook Functions], page 148.

double copysign (double value, double sign)

‘math.h’ (BSD): Section 14.4 [Normalization Functions], page 307.

double cosh (double x)

‘math.h’ (ISO): Section 13.5 [Hyperbolic Functions], page 301.

double cos (double x)

‘math.h’ (ISO): Section 13.2 [Trigonometric Functions], page 298.

tcflag_t CREAD

‘termios.h’ (POSIX.1): Section 12.4.6 [Control Modes], page 280.

int creat (const char *filename, mode_t mode)

‘fcntl.h’ (POSIX.1): Section 8.1 [Opening and Closing Files], page 151.

tcflag_t CRTS_IFLOW

‘termios.h’ (BSD): Section 12.4.6 [Control Modes], page 280.

tcflag_t CS5

‘termios.h’ (POSIX.1): Section 12.4.6 [Control Modes], page 280.

tcflag_t CS6

‘termios.h’ (POSIX.1): Section 12.4.6 [Control Modes], page 280.

tcflag_t CS7

‘termios.h’ (POSIX.1): Section 12.4.6 [Control Modes], page 280.

tcflag_t CS8

‘termios.h’ (POSIX.1): Section 12.4.6 [Control Modes], page 280.

tcflag_t CSIZE

‘termios.h’ (POSIX.1): Section 12.4.6 [Control Modes], page 280.

_CS_PATH

‘unistd.h’ (POSIX.2): Section 28.12 [String-Valued Parameters], page 551.

tcflag_t CSTOPB

‘termios.h’ (POSIX.1): Section 12.4.6 [Control Modes], page 280.

char * ctermid (char *string)

‘stdio.h’ (POSIX.1): Section 24.7.1 [Identifying the Controlling Terminal], page 497.

char * ctime (const time_t *time)

‘time.h’ (ISO): Section 17.2.4 [Formatting Date and Time], page 347.

char * cuserid (char *string)

‘stdio.h’ (POSIX.1): Section 26.11 [Identifying Who Logged In], page 520.

int daylight

‘time.h’ (SVID): Section 17.2.6 [Functions and Variables for Time Zones], page 354.

DBL_DIG

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

DBL_EPSILON

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

DBL_MANT_DIG

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

DBL_MAX_10_EXP

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

DBL_MAX_EXP

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

DBL_MAX

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

DBL_MIN_10_EXP

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

DBL_MIN_EXP

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

DBL_MIN

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

dev_t

‘sys/types.h’ (POSIX.1): Section 9.8.1 [What the File Attribute Values Mean], page 195.

double difftime (time_t time1, time_t time0)

‘time.h’ (ISO): Section 17.2.1 [Simple Calendar Time], page 342.

DIR

‘dirent.h’ (POSIX.1): Section 9.2.2 [Opening a Directory Stream], page 184.

div_t div (int numerator, int denominator)

‘stdlib.h’ (ISO): Section 14.6 [Integer Division], page 309.

div_t

‘stdlib.h’ (ISO): Section 14.6 [Integer Division], page 309.

double drem (double numerator, double denominator)

‘math.h’ (BSD): Section 14.5 [Rounding and Remainder Functions], page 308.

int dup2 (int old, int new)

‘unistd.h’ (POSIX.1): Section 8.8 [Duplicating Descriptors], page 166.

int dup (int old)

‘unistd.h’ (POSIX.1): Section 8.8 [Duplicating Descriptors], page 166.

int E2BIG

‘errno.h’ (POSIX.1: Argument list too long): Section 2.2 [Error Codes], page 16.

int EACCES

‘errno.h’ (POSIX.1: Permission denied): Section 2.2 [Error Codes], page 16.

int EADDRINUSE

‘errno.h’ (BSD: Address already in use): Section 2.2 [Error Codes], page 16.

int EADDRNOTAVAIL

‘errno.h’ (BSD: Cannot assign requested address): Section 2.2 [Error Codes], page 16.

int EADV

‘errno.h’ (Linux???: Advertise error): Section 2.2 [Error Codes], page 16.

int EAFNOSUPPORT

‘errno.h’ (BSD: Address family not supported by protocol): Section 2.2 [Error Codes], page 16.

int EAGAIN

‘errno.h’ (POSIX.1: Resource temporarily unavailable): Section 2.2 [Error Codes], page 16.

int EALREADY

‘errno.h’ (BSD: Operation already in progress): Section 2.2 [Error Codes], page 16.

int EAUTH

‘errno.h’ (BSD: Authentication error): Section 2.2 [Error Codes], page 16.

int EBACKGROUND

‘errno.h’ (GNU: Inappropriate operation for background process): Section 2.2 [Error Codes], page 16.

int EBADE

‘errno.h’ (Linux???: Invalid exchange): Section 2.2 [Error Codes], page 16.

int EBADFD

‘errno.h’ (Linux???: File descriptor in bad state): Section 2.2 [Error Codes], page 16.

int EBADF

‘errno.h’ (POSIX.1: Bad file descriptor): Section 2.2 [Error Codes], page 16.

int EBADMSG

‘errno.h’ (XOPEN: Bad message): Section 2.2 [Error Codes], page 16.

int EBADR

‘errno.h’ (Linux???: Invalid request descriptor): Section 2.2 [Error Codes], page 16.

int EBADRPC

‘errno.h’ (BSD: RPC struct is bad): Section 2.2 [Error Codes], page 16.

int EBADRQC

‘errno.h’ (Linux???: Invalid request code): Section 2.2 [Error Codes], page 16.

int EBADSLT

‘errno.h’ (Linux???: Invalid slot): Section 2.2 [Error Codes], page 16.

int EBFONT

‘errno.h’ (Linux???: Bad font file format): Section 2.2 [Error Codes], page 16.

int EBUSY

‘errno.h’ (POSIX.1: Device or resource busy): Section 2.2 [Error Codes], page 16.

int ECHILD

‘errno.h’ (POSIX.1: No child processes): Section 2.2 [Error Codes], page 16.

tcflag_t ECHOCTL

‘termios.h’ (BSD): Section 12.4.7 [Local Modes], page 282.

tcflag_t ECHOE

‘termios.h’ (POSIX.1): Section 12.4.7 [Local Modes], page 282.

tcflag_t ECHO

‘termios.h’ (POSIX.1): Section 12.4.7 [Local Modes], page 282.

tcflag_t ECHOKE

‘termios.h’ (BSD): Section 12.4.7 [Local Modes], page 282.

tcflag_t ECHOK

‘termios.h’ (POSIX.1): Section 12.4.7 [Local Modes], page 282.

tcflag_t ECHONL

‘termios.h’ (POSIX.1): Section 12.4.7 [Local Modes], page 282.

tcflag_t ECHOPRT

‘termios.h’ (BSD): Section 12.4.7 [Local Modes], page 282.

int ECHRNG

‘errno.h’ (Linux???: Channel number out of range): Section 2.2 [Error Codes], page 16.

int ECOMM

‘errno.h’ (Linux???: Communication error on send): Section 2.2 [Error Codes], page 16.

int ECONNABORTED

‘errno.h’ (BSD: Software caused connection abort): Section 2.2 [Error Codes], page 16.

int ECONNREFUSED

‘errno.h’ (BSD: Connection refused): Section 2.2 [Error Codes], page 16.

int ECONNRESET

‘errno.h’ (BSD: Connection reset by peer): Section 2.2 [Error Codes], page 16.

int EDEADLK

‘errno.h’ (POSIX.1: Resource deadlock avoided): Section 2.2 [Error Codes], page 16.

int EDEADLOCK

‘errno.h’ (Linux???: File locking deadlock error): Section 2.2 [Error Codes], page 16.

int EDESTADDRREQ

‘errno.h’ (BSD: Destination address required): Section 2.2 [Error Codes], page 16.

int EDIED

‘errno.h’ (GNU: Translator died): Section 2.2 [Error Codes], page 16.

int ED

‘errno.h’ (GNU: ?): Section 2.2 [Error Codes], page 16.

int EDOM

‘errno.h’ (ISO: Numerical argument out of domain): Section 2.2 [Error Codes], page 16.

int EDOTDOT

‘errno.h’ (Linux???: RFS specific error): Section 2.2 [Error Codes], page 16.

int EDQUOT

‘errno.h’ (BSD: Disc quota exceeded): Section 2.2 [Error Codes], page 16.

int EEXIST

‘errno.h’ (POSIX.1: File exists): Section 2.2 [Error Codes], page 16.

int EFAULT

‘errno.h’ (POSIX.1: Bad address): Section 2.2 [Error Codes], page 16.

int EFBIG

‘errno.h’ (POSIX.1: File too large): Section 2.2 [Error Codes], page 16.

int EFTYPE

‘errno.h’ (BSD: Inappropriate file type or format): Section 2.2 [Error Codes], page 16.

int EGRATUITOUS

‘errno.h’ (GNU: Gratuitous error): Section 2.2 [Error Codes], page 16.

int EGREGIOUS

‘errno.h’ (GNU: You really blew it this time): Section 2.2 [Error Codes], page 16.

int EHOSTDOWN

‘errno.h’ (BSD: Host is down): Section 2.2 [Error Codes], page 16.

int EHOSTUNREACH

‘errno.h’ (BSD: No route to host): Section 2.2 [Error Codes], page 16.

int EIDRM

‘errno.h’ (XOPEN: Identifier removed): Section 2.2 [Error Codes], page 16.

int EIEIO

‘errno.h’ (GNU: Computer bought the farm): Section 2.2 [Error Codes], page 16.

int EILSEQ

‘errno.h’ (ISO: Invalid or incomplete multibyte or wide character): Section 2.2 [Error Codes], page 16.

int EINPROGRESS

‘errno.h’ (BSD: Operation now in progress): Section 2.2 [Error Codes], page 16.

int EINTR

‘errno.h’ (POSIX.1: Interrupted system call): Section 2.2 [Error Codes], page 16.

int EINVAL

‘errno.h’ (POSIX.1: Invalid argument): Section 2.2 [Error Codes], page 16.

int EIO

‘errno.h’ (POSIX.1: Input/output error): Section 2.2 [Error Codes], page 16.

int EISCONN

‘errno.h’ (BSD: Transport endpoint is already connected): Section 2.2 [Error Codes], page 16.

int EISDIR

‘errno.h’ (POSIX.1: Is a directory): Section 2.2 [Error Codes], page 16.

int EISNAM

‘errno.h’ (Linux???: Is a named type file): Section 2.2 [Error Codes], page 16.

int EL2HLT

‘errno.h’ (Linux???: Level 2 halted): Section 2.2 [Error Codes], page 16.

int EL2NSYNC

‘errno.h’ (Linux???: Level 2 not synchronized): Section 2.2 [Error Codes], page 16.

int EL3HLT

‘errno.h’ (Linux???: Level 3 halted): Section 2.2 [Error Codes], page 16.

int EL3RST

‘errno.h’ (Linux???: Level 3 reset): Section 2.2 [Error Codes], page 16.

int ELIBACC

‘errno.h’ (Linux???: Can not access a needed shared library): Section 2.2 [Error Codes], page 16.

int ELIBBAD

‘errno.h’ (Linux???: Accessing a corrupted shared library): Section 2.2 [Error Codes], page 16.

int ELIBEXEC

‘errno.h’ (Linux???: Cannot exec a shared library directly): Section 2.2 [Error Codes], page 16.

int ELIBMAX

‘errno.h’ (Linux???: Attempting to link in too many shared libraries): Section 2.2 [Error Codes], page 16.

int ELIBSCN

‘errno.h’ (Linux???: .lib section in a.out corrupted): Section 2.2 [Error Codes], page 16.

int ELNRNG

‘errno.h’ (Linux???: Link number out of range): Section 2.2 [Error Codes], page 16.

int ELOOP

‘errno.h’ (BSD: Too many levels of symbolic links): Section 2.2 [Error Codes], page 16.

int EMEDIUMTYPE

‘errno.h’ (Linux???: Wrong medium type): Section 2.2 [Error Codes], page 16.

int EMFILE

‘errno.h’ (POSIX.1: Too many open files): Section 2.2 [Error Codes], page 16.

int EMLINK

‘errno.h’ (POSIX.1: Too many links): Section 2.2 [Error Codes], page 16.

int EMSGSIZE

‘errno.h’ (BSD: Message too long): Section 2.2 [Error Codes], page 16.

int EMULTIHOP

‘errno.h’ (XOPEN: Multihop attempted): Section 2.2 [Error Codes], page 16.

int ENAMETOOLONG

‘errno.h’ (POSIX.1: File name too long): Section 2.2 [Error Codes], page 16.

int ENAVAIL

‘errno.h’ (Linux???: No XENIX semaphores available): Section 2.2 [Error Codes], page 16.

void endgrent (void)

‘grp.h’ (SVID, BSD): Section 26.13.3 [Scanning the List of All Groups], page 527.

void endhostent ()

‘netdb.h’ (BSD): Section 11.5.2.4 [Host Names], page 235.

void endnetent (void)

‘netdb.h’ (BSD): Section 11.12 [Networks Database], page 269.

void endnetgrent (void)

‘netdb.h’ (netdb.h): Section 26.14.2 [Looking up one Netgroup], page 529.

void endprotoent (void)

‘netdb.h’ (BSD): Section 11.5.6 [Protocols Database], page 240.

void endpwent (void)

‘pwd.h’ (SVID, BSD): Section 26.12.3 [Scanning the List of All Users], page 523.

void endservent (void)

‘netdb.h’ (BSD): Section 11.5.4 [The Services Database], page 238.

int ENEEDAUTH

‘errno.h’ (BSD: Need authenticator): Section 2.2 [Error Codes], page 16.

int ENETDOWN

‘errno.h’ (BSD: Network is down): Section 2.2 [Error Codes], page 16.

int ENETRESET

‘errno.h’ (BSD: Network dropped connection on reset): Section 2.2 [Error Codes], page 16.

int ENETUNREACH

‘errno.h’ (BSD: Network is unreachable): Section 2.2 [Error Codes], page 16.

int ENFILE

‘errno.h’ (POSIX.1: Too many open files in system): Section 2.2 [Error Codes], page 16.

int ENOANO

‘errno.h’ (Linux???: No anode): Section 2.2 [Error Codes], page 16.

int ENOBUFS

‘errno.h’ (BSD: No buffer space available): Section 2.2 [Error Codes], page 16.

int ENOCSI

‘errno.h’ (Linux???: No CSI structure available): Section 2.2 [Error Codes], page 16.

int ENODATA

‘errno.h’ (XOPEN: No data available): Section 2.2 [Error Codes], page 16.

int ENODEV

‘errno.h’ (POSIX.1: Operation not supported by device): Section 2.2 [Error Codes], page 16.

int ENOENT

‘errno.h’ (POSIX.1: No such file or directory): Section 2.2 [Error Codes], page 16.

int ENOEXEC

‘errno.h’ (POSIX.1: Exec format error): Section 2.2 [Error Codes], page 16.

int ENOLCK

‘errno.h’ (POSIX.1: No locks available): Section 2.2 [Error Codes], page 16.

int ENOLINK

‘errno.h’ (XOPEN: Link has been severed): Section 2.2 [Error Codes], page 16.

int ENOMEDIUM

‘errno.h’ (Linux???: No medium found): Section 2.2 [Error Codes], page 16.

int ENOMEM

‘errno.h’ (POSIX.1: Cannot allocate memory): Section 2.2 [Error Codes], page 16.

int ENOMSG

‘errno.h’ (XOPEN: No message of desired type): Section 2.2 [Error Codes], page 16.

int ENONET

‘errno.h’ (Linux???: Machine is not on the network): Section 2.2 [Error Codes], page 16.

int ENOPKG

‘errno.h’ (Linux???: Package not installed): Section 2.2 [Error Codes], page 16.

int ENOPROTOOPT

‘errno.h’ (BSD: Protocol not available): Section 2.2 [Error Codes], page 16.

int ENOSPC

‘errno.h’ (POSIX.1: No space left on device): Section 2.2 [Error Codes], page 16.

int ENOSR

‘errno.h’ (XOPEN: Out of streams resources): Section 2.2 [Error Codes], page 16.

int ENOSTR

‘errno.h’ (XOPEN: Device not a stream): Section 2.2 [Error Codes], page 16.

int ENOSYS

‘errno.h’ (POSIX.1: Function not implemented): Section 2.2 [Error Codes], page 16.

int ENOTBLK

‘errno.h’ (BSD: Block device required): Section 2.2 [Error Codes], page 16.

int ENOTCONN

‘errno.h’ (BSD: Transport endpoint is not connected): Section 2.2 [Error Codes], page 16.

int ENOTDIR

‘errno.h’ (POSIX.1: Not a directory): Section 2.2 [Error Codes], page 16.

int ENOTEMPTY

‘errno.h’ (POSIX.1: Directory not empty): Section 2.2 [Error Codes], page 16.

int ENOTNAM

‘errno.h’ (Linux???: Not a XENIX named type file): Section 2.2 [Error Codes], page 16.

int ENOTSOCK

‘errno.h’ (BSD: Socket operation on non-socket): Section 2.2 [Error Codes], page 16.

int ENOTTY

‘errno.h’ (POSIX.1: Inappropriate ioctl for device): Section 2.2 [Error Codes], page 16.

int ENOTUNIQ

‘errno.h’ (Linux???: Name not unique on network): Section 2.2 [Error Codes], page 16.

char ** environ

‘unistd.h’ (POSIX.1): Section 22.2.1 [Environment Access], page 459.

int ENXIO

‘errno.h’ (POSIX.1: Device not configured): Section 2.2 [Error Codes], page 16.

int EOF

‘stdio.h’ (ISO): Section 7.13 [End-Of-File and Errors], page 135.

int EOPNOTSUPP

‘errno.h’ (BSD: Operation not supported): Section 2.2 [Error Codes], page 16.

int EOVERFLOW

‘errno.h’ (XOPEN: Value too large for defined data type): Section 2.2 [Error Codes], page 16.

int EPERM

‘errno.h’ (POSIX.1: Operation not permitted): Section 2.2 [Error Codes], page 16.

int EPFNOSUPPORT

‘errno.h’ (BSD: Protocol family not supported): Section 2.2 [Error Codes], page 16.

int EPIPE

‘errno.h’ (POSIX.1: Broken pipe): Section 2.2 [Error Codes], page 16.

int EPROCLIM

‘errno.h’ (BSD: Too many processes): Section 2.2 [Error Codes], page 16.

int EPROCUNAVAIL

‘errno.h’ (BSD: RPC bad procedure for program): Section 2.2 [Error Codes], page 16.

int EPROGMISMATCH

‘errno.h’ (BSD: RPC program version wrong): Section 2.2 [Error Codes], page 16.

int EPROGUNAVAIL

‘errno.h’ (BSD: RPC program not available): Section 2.2 [Error Codes], page 16.

int EPROTO

‘errno.h’ (XOPEN: Protocol error): Section 2.2 [Error Codes], page 16.

int EPROTONOSUPPORT

‘errno.h’ (BSD: Protocol not supported): Section 2.2 [Error Codes], page 16.

int EPROTOTYPE

‘errno.h’ (BSD: Protocol wrong type for socket): Section 2.2 [Error Codes], page 16.

int EQUIV_CLASS_MAX

‘limits.h’ (POSIX.2): Section 28.10 [Utility Program Capacity Limits], page 549.

int ERANGE

‘errno.h’ (ISO: Numerical result out of range): Section 2.2 [Error Codes], page 16.

int EREMCHG

‘errno.h’ (Linux???: Remote address changed): Section 2.2 [Error Codes], page 16.

int EREMOTEIO

‘errno.h’ (Linux???: Remote I/O error): Section 2.2 [Error Codes], page 16.

int EREMOTE

‘errno.h’ (BSD: Object is remote): Section 2.2 [Error Codes], page 16.

int ERESTART

‘errno.h’ (Linux???: Interrupted system call should be restarted): Section 2.2 [Error Codes], page 16.

int EROFS

‘errno.h’ (POSIX.1: Read-only file system): Section 2.2 [Error Codes], page 16.

int ERPCMISMATCH

‘errno.h’ (BSD: RPC version wrong): Section 2.2 [Error Codes], page 16.

volatile int errno

‘errno.h’ (ISO): Section 2.1 [Checking for Errors], page 15.

int ESHUTDOWN

‘errno.h’ (BSD: Cannot send after transport endpoint shutdown): Section 2.2 [Error Codes], page 16.

int ESOCKTNOSUPPORT

‘errno.h’ (BSD: Socket type not supported): Section 2.2 [Error Codes], page 16.

int ESPIPE

‘errno.h’ (POSIX.1: Illegal seek): Section 2.2 [Error Codes], page 16.

int ESRCH

‘errno.h’ (POSIX.1: No such process): Section 2.2 [Error Codes], page 16.

int ESRMNT

‘errno.h’ (Linux???: Srmount error): Section 2.2 [Error Codes], page 16.

int ESTALE

‘errno.h’ (BSD: Stale NFS file handle): Section 2.2 [Error Codes], page 16.

int ESTRPIPE

‘errno.h’ (Linux???: Streams pipe error): Section 2.2 [Error Codes], page 16.

int ETIMEDOUT

‘errno.h’ (BSD: Connection timed out): Section 2.2 [Error Codes], page 16.

int ETIME

‘errno.h’ (XOPEN: Timer expired): Section 2.2 [Error Codes], page 16.

int ETOOMANYREFS

‘errno.h’ (BSD: Too many references: cannot splice): Section 2.2 [Error Codes], page 16.

int ETXTBSY

‘errno.h’ (BSD: Text file busy): Section 2.2 [Error Codes], page 16.

int EUCLEAN

‘errno.h’ (Linux???: Structure needs cleaning): Section 2.2 [Error Codes], page 16.

int EUNATCH

‘errno.h’ (Linux???: Protocol driver not attached): Section 2.2 [Error Codes], page 16.

int EUSERS

‘errno.h’ (BSD: Too many users): Section 2.2 [Error Codes], page 16.

int EWOULDBLOCK

‘errno.h’ (BSD: Operation would block): Section 2.2 [Error Codes], page 16.

int EXDEV

‘errno.h’ (POSIX.1: Invalid cross-device link): Section 2.2 [Error Codes], page 16.

int execle (const char *filename, const char *arg0, char *const

env[], …)

‘unistd.h’ (POSIX.1): Section 23.5 [Executing a File], page 470.

int execl (const char *filename, const char *arg0, . . .)

‘unistd.h’ (POSIX.1): Section 23.5 [Executing a File], page 470.

int execlp (const char *filename, const char *arg0, . . .)

‘unistd.h’ (POSIX.1): Section 23.5 [Executing a File], page 470.

int execve (const char *filename, char *const argv[], char *const
env[])

‘unistd.h’ (POSIX.1): Section 23.5 [Executing a File], page 470.

int execv (const char *filename, char *const argv[])

‘unistd.h’ (POSIX.1): Section 23.5 [Executing a File], page 470.

int execvp (const char *filename, char *const argv[])

‘unistd.h’ (POSIX.1): Section 23.5 [Executing a File], page 470.

int EXFULL

‘errno.h’ (Linux???: Exchange full): Section 2.2 [Error Codes], page 16.

int EXIT_FAILURE

‘stdlib.h’ (ISO): Section 22.3.2 [Exit Status], page 463.

void _exit (int status)

‘unistd.h’ (POSIX.1): Section 22.3.5 [Termination Internals], page 465.

void exit (int status)

‘stdlib.h’ (ISO): Section 22.3.1 [Normal Termination], page 462.

int EXIT_SUCCESS

‘stdlib.h’ (ISO): Section 22.3.2 [Exit Status], page 463.

double exp (double x)

‘math.h’ (ISO): Section 13.4 [Exponentiation and Logarithms], page 300.

double expm1 (double x)

‘math.h’ (BSD): Section 13.4 [Exponentiation and Logarithms], page 300.

int EXPR_NEST_MAX

‘limits.h’ (POSIX.2): Section 28.10 [Utility Program Capacity Limits], page 549.

double fabs (double number)

‘math.h’ (ISO): Section 14.3 [Absolute Value], page 306.

int fchmod (int filedes, int mode)

‘sys/stat.h’ (BSD): Section 9.8.7 [Assigning File Permissions], page 204.

int fchown (int filedes, int owner, int group)

‘unistd.h’ (BSD): Section 9.8.4 [File Owner], page 200.

int fclean (FILE *stream)

‘stdio.h’ (GNU): Section 8.5.3 [Cleaning Streams], page 161.

int fcloseall (void)

‘stdio.h’ (GNU): Section 7.4 [Closing Streams], page 96.

int fclose (FILE *stream)

‘stdio.h’ (ISO): Section 7.4 [Closing Streams], page 96.

int fcntl (int filedes, int command, . . .)

‘fcntl.h’ (POSIX.1): Section 8.7 [Control Operations on Files], page 165.

int FD_CLOEXEC

‘fcntl.h’ (POSIX.1): Section 8.9 [File Descriptor Flags], page 168.

void FD_CLR (int filedes, fd_set *set)

‘sys/types.h’ (BSD): Section 8.6 [Waiting for Input or Output], page 162.

int FD_ISSET (int filedes, fd_set *set)

‘sys/types.h’ (BSD): Section 8.6 [Waiting for Input or Output], page 162.

FILE * fdopen (int filedes, const char *opentype)

‘stdio.h’ (POSIX.1): Section 8.4 [Descriptors and Streams], page 159.

void FD_SET (int filedes, fd_set *set)

‘sys/types.h’ (BSD): Section 8.6 [Waiting for Input or Output], page 162.

fd_set

‘sys/types.h’ (BSD): Section 8.6 [Waiting for Input or Output], page 162.

int FD_SETSIZE

‘sys/types.h’ (BSD): Section 8.6 [Waiting for Input or Output], page 162.

int F_DUPFD

‘fcntl.h’ (POSIX.1): Section 8.8 [Duplicating Descriptors], page 166.

void FD_ZERO (fd_set *set)

‘sys/types.h’ (BSD): Section 8.6 [Waiting for Input or Output], page 162.

int feof (FILE *stream)

‘stdio.h’ (ISO): Section 7.13 [End-Of-File and Errors], page 135.

int ferror (FILE *stream)

‘stdio.h’ (ISO): Section 7.13 [End-Of-File and Errors], page 135.

int fflush (FILE *stream)

‘stdio.h’ (ISO): Section 7.17.2 [Flushing Buffers], page 141.

int fgetc (FILE *stream)

‘stdio.h’ (ISO): Section 7.6 [Character Input], page 98.

int F_GETFD

‘fcntl.h’ (POSIX.1): Section 8.9 [File Descriptor Flags], page 168.

int F_GETFL

‘fcntl.h’ (POSIX.1): Section 8.10.4 [Getting and Setting File Status Flags], page 174.

struct group * fgetgrent (FILE *stream)

‘grp.h’ (SVID): Section 26.13.3 [Scanning the List of All Groups], page 527.

int fgetgrent_r (FILE *stream, struct group *result buf, char
*buffer, size_t buflen, struct group **result)

‘grp.h’ (GNU): Section 26.13.3 [Scanning the List of All Groups], page 527.

int F_GETLK

‘fcntl.h’ (POSIX.1): Section 8.11 [File Locks], page 175.

int F_GETOWN

‘fcntl.h’ (BSD): Section 8.12 [Interrupt-Driven Input], page 179.

int fgetpos (FILE *stream, fpos_t *position)

‘stdio.h’ (ISO): Section 7.16 [Portable File-Position Functions], page 138.

struct passwd * fgetpwent (FILE *stream)

‘pwd.h’ (SVID): Section 26.12.3 [Scanning the List of All Users], page 523.

int fgetpwent_r (FILE *stream, struct passwd *result buf, char
*buffer, size_t buflen, struct passwd **result)

‘pwd.h’ (GNU): Section 26.12.3 [Scanning the List of All Users], page 523.

char * fgets (char *s, int count, FILE *stream)

‘stdio.h’ (ISO): Section 7.7 [Line-Oriented Input], page 100.

FILE

‘stdio.h’ (ISO): Section 7.1 [Streams], page 93.

int FILENAME_MAX

‘stdio.h’ (ISO): Section 28.6 [Limits on File System Capacity], page 545.

int fileno (FILE *stream)

‘stdio.h’ (POSIX.1): Section 8.4 [Descriptors and Streams], page 159.

int finite (double x)

‘math.h’ (BSD): Section 14.2 [Predicates on Floats], page 305.

double floor (double x)

‘math.h’ (ISO): Section 14.5 [Rounding and Remainder Functions], page 308.

FLT_DIG

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

FLT_EPSILON

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

FLT_MANT_DIG

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

FLT_MAX_10_EXP

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

FLT_MAX_EXP

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

FLT_MAX

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

FLT_MIN_10_EXP

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

FLT_MIN_EXP

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

FLT_MIN

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

FLT_RADIX

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

FLT_ROUNDS

‘float.h’ (ISO): Section A.5.3.2 [Floating Point Parameters], page 566.

tcflag_t FLUSHO

‘termios.h’ (BSD): Section 12.4.7 [Local Modes], page 282.

FILE * fmemopen (void *buf, size_t size, const char *opentype)

‘stdio.h’ (GNU): Section 7.18.1 [String Streams], page 144.

double fmod (double numerator, double denominator)

‘math.h’ (ISO): Section 14.5 [Rounding and Remainder Functions], page 308.

int fnmatch (const char *pattern, const char *string, int flags)

‘fnmatch.h’ (POSIX.2): Section 16.1 [Wildcard Matching], page 323.

FNM_CASEFOLD

‘fnmatch.h’ (GNU): Section 16.1 [Wildcard Matching], page 323.

FNM_FILE_NAME

‘fnmatch.h’ (GNU): Section 16.1 [Wildcard Matching], page 323.

FNM_LEADING_DIR

‘fnmatch.h’ (GNU): Section 16.1 [Wildcard Matching], page 323.

FNM_NOESCAPE

‘fnmatch.h’ (POSIX.2): Section 16.1 [Wildcard Matching], page 323.

FNM_PATHNAME

‘fnmatch.h’ (POSIX.2): Section 16.1 [Wildcard Matching], page 323.

FNM_PERIOD

‘fnmatch.h’ (POSIX.2): Section 16.1 [Wildcard Matching], page 323.

int F_OK

‘unistd.h’ (POSIX.1): Section 9.8.8 [Testing Permission to Access a File], page 206.

FILE * fopencookie (void *cookie, const char *opentype,
cookie_io_functions_t io-functions)

‘stdio.h’ (GNU): Section 7.18.3.1 [Custom Streams and Cookies], page 147.

FILE * fopen (const char *filename, const char *opentype)

‘stdio.h’ (ISO): Section 7.3 [Opening Streams], page 94.

int FOPEN_MAX

‘stdio.h’ (ISO): Section 7.3 [Opening Streams], page 94.

pid_t fork (void)

‘unistd.h’ (POSIX.1): Section 23.4 [Creating a Process], page 469.

long int fpathconf (int filedes, int parameter)

‘unistd.h’ (POSIX.1): Section 28.9 [Using “codepathconf], page 548.

FPE_DECOVF_TRAP

‘signal.h’ (BSD): Section 21.2.1 [Program Error Signals], page 396.

FPE_FLTDIV_FAULT

‘signal.h’ (BSD): Section 21.2.1 [Program Error Signals], page 396.

FPE_FLTDIV_TRAP

‘signal.h’ (BSD): Section 21.2.1 [Program Error Signals], page 396.

FPE_FLTOVF_FAULT

‘signal.h’ (BSD): Section 21.2.1 [Program Error Signals], page 396.

FPE_FLTOVF_TRAP

‘signal.h’ (BSD): Section 21.2.1 [Program Error Signals], page 396.

FPE_FLTUND_FAULT

‘signal.h’ (BSD): Section 21.2.1 [Program Error Signals], page 396.

FPE_FLTUND_TRAP

‘signal.h’ (BSD): Section 21.2.1 [Program Error Signals], page 396.

FPE_INTDIV_TRAP

‘signal.h’ (BSD): Section 21.2.1 [Program Error Signals], page 396.

FPE_INTOVF_TRAP

‘signal.h’ (BSD): Section 21.2.1 [Program Error Signals], page 396.

FPE_SUBRNG_TRAP

‘signal.h’ (BSD): Section 21.2.1 [Program Error Signals], page 396.

fpos_t

‘stdio.h’ (ISO): Section 7.16 [Portable File-Position Functions], page 138.

int fprintf (FILE *stream, const char *template, . . .)

‘stdio.h’ (ISO): Section 7.10.7 [Formatted Output Functions], page 113.

int fputc (int c, FILE *stream)

‘stdio.h’ (ISO): Section 7.5 [Simple Output by Characters or Lines], page 97.

int fputs (const char *s, FILE *stream)

‘stdio.h’ (ISO): Section 7.5 [Simple Output by Characters or Lines], page 97.

F_RDLCK

‘fcntl.h’ (POSIX.1): Section 8.11 [File Locks], page 175.

size_t fread (void *data, size_t size, size_t count, FILE *stream)

‘stdio.h’ (ISO): Section 7.9 [Block Input/Output], page 104.

__free_hook

‘malloc.h’ (GNU): Section 3.3.10 [Storage Allocation Hooks], page 40.

void free (void *ptr)

‘malloc.h’, ‘stdlib.h’ (ISO): Section 3.3.3 [Freeing Memory Allocated with “codemalloc], page 34.

FILE * freopen (const char *filename, const char *opentype, FILE

*stream)

‘stdio.h’ (ISO): Section 7.3 [Opening Streams], page 94.

double frexp (double value, int *exponent)

‘math.h’ (ISO): Section 14.4 [Normalization Functions], page 307.

int fscanf (FILE *stream, const char *template, . . .)

‘stdio.h’ (ISO): Section 7.12.8 [Formatted Input Functions], page 134.

int fseek (FILE *stream, long int offset, int whence)

‘stdio.h’ (ISO): Section 7.15 [File Positioning], page 137.

int F_SETFD

‘fcntl.h’ (POSIX.1): Section 8.9 [File Descriptor Flags], page 168.

int F_SETFL

‘fcntl.h’ (POSIX.1): Section 8.10.4 [Getting and Setting File Status Flags], page 174.

int F_SETLK

‘fcntl.h’ (POSIX.1): Section 8.11 [File Locks], page 175.

int F_SETLKW

‘fcntl.h’ (POSIX.1): Section 8.11 [File Locks], page 175.

int F_SETOWN

‘fcntl.h’ (BSD): Section 8.12 [Interrupt-Driven Input], page 179.

int fsetpos (FILE *stream, const fpos_t position)

‘stdio.h’ (ISO): Section 7.16 [Portable File-Position Functions], page 138.

int fstat (int filedes, struct stat *buf )

‘sys/stat.h’ (POSIX.1): Section 9.8.2 [Reading the Attributes of a File], page 198.

long int ftell (FILE *stream)

‘stdio.h’ (ISO): Section 7.15 [File Positioning], page 137.

F_UNLCK

‘fcntl.h’ (POSIX.1): Section 8.11 [File Locks], page 175.

size_t fwrite (const void *data, size_t size, size_t count, FILE

*stream)

‘stdio.h’ (ISO): Section 7.9 [Block Input/Output], page 104.

F_WRLCK

‘fcntl.h’ (POSIX.1): Section 8.11 [File Locks], page 175.

int getchar (void)

‘stdio.h’ (ISO): Section 7.6 [Character Input], page 98.

int getc (FILE *stream)

‘stdio.h’ (ISO): Section 7.6 [Character Input], page 98.

char * getcwd (char *buffer, size_t size)

‘unistd.h’ (POSIX.1): Section 9.1 [Working Directory], page 181.

ssize_t getdelim (char **lineptr, size_t *n, int delimiter, FILE

*stream)

‘stdio.h’ (GNU): Section 7.7 [Line-Oriented Input], page 100.

gid_t getegid (void)

‘unistd.h’ (POSIX.1): Section 26.5 [Reading the Persona of a Process], page 513.

char * getenv (const char *name)

‘stdlib.h’ (ISO): Section 22.2.1 [Environment Access], page 459.

uid_t geteuid (void)

‘unistd.h’ (POSIX.1): Section 26.5 [Reading the Persona of a Process], page 513.

gid_t getgid (void)

‘unistd.h’ (POSIX.1): Section 26.5 [Reading the Persona of a Process], page 513.

struct group * getgrent (void)

‘grp.h’ (SVID, BSD): Section 26.13.3 [Scanning the List of All Groups], page 527.

int getgrent_r (struct group *result buf, char *buffer, size_t
buflen, struct group **result)

‘grp.h’ (GNU): Section 26.13.3 [Scanning the List of All Groups], page 527.

struct group * getgrgid (gid_t gid)

‘grp.h’ (POSIX.1): Section 26.13.2 [Looking Up One Group], page 525.

int getgrgid_r (gid_t gid, struct group *result buf, char *buffer,
size_t buflen, struct group **result)

‘grp.h’ (POSIX.1c): Section 26.13.2 [Looking Up One Group], page 525.

struct group * getgrnam (const char *name)

‘grp.h’ (SVID, BSD): Section 26.13.2 [Looking Up One Group], page 525.

int getgrnam_r (const char *name, struct group *result buf, char
*buffer, size_t buflen, struct group **result)

‘grp.h’ (POSIX.1c): Section 26.13.2 [Looking Up One Group], page 525.

int getgroups (int count, gid_t *groups)

‘unistd.h’ (POSIX.1): Section 26.5 [Reading the Persona of a Process], page 513.

struct hostent * gethostbyaddr (const char *addr, int length, int

format)

‘netdb.h’ (BSD): Section 11.5.2.4 [Host Names], page 235.

struct hostent * gethostbyname2 (const char *name, int af )

‘netdb.h’ (IPv6 Basic API): Section 11.5.2.4 [Host Names], page 235.

struct hostent * gethostbyname (const char *name)

‘netdb.h’ (BSD): Section 11.5.2.4 [Host Names], page 235.

struct hostent * gethostent ()

‘netdb.h’ (BSD): Section 11.5.2.4 [Host Names], page 235.

long int gethostid (void)

‘unistd.h’ (BSD): Section 27.1 [Host Identification], page 533.

int gethostname (char *name, size_t size)

‘unistd.h’ (BSD): Section 27.1 [Host Identification], page 533.

int getitimer (int which, struct itimerval *old)

‘sys/time.h’ (BSD): Section 17.3 [Setting an Alarm], page 356.

ssize_t getline (char **lineptr, size_t *n, FILE *stream)

‘stdio.h’ (GNU): Section 7.7 [Line-Oriented Input], page 100.

char * getlogin (void)

‘unistd.h’ (POSIX.1): Section 26.11 [Identifying Who Logged In], page 520.

struct netent * getnetbyaddr (long net, int type)

‘netdb.h’ (BSD): Section 11.12 [Networks Database], page 269.

struct netent * getnetbyname (const char *name)

‘netdb.h’ (BSD): Section 11.12 [Networks Database], page 269.

struct netent * getnetent (void)

‘netdb.h’ (BSD): Section 11.12 [Networks Database], page 269.

int getnetgrent (char **hostp, char **userp, char **domainp)

‘netdb.h’ (netdb.h): Section 26.14.2 [Looking up one Netgroup], page 529.

int getnetgrent_r (char **hostp, char **userp, char **domainp,
char *buffer, int buflen)

‘netdb.h’ (netdb.h): Section 26.14.2 [Looking up one Netgroup], page 529.

int getopt (int argc, char **argv, const char *options)

‘unistd.h’ (POSIX.2): Section 22.1.2 [Parsing Program Options], page 449.

int getopt_long (int argc, char **argv, const char *shortopts,
struct option *longopts, int *indexptr)

‘getopt.h’ (GNU): Section 22.1.4 [Parsing Long Options], page 452.

int getpeername (int socket, struct sockaddr *addr, size_t

*length-ptr)

‘sys/socket.h’ (BSD): Section 11.8.4 [Who is Connected to Me?], page 250.

pid_t getpgrp (pid_t pid)

‘unistd.h’ (BSD): Section 24.7.2 [Process Group Functions], page 498.

pid_t getpgrp (void)

‘unistd.h’ (POSIX.1): Section 24.7.2 [Process Group Functions], page 498.

pid_t getpid (void)

‘unistd.h’ (POSIX.1): Section 23.3 [Process Identification], page 468.

pid_t getppid (void)

‘unistd.h’ (POSIX.1): Section 23.3 [Process Identification], page 468.

int getpriority (int class, int id)

‘sys/resource.h’ (BSD): Section 17.7 [Process Priority], page 364.

struct protoent * getprotobyname (const char *name)

‘netdb.h’ (BSD): Section 11.5.6 [Protocols Database], page 240.

struct protoent * getprotobynumber (int protocol)

‘netdb.h’ (BSD): Section 11.5.6 [Protocols Database], page 240.

struct protoent * getprotoent (void)

‘netdb.h’ (BSD): Section 11.5.6 [Protocols Database], page 240.

struct passwd * getpwent (void)

‘pwd.h’ (POSIX.1): Section 26.12.3 [Scanning the List of All Users], page 523.

int getpwent_r (struct passwd *result buf, char *buffer, int buflen,
struct passwd **result)

‘pwd.h’ (GNU): Section 26.12.3 [Scanning the List of All Users], page 523.

struct passwd * getpwnam (const char *name)

‘pwd.h’ (POSIX.1): Section 26.12.2 [Looking Up One User], page 522.

int getpwnam_r (const char *name, struct passwd *result buf, char
*buffer, size_t buflen, struct passwd **result)

‘pwd.h’ (POSIX.1c): Section 26.12.2 [Looking Up One User], page 522.

struct passwd * getpwuid (uid_t uid)

‘pwd.h’ (POSIX.1): Section 26.12.2 [Looking Up One User], page 522.

int getpwuid_r (uid_t uid, struct passwd *result buf, char *buffer,
size_t buflen, struct passwd **result)

‘pwd.h’ (POSIX.1c): Section 26.12.2 [Looking Up One User], page 522.

int getrlimit (int resource, struct rlimit *rlp)

‘sys/resource.h’ (BSD): Section 17.6 [Limiting Resource Usage], page 362.

int getrusage (int processes, struct rusage *rusage)

‘sys/resource.h’ (BSD): Section 17.5 [Resource Usage], page 360.

struct servent * getservbyname (const char *name, const char

*proto)

‘netdb.h’ (BSD): Section 11.5.4 [The Services Database], page 238.

struct servent * getservbyport (int port, const char *proto)

‘netdb.h’ (BSD): Section 11.5.4 [The Services Database], page 238.

struct servent * getservent (void)

‘netdb.h’ (BSD): Section 11.5.4 [The Services Database], page 238.

char * gets (char *s)

‘stdio.h’ (ISO): Section 7.7 [Line-Oriented Input], page 100.

int getsockname (int socket, struct sockaddr *addr, socklen_t

*length-ptr)

‘sys/socket.h’ (BSD): Section 11.3.3 [Reading the Address of a Socket], page 226.

int getsockopt (int socket, int level, int optname, void *optval,
socklen_t *optlen-ptr)

‘sys/socket.h’ (BSD): Section 11.11.1 [Socket Option Functions], page 266.

int getsubopt (char **optionp, const char* const *tokens, char

**valuep)

‘stdlib.h’ (stdlib.h): Section 22.1.6 [Parsing of Suboptions], page 456.

int gettimeofday (struct timeval *tp, struct timezone *tzp)

‘sys/time.h’ (BSD): Section 17.2.2 [High-Resolution Calendar], page 342.

uid_t getuid (void)

‘unistd.h’ (POSIX.1): Section 26.5 [Reading the Persona of a Process], page 513.

mode_t getumask (void)

‘sys/stat.h’ (GNU): Section 9.8.7 [Assigning File Permissions], page 204.

char * getwd (char *buffer)

‘unistd.h’ (BSD): Section 9.1 [Working Directory], page 181.

int getw (FILE *stream)

‘stdio.h’ (SVID): Section 7.6 [Character Input], page 98.

gid_t

‘sys/types.h’ (POSIX.1): Section 26.5 [Reading the Persona of a Process], page 513.

GLOB_ABORTED

‘glob.h’ (POSIX.2): Section 16.2.1 [Calling “codeglob], page 324.

GLOB_APPEND

‘glob.h’ (POSIX.2): Section 16.2.2 [Flags for Globbing], page 326.

GLOB_DOOFFS

‘glob.h’ (POSIX.2): Section 16.2.2 [Flags for Globbing], page 326.

GLOB_ERR

‘glob.h’ (POSIX.2): Section 16.2.2 [Flags for Globbing], page 326.

int glob (const char *pattern, int flags, int (*errfunc) (const char

*filename, int error-code), glob_t *vector-ptr)

‘glob.h’ (POSIX.2): Section 16.2.1 [Calling “codeglob], page 324.

GLOB_MARK

‘glob.h’ (POSIX.2): Section 16.2.2 [Flags for Globbing], page 326.

GLOB_NOCHECK

‘glob.h’ (POSIX.2): Section 16.2.2 [Flags for Globbing], page 326.

GLOB_NOESCAPE

‘glob.h’ (POSIX.2): Section 16.2.2 [Flags for Globbing], page 326.

GLOB_NOMATCH

‘glob.h’ (POSIX.2): Section 16.2.1 [Calling “codeglob], page 324.

GLOB_NOSORT

‘glob.h’ (POSIX.2): Section 16.2.2 [Flags for Globbing], page 326.

GLOB_NOSPACE

‘glob.h’ (POSIX.2): Section 16.2.1 [Calling “codeglob], page 324.

glob_t

‘glob.h’ (POSIX.2): Section 16.2.1 [Calling “codeglob], page 324.

struct tm * gmtime (const time_t *time)

‘time.h’ (ISO): Section 17.2.3 [Broken-down Time], page 345.

_GNU_SOURCE

(GNU): Section 1.3.4 [Feature Test Macros], page 8.

int gsignal (int signum)

‘signal.h’ (SVID): Section 21.6.1 [Signaling Yourself], page 426.

HOST_NOT_FOUND

‘netdb.h’ (BSD): Section 11.5.2.4 [Host Names], page 235.

unsigned long int htonl (unsigned long int hostlong)

‘netinet/in.h’ (BSD): Section 11.5.5 [Byte Order Conversion], page 239.

unsigned short int htons (unsigned short int hostshort)

‘netinet/in.h’ (BSD): Section 11.5.5 [Byte Order Conversion], page 239.

float HUGE_VALf

‘math.h’ (GNU): Section 13.1 [Domain and Range Errors], page 297.

double HUGE_VAL

‘math.h’ (ISO): Section 13.1 [Domain and Range Errors], page 297.

long double HUGE_VALl

‘math.h’ (GNU): Section 13.1 [Domain and Range Errors], page 297.

tcflag_t HUPCL

‘termios.h’ (POSIX.1): Section 12.4.6 [Control Modes], page 280.

double hypot (double x, double y)

‘math.h’ (BSD): Section 13.4 [Exponentiation and Logarithms], page 300.

tcflag_t ICANON

‘termios.h’ (POSIX.1): Section 12.4.7 [Local Modes], page 282.

tcflag_t ICRNL

‘termios.h’ (POSIX.1): Section 12.4.4 [Input Modes], page 277.

tcflag_t IEXTEN

‘termios.h’ (POSIX.1): Section 12.4.7 [Local Modes], page 282.

tcflag_t IGNBRK

‘termios.h’ (POSIX.1): Section 12.4.4 [Input Modes], page 277.

tcflag_t IGNCR

‘termios.h’ (POSIX.1): Section 12.4.4 [Input Modes], page 277.

tcflag_t IGNPAR

‘termios.h’ (POSIX.1): Section 12.4.4 [Input Modes], page 277.

tcflag_t IMAXBEL

‘termios.h’ (BSD): Section 12.4.4 [Input Modes], page 277.

struct in6_addr in6addr_any

‘netinet/in.h’ (IPv6 basic API): Section 11.5.2.2 [Host Address Data Type], page 232.

struct in6_addr in6addr_loopback.

‘netinet/in.h’ (IPv6 basic API): Section 11.5.2.2 [Host Address Data Type], page 232.

unsigned int INADDR_ANY

‘netinet/in.h’ (BSD): Section 11.5.2.2 [Host Address Data Type], page 232.

unsigned int INADDR_BROADCAST

‘netinet/in.h’ (BSD): Section 11.5.2.2 [Host Address Data Type], page 232.

unsigned int INADDR_LOOPBACK

‘netinet/in.h’ (BSD): Section 11.5.2.2 [Host Address Data Type], page 232.

unsigned int INADDR_NONE

‘netinet/in.h’ (BSD): Section 11.5.2.2 [Host Address Data Type], page 232.

char * index (const char *string, int c)

‘string.h’ (BSD): Section 5.7 [Search Functions], page 78.

unsigned long int inet_addr (const char *name)

‘arpa/inet.h’ (BSD): Section 11.5.2.3 [Host Address Functions], page 233.

int inet_aton (const char *name, struct in_addr *addr)

‘arpa/inet.h’ (BSD): Section 11.5.2.3 [Host Address Functions], page 233.

int inet_lnaof (struct in_addr addr)

‘arpa/inet.h’ (BSD): Section 11.5.2.3 [Host Address Functions], page 233.

struct in_addr inet_makeaddr (int net, int local)

‘arpa/inet.h’ (BSD): Section 11.5.2.3 [Host Address Functions], page 233.

int inet_netof (struct in_addr addr)

‘arpa/inet.h’ (BSD): Section 11.5.2.3 [Host Address Functions], page 233.

unsigned long int inet_network (const char *name)

‘arpa/inet.h’ (BSD): Section 11.5.2.3 [Host Address Functions], page 233.

char * inet_ntoa (struct in_addr addr)

‘arpa/inet.h’ (BSD): Section 11.5.2.3 [Host Address Functions], page 233.

char * inet_ntop (int af, const void *cp, char *buf, size_t len)

‘arpa/inet.h’ (IPv6 basic API): Section 11.5.2.3 [Host Address Functions], page 233.

int inet_pton (int af, const char *cp, void *buf )

‘arpa/inet.h’ (IPv6 basic API): Section 11.5.2.3 [Host Address Functions], page 233.

double infnan (int error)

‘math.h’ (BSD): Section 14.2 [Predicates on Floats], page 305.

int initgroups (const char *user, gid_t gid)

‘grp.h’ (BSD): Section 26.7 [Setting the Group IDs], page 515.

void * initstate (unsigned int seed, void *state, size_t size)

‘stdlib.h’ (BSD): Section 13.6.2 [BSD Random Number Functions], page 303.

tcflag_t INLCR

‘termios.h’ (POSIX.1): Section 12.4.4 [Input Modes], page 277.

int innetgr (const char *netgroup, const char *host, const char
*user, const char *domain)

‘netdb.h’ (netdb.h): Section 26.14.3 [Testing for Netgroup Membership], page 530.

ino_t

‘sys/types.h’ (POSIX.1): Section 9.8.1 [What the File Attribute Values Mean], page 195.

tcflag_t INPCK

‘termios.h’ (POSIX.1): Section 12.4.4 [Input Modes], page 277.

int RLIM_INFINITY

‘sys/resource.h’ (BSD): Section 17.6 [Limiting Resource Usage], page 362.

INT_MAX

‘limits.h’ (ISO): Section A.5.2 [Range of an Integer Type], page 563.

INT_MIN

‘limits.h’ (ISO): Section A.5.2 [Range of an Integer Type], page 563.

int _IOFBF

‘stdio.h’ (ISO): Section 7.17.3 [Controlling Which Kind of Buffering], page 141.

int _IOLBF

‘stdio.h’ (ISO): Section 7.17.3 [Controlling Which Kind of Buffering], page 141.

int _IONBF

‘stdio.h’ (ISO): Section 7.17.3 [Controlling Which Kind of Buffering], page 141.

int IPPORT_RESERVED

‘netinet/in.h’ (BSD): Section 11.5.3 [Internet Ports], page 237.

int IPPORT_USERRESERVED

‘netinet/in.h’ (BSD): Section 11.5.3 [Internet Ports], page 237.

int isalnum (int c)

‘ctype.h’ (ISO): Section 4.1 [Classification of Characters], page 61.

int isalpha (int c)

‘ctype.h’ (ISO): Section 4.1 [Classification of Characters], page 61.

int isascii (int c)

‘ctype.h’ (SVID, BSD): Section 4.1 [Classification of Characters], page 61.

int isatty (int filedes)

‘unistd.h’ (POSIX.1): Section 12.1 [Identifying Terminals], page 271.

int isblank (int c)

‘ctype.h’ (GNU): Section 4.1 [Classification of Characters], page 61.

int iscntrl (int c)

‘ctype.h’ (ISO): Section 4.1 [Classification of Characters], page 61.

int isdigit (int c)

‘ctype.h’ (ISO): Section 4.1 [Classification of Characters], page 61.

int isgraph (int c)

‘ctype.h’ (ISO): Section 4.1 [Classification of Characters], page 61.

tcflag_t ISIG

‘termios.h’ (POSIX.1): Section 12.4.7 [Local Modes], page 282.

int isinf (double x)

‘math.h’ (BSD): Section 14.2 [Predicates on Floats], page 305.

int islower (int c)

‘ctype.h’ (ISO): Section 4.1 [Classification of Characters], page 61.

int isnan (double x)

‘math.h’ (BSD): Section 14.2 [Predicates on Floats], page 305.

int isprint (int c)

‘ctype.h’ (ISO): Section 4.1 [Classification of Characters], page 61.

int ispunct (int c)

‘ctype.h’ (ISO): Section 4.1 [Classification of Characters], page 61.

int isspace (int c)

‘ctype.h’ (ISO): Section 4.1 [Classification of Characters], page 61.

tcflag_t ISTRIP

‘termios.h’ (POSIX.1): Section 12.4.4 [Input Modes], page 277.

int isupper (int c)

‘ctype.h’ (ISO): Section 4.1 [Classification of Characters], page 61.

int isxdigit (int c)

‘ctype.h’ (ISO): Section 4.1 [Classification of Characters], page 61.

ITIMER_PROF

‘sys/time.h’ (BSD): Section 17.3 [Setting an Alarm], page 356.

ITIMER_REAL

‘sys/time.h’ (BSD): Section 17.3 [Setting an Alarm], page 356.

ITIMER_VIRTUAL

‘sys/time.h’ (BSD): Section 17.3 [Setting an Alarm], page 356.

tcflag_t IXANY

‘termios.h’ (BSD): Section 12.4.4 [Input Modes], page 277.

tcflag_t IXOFF

‘termios.h’ (POSIX.1): Section 12.4.4 [Input Modes], page 277.

tcflag_t IXON

‘termios.h’ (POSIX.1): Section 12.4.4 [Input Modes], page 277.

jmp_buf

‘setjmp.h’ (ISO): Section 20.2 [Details of Non-Local Exits], page 391.

int kill (pid_t pid, int signum)

‘signal.h’ (POSIX.1): Section 21.6.2 [Signaling Another Process], page 428.

int killpg (int pgid, int signum)

‘signal.h’ (BSD): Section 21.6.2 [Signaling Another Process], page 428.

long int labs (long int number)

‘stdlib.h’ (ISO): Section 14.3 [Absolute Value], page 306.

LANG

‘locale.h’ (ISO): Section 19.3 [Categories of Activities that Locales Affect], page 380.

LC_ALL

‘locale.h’ (ISO): Section 19.3 [Categories of Activities that Locales Affect], page 380.

LC_COLLATE

‘locale.h’ (ISO): Section 19.3 [Categories of Activities that Locales Affect], page 380.

LC_CTYPE

‘locale.h’ (ISO): Section 19.3 [Categories of Activities that Locales Affect], page 380.

LC_MESSAGES

‘locale.h’ (XOPEN): Section 19.3 [Categories of Activities that Locales Affect], page 380.

LC_MONETARY

‘locale.h’ (ISO): Section 19.3 [Categories of Activities that Locales Affect], page 380.

LC_NUMERIC

‘locale.h’ (ISO): Section 19.3 [Categories of Activities that Locales Affect], page 380.

int L_ctermid

‘stdio.h’ (POSIX.1): Section 24.7.1 [Identifying the Controlling Terminal], page 497.

LC_TIME

‘locale.h’ (ISO): Section 19.3 [Categories of Activities that Locales Affect], page 380.

int L_cuserid

‘stdio.h’ (POSIX.1): Section 26.11 [Identifying Who Logged In], page 520.

double ldexp (double value, int exponent)

‘math.h’ (ISO): Section 14.4 [Normalization Functions], page 307.

ldiv_t ldiv (long int numerator, long int denominator)

‘stdlib.h’ (ISO): Section 14.6 [Integer Division], page 309.

ldiv_t

‘stdlib.h’ (ISO): Section 14.6 [Integer Division], page 309.

L_INCR

‘sys/file.h’ (BSD): Section 7.15 [File Positioning], page 137.

int LINE_MAX

‘limits.h’ (POSIX.2): Section 28.10 [Utility Program Capacity Limits], page 549.

int link (const char *oldname, const char *newname)

‘unistd.h’ (POSIX.1): Section 9.3 [Hard Links], page 189.

int LINK_MAX

‘limits.h’ (POSIX.1): Section 28.6 [Limits on File System Capacity], page 545.

int listen (int socket, unsigned int n)

‘sys/socket.h’ (BSD): Section 11.8.2 [Listening for Connections], page 248.

struct lconv * localeconv (void)

‘locale.h’ (ISO): Section 19.6 [Numeric Formatting], page 383.

struct tm * localtime (const time_t *time)

‘time.h’ (ISO): Section 17.2.3 [Broken-down Time], page 345.

double log10 (double x)

‘math.h’ (ISO): Section 13.4 [Exponentiation and Logarithms], page 300.

double log1p (double x)

‘math.h’ (BSD): Section 13.4 [Exponentiation and Logarithms], page 300.

double logb (double x)

‘math.h’ (BSD): Section 14.4 [Normalization Functions], page 307.

double log (double x)

‘math.h’ (ISO): Section 13.4 [Exponentiation and Logarithms], page 300.

void longjmp (jmp_buf state, int value)

‘setjmp.h’ (ISO): Section 20.2 [Details of Non-Local Exits], page 391.

LONG_LONG_MAX

‘limits.h’ (GNU): Section A.5.2 [Range of an Integer Type], page 563.

LONG_LONG_MIN

‘limits.h’ (GNU): Section A.5.2 [Range of an Integer Type], page 563.

LONG_MAX

‘limits.h’ (ISO): Section A.5.2 [Range of an Integer Type], page 563.

LONG_MIN

‘limits.h’ (ISO): Section A.5.2 [Range of an Integer Type], page 563.

off_t lseek (int filedes, off_t offset, int whence)

‘unistd.h’ (POSIX.1): Section 8.3 [Setting the File Position of a Descriptor], page 156.

L_SET

‘sys/file.h’ (BSD): Section 7.15 [File Positioning], page 137.

int lstat (const char *filename, struct stat *buf )

‘sys/stat.h’ (BSD): Section 9.8.2 [Reading the Attributes of a File], page 198.

int L_tmpnam

‘stdio.h’ (ISO): Section 9.10 [Temporary Files], page 210.

L_XTND

‘sys/file.h’ (BSD): Section 7.15 [File Positioning], page 137.

struct mallinfo mallinfo (void)

‘malloc.h’ (SVID): Section 3.3.11 [Statistics for Storage Allocation with “codemalloc], page 41.

__malloc_hook

‘malloc.h’ (GNU): Section 3.3.10 [Storage Allocation Hooks], page 40.

void * malloc (size_t size)

‘malloc.h’, ‘stdlib.h’ (ISO): Section 3.3.1 [Basic Storage Allocation], page 32.

int MAX_CANON

‘limits.h’ (POSIX.1): Section 28.6 [Limits on File System Capacity], page 545.

int MAX_INPUT

‘limits.h’ (POSIX.1): Section 28.6 [Limits on File System Capacity], page 545.

int MAXNAMLEN

‘dirent.h’ (BSD): Section 28.6 [Limits on File System Capacity], page 545.

int MB_CUR_MAX

‘stdlib.h’ (ISO): Section 18.3 [Multibyte Characters], page 368.

int mblen (const char *string, size_t size)

‘stdlib.h’ (ISO): Section 18.6 [Multibyte Character Length], page 373.

int MB_LEN_MAX

‘limits.h’ (ISO): Section 18.3 [Multibyte Characters], page 368.

size_t mbstowcs (wchar_t *wstring, const char *string, size_t size)

‘stdlib.h’ (ISO): Section 18.5 [Conversion of Extended Strings], page 372.

int mbtowc (wchar_t *result, const char *string, size_t size)

‘stdlib.h’ (ISO): Section 18.7 [Conversion of Extended Characters One by One], page 374.

int mcheck (void (*abortfn) (enum mcheck_status status))

‘malloc.h’ (GNU): Section 3.3.9 [Heap Consistency Checking], page 38.

tcflag_t MDMBUF

‘termios.h’ (BSD): Section 12.4.6 [Control Modes], page 280.

void * memalign (size_t boundary, size_t size)

‘malloc.h’, ‘stdlib.h’ (BSD): Section 3.3.7 [Allocating Aligned Memory Blocks], page 37.

void * memccpy (void *to, const void *from, int c, size_t size)

‘string.h’ (SVID): Section 5.4 [Copying and Concatenation], page 67.

void * memchr (const void *block, int c, size_t size)

‘string.h’ (ISO): Section 5.7 [Search Functions], page 78.

int memcmp (const void *a1, const void *a2, size_t size)

‘string.h’ (ISO): Section 5.5 [String/Array Comparison], page 72.

void * memcpy (void *to, const void *from, size_t size)

‘string.h’ (ISO): Section 5.4 [Copying and Concatenation], page 67.

void * memmem (const void *haystack, size_t haystack-len,
const void *needle, size_t needle-len)

‘string.h’ (GNU): Section 5.7 [Search Functions], page 78.

void * memmove (void *to, const void *from, size_t size)

‘string.h’ (ISO): Section 5.4 [Copying and Concatenation], page 67.

void * memset (void *block, int c, size_t size)

‘string.h’ (ISO): Section 5.4 [Copying and Concatenation], page 67.

int mkdir (const char *filename, mode_t mode)

‘sys/stat.h’ (POSIX.1): Section 9.7 [Creating Directories], page 194.

int mkfifo (const char *filename, mode_t mode)

‘sys/stat.h’ (POSIX.1): Section 10.3 [FIFO Special Files], page 219.

int mknod (const char *filename, int mode, int dev)

‘sys/stat.h’ (BSD): Section 9.9 [Making Special Files], page 210.

int mkstemp (char *template)

‘unistd.h’ (BSD): Section 9.10 [Temporary Files], page 210.

char * mktemp (char *template)

‘unistd.h’ (Unix): Section 9.10 [Temporary Files], page 210.

time_t mktime (struct tm *brokentime)

‘time.h’ (ISO): Section 17.2.3 [Broken-down Time], page 345.

mode_t

‘sys/types.h’ (POSIX.1): Section 9.8.1 [What the File Attribute Values Mean], page 195.

double modf (double value, double *integer-part)

‘math.h’ (ISO): Section 14.5 [Rounding and Remainder Functions], page 308.

int MSG_DONTROUTE

‘sys/socket.h’ (BSD): Section 11.8.5.3 [Socket Data Options], page 252.

int MSG_OOB

‘sys/socket.h’ (BSD): Section 11.8.5.3 [Socket Data Options], page 252.

int MSG_PEEK

‘sys/socket.h’ (BSD): Section 11.8.5.3 [Socket Data Options], page 252.

int NAME_MAX

‘limits.h’ (POSIX.1): Section 28.6 [Limits on File System Capacity], page 545.

double NAN

‘math.h’ (GNU): Section 14.1 [“Not a NumberV́alues], page 305.

int NCCS

‘termios.h’ (POSIX.1): Section 12.4.1 [Terminal Mode Data Types], page 273.

int NGROUPS_MAX

‘limits.h’ (POSIX.1): Section 28.1 [General Capacity Limits], page 537.

int nice (int increment)

‘dunno.h’ (dunno.h): Section 17.7 [Process Priority], page 364.

nlink_t

‘sys/types.h’ (POSIX.1): Section 9.8.1 [What the File Attribute Values Mean], page 195.

NO_ADDRESS

‘netdb.h’ (BSD): Section 11.5.2.4 [Host Names], page 235.

tcflag_t NOFLSH

‘termios.h’ (POSIX.1): Section 12.4.7 [Local Modes], page 282.

tcflag_t NOKERNINFO

‘termios.h’ (BSD): Section 12.4.7 [Local Modes], page 282.

NO_RECOVERY

‘netdb.h’ (BSD): Section 11.5.2.4 [Host Names], page 235.

int NSIG

‘signal.h’ (BSD): Section 21.2 [Standard Signals], page 395.

unsigned long int ntohl (unsigned long int netlong)

‘netinet/in.h’ (BSD): Section 11.5.5 [Byte Order Conversion], page 239.

unsigned short int ntohs (unsigned short int netshort)

‘netinet/in.h’ (BSD): Section 11.5.5 [Byte Order Conversion], page 239.

void * NULL

‘stddef.h’ (ISO): Section A.3 [Null Pointer Constant], page 561.

int O_ACCMODE

‘fcntl.h’ (POSIX.1): Section 8.10.1 [File Access Modes], page 170.

int O_APPEND

‘fcntl.h’ (POSIX.1): Section 8.10.3 [I/O Operating Modes], page 173.

int O_ASYNC

‘fcntl.h’ (BSD): Section 8.10.3 [I/O Operating Modes], page 173.

void obstack_1grow_fast (struct obstack *obstack-ptr, char c)

‘obstack.h’ (GNU): Section 3.4.7 [Extra Fast Growing Objects], page 49.

void obstack_1grow (struct obstack *obstack-ptr, char c)

‘obstack.h’ (GNU): Section 3.4.6 [Growing Objects], page 47.

int obstack_alignment_mask (struct obstack *obstack-ptr)

‘obstack.h’ (GNU): Section 3.4.9 [Alignment of Data in Obstacks], page 52.

void * obstack_alloc (struct obstack *obstack-ptr, int size)

‘obstack.h’ (GNU): Section 3.4.3 [Allocation in an Obstack], page 45.

void * obstack_base (struct obstack *obstack-ptr)

‘obstack.h’ (GNU): Section 3.4.8 [Status of an Obstack], page 51.

void obstack_blank_fast (struct obstack *obstack-ptr, int size)

‘obstack.h’ (GNU): Section 3.4.7 [Extra Fast Growing Objects], page 49.

void obstack_blank (struct obstack *obstack-ptr, int size)

‘obstack.h’ (GNU): Section 3.4.6 [Growing Objects], page 47.

int obstack_chunk_size (struct obstack *obstack-ptr)

‘obstack.h’ (GNU): Section 3.4.10 [Obstack Chunks], page 52.

void * obstack_copy0 (struct obstack *obstack-ptr, void *address,
int size)

‘obstack.h’ (GNU): Section 3.4.3 [Allocation in an Obstack], page 45.

void * obstack_copy (struct obstack *obstack-ptr, void *address, int

size)

‘obstack.h’ (GNU): Section 3.4.3 [Allocation in an Obstack], page 45.

void * obstack_finish (struct obstack *obstack-ptr)

‘obstack.h’ (GNU): Section 3.4.6 [Growing Objects], page 47.

void obstack_free (struct obstack *obstack-ptr, void *object)

‘obstack.h’ (GNU): Section 3.4.4 [Freeing Objects in an Obstack], page 46.

void obstack_grow0 (struct obstack *obstack-ptr, void *data, int

size)

‘obstack.h’ (GNU): Section 3.4.6 [Growing Objects], page 47.

void obstack_grow (struct obstack *obstack-ptr, void *data, int

size)

‘obstack.h’ (GNU): Section 3.4.6 [Growing Objects], page 47.

int obstack_init (struct obstack *obstack-ptr)

‘obstack.h’ (GNU): Section 3.4.2 [Preparing for Using Obstacks], page 44.

void obstack_int_grow_fast (struct obstack *obstack-ptr, int data)

‘obstack.h’ (GNU): Section 3.4.7 [Extra Fast Growing Objects], page 49.

void obstack_int_grow (struct obstack *obstack-ptr, int data)

‘obstack.h’ (GNU): Section 3.4.6 [Growing Objects], page 47.

void * obstack_next_free (struct obstack *obstack-ptr)

‘obstack.h’ (GNU): Section 3.4.8 [Status of an Obstack], page 51.

int obstack_object_size (struct obstack *obstack-ptr)

‘obstack.h’ (GNU): Section 3.4.6 [Growing Objects], page 47.

int obstack_object_size (struct obstack *obstack-ptr)

‘obstack.h’ (GNU): Section 3.4.8 [Status of an Obstack], page 51.

int obstack_printf (struct obstack *obstack, const char *template,

…)

‘stdio.h’ (GNU): Section 7.10.8 [Dynamically Allocating Formatted Output], page 114.

void obstack_ptr_grow_fast (struct obstack *obstack-ptr, void

*data)

‘obstack.h’ (GNU): Section 3.4.7 [Extra Fast Growing Objects], page 49.

void obstack_ptr_grow (struct obstack *obstack-ptr, void *data)

‘obstack.h’ (GNU): Section 3.4.6 [Growing Objects], page 47.

int obstack_room (struct obstack *obstack-ptr)

‘obstack.h’ (GNU): Section 3.4.7 [Extra Fast Growing Objects], page 49.

int obstack_vprintf (struct obstack *obstack, const char
*template, va_list ap)

‘stdio.h’ (GNU): Section 7.10.9 [Variable Arguments Output Functions], page 115.

int O_CREAT

‘fcntl.h’ (POSIX.1): Section 8.10.2 [Open-time Flags], page 171.

int O_EXCL

‘fcntl.h’ (POSIX.1): Section 8.10.2 [Open-time Flags], page 171.

int O_EXEC

‘fcntl.h’ (GNU): Section 8.10.1 [File Access Modes], page 170.

int O_EXLOCK

‘fcntl.h’ (BSD): Section 8.10.2 [Open-time Flags], page 171.

size_t offsetof (type, member)

‘stddef.h’ (ISO): Section A.5.4 [Structure Field Offset Measurement], page 570.

off_t

‘sys/types.h’ (POSIX.1): Section 8.3 [Setting the File Position of a Descriptor], page 156.

int O_FSYNC

‘fcntl.h’ (BSD): Section 8.10.3 [I/O Operating Modes], page 173.

int O_IGNORE_CTTY

‘fcntl.h’ (GNU): Section 8.10.2 [Open-time Flags], page 171.

int O_NDELAY

‘fcntl.h’ (BSD): Section 8.10.3 [I/O Operating Modes], page 173.

int on_exit (void (*function)(int status, void *arg), void *arg)

‘stdlib.h’ (SunOS): Section 22.3.3 [Cleanups on Exit], page 464.

tcflag_t ONLCR

‘termios.h’ (BSD): Section 12.4.5 [Output Modes], page 279.

int O_NOATIME

‘fcntl.h’ (GNU): Section 8.10.3 [I/O Operating Modes], page 173.

int O_NOCTTY

‘fcntl.h’ (POSIX.1): Section 8.10.2 [Open-time Flags], page 171.

tcflag_t ONOEOT

‘termios.h’ (BSD): Section 12.4.5 [Output Modes], page 279.

int O_NOLINK

‘fcntl.h’ (GNU): Section 8.10.2 [Open-time Flags], page 171.

int O_NONBLOCK

‘fcntl.h’ (POSIX.1): Section 8.10.2 [Open-time Flags], page 171.

int O_NONBLOCK

‘fcntl.h’ (POSIX.1): Section 8.10.3 [I/O Operating Modes], page 173.

int O_NOTRANS

‘fcntl.h’ (GNU): Section 8.10.2 [Open-time Flags], page 171.

DIR * opendir (const char *dirname)

‘dirent.h’ (POSIX.1): Section 9.2.2 [Opening a Directory Stream], page 184.

int open (const char *filename, int flags[, mode_t mode])

‘fcntl.h’ (POSIX.1): Section 8.1 [Opening and Closing Files], page 151.

int OPEN_MAX

‘limits.h’ (POSIX.1): Section 28.1 [General Capacity Limits], page 537.

FILE * open_memstream (char **ptr, size_t *sizeloc)

‘stdio.h’ (GNU): Section 7.18.1 [String Streams], page 144.

FILE * open_obstack_stream (struct obstack *obstack)

‘stdio.h’ (GNU): Section 7.18.2 [Obstack Streams], page 146.

tcflag_t OPOST

‘termios.h’ (POSIX.1): Section 12.4.5 [Output Modes], page 279.

char * optarg

‘unistd.h’ (POSIX.2): Section 22.1.2 [Parsing Program Options], page 449.

int opterr

‘unistd.h’ (POSIX.2): Section 22.1.2 [Parsing Program Options], page 449.

int optind

‘unistd.h’ (POSIX.2): Section 22.1.2 [Parsing Program Options], page 449.

int optopt

‘unistd.h’ (POSIX.2): Section 22.1.2 [Parsing Program Options], page 449.

int O_RDONLY

‘fcntl.h’ (POSIX.1): Section 8.10.1 [File Access Modes], page 170.

int O_RDWR

‘fcntl.h’ (POSIX.1): Section 8.10.1 [File Access Modes], page 170.

int O_READ

‘fcntl.h’ (GNU): Section 8.10.1 [File Access Modes], page 170.

int O_SHLOCK

‘fcntl.h’ (BSD): Section 8.10.2 [Open-time Flags], page 171.

int O_SYNC

‘fcntl.h’ (BSD): Section 8.10.3 [I/O Operating Modes], page 173.

int O_TRUNC

‘fcntl.h’ (POSIX.1): Section 8.10.2 [Open-time Flags], page 171.

int O_WRITE

‘fcntl.h’ (GNU): Section 8.10.1 [File Access Modes], page 170.

int O_WRONLY

‘fcntl.h’ (POSIX.1): Section 8.10.1 [File Access Modes], page 170.

tcflag_t OXTABS

‘termios.h’ (BSD): Section 12.4.5 [Output Modes], page 279.

PA_CHAR

‘printf.h’ (GNU): Section 7.10.10 [Parsing a Template String], page 118.

PA_DOUBLE

‘printf.h’ (GNU): Section 7.10.10 [Parsing a Template String], page 118.

PA_FLAG_LONG_DOUBLE

‘printf.h’ (GNU): Section 7.10.10 [Parsing a Template String], page 118.

PA_FLAG_LONG

‘printf.h’ (GNU): Section 7.10.10 [Parsing a Template String], page 118.

PA_FLAG_LONG_LONG

‘printf.h’ (GNU): Section 7.10.10 [Parsing a Template String], page 118.

int PA_FLAG_MASK

‘printf.h’ (GNU): Section 7.10.10 [Parsing a Template String], page 118.

PA_FLAG_PTR

‘printf.h’ (GNU): Section 7.10.10 [Parsing a Template String], page 118.

PA_FLAG_SHORT

‘printf.h’ (GNU): Section 7.10.10 [Parsing a Template String], page 118.

PA_FLOAT

‘printf.h’ (GNU): Section 7.10.10 [Parsing a Template String], page 118.

PA_INT

‘printf.h’ (GNU): Section 7.10.10 [Parsing a Template String], page 118.

PA_LAST

‘printf.h’ (GNU): Section 7.10.10 [Parsing a Template String], page 118.

PA_POINTER

‘printf.h’ (GNU): Section 7.10.10 [Parsing a Template String], page 118.

tcflag_t PARENB

‘termios.h’ (POSIX.1): Section 12.4.6 [Control Modes], page 280.

tcflag_t PARMRK

‘termios.h’ (POSIX.1): Section 12.4.4 [Input Modes], page 277.

tcflag_t PARODD

‘termios.h’ (POSIX.1): Section 12.4.6 [Control Modes], page 280.

size_t parse_printf_format (const char *template, size_t n, int

*argtypes)

‘printf.h’ (GNU): Section 7.10.10 [Parsing a Template String], page 118.

PA_STRING

‘printf.h’ (GNU): Section 7.10.10 [Parsing a Template String], page 118.

long int pathconf (const char *filename, int parameter)

‘unistd.h’ (POSIX.1): Section 28.9 [Using “codepathconf], page 548.

int PATH_MAX

‘limits.h’ (POSIX.1): Section 28.6 [Limits on File System Capacity], page 545.

int pause ()

‘unistd.h’ (POSIX.1): Section 21.8.1 [Using “codepause], page 439.

_PC_CHOWN_RESTRICTED

‘unistd.h’ (POSIX.1): Section 28.9 [Using “codepathconf], page 548.

_PC_LINK_MAX

‘unistd.h’ (POSIX.1): Section 28.9 [Using “codepathconf], page 548.

int pclose (FILE *stream)

‘stdio.h’ (POSIX.2, SVID, BSD): Section 10.2 [Pipe to a Subprocess], page 217.

_PC_MAX_CANON

‘unistd.h’ (POSIX.1): Section 28.9 [Using “codepathconf], page 548.

_PC_MAX_INPUT

‘unistd.h’ (POSIX.1): Section 28.9 [Using “codepathconf], page 548.

_PC_NAME_MAX

‘unistd.h’ (POSIX.1): Section 28.9 [Using “codepathconf], page 548.

_PC_NO_TRUNC

‘unistd.h’ (POSIX.1): Section 28.9 [Using “codepathconf], page 548.

_PC_PATH_MAX

‘unistd.h’ (POSIX.1): Section 28.9 [Using “codepathconf], page 548.

_PC_PIPE_BUF

‘unistd.h’ (POSIX.1): Section 28.9 [Using “codepathconf], page 548.

_PC_VDISABLE

‘unistd.h’ (POSIX.1): Section 28.9 [Using “codepathconf], page 548.

tcflag_t PENDIN

‘termios.h’ (BSD): Section 12.4.7 [Local Modes], page 282.

void perror (const char *message)

‘stdio.h’ (ISO): Section 2.3 [Error Messages], page 27.

int PF_FILE

‘sys/socket.h’ (GNU): Section 11.4.2 [Details of File Namespace], page 227.

int PF_INET

‘sys/socket.h’ (BSD): Section 11.5 [The Internet Namespace], page 229.

int PF_UNIX

‘sys/socket.h’ (BSD): Section 11.4.2 [Details of File Namespace], page 227.

pid_t

‘sys/types.h’ (POSIX.1): Section 23.3 [Process Identification], page 468.

int PIPE_BUF

‘limits.h’ (POSIX.1): Section 28.6 [Limits on File System Capacity], page 545.

int pipe (int filedes[2])

‘unistd.h’ (POSIX.1): Section 10.1 [Creating a Pipe], page 215.

FILE * popen (const char *command, const char *mode)

‘stdio.h’ (POSIX.2, SVID, BSD): Section 10.2 [Pipe to a Subprocess], page 217.

_POSIX2_BC_BASE_MAX

‘limits.h’ (POSIX.2): Section 28.11 [Minimum Values for Utility Limits], page 550.

_POSIX2_BC_DIM_MAX

‘limits.h’ (POSIX.2): Section 28.11 [Minimum Values for Utility Limits], page 550.

_POSIX2_BC_SCALE_MAX

‘limits.h’ (POSIX.2): Section 28.11 [Minimum Values for Utility Limits], page 550.

_POSIX2_BC_STRING_MAX

‘limits.h’ (POSIX.2): Section 28.11 [Minimum Values for Utility Limits], page 550.

int _POSIX2_C_DEV

‘unistd.h’ (POSIX.2): Section 28.2 [Overall System Options], page 538.

_POSIX2_COLL_WEIGHTS_MAX

‘limits.h’ (POSIX.2): Section 28.11 [Minimum Values for Utility Limits], page 550.

long int _POSIX2_C_VERSION

‘unistd.h’ (POSIX.2): Section 28.3 [Which Version of POSIX is Supported], page 540.

_POSIX2_EQUIV_CLASS_MAX

‘limits.h’ (POSIX.2): Section 28.11 [Minimum Values for Utility Limits], page 550.

_POSIX2_EXPR_NEST_MAX

‘limits.h’ (POSIX.2): Section 28.11 [Minimum Values for Utility Limits], page 550.

int _POSIX2_FORT_DEV

‘unistd.h’ (POSIX.2): Section 28.2 [Overall System Options], page 538.

int _POSIX2_FORT_RUN

‘unistd.h’ (POSIX.2): Section 28.2 [Overall System Options], page 538.

_POSIX2_LINE_MAX

‘limits.h’ (POSIX.2): Section 28.11 [Minimum Values for Utility Limits], page 550.

int _POSIX2_LOCALEDEF

‘unistd.h’ (POSIX.2): Section 28.2 [Overall System Options], page 538.

_POSIX2_RE_DUP_MAX

‘limits.h’ (POSIX.2): Section 28.5 [Minimum Values for General Capacity Limits], page 544.

int _POSIX2_SW_DEV

‘unistd.h’ (POSIX.2): Section 28.2 [Overall System Options], page 538.

_POSIX_ARG_MAX

‘limits.h’ (POSIX.1): Section 28.5 [Minimum Values for General Capacity Limits], page 544.

_POSIX_CHILD_MAX

‘limits.h’ (POSIX.1): Section 28.5 [Minimum Values for General Capacity Limits], page 544.

int _POSIX_CHOWN_RESTRICTED

‘unistd.h’ (POSIX.1): Section 28.7 [Optional Features in File Support], page 546.

_POSIX_C_SOURCE

(POSIX.2): Section 1.3.4 [Feature Test Macros], page 8.

int _POSIX_JOB_CONTROL

‘unistd.h’ (POSIX.1): Section 28.2 [Overall System Options], page 538.

_POSIX_LINK_MAX

‘limits.h’ (POSIX.1): Section 28.8 [Minimum Values for File System Limits], page 547.

_POSIX_MAX_CANON

‘limits.h’ (POSIX.1): Section 28.8 [Minimum Values for File System Limits], page 547.

_POSIX_MAX_INPUT

‘limits.h’ (POSIX.1): Section 28.8 [Minimum Values for File System Limits], page 547.

_POSIX_NAME_MAX

‘limits.h’ (POSIX.1): Section 28.8 [Minimum Values for File System Limits], page 547.

_POSIX_NGROUPS_MAX

‘limits.h’ (POSIX.1): Section 28.5 [Minimum Values for General Capacity Limits], page 544.

int _POSIX_NO_TRUNC

‘unistd.h’ (POSIX.1): Section 28.7 [Optional Features in File Support], page 546.

_POSIX_OPEN_MAX

‘limits.h’ (POSIX.1): Section 28.5 [Minimum Values for General Capacity Limits], page 544.

_POSIX_PATH_MAX

‘limits.h’ (POSIX.1): Section 28.8 [Minimum Values for File System Limits], page 547.

_POSIX_PIPE_BUF

‘limits.h’ (POSIX.1): Section 28.8 [Minimum Values for File System Limits], page 547.

int _POSIX_SAVED_IDS

‘unistd.h’ (POSIX.1): Section 28.2 [Overall System Options], page 538.

_POSIX_SOURCE

(POSIX.1): Section 1.3.4 [Feature Test Macros], page 8.

_POSIX_SSIZE_MAX

‘limits.h’ (POSIX.1): Section 28.5 [Minimum Values for General Capacity Limits], page 544.

_POSIX_STREAM_MAX

‘limits.h’ (POSIX.1): Section 28.5 [Minimum Values for General Capacity Limits], page 544.

_POSIX_TZNAME_MAX

‘limits.h’ (POSIX.1): Section 28.5 [Minimum Values for General Capacity Limits], page 544.

unsigned char _POSIX_VDISABLE

‘unistd.h’ (POSIX.1): Section 28.7 [Optional Features in File Support], page 546.

long int _POSIX_VERSION

‘unistd.h’ (POSIX.1): Section 28.3 [Which Version of POSIX is Supported], page 540.

double pow (double base, double power)

‘math.h’ (ISO): Section 13.4 [Exponentiation and Logarithms], page 300.

printf_arginfo_function

‘printf.h’ (GNU): Section 7.11.3 [Defining the Output Handler], page 123.

printf_function

‘printf.h’ (GNU): Section 7.11.3 [Defining the Output Handler], page 123.

int printf (const char *template, . . .)

‘stdio.h’ (ISO): Section 7.10.7 [Formatted Output Functions], page 113.

PRIO_MAX

‘sys/resource.h’ (BSD): Section 17.7 [Process Priority], page 364.

PRIO_MIN

‘sys/resource.h’ (BSD): Section 17.7 [Process Priority], page 364.

PRIO_PGRP

‘sys/resource.h’ (BSD): Section 17.7 [Process Priority], page 364.

PRIO_PROCESS

‘sys/resource.h’ (BSD): Section 17.7 [Process Priority], page 364.

PRIO_USER

‘sys/resource.h’ (BSD): Section 17.7 [Process Priority], page 364.

char * program_invocation_name

‘errno.h’ (GNU): Section 2.3 [Error Messages], page 27.

char * program_invocation_short_name

‘errno.h’ (GNU): Section 2.3 [Error Messages], page 27.

void psignal (int signum, const char *message)

‘signal.h’ (BSD): Section 21.2.8 [Signal Messages], page 405.

char * P_tmpdir

‘stdio.h’ (SVID): Section 9.10 [Temporary Files], page 210.

ptrdiff_t

‘stddef.h’ (ISO): Section A.4 [Important Data Types], page 561.

int putchar (int c)

‘stdio.h’ (ISO): Section 7.5 [Simple Output by Characters or Lines], page 97.

int putc (int c, FILE *stream)

‘stdio.h’ (ISO): Section 7.5 [Simple Output by Characters or Lines], page 97.

int putenv (const char *string)

‘stdlib.h’ (SVID): Section 22.2.1 [Environment Access], page 459.

int putpwent (const struct passwd *p, FILE *stream)

‘pwd.h’ (SVID): Section 26.12.4 [Writing a User Entry], page 525.

int puts (const char *s)

‘stdio.h’ (ISO): Section 7.5 [Simple Output by Characters or Lines], page 97.

int putw (int w, FILE *stream)

‘stdio.h’ (SVID): Section 7.5 [Simple Output by Characters or Lines], page 97.

void qsort (void *array, size_t count, size_t size, comparison_fn_t

compare)

‘stdlib.h’ (ISO): Section 15.3 [Array Sort Function], page 318.

int raise (int signum)

‘signal.h’ (ISO): Section 21.6.1 [Signaling Yourself], page 426.

void r_alloc_free (void **handleptr)

‘malloc.h’ (GNU): Section 3.6.2 [Allocating and Freeing Relocatable Blocks], page 58.

void * r_alloc (void **handleptr, size_t size)

‘malloc.h’ (GNU): Section 3.6.2 [Allocating and Freeing Relocatable Blocks], page 58.

int rand ()

‘stdlib.h’ (ISO): Section 13.6.1 [ISO C Random Number Functions], page 302.

int RAND_MAX

‘stdlib.h’ (ISO): Section 13.6.1 [ISO C Random Number Functions], page 302.

long int random ()

‘stdlib.h’ (BSD): Section 13.6.2 [BSD Random Number Functions], page 303.

struct dirent * readdir (DIR *dirstream)

‘dirent.h’ (POSIX.1): Section 9.2.3 [Reading and Closing a Directory Stream], page 185.

int readdir_r (DIR *dirstream, struct dirent *entry, struct dirent

**result)

‘dirent.h’ (GNU): Section 9.2.3 [Reading and Closing a Directory Stream], page 185.

ssize_t read (int filedes, void *buffer, size_t size)

‘unistd.h’ (POSIX.1): Section 8.2 [Input and Output Primitives], page 153.

int readlink (const char *filename, char *buffer, size_t size)

‘unistd.h’ (BSD): Section 9.4 [Symbolic Links], page 190.

__realloc_hook

‘malloc.h’ (GNU): Section 3.3.10 [Storage Allocation Hooks], page 40.

void * realloc (void *ptr, size_t newsize)

‘malloc.h’, ‘stdlib.h’ (ISO): Section 3.3.4 [Changing the Size of a Block], page 35.

int recvfrom (int socket, void *buffer, size_t size, int flags, struct
sockaddr *addr, socklen_t *length-ptr)

‘sys/socket.h’ (BSD): Section 11.9.2 [Receiving Datagrams], page 261.

int recv (int socket, void *buffer, size_t size, int flags)

‘sys/socket.h’ (BSD): Section 11.8.5.2 [Receiving Data], page 251.

int recvmsg (int socket, struct msghdr *message, int flags)

‘sys/socket.h’ (BSD): Section 11.9.2 [Receiving Datagrams], page 261.

int RE_DUP_MAX

‘limits.h’ (POSIX.2): Section 28.1 [General Capacity Limits], page 537.

_REENTRANT

(GNU): Section 1.3.4 [Feature Test Macros], page 8.

REG_BADBR

‘regex.h’ (POSIX.2): Section 16.3.1 [POSIX Regular Expression Compilation], page 327.

REG_BADPAT

‘regex.h’ (POSIX.2): Section 16.3.1 [POSIX Regular Expression Compilation], page 327.

REG_BADRPT

‘regex.h’ (POSIX.2): Section 16.3.1 [POSIX Regular Expression Compilation], page 327.

int regcomp (regex_t *compiled, const char *pattern, int cflags)

‘regex.h’ (POSIX.2): Section 16.3.1 [POSIX Regular Expression Compilation], page 327.

REG_EBRACE

‘regex.h’ (POSIX.2): Section 16.3.1 [POSIX Regular Expression Compilation], page 327.

REG_EBRACK

‘regex.h’ (POSIX.2): Section 16.3.1 [POSIX Regular Expression Compilation], page 327.

REG_ECOLLATE

‘regex.h’ (POSIX.2): Section 16.3.1 [POSIX Regular Expression Compilation], page 327.

REG_ECTYPE

‘regex.h’ (POSIX.2): Section 16.3.1 [POSIX Regular Expression Compilation], page 327.

REG_EESCAPE

‘regex.h’ (POSIX.2): Section 16.3.1 [POSIX Regular Expression Compilation], page 327.

REG_EPAREN

‘regex.h’ (POSIX.2): Section 16.3.1 [POSIX Regular Expression Compilation], page 327.

REG_ERANGE

‘regex.h’ (POSIX.2): Section 16.3.1 [POSIX Regular Expression Compilation], page 327.

size_t regerror (int errcode, regex_t *compiled, char *buffer,
size_t length)

‘regex.h’ (POSIX.2): Section 16.3.6 [POSIX Regexp Matching Cleanup], page 332.

REG_ESPACE

‘regex.h’ (POSIX.2): Section 16.3.3 [Matching a Compiled POSIX Regular Expression], page 330.

REG_ESPACE

‘regex.h’ (POSIX.2): Section 16.3.1 [POSIX Regular Expression Compilation], page 327.

REG_ESUBREG

‘regex.h’ (POSIX.2): Section 16.3.1 [POSIX Regular Expression Compilation], page 327.

int regexec (regex_t *compiled, char *string, size_t nmatch,
regmatch_t matchptr [], int eflags)

‘regex.h’ (POSIX.2): Section 16.3.3 [Matching a Compiled POSIX Regular Expression], page 330.

REG_EXTENDED

‘regex.h’ (POSIX.2): Section 16.3.2 [Flags for POSIX Regular Expressions], page 329.

regex_t

‘regex.h’ (POSIX.2): Section 16.3.1 [POSIX Regular Expression Compilation], page 327.

void regfree (regex_t *compiled)

‘regex.h’ (POSIX.2): Section 16.3.6 [POSIX Regexp Matching Cleanup], page 332.

REG_ICASE

‘regex.h’ (POSIX.2): Section 16.3.2 [Flags for POSIX Regular Expressions], page 329.

int register_printf_function (int spec, printf_function

handler-function, printf_arginfo_function arginfo-function)

‘printf.h’ (GNU): Section 7.11.1 [Registering New Conversions], page 121.

regmatch_t

‘regex.h’ (POSIX.2): Section 16.3.4 [Match Results with Subexpressions], page 331.

REG_NEWLINE

‘regex.h’ (POSIX.2): Section 16.3.2 [Flags for POSIX Regular Expressions], page 329.

REG_NOMATCH

‘regex.h’ (POSIX.2): Section 16.3.3 [Matching a Compiled POSIX Regular Expression], page 330.

REG_NOSUB

‘regex.h’ (POSIX.2): Section 16.3.2 [Flags for POSIX Regular Expressions], page 329.

REG_NOTBOL

‘regex.h’ (POSIX.2): Section 16.3.3 [Matching a Compiled POSIX Regular Expression], page 330.

REG_NOTEOL

‘regex.h’ (POSIX.2): Section 16.3.3 [Matching a Compiled POSIX Regular Expression], page 330.

regoff_t

‘regex.h’ (POSIX.2): Section 16.3.4 [Match Results with Subexpressions], page 331.

int remove (const char *filename)

‘stdio.h’ (ISO): Section 9.5 [Deleting Files], page 192.

int rename (const char *oldname, const char *newname)

‘stdio.h’ (ISO): Section 9.6 [Renaming Files], page 193.

void rewinddir (DIR *dirstream)

‘dirent.h’ (POSIX.1): Section 9.2.5 [Random Access in a Directory Stream], page 187.

void rewind (FILE *stream)

‘stdio.h’ (ISO): Section 7.15 [File Positioning], page 137.

char * rindex (const char *string, int c)

‘string.h’ (BSD): Section 5.7 [Search Functions], page 78.

double rint (double x)

‘math.h’ (BSD): Section 14.5 [Rounding and Remainder Functions], page 308.

RLIMIT_CORE

‘sys/resource.h’ (BSD): Section 17.6 [Limiting Resource Usage], page 362.

RLIMIT_CPU

‘sys/resource.h’ (BSD): Section 17.6 [Limiting Resource Usage], page 362.

RLIMIT_DATA

‘sys/resource.h’ (BSD): Section 17.6 [Limiting Resource Usage], page 362.

RLIMIT_FSIZE

‘sys/resource.h’ (BSD): Section 17.6 [Limiting Resource Usage], page 362.

RLIMIT_MEMLOCK

‘sys/resource.h’ (BSD): Section 17.6 [Limiting Resource Usage], page 362.

RLIMIT_NOFILE

‘sys/resource.h’ (BSD): Section 17.6 [Limiting Resource Usage], page 362.

RLIMIT_NPROC

‘sys/resource.h’ (BSD): Section 17.6 [Limiting Resource Usage], page 362.

RLIMIT_RSS

‘sys/resource.h’ (BSD): Section 17.6 [Limiting Resource Usage], page 362.

RLIMIT_STACK

‘sys/resource.h’ (BSD): Section 17.6 [Limiting Resource Usage], page 362.

RLIM_NLIMITS

‘sys/resource.h’ (BSD): Section 17.6 [Limiting Resource Usage], page 362.

int rmdir (const char *filename)

‘unistd.h’ (POSIX.1): Section 9.5 [Deleting Files], page 192.

int R_OK

‘unistd.h’ (POSIX.1): Section 9.8.8 [Testing Permission to Access a File], page 206.

void * r_re_alloc (void **handleptr, size_t size)

‘malloc.h’ (GNU): Section 3.6.2 [Allocating and Freeing Relocatable Blocks], page 58.

RUSAGE_CHILDREN

‘sys/resource.h’ (BSD): Section 17.5 [Resource Usage], page 360.

RUSAGE_SELF

‘sys/resource.h’ (BSD): Section 17.5 [Resource Usage], page 360.

int SA_NOCLDSTOP

‘signal.h’ (POSIX.1): Section 21.3.5 [Flags for “codesigaction], page 412.

int SA_ONSTACK

‘signal.h’ (BSD): Section 21.3.5 [Flags for “codesigaction], page 412.

int SA_RESTART

‘signal.h’ (BSD): Section 21.3.5 [Flags for “codesigaction], page 412.

_SC_2_C_DEV

‘unistd.h’ (POSIX.2): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_2_FORT_DEV

‘unistd.h’ (POSIX.2): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_2_FORT_RUN

‘unistd.h’ (POSIX.2): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_2_LOCALEDEF

‘unistd.h’ (POSIX.2): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_2_SW_DEV

‘unistd.h’ (POSIX.2): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_2_VERSION

‘unistd.h’ (POSIX.2): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

double scalb (double value, int exponent)

‘math.h’ (BSD): Section 14.4 [Normalization Functions], page 307.

int scanf (const char *template, . . .)

‘stdio.h’ (ISO): Section 7.12.8 [Formatted Input Functions], page 134.

_SC_ARG_MAX

‘unistd.h’ (POSIX.1): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_BC_BASE_MAX

‘unistd.h’ (POSIX.2): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_BC_DIM_MAX

‘unistd.h’ (POSIX.2): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_BC_SCALE_MAX

‘unistd.h’ (POSIX.2): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_BC_STRING_MAX

‘unistd.h’ (POSIX.2): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_CHILD_MAX

‘unistd.h’ (POSIX.1): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_CLK_TCK

‘unistd.h’ (POSIX.1): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_COLL_WEIGHTS_MAX

‘unistd.h’ (POSIX.2): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_EQUIV_CLASS_MAX

‘unistd.h’ (POSIX.2): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_EXPR_NEST_MAX

‘unistd.h’ (POSIX.2): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

SCHAR_MAX

‘limits.h’ (ISO): Section A.5.2 [Range of an Integer Type], page 563.

SCHAR_MIN

‘limits.h’ (ISO): Section A.5.2 [Range of an Integer Type], page 563.

_SC_JOB_CONTROL

‘unistd.h’ (POSIX.1): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_LINE_MAX

‘unistd.h’ (POSIX.2): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_NGROUPS_MAX

‘unistd.h’ (POSIX.1): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_OPEN_MAX

‘unistd.h’ (POSIX.1): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_PAGESIZE

‘unistd.h’ (GNU): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_SAVED_IDS

‘unistd.h’ (POSIX.1): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_STREAM_MAX

‘unistd.h’ (POSIX.1): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_TZNAME_MAX

‘unistd.h’ (POSIX.1): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_VERSION

‘unistd.h’ (POSIX.1): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

_SC_VERSION

‘unistd.h’ (POSIX.2): Section 28.4.2 [Constants for “codesysconf Parameters], page 541.

int SEEK_CUR

‘stdio.h’ (ISO): Section 7.15 [File Positioning], page 137.

void seekdir (DIR *dirstream, off_t pos)

‘dirent.h’ (BSD): Section 9.2.5 [Random Access in a Directory Stream], page 187.

int SEEK_END

‘stdio.h’ (ISO): Section 7.15 [File Positioning], page 137.

int SEEK_SET

‘stdio.h’ (ISO): Section 7.15 [File Positioning], page 137.

int select (int nfds, fd_set *read-fds, fd_set *write-fds, fd_set
*except-fds, struct timeval *timeout)

‘sys/types.h’ (BSD): Section 8.6 [Waiting for Input or Output], page 162.

int send (int socket, void *buffer, size_t size, int flags)

‘sys/socket.h’ (BSD): Section 11.8.5.1 [Sending Data], page 250.

int sendmsg (int socket, const struct msghdr *message, int flags)

‘sys/socket.h’ (BSD): Section 11.9.2 [Receiving Datagrams], page 261.

int sendto (int socket, void *buffer. size_t size, int flags, struct
sockaddr *addr, socklen_t length)

‘sys/socket.h’ (BSD): Section 11.9.1 [Sending Datagrams], page 260.

void setbuffer (FILE *stream, char *buf, size_t size)

‘stdio.h’ (BSD): Section 7.17.3 [Controlling Which Kind of Buffering], page 141.

void setbuf (FILE *stream, char *buf )

‘stdio.h’ (ISO): Section 7.17.3 [Controlling Which Kind of Buffering], page 141.

int setgid (gid_t newgid)

‘unistd.h’ (POSIX.1): Section 26.7 [Setting the Group IDs], page 515.

void setgrent (void)

‘grp.h’ (SVID, BSD): Section 26.13.3 [Scanning the List of All Groups], page 527.

int setgroups (size_t count, gid_t *groups)

‘grp.h’ (BSD): Section 26.7 [Setting the Group IDs], page 515.

void sethostent (int stayopen)

‘netdb.h’ (BSD): Section 11.5.2.4 [Host Names], page 235.

int sethostid (long int id)

‘unistd.h’ (BSD): Section 27.1 [Host Identification], page 533.

int sethostname (const char *name, size_t length)

‘unistd.h’ (BSD): Section 27.1 [Host Identification], page 533.

int setitimer (int which, struct itimerval *new, struct itimerval

*old)

‘sys/time.h’ (BSD): Section 17.3 [Setting an Alarm], page 356.

int setjmp (jmp_buf state)

‘setjmp.h’ (ISO): Section 20.2 [Details of Non-Local Exits], page 391.

void setlinebuf (FILE *stream)

‘stdio.h’ (BSD): Section 7.17.3 [Controlling Which Kind of Buffering], page 141.

char * setlocale (int category, const char *locale)

‘locale.h’ (ISO): Section 19.4 [How Programs Set the Locale], page 381.

void setnetent (int stayopen)

‘netdb.h’ (BSD): Section 11.12 [Networks Database], page 269.

int setnetgrent (const char *netgroup)

‘netdb.h’ (netdb.h): Section 26.14.2 [Looking up one Netgroup], page 529.

int setpgid (pid_t pid, pid_t pgid)

‘unistd.h’ (POSIX.1): Section 24.7.2 [Process Group Functions], page 498.

int setpgrp (pid_t pid, pid_t pgid)

‘unistd.h’ (BSD): Section 24.7.2 [Process Group Functions], page 498.

int setpriority (int class, int id, int priority)

‘sys/resource.h’ (BSD): Section 17.7 [Process Priority], page 364.

void setprotoent (int stayopen)

‘netdb.h’ (BSD): Section 11.5.6 [Protocols Database], page 240.

void setpwent (void)

‘pwd.h’ (SVID, BSD): Section 26.12.3 [Scanning the List of All Users], page 523.

int setregid (gid_t rgid, fid_t egid)

‘unistd.h’ (BSD): Section 26.7 [Setting the Group IDs], page 515.

int setreuid (uid_t ruid, uid_t euid)

‘unistd.h’ (BSD): Section 26.6 [Setting the User ID], page 514.

int setrlimit (int resource, struct rlimit *rlp)

‘sys/resource.h’ (BSD): Section 17.6 [Limiting Resource Usage], page 362.

void setservent (int stayopen)

‘netdb.h’ (BSD): Section 11.5.4 [The Services Database], page 238.

pid_t setsid (void)

‘unistd.h’ (POSIX.1): Section 24.7.2 [Process Group Functions], page 498.

int setsockopt (int socket, int level, int optname, void *optval,
socklen_t optlen)

‘sys/socket.h’ (BSD): Section 11.11.1 [Socket Option Functions], page 266.

void * setstate (void *state)

‘stdlib.h’ (BSD): Section 13.6.2 [BSD Random Number Functions], page 303.

int settimeofday (const struct timeval *tp, const struct timezone

*tzp)

‘sys/time.h’ (BSD): Section 17.2.2 [High-Resolution Calendar], page 342.

int setuid (uid_t newuid)

‘unistd.h’ (POSIX.1): Section 26.6 [Setting the User ID], page 514.

int setvbuf (FILE *stream, char *buf, int mode, size_t size)

‘stdio.h’ (ISO): Section 7.17.3 [Controlling Which Kind of Buffering], page 141.

SHRT_MAX

‘limits.h’ (ISO): Section A.5.2 [Range of an Integer Type], page 563.

SHRT_MIN

‘limits.h’ (ISO): Section A.5.2 [Range of an Integer Type], page 563.

int shutdown (int socket, int how)

‘sys/socket.h’ (BSD): Section 11.7.2 [Closing a Socket], page 244.

S_IEXEC

‘sys/stat.h’ (BSD): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

S_IFBLK

‘sys/stat.h’ (BSD): Section 9.8.3 [Testing the Type of a File], page 199.

S_IFCHR

‘sys/stat.h’ (BSD): Section 9.8.3 [Testing the Type of a File], page 199.

S_IFDIR

‘sys/stat.h’ (BSD): Section 9.8.3 [Testing the Type of a File], page 199.

S_IFIFO

‘sys/stat.h’ (BSD): Section 9.8.3 [Testing the Type of a File], page 199.

S_IFLNK

‘sys/stat.h’ (BSD): Section 9.8.3 [Testing the Type of a File], page 199.

int S_IFMT

‘sys/stat.h’ (BSD): Section 9.8.3 [Testing the Type of a File], page 199.

S_IFREG

‘sys/stat.h’ (BSD): Section 9.8.3 [Testing the Type of a File], page 199.

S_IFSOCK

‘sys/stat.h’ (BSD): Section 9.8.3 [Testing the Type of a File], page 199.

int SIGABRT

‘signal.h’ (ISO): Section 21.2.1 [Program Error Signals], page 396.

int sigaction (int signum, const struct sigaction *action, struct
sigaction *old-action)

‘signal.h’ (POSIX.1): Section 21.3.2 [Advanced Signal Handling], page 408.

int sigaddset (sigset_t *set, int signum)

‘signal.h’ (POSIX.1): Section 21.7.2 [Signal Sets], page 432.

int SIGALRM

‘signal.h’ (POSIX.1): Section 21.2.3 [Alarm Signals], page 401.

int sigaltstack (const struct sigaltstack *stack, struct
sigaltstack *oldstack)

‘signal.h’ (BSD): Section 21.9 [Using a Separate Signal Stack], page 442.

sig_atomic_t

‘signal.h’ (ISO): Section 21.4.7.2 [Atomic Types], page 424.

SIG_BLOCK

‘signal.h’ (POSIX.1): Section 21.7.3 [Process Signal Mask], page 433.

int sigblock (int mask)

‘signal.h’ (BSD): Section 21.10.2 [BSD Functions for Blocking Signals], page 446.

int SIGBUS

‘signal.h’ (BSD): Section 21.2.1 [Program Error Signals], page 396.

int SIGCHLD

‘signal.h’ (POSIX.1): Section 21.2.5 [Job Control Signals], page 402.

int SIGCLD

‘signal.h’ (SVID): Section 21.2.5 [Job Control Signals], page 402.

int SIGCONT

‘signal.h’ (POSIX.1): Section 21.2.5 [Job Control Signals], page 402.

int sigdelset (sigset_t *set, int signum)

‘signal.h’ (POSIX.1): Section 21.7.2 [Signal Sets], page 432.

int sigemptyset (sigset_t *set)

‘signal.h’ (POSIX.1): Section 21.7.2 [Signal Sets], page 432.

int SIGEMT

‘signal.h’ (BSD): Section 21.2.1 [Program Error Signals], page 396.

sighandler_t SIG_ERR

‘signal.h’ (ISO): Section 21.3.1 [Basic Signal Handling], page 406.

int sigfillset (sigset_t *set)

‘signal.h’ (POSIX.1): Section 21.7.2 [Signal Sets], page 432.

int SIGFPE

‘signal.h’ (ISO): Section 21.2.1 [Program Error Signals], page 396.

sighandler_t

‘signal.h’ (GNU): Section 21.3.1 [Basic Signal Handling], page 406.

int SIGHUP

‘signal.h’ (POSIX.1): Section 21.2.2 [Termination Signals], page 399.

int SIGILL

‘signal.h’ (ISO): Section 21.2.1 [Program Error Signals], page 396.

int SIGINFO

‘signal.h’ (BSD): Section 21.2.7 [Miscellaneous Signals], page 404.

int siginterrupt (int signum, int failflag)

‘signal.h’ (BSD): Section 21.10.1 [BSD Function to Establish a Handler], page 445.

int SIGINT

‘signal.h’ (ISO): Section 21.2.2 [Termination Signals], page 399.

int SIGIO

‘signal.h’ (BSD): Section 21.2.4 [Asynchronous I/O Signals], page 401.

int SIGIOT

‘signal.h’ (Unix): Section 21.2.1 [Program Error Signals], page 396.

int sigismember (const sigset_t *set, int signum)

‘signal.h’ (POSIX.1): Section 21.7.2 [Signal Sets], page 432.

sigjmp_buf

‘setjmp.h’ (POSIX.1): Section 20.3 [Non-Local Exits and Signals], page 392.

int SIGKILL

‘signal.h’ (POSIX.1): Section 21.2.2 [Termination Signals], page 399.

void siglongjmp (sigjmp_buf state, int value)

‘setjmp.h’ (POSIX.1): Section 20.3 [Non-Local Exits and Signals], page 392.

int SIGLOST

‘signal.h’ (GNU): Section 21.2.6 [Operation Error Signals], page 403.

int sigmask (int signum)

‘signal.h’ (BSD): Section 21.10.2 [BSD Functions for Blocking Signals], page 446.

sighandler_t signal (int signum, sighandler_t action)

‘signal.h’ (ISO): Section 21.3.1 [Basic Signal Handling], page 406.

int sigpause (int mask)

‘signal.h’ (BSD): Section 21.10.2 [BSD Functions for Blocking Signals], page 446.

int sigpending (sigset_t *set)

‘signal.h’ (POSIX.1): Section 21.7.6 [Checking for Pending Signals], page 437.

int SIGPIPE

‘signal.h’ (POSIX.1): Section 21.2.6 [Operation Error Signals], page 403.

int SIGPOLL

‘signal.h’ (SVID): Section 21.2.4 [Asynchronous I/O Signals], page 401.

int sigprocmask (int how, const sigset_t *set, sigset_t *oldset)

‘signal.h’ (POSIX.1): Section 21.7.3 [Process Signal Mask], page 433.

int SIGPROF

‘signal.h’ (BSD): Section 21.2.3 [Alarm Signals], page 401.

int SIGQUIT

‘signal.h’ (POSIX.1): Section 21.2.2 [Termination Signals], page 399.

int SIGSEGV

‘signal.h’ (ISO): Section 21.2.1 [Program Error Signals], page 396.

int sigsetjmp (sigjmp_buf state, int savesigs)

‘setjmp.h’ (POSIX.1): Section 20.3 [Non-Local Exits and Signals], page 392.

SIG_SETMASK

‘signal.h’ (POSIX.1): Section 21.7.3 [Process Signal Mask], page 433.

int sigsetmask (int mask)

‘signal.h’ (BSD): Section 21.10.2 [BSD Functions for Blocking Signals], page 446.

sigset_t

‘signal.h’ (POSIX.1): Section 21.7.2 [Signal Sets], page 432.

int sigstack (const struct sigstack *stack, struct sigstack

*oldstack)

‘signal.h’ (BSD): Section 21.9 [Using a Separate Signal Stack], page 442.

int SIGSTOP

‘signal.h’ (POSIX.1): Section 21.2.5 [Job Control Signals], page 402.

int sigsuspend (const sigset_t *set)

‘signal.h’ (POSIX.1): Section 21.8.3 [Using “codesigsuspend], page 441.

int SIGSYS

‘signal.h’ (Unix): Section 21.2.1 [Program Error Signals], page 396.

int SIGTERM

‘signal.h’ (ISO): Section 21.2.2 [Termination Signals], page 399.

int SIGTRAP

‘signal.h’ (BSD): Section 21.2.1 [Program Error Signals], page 396.

int SIGTSTP

‘signal.h’ (POSIX.1): Section 21.2.5 [Job Control Signals], page 402.

int SIGTTIN

‘signal.h’ (POSIX.1): Section 21.2.5 [Job Control Signals], page 402.

int SIGTTOU

‘signal.h’ (POSIX.1): Section 21.2.5 [Job Control Signals], page 402.

SIG_UNBLOCK

‘signal.h’ (POSIX.1): Section 21.7.3 [Process Signal Mask], page 433.

int SIGURG

‘signal.h’ (BSD): Section 21.2.4 [Asynchronous I/O Signals], page 401.

int SIGUSR1

‘signal.h’ (POSIX.1): Section 21.2.7 [Miscellaneous Signals], page 404.

int SIGUSR2

‘signal.h’ (POSIX.1): Section 21.2.7 [Miscellaneous Signals], page 404.

int sigvec (int signum, const struct sigvec *action,struct sigvec

*old-action)

‘signal.h’ (BSD): Section 21.10.1 [BSD Function to Establish a Handler], page 445.

int SIGVTALRM

‘signal.h’ (BSD): Section 21.2.3 [Alarm Signals], page 401.

int SIGWINCH

‘signal.h’ (BSD): Section 21.2.7 [Miscellaneous Signals], page 404.

int SIGXCPU

‘signal.h’ (BSD): Section 21.2.6 [Operation Error Signals], page 403.

int SIGXFSZ

‘signal.h’ (BSD): Section 21.2.6 [Operation Error Signals], page 403.

double sinh (double x)

‘math.h’ (ISO): Section 13.5 [Hyperbolic Functions], page 301.

double sin (double x)

‘math.h’ (ISO): Section 13.2 [Trigonometric Functions], page 298.

S_IREAD

‘sys/stat.h’ (BSD): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

S_IRGRP

‘sys/stat.h’ (POSIX.1): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

S_IROTH

‘sys/stat.h’ (POSIX.1): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

S_IRUSR

‘sys/stat.h’ (POSIX.1): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

S_IRWXG

‘sys/stat.h’ (POSIX.1): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

S_IRWXO

‘sys/stat.h’ (POSIX.1): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

S_IRWXU

‘sys/stat.h’ (POSIX.1): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

int S_ISBLK (mode_t m)

‘sys/stat.h’ (POSIX): Section 9.8.3 [Testing the Type of a File], page 199.

int S_ISCHR (mode_t m)

‘sys/stat.h’ (POSIX): Section 9.8.3 [Testing the Type of a File], page 199.

int S_ISDIR (mode_t m)

‘sys/stat.h’ (POSIX): Section 9.8.3 [Testing the Type of a File], page 199.

int S_ISFIFO (mode_t m)

‘sys/stat.h’ (POSIX): Section 9.8.3 [Testing the Type of a File], page 199.

S_ISGID

‘sys/stat.h’ (POSIX): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

int S_ISLNK (mode_t m)

‘sys/stat.h’ (GNU): Section 9.8.3 [Testing the Type of a File], page 199.

int S_ISREG (mode_t m)

‘sys/stat.h’ (POSIX): Section 9.8.3 [Testing the Type of a File], page 199.

int S_ISSOCK (mode_t m)

‘sys/stat.h’ (GNU): Section 9.8.3 [Testing the Type of a File], page 199.

S_ISUID

‘sys/stat.h’ (POSIX): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

S_ISVTX

‘sys/stat.h’ (BSD): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

S_IWGRP

‘sys/stat.h’ (POSIX.1): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

S_IWOTH

‘sys/stat.h’ (POSIX.1): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

S_IWRITE

‘sys/stat.h’ (BSD): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

S_IWUSR

‘sys/stat.h’ (POSIX.1): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

S_IXGRP

‘sys/stat.h’ (POSIX.1): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

S_IXOTH

‘sys/stat.h’ (POSIX.1): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

S_IXUSR

‘sys/stat.h’ (POSIX.1): Section 9.8.5 [The Mode Bits for Access Permission], page 202.

size_t

‘stddef.h’ (ISO): Section A.4 [Important Data Types], page 561.

unsigned int sleep (unsigned int seconds)

‘unistd.h’ (POSIX.1): Section 17.4 [Sleeping], page 359.

int snprintf (char *s, size_t size, const char *template, . . .)

‘stdio.h’ (GNU): Section 7.10.7 [Formatted Output Functions], page 113.

SO_BROADCAST

‘sys/socket.h’ (BSD): Section 11.11.2 [Socket-Level Options], page 267.

int SOCK_DGRAM

‘sys/socket.h’ (BSD): Section 11.2 [Communication Styles], page 222.

int socket (int namespace, int style, int protocol)

‘sys/socket.h’ (BSD): Section 11.7.1 [Creating a Socket], page 244.

int socketpair (int namespace, int style, int protocol, int

filedes[2])

‘sys/socket.h’ (BSD): Section 11.7.3 [Socket Pairs], page 245.

int SOCK_RAW

‘sys/socket.h’ (BSD): Section 11.2 [Communication Styles], page 222.

int SOCK_RDM

‘sys/socket.h’ (BSD): Section 11.2 [Communication Styles], page 222.

int SOCK_SEQPACKET

‘sys/socket.h’ (BSD): Section 11.2 [Communication Styles], page 222.

int SOCK_STREAM

‘sys/socket.h’ (BSD): Section 11.2 [Communication Styles], page 222.

SO_DEBUG

‘sys/socket.h’ (BSD): Section 11.11.2 [Socket-Level Options], page 267.

SO_DONTROUTE

‘sys/socket.h’ (BSD): Section 11.11.2 [Socket-Level Options], page 267.

SO_ERROR

‘sys/socket.h’ (BSD): Section 11.11.2 [Socket-Level Options], page 267.

SO_KEEPALIVE

‘sys/socket.h’ (BSD): Section 11.11.2 [Socket-Level Options], page 267.

SO_LINGER

‘sys/socket.h’ (BSD): Section 11.11.2 [Socket-Level Options], page 267.

int SOL_SOCKET

‘sys/socket.h’ (BSD): Section 11.11.2 [Socket-Level Options], page 267.

SO_OOBINLINE

‘sys/socket.h’ (BSD): Section 11.11.2 [Socket-Level Options], page 267.

SO_RCVBUF

‘sys/socket.h’ (BSD): Section 11.11.2 [Socket-Level Options], page 267.

SO_REUSEADDR

‘sys/socket.h’ (BSD): Section 11.11.2 [Socket-Level Options], page 267.

SO_SNDBUF

‘sys/socket.h’ (BSD): Section 11.11.2 [Socket-Level Options], page 267.

SO_STYLE

‘sys/socket.h’ (GNU): Section 11.11.2 [Socket-Level Options], page 267.

SO_TYPE

‘sys/socket.h’ (BSD): Section 11.11.2 [Socket-Level Options], page 267.

speed_t

‘termios.h’ (POSIX.1): Section 12.4.8 [Line Speed], page 284.

int sprintf (char *s, const char *template, . . .)

‘stdio.h’ (ISO): Section 7.10.7 [Formatted Output Functions], page 113.

double sqrt (double x)

‘math.h’ (ISO): Section 13.4 [Exponentiation and Logarithms], page 300.

void srand (unsigned int seed)

‘stdlib.h’ (ISO): Section 13.6.1 [ISO C Random Number Functions], page 302.

void srandom (unsigned int seed)

‘stdlib.h’ (BSD): Section 13.6.2 [BSD Random Number Functions], page 303.

int sscanf (const char *s, const char *template, . . .)

‘stdio.h’ (ISO): Section 7.12.8 [Formatted Input Functions], page 134.

sighandler_t ssignal (int signum, sighandler_t action)

‘signal.h’ (SVID): Section 21.3.1 [Basic Signal Handling], page 406.

int SSIZE_MAX

‘limits.h’ (POSIX.1): Section 28.1 [General Capacity Limits], page 537.

ssize_t

‘unistd.h’ (POSIX.1): Section 8.2 [Input and Output Primitives], page 153.

int stat (const char *filename, struct stat *buf )

‘sys/stat.h’ (POSIX.1): Section 9.8.2 [Reading the Attributes of a File], page 198.

STDERR_FILENO

‘unistd.h’ (POSIX.1): Section 8.4 [Descriptors and Streams], page 159.

FILE * stderr

‘stdio.h’ (ISO): Section 7.2 [Standard Streams], page 93.

STDIN_FILENO

‘unistd.h’ (POSIX.1): Section 8.4 [Descriptors and Streams], page 159.

FILE * stdin

‘stdio.h’ (ISO): Section 7.2 [Standard Streams], page 93.

STDOUT_FILENO

‘unistd.h’ (POSIX.1): Section 8.4 [Descriptors and Streams], page 159.

FILE * stdout

‘stdio.h’ (ISO): Section 7.2 [Standard Streams], page 93.

char * stpcpy (char *to, const char *from)

‘string.h’ (Unknown origin): Section 5.4 [Copying and Concatenation], page 67.

char * stpncpy (char *to, const char *from, size_t size)

‘string.h’ (GNU): Section 5.4 [Copying and Concatenation], page 67.

int strcasecmp (const char *s1, const char *s2)

‘string.h’ (BSD): Section 5.5 [String/Array Comparison], page 72.

char * strcat (char *to, const char *from)

‘string.h’ (ISO): Section 5.4 [Copying and Concatenation], page 67.

char * strchr (const char *string, int c)

‘string.h’ (ISO): Section 5.7 [Search Functions], page 78.

int strcmp (const char *s1, const char *s2)

‘string.h’ (ISO): Section 5.5 [String/Array Comparison], page 72.

int strcoll (const char *s1, const char *s2)

‘string.h’ (ISO): Section 5.6 [Collation Functions], page 75.

char * strcpy (char *to, const char *from)

‘string.h’ (ISO): Section 5.4 [Copying and Concatenation], page 67.

size_t strcspn (const char *string, const char *stopset)

‘string.h’ (ISO): Section 5.7 [Search Functions], page 78.

char * strdupa (const char *s)

‘string.h’ (GNU): Section 5.4 [Copying and Concatenation], page 67.

char * strdup (const char *s)

‘string.h’ (SVID): Section 5.4 [Copying and Concatenation], page 67.

int STREAM_MAX

‘limits.h’ (POSIX.1): Section 28.1 [General Capacity Limits], page 537.

char * strerror (int errnum)

‘string.h’ (ISO): Section 2.3 [Error Messages], page 27.

char * strerror_r (int errnum, char *buf, size_t n)

‘string.h’ (GNU): Section 2.3 [Error Messages], page 27.

size_t strftime (char *s, size_t size, const char *template, const
struct tm *brokentime)

‘time.h’ (POSIX.2): Section 17.2.4 [Formatting Date and Time], page 347.

size_t strlen (const char *s)

‘string.h’ (ISO): Section 5.3 [String Length], page 67.

int strncasecmp (const char *s1, const char *s2, size_t n)

‘string.h’ (BSD): Section 5.5 [String/Array Comparison], page 72.

char * strncat (char *to, const char *from, size_t size)

‘string.h’ (ISO): Section 5.4 [Copying and Concatenation], page 67.

int strncmp (const char *s1, const char *s2, size_t size)

‘string.h’ (ISO): Section 5.5 [String/Array Comparison], page 72.

char * strncpy (char *to, const char *from, size_t size)

‘string.h’ (ISO): Section 5.4 [Copying and Concatenation], page 67.

char * strndupa (const char *s, size_t size)

‘string.h’ (GNU): Section 5.4 [Copying and Concatenation], page 67.

char * strndup (const char *s, size_t size)

‘string.h’ (GNU): Section 5.4 [Copying and Concatenation], page 67.

char * strpbrk (const char *string, const char *stopset)

‘string.h’ (ISO): Section 5.7 [Search Functions], page 78.

char * strrchr (const char *string, int c)

‘string.h’ (ISO): Section 5.7 [Search Functions], page 78.

char * strsep (char **string ptr, const char *delimiter)

‘string.h’ (BSD): Section 5.8 [Finding Tokens in a String], page 80.

char * strsignal (int signum)

‘string.h’ (GNU): Section 21.2.8 [Signal Messages], page 405.

size_t strspn (const char *string, const char *skipset)

‘string.h’ (ISO): Section 5.7 [Search Functions], page 78.

char * strstr (const char *haystack, const char *needle)

‘string.h’ (ISO): Section 5.7 [Search Functions], page 78.

double strtod (const char *string, char **tailptr)

‘stdlib.h’ (ISO): Section 14.7.2 [Parsing of Floats], page 314.

float strtof (const char *string, char **tailptr)

‘stdlib.h’ (GNU): Section 14.7.2 [Parsing of Floats], page 314.

char * strtok (char *newstring, const char *delimiters)

‘string.h’ (ISO): Section 5.8 [Finding Tokens in a String], page 80.

char * strtok_r (char *newstring, const char *delimiters, char

**save ptr)

‘string.h’ (POSIX): Section 5.8 [Finding Tokens in a String], page 80.

long double strtold (const char *string, char **tailptr)

‘stdlib.h’ (GNU): Section 14.7.2 [Parsing of Floats], page 314.

long int strtol (const char *string, char **tailptr, int base)

‘stdlib.h’ (ISO): Section 14.7.1 [Parsing of Integers], page 310.

long long int strtoll (const char *string, char **tailptr, int base)

‘stdlib.h’ (GNU): Section 14.7.1 [Parsing of Integers], page 310.

long long int strtoq (const char *string, char **tailptr, int base)

‘stdlib.h’ (BSD): Section 14.7.1 [Parsing of Integers], page 310.

unsigned long int strtoul (const char *string, char **tailptr, int

base)

‘stdlib.h’ (ISO): Section 14.7.1 [Parsing of Integers], page 310.

unsigned long long int strtoull (const char *string, char **tailptr,
int base)

‘stdlib.h’ (GNU): Section 14.7.1 [Parsing of Integers], page 310.

unsigned long long int strtouq (const char *string, char **tailptr,
int base)

‘stdlib.h’ (BSD): Section 14.7.1 [Parsing of Integers], page 310.

struct dirent

‘dirent.h’ (POSIX.1): Section 9.2.1 [Format of a Directory Entry], page 183.

struct flock

‘fcntl.h’ (POSIX.1): Section 8.11 [File Locks], page 175.

struct group

‘grp.h’ (POSIX.1): Section 26.13.1 [The Data Structure for a Group], page 525.

struct hostent

‘netdb.h’ (BSD): Section 11.5.2.4 [Host Names], page 235.

struct in6_addr

‘netinet/in.h’ (IPv6 basic API): Section 11.5.2.2 [Host Address Data Type], page 232.

struct in_addr

‘netinet/in.h’ (BSD): Section 11.5.2.2 [Host Address Data Type], page 232.

struct itimerval

‘sys/time.h’ (BSD): Section 17.3 [Setting an Alarm], page 356.

struct lconv

‘locale.h’ (ISO): Section 19.6 [Numeric Formatting], page 383.

struct linger

‘sys/socket.h’ (BSD): Section 11.11.2 [Socket-Level Options], page 267.

struct mallinfo

‘malloc.h’ (GNU): Section 3.3.11 [Statistics for Storage Allocation with “codemalloc], page 41.

struct msghdr

‘sys/socket.h’ (BSD): Section 11.9.2 [Receiving Datagrams], page 261.

struct netent

‘netdb.h’ (BSD): Section 11.12 [Networks Database], page 269.

struct obstack

‘obstack.h’ (GNU): Section 3.4.1 [Creating Obstacks], page 43.

struct option

‘getopt.h’ (GNU): Section 22.1.4 [Parsing Long Options], page 452.

struct passwd

‘pwd.h’ (POSIX.1): Section 26.12.1 [The Data Structure that Describes a User], page 521.

struct printf_info

‘printf.h’ (GNU): Section 7.11.2 [Conversion Specifier Options], page 122.

struct protoent

‘netdb.h’ (BSD): Section 11.5.6 [Protocols Database], page 240.

struct rlimit

‘sys/resource.h’ (BSD): Section 17.6 [Limiting Resource Usage], page 362.

struct rusage

‘sys/resource.h’ (BSD): Section 17.5 [Resource Usage], page 360.

struct servent

‘netdb.h’ (BSD): Section 11.5.4 [The Services Database], page 238.

struct sigaction

‘signal.h’ (POSIX.1): Section 21.3.2 [Advanced Signal Handling], page 408.

struct sigaltstack

‘signal.h’ (BSD): Section 21.9 [Using a Separate Signal Stack], page 442.

struct sigstack

‘signal.h’ (BSD): Section 21.9 [Using a Separate Signal Stack], page 442.

struct sigvec

‘signal.h’ (BSD): Section 21.10.1 [BSD Function to Establish a Handler], page 445.

struct sockaddr

‘sys/socket.h’ (BSD): Section 11.3.1 [Address Formats], page 224.

struct sockaddr_in

‘netinet/in.h’ (BSD): Section 11.5.1 [Internet Socket Address Formats], page 230.

struct sockaddr_un

‘sys/un.h’ (BSD): Section 11.4.2 [Details of File Namespace], page 227.

struct stat

‘sys/stat.h’ (POSIX.1): Section 9.8.1 [What the File Attribute Values Mean], page 195.

struct termios

‘termios.h’ (POSIX.1): Section 12.4.1 [Terminal Mode Data Types], page 273.

struct timeval

‘sys/time.h’ (BSD): Section 17.2.2 [High-Resolution Calendar], page 342.

struct timezone

‘sys/time.h’ (BSD): Section 17.2.2 [High-Resolution Calendar], page 342.

struct tm

‘time.h’ (ISO): Section 17.2.3 [Broken-down Time], page 345.

struct tms

‘sys/times.h’ (POSIX.1): Section 17.1.2 [Detailed Elapsed CPU Time Inquiry], page 340.

struct utimbuf

‘time.h’ (POSIX.1): Section 9.8.9 [File Times], page 208.

struct utsname

‘sys/utsname.h’ (POSIX.1): Section 27.2 [Hardware/Software Type Identification], page 534.

size_t strxfrm (char *to, const char *from, size_t size)

‘string.h’ (ISO): Section 5.6 [Collation Functions], page 75.

_SVID_SOURCE

(GNU): Section 1.3.4 [Feature Test Macros], page 8.

int SV_INTERRUPT

‘signal.h’ (BSD): Section 21.10.1 [BSD Function to Establish a Handler], page 445.

int SV_ONSTACK

‘signal.h’ (BSD): Section 21.10.1 [BSD Function to Establish a Handler], page 445.

int SV_RESETHAND

‘signal.h’ (Sun): Section 21.10.1 [BSD Function to Establish a Handler], page 445.

int symlink (const char *oldname, const char *newname)

‘unistd.h’ (BSD): Section 9.4 [Symbolic Links], page 190.

long int sysconf (int parameter)

‘unistd.h’ (POSIX.1): Section 28.4.1 [Definition of “codesysconf], page 541.

int system (const char *command)

‘stdlib.h’ (ISO): Section 23.1 [Running a Command], page 467.

double tanh (double x)

‘math.h’ (ISO): Section 13.5 [Hyperbolic Functions], page 301.

double tan (double x)

‘math.h’ (ISO): Section 13.2 [Trigonometric Functions], page 298.

int tcdrain (int filedes)

‘termios.h’ (POSIX.1): Section 12.5 [Line Control Functions], page 293.

tcflag_t

‘termios.h’ (POSIX.1): Section 12.4.1 [Terminal Mode Data Types], page 273.

int tcflow (int filedes, int action)

‘termios.h’ (POSIX.1): Section 12.5 [Line Control Functions], page 293.

int tcflush (int filedes, int queue)

‘termios.h’ (POSIX.1): Section 12.5 [Line Control Functions], page 293.

int tcgetattr (int filedes, struct termios *termios-p)

‘termios.h’ (POSIX.1): Section 12.4.2 [Terminal Mode Functions], page 274.

pid_t tcgetpgrp (int filedes)

‘unistd.h’ (POSIX.1): Section 24.7.3 [Functions for Controlling Terminal Access], page 499.

TCSADRAIN

‘termios.h’ (POSIX.1): Section 12.4.2 [Terminal Mode Functions], page 274.

TCSAFLUSH

‘termios.h’ (POSIX.1): Section 12.4.2 [Terminal Mode Functions], page 274.

TCSANOW

‘termios.h’ (POSIX.1): Section 12.4.2 [Terminal Mode Functions], page 274.

TCSASOFT

‘termios.h’ (BSD): Section 12.4.2 [Terminal Mode Functions], page 274.

int tcsendbreak (int filedes, int duration)

‘termios.h’ (POSIX.1): Section 12.5 [Line Control Functions], page 293.

int tcsetattr (int filedes, int when, const struct termios

*termios-p)

‘termios.h’ (POSIX.1): Section 12.4.2 [Terminal Mode Functions], page 274.

int tcsetpgrp (int filedes, pid_t pgid)

‘unistd.h’ (POSIX.1): Section 24.7.3 [Functions for Controlling Terminal Access], page 499.

off_t telldir (DIR *dirstream)

‘dirent.h’ (BSD): Section 9.2.5 [Random Access in a Directory Stream], page 187.

TEMP_FAILURE_RETRY (expression)

‘unistd.h’ (GNU): Section 21.5 [Primitives Interrupted by Signals], page 425.

char * tempnam (const char *dir, const char *prefix)

‘stdio.h’ (SVID): Section 9.10 [Temporary Files], page 210.

time_t time (time_t *result)

‘time.h’ (ISO): Section 17.2.1 [Simple Calendar Time], page 342.

clock_t times (struct tms *buffer)

‘sys/times.h’ (POSIX.1): Section 17.1.2 [Detailed Elapsed CPU Time Inquiry], page 340.

time_t

‘time.h’ (ISO): Section 17.2.1 [Simple Calendar Time], page 342.

long int timezone

‘time.h’ (SVID): Section 17.2.6 [Functions and Variables for Time Zones], page 354.

FILE * tmpfile (void)

‘stdio.h’ (ISO): Section 9.10 [Temporary Files], page 210.

int TMP_MAX

‘stdio.h’ (ISO): Section 9.10 [Temporary Files], page 210.

char * tmpnam (char *result)

‘stdio.h’ (ISO): Section 9.10 [Temporary Files], page 210.

char * tmpnam_r (char *result)

‘stdio.h’ (GNU): Section 9.10 [Temporary Files], page 210.

int toascii (int c)

‘ctype.h’ (SVID, BSD): Section 4.2 [Case Conversion], page 63.

int _tolower (int c)

‘ctype.h’ (SVID): Section 4.2 [Case Conversion], page 63.

int tolower (int c)

‘ctype.h’ (ISO): Section 4.2 [Case Conversion], page 63.

tcflag_t TOSTOP

‘termios.h’ (POSIX.1): Section 12.4.7 [Local Modes], page 282.

int _toupper (int c)

‘ctype.h’ (SVID): Section 4.2 [Case Conversion], page 63.

int toupper (int c)

‘ctype.h’ (ISO): Section 4.2 [Case Conversion], page 63.

TRY_AGAIN

‘netdb.h’ (BSD): Section 11.5.2.4 [Host Names], page 235.

char * ttyname (int filedes)

‘unistd.h’ (POSIX.1): Section 12.1 [Identifying Terminals], page 271.

char * tzname [2]

‘time.h’ (POSIX.1): Section 17.2.6 [Functions and Variables for Time Zones], page 354.

int TZNAME_MAX

‘limits.h’ (POSIX.1): Section 28.1 [General Capacity Limits], page 537.

void tzset (void)

‘time.h’ (POSIX.1): Section 17.2.6 [Functions and Variables for Time Zones], page 354.

UCHAR_MAX

‘limits.h’ (ISO): Section A.5.2 [Range of an Integer Type], page 563.

uid_t

‘sys/types.h’ (POSIX.1): Section 26.5 [Reading the Persona of a Process], page 513.

UINT_MAX

‘limits.h’ (ISO): Section A.5.2 [Range of an Integer Type], page 563.

ULONG_LONG_MAX

‘limits.h’ (ISO): Section A.5.2 [Range of an Integer Type], page 563.

ULONG_MAX

‘limits.h’ (ISO): Section A.5.2 [Range of an Integer Type], page 563.

mode_t umask (mode_t mask)

‘sys/stat.h’ (POSIX.1): Section 9.8.7 [Assigning File Permissions], page 204.

int uname (struct utsname *info)

‘sys/utsname.h’ (POSIX.1): Section 27.2 [Hardware/Software Type Identification], page 534.

int ungetc (int c, FILE *stream)

‘stdio.h’ (ISO): Section 7.8.2 [Using “codeungetc To Do Unreading], page 102.

union wait

‘sys/wait.h’ (BSD): Section 23.8 [BSD Process Wait Functions], page 476.

int unlink (const char *filename)

‘unistd.h’ (POSIX.1): Section 9.5 [Deleting Files], page 192.

USHRT_MAX

‘limits.h’ (ISO): Section A.5.2 [Range of an Integer Type], page 563.

int utime (const char *filename, const struct utimbuf *times)

‘time.h’ (POSIX.1): Section 9.8.9 [File Times], page 208.

int utimes (const char *filename, struct timeval tvp[2])

‘sys/time.h’ (BSD): Section 9.8.9 [File Times], page 208.

va_alist

‘varargs.h’ (Unix): Section A.2.3.1 [Old-Style Variadic Functions], page 560.

type va_arg (va_list ap, type)

‘stdarg.h’ (ISO): Section A.2.2.5 [Argument Access Macros], page 559.

va_dcl

‘varargs.h’ (Unix): Section A.2.3.1 [Old-Style Variadic Functions], page 560.

void va_end (va_list ap)

‘stdarg.h’ (ISO): Section A.2.2.5 [Argument Access Macros], page 559.

va_list

‘stdarg.h’ (ISO): Section A.2.2.5 [Argument Access Macros], page 559.

void * valloc (size_t size)

‘malloc.h’, ‘stdlib.h’ (BSD): Section 3.3.7 [Allocating Aligned Memory Blocks], page 37.

int vasprintf (char **ptr, const char *template, va_list ap)

‘stdio.h’ (GNU): Section 7.10.9 [Variable Arguments Output Functions], page 115.

void va_start (va_list ap)

‘varargs.h’ (Unix): Section A.2.3.1 [Old-Style Variadic Functions], page 560.

void va_start (va_list ap, last-required)

‘stdarg.h’ (ISO): Section A.2.2.5 [Argument Access Macros], page 559.

int VDISCARD

‘termios.h’ (BSD): Section 12.4.9.4 [Other Special Characters], page 290.

int VDSUSP

‘termios.h’ (BSD): Section 12.4.9.2 [Characters that Cause Signals], page 288.

int VEOF

‘termios.h’ (POSIX.1): Section 12.4.9.1 [Characters for Input Editing], page 287.

int VEOL2

‘termios.h’ (BSD): Section 12.4.9.1 [Characters for Input Editing], page 287.

int VEOL

‘termios.h’ (POSIX.1): Section 12.4.9.1 [Characters for Input Editing], page 287.

int VERASE

‘termios.h’ (POSIX.1): Section 12.4.9.1 [Characters for Input Editing], page 287.

pid_t vfork (void)

‘unistd.h’ (BSD): Section 23.4 [Creating a Process], page 469.

int vfprintf (FILE *stream, const char *template, va_list ap)

‘stdio.h’ (ISO): Section 7.10.9 [Variable Arguments Output Functions], page 115.

int vfscanf (FILE *stream, const char *template, va_list ap)

‘stdio.h’ (GNU): Section 7.12.9 [Variable Arguments Input Functions], page 134.

int VINTR

‘termios.h’ (POSIX.1): Section 12.4.9.2 [Characters that Cause Signals], page 288.

int VKILL

‘termios.h’ (POSIX.1): Section 12.4.9.1 [Characters for Input Editing], page 287.

int VLNEXT

‘termios.h’ (BSD): Section 12.4.9.4 [Other Special Characters], page 290.

int VMIN

‘termios.h’ (POSIX.1): Section 12.4.10 [Noncanonical Input], page 291.

int vprintf (const char *template, va_list ap)

‘stdio.h’ (ISO): Section 7.10.9 [Variable Arguments Output Functions], page 115.

int VQUIT

‘termios.h’ (POSIX.1): Section 12.4.9.2 [Characters that Cause Signals], page 288.

int VREPRINT

‘termios.h’ (BSD): Section 12.4.9.1 [Characters for Input Editing], page 287.

int vscanf (const char *template, va_list ap)

‘stdio.h’ (GNU): Section 7.12.9 [Variable Arguments Input Functions], page 134.

int vsnprintf (char *s, size_t size, const char *template, va_list

ap)

‘stdio.h’ (GNU): Section 7.10.9 [Variable Arguments Output Functions], page 115.

int vsprintf (char *s, const char *template, va_list ap)

‘stdio.h’ (ISO): Section 7.10.9 [Variable Arguments Output Functions], page 115.

int vsscanf (const char *s, const char *template, va_list ap)

‘stdio.h’ (GNU): Section 7.12.9 [Variable Arguments Input Functions], page 134.

int VSTART

‘termios.h’ (POSIX.1): Section 12.4.9.3 [Special Characters for Flow Control], page 290.

int VSTATUS

‘termios.h’ (BSD): Section 12.4.9.4 [Other Special Characters], page 290.

int VSTOP

‘termios.h’ (POSIX.1): Section 12.4.9.3 [Special Characters for Flow Control], page 290.

int VSUSP

‘termios.h’ (POSIX.1): Section 12.4.9.2 [Characters that Cause Signals], page 288.

int VTIME

‘termios.h’ (POSIX.1): Section 12.4.10 [Noncanonical Input], page 291.

int VWERASE

‘termios.h’ (BSD): Section 12.4.9.1 [Characters for Input Editing], page 287.

pid_t wait3 (union wait *status-ptr, int options, struct rusage

*usage)

‘sys/wait.h’ (BSD): Section 23.8 [BSD Process Wait Functions], page 476.

pid_t wait4 (pid_t pid, int *status-ptr, int options, struct rusage

*usage)

‘sys/wait.h’ (BSD): Section 23.6 [Process Completion], page 473.

pid_t wait (int *status-ptr)

‘sys/wait.h’ (POSIX.1): Section 23.6 [Process Completion], page 473.

pid_t waitpid (pid_t pid, int *status-ptr, int options)

‘sys/wait.h’ (POSIX.1): Section 23.6 [Process Completion], page 473.

WCHAR_MAX

‘limits.h’ (GNU): Section A.5.2 [Range of an Integer Type], page 563.

wchar_t

‘stddef.h’ (ISO): Section 18.4 [Wide Character Introduction], page 371.

int WCOREDUMP (int status)

‘sys/wait.h’ (BSD): Section 23.7 [Process Completion Status], page 475.

size_t wcstombs (char *string, const wchar_t wstring, size_t size)

‘stdlib.h’ (ISO): Section 18.5 [Conversion of Extended Strings], page 372.

int wctomb (char *string, wchar_t wchar)

‘stdlib.h’ (ISO): Section 18.7 [Conversion of Extended Characters One by One], page 374.

int WEXITSTATUS (int status)

‘sys/wait.h’ (POSIX.1): Section 23.7 [Process Completion Status], page 475.

int WIFEXITED (int status)

‘sys/wait.h’ (POSIX.1): Section 23.7 [Process Completion Status], page 475.

int WIFSIGNALED (int status)

‘sys/wait.h’ (POSIX.1): Section 23.7 [Process Completion Status], page 475.

int WIFSTOPPED (int status)

‘sys/wait.h’ (POSIX.1): Section 23.7 [Process Completion Status], page 475.

int W_OK

‘unistd.h’ (POSIX.1): Section 9.8.8 [Testing Permission to Access a File], page 206.

int wordexp (const char *words, wordexp_t *word-vector-ptr, int

flags)

‘wordexp.h’ (POSIX.2): Section 16.4.2 [Calling “codewordexp], page 334.

wordexp_t

‘wordexp.h’ (POSIX.2): Section 16.4.2 [Calling “codewordexp], page 334.

void wordfree (wordexp_t *word-vector-ptr)

‘wordexp.h’ (POSIX.2): Section 16.4.2 [Calling “codewordexp], page 334.

WRDE_APPEND

‘wordexp.h’ (POSIX.2): Section 16.4.3 [Flags for Word Expansion], page 336.

WRDE_BADCHAR

‘wordexp.h’ (POSIX.2): Section 16.4.2 [Calling “codewordexp], page 334.

WRDE_BADVAL

‘wordexp.h’ (POSIX.2): Section 16.4.2 [Calling “codewordexp], page 334.

WRDE_CMDSUB

‘wordexp.h’ (POSIX.2): Section 16.4.2 [Calling “codewordexp], page 334.

WRDE_DOOFFS

‘wordexp.h’ (POSIX.2): Section 16.4.3 [Flags for Word Expansion], page 336.

WRDE_NOCMD

‘wordexp.h’ (POSIX.2): Section 16.4.3 [Flags for Word Expansion], page 336.

WRDE_NOSPACE

‘wordexp.h’ (POSIX.2): Section 16.4.2 [Calling “codewordexp], page 334.

WRDE_REUSE

‘wordexp.h’ (POSIX.2): Section 16.4.3 [Flags for Word Expansion], page 336.

WRDE_SHOWERR

‘wordexp.h’ (POSIX.2): Section 16.4.3 [Flags for Word Expansion], page 336.

WRDE_SYNTAX

‘wordexp.h’ (POSIX.2): Section 16.4.2 [Calling “codewordexp], page 334.

WRDE_UNDEF

‘wordexp.h’ (POSIX.2): Section 16.4.3 [Flags for Word Expansion], page 336.

ssize_t write (int filedes, const void *buffer, size_t size)

‘unistd.h’ (POSIX.1): Section 8.2 [Input and Output Primitives], page 153.

int WSTOPSIG (int status)

‘sys/wait.h’ (POSIX.1): Section 23.7 [Process Completion Status], page 475.

int WTERMSIG (int status)

‘sys/wait.h’ (POSIX.1): Section 23.7 [Process Completion Status], page 475.

int X_OK

‘unistd.h’ (POSIX.1): Section 9.8.8 [Testing Permission to Access a File], page 206.

_XOPEN_SOURCE

(XOPEN): Section 1.3.4 [Feature Test Macros], page 8.

Appendix C Library Maintenance

C.1 How to Install the GNU C Library

Installation of the GNU C library is relatively simple, but usually requires several GNU tools to be installed already. (see Section C.1.1 [Recommended Tools to Install the GNU C Library], page 669, below.)

To configure the GNU C library for your system, run the shell script ‘configure’ with sh. Use an argument which is the conventional GNU name for your system configuration—for example, ‘sparc-sun-sunos4.1’, for a Sun 4 running SunOS 4.1. See section “Installing GNU CC” in Using and Porting GNU CC, for a full description of standard GNU configuration names. If you omit the configuration name, ‘configure’ will try to guess one for you by inspecting the system it is running on. It may or may not be able to come up with a guess, and the its guess might be wrong. ‘configure’ will tell you the canonical name of the chosen configuration before proceeding.

Here are some options that you should specify (if appropriate) when you

run configure:
‘--with-gnu-ld’

Use this option if you plan to use GNU ld to link programs with the GNU C Library. (We strongly recommend that you do.) This option enables use of features that exist only in GNU ld; so if you configure for GNU ld you must use GNU ld every time you link with the GNU C Library, and when building it.

‘--with-gnu-as’

Use this option if you plan to use the GNU assembler, gas, when building the GNU C Library. On some systems, the library may not build properly if you do not use gas.

‘--with-gnu-binutils’

This option implies both ‘--with-gnu-ld’ and ‘--with-gnu-as’. On systems where GNU tools are the system tools, there is no need to specify this option. These include GNU, GNU/Linux, and free BSD systems.

‘--without-fp’
‘--nfp’

Use this option if your computer lacks hardware floating-point support.

‘--prefix=directory’ Install machine-independent data files in subdirectories of ‘directory’. (You can also set this in ‘configparms’; see below.)

‘--exec-prefix=directory’

Install the library and other machine-dependent files in subdirectories of ‘directory’. (You can also set this in ‘configparms’; see below.)

‘--enable-shared’
‘--disable-shared’

Enable or disable building of an ELF shared library on systems that support it. The default is to build the shared library on systems using ELF when the GNU binutils are available.

‘--enable-profile’
‘--disable-profile’

Enable or disable building of the profiled C library, ‘-lc_p’. The default is to build the profiled library. You may wish to disable it if you don’t plan to do profiling, because it doubles the build time of compiling just the unprofiled static library.

‘--enable-omitfp’

Enable building a highly-optimized but possibly undebuggable static C library. This causes the normal static and shared (if enabled) C libraries to be compiled with maximal optimization, including the ‘-fomit-frame-pointer’ switch that makes debugging impossible on many machines, and without debugging information (which makes the binaries substantially smaller). An additional static library is compiled with no optimization and full debugging information, and installed as ‘-lc_g’.

The simplest way to run configure is to do it in the directory that contains the library sources. This prepares to build the library in that very directory.

You can prepare to build the library in some other directory by going to that other directory to run configure. In order to run configure, you will have to specify a directory for it, like this:

     mkdir sun4
     cd sun4
     ../configure sparc-sun-sunos4.1

configure looks for the sources in whatever directory you specified for finding configure itself. It does not matter where in the file system the source and build directories are—as long as you specify the source directory when you run configure, you will get the proper results.

This feature lets you keep sources and binaries in different directories, and that makes it easy to build the library for several different machines from the same set of sources. Simply create a build directory for each target machine, and run configure in that directory specifying the target machine’s configuration name.

The library has a number of special-purpose configuration parameters. These are defined in the file ‘Makeconfig’; see the comments in that file for the details.

But don’t edit the file ‘Makeconfig’ yourself—instead, create a file ‘configparms’ in the directory where you are building the library, and define in that file the parameters you want to specify. ‘configparms’ should not be an edited copy of ‘Makeconfig’; specify only the parameters that you want to override. To see how to set these parameters, find the section of ‘Makeconfig’ that says “These are the configuration variables.” Then for each parameter that you want to change, copy the definition from ‘Makeconfig’ to your new ‘configparms’ file, and change the value as appropriate for your system.

It is easy to configure the GNU C library for cross-compilation by setting a few variables in ‘configparms’. Set CC to the cross-compiler for the target you configured the library for; it is important to use this same CC value when running configure, like this: ‘CC=target-gcc configure target’. Set BUILD_CC to the compiler to use for for programs run on the build system as part of compiling the library. You may need to set AR and RANLIB to crosscompiling versions of ar and ranlib if the native tools are not configured to work with object files for the target you configured for.

Some of the machine-dependent code for some machines uses extensions in the GNU C compiler, so you may need to compile the library with GCC. (In fact, all of the existing complete ports require GCC.)

To build the library and related programs, type make. This will produce a lot of output, some of which may look like errors from make (but isn’t). Look for error messages from make containing ‘***’. Those indicate that something is really wrong.

To build and run some test programs which exercise some of the library facilities, type make check. This will produce several files with names like ‘program.out’.

To format the GNU C Library Reference Manual for printing, type make dvi. You need a working TeX installation to do this.

To install the library and its header files, and the Info files of the manual, type make install. This will build things if necessary, before installing them. If you want to install the files in a different place than the one specified at configuration time you can specify a value for the Makefile variable install_root on the command line. This is useful to create chroot’ed environment or to prepare binary releases.

C.1.1 Recommended Tools to Install the GNU C Library

We recommend installing the following GNU tools before attempting to build the GNU C library:

• make 3.75 You need the latest version of GNU make. Modifying the GNU C Library to work with other make programs would be so hard that we recommend you port GNU make instead. Really. We recommend version GNU make version 3.75. Versions 3.76 and 3.76.1 are known to have bugs which only show up in big projects like GNU libc.

• GCC 2.7.2 On most platforms, the GNU C library can only be compiled with the GNU C compiler. We recommend GCC version 2.7.2 or later; earlier versions may have problems.

• binutils 2.8.1

Using the GNU binutils (assembler, linker, and related tools) is preferable when possible, and they are required to build an ELF shared C library. We recommend binutils version 2.8 or later; earlier versions are known to have problems or to not support all architectures.

• texinfo 3.11

To correctly translate and install the Texinfo documentation you need this version of the texinfo package. Former versions did not understand all the tags used in the document and also the installation mechanisms for the info files was not present or worked differently. On some Debian Linux based systems the used install-info program works differently. Here you have to run make like this:

     make INSTALL_INFO=/path/to/GNU/install-info install

C.1.2 Supported Configurations

The GNU C Library currently supports configurations that match the following patterns:

alpha-anything-linux

ix86-anything-gnu

ix86-anything-linux
m68k-anything-linux

Former releases of this library (version 1.09.1 and perhaps earlier versions) used to run on the following configurations:

alpha-dec-osf1
ix86-anything-bsd4.3
ix86-anything-isc2.2
ix86-anything-isc3.n
ix86-anything-sco3.2
ix86-anything-sco3.2v4
ix86-anything-sysv
ix86-anything-sysv4

     ix86-force_cpu386-none
     ix86-sequent-bsd
     i960-nindy960-none
     m68k-hp-bsd4.3
     m68k-mvme135-none
     m68k-mvme136-none
     m68k-sony-newsos3
     m68k-sony-newsos4
     m68k-sun-sunos4.n
     mips-dec-ultrix4.n
     mips-sgi-irix4.n
     sparc-sun-solaris2.n
     sparc-sun-sunos4.n

Since no one has volunteered to test and fix the above configurations, these are not supported at the moment. It’s expected that these don’t work anymore. Porting the library is not hard. If you are interested in doing a port, please contact the glibc maintainers by sending electronic mail to

bug-glibc@prep.ai.mit.edu.

Each case of ‘ix86’ can be ‘i386’, ‘i486’, ‘i586’, or ‘i686’. All of those configurations produce a library that can run on any of these processors. The library will be optimized for the specified processor, but will not use instructions not available on all of them.

While no other configurations are supported, there are handy aliases for these few. (These aliases work in other GNU software as well.)

     decstation
     hp320-bsd4.3 hp300bsd
     i486-gnu
     i586-linux
     i386-sco
     i386-sco3.2v4
     i386-sequent-dynix
     i386-svr4
     news
     sun3-sunos4.n sun3
     sun4-solaris2.n sun4-sunos5.n
     sun4-sunos4.n sun4

C.2 Reporting Bugs

There are probably bugs in the GNU C library. There are certainly errors and omissions in this manual. If you report them, they will get fixed. If you don’t, no one will ever know about them and they will remain unfixed for all eternity, if not longer.

To report a bug, first you must find it. Hopefully, this will be the hard part. Once you’ve found a bug, make sure it’s really a bug. A good way to do this is to see if the GNU C library behaves the same way some other C library does. If so, probably you are wrong and the libraries are right (but not necessarily). If not, one of the libraries is probably wrong.

Once you’re sure you’ve found a bug, try to narrow it down to the smallest test case that reproduces the problem. In the case of a C library, you really only need to narrow it down to one library function call, if possible. This should not be too difficult.

The final step when you have a simple test case is to report the bug. When reporting a bug, send your test case, the results you got, the results you expected, what you think the problem might be (if you’ve thought of anything), your system type, and the version of the GNU C library which you are using. Also include the files ‘config.status’ and ‘config.make’ which are created by running ‘configure’; they will be in whatever directory was current when you ran ‘configure’.

If you think you have found some way in which the GNU C library does not conform to the ISO and POSIX standards (see Section 1.2 [Standards and Portability], page 1), that is definitely a bug. Report it!

Send bug reports to the Internet address bug-glibc@prep.ai.mit.edu

or the UUCP path mit-eddie!prep.ai.mit.edu!bug-glibc. If you have

other problems with installation or use, please report those as well.

If you are not sure how a function should behave, and this manual doesn’t tell you, that’s a bug in the manual. Report that too! If the function’s behavior disagrees with the manual, then either the library or the manual has a bug, so report the disagreement. If you find any errors or omissions in this manual, please report them to the In-

ternet address bug-glibc-manual@prep.ai.mit.edu or the UUCP path
mit-eddie!prep.ai.mit.edu!bug-glibc-manual.

C.3 Adding New Functions

The process of building the library is driven by the makefiles, which make heavy use of special features of GNU make. The makefiles are very complex, and you probably don’t want to try to understand them. But what they do is fairly straightforward, and only requires that you define a few variables in the right places.

The library sources are divided into subdirectories, grouped by topic. The ‘string’ subdirectory has all the string-manipulation functions,

‘math’ has all the mathematical functions, etc.

Each subdirectory contains a simple makefile, called ‘Makefile’, which defines a few make variables and then includes the global makefile ‘Rules’ with a line like:

include ../Rules

The basic variables that a subdirectory makefile defines are:

subdir The name of the subdirectory, for example ‘stdio’. This variable must be defined.

headers The names of the header files in this section of the library, such

as ‘stdio.h’.
routines

aux The names of the modules (source files) in this section of the library. These should be simple names, such as ‘strlen’ (rather than complete file names, such as ‘strlen.c’). Use routines for modules that define functions in the library, and aux for auxiliary modules containing things like data definitions. But the values of routines and aux are just concatenated, so there really is no practical difference.

tests The names of test programs for this section of the library. These should be simple names, such as ‘tester’ (rather than complete file names, such as ‘tester.c’). ‘make tests’ will build and run all the test programs. If a test program needs input, put the test data in a file called ‘test-program.input’; it will be given to the test program on its standard input. If a test program wants to be run with arguments, put the arguments (all on a single line) in a file called ‘test-program.args’. Test programs should exit with zero status when the test passes, and nonzero status when the test indicates a bug in the library or error in building.

others The names of “other” programs associated with this section of the library. These are programs which are not tests per se, but are other small programs included with the library. They are

built by ‘make others’.
install-lib
install-data

install Files to be installed by ‘make install’. Files listed in ‘install-lib’ are installed in the directory specified by

‘libdir’ in ‘configparms’ or ‘Makeconfig’ (see Section C.1

[How to Install the GNU C Library], page 667). Files listed in install-data are installed in the directory specified by

‘datadir’ in ‘configparms’ or ‘Makeconfig’. Files listed in

install are installed in the directory specified by ‘bindir’ in

‘configparms’ or ‘Makeconfig’.
distribute

Other files from this subdirectory which should be put into a distribution tar file. You need not list here the makefile itself or the source and header files listed in the other standard variables. Only define distribute if there are files used in an unusual way that should go into the distribution.

generated

Files which are generated by ‘Makefile’ in this subdirectory. These files will be removed by ‘make clean’, and they will never go into a distribution.

extra-objs

Extra object files which are built by ‘Makefile’ in this subdirectory. This should be a list of file names like ‘foo.o’; the files will actually be found in whatever directory object files are being built in. These files will be removed by ‘make clean’. This variable is used for secondary object files needed to build others

          or tests.

C.4 Porting the GNU C Library

The GNU C library is written to be easily portable to a variety of machines and operating systems. Machine- and operating system-dependent functions are well separated to make it easy to add implementations for new machines or operating systems. This section describes the layout of the library source tree and explains the mechanisms used to select machinedependent code to use.

All the machine-dependent and operating system-dependent files in the library are in the subdirectory ‘sysdeps’ under the top-level library source directory. This directory contains a hierarchy of subdirectories (see Section C.4.1 [Layout of the “filesysdeps Directory Hierarchy], page 677).

Each subdirectory of ‘sysdeps’ contains source files for a particular machine or operating system, or for a class of machine or operating system (for example, systems by a particular vendor, or all machines that use IEEE 754 floating-point format). A configuration specifies an ordered list of these subdirectories. Each subdirectory implicitly appends its parent directory to the list. For example, specifying the list ‘unix/bsd/vax’ is equivalent to specifying the list ‘unix/bsd/vax unix/bsd unix’. A subdirectory can also specify that it implies other subdirectories which are not directly above it in the directory hierarchy. If the file ‘Implies’ exists in a subdirectory, it lists other subdirectories of ‘sysdeps’ which are appended to the list, appearing after the subdirectory containing the ‘Implies’ file. Lines in an ‘Implies’ file that begin with a ‘#’ character are ignored as comments. For example,

‘unix/bsd/Implies’ contains:
     # BSD has Internet-related things.
     unix/inet

and ‘unix/Implies’ contains:

     posix
So the final list is ‘unix/bsd/vax unix/bsd unix/inet unix posix’.

‘sysdeps’ has two “special” subdirectories, called ‘generic’ and ‘stub’. These two are always implicitly appended to the list of subdirectories (in that order), so you needn’t put them in an ‘Implies’ file, and you should not create any subdirectories under them intended to be new specific categories. ‘generic’ is for things that can be implemented in machine-independent C, using only other machine-independent functions in the C library. ‘stub’ is for stub versions of functions which cannot be implemented on a particular machine or operating system. The stub functions always return an error, and set errno to ENOSYS (Function not implemented). See Chapter 2 [Error Reporting], page 15.

A source file is known to be system-dependent by its having a version in ‘generic’ or ‘stub’; every generally-available function whose implementation is system-dependent in should have either a generic or stub implementation (there is no point in having both). Some rare functions are only useful on specific systems and aren’t defined at all on others; these do not appear anywhere in the system-independent source code or makefiles (including the ‘generic’ and ‘stub’ directories), only in the system-dependent ‘Makefile’ in the specific system’s subdirectory.

If you come across a file that is in one of the main source directories (‘string’, ‘stdio’, etc.), and you want to write a machine- or operating system-dependent version of it, move the file into ‘sysdeps/generic’ and write your new implementation in the appropriate system-specific subdirectory. Note that if a file is to be system-dependent, it must not appear in one of the main source directories.

There are a few special files that may exist in each subdirectory of

‘sysdeps’:
‘Makefile’

A makefile for this machine or operating system, or class of machine or operating system. This file is included by the library makefile ‘Makerules’, which is used by the top-level makefile and the subdirectory makefiles. It can change the variables set in the including makefile or add new rules. It can use GNU make conditional directives based on the variable ‘subdir’ (see above) to select different sets of variables and rules for different sections of the library. It can also set the make variable ‘sysdep-routines’, to specify extra modules to be included in the library. You should use ‘sysdep-routines’ rather than adding modules to ‘routines’ because the latter is used in determining what to distribute for each subdirectory of the main source tree.

Each makefile in a subdirectory in the ordered list of subdirectories to be searched is included in order. Since several systemdependent makefiles may be included, each should append to ‘sysdep-routines’ rather than simply setting it:

     sysdep-routines := $(sysdep-routines) foo bar
‘Subdirs’

This file contains the names of new whole subdirectories under the top-level library source tree that should be included for this system. These subdirectories are treated just like the systemindependent subdirectories in the library source tree, such as

‘stdio’ and ‘math’.

Use this when there are completely new sets of functions and header files that should go into the library for the system this subdirectory of ‘sysdeps’ implements. For example,

‘sysdeps/unix/inet/Subdirs’ contains ‘inet’; the ‘inet’ di-

rectory contains various network-oriented operations which only make sense to put in the library on systems that support the Internet.

‘Dist’

This file contains the names of files (relative to the subdirectory of ‘sysdeps’ in which it appears) which should be included in the distribution. List any new files used by rules in the ‘Makefile’ in the same directory, or header files used by the source files in that directory. You don’t need to list files that are implementations (either C or assembly source) of routines whose names are given in the machine-independent makefiles in the main source tree.

‘configure’

This file is a shell script fragment to be run at configuration time. The top-level ‘configure’ script uses the shell . command to read the ‘configure’ file in each system-dependent directory chosen, in order. The ‘configure’ files are often generated from ‘configure.in’ files using Autoconf. A system-dependent ‘configure’ script will usually add things to the shell variables ‘DEFS’ and ‘config_vars’; see the toplevel ‘configure’ script for details. The script can check for ‘--with-package’ options that were passed to the toplevel ‘configure’. For an option ‘--with-package=value’ ‘configure’ sets the shell variable ‘with_package’ (with any dashes in package converted to underscores) to value; if the option is just ‘--with-package’ (no argument), then it sets ‘with_package’ to ‘yes’.

‘configure.in’

This file is an Autoconf input fragment to be processed into the file ‘configure’ in this subdirectory. See section “Introduction” in Autoconf: Generating Automatic Configuration Scripts, for a description of Autoconf. You should write either ‘configure’ or ‘configure.in’, but not both. The first line of ‘configure.in’ should invoke the m4 macro ‘GLIBC_PROVIDES’. This macro does several AC_PROVIDE calls for Autoconf macros which are used by the top-level ‘configure’ script; without this, those macros might be invoked again unnecessarily by Autoconf.

That is the general system for how system-dependencies are isolated. The next section explains how to decide what directories in ‘sysdeps’ to use. Section C.4.2 [Porting the GNU C Library to Unix Systems], page 680, has some tips on porting the library to Unix variants.

C.4.1 Layout of the ‘sysdeps’ Directory Hierarchy

A GNU configuration name has three parts: the CPU type, the manufacturer’s name, and the operating system. ‘configure’ uses these to pick the list of system-dependent directories to look for. If the ‘--nfp’ option is not passed to ‘configure’, the directory ‘machine/fpu’ is also used. The operating system often has a base operating system; for example, if the operating system is ‘sunos4.1’, the base operating system is ‘unix/bsd’. The algorithm used to pick the list of directories is simple: ‘configure’ makes a list of the base operating system, manufacturer, CPU type, and operating system, in that order. It then concatenates all these together with slashes in between, to produce a directory name; for example, the configuration ‘sparc-sun-sunos4.1’ results in

‘unix/bsd/sun/sparc/sunos4.1’. ‘configure’ then tries removing each el-

ement of the list in turn, so ‘unix/bsd/sparc’ and ‘sun/sparc’ are also tried, among others. Since the precise version number of the operating system is often not important, and it would be very inconvenient, for example, to have identical ‘sunos4.1.1’ and ‘sunos4.1.2’ directories, ‘configure’ tries successively less specific operating system names by removing trailing suffixes starting with a period.

As an example, here is the complete list of directories that would be tried for the configuration ‘sparc-sun-sunos4.1’ (without the ‘--nfp’ option):

     sparc/fpu
     unix/bsd/sun/sunos4.1/sparc
     unix/bsd/sun/sunos4.1
     unix/bsd/sun/sunos4/sparc
     unix/bsd/sun/sunos4
     unix/bsd/sun/sunos/sparc
     unix/bsd/sun/sunos

unix/bsd/sun/sparc
unix/bsd/sun
unix/bsd/sunos4.1/sparc
unix/bsd/sunos4.1
unix/bsd/sunos4/sparc
unix/bsd/sunos4
unix/bsd/sunos/sparc
unix/bsd/sunos
unix/bsd/sparc
unix/bsd
unix/sun/sunos4.1/sparc
unix/sun/sunos4.1
unix/sun/sunos4/sparc
unix/sun/sunos4
unix/sun/sunos/sparc
unix/sun/sunos
unix/sun/sparc
unix/sun
unix/sunos4.1/sparc
unix/sunos4.1
unix/sunos4/sparc
unix/sunos4
unix/sunos/sparc
unix/sunos
unix/sparc
unix
sun/sunos4.1/sparc
sun/sunos4.1
sun/sunos4/sparc
sun/sunos4
sun/sunos/sparc
sun/sunos
sun/sparc
sun
sunos4.1/sparc
sunos4.1
sunos4/sparc
sunos4
sunos/sparc
sunos
sparc

Different machine architectures are conventionally subdirectories at the top level of the ‘sysdeps’ directory tree. For example, ‘sysdeps/sparc’ and ‘sysdeps/m68k’. These contain files specific to those machine architectures, but not specific to any particular operating system. There might be subdirectories for specializations of those architectures, such as ‘sysdeps/m68k/68020’. Code which is specific to the floating-point coprocessor used with a particular machine should go in ‘sysdeps/machine/fpu’.

There are a few directories at the top level of the ‘sysdeps’ hierarchy that are not for particular machine architectures.

‘generic’

‘stub’ As described above (see Section C.4 [Porting the GNU C Library], page 674), these are the two subdirectories that every configuration implicitly uses after all others.

‘ieee754’ This directory is for code using the IEEE 754 floating-point format, where the C type float is IEEE 754 single-precision format, and double is IEEE 754 double-precision format. Usually this directory is referred to in the ‘Implies’ file in a machine architecture-specific directory, such as ‘m68k/Implies’.

‘posix’ This directory contains implementations of things in the library in terms of POSIX.1 functions. This includes some of the POSIX.1 functions themselves. Of course, POSIX.1 cannot be completely implemented in terms of itself, so a configuration using just ‘posix’ cannot be complete.

‘unix’ This is the directory for Unix-like things. See Section C.4.2 [Porting the GNU C Library to Unix Systems], page 680. ‘unix’ implies ‘posix’. There are some special-purpose subdirectories of ‘unix’:

‘unix/common’

This directory is for things common to both BSD and System V release 4. Both ‘unix/bsd’ and

          ‘unix/sysv/sysv4’ imply ‘unix/common’.
‘unix/inet’

This directory is for socket and related functions on Unix systems. The ‘inet’ top-level subdirectory

          is enabled by ‘unix/inet/Subdirs’. ‘unix/common’

implies ‘unix/inet’.

‘mach’ This is the directory for things based on the Mach microkernel from CMU (including the GNU operating system). Other basic operating systems (VMS, for example) would have their own directories at the top level of the ‘sysdeps’ hierarchy, parallel to ‘unix’ and ‘mach’.

C.4.2 Porting the GNU C Library to Unix Systems

Most Unix systems are fundamentally very similar. There are variations between different machines, and variations in what facilities are provided by the kernel. But the interface to the operating system facilities is, for the most part, pretty uniform and simple.

The code for Unix systems is in the directory ‘unix’, at the top level of the ‘sysdeps’ hierarchy. This directory contains subdirectories (and subdirectory trees) for various Unix variants.

The functions which are system calls in most Unix systems are implemented in assembly code in files in ‘sysdeps/unix’. These files are named with a suffix of ‘.S’; for example, ‘__open.S’. Files ending in ‘.S’ are run through the C preprocessor before being fed to the assembler.

These files all use a set of macros that should be defined in ‘sysdep.h’. The ‘sysdep.h’ file in ‘sysdeps/unix’ partially defines them; a ‘sysdep.h’ file in another directory must finish defining them for the particular machine and operating system variant. See ‘sysdeps/unix/sysdep.h’ and the machine-specific ‘sysdep.h’ implementations to see what these macros are and what they should do.

The system-specific makefile for the ‘unix’ directory (that is, the file ‘sysdeps/unix/Makefile’) gives rules to generate several files from the Unix system you are building the library on (which is assumed to be the target system you are building the library for). All the generated files are put in the directory where the object files are kept; they should not affect the source tree itself. The files generated are ‘ioctls.h’, ‘errnos.h’, ‘sys/param.h’, and ‘errlist.c’ (for the ‘stdio’ section of the library).

C.5 Contributors to the GNU C Library

The GNU C library was written originally by Roland McGrath. Some parts of the library were contributed or worked on by other people.

• The getopt function and related code were written by Richard Stallman, David J. MacKenzie, and Roland McGrath.

• The merge sort function qsort was written by Michael J. Haertel.

• The quick sort function used as a fallback by qsort was written by Douglas C. Schmidt.

• The memory allocation functions malloc, realloc and free and related code were written by Michael J. Haertel.

• Fast implementations of many of the string functions (memcpy, strlen, etc.) were written by Torbjörn Granlund.

• The ‘tar.h’ header file was written by David J. MacKenzie.

• The port to the MIPS DECStation running Ultrix 4 (mips-dec-ultrix4) was contributed by Brendan Kehoe and Ian Lance Taylor.

• The DES encryption function crypt and related functions were contributed by Michael Glad.

• The ftw function was contributed by Ian Lance Taylor.

• The startup code to support SunOS shared libraries was contributed by Tom Quinn.

• The mktime function was contributed by Paul Eggert.

• The port to the Sequent Symmetry running Dynix version 3 (i386sequent-bsd) was contributed by Jason Merrill.

• The timezone support code is derived from the public-domain timezone package by Arthur David Olson and his many contributors.

• The port to the DEC Alpha running OSF/1 (alpha-dec-osf1) was contributed by Brendan Kehoe, using some code written by Roland McGrath.

• The port to SGI machines running Irix 4 (mips-sgi-irix4) was contributed by Tom Quinn.

• The port of the Mach and Hurd code to the MIPS architecture (mipsanything-gnu) was contributed by Kazumoto Kojima.

• The floating-point printing function used by printf and friends and the floating-point reading function used by scanf, strtod and friends were written by Ulrich Drepper. The multi-precision integer functions used in those functions are taken from GNU MP, which was contributed by Torbjörn Granlund.

• The internationalization support in the library, and the support programs locale and localedef, were written by Ulrich Drepper. Ulrich Drepper adapted the support code for message catalogs (‘libintl.h’, etc.) from the GNU gettext package, which he also wrote. He also contributed the catgets support and the entire suite of multi-byte and wide-character support functions (‘wctype.h’, ‘wchar.h’, etc.).

• The implementations of the ‘nsswitch.conf’ mechanism and the files and DNS backends for it were designed and written by Ulrich Drepper and Roland McGrath, based on a backend interface defined by Peter Eriksson.

• The port to Linux i386/ELF (i386-anything-linux) was contributed by Ulrich Drepper, based in large part on work done in Hongjiu Lu’s Linux version of the GNU C Library.

• The port to Linux/m68k (m68k-anything-linux) was contributed by Andreas Schwab.

• Richard Henderson contributed the ELF dynamic linking code and other support for the Alpha processor.

• David Mosberger-Tang contributed the port to Linux/Alpha (alphaanything-linux).

• Stephen R. van den Berg contributed a highly-optimized strstr function.

• Ulrich Drepper contributed the hsearch and drand48 families of functions; reentrant ‘…_r’ versions of the random family; System V shared memory and IPC support code; and several highly-optimized string functions for ix86 processors.

• The math functions are taken from fdlibm-5.1 by Sun Microsystems, as modified by J.T. Conklin, Ian Lance Taylor, Ulrich Drepper, Andreas Schwab, and Roland McGrath.

• The libio library used to implement stdio functions on some platforms was written by Per Bothner and modified by Ulrich Drepper.

• The Internet-related code (most of the ‘inet’ subdirectory) and several other miscellaneous functions and header files have been included from 4.4 BSD with little or no modification. All code incorporated from 4.4 BSD is under the following copyright:

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1991 Regents of the University of California. All rights reserved.

Redistribution and use in source and binary forms, with or without modification, are permitted provided that the following conditions are met:

1. Redistributions of source code must retain the above copyright notice, this list of conditions and the following disclaimer.

2. Redistributions in binary form must reproduce the above copyright notice, this list of conditions and the following disclaimer in the documentation and/or other materials provided with the distribution.

3. All advertising materials mentioning features or use of this

software must display the following acknowledgement: This product includes software developed by the University of California, Berkeley and its contributors.

4. Neither the name of the University nor the names of its contributors may be used to endorse or promote products derived from this software without specific prior written permission.

this software is provided by the regents and contributors “as is” and any express or implied warranties, including, but not limited to, the implied warranties of merchantability and fitness for a particular purpose are disclaimed. in no event shall the regents or contributors be liable for any direct, indirect, incidental, special, exemplary, or consequential damages (including, but not limited to, procurement of substitute goods or services; loss of use, data, or profits; or business interruption) however caused and on any theory of liability, whether in contract, strict liability, or tort (including negligence or otherwise) arising in any way out of the use of this software, even if advised of the possibility of such damage.

• The random number generation functions random, srandom, setstate and initstate, which are also the basis for the rand and srand functions, were written by Earl T. Cohen for the University of California at Berkeley and are copyrighted by the Regents of the University of California. They have undergone minor changes to fit into the GNU C library and to fit the ISO C standard, but the functional code is Berkeley’s.

• The Internet resolver code is taken directly from BIND 4.9.5, which is

under both the Berkeley copyright above and also:

Portions Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993 by Digital Equipment Corporation. Permission to use, copy, modify, and distribute this software for any purpose with or without fee is hereby granted, provided that the above copyright notice and this permission notice appear in all copies, and that the name of Digital Equipment Corporation not be used in advertising or publicity pertaining to distribution of the document or software without specific, written prior permission. the software is provided “as is” and digital equipment corp. disclaims all warranties with regard to this software, including all implied warranties of merchantability and fitness. in no event shall digital equipment corporation be liable for any special, direct, indirect, or consequential damages or any damages whatsoever resulting from loss of use, data or profits, whether in an action of contract, negligence or other tortious action, arising out of or in connection with the use or performance of this software.

• The code to support Sun RPC is taken verbatim from Sun’s rpcsrc-4.0

distribution, and is covered by this copyright:

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1984, Sun Microsystems, Inc. Sun RPC is a product of Sun Microsystems, Inc. and is provided for unrestricted use provided that this legend is included on all tape media and as a part of the software program in whole or part. Users may copy or modify Sun RPC without charge, but are not authorized to license or distribute it to anyone else except as part of a product or program developed by the user. sun rpc is provided as is with no warranties of any kind including the warranties of design, merchantibility and fitness for a particular purpose, or arising from a course of dealing, usage or trade practice. Sun RPC is provided with no support and without any obligation on the part of Sun Microsystems, Inc. to assist in its use, correction, modification or enhancement. sun microsystems, inc. shall have no liability with respect to the infringement of copyrights, trade secrets or any patents by sun rpc or any part thereof. In no event will Sun Microsystems, Inc. be liable for any lost revenue or profits or other special, indirect and consequential damages, even if Sun has been advised of the possibility of such damages.

Sun Microsystems, Inc. 2550 Garcia Avenue Mountain View, California 94043

• Some of the support code for Mach is taken from Mach 3.0 by CMU,

and is under the following copyright terms: Mach Operating System Copyright \begin{matrix}c \\ ◯ \end{matrix} 1991,1990,1989 Carnegie Mellon University All Rights Reserved.

Permission to use, copy, modify and distribute this software and its documentation is hereby granted, provided that both the copyright notice and this permission notice appear in all copies of the software, derivative works or modified versions, and any portions thereof, and that both notices appear in supporting documentation. carnegie mellon allows free use of this software in its “as is” condition. carnegie mellon disclaims any liability of any kind for any damages whatsoever resulting from the use of this software. Carnegie Mellon requests users of this software to return to

Software Distribution Coordinator School of Computer Science Carnegie Mellon University Pittsburgh PA 15213-3890

or Software.Distribution@CS.CMU.EDU any improvements

or extensions that they make and grant Carnegie Mellon the rights to redistribute these changes.

686 The GNU C Library

Appendix D GNU LIBRARY GENERAL PUBLIC LICENSE

Version 2, June 1991

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1991 Free Software Foundation, Inc.

675 Mass Ave, Cambridge, MA 02139, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

[This is the first released version of the library GPL. It is numbered 2 because it goes with version 2 of the ordinary GPL.]

Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public Licenses are intended to guarantee your freedom to share and change free software—to make sure the software is free for all its users.

This license, the Library General Public License, applies to some specially designated Free Software Foundation software, and to any other libraries whose authors decide to use it. You can use it for your libraries, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs; and that you know you can do these things.

To protect your rights, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. These restrictions translate to certain responsibilities for you if you distribute copies of the library, or if you modify it.

For example, if you distribute copies of the library, whether gratis or for a fee, you must give the recipients all the rights that we gave you. You must make sure that they, too, receive or can get the source code. If you link a program with the library, you must provide complete object files to the recipients so that they can relink them with the library, after making changes to the library and recompiling it. And you must show them these terms so they know their rights.

Our method of protecting your rights has two steps: (1) copyright the library, and (2) offer you this license which gives you legal permission to copy, distribute and/or modify the library.

Also, for each distributor’s protection, we want to make certain that everyone understands that there is no warranty for this free library. If the library is modified by someone else and passed on, we want its recipients to know that what they have is not the original version, so that any problems introduced by others will not reflect on the original authors’ reputations.

Finally, any free program is threatened constantly by software patents. We wish to avoid the danger that companies distributing free software will individually obtain patent licenses, thus in effect transforming the program into proprietary software. To prevent this, we have made it clear that any patent must be licensed for everyone’s free use or not licensed at all.

Most GNU software, including some libraries, is covered by the ordinary GNU General Public License, which was designed for utility programs. This license, the GNU Library General Public License, applies to certain designated libraries. This license is quite different from the ordinary one; be sure to read it in full, and don’t assume that anything in it is the same as in the ordinary license.

The reason we have a separate public license for some libraries is that they blur the distinction we usually make between modifying or adding to a program and simply using it. Linking a program with a library, without changing the library, is in some sense simply using the library, and is analogous to running a utility program or application program. However, in a textual and legal sense, the linked executable is a combined work, a derivative of the original library, and the ordinary General Public License treats it as such.

Because of this blurred distinction, using the ordinary General Public License for libraries did not effectively promote software sharing, because most developers did not use the libraries. We concluded that weaker conditions might promote sharing better.

However, unrestricted linking of non-free programs would deprive the users of those programs of all benefit from the free status of the libraries themselves. This Library General Public License is intended to permit developers of non-free programs to use free libraries, while preserving your freedom as a user of such programs to change the free libraries that are incorporated in them. (We have not seen how to achieve this as regards changes in header files, but we have achieved it as regards changes in the actual functions of the Library.) The hope is that this will lead to faster development of free libraries.

The precise terms and conditions for copying, distribution and modification follow. Pay close attention to the difference between a “work based on the library” and a “work that uses the library”. The former contains code derived from the library, while the latter only works together with the library.

Note that it is possible for a library to be covered by the ordinary General Public License rather than by this special one.

TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND MODIFICATION

0. This License Agreement applies to any software library which contains a notice placed by the copyright holder or other authorized party saying it may be distributed under the terms of this Library General Public License (also called “this License”). Each licensee is addressed as “you”. A “library” means a collection of software functions and/or data prepared so as to be conveniently linked with application programs (which use some of those functions and data) to form executables. The “Library”, below, refers to any such software library or work which has been distributed under these terms. A “work based on the Library” means either the Library or any derivative work under copyright law: that is to say, a work containing the Library or a portion of it, either verbatim or with modifications and/or translated straightforwardly into another language. (Hereinafter, translation is included without limitation in the term “modification”.) “Source code” for a work means the preferred form of the work for making modifications to it. For a library, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the library. Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running a program using the Library is not restricted, and output from such a program is covered only if its contents constitute a work based on the Library (independent of the use of the Library in a tool for writing it). Whether that is true depends on what the Library does and what the program that uses the Library does.

1. You may copy and distribute verbatim copies of the Library’s complete source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and distribute a copy of this License along with the Library. You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

2. You may modify your copy or copies of the Library or any portion of it, thus forming a work based on the Library, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions: a. The modified work must itself be a software library. b. You must cause the files modified to carry prominent notices stating

that you changed the files and the date of any change.

c. You must cause the whole of the work to be licensed at no charge to all third parties under the terms of this License.

d. If a facility in the modified Library refers to a function or a table of data to be supplied by an application program that uses the facility, other than as an argument passed when the facility is invoked, then you must make a good faith effort to ensure that, in the event an application does not supply such function or table, the facility still operates, and performs whatever part of its purpose remains meaningful.

(For example, a function in a library to compute square roots has a purpose that is entirely well-defined independent of the application. Therefore, Subsection 2d requires that any applicationsupplied function or table used by this function must be optional: if the application does not supply it, the square root function must still compute square roots.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Library, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Library, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it.

Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Library.

In addition, mere aggregation of another work not based on the Library with the Library (or with a work based on the Library) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

3. You may opt to apply the terms of the ordinary GNU General Public License instead of this License to a given copy of the Library. To do this, you must alter all the notices that refer to this License, so that they refer to the ordinary GNU General Public License, version 2, instead of to this License. (If a newer version than version 2 of the ordinary GNU General Public License has appeared, then you can specify that version instead if you wish.) Do not make any other change in these notices.

Once this change is made in a given copy, it is irreversible for that copy, so the ordinary GNU General Public License applies to all subsequent copies and derivative works made from that copy. This option is useful when you wish to copy part of the code of the Library into a program that is not a library.

4. You may copy and distribute the Library (or a portion or derivative of it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you accompany it with the complete corresponding machine-readable source code, which must be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange. If distribution of object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place satisfies the requirement to distribute the source code, even though third parties are not compelled to copy the source along with the object code.

5. A program that contains no derivative of any portion of the Library, but is designed to work with the Library by being compiled or linked with it, is called a “work that uses the Library”. Such a work, in isolation, is not a derivative work of the Library, and therefore falls outside the scope of this License. However, linking a “work that uses the Library” with the Library creates an executable that is a derivative of the Library (because it contains portions of the Library), rather than a “work that uses the library”. The executable is therefore covered by this License. Section 6 states terms for distribution of such executables. When a “work that uses the Library” uses material from a header file that is part of the Library, the object code for the work may be a derivative work of the Library even though the source code is not. Whether this is true is especially significant if the work can be linked without the Library, or if the work is itself a library. The threshold for this to be true is not precisely defined by law. If such an object file uses only numerical parameters, data structure layouts and accessors, and small macros and small inline functions (ten lines or less in length), then the use of the object file is unrestricted, regardless of whether it is legally a derivative work. (Executables containing this object code plus portions of the Library will still fall under Section 6.) Otherwise, if the work is a derivative of the Library, you may distribute the object code for the work under the terms of Section 6. Any executables containing that work also fall under Section 6, whether or not they are linked directly with the Library itself.

6. As an exception to the Sections above, you may also compile or link a “work that uses the Library” with the Library to produce a work containing portions of the Library, and distribute that work under terms of your choice, provided that the terms permit modification of the work for the customer’s own use and reverse engineering for debugging such modifications. You must give prominent notice with each copy of the work that the Library is used in it and that the Library and its use are covered by this License. You must supply a copy of this License. If the work during execution displays copyright notices, you must include the copyright notice for the Library among them, as well as a reference directing the user to the copy of this License. Also, you must do one of these things: a. Accompany the work with the complete corresponding machine-

readable source code for the Library including whatever changes were used in the work (which must be distributed under Sections 1 and 2 above); and, if the work is an executable linked with the Library, with the complete machine-readable “work that uses the Library”, as object code and/or source code, so that the user can modify the Library and then relink to produce a modified executable containing the modified Library. (It is understood that the user who changes the contents of definitions files in the Library will not necessarily be able to recompile the application to use the modified definitions.)

b. Accompany the work with a written offer, valid for at least three years, to give the same user the materials specified in Subsection 6a, above, for a charge no more than the cost of performing this distribution.

c. If distribution of the work is made by offering access to copy from a designated place, offer equivalent access to copy the above specified materials from the same place.

d. Verify that the user has already received a copy of these materials or that you have already sent this user a copy.

For an executable, the required form of the “work that uses the Library” must include any data and utility programs needed for reproducing the executable from it. However, as a special exception, the source code distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable. It may happen that this requirement contradicts the license restrictions of other proprietary libraries that do not normally accompany the operating system. Such a contradiction means you cannot use both them and the Library together in an executable that you distribute.

7. You may place library facilities that are a work based on the Library side-by-side in a single library together with other library facilities not covered by this License, and distribute such a combined library, provided that the separate distribution of the work based on the Library and of the other library facilities is otherwise permitted, and provided that you do these two things:

a. Accompany the combined library with a copy of the same work based on the Library, uncombined with any other library facilities. This must be distributed under the terms of the Sections above.

b. Give prominent notice with the combined library of the fact that part of it is a work based on the Library, and explaining where to find the accompanying uncombined form of the same work.

8. You may not copy, modify, sublicense, link with, or distribute the Library except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, link with, or distribute the Library is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

9. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Library or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Library (or any work based on the Library), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Library or works based on it.

10. Each time you redistribute the Library (or any work based on the Library), the recipient automatically receives a license from the original licensor to copy, distribute, link with or modify the Library subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties to this License.

11. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Library at all. For example, if a patent license would not permit royaltyfree redistribution of the Library by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Library.

If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply, and the section as a whole is intended to apply in other circumstances.

It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice.

This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

12. If the distribution and/or use of the Library is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Library under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

13. The Free Software Foundation may publish revised and/or new versions of the Library General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

Each version is given a distinguishing version number. If the Library specifies a version number of this License which applies to it and “any later version”, you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Library does not specify a license version number, you may choose any version ever published by the Free Software Foundation.

14. If you wish to incorporate parts of the Library into other free programs whose distribution conditions are incompatible with these, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

15. BECAUSE THE LIBRARY IS LICENSED FREE OF CHARGE, THERE IS NO WARRANTY FOR THE LIBRARY, TO THE EX-TENT PERMITTED BY APPLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLD-ERS AND/OR OTHER PARTIES PROVIDE THE LIBRARY “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUAL-ITY AND PERFORMANCE OF THE LIBRARY IS WITH YOU. SHOULD THE LIBRARY PROVE DEFECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORREC-TION.

16. IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY MODIFY AND/OR REDIS-TRIBUTE THE LIBRARY AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE LIBRARY (INCLUD-ING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE LIBRARY TO OPER-ATE WITH ANY OTHER SOFTWARE), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAMAGES.

END OF TERMS AND CONDITIONS How to Apply These Terms to Your New Libraries

If you develop a new library, and you want it to be of the greatest possible use to the public, we recommend making it free software that everyone can redistribute and change. You can do so by permitting redistribution under these terms (or, alternatively, under the terms of the ordinary General Public License).

To apply these terms, attach the following notices to the library. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

one line to give the library’s name and an idea of what it does. Copyright (C) year name of author

This library is free software; you can redistribute it and/or modify it
under the terms of the GNU Library General Public License as published
by the Free Software Foundation; either version 2 of the License, or (at
your option) any later version.
This library is distributed in the hope that it will be useful, but
WITHOUT ANY WARRANTY; without even the implied warranty of
MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the GNU
Library General Public License for more details.
You should have received a copy of the GNU Library General Public
License along with this library; if not, write to the Free Software
Foundation, Inc., 675 Mass Ave, Cambridge, MA 02139, USA.

Also add information on how to contact you by electronic and paper mail. You should also get your employer (if you work as a programmer) or your

school, if any, to sign a “copyright disclaimer” for the library, if necessary. Here is a sample; alter the names:

Yoyodyne, Inc., hereby disclaims all copyright interest in the library
`Frob' (a library for tweaking knobs) written by James Random Hacker.

signature of Ty Coon, 1 April 1990

Ty Coon, President of Vice

That’s all there is to it!

Concept Index 697

Concept Index

/

‘/etc/nsswitch.conf’. . . . . . . . . . . . . . . . 502

POSIX OPTION ORDER environment variable.. . . . . . . . . . . . . . . . . . . . . . . . . 462

4

4.n BSD Unix . . . . . . . . . . . . . . . . . . . . . . . . . . 3

A

abort signal . . . . . . . . . . . . . . . . . . . . . . . . . . 398 aborting a program . . . . . . . . . . . . . . . . . . 465 absolute file name . . . . . . . . . . . . . . . . . . . . . 88 absolute value functions . . . . . . . . . . . . . . 306 accepting connections . . . . . . . . . . . . . . . . 248 access permission for a file . . . . . . . . . . . 204 access, testing for . . . . . . . . . . . . . . . . . . . . 206 accessing directories . . . . . . . . . . . . . . . . . . 183 address of socket . . . . . . . . . . . . . . . . . . . . . 223 alarm signal . . . . . . . . . . . . . . . . . . . . . . . . . 401 alarms, setting . . . . . . . . . . . . . . . . . . . . . . . 356 alignment (in obstacks). . . . . . . . . . . . . . . . 52 alignment (with malloc) . . . . . . . . . . . . . . 37 alloca disadvantages . . . . . . . . . . . . . . . . . 57 alloca function . . . . . . . . . . . . . . . . . . . . . . . 55 allocation (obstacks) . . . . . . . . . . . . . . . . . . 45 allocation hooks, for malloc . . . . . . . . . . . 40 allocation of memory with malloc . . . . . 32 allocation size of string . . . . . . . . . . . . . . . . 66 allocation statistics. . . . . . . . . . . . . . . . . . . . 41 alphabetic character. . . . . . . . . . . . . . . . . . . 61 alphanumeric character. . . . . . . . . . . . . . . . 62 append-access files . . . . . . . . . . . . . . . . . . . . 87 argc (program argument count) . . . . . . 447 argument promotion . . . . . . . . . . . . . . . . . 558 arguments (variadic functions) . . . . . . . 556 arguments, how many . . . . . . . . . . . . . . . . 557 arguments, to program . . . . . . . . . . . . . . . 447 argv (program argument vector). . . . . . 447 arithmetic expansion . . . . . . . . . . . . . . . . . 334 array comparison functions . . . . . . . . . . . . 72 array copy functions. . . . . . . . . . . . . . . . . . . 67 array search function . . . . . . . . . . . . . . . . . 317 array sort function . . . . . . . . . . . . . . . . . . . 318 ASCII character. . . . . . . . . . . . . . . . . . . . . . . 63 assertions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 attributes of a file . . . . . . . . . . . . . . . . . . . . 195 automatic allocation . . . . . . . . . . . . . . . . . . 31 automatic freeing . . . . . . . . . . . . . . . . . . . . . 55 automatic storage with variable size … 55

B

background job . . . . . . . . . . . . . . . . . . . . . . 480 background job, launching . . . . . . . . . . . 491 base (of floating point number) . . . . . . . 565 basic byte sequence . . . . . . . . . . . . . . . . . . 368 baud rate . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 Berkeley Unix . . . . . . . . . . . . . . . . . . . . . . . . . . 3 bias (of floating point number exponent)

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 565 big-endian . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 binary I/O to a stream . . . . . . . . . . . . . . . 104 binary search function (for arrays)…. 317 binary stream . . . . . . . . . . . . . . . . . . . . . . . . 136 binding a socket address . . . . . . . . . . . . . 223 blank character . . . . . . . . . . . . . . . . . . . . . . . 62 block I/O to a stream . . . . . . . . . . . . . . . . 104 blocked signals . . . . . . . . . . . . . . . . . . . . . . . 394 blocked signals, checking for . . . . . . . . . . 437 blocking signals . . . . . . . . . . . . . . . . . . . . . . 431 blocking signals, in a handler . . . . . . . . . 436 bootstrapping, and services. . . . . . . . . . . 504 break condition, detecting . . . . . . . . . . . . 278 break condition, generating. . . . . . . . . . . 293 breaking a string into tokens . . . . . . . . . . 80 broken pipe signal. . . . . . . . . . . . . . . . . . . . 404 broken-down time . . . . . . . . . . . . . . . 341, 345 BSD compatibility library . . . . . . . . . . . . 498 BSD compatibility library.. . . . . . . . . . . . . . 9 BSD Unix. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 buffering of streams . . . . . . . . . . . . . . . . . . 140 buffering, controlling . . . . . . . . . . . . . . . . . 141 bugs, reporting. . . . . . . . . . . . . . . . . . . . . . . 671 bus error. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 398 byte order conversion, for socket. . . . . . 239 byte stream . . . . . . . . . . . . . . . . . . . . . . . . . . 221

C

calendar time . . . . . . . . . . . . . . . . . . . . . . . . 341 calendar time and broken-down time

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 calling variadic functions . . . . . . . . . . . . . 558 canonical input processing. . . . . . . . . . . . 272 capacity limits, POSIX . . . . . . . . . . . . . . . 537 carrier detect . . . . . . . . . . . . . . . . . . . . . . . . 280 case conversion of characters . . . . . . . . . . 63 catching signals . . . . . . . . . . . . . . . . . . . . . . 395 categories for locales . . . . . . . . . . . . . . . . . 380 change working directory . . . . . . . . . . . . . 181 changing the locale. . . . . . . . . . . . . . . . . . . 381 changing the size of a block (malloc) . . 35 changing the size of a block (obstacks)

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 channels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 character case conversion . . . . . . . . . . . . . . 63 character code . . . . . . . . . . . . . . . . . . . . . . . 371 character predicates . . . . . . . . . . . . . . . . . . . 61 character testing . . . . . . . . . . . . . . . . . . . . . . 61 checking for pending signals . . . . . . . . . . 437 child process . . . . . . . . . . . . . . . . . . . . 467, 468 child process signal. . . . . . . . . . . . . . . . . . . 402 chunks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 classification of characters . . . . . . . . . . . . . 61 cleaning up a stream . . . . . . . . . . . . . . . . . 160 clearing terminal input queue . . . . . . . . 294 client . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 clock ticks . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 close-on-exec (file descriptor flag). . . . . 169 closing a file descriptor . . . . . . . . . . . . . . . 151 closing a socket . . . . . . . . . . . . . . . . . . . . . . 244 closing a stream . . . . . . . . . . . . . . . . . . . . . . . 96 code, character. . . . . . . . . . . . . . . . . . . . . . . 371 collating strings . . . . . . . . . . . . . . . . . . . . . . . 75 combining locales . . . . . . . . . . . . . . . . . . . . 380 command argument syntax . . . . . . . . . . . 448 command arguments, parsing. . . . . . . . . 449 command line arguments . . . . . . . . . . . . . 447 command substitution . . . . . . . . . . . . . . . 334 communication style (of a socket) …. 221 comparing strings and arrays . . . . . . . . . . 72 Comparison Function . . . . . . . . . . . . . . . . 317 concatenating strings . . . . . . . . . . . . . . . . . . 67 configurations, all supported . . . . . . . . . 670 connecting a socket . . . . . . . . . . . . . . . . . . 246 connection . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 consistency checking . . . . . . . . . . . . . . . . . 553 consistency checking, of heap . . . . . . . . . . 38 continue signal . . . . . . . . . . . . . . . . . . . . . . . 402 control character . . . . . . . . . . . . . . . . . . . . . . 62 control operations on files . . . . . . . . . . . . 165 controlling process . . . . . . . . . . . . . . . . . . . 481 controlling terminal . . . . . . . . . . . . . . . . . . 479 controlling terminal, access to . . . . . . . . 481 controlling terminal, determining . . . . . 497 controlling terminal, setting . . . . . . . . . . 172 conversion specifications (printf) …. 105 conversion specifications (scanf) . . . . . 127 converting byte order . . . . . . . . . . . . . . . . 239 converting case of characters . . . . . . . . . . 63 converting extended characters . . . . . . . 374 converting extended strings . . . . . . . . . . 372 converting file descriptor to stream … 159 converting floats to integers . . . . . . . . . . 308 converting group ID to group name… 525 converting group name to group ID… 525 converting host address to name. . . . . . 235 converting host name to address. . . . . . 235 converting network name to network

number. . . . . . . . . . . . . . . . . . . . . . . . . . 269 converting network number to network

name . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 converting port number to service name

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 converting service name to port number

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 converting string to collation order …. 76 converting strings to numbers . . . . . . . . 310 converting user ID to user name . . . . . . 522 converting user name to user ID . . . . . . 522 cookie, for custom stream . . . . . . . . . . . . 147 copying strings and arrays . . . . . . . . . . . . . 67 CPU time. . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 create on open (file status flag) . . . . . . . 171 creating a directory . . . . . . . . . . . . . . . . . . 194 creating a FIFO special file. . . . . . . . . . . 219 creating a pipe . . . . . . . . . . . . . . . . . . . . . . . 215 creating a pipe to a subprocess . . . . . . . 217 creating a process . . . . . . . . . . . . . . . . . . . . 468 creating a socket . . . . . . . . . . . . . . . . . . . . . 244 creating a socket pair . . . . . . . . . . . . . . . . 245 creating special files . . . . . . . . . . . . . . . . . . 210 cube root function . . . . . . . . . . . . . . . . . . . 300 currency symbols . . . . . . . . . . . . . . . . . . . . 385 current working directory . . . . . . . . . . . . 181 custom streams . . . . . . . . . . . . . . . . . . . . . . 147 customizing printf . . . . . . . . . . . . . . . . . . 121

D

data loss on sockets . . . . . . . . . . . . . . . . . . 221 databases . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 datagram socket . . . . . . . . . . . . . . . . . . . . . 260 datagrams, transmitting. . . . . . . . . . . . . . 260 date and time . . . . . . . . . . . . . . . . . . . . . . . . 341 Daylight Saving Time . . . . . . . . . . . . . . . . 346 decimal digit character . . . . . . . . . . . . . . . . 61 decimal-point separator . . . . . . . . . . . . . . 384 declaration (compared to definition) …. 4 declaring variadic functions . . . . . . . . . . 558 default action (for a signal) . . . . . . . . . . 395 default action for a signal . . . . . . . . . . . . 407 default argument promotions . . . . . . . . . 558 default value, and NSS . . . . . . . . . . . . . . . 504 defining new printf conversions. . . . . . 121 definition (compared to declaration) …. 4 delayed suspend character . . . . . . . . . . . . 289 deleting a directory . . . . . . . . . . . . . . . . . . 193 deleting a file . . . . . . . . . . . . . . . . . . . . . . . . 192 delivery of signals . . . . . . . . . . . . . . . . . . . . 394 descriptors and streams . . . . . . . . . . . . . . 160 digit character . . . . . . . . . . . . . . . . . . . . . . . . 61 directories, accessing . . . . . . . . . . . . . . . . . 183 directories, creating . . . . . . . . . . . . . . . . . . 194 directories, deleting . . . . . . . . . . . . . . . . . . 193 directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 directory entry . . . . . . . . . . . . . . . . . . . . . . . . 88 directory stream . . . . . . . . . . . . . . . . . . . . . 183 disadvantages of alloca . . . . . . . . . . . . . . . 57 DISCARD character . . . . . . . . . . . . . . . . . 291 DNS server unavailable. . . . . . . . . . . . . . . 503 domain (of socket) . . . . . . . . . . . . . . . . . . . 221 domain error . . . . . . . . . . . . . . . . . . . . . . . . . 297 dot notation, for Internet addresses… 232 DSUSP character . . . . . . . . . . . . . . . . . . . . 289 duplicating file descriptors . . . . . . . . . . . 166 dynamic allocation . . . . . . . . . . . . . . . . . . . . 31

E

echo of terminal input. . . . . . . . . . . . . . . . 282 effective group ID . . . . . . . . . . . . . . . . . . . . 511 effective user ID. . . . . . . . . . . . . . . . . . . . . . 511 efficiency and malloc. . . . . . . . . . . . . . . . . . 37 efficiency and obstacks . . . . . . . . . . . . . . . . 49 efficiency of chunks. . . . . . . . . . . . . . . . . . . . 52 EINTR, and restarting interrupted

primitives . . . . . . . . . . . . . . . . . . . . . . . 426 end of file, on a stream . . . . . . . . . . . . . . . 135 end-of-file, on a file descriptor . . . . . . . . 153 environment . . . . . . . . . . . . . . . . . . . . . . . . . 459 environment access. . . . . . . . . . . . . . . . . . . 459 environment representation. . . . . . . . . . . 459 environment variable . . . . . . . . . . . . . . . . . 458 EOF character . . . . . . . . . . . . . . . . . . . . . . . 287 EOL character . . . . . . . . . . . . . . . . . . . . . . . 287 EOL2 character . . . . . . . . . . . . . . . . . . . . . . 287 epoch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 342 ERASE character . . . . . . . . . . . . . . . . . . . . 287 error codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 error reporting . . . . . . . . . . . . . . . . . . . . . . . . 15 establishing a handler . . . . . . . . . . . . . . . . 406 ethers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 exception . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 exclusive lock . . . . . . . . . . . . . . . . . . . . . . . . 175 exec functions . . . . . . . . . . . . . . . . . . . . . . . 470 executing a file. . . . . . . . . . . . . . . . . . . . . . . 470 exit status . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 exit status value . . . . . . . . . . . . . . . . . . . . . 462 expansion of shell words. . . . . . . . . . . . . . 333 exponent (of floating point number) . . 565 exponentiation functions . . . . . . . . . . . . . 300 extended character sets . . . . . . . . . . . . . . 367 extended characters, converting . . . . . . 374 extended strings, converting

representations . . . . . . . . . . . . . . . . . . 372 extending printf . . . . . . . . . . . . . . . . . . . . 121 extracting file descriptor from stream

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159

F

fcntl function . . . . . . . . . . . . . . . . . . . . . . . 165 feature test macros . . . . . . . . . . . . . . . . . . . . . 8 field splitting . . . . . . . . . . . . . . . . . . . . . . . . 334 FIFO special file . . . . . . . . . . . . . . . . . . . . . 215 file access permission . . . . . . . . . . . . . . . . . 204 file access time . . . . . . . . . . . . . . . . . . . . . . . 208 file attribute modification time . . . . . . . 208 file attributes . . . . . . . . . . . . . . . . . . . . . . . . 195 file creation mask . . . . . . . . . . . . . . . . . . . . 204 file descriptor flags . . . . . . . . . . . . . . . . . . . 168 file descriptor sets, for select . . . . . . . . 162 file descriptors, standard . . . . . . . . . . . . . 159 file locks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 file modification time. . . . . . . . . . . . . . . . . 208 file name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 file name component . . . . . . . . . . . . . . . . . . 88 file name errors . . . . . . . . . . . . . . . . . . . . . . . 89 file name resolution . . . . . . . . . . . . . . . . . . . 88 file name translation flags . . . . . . . . . . . . 171 file names, multiple . . . . . . . . . . . . . . . . . . 189 file namespace, for sockets. . . . . . . . . . . . 227 file owner . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 file permission bits . . . . . . . . . . . . . . . . . . . 202 file pointer . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 file position . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 file positioning on a file descriptor …. 156 file positioning on a stream. . . . . . . . . . . 137 file status flags . . . . . . . . . . . . . . . . . . . . . . . 170 filtering i/o through subprocess. . . . . . . 217 flag character (printf) . . . . . . . . . . . . . . . 106 flag character (scanf) . . . . . . . . . . . . . . . . 128 flags for sigaction. . . . . . . . . . . . . . . . . . . 412 flags, file name translation . . . . . . . . . . . 171 flags, open-time action . . . . . . . . . . . . . . . 171 floating point, IEEE . . . . . . . . . . . . . . . . . 570 floating type measurements . . . . . . . . . . 565 floating-point exception . . . . . . . . . . . . . . 397 flow control, terminal . . . . . . . . . . . . . . . . 294 flushing a stream. . . . . . . . . . . . . . . . . . . . . 141 flushing terminal output queue . . . . . . . 293 foreground job . . . . . . . . . . . . . . . . . . . . . . . 480 foreground job, launching . . . . . . . . . . . . 490 forking a process . . . . . . . . . . . . . . . . . . . . . 468 format string, for printf . . . . . . . . . . . . . 104 format string, for scanf . . . . . . . . . . . . . . 126 formatted input from a stream . . . . . . . 126 formatted output to a stream . . . . . . . . 104 freeing (obstacks) . . . . . . . . . . . . . . . . . . . . . 46 freeing memory allocated with malloc

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 fully buffered stream . . . . . . . . . . . . . . . . . 140 function prototypes (variadic) . . . . . . . . 556

G

generation of signals . . . . . . . . . . . . . . . . . 394 globbing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 graphic character. . . . . . . . . . . . . . . . . . . . . . 62 Gregorian calendar . . . . . . . . . . . . . . . . . . . 341 group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 group database . . . . . . . . . . . . . . . . . . . . . . 525 group ID. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 group name . . . . . . . . . . . . . . . . . . . . . . . . . . 511 group owner of a file . . . . . . . . . . . . . . . . . 200 grouping of digits . . . . . . . . . . . . . . . . . . . . 384 growing objects (in obstacks) . . . . . . . . . . 47

H

handle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 handling multiple signals . . . . . . . . . . . . . 418 hangup signal . . . . . . . . . . . . . . . . . . . . . . . . 400 hard limit. . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 hard link. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 header files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 heap consistency checking . . . . . . . . . . . . . 38 heap, dynamic allocation from. . . . . . . . . 32 heap, freeing memory from . . . . . . . . . . . . 34 hexadecimal digit character . . . . . . . . . . . 62 hidden bit (of floating point number

mantissa). . . . . . . . . . . . . . . . . . . . . . . . 566 high-priority data . . . . . . . . . . . . . . . . . . . . 257 high-resolution time . . . . . . . . . . . . . . . . . . 341 holes in files . . . . . . . . . . . . . . . . . . . . . . . . . 157 home directory. . . . . . . . . . . . . . . . . . . . . . . 460 HOME environment variable . . . . . . . . . 460 hook functions (of custom streams)… 148 host address, Internet . . . . . . . . . . . . . . . . 231 hosts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 hosts database . . . . . . . . . . . . . . . . . . . . . . . 235 how many arguments . . . . . . . . . . . . . . . . 557 hyperbolic functions. . . . . . . . . . . . . . . . . . 301

I

identifying terminals . . . . . . . . . . . . . . . . . 271 IEEE floating point . . . . . . . . . . . . . . . . . . 305 IEEE floating point representation … 570 IEEE Std 1003.1 . . . . . . . . . . . . . . . . . . . . . . . 2 IEEE Std 1003.2 . . . . . . . . . . . . . . . . . . . . . . . 2 ignore action for a signal . . . . . . . . . . . . . 407 illegal instruction . . . . . . . . . . . . . . . . . . . . 398 impossible events . . . . . . . . . . . . . . . . . . . . 553 independent channels . . . . . . . . . . . . . . . . 160 initial signal actions . . . . . . . . . . . . . . . . . . 413 inode number . . . . . . . . . . . . . . . . . . . . . . . . 198 input available signal. . . . . . . . . . . . . . . . . 401 input conversions, for scanf . . . . . . . . . . 129 input from multiple files. . . . . . . . . . . . . . 162 installation tools . . . . . . . . . . . . . . . . . . . . . 669 installing the library . . . . . . . . . . . . . . . . . 667 integer division functions . . . . . . . . . . . . . 309 integer type range. . . . . . . . . . . . . . . . . . . . 563 integer type width . . . . . . . . . . . . . . . . . . . 563 interactive signals, from terminal . . . . . 283 interactive stop signal . . . . . . . . . . . . . . . . 403 internationalization . . . . . . . . . . . . . . . . . . 379 Internet host address . . . . . . . . . . . . . . . . . 231 Internet namespace, for sockets. . . . . . . 229 interprocess communication, with FIFO

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 interprocess communication, with pipes

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 interprocess communication, with signals

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 430 interprocess communication, with sockets

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 interrupt character . . . . . . . . . . . . . . . . . . . 288 interrupt signal . . . . . . . . . . . . . . . . . . . . . . 399 interrupt-driven input . . . . . . . . . . . . . . . . 179 interrupting primitives . . . . . . . . . . . . . . . 426 interval timer, setting . . . . . . . . . . . . . . . . 356 INTR character . . . . . . . . . . . . . . . . . . . . . . 288 inverse hyperbolic functions . . . . . . . . . . 301 inverse trigonometric functions . . . . . . . 299 invocation of program . . . . . . . . . . . . . . . . 447 ISO C. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

J

job . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 job control . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 job control functions . . . . . . . . . . . . . . . . . 497 job control is optional . . . . . . . . . . . . . . . . 480 job control signals. . . . . . . . . . . . . . . . . . . . 402 job control, enabling . . . . . . . . . . . . . . . . . 484

K

Kermit the frog . . . . . . . . . . . . . . . . . . . . . . 321 KILL character . . . . . . . . . . . . . . . . . . . . . . 288 kill signal . . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 killing a process . . . . . . . . . . . . . . . . . . . . . . 428

L

LANG environment variable. . . . . . . . . . 461 launching jobs . . . . . . . . . . . . . . . . . . . . . . . 486 LC COLLATE environment variable

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 461 LC CTYPE environment variable …. 461 LC MONETARY environment variable

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 461 LC NUMERIC environment variable

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 461 LC TIME environment variable . . . . . . 462 leap second . . . . . . . . . . . . . . . . . . . . . . . . . . 345 length of multibyte character . . . . . . . . . 373 length of string. . . . . . . . . . . . . . . . . . . . . . . . 66 level, for socket options . . . . . . . . . . . . . . 266 library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 limits on resource usage . . . . . . . . . . . . . . 362 limits, file name length . . . . . . . . . . . . . . . 545 limits, floating types . . . . . . . . . . . . . . . . . 565 limits, integer types . . . . . . . . . . . . . . . . . . 563 limits, link count of files. . . . . . . . . . . . . . 545 limits, number of open files. . . . . . . . . . . 537 limits, number of processes . . . . . . . . . . . 537 limits, number of supplementary group

IDs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 limits, pipe buffer size . . . . . . . . . . . . . . . . 546 limits, POSIX. . . . . . . . . . . . . . . . . . . . . . . . 537 limits, program argument size . . . . . . . . 537 limits, terminal input queue . . . . . . . . . . 545 limits, time zone name length . . . . . . . . 538 line buffered stream . . . . . . . . . . . . . . . . . . 140 line speed . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 lines (in a text file). . . . . . . . . . . . . . . . . . . 136 link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 link, hard . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 link, soft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 link, symbolic . . . . . . . . . . . . . . . . . . . . . . . . 190 linked channels. . . . . . . . . . . . . . . . . . . . . . . 160 listening (sockets) . . . . . . . . . . . . . . . . . . . . 248 little-endian. . . . . . . . . . . . . . . . . . . . . . . . . . 239 LNEXT character . . . . . . . . . . . . . . . . . . . . 290 local network address number . . . . . . . . 231 local time . . . . . . . . . . . . . . . . . . . . . . . . . . . . 341 locale categories. . . . . . . . . . . . . . . . . . . . . . 380 locale, changing . . . . . . . . . . . . . . . . . . . . . . 381 locales . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 logarithm functions . . . . . . . . . . . . . . . . . . 300 login name . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 login name, determining. . . . . . . . . . . . . . 520 LOGNAME environment variable …. 460 long jumps. . . . . . . . . . . . . . . . . . . . . . . . . . . 389 long-named options . . . . . . . . . . . . . . . . . . 448 longjmp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 loss of data on sockets. . . . . . . . . . . . . . . . 221 lost resource signal . . . . . . . . . . . . . . . . . . . 404 lower-case character . . . . . . . . . . . . . . . . . . . 61

M

macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 main function . . . . . . . . . . . . . . . . . . . . . . . . 447 malloc function . . . . . . . . . . . . . . . . . . . . . . . 32 mantissa (of floating point number) . . 566 matching failure, in scanf . . . . . . . . . . . . 127 maximum field width (scanf) . . . . . . . . 128 measurements of floating types . . . . . . . 565 memory allocation . . . . . . . . . . . . . . . . . . . . 31 merging of signals . . . . . . . . . . . . . . . . . . . . 418 MIN termios slot. . . . . . . . . . . . . . . . . . . . . 291 minimum field width (printf). . . . . . . . 106 mixing descriptors and streams . . . . . . . 160 modem disconnect . . . . . . . . . . . . . . . . . . . 280 modem status lines. . . . . . . . . . . . . . . . . . . 280 monetary value formatting . . . . . . . . . . . 383 multibyte character, length of . . . . . . . . 373 multibyte characters . . . . . . . . . . . . . . . . . 368 multiple names for one file . . . . . . . . . . . 189 multiplexing input . . . . . . . . . . . . . . . . . . . 162

N

name of running program . . . . . . . . . . . . . 28 name of socket . . . . . . . . . . . . . . . . . . . . . . . 223 Name Service Switch . . . . . . . . . . . . . . . . . 501 name space . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 names of signals. . . . . . . . . . . . . . . . . . . . . . 395 namespace (of socket) . . . . . . . . . . . . . . . . 221 NaN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 netgroup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 Netgroup . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 network. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 network byte order . . . . . . . . . . . . . . . . . . . 239 network number . . . . . . . . . . . . . . . . . . . . . 231 network protocol . . . . . . . . . . . . . . . . . . . . . 222 networks database . . . . . . . . . . . . . . . . . . . 269 nisplus, and booting . . . . . . . . . . . . . . . . . 504 nisplus, and completeness . . . . . . . . . . . . 504 non-blocking open . . . . . . . . . . . . . . . . . . . 171 non-local exit, from signal handler. … 416 non-local exits . . . . . . . . . . . . . . . . . . . . . . . 389 noncanonical input processing . . . . . . . . 272 normalization functions (floating-point)

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 normalized floating point number. . . . . 566 not a number . . . . . . . . . . . . . . . . . . . . . . . . 305 NSS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 ‘nsswitch.conf’ . . . . . . . . . . . . . . . . . . . . . 502 null character . . . . . . . . . . . . . . . . . . . . . . . . . 65 null pointer constant . . . . . . . . . . . . . . . . . 561 number of arguments passed . . . . . . . . . 557 number syntax, parsing . . . . . . . . . . . . . . 310 numeric value formatting. . . . . . . . . . . . . 383

O

obstack status. . . . . . . . . . . . . . . . . . . . . . . . . 51 obstacks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 open-time action flags . . . . . . . . . . . . . . . . 171 opening a file . . . . . . . . . . . . . . . . . . . . . . . . . 85 opening a file descriptor . . . . . . . . . . . . . . 151 opening a pipe . . . . . . . . . . . . . . . . . . . . . . . 215 opening a socket . . . . . . . . . . . . . . . . . . . . . 244 opening a socket pair. . . . . . . . . . . . . . . . . 245 opening a stream. . . . . . . . . . . . . . . . . . . . . . 94 optimizing NSS . . . . . . . . . . . . . . . . . . . . . . 505 optional arguments. . . . . . . . . . . . . . . . . . . 555 optional POSIX features . . . . . . . . . . . . . 538 orphaned process group . . . . . . . . . . . . . . 481 out-of-band data . . . . . . . . . . . . . . . . . . . . . 257 output conversions, for printf . . . . . . . 107 output possible signal . . . . . . . . . . . . . . . . 401 owner of a file. . . . . . . . . . . . . . . . . . . . . . . . 200

P

packet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 page boundary . . . . . . . . . . . . . . . . . . . . . . . . 37 parent directory . . . . . . . . . . . . . . . . . . . . . . . 89 parent process. . . . . . . . . . . . . . . . . . . 467, 468 parity checking. . . . . . . . . . . . . . . . . . . . . . . 277 parsing a template string . . . . . . . . . . . . . 118 parsing numbers (in formatted input)

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 parsing program arguments . . . . . . . . . . 449 parsing tokens from a string . . . . . . . . . . . 80 passwd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 password database . . . . . . . . . . . . . . . . . . . 521 PATH environment variable . . . . . . . . . . 460 pause function . . . . . . . . . . . . . . . . . . . . . . . 439 peeking at input . . . . . . . . . . . . . . . . . . . . . 102 pending signals . . . . . . . . . . . . . . . . . . . . . . 394 pending signals, checking for . . . . . . . . . 437 permission to access a file . . . . . . . . . . . . 204 persona . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 pi (trigonometric constant) . . . . . . . . . . . 298 pipe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 pipe signal . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 pipe to a subprocess . . . . . . . . . . . . . . . . . 217 port number . . . . . . . . . . . . . . . . . . . . . . . . . 237 positioning a file descriptor. . . . . . . . . . . 156 positioning a stream . . . . . . . . . . . . . . . . . 137 POSIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 POSIX capacity limits . . . . . . . . . . . . . . . 537 POSIX optional features . . . . . . . . . . . . . 538 POSIX.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 POSIX.2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 power functions . . . . . . . . . . . . . . . . . . . . . . 300 precision (of floating point number) . . 566 precision (printf) . . . . . . . . . . . . . . . . . . . 106 predicates on arrays . . . . . . . . . . . . . . . . . . . 72 predicates on characters . . . . . . . . . . . . . . . 61 predicates on strings . . . . . . . . . . . . . . . . . . 72 primitives, interrupting. . . . . . . . . . . . . . . 426 printing character . . . . . . . . . . . . . . . . . . . . . 62 priority of a process . . . . . . . . . . . . . . . . . . 364 process. . . . . . . . . . . . . . . . . . . . . . . . . . 447, 467 process completion . . . . . . . . . . . . . . . . . . . 473 process group functions . . . . . . . . . . . . . . 497 process group ID . . . . . . . . . . . . . . . . . . . . . 486 process group leader . . . . . . . . . . . . . . . . . 486 process groups . . . . . . . . . . . . . . . . . . . . . . . 479 process ID . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 process image . . . . . . . . . . . . . . . . . . . . . . . . 468 process lifetime . . . . . . . . . . . . . . . . . . . . . . 468 process priority . . . . . . . . . . . . . . . . . . . . . . 364 process signal mask . . . . . . . . . . . . . . . . . . 433 process termination . . . . . . . . . . . . . . . . . . 462 processor time . . . . . . . . . . . . . . . . . . . . . . . 339 profiling alarm signal. . . . . . . . . . . . . . . . . 401 profiling timer . . . . . . . . . . . . . . . . . . . . . . . 356 program argument syntax . . . . . . . . . . . . 448 program arguments . . . . . . . . . . . . . . . . . . 447 program arguments, parsing . . . . . . . . . . 449 program error signals . . . . . . . . . . . . . . . . 396 program name . . . . . . . . . . . . . . . . . . . . . . . . 28 program startup . . . . . . . . . . . . . . . . . . . . . 447 program termination . . . . . . . . . . . . . . . . . 462 program termination signals . . . . . . . . . . 399 programming your own streams . . . . . . 147 protocol (of socket) . . . . . . . . . . . . . . . . . . 222 protocol family. . . . . . . . . . . . . . . . . . . . . . . 222 protocols. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 protocols database . . . . . . . . . . . . . . . . . . . 240 prototypes for variadic functions . . . . . 556 pseudo-random numbers . . . . . . . . . . . . . 302 punctuation character . . . . . . . . . . . . . . . . . 62 pushing input back. . . . . . . . . . . . . . . . . . . 102

Q

quick sort function (for arrays) . . . . . . . 318 QUIT character . . . . . . . . . . . . . . . . . . . . . . 289 quit signal . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 quote removal. . . . . . . . . . . . . . . . . . . . . . . . 334

R

race conditions, relating to job control

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 race conditions, relating to signals …. 417 radix (of floating point number) . . . . . . 565 raising signals . . . . . . . . . . . . . . . . . . . . . . . . 426 random numbers . . . . . . . . . . . . . . . . . . . . . 302 random-access files . . . . . . . . . . . . . . . . . . . . 87 range error. . . . . . . . . . . . . . . . . . . . . . . . . . . 297 range of integer type . . . . . . . . . . . . . . . . . 563 read lock. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 reading from a directory . . . . . . . . . . . . . 183 reading from a file descriptor . . . . . . . . . 153 reading from a socket . . . . . . . . . . . . . . . . 250 reading from a stream, by blocks . . . . . 104 reading from a stream, by characters . . 98 reading from a stream, formatted. . . . . 126 real group ID . . . . . . . . . . . . . . . . . . . . . . . . 511 real user ID . . . . . . . . . . . . . . . . . . . . . . . . . . 511 real-time timer . . . . . . . . . . . . . . . . . . . . . . . 356 receiving datagrams . . . . . . . . . . . . . . . . . . 261 record locking. . . . . . . . . . . . . . . . . . . . . . . . 175 redirecting input and output . . . . . . . . . 166 reentrant functions . . . . . . . . . . . . . . . . . . . 421 reentrant NSS functions . . . . . . . . . . . . . . 506 relative file name . . . . . . . . . . . . . . . . . . . . . . 88 relocating memory allocator . . . . . . . . . . . 58 remainder functions . . . . . . . . . . . . . . . . . . 308 removal of quotes . . . . . . . . . . . . . . . . . . . . 334 removing a file . . . . . . . . . . . . . . . . . . . . . . . 192 removing macros that shadow functions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 renaming a file . . . . . . . . . . . . . . . . . . . . . . . 193 reporting bugs . . . . . . . . . . . . . . . . . . . . . . . 671 reporting errors . . . . . . . . . . . . . . . . . . . . . . . 15 REPRINT character . . . . . . . . . . . . . . . . . 288 reserved names . . . . . . . . . . . . . . . . . . . . . . . . . 6 resource limits . . . . . . . . . . . . . . . . . . . . . . . 362 restarting interrupted primitives. . . . . . 426 restrictions on signal handler functions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 421 root directory . . . . . . . . . . . . . . . . . . . . . . . . . 88 rounding functions . . . . . . . . . . . . . . . . . . . 308 rpc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 running a command . . . . . . . . . . . . . . . . . . 467

S

scanning the group list . . . . . . . . . . . . . . . 527 scanning the user list. . . . . . . . . . . . . . . . . 523 search function (for arrays). . . . . . . . . . . 317 search functions (for strings) . . . . . . . . . . 78 seed (for random numbers) . . . . . . . . . . . 302 seeking on a file descriptor . . . . . . . . . . . 156 seeking on a stream . . . . . . . . . . . . . . . . . . 137 segmentation violation . . . . . . . . . . . . . . . 398 sending a datagram . . . . . . . . . . . . . . . . . . 260 sending signals . . . . . . . . . . . . . . . . . . . . . . . 426 sequential-access files . . . . . . . . . . . . . . . . . . 87 server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 services database. . . . . . . . . . . . . . . . . . . . . 238 session . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 session leader . . . . . . . . . . . . . . . . . . . . . . . . 479 setting an alarm . . . . . . . . . . . . . . . . . . . . . 356 setuid programs . . . . . . . . . . . . . . . . . . . . . 513 setuid programs and file access . . . . . . . 206 shadow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 shadowing functions with macros . . . . . . . 5 shared lock . . . . . . . . . . . . . . . . . . . . . . . . . . 176 shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 shrinking objects . . . . . . . . . . . . . . . . . . . . . . 49 shutting down a socket . . . . . . . . . . . . . . . 244 sigaction flags . . . . . . . . . . . . . . . . . . . . . . 412 sigaction function . . . . . . . . . . . . . . . . . . 408 SIGCHLD, handling of . . . . . . . . . . . . . . . . . 492 sign (of floating point number) . . . . . . . 565 signal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 393 signal action . . . . . . . . . . . . . . . . . . . . . . . . . 395 signal actions . . . . . . . . . . . . . . . . . . . . . . . . 406 signal flags . . . . . . . . . . . . . . . . . . . . . . . . . . . 412 signal function . . . . . . . . . . . . . . . . . . . . . . 406 signal handler function . . . . . . . . . . . . . . . 413 signal mask . . . . . . . . . . . . . . . . . . . . . . . . . . 433 signal messages . . . . . . . . . . . . . . . . . . . . . . 405 signal names . . . . . . . . . . . . . . . . . . . . . . . . . 395 signal number. . . . . . . . . . . . . . . . . . . . . . . . 395 signal set . . . . . . . . . . . . . . . . . . . . . . . . . . . . 432 signals, generating . . . . . . . . . . . . . . . . . . . 426 significand (of floating point number)

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 566 SIGTTIN, from background job. . . . . . . . 481 SIGTTOU, from background job. . . . . . . . 481 size of string . . . . . . . . . . . . . . . . . . . . . . . . . . 66 socket . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 socket address (name) binding. . . . . . . . 223 socket domain . . . . . . . . . . . . . . . . . . . . . . . 221 socket namespace . . . . . . . . . . . . . . . . . . . . 221 socket option level . . . . . . . . . . . . . . . . . . . 266 socket options. . . . . . . . . . . . . . . . . . . . . . . . 266 socket pair . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 socket protocol. . . . . . . . . . . . . . . . . . . . . . . 222 socket shutdown . . . . . . . . . . . . . . . . . . . . . 244 socket, client actions . . . . . . . . . . . . . . . . . 246 socket, closing . . . . . . . . . . . . . . . . . . . . . . . 244 socket, connecting. . . . . . . . . . . . . . . . . . . . 246 socket, creating . . . . . . . . . . . . . . . . . . . . . . 244 socket, initiating a connection . . . . . . . . 246 sockets, accepting connections . . . . . . . . 248 sockets, listening . . . . . . . . . . . . . . . . . . . . . 248 sockets, server actions . . . . . . . . . . . . . . . . 248 soft limit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 soft link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 sort function (for arrays) . . . . . . . . . . . . . 318 sparse files . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 special files . . . . . . . . . . . . . . . . . . . . . . . . . . 210 specified action (for a signal) . . . . . . . . . 395 square root function. . . . . . . . . . . . . . . . . . 300 stable sorting . . . . . . . . . . . . . . . . . . . . . . . . 318 standard dot notation, for Internet

addresses . . . . . . . . . . . . . . . . . . . . . . . . 232 standard environment variables. . . . . . . 460 standard error file descriptor . . . . . . . . . 160 standard error stream . . . . . . . . . . . . . . . . . 94 standard file descriptors . . . . . . . . . . . . . . 159 standard input file descriptor . . . . . . . . . 159 standard input stream. . . . . . . . . . . . . . . . . 93 standard output file descriptor . . . . . . . 160 standard output stream . . . . . . . . . . . . . . . 93 standard streams . . . . . . . . . . . . . . . . . . . . . . 93 standards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 START character . . . . . . . . . . . . . . . . . . . . 290 startup of program . . . . . . . . . . . . . . . . . . . 447 static allocation . . . . . . . . . . . . . . . . . . . . . . . 31 STATUS character . . . . . . . . . . . . . . . . . . . 291 status codes. . . . . . . . . . . . . . . . . . . . . . . . . . . 15 status of a file. . . . . . . . . . . . . . . . . . . . . . . . 195 status of obstack . . . . . . . . . . . . . . . . . . . . . . 51 sticky bit . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 STOP character. . . . . . . . . . . . . . . . . . . . . . 290 stop signal . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 stopped job . . . . . . . . . . . . . . . . . . . . . . . . . . 480 stopped jobs, continuing . . . . . . . . . . . . . 496 stopped jobs, detecting . . . . . . . . . . . . . . . 492 storage allocation . . . . . . . . . . . . . . . . . . . . . 31 stream (sockets) . . . . . . . . . . . . . . . . . . . . . 221 stream, for I/O to a string . . . . . . . . . . . 144 streams and descriptors . . . . . . . . . . . . . . 160 streams, and file descriptors . . . . . . . . . . 159 streams, standard . . . . . . . . . . . . . . . . . . . . . 93 string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 string allocation. . . . . . . . . . . . . . . . . . . . . . . 66 string collation functions . . . . . . . . . . . . . . 75 string comparison functions . . . . . . . . . . . 72 string concatenation functions . . . . . . . . . 67 string copy functions . . . . . . . . . . . . . . . . . . 67 string length . . . . . . . . . . . . . . . . . . . . . . . . . . 66 string literal . . . . . . . . . . . . . . . . . . . . . . . . . . 65 string search functions . . . . . . . . . . . . . . . . 78 string stream . . . . . . . . . . . . . . . . . . . . . . . . 144 string, representation of . . . . . . . . . . . . . . . 65 style of communication (of a socket) . . 221 subshell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 substitution of variables and commands

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 successive signals. . . . . . . . . . . . . . . . . . . . . 418 summer time. . . . . . . . . . . . . . . . . . . . . . . . . 346 SunOS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 supplementary group IDs. . . . . . . . . . . . . 512 SUSP character . . . . . . . . . . . . . . . . . . . . . . 289 suspend character . . . . . . . . . . . . . . . . . . . . 289 SVID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 symbolic link . . . . . . . . . . . . . . . . . . . . . . . . 190 symbolic link, opening . . . . . . . . . . . . . . . 172 syntax, for program arguments . . . . . . . 448 syntax, for reading numbers . . . . . . . . . . 310 System V Unix . . . . . . . . . . . . . . . . . . . . . . . . . 3

T

TCP (Internet protocol). . . . . . . . . . . . . . 240 template, for printf . . . . . . . . . . . . . . . . . 104 template, for scanf . . . . . . . . . . . . . . . . . . 126 TERM environment variable . . . . . . . . . 461 terminal flow control . . . . . . . . . . . . . . . . . 294 terminal identification. . . . . . . . . . . . . . . . 271 terminal input queue . . . . . . . . . . . . . . . . . 271 terminal input queue, clearing . . . . . . . . 294 terminal input signal . . . . . . . . . . . . . . . . . 403 terminal line control functions . . . . . . . . 293 terminal line speed . . . . . . . . . . . . . . . . . . . 284 terminal mode data types . . . . . . . . . . . . 273 terminal mode functions . . . . . . . . . . . . . 274 terminal output queue . . . . . . . . . . . . . . . 272 terminal output queue, flushing . . . . . . 293 terminal output signal . . . . . . . . . . . . . . . 403 terminated jobs, detecting. . . . . . . . . . . . 492 termination signal. . . . . . . . . . . . . . . . . . . . 399 testing access permission . . . . . . . . . . . . . 206 testing exit status of child process …. 473 text stream . . . . . . . . . . . . . . . . . . . . . . . . . . 136 ticks, clock. . . . . . . . . . . . . . . . . . . . . . . . . . . 339 tilde expansion. . . . . . . . . . . . . . . . . . . . . . . 334 TIME termios slot . . . . . . . . . . . . . . . . . . . 291 time zone . . . . . . . . . . . . . . . . . . . . . . . . . . . . 352 time zone database. . . . . . . . . . . . . . . . . . . 354 time, calendar . . . . . . . . . . . . . . . . . . . . . . . 341 time, elapsed CPU . . . . . . . . . . . . . . . . . . . 339 timer, profiling . . . . . . . . . . . . . . . . . . . . . . . 356 timer, real-time . . . . . . . . . . . . . . . . . . . . . . 356 timer, virtual . . . . . . . . . . . . . . . . . . . . . . . . 356 timers, setting . . . . . . . . . . . . . . . . . . . . . . . 356 timing error in signal handling . . . . . . . 439 TMPDIR environment variable . . . . . . . 212 tokenizing strings . . . . . . . . . . . . . . . . . . . . . 80 tools, for installing library. . . . . . . . . . . . 669 transmitting datagrams . . . . . . . . . . . . . . 260 trigonometric functions. . . . . . . . . . . . . . . 298 type measurements, floating . . . . . . . . . . 565 type measurements, integer. . . . . . . . . . . 563 type modifier character (printf) . . . . . 107 type modifier character (scanf) . . . . . . 128 typeahead buffer . . . . . . . . . . . . . . . . . . . . . 271 TZ environment variable . . . . . . . . . . . . . 461

U

umask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 unbuffered stream . . . . . . . . . . . . . . . . . . . . 140 unconstrained storage allocation . . . . . . . 32 undefining macros that shadow functions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Unix, Berkeley . . . . . . . . . . . . . . . . . . . . . . . . . 3 Unix, System V . . . . . . . . . . . . . . . . . . . . . . . . 3 unlinking a file . . . . . . . . . . . . . . . . . . . . . . . 192 unreading characters . . . . . . . . . . . . . . . . . 102 upper-case character . . . . . . . . . . . . . . . . . . 61 urgent data signal. . . . . . . . . . . . . . . . . . . . 401 urgent socket condition. . . . . . . . . . . . . . . 257 usage limits . . . . . . . . . . . . . . . . . . . . . . . . . . 362 user database . . . . . . . . . . . . . . . . . . . . . . . . 521 user ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 user ID, determining . . . . . . . . . . . . . . . . . 520 user name . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 user signals . . . . . . . . . . . . . . . . . . . . . . . . . . 404 usual file name errors . . . . . . . . . . . . . . . . . 89

V

variable number of arguments . . . . . . . . 555 variable substitution . . . . . . . . . . . . . . . . . 334 variable-sized arrays. . . . . . . . . . . . . . . . . . . 57 variadic function argument access …. 556 variadic function prototypes . . . . . . . . . . 556 variadic functions . . . . . . . . . . . . . . . . . . . . 555 variadic functions, calling . . . . . . . . . . . . 558 virtual time alarm signal . . . . . . . . . . . . . 401 virtual timer . . . . . . . . . . . . . . . . . . . . . . . . . 356 volatile declarations . . . . . . . . . . . . . . . . 421

W

waiting for a signal. . . . . . . . . . . . . . . . . . . 439 waiting for completion of child process

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 waiting for input or output . . . . . . . . . . . 162 WERASE character . . . . . . . . . . . . . . . . . . 288 whitespace character . . . . . . . . . . . . . . . . . . 62 wide characters . . . . . . . . . . . . . . . . . . . . . . 367 width of integer type . . . . . . . . . . . . . . . . . 563 wildcard expansion. . . . . . . . . . . . . . . . . . . 334 word expansion . . . . . . . . . . . . . . . . . . . . . . 333 working directory . . . . . . . . . . . . . . . . . . . . 181 write lock . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 writing to a file descriptor . . . . . . . . . . . . 155 writing to a socket . . . . . . . . . . . . . . . . . . . 250 writing to a stream, by blocks . . . . . . . . 104 writing to a stream, by characters . . . . . 97 writing to a stream, formatted . . . . . . . 104

708 The GNU C Library Type Index 709

Type Index

C

cc_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 clock_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 340 comparison_fn_t . . . . . . . . . . . . . . . . . . . . 317

cookie_close_function . . . . . . . . . . . . . 149
cookie_io_functions_t . . . . . . . . . . . . . 147
cookie_read_function. . . . . . . . . . . . . . . 149
cookie_seek_function. . . . . . . . . . . . . . . 149
cookie_write_function . . . . . . . . . . . . . 149

D

dev_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 DIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 div_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309

E

enum mcheck_status . . . . . . . . . . . . . . . . . . 39

F

fd_set. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 FILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 fpos_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139

G

gid_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 glob_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324

I

ino_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198

J

jmp_buf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391

L

ldiv_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310

M

mode_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197

N

nlink_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198

O

off_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158

P

pid_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468

printf_arginfo_function . . . . . . . . . . . 124

printf_function . . . . . . . . . . . . . . . . . . . . 124 ptrdiff_t . . . . . . . . . . . . . . . . . . . . . . . . . . . 562

R

regex_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 328 regmatch_t . . . . . . . . . . . . . . . . . . . . . . . . . . 331 regoff_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . 331

S

sig_atomic_t . . . . . . . . . . . . . . . . . . . . . . . . 424 sighandler_t . . . . . . . . . . . . . . . . . . . . . . . . 406 sigjmp_buf . . . . . . . . . . . . . . . . . . . . . . . . . . 392 sigset_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . 432 size_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 562 speed_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 ssize_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 struct dirent . . . . . . . . . . . . . . . . . . . . . . . 183 struct flock . . . . . . . . . . . . . . . . . . . . . . . . 176 struct group . . . . . . . . . . . . . . . . . . . . . . . . 525 struct hostent . . . . . . . . . . . . . . . . . . . . . . 235 struct in_addr . . . . . . . . . . . . . . . . . . . . . . 232 struct in6_addr. . . . . . . . . . . . . . . . . . . . . 233 struct itimerval. . . . . . . . . . . . . . . . . . . . 357 struct lconv . . . . . . . . . . . . . . . . . . . . . . . . 384 struct linger . . . . . . . . . . . . . . . . . . . . . . . 268 struct mallinfo . . . . . . . . . . . . . . . . . . . . . . 41 struct netent . . . . . . . . . . . . . . . . . . . . . . . 269 struct obstack . . . . . . . . . . . . . . . . . . . . . . . 43 struct option . . . . . . . . . . . . . . . . . . . . . . . 452 struct passwd . . . . . . . . . . . . . . . . . . . . . . . 521 struct printf_info . . . . . . . . . . . . . . . . . 122 struct protoent. . . . . . . . . . . . . . . . . . . . . 241 struct rlimit . . . . . . . . . . . . . . . . . . . . . . . 362 struct rusage . . . . . . . . . . . . . . . . . . . . . . . 360 struct servent . . . . . . . . . . . . . . . . . . . . . . 238 struct sigaction. . . . . . . . . . . . . . . . . . . . 408 struct sigaltstack . . . . . . . . . . . . . . . . . 442 struct sigstack. . . . . . . . . . . . . . . . . . . . . 444 struct sigvec . . . . . . . . . . . . . . . . . . . . . . . 445 struct sockaddr. . . . . . . . . . . . . . . . . . . . . 224 struct sockaddr_in . . . . . . . . . . . . . . . . . 230 struct sockaddr_in6 . . . . . . . . . . . . . . . . 230 struct sockaddr_un . . . . . . . . . . . . . . . . . 228 struct stat . . . . . . . . . . . . . . . . . . . . . . . . . 196 struct termios . . . . . . . . . . . . . . . . . . . . . . 273 struct timeval . . . . . . . . . . . . . . . . . . . . . . 342 struct timezone. . . . . . . . . . . . . . . . . . . . . 343 struct tm. . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 struct tms . . . . . . . . . . . . . . . . . . . . . . . . . . 340 struct utimbuf . . . . . . . . . . . . . . . . . . . . . . 208 struct utsname . . . . . . . . . . . . . . . . . . . . . . 534

T

tcflag_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 time_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 342

U

uid_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 union wait . . . . . . . . . . . . . . . . . . . . . . . . . . 476

V

va_list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559

W

wchar_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 wordexp_t . . . . . . . . . . . . . . . . . . . . . . . . . . . 335

Function and Macro Index 711

Function and Macro Index

_exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 465 _tolower . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 _toupper . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

A

abort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 465 abs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 accept. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 access. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 acos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 acosh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 adjtime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 alarm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 358 alloca . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 alphasort . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 asctime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 347 asin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 asinh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 asprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 assert. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 assert_perror. . . . . . . . . . . . . . . . . . . . . . . 554 atan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 atan2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 atanh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 atexit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 464 atof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 atoi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 atol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312

B

bcmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 bcopy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 bind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 bsearch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 bzero . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

C

cabs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 calloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 cbrt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 ceil . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 cfgetispeed . . . . . . . . . . . . . . . . . . . . . . . . . 285 cfgetospeed . . . . . . . . . . . . . . . . . . . . . . . . . 285 cfmakeraw . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 cfree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 cfsetispeed . . . . . . . . . . . . . . . . . . . . . . . . . 285 cfsetospeed . . . . . . . . . . . . . . . . . . . . . . . . . 285 cfsetspeed . . . . . . . . . . . . . . . . . . . . . . . . . . 285 chdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 chmod . . . . . . . . . . . . . . . . . . . . . . . . . . . 204, 205 chown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 clearerr . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 clock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 340 close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 closedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 confstr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 551 connect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 copysign . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 cos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 cosh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 creat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 ctermid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497 ctime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 348 cuserid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 521

D

difftime . . . . . . . . . . . . . . . . . . . . . . . . . . . . 342 div . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 drem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 DTTOIF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 dup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 dup2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167

E

endgrent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 endhostent . . . . . . . . . . . . . . . . . . . . . . . . . . 237 endnetent . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 endnetgrent . . . . . . . . . . . . . . . . . . . . . . . . . 530 endprotoent . . . . . . . . . . . . . . . . . . . . . . . . . 242 endpwent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 524 endservent . . . . . . . . . . . . . . . . . . . . . . . . . . 239 execl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471 execle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471 execlp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471 execv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 470 execve. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471 execvp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471 exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 462 exp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 expm1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301

F

fabs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 fchmod. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 fchown. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 fclean. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 fclose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 fcloseall . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 fcntl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 FD_CLR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 FD_ISSET . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 FD_SET. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 FD_ZERO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 fdopen. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 feof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 ferror. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 fflush. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 fgetc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 fgetgrent . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 fgetgrent_r . . . . . . . . . . . . . . . . . . . . . . . . . 527 fgetpos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 fgetpwent . . . . . . . . . . . . . . . . . . . . . . . . . . . 523 fgetpwent_r . . . . . . . . . . . . . . . . . . . . . . . . . 523 fgets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 fileno. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 finite. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 floor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 fmemopen . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 fmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 fnmatch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 fopen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 fopencookie . . . . . . . . . . . . . . . . . . . . . . . . . 148 fork . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 469 fpathconf . . . . . . . . . . . . . . . . . . . . . . . . . . . 548 fprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 fputc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 fputs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 fread . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 freopen. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 frexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 fscanf. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 fseek . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 fsetpos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 fstat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 ftell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 fwrite. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

G

getc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 getchar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 getcwd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 getdelim . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 getegid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 getenv. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 459 geteuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 getgid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 getgrent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 getgrent_r . . . . . . . . . . . . . . . . . . . . . . . . . . 527 getgrgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . 526 getgrgid_r . . . . . . . . . . . . . . . . . . . . . . . . . . 526 getgrnam . . . . . . . . . . . . . . . . . . . . . . . . . . . . 526 getgrnam_r . . . . . . . . . . . . . . . . . . . . . . . . . . 526 getgroups . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 gethostbyaddr. . . . . . . . . . . . . . . . . . . . . . . 236 gethostbyname. . . . . . . . . . . . . . . . . . . . . . . 236 gethostbyname2 . . . . . . . . . . . . . . . . . . . . . 236 gethostent . . . . . . . . . . . . . . . . . . . . . . . . . . 237 gethostid . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 gethostname . . . . . . . . . . . . . . . . . . . . . . . . . 533 getitimer . . . . . . . . . . . . . . . . . . . . . . . . . . . 357 getline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 getlogin . . . . . . . . . . . . . . . . . . . . . . . . . . . . 520 getnetbyaddr . . . . . . . . . . . . . . . . . . . . . . . . 269 getnetbyname . . . . . . . . . . . . . . . . . . . . . . . . 269 getnetent . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 getnetgrent . . . . . . . . . . . . . . . . . . . . . . . . . 529 getnetgrent_r. . . . . . . . . . . . . . . . . . . . . . . 529 getopt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 449 getopt_long . . . . . . . . . . . . . . . . . . . . . . . . . 453 getpeername . . . . . . . . . . . . . . . . . . . . . . . . . 250 getpgrp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 getpid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 getppid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 469 getpriority . . . . . . . . . . . . . . . . . . . . . . . . . 364 getprotobyname . . . . . . . . . . . . . . . . . . . . . 241 getprotobynumber . . . . . . . . . . . . . . . . . . . 241 getprotoent . . . . . . . . . . . . . . . . . . . . . . . . . 242 getpwent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 524 getpwent_r . . . . . . . . . . . . . . . . . . . . . . . . . . 524 getpwnam . . . . . . . . . . . . . . . . . . . . . . . . . . . . 523 getpwnam_r . . . . . . . . . . . . . . . . . . . . . . . . . . 523 getpwuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . 522 getpwuid_r . . . . . . . . . . . . . . . . . . . . . . . . . . 522 getrlimit . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 getrusage . . . . . . . . . . . . . . . . . . . . . . . . . . . 360 gets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 getservbyname. . . . . . . . . . . . . . . . . . . . . . . 239 getservbyport. . . . . . . . . . . . . . . . . . . . . . . 239 getservent . . . . . . . . . . . . . . . . . . . . . . . . . . 239 getsockname . . . . . . . . . . . . . . . . . . . . . . . . . 226 getsockopt . . . . . . . . . . . . . . . . . . . . . . . . . . 266 getsubopt . . . . . . . . . . . . . . . . . . . . . . . . . . . 456 gettimeofday . . . . . . . . . . . . . . . . . . . . . . . . 344 getuid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 getumask . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 getw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 getwd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 glob . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 gmtime. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 347 gsignal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 427

H

htonl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 htons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 hypot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301

I

IFTODT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 inet_addr . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 inet_aton . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 inet_lnaof . . . . . . . . . . . . . . . . . . . . . . . . . . 234 inet_makeaddr. . . . . . . . . . . . . . . . . . . . . . . 234 inet_netof . . . . . . . . . . . . . . . . . . . . . . . . . . 234 inet_network . . . . . . . . . . . . . . . . . . . . . . . . 234 inet_ntoa . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 inet_ntop . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 inet_pton . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 infnan. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 initgroups . . . . . . . . . . . . . . . . . . . . . . . . . . 516 initstate . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 innetgr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 isalnum. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 isalpha. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 isascii. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 isatty. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 isblank. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 iscntrl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 isdigit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 isgraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 isinf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 islower. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 isnan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 isprint. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 ispunct. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 isspace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 isupper. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 isxdigit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 ITIMER_PROF . . . . . . . . . . . . . . . . . . . . . . . . . 358 ITIMER_REAL . . . . . . . . . . . . . . . . . . . . . . . . . 358 ITIMER_VIRTUAL . . . . . . . . . . . . . . . . . . . . . 358

K

kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428 killpg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429

L

labs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 ldexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 ldiv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 listen. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 localeconv . . . . . . . . . . . . . . . . . . . . . . . . . . 383 localtime . . . . . . . . . . . . . . . . . . . . . . . . . . . 346 log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 log10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 log1p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 logb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 longjmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391 lseek . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 lstat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198

M

main . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447 mallinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 malloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 mallopt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 mblen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 373 mbstowcs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 372 mbtowc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 374 mcheck . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 memalign . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 memccpy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 memchr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 memcmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 memcpy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 memmem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 memmove. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 memset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 mkdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 mkfifo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 mknod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 mkstemp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 mktemp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 mktime. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 347 modf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 mprobe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

N

nice . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 365 notfound . . . . . . . . . . . . . . . . . . . . . . . . . . . . 503

NSS_STATUS_NOTFOUND. . . . . . . . . . . . . . . . 507

NSS_STATUS_SUCCESS . . . . . . . . . . . . . . . . . 507

NSS_STATUS_TRYAGAIN. . . . . . . . . . . . . . . . 507

NSS_STATUS_UNAVAIL . . . . . . . . . . . . . . . . . 507 ntohl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 ntohs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240

O

obstack_1grow. . . . . . . . . . . . . . . . . . . . . . . . 48 obstack_1grow_fast . . . . . . . . . . . . . . . . . . 50

obstack_alignment_mask . . . . . . . . . . . . . 52

obstack_alloc. . . . . . . . . . . . . . . . . . . . . . . . 45 obstack_base . . . . . . . . . . . . . . . . . . . . . . . . . 51 obstack_blank. . . . . . . . . . . . . . . . . . . . . . . . 48 obstack_blank_fast . . . . . . . . . . . . . . . . . . 50

obstack_chunk_alloc . . . . . . . . . . . . . . . . . 44

obstack_chunk_free . . . . . . . . . . . . . . . . . . 44 obstack_chunk_size . . . . . . . . . . . . . . . . . . 53 obstack_copy . . . . . . . . . . . . . . . . . . . . . . . . . 46 obstack_copy0. . . . . . . . . . . . . . . . . . . . . . . . 46 obstack_finish. . . . . . . . . . . . . . . . . . . . . . . 49 obstack_free . . . . . . . . . . . . . . . . . . . . . . . . . 46 obstack_grow . . . . . . . . . . . . . . . . . . . . . . . . . 48 obstack_grow0. . . . . . . . . . . . . . . . . . . . . . . . 48 obstack_init . . . . . . . . . . . . . . . . . . . . . . . . . 44 obstack_int_grow . . . . . . . . . . . . . . . . . . . . 48

obstack_int_grow_fast . . . . . . . . . . . . . . 50

obstack_next_free . . . . . . . . . . . . . . . . . . . 51

obstack_object_size . . . . . . . . . . . . . 49, 52

obstack_printf . . . . . . . . . . . . . . . . . . . . . 115 obstack_ptr_grow . . . . . . . . . . . . . . . . . . . . 48

obstack_ptr_grow_fast . . . . . . . . . . . . . . 50

obstack_room . . . . . . . . . . . . . . . . . . . . . . . . . 50 obstack_vprintf . . . . . . . . . . . . . . . . . . . . 117 offsetof . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 on_exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 464 open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 open_memstream . . . . . . . . . . . . . . . . . . . . . 145

open_obstack_stream. . . . . . . . . . . . . . . . 146

opendir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184

P

parse_printf_format. . . . . . . . . . . . . . . . 118

pathconf . . . . . . . . . . . . . . . . . . . . . . . . . . . . 548 pause . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 pclose. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 perror . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 pipe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 popen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 pow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 printf. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 psignal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 405 putc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 putchar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 putenv. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 459 putpwent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525 puts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 putw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

Q

qsort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318

R

r_alloc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 r_alloc_free . . . . . . . . . . . . . . . . . . . . . . . . . 58 r_re_alloc . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 raise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 rand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 random. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 readdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 readdir_r . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 readlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 realloc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 recv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 recvfrom . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 regcomp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 328 regerror . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 regexec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 330 regfree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333

register_printf_function . . . . . . . . . . 121

remove. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 rename. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 rewind. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 rewinddir . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 rindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 rint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 rmdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193

S

S_ISBLK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 S_ISCHR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 S_ISDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 S_ISFIFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 S_ISLNK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 S_ISREG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 S_ISSOCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 scalb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 scandir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 scanf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 seekdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 select. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 send . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 sendto. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 setbuf. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 setbuffer . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 setgid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 setgrent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 setgroups . . . . . . . . . . . . . . . . . . . . . . . . . . . 516 sethostent . . . . . . . . . . . . . . . . . . . . . . . . . . 237 sethostid . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 sethostname . . . . . . . . . . . . . . . . . . . . . . . . . 533 setitimer . . . . . . . . . . . . . . . . . . . . . . . . . . . 357 setjmp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391 setlinebuf . . . . . . . . . . . . . . . . . . . . . . . . . . 143 setlocale . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 setnetent . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 setnetgrent . . . . . . . . . . . . . . . . . . . . . . . . . 529 setpgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 setpgrp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 setpriority . . . . . . . . . . . . . . . . . . . . . . . . . 364 setprotoent . . . . . . . . . . . . . . . . . . . . . . . . . 241 setpwent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 524 setregid . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 setreuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 setrlimit . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 setservent . . . . . . . . . . . . . . . . . . . . . . . . . . 239 setsid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 setsockopt . . . . . . . . . . . . . . . . . . . . . . . . . . 267 setstate . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 settimeofday . . . . . . . . . . . . . . . . . . . . . . . . 344 setuid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 setvbuf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 shutdown . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 sigaction . . . . . . . . . . . . . . . . . . . . . . . . . . . 409 sigaddset . . . . . . . . . . . . . . . . . . . . . . . . . . . 433 sigaltstack . . . . . . . . . . . . . . . . . . . . . . . . . 443 sigblock . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 sigdelset . . . . . . . . . . . . . . . . . . . . . . . . . . . 433 sigemptyset . . . . . . . . . . . . . . . . . . . . . . . . . 433 sigfillset . . . . . . . . . . . . . . . . . . . . . . . . . . 433 siginterrupt . . . . . . . . . . . . . . . . . . . . . . . . 446 sigismember . . . . . . . . . . . . . . . . . . . . . . . . . 433 siglongjmp . . . . . . . . . . . . . . . . . . . . . . . . . . 392 sigmask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 signal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 sigpause . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 sigpending . . . . . . . . . . . . . . . . . . . . . . . . . . 437 sigprocmask . . . . . . . . . . . . . . . . . . . . . . . . . 434 sigsetjmp . . . . . . . . . . . . . . . . . . . . . . . . . . . 392 sigsetmask . . . . . . . . . . . . . . . . . . . . . . . . . . 446 sigstack . . . . . . . . . . . . . . . . . . . . . . . . . . . . 444 sigsuspend . . . . . . . . . . . . . . . . . . . . . . . . . . 441 sigvec. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 sin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 sinh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 sleep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 snprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 socket. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 socketpair . . . . . . . . . . . . . . . . . . . . . . . . . . 245 sprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 sqrt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 srand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 srandom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 sscanf. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 ssignal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 408 stat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 stpcpy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 stpncpy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 strcasecmp . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 strcat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 strchr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 strcmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 strcoll. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 strcpy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 strcspn. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 strdup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 strdupa. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 strerror . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 strerror_r . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 strftime . . . . . . . . . . . . . . . . . . . . . . . . . . . . 348 strlen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 strncasecmp . . . . . . . . . . . . . . . . . . . . . . . . . . 74 strncat. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 strncmp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 strncpy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 strndup. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 strndupa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 strpbrk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 strrchr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 strsep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 strsignal . . . . . . . . . . . . . . . . . . . . . . . . . . . 405 strspn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 strstr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 strtod. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 strtof. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 strtok . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 strtok_r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 strtol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 strtold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 strtoll . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 strtoq. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 strtoul . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 strtoull . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 strtouq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 strxfrm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 success . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 503 symlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 sysconf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541 system. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 467

T

tan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 tanh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 tcdrain . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 tcflow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 tcflush . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 tcgetattr . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 tcgetpgrp . . . . . . . . . . . . . . . . . . . . . . . . . . . 500 tcsendbreak . . . . . . . . . . . . . . . . . . . . . . . . . 293 tcsetattr . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 tcsetpgrp . . . . . . . . . . . . . . . . . . . . . . . . . . . 500 telldir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 TEMP_FAILURE_RETRY . . . . . . . . . . . . . . . . . 425 tempnam . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 342 times . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 341 tmpfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 tmpnam. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 tmpnam_r . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 toascii. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 tolower. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 toupper. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 tryagain . . . . . . . . . . . . . . . . . . . . . . . . . . . . 504 ttyname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 tzset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 355

U

umask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 uname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 535 unavail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 503 ungetc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 unlink. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 utime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 utimes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209

V

va_alist . . . . . . . . . . . . . . . . . . . . . . . . . . . . 561 va_arg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559 va_dcl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 561 va_end. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559 va_start. . . . . . . . . . . . . . . . . . . . . . . . 559, 561 valloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 vasprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 vfork . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 470 vfprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 vfscanf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 vprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 vscanf. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 vsnprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 vsprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 vsscanf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

W

wait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 wait3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 wait4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 waitpid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 WCOREDUMP . . . . . . . . . . . . . . . . . . . . . . . . . . . 476 wcstombs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 372 wctomb. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 374 WEXITSTATUS . . . . . . . . . . . . . . . . . . . . . . . . . 476 WIFEXITED . . . . . . . . . . . . . . . . . . . . . . . . . . . 476 WIFSIGNALED . . . . . . . . . . . . . . . . . . . . . . . . . 476 WIFSTOPPED . . . . . . . . . . . . . . . . . . . . . . . . . . 476 wordexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 335 wordfree . . . . . . . . . . . . . . . . . . . . . . . . . . . . 336 write . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 WSTOPSIG . . . . . . . . . . . . . . . . . . . . . . . . . . . . 476 WTERMSIG . . . . . . . . . . . . . . . . . . . . . . . . . . . . 476

718 The GNU C Library

Variable and Constant Macro Index 719

Variable and Constant Macro Index

__free_hook . . . . . . . . . . . . . . . . . . . . . . . . . . 40 __malloc_hook. . . . . . . . . . . . . . . . . . . . . . . . 40 __realloc_hook. . . . . . . . . . . . . . . . . . . . . . . 40 _BSD_SOURCE . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 _GNU_SOURCE . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 _IOFBF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 _IOLBF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 _IONBF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 _POSIX_C_SOURCE. . . . . . . . . . . . . . . . . . . . . . . 8

_POSIX_CHOWN_RESTRICTED . . . . . . . . . . . 547

_POSIX_JOB_CONTROL . . . . . . . . . . . . . . . . . 539 _POSIX_NO_TRUNC . . . . . . . . . . . . . . . . . . . . 547 _POSIX_SAVED_IDS . . . . . . . . . . . . . . . . . . . 539 _POSIX_SOURCE . . . . . . . . . . . . . . . . . . . . . . . . . 8 _POSIX_VDISABLE. . . . . . . . . . . . . . . . 286, 547 _POSIX_VERSION . . . . . . . . . . . . . . . . . . . . . 540 _POSIX2_C_DEV. . . . . . . . . . . . . . . . . . . . . . . 539 _POSIX2_C_VERSION . . . . . . . . . . . . . . . . . . 540 _POSIX2_FORT_DEV . . . . . . . . . . . . . . . . . . . 539 _POSIX2_FORT_RUN . . . . . . . . . . . . . . . . . . . 539 _POSIX2_LOCALEDEF . . . . . . . . . . . . . . . . . . 539 _POSIX2_SW_DEV . . . . . . . . . . . . . . . . . . . . . 540 _REENTRANT . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 _SVID_SOURCE . . . . . . . . . . . . . . . . . . . . . . . . . . 9 _THREAD_SAFE . . . . . . . . . . . . . . . . . . . . . . . . . 10 _XOPEN_SOURCE . . . . . . . . . . . . . . . . . . . . . . . . . 9

A

AF_FILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 AF_INET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 AF_UNIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 AF_UNSPEC . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 aliases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 ALTWERASE . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 ARG_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 537

B

B0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 B110 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 B1200 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 B134 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 B150 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 B1800 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 B19200. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 B200 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 B2400 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 B300 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 B38400. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 B4800 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 B50 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 B600 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 B75 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 B9600 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 BC_BASE_MAX . . . . . . . . . . . . . . . . . . . . . . . . . 549 BC_DIM_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . 550 BC_SCALE_MAX . . . . . . . . . . . . . . . . . . . . . . . . 549 BC_STRING_MAX. . . . . . . . . . . . . . . . . . . . . . . 550 BRKINT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 BUFSIZ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142

C

CCTS_OFLOW . . . . . . . . . . . . . . . . . . . . . . . . . . 281 CHILD_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . . 537 CIGNORE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 CLK_TCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 340 CLOCAL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 CLOCKS_PER_SEC . . . . . . . . . . . . . . . . . . . . . 340 COLL_WEIGHTS_MAX . . . . . . . . . . . . . . . . . . . 550 COREFILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . 396 CREAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 CRTS_IFLOW . . . . . . . . . . . . . . . . . . . . . . . . . . 281 CS5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 CS6 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 CS7 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 CS8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 CSIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 CSTOPB. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280

D

daylight . . . . . . . . . . . . . . . . . . . . . . . . . . . . 355

E

E2BIG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 EACCES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 EADDRINUSE . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 EADDRNOTAVAIL. . . . . . . . . . . . . . . . . . . . . . . . 22 EADV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 EAFNOSUPPORT . . . . . . . . . . . . . . . . . . . . . . . . . 22 EAGAIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 EALREADY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 EAUTH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 EBACKGROUND . . . . . . . . . . . . . . . . . . . . . . . . . . 25 EBADE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 EBADF . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17, 295 EBADFD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 EBADMSG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 EBADR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 EBADRPC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 EBADRQC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 EBADSLT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 EBFONT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 EBUSY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ECHILD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ECHO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 ECHOCTL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 ECHOE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 ECHOK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 ECHOKE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 ECHONL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 ECHOPRT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 ECHRNG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ECOMM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ECONNABORTED . . . . . . . . . . . . . . . . . . . . . . . . . 22 ECONNREFUSED . . . . . . . . . . . . . . . . . . . . . . . . . 23 ECONNRESET . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 ED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 EDEADLK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 EDEADLOCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 EDESTADDRREQ . . . . . . . . . . . . . . . . . . . . . . . . . 23 EDIED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 EDOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 EDOTDOT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 EDQUOT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 EEXIST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 EFAULT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 EFBIG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 EFTYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 EGRATUITOUS . . . . . . . . . . . . . . . . . . . . . . . . . . 25 EGREGIOUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 EHOSTDOWN . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 EHOSTUNREACH . . . . . . . . . . . . . . . . . . . . . . . . . 23 EIDRM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 EIEIO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 EILSEQ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 EINPROGRESS . . . . . . . . . . . . . . . . . . . . . . . . . . 20 EINTR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 EINVAL . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 295 EIO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 EISCONN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 EISDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 EISNAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 EL2HLT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 EL2NSYNC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 EL3HLT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 EL3RST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ELIBACC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ELIBBAD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ELIBEXEC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ELIBMAX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ELIBSCN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ELNRNG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ELOOP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 EMEDIUMTYPE . . . . . . . . . . . . . . . . . . . . . . . . . . 27 EMFILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 EMLINK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 EMSGSIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 EMULTIHOP . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ENAMETOOLONG . . . . . . . . . . . . . . . . . . . . . . . . . 23 ENAVAIL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ENEEDAUTH . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ENETDOWN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 ENETRESET . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 ENETUNREACH . . . . . . . . . . . . . . . . . . . . . . . . . . 22 ENFILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ENOANO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ENOBUFS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 ENOCSI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ENODATA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ENODEV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ENOENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 ENOEXEC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ENOLCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ENOLINK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ENOMEDIUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ENOMEM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ENOMSG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ENONET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ENOPKG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ENOPROTOOPT . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ENOSPC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ENOSR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ENOSTR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ENOSYS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ENOTBLK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ENOTCONN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 ENOTDIR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ENOTEMPTY . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 ENOTNAM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ENOTSOCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ENOTTY . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 295 ENOTUNIQ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 environ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 ENXIO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 EOF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 EOPNOTSUPP . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 EOVERFLOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 EPERM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 EPFNOSUPPORT . . . . . . . . . . . . . . . . . . . . . . . . . 21 EPIPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 EPROCLIM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 EPROCUNAVAIL . . . . . . . . . . . . . . . . . . . . . . . . . 24 EPROGMISMATCH. . . . . . . . . . . . . . . . . . . . . . . . 24 EPROGUNAVAIL . . . . . . . . . . . . . . . . . . . . . . . . . 24 EPROTO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 EPROTONOSUPPORT . . . . . . . . . . . . . . . . . . . . . 21 EPROTOTYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 EQUIV_CLASS_MAX . . . . . . . . . . . . . . . . . . . . 550 ERANGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 EREMCHG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 EREMOTE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 EREMOTEIO . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ERESTART . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 EROFS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ERPCMISMATCH . . . . . . . . . . . . . . . . . . . . . . . . . 24 errno . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ESHUTDOWN . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 ESOCKTNOSUPPORT . . . . . . . . . . . . . . . . . . . . . 21 ESPIPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ESRCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ESRMNT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ESTALE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ESTRPIPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ethers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 ETIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ETIMEDOUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 ETOOMANYREFS . . . . . . . . . . . . . . . . . . . . . . . . . 23 ETXTBSY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 EUCLEAN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 EUNATCH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 EUSERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 EWOULDBLOCK . . . . . . . . . . . . . . . . . . . . . . . . . . 20 EXDEV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 EXFULL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 EXIT_FAILURE . . . . . . . . . . . . . . . . . . . . . . . . 463 EXIT_SUCCESS . . . . . . . . . . . . . . . . . . . . . . . . 463 EXPR_NEST_MAX. . . . . . . . . . . . . . . . . . . . . . . 550 EXTA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 EXTB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286

F

F_DUPFD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 F_GETFD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 F_GETFL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 F_GETLK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 F_GETOWN . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 F_OK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 F_RDLCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 F_SETFD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 F_SETFL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 F_SETLK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 F_SETLKW . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 F_SETOWN . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 F_UNLCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 F_WRLCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 FD_CLOEXEC . . . . . . . . . . . . . . . . . . . . . . . . . . 169 FD_SETSIZE . . . . . . . . . . . . . . . . . . . . . . . . . . 163 FILENAME_MAX . . . . . . . . . . . . . . . . . . . . . . . . 546 FLUSHO. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 FOPEN_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 FPE_DECOVF_TRAP . . . . . . . . . . . . . . . . . . . . 397 FPE_FLTDIV_TRAP . . . . . . . . . . . . . . . . . . . . 397 FPE_FLTOVF_TRAP . . . . . . . . . . . . . . . . . . . . 397 FPE_FLTUND_TRAP . . . . . . . . . . . . . . . . . . . . 397 FPE_INTDIV_TRAP . . . . . . . . . . . . . . . . . . . . 397 FPE_INTOVF_TRAP . . . . . . . . . . . . . . . . . . . . 397 FPE_SUBRNG_TRAP . . . . . . . . . . . . . . . . . . . . 397

G

group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501

H

h_errno . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 HOST_NOT_FOUND . . . . . . . . . . . . . . . . . . . . . 236 hosts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 HUGE_VAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 HUGE_VALf . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 HUGE_VALl . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 HUPCL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280

I

ICANON. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 ICRNL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 IEXTEN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 IGNBRK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 IGNCR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 IGNPAR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 IMAXBEL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 in6addr_any . . . . . . . . . . . . . . . . . . . . . . . . . 233 in6addr_loopback. . . . . . . . . . . . . . . . . . . 233 INADDR_ANY . . . . . . . . . . . . . . . . . . . . . . . . . . 233 INADDR_BROADCAST . . . . . . . . . . . . . . . . . . . 233 INADDR_LOOPBACK . . . . . . . . . . . . . . . . . . . . 232 INADDR_NONE . . . . . . . . . . . . . . . . . . . . . . . . . 233 INLCR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 INPCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 IPPORT_RESERVED . . . . . . . . . . . . . . . . . . . . 238

IPPORT_USERRESERVED. . . . . . . . . . . . . . . . 238

ISIG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 ISTRIP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 IXANY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 IXOFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 IXON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278

L

L_ctermid . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 L_cuserid . . . . . . . . . . . . . . . . . . . . . . . . . . . 521 L_INCR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 L_SET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 L_tmpnam . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 L_XTND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 LANG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 LC_ALL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 LC_COLLATE . . . . . . . . . . . . . . . . . . . . . . . . . . 380 LC_CTYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 LC_MESSAGES . . . . . . . . . . . . . . . . . . . . . . . . . 381 LC_MONETARY . . . . . . . . . . . . . . . . . . . . . . . . . 381 LC_NUMERIC . . . . . . . . . . . . . . . . . . . . . . . . . . 381 LC_TIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 LINE_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . 550 LINK_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . 545

M

MAX_CANON . . . . . . . . . . . . . . . . . . . . . . . . . . . 545 MAX_INPUT . . . . . . . . . . . . . . . . . . . . . . . . . . . 545 MAXNAMLEN . . . . . . . . . . . . . . . . . . . . . . . . . . . 546 MB_CUR_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . 370 MB_LEN_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . 370 MDMBUF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 MINSIGSTKSZ . . . . . . . . . . . . . . . . . . . . . . . . . 443 MSG_DONTROUTE. . . . . . . . . . . . . . . . . . . . . . . 252 MSG_OOB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252 MSG_PEEK . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252

N

NAME_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . 545 NAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 NCCS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 NDEBUG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 netgroup . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 network . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 NGROUPS_MAX . . . . . . . . . . . . . . . . . . . . . . . . . 538 NO_ADDRESS . . . . . . . . . . . . . . . . . . . . . . . . . . 237 NO_RECOVERY . . . . . . . . . . . . . . . . . . . . . . . . . 237 NOFLSH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 NOKERNINFO . . . . . . . . . . . . . . . . . . . . . . . . . . 284 NSIG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 396

NSS_STATUS_NOTFOUND. . . . . . . . . . . . . . . . 507

NSS_STATUS_SUCCESS . . . . . . . . . . . . . . . . . 507

NSS_STATUS_TRYAGAIN. . . . . . . . . . . . . . . . 507

NSS_STATUS_UNAVAIL . . . . . . . . . . . . . . . . . 507 NULL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 561

O

O_ACCMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 O_APPEND . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 O_ASYNC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 O_CREAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 O_EXCL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 O_EXEC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 O_EXLOCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 O_FSYNC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 O_IGNORE_CTTY. . . . . . . . . . . . . . . . . . . . . . . 172 O_NDELAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 O_NOATIME . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 O_NOCTTY . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 O_NOLINK . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 O_NONBLOCK . . . . . . . . . . . . . . . . . . . . . 171, 173 O_NOTRANS . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 O_RDONLY . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 O_RDWR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 O_READ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 O_SHLOCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 O_SYNC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 O_TRUNC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 O_WRITE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 O_WRONLY . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 ONLCR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 ONOEOT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 OPEN_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . 537 OPOST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 optarg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 449 opterr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 449 optind. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 449 optopt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 449 OXTABS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279

P

P_tmpdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 PA_CHAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 PA_DOUBLE . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 PA_FLAG_LONG . . . . . . . . . . . . . . . . . . . . . . . . 119

PA_FLAG_LONG_DOUBLE. . . . . . . . . . . . . . . . 119

PA_FLAG_LONG_LONG . . . . . . . . . . . . . . . . . . 119 PA_FLAG_MASK . . . . . . . . . . . . . . . . . . . . . . . . 118 PA_FLAG_PTR . . . . . . . . . . . . . . . . . . . . . . . . . 119 PA_FLAG_SHORT. . . . . . . . . . . . . . . . . . . . . . . 119 PA_FLOAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 PA_INT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 PA_LAST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 PA_POINTER . . . . . . . . . . . . . . . . . . . . . . . . . . 118 PA_STRING . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 PARENB. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 PARMRK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 PARODD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 passwd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 PATH_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . 546 PENDIN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 PF_CCITT . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 PF_FILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 PF_IMPLINK . . . . . . . . . . . . . . . . . . . . . . . . . . 243 PF_INET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 PF_ISO. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 PF_NS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 PF_ROUTE . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 PF_UNIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 PIPE_BUF . . . . . . . . . . . . . . . . . . . . . . . . . . . . 546 PRIO_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . 364 PRIO_MIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . 364 PRIO_PGRP . . . . . . . . . . . . . . . . . . . . . . . . . . . 365 PRIO_PROCESS . . . . . . . . . . . . . . . . . . . . . . . . 365 PRIO_USER . . . . . . . . . . . . . . . . . . . . . . . . . . . 365

program_invocation_name . . . . . . . . . . . . 28
program_invocation_short_name . . . . . 28

protocols . . . . . . . . . . . . . . . . . . . . . . . . . . . 502

R

R_OK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 RAND_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 RE_DUP_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . 538 RLIM_NLIMITS . . . . . . . . . . . . . . . . . . . . . . . . 363 RLIMIT_CORE . . . . . . . . . . . . . . . . . . . . . . . . . 363 RLIMIT_CPU . . . . . . . . . . . . . . . . . . . . . . . . . . 362 RLIMIT_DATA . . . . . . . . . . . . . . . . . . . . . . . . . 363 RLIMIT_FSIZE . . . . . . . . . . . . . . . . . . . . . . . . 363 RLIMIT_NOFILE. . . . . . . . . . . . . . . . . . . . . . . 363 RLIMIT_OFILE . . . . . . . . . . . . . . . . . . . . . . . . 363 RLIMIT_RSS . . . . . . . . . . . . . . . . . . . . . . . . . . 363 RLIMIT_STACK . . . . . . . . . . . . . . . . . . . . . . . . 363 rpc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502

S

S_IEXEC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 S_IFBLK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 S_IFCHR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 S_IFDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 S_IFIFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 S_IFLNK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 S_IFMT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 S_IFREG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 S_IFSOCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 S_IREAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 S_IRGRP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 S_IROTH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 S_IRUSR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 S_IRWXG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 S_IRWXO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 S_IRWXU . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 S_ISGID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 S_ISUID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 S_ISVTX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 S_IWGRP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 S_IWOTH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 S_IWRITE . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 S_IWUSR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 S_IXGRP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 S_IXOTH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 S_IXUSR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 SA_DISABLE . . . . . . . . . . . . . . . . . . . . . . . . . . 443 SA_NOCLDSTOP . . . . . . . . . . . . . . . . . . . . . . . . 412 SA_ONSTACK . . . . . . . . . . . . . . . . . . . . . 412, 443 SA_RESTART . . . . . . . . . . . . . . . . . . . . . . . . . . 412 SEEK_CUR . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 SEEK_END . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 SEEK_SET . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 services . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 shadow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 502 SIG_BLOCK . . . . . . . . . . . . . . . . . . . . . . . . . . . 434 SIG_DFL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 SIG_ERR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 408 SIG_IGN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 SIG_SETMASK . . . . . . . . . . . . . . . . . . . . . . . . . 434 SIG_UNBLOCK . . . . . . . . . . . . . . . . . . . . . . . . . 434 SIGABRT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 398 SIGALRM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 SIGBUS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 398 SIGCHLD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 SIGCLD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 SIGCONT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 SIGEMT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 399 SIGFPE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 396 SIGHUP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 SIGILL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 398 SIGINFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 405 SIGINT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 399 SIGIO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 SIGIOT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 398 SIGKILL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 SIGLOST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 SIGPIPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 SIGPOLL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 SIGPROF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 SIGQUIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 SIGSEGV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 398 SIGSTKSZ . . . . . . . . . . . . . . . . . . . . . . . . . . . . 443 SIGSTOP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 SIGSYS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 399 SIGTERM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 399 SIGTRAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 399 SIGTSTP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 SIGTTIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 SIGTTOU . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 SIGURG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 SIGUSR1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 SIGUSR2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 SIGVTALRM . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 SIGWINCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . 405 SIGXCPU . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 SIGXFSZ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 SOCK_DGRAM . . . . . . . . . . . . . . . . . . . . . . . . . . 223 SOCK_RAW . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 SOCK_STREAM . . . . . . . . . . . . . . . . . . . . . . . . . 223 SOL_SOCKET . . . . . . . . . . . . . . . . . . . . . . . . . . 267 SSIZE_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 stderr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 STDERR_FILENO. . . . . . . . . . . . . . . . . . . . . . . 160 stdin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 STDIN_FILENO . . . . . . . . . . . . . . . . . . . . . . . . 159 stdout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 STDOUT_FILENO. . . . . . . . . . . . . . . . . . . . . . . 160 STREAM_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . 537 SV_INTERRUPT . . . . . . . . . . . . . . . . . . . . . . . . 445 SV_ONSTACK . . . . . . . . . . . . . . . . . . . . . . . . . . 445 SV_RESETHAND . . . . . . . . . . . . . . . . . . . . . . . . 446 sys_siglist . . . . . . . . . . . . . . . . . . . . . . . . . 406

T

TCIFLUSH . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 TCIOFF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 TCIOFLUSH . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 TCION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 TCOFLUSH . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 TCOOFF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 TCOON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 TCSADRAIN . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 TCSAFLUSH . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 TCSANOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 TCSASOFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 timezone . . . . . . . . . . . . . . . . . . . . . . . . . . . . 355 TMP_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 TOSTOP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 TRY_AGAIN . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 tzname. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 354 TZNAME_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . 538

V

VDISCARD . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 VDSUSP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 VEOF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 VEOL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 VEOL2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 VERASE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 VINTR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 VKILL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 VLNEXT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 VMIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 VQUIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 VREPRINT . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 VSTART. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 VSTATUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 VSTOP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 VSUSP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 VTIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 VWERASE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288

W

W_OK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207

X

X_OK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208

726 The GNU C Library Program and File Index 727

Program and File Index

-

-lbsd-compat . . . . . . . . . . . . . . . . . . . . . 9, 498

/

/etc/group . . . . . . . . . . . . . . . . . . . . . . . . . . 525 /etc/hosts . . . . . . . . . . . . . . . . . . . . . . . . . . 235 /etc/localtime . . . . . . . . . . . . . . . . . . . . . 354 /etc/networks. . . . . . . . . . . . . . . . . . . . . . . 269 /etc/passwd . . . . . . . . . . . . . . . . . . . . . . . . . 521 /etc/protocols . . . . . . . . . . . . . . . . . . . . . 240 /etc/services. . . . . . . . . . . . . . . . . . . . . . . 238

/share/lib/zoneinfo. . . . . . . . . . . . . . . . 354

A

arpa/inet.h . . . . . . . . . . . . . . . . . . . . . . . . . 233 assert.h . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553

B

bsd-compat. . . . . . . . . . . . . . . . . . . . . . . . 9, 498

C

cd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 chgrp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 chown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 ctype.h . . . . . . . . . . . . . . . . . . . . . . . . . . . 61, 63

D

dirent.h. . . . . . . . . . . . 7, 183, 184, 185, 187

E

errno.h . . . . . . . . . . . . . . . . . . . . . . . . . . . 15, 16

F

fcntl.h… 8, 151, 165, 166, 168, 170, 176,

179 float.h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 566 fnmatch.h . . . . . . . . . . . . . . . . . . . . . . . . . . . 323

G

gcc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 grp.h. . . . . . . . . . . . . . . . . . . . . . . . . 8, 516, 525

H

hostid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 533 hostname . . . . . . . . . . . . . . . . . . . . . . . . . . . . 533

K

kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 399

L

limits.h. . . . . . . . . . . . 8, 370, 537, 545, 563 locale.h. . . . . . . . . . . . . . . . . . . . . . . . 381, 383 localtime . . . . . . . . . . . . . . . . . . . . . . . . . . . 354 ls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195

M

malloc.h. . . . . . . . . . . . . . . . . . . 38, 40, 41, 58 math.h. . . . . . . . . . . . 297, 305, 306, 307, 308 mkdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194

N

netdb.h . . . . . . . . . . . . . . . 235, 238, 241, 269 netinet/in.h . . . . . . . . . 230, 232, 238, 240

O

obstack.h . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

P

printf.h. . . . . . . . . . . . . . . . . . . . . . . . 121, 122 pwd.h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 521

S

setjmp.h. . . . . . . . . . . . . . . . . . . . . . . . 391, 392 sh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 467 signal.h . . . . . . 8, 395, 406, 408, 412, 426,

428, 432, 434, 437, 445 stdarg.h. . . . . . . . . . . . . . . . . . . . . . . . 556, 559 stddef.h. . . . . . . . . . . . . . . . . . . . . . . . 371, 562 stdio.h . . . . . 93, 94, 97, 98, 104, 113, 116,

134, 135, 137, 139, 141, 144, 147, 159, 193, 211, 406, 497, 520

stdlib.h . . . . . . 32, 34, 35, 36, 37, 55, 302, 303, 306, 309, 310, 314, 317, 318, 370, 372, 373, 374, 459, 463, 465, 467

string.h . . . . . . . . . . 67, 73, 75, 78, 80, 405 sys/param.h . . . . . . . . . . . . . . . . . . . . . . . . . 533 sys/resource.h . . . . . . . . . . . . 360, 362, 364 sys/socket.h …. 223, 224, 225, 226, 227,

229, 244, 245, 250, 251, 252, 260, 266, 267

sys/stat.h …. 8, 195, 196, 199, 202, 205,

210, 219 sys/time.h. . . . . . . . . . . . . . . . . 209, 342, 357 sys/times.h . . . . . . . . . . . . . . . . . . . . . . 8, 340 sys/types.h . . . . . 162, 468, 498, 499, 513,

514, 515 sys/un.h . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 sys/utsname.h. . . . . . . . . . . . . . . . . . . . . . . 534 sys/wait.h . . . . . . . . . . . . . . . . . . . . . 473, 476

T

termios.h. . . . . . . . . . . . . . . . . . . . . . . . . 8, 273 time.h. . . . . . . . . . . . 208, 339, 342, 347, 352

U

umask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 unistd.h …. 151, 153, 159, 166, 181, 189, 191, 192, 193, 201, 207, 215, 271, 357, 449, 465, 468, 469, 470, 498, 499, 513, 514, 515, 520, 533, 538, 546, 547 utime.h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208

V

varargs.h . . . . . . . . . . . . . . . . . . . . . . . . . . . 560

Z

zoneinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . 354

Short Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Error Reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3 Memory Allocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 4 Character Handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 5 String and Array Utilities. . . . . . . . . . . . . . . . . . . . . . . . . 65 6 Input/Output Overview . . . . . . . . . . . . . . . . . . . . . . . . . . 85 7 Input/Output on Streams . . . . . . . . . . . . . . . . . . . . . . . . 93 8 Low-Level Input/Output . . . . . . . . . . . . . . . . . . . . . . . . 151 9 File System Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 10 Pipes and FIFOs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 11 Sockets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 12 Low-Level Terminal Interface . . . . . . . . . . . . . . . . . . . . . 271 13 Mathematics. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 14 Low-Level Arithmetic Functions . . . . . . . . . . . . . . . . . . . 305 15 Searching and Sorting . . . . . . . . . . . . . . . . . . . . . . . . . . 317 16 Pattern Matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 17 Date and Time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 18 Extended Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . 367 19 Locales and Internationalization . . . . . . . . . . . . . . . . . . . 379 20 Non-Local Exits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 389 21 Signal Handling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 393 22 Process Startup and Termination. . . . . . . . . . . . . . . . . . . 447 23 Processes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 467 24 Job Control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 25 System Databases and Name Service Switch . . . . . . . . . . . 501 26 Users and Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 27 System Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . 533 28 System Configuration Parameters . . . . . . . . . . . . . . . . . . 537 Appendix A C Language Facilities in the Library . . . . . . . . . . 553 Appendix B Summary of Library Facilities . . . . . . . . . . . . . . 573

ii The GNU C Library

Appendix C Library Maintenance . . . . . . . . . . . . . . . . . . . . 667 Appendix D GNU LIBRARY GENERAL PUBLIC LICENSE . 687 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 697 Type Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 709 Function and Macro Index . . . . . . . . . . . . . . . . . . . . . . . . . . 711 Variable and Constant Macro Index . . . . . . . . . . . . . . . . . . . . 719 Program and File Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . 727

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Getting Started . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Standards and Portability. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.2.1 ISO C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2.2 POSIX (The Portable Operating System Interface)

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2.3 Berkeley Unix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2.4 SVID (The System V Interface Description) . . . . . . 3

1.3 Using the Library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3.1 Header Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3.2 Macro Definitions of Functions. . . . . . . . . . . . . . . . . . . 5 1.3.3 Reserved Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.3.4 Feature Test Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

1.4 Roadmap to the Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

2 Error Reporting . . . . . . . . . . . . . . . . . . . . . . . . . . 15

2.1 Checking for Errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.2 Error Codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.3 Error Messages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

3 Memory Allocation . . . . . . . . . . . . . . . . . . . . . . . 31

3.1 Dynamic Memory Allocation Concepts. . . . . . . . . . . . . . . . . . . 31 3.2 Dynamic Allocation and C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.3 Unconstrained Allocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

3.3.1 Basic Storage Allocation . . . . . . . . . . . . . . . . . . . . . . . 32 3.3.2 Examples of malloc . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 3.3.3 Freeing Memory Allocated with malloc . . . . . . . . . 34 3.3.4 Changing the Size of a Block . . . . . . . . . . . . . . . . . . . 35 3.3.5 Allocating Cleared Space . . . . . . . . . . . . . . . . . . . . . . . 36 3.3.6 Efficiency Considerations for malloc . . . . . . . . . . . . 37 3.3.7 Allocating Aligned Memory Blocks. . . . . . . . . . . . . . 37 3.3.8 Malloc Tunable Parameters. . . . . . . . . . . . . . . . . . . . . 38 3.3.9 Heap Consistency Checking . . . . . . . . . . . . . . . . . . . . 38 3.3.10 Storage Allocation Hooks . . . . . . . . . . . . . . . . . . . . . 40 3.3.11 Statistics for Storage Allocation with malloc … 41 3.3.12 Summary of malloc-Related Functions . . . . . . . . . 42

3.4 Obstacks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 3.4.1 Creating Obstacks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 3.4.2 Preparing for Using Obstacks . . . . . . . . . . . . . . . . . . . 44 3.4.3 Allocation in an Obstack . . . . . . . . . . . . . . . . . . . . . . . 45 3.4.4 Freeing Objects in an Obstack . . . . . . . . . . . . . . . . . . 46 3.4.5 Obstack Functions and Macros . . . . . . . . . . . . . . . . . 47 3.4.6 Growing Objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 3.4.7 Extra Fast Growing Objects . . . . . . . . . . . . . . . . . . . . 49 3.4.8 Status of an Obstack. . . . . . . . . . . . . . . . . . . . . . . . . . . 51 3.4.9 Alignment of Data in Obstacks . . . . . . . . . . . . . . . . . 52 3.4.10 Obstack Chunks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 3.4.11 Summary of Obstack Functions . . . . . . . . . . . . . . . . 53

3.5 Automatic Storage with Variable Size . . . . . . . . . . . . . . . . . . . . 55 3.5.1 alloca Example. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 3.5.2 Advantages of alloca. . . . . . . . . . . . . . . . . . . . . . . . . . 56 3.5.3 Disadvantages of alloca . . . . . . . . . . . . . . . . . . . . . . . 57 3.5.4 GNU C Variable-Size Arrays . . . . . . . . . . . . . . . . . . . 57

3.6 Relocating Allocator. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 3.6.1 Concepts of Relocating Allocation . . . . . . . . . . . . . . 58 3.6.2 Allocating and Freeing Relocatable Blocks . . . . . . . 58

4 Character Handling . . . . . . . . . . . . . . . . . . . . . . . 61

4.1 Classification of Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 4.2 Case Conversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

5 String and Array Utilities . . . . . . . . . . . . . . . . . 65

5.1 Representation of Strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 5.2 String and Array Conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 5.3 String Length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 5.4 Copying and Concatenation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 5.5 String/Array Comparison . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 5.6 Collation Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 5.7 Search Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 5.8 Finding Tokens in a String . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

6 Input/Output Overview . . . . . . . . . . . . . . . . . . . 85

6.1 Input/Output Concepts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 6.1.1 Streams and File Descriptors . . . . . . . . . . . . . . . . . . . 85 6.1.2 File Position . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

6.2 File Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 6.2.1 Directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 6.2.2 File Name Resolution . . . . . . . . . . . . . . . . . . . . . . . . . . 88 6.2.3 File Name Errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 6.2.4 Portability of File Names. . . . . . . . . . . . . . . . . . . . . . . 90

7 Input/Output on Streams . . . . . . . . . . . . . . . . . 93

7.1 Streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 7.2 Standard Streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 7.3 Opening Streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 7.4 Closing Streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 7.5 Simple Output by Characters or Lines . . . . . . . . . . . . . . . . . . . 97 7.6 Character Input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 7.7 Line-Oriented Input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 7.8 Unreading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

7.8.1 What Unreading Means . . . . . . . . . . . . . . . . . . . . . . . 102 7.8.2 Using ungetc To Do Unreading. . . . . . . . . . . . . . . . 102

7.9 Block Input/Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 7.10 Formatted Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

7.10.1 Formatted Output Basics . . . . . . . . . . . . . . . . . . . . 105 7.10.2 Output Conversion Syntax . . . . . . . . . . . . . . . . . . . 106 7.10.3 Table of Output Conversions . . . . . . . . . . . . . . . . . 107 7.10.4 Integer Conversions . . . . . . . . . . . . . . . . . . . . . . . . . . 108 7.10.5 Floating-Point Conversions . . . . . . . . . . . . . . . . . . . 110 7.10.6 Other Output Conversions . . . . . . . . . . . . . . . . . . . 112 7.10.7 Formatted Output Functions . . . . . . . . . . . . . . . . . 113 7.10.8 Dynamically Allocating Formatted Output …. 114 7.10.9 Variable Arguments Output Functions . . . . . . . . 115 7.10.10 Parsing a Template String. . . . . . . . . . . . . . . . . . . 118 7.10.11 Example of Parsing a Template String . . . . . . . 119

7.11 Customizing printf. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 7.11.1 Registering New Conversions . . . . . . . . . . . . . . . . . 121 7.11.2 Conversion Specifier Options . . . . . . . . . . . . . . . . . 122 7.11.3 Defining the Output Handler . . . . . . . . . . . . . . . . . 123 7.11.4 printf Extension Example. . . . . . . . . . . . . . . . . . . 124

7.12 Formatted Input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 7.12.1 Formatted Input Basics . . . . . . . . . . . . . . . . . . . . . . 126 7.12.2 Input Conversion Syntax . . . . . . . . . . . . . . . . . . . . . 127 7.12.3 Table of Input Conversions . . . . . . . . . . . . . . . . . . . 129 7.12.4 Numeric Input Conversions. . . . . . . . . . . . . . . . . . . 130 7.12.5 String Input Conversions . . . . . . . . . . . . . . . . . . . . . 131 7.12.6 Dynamically Allocating String Conversions …. 133 7.12.7 Other Input Conversions . . . . . . . . . . . . . . . . . . . . . 133 7.12.8 Formatted Input Functions . . . . . . . . . . . . . . . . . . . 134 7.12.9 Variable Arguments Input Functions . . . . . . . . . . 134

7.13 End-Of-File and Errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 7.14 Text and Binary Streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 7.15 File Positioning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 7.16 Portable File-Position Functions. . . . . . . . . . . . . . . . . . . . . . . 138 7.17 Stream Buffering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140

7.17.1 Buffering Concepts . . . . . . . . . . . . . . . . . . . . . . . . . . 140 7.17.2 Flushing Buffers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 7.17.3 Controlling Which Kind of Buffering . . . . . . . . . . 141

7.18 Other Kinds of Streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 7.18.1 String Streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 7.18.2 Obstack Streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 7.18.3 Programming Your Own Custom Streams . . . . . 147

7.18.3.1 Custom Streams and Cookies . . . . . . . . 147 7.18.3.2 Custom Stream Hook Functions. . . . . . 148

8 Low-Level Input/Output . . . . . . . . . . . . . . . . . 151

8.1 Opening and Closing Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 8.2 Input and Output Primitives . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 8.3 Setting the File Position of a Descriptor . . . . . . . . . . . . . . . . 156 8.4 Descriptors and Streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 8.5 Dangers of Mixing Streams and Descriptors . . . . . . . . . . . . . 160

8.5.1 Linked Channels. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 8.5.2 Independent Channels . . . . . . . . . . . . . . . . . . . . . . . . 160 8.5.3 Cleaning Streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161

8.6 Waiting for Input or Output . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 8.7 Control Operations on Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 8.8 Duplicating Descriptors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 8.9 File Descriptor Flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 8.10 File Status Flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170

8.10.1 File Access Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 8.10.2 Open-time Flags. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 8.10.3 I/O Operating Modes . . . . . . . . . . . . . . . . . . . . . . . . 173 8.10.4 Getting and Setting File Status Flags . . . . . . . . . 174

8.11 File Locks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 8.12 Interrupt-Driven Input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179

9 File System Interface . . . . . . . . . . . . . . . . . . . . 181

9.1 Working Directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 9.2 Accessing Directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183

9.2.1 Format of a Directory Entry. . . . . . . . . . . . . . . . . . . 183 9.2.2 Opening a Directory Stream . . . . . . . . . . . . . . . . . . . 184 9.2.3 Reading and Closing a Directory Stream . . . . . . . 185 9.2.4 Simple Program to List a Directory . . . . . . . . . . . . 186 9.2.5 Random Access in a Directory Stream. . . . . . . . . . 187 9.2.6 Scanning the Content of a Directory . . . . . . . . . . . 187 9.2.7 Simple Program to List a Directory, Mark II. … 188

9.3 Hard Links . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 9.4 Symbolic Links . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 9.5 Deleting Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 9.6 Renaming Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 9.7 Creating Directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 9.8 File Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195

9.8.1 What the File Attribute Values Mean . . . . . . . . . . 195 9.8.2 Reading the Attributes of a File . . . . . . . . . . . . . . . 198 9.8.3 Testing the Type of a File . . . . . . . . . . . . . . . . . . . . . 199 9.8.4 File Owner. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 9.8.5 The Mode Bits for Access Permission . . . . . . . . . . 202 9.8.6 How Your Access to a File is Decided . . . . . . . . . . 204 9.8.7 Assigning File Permissions . . . . . . . . . . . . . . . . . . . . 204 9.8.8 Testing Permission to Access a File . . . . . . . . . . . . 206 9.8.9 File Times . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208

9.9 Making Special Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 9.10 Temporary Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210

10 Pipes and FIFOs . . . . . . . . . . . . . . . . . . . . . . . 215

10.1 Creating a Pipe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 10.2 Pipe to a Subprocess . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 10.3 FIFO Special Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 10.4 Atomicity of Pipe I/O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220

11 Sockets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221

11.1 Socket Concepts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 11.2 Communication Styles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 11.3 Socket Addresses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223

11.3.1 Address Formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 11.3.2 Setting the Address of a Socket . . . . . . . . . . . . . . . 225 11.3.3 Reading the Address of a Socket . . . . . . . . . . . . . . 226

11.4 The File Namespace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 11.4.1 File Namespace Concepts . . . . . . . . . . . . . . . . . . . . 227 11.4.2 Details of File Namespace . . . . . . . . . . . . . . . . . . . . 227 11.4.3 Example of File-Namespace Sockets . . . . . . . . . . . 228

11.5 The Internet Namespace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 11.5.1 Internet Socket Address Formats. . . . . . . . . . . . . . 230 11.5.2 Host Addresses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231

11.5.2.1 Internet Host Addresses . . . . . . . . . . . . . 231 11.5.2.2 Host Address Data Type . . . . . . . . . . . . 232 11.5.2.3 Host Address Functions . . . . . . . . . . . . . 233 11.5.2.4 Host Names . . . . . . . . . . . . . . . . . . . . . . . . 235

11.5.3 Internet Ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 11.5.4 The Services Database . . . . . . . . . . . . . . . . . . . . . . . 238 11.5.5 Byte Order Conversion. . . . . . . . . . . . . . . . . . . . . . . 239 11.5.6 Protocols Database . . . . . . . . . . . . . . . . . . . . . . . . . . 240 11.5.7 Internet Socket Example . . . . . . . . . . . . . . . . . . . . . 242

11.6 Other Namespaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 11.7 Opening and Closing Sockets. . . . . . . . . . . . . . . . . . . . . . . . . . 244

11.7.1 Creating a Socket. . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 11.7.2 Closing a Socket. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 11.7.3 Socket Pairs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245

11.8 Using Sockets with Connections . . . . . . . . . . . . . . . . . . . . . . . 246 11.8.1 Making a Connection . . . . . . . . . . . . . . . . . . . . . . . . 246 11.8.2 Listening for Connections . . . . . . . . . . . . . . . . . . . . 248 11.8.3 Accepting Connections . . . . . . . . . . . . . . . . . . . . . . . 248 11.8.4 Who is Connected to Me? . . . . . . . . . . . . . . . . . . . . 250 11.8.5 Transferring Data . . . . . . . . . . . . . . . . . . . . . . . . . . . 250

11.8.5.1 Sending Data . . . . . . . . . . . . . . . . . . . . . . . 250 11.8.5.2 Receiving Data . . . . . . . . . . . . . . . . . . . . . 251 11.8.5.3 Socket Data Options . . . . . . . . . . . . . . . . 252

11.8.6 Byte Stream Socket Example . . . . . . . . . . . . . . . . . 252 11.8.7 Byte Stream Connection Server Example . . . . . . 254 11.8.8 Out-of-Band Data . . . . . . . . . . . . . . . . . . . . . . . . . . . 257

11.9 Datagram Socket Operations . . . . . . . . . . . . . . . . . . . . . . . . . . 260 11.9.1 Sending Datagrams . . . . . . . . . . . . . . . . . . . . . . . . . . 260 11.9.2 Receiving Datagrams . . . . . . . . . . . . . . . . . . . . . . . . 261 11.9.3 Datagram Socket Example . . . . . . . . . . . . . . . . . . . 261 11.9.4 Example of Reading Datagrams . . . . . . . . . . . . . . 263

11.10 The inetd Daemon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 11.10.1 inetd Servers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 11.10.2 Configuring inetd . . . . . . . . . . . . . . . . . . . . . . . . . . 265

11.11 Socket Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 11.11.1 Socket Option Functions . . . . . . . . . . . . . . . . . . . . 266 11.11.2 Socket-Level Options . . . . . . . . . . . . . . . . . . . . . . . 267

11.12 Networks Database. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269

12 Low-Level Terminal Interface. . . . . . . . . . . . 271

12.1 Identifying Terminals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 12.2 I/O Queues. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 12.3 Two Styles of Input: Canonical or Not . . . . . . . . . . . . . . . . . 272 12.4 Terminal Modes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273

12.4.1 Terminal Mode Data Types . . . . . . . . . . . . . . . . . . 273 12.4.2 Terminal Mode Functions . . . . . . . . . . . . . . . . . . . . 274 12.4.3 Setting Terminal Modes Properly . . . . . . . . . . . . . 276 12.4.4 Input Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 12.4.5 Output Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 12.4.6 Control Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 12.4.7 Local Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 12.4.8 Line Speed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 12.4.9 Special Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . 286

12.4.9.1 Characters for Input Editing . . . . . . . . . 287 12.4.9.2 Characters that Cause Signals . . . . . . . 288 12.4.9.3 Special Characters for Flow Control . . 290 12.4.9.4 Other Special Characters . . . . . . . . . . . . 290

12.4.10 Noncanonical Input. . . . . . . . . . . . . . . . . . . . . . . . . 291 12.5 Line Control Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 12.6 Noncanonical Mode Example. . . . . . . . . . . . . . . . . . . . . . . . . . 295

13 Mathematics . . . . . . . . . . . . . . . . . . . . . . . . . . . 297

13.1 Domain and Range Errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 13.2 Trigonometric Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 13.3 Inverse Trigonometric Functions . . . . . . . . . . . . . . . . . . . . . . . 299 13.4 Exponentiation and Logarithms . . . . . . . . . . . . . . . . . . . . . . . 300 13.5 Hyperbolic Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 13.6 Pseudo-Random Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302

13.6.1 ISO C Random Number Functions . . . . . . . . . . . . 302 13.6.2 BSD Random Number Functions . . . . . . . . . . . . . 303

14 Low-Level Arithmetic Functions . . . . . . . . . 305

14.1 “Not a Number” Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 14.2 Predicates on Floats. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 14.3 Absolute Value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 14.4 Normalization Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 14.5 Rounding and Remainder Functions . . . . . . . . . . . . . . . . . . . 308 14.6 Integer Division . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 14.7 Parsing of Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310

14.7.1 Parsing of Integers. . . . . . . . . . . . . . . . . . . . . . . . . . . 310 14.7.2 Parsing of Floats . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314

15 Searching and Sorting . . . . . . . . . . . . . . . . . . 317

15.1 Defining the Comparison Function. . . . . . . . . . . . . . . . . . . . . 317 15.2 Array Search Function. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 15.3 Array Sort Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 15.4 Searching and Sorting Example . . . . . . . . . . . . . . . . . . . . . . . 319

16 Pattern Matching . . . . . . . . . . . . . . . . . . . . . . 323

16.1 Wildcard Matching. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 16.2 Globbing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324

16.2.1 Calling glob . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 16.2.2 Flags for Globbing. . . . . . . . . . . . . . . . . . . . . . . . . . . 326

16.3 Regular Expression Matching . . . . . . . . . . . . . . . . . . . . . . . . . 327 16.3.1 POSIX Regular Expression Compilation. . . . . . . 327 16.3.2 Flags for POSIX Regular Expressions . . . . . . . . . 329 16.3.3 Matching a Compiled POSIX Regular Expression

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 330 16.3.4 Match Results with Subexpressions . . . . . . . . . . . 331 16.3.5 Complications in Subexpression Matching . . . . . 332 16.3.6 POSIX Regexp Matching Cleanup . . . . . . . . . . . . 332

16.4 Shell-Style Word Expansion . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 16.4.1 The Stages of Word Expansion . . . . . . . . . . . . . . . 334 16.4.2 Calling wordexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 16.4.3 Flags for Word Expansion . . . . . . . . . . . . . . . . . . . . 336 16.4.4 wordexp Example . . . . . . . . . . . . . . . . . . . . . . . . . . . 337

17 Date and Time . . . . . . . . . . . . . . . . . . . . . . . . . 339

17.1 Processor Time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 17.1.1 Basic CPU Time Inquiry . . . . . . . . . . . . . . . . . . . . . 339 17.1.2 Detailed Elapsed CPU Time Inquiry . . . . . . . . . . 340

17.2 Calendar Time. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 341 17.2.1 Simple Calendar Time . . . . . . . . . . . . . . . . . . . . . . . 342 17.2.2 High-Resolution Calendar . . . . . . . . . . . . . . . . . . . . 342 17.2.3 Broken-down Time . . . . . . . . . . . . . . . . . . . . . . . . . . 345 17.2.4 Formatting Date and Time . . . . . . . . . . . . . . . . . . . 347 17.2.5 Specifying the Time Zone with TZ . . . . . . . . . . . . 352 17.2.6 Functions and Variables for Time Zones . . . . . . . 354 17.2.7 Time Functions Example. . . . . . . . . . . . . . . . . . . . . 355

17.3 Setting an Alarm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 356 17.4 Sleeping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 17.5 Resource Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 360 17.6 Limiting Resource Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 17.7 Process Priority . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 364

18 Extended Characters . . . . . . . . . . . . . . . . . . . 367

18.1 Introduction to Extended Characters . . . . . . . . . . . . . . . . . . 367 18.2 Locales and Extended Characters . . . . . . . . . . . . . . . . . . . . . 368 18.3 Multibyte Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 368 18.4 Wide Character Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . 371 18.5 Conversion of Extended Strings . . . . . . . . . . . . . . . . . . . . . . . 372 18.6 Multibyte Character Length . . . . . . . . . . . . . . . . . . . . . . . . . . 373 18.7 Conversion of Extended Characters One by One . . . . . . . . 374 18.8 Character-by-Character Conversion Example . . . . . . . . . . . 375 18.9 Multibyte Codes Using Shift Sequences . . . . . . . . . . . . . . . . 377

19 Locales and Internationalization . . . . . . . . . 379

19.1 What Effects a Locale Has . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 19.2 Choosing a Locale. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 19.3 Categories of Activities that Locales Affect. . . . . . . . . . . . . 380 19.4 How Programs Set the Locale . . . . . . . . . . . . . . . . . . . . . . . . . 381 19.5 Standard Locales. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383 19.6 Numeric Formatting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383

19.6.1 Generic Numeric Formatting Parameters . . . . . . 384 19.6.2 Printing the Currency Symbol . . . . . . . . . . . . . . . . 385 19.6.3 Printing the Sign of an Amount of Money . . . . . 387

20 Non-Local Exits . . . . . . . . . . . . . . . . . . . . . . . . 389

20.1 Introduction to Non-Local Exits. . . . . . . . . . . . . . . . . . . . . . . 389 20.2 Details of Non-Local Exits . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391 20.3 Non-Local Exits and Signals . . . . . . . . . . . . . . . . . . . . . . . . . . 392

21 Signal Handling . . . . . . . . . . . . . . . . . . . . . . . . 393

21.1 Basic Concepts of Signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 393 21.1.1 Some Kinds of Signals . . . . . . . . . . . . . . . . . . . . . . . 393 21.1.2 Concepts of Signal Generation . . . . . . . . . . . . . . . . 394 21.1.3 How Signals Are Delivered . . . . . . . . . . . . . . . . . . . 394

21.2 Standard Signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 21.2.1 Program Error Signals . . . . . . . . . . . . . . . . . . . . . . . 396 21.2.2 Termination Signals . . . . . . . . . . . . . . . . . . . . . . . . . 399 21.2.3 Alarm Signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 21.2.4 Asynchronous I/O Signals . . . . . . . . . . . . . . . . . . . . 401 21.2.5 Job Control Signals . . . . . . . . . . . . . . . . . . . . . . . . . . 402 21.2.6 Operation Error Signals . . . . . . . . . . . . . . . . . . . . . . 403 21.2.7 Miscellaneous Signals . . . . . . . . . . . . . . . . . . . . . . . . 404 21.2.8 Signal Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 405

21.3 Specifying Signal Actions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 21.3.1 Basic Signal Handling. . . . . . . . . . . . . . . . . . . . . . . . 406 21.3.2 Advanced Signal Handling . . . . . . . . . . . . . . . . . . . 408 21.3.3 Interaction of signal and sigaction . . . . . . . . . 410 21.3.4 sigaction Function Example . . . . . . . . . . . . . . . . 410 21.3.5 Flags for sigaction . . . . . . . . . . . . . . . . . . . . . . . . . 412 21.3.6 Initial Signal Actions . . . . . . . . . . . . . . . . . . . . . . . . 413

21.4 Defining Signal Handlers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413 21.4.1 Signal Handlers that Return. . . . . . . . . . . . . . . . . . 414 21.4.2 Handlers That Terminate the Process . . . . . . . . . 415 21.4.3 Nonlocal Control Transfer in Handlers . . . . . . . . 416 21.4.4 Signals Arriving While a Handler Runs. . . . . . . . 417 21.4.5 Signals Close Together Merge into One . . . . . . . . 418 21.4.6 Signal Handling and Nonreentrant Functions . . 421 21.4.7 Atomic Data Access and Signal Handling. . . . . . 423

21.4.7.1 Problems with Non-Atomic Access … 423 21.4.7.2 Atomic Types . . . . . . . . . . . . . . . . . . . . . . 424 21.4.7.3 Atomic Usage Patterns . . . . . . . . . . . . . . 425

21.5 Primitives Interrupted by Signals . . . . . . . . . . . . . . . . . . . . . . 425 21.6 Generating Signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 426

21.6.1 Signaling Yourself . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 21.6.2 Signaling Another Process. . . . . . . . . . . . . . . . . . . . 428 21.6.3 Permission for using kill . . . . . . . . . . . . . . . . . . . . 429 21.6.4 Using kill for Communication . . . . . . . . . . . . . . . 430

21.7 Blocking Signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 431 21.7.1 Why Blocking Signals is Useful . . . . . . . . . . . . . . . 432 21.7.2 Signal Sets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 432 21.7.3 Process Signal Mask . . . . . . . . . . . . . . . . . . . . . . . . . 433 21.7.4 Blocking to Test for Delivery of a Signal. . . . . . . 435 21.7.5 Blocking Signals for a Handler . . . . . . . . . . . . . . . . 436 21.7.6 Checking for Pending Signals . . . . . . . . . . . . . . . . . 437 21.7.7 Remembering a Signal to Act On Later . . . . . . . 438

21.8 Waiting for a Signal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 439 21.8.1 Using pause . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 439 21.8.2 Problems with pause . . . . . . . . . . . . . . . . . . . . . . . . 440 21.8.3 Using sigsuspend . . . . . . . . . . . . . . . . . . . . . . . . . . . 441

21.9 Using a Separate Signal Stack . . . . . . . . . . . . . . . . . . . . . . . . . 442 21.10 BSD Signal Handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 444

21.10.1 BSD Function to Establish a Handler . . . . . . . . 445 21.10.2 BSD Functions for Blocking Signals . . . . . . . . . . 446

22 Process Startup and Termination . . . . . . . . 447

22.1 Program Arguments. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447 22.1.1 Program Argument Syntax Conventions . . . . . . . 448 22.1.2 Parsing Program Options . . . . . . . . . . . . . . . . . . . . 449 22.1.3 Example of Parsing Arguments with getopt … 450 22.1.4 Parsing Long Options . . . . . . . . . . . . . . . . . . . . . . . . 452 22.1.5 Example of Parsing Long Options. . . . . . . . . . . . . 454 22.1.6 Parsing of Suboptions. . . . . . . . . . . . . . . . . . . . . . . . 456 22.1.7 Parsing of Suboptions Example . . . . . . . . . . . . . . . 457

22.2 Environment Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 458 22.2.1 Environment Access . . . . . . . . . . . . . . . . . . . . . . . . . 459 22.2.2 Standard Environment Variables. . . . . . . . . . . . . . 460

22.3 Program Termination . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 462 22.3.1 Normal Termination . . . . . . . . . . . . . . . . . . . . . . . . . 462 22.3.2 Exit Status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 22.3.3 Cleanups on Exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . 464 22.3.4 Aborting a Program . . . . . . . . . . . . . . . . . . . . . . . . . 465 22.3.5 Termination Internals . . . . . . . . . . . . . . . . . . . . . . . . 465

23 Processes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 467

23.1 Running a Command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 467 23.2 Process Creation Concepts . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 23.3 Process Identification. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 23.4 Creating a Process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 469 23.5 Executing a File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 470 23.6 Process Completion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 23.7 Process Completion Status. . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 23.8 BSD Process Wait Functions . . . . . . . . . . . . . . . . . . . . . . . . . . 476 23.9 Process Creation Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477

24 Job Control . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479

24.1 Concepts of Job Control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 24.2 Job Control is Optional. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 24.3 Controlling Terminal of a Process . . . . . . . . . . . . . . . . . . . . . 480 24.4 Access to the Controlling Terminal . . . . . . . . . . . . . . . . . . . . 481 24.5 Orphaned Process Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 24.6 Implementing a Job Control Shell . . . . . . . . . . . . . . . . . . . . . 482

24.6.1 Data Structures for the Shell . . . . . . . . . . . . . . . . . 482 24.6.2 Initializing the Shell . . . . . . . . . . . . . . . . . . . . . . . . . 484 24.6.3 Launching Jobs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 24.6.4 Foreground and Background. . . . . . . . . . . . . . . . . . 490 24.6.5 Stopped and Terminated Jobs . . . . . . . . . . . . . . . . 492 24.6.6 Continuing Stopped Jobs. . . . . . . . . . . . . . . . . . . . . 496 24.6.7 The Missing Pieces . . . . . . . . . . . . . . . . . . . . . . . . . . 496 24.7 Functions for Job Control. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497

24.7.1 Identifying the Controlling Terminal . . . . . . . . . . 497 24.7.2 Process Group Functions . . . . . . . . . . . . . . . . . . . . . 498 24.7.3 Functions for Controlling Terminal Access . . . . . 499

25 System Databases and Name Service Switch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501

25.1 NSS Basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 25.2 The NSS Configuration File. . . . . . . . . . . . . . . . . . . . . . . . . . . 502

25.2.1 Services in the NSS configuration File . . . . . . . . . 503 25.2.2 Actions in the NSS configuration . . . . . . . . . . . . . 503 25.2.3 Notes on the NSS Configuration File . . . . . . . . . . 504

25.3 NSS Module Internals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 505 25.3.1 The Naming Scheme of the NSS Modules. . . . . . 505 25.3.2 The Interface of the Function in NSS Modules. . 506

25.4 Extending NSS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 25.4.1 Adding another Service to NSS . . . . . . . . . . . . . . . 508 25.4.2 Internals of the NSS Module Functions . . . . . . . . 509

26 Users and Groups . . . . . . . . . . . . . . . . . . . . . . 511

26.1 User and Group IDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 26.2 The Persona of a Process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 26.3 Why Change the Persona of a Process? . . . . . . . . . . . . . . . . 512 26.4 How an Application Can Change Persona . . . . . . . . . . . . . . 513 26.5 Reading the Persona of a Process. . . . . . . . . . . . . . . . . . . . . . 513 26.6 Setting the User ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 26.7 Setting the Group IDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 26.8 Enabling and Disabling Setuid Access . . . . . . . . . . . . . . . . . 516 26.9 Setuid Program Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 26.10 Tips for Writing Setuid Programs . . . . . . . . . . . . . . . . . . . . 520 26.11 Identifying Who Logged In . . . . . . . . . . . . . . . . . . . . . . . . . . 520 26.12 User Database . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 521

26.12.1 The Data Structure that Describes a User …. 521 26.12.2 Looking Up One User. . . . . . . . . . . . . . . . . . . . . . . 522 26.12.3 Scanning the List of All Users . . . . . . . . . . . . . . . 523 26.12.4 Writing a User Entry . . . . . . . . . . . . . . . . . . . . . . . 525

26.13 Group Database . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525 26.13.1 The Data Structure for a Group . . . . . . . . . . . . . 525 26.13.2 Looking Up One Group . . . . . . . . . . . . . . . . . . . . . 525 26.13.3 Scanning the List of All Groups . . . . . . . . . . . . . 527

26.14 Netgroup Database. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 26.14.1 Netgroup Data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 26.14.2 Looking up one Netgroup . . . . . . . . . . . . . . . . . . . 529

26.14.3 Testing for Netgroup Membership. . . . . . . . . . . . 530 26.15 User and Group Database Example. . . . . . . . . . . . . . . . . . . 530

27 System Information. . . . . . . . . . . . . . . . . . . . . 533

27.1 Host Identification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 533 27.2 Hardware/Software Type Identification . . . . . . . . . . . . . . . . 534

28 System Configuration Parameters . . . . . . . 537

28.1 General Capacity Limits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 537 28.2 Overall System Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 28.3 Which Version of POSIX is Supported . . . . . . . . . . . . . . . . . 540 28.4 Using sysconf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541

28.4.1 Definition of sysconf . . . . . . . . . . . . . . . . . . . . . . . . 541 28.4.2 Constants for sysconf Parameters . . . . . . . . . . . . 541 28.4.3 Examples of sysconf . . . . . . . . . . . . . . . . . . . . . . . . 543

28.5 Minimum Values for General Capacity Limits . . . . . . . . . . 544 28.6 Limits on File System Capacity . . . . . . . . . . . . . . . . . . . . . . . 545 28.7 Optional Features in File Support . . . . . . . . . . . . . . . . . . . . . 546 28.8 Minimum Values for File System Limits. . . . . . . . . . . . . . . . 547 28.9 Using pathconf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 548 28.10 Utility Program Capacity Limits . . . . . . . . . . . . . . . . . . . . . 549 28.11 Minimum Values for Utility Limits . . . . . . . . . . . . . . . . . . . 550 28.12 String-Valued Parameters. . . . . . . . . . . . . . . . . . . . . . . . . . . . 551

Appendix A C Language Facilities in the Library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553

A.1 Explicitly Checking Internal Consistency . . . . . . . . . . . . . . . 553 A.2 Variadic Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555

A.2.1 Why Variadic Functions are Used . . . . . . . . . . . . . 555 A.2.2 How Variadic Functions are Defined and Used . . 556

A.2.2.1 Syntax for Variable Arguments . . . . . . . 556 A.2.2.2 Receiving the Argument Values. . . . . . . 556 A.2.2.3 How Many Arguments Were Supplied

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 557 A.2.2.4 Calling Variadic Functions . . . . . . . . . . . 558 A.2.2.5 Argument Access Macros . . . . . . . . . . . . 559

A.2.3 Example of a Variadic Function . . . . . . . . . . . . . . . 559

A.2.3.1 Old-Style Variadic Functions . . . . . . . . . 560

A.3 Null Pointer Constant. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 561 A.4 Important Data Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 561 A.5 Data Type Measurements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 562

A.5.1 Computing the Width of an Integer Data Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 563 A.5.2 Range of an Integer Type . . . . . . . . . . . . . . . . . . . . . 563 A.5.3 Floating Type Macros . . . . . . . . . . . . . . . . . . . . . . . . 565 A.5.3.1 Floating Point Representation Concepts

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 565 A.5.3.2 Floating Point Parameters . . . . . . . . . . . 566 A.5.3.3 IEEE Floating Point . . . . . . . . . . . . . . . . . 570

A.5.4 Structure Field Offset Measurement . . . . . . . . . . . 570

Appendix B Summary of Library Facilities . . 573

Appendix C Library Maintenance . . . . . . . . . . 667

C.1 How to Install the GNU C Library . . . . . . . . . . . . . . . . . . . . . 667

C.1.1 Recommended Tools to Install the GNU C Library

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 669 C.1.2 Supported Configurations . . . . . . . . . . . . . . . . . . . . . 670

C.2 Reporting Bugs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 C.3 Adding New Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 672 C.4 Porting the GNU C Library . . . . . . . . . . . . . . . . . . . . . . . . . . . 674

C.4.1 Layout of the ‘sysdeps’ Directory Hierarchy … 677 C.4.2 Porting the GNU C Library to Unix Systems… 680

C.5 Contributors to the GNU C Library . . . . . . . . . . . . . . . . . . . . 680

Appendix D GNU LIBRARY GENERAL PUBLIC LICENSE . . . . . . . . . . . . . . . . . . . . . 687

Preamble . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 687 TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION

AND MODIFICATION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 689 How to Apply These Terms to Your New Libraries . . . . . . . . . . . 696

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 697

Type Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 709

Function and Macro Index . . . . . . . . . . . . . . . . . . 711

Variable and Constant Macro Index. . . . . . . . . . 719

Program and File Index . . . . . . . . . . . . . . . . . . . . . 727